A Complete Bibliography of Publications in the Journal of Network and Computer Applications

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

06 August 2024 Version 1.44

Title word cross-reference

(t, n) [3532]. 2 [994]. 3 [40, 655, 695, 1071, 1250, 1585, 2142, 3555]. 5 [2271, 2274, 2287, 2331, 2339, 2358, 2361, 2443, 2461, 2486, 2487, 2497, 2505, 2521, 2581]. ^{+} [2649]. ^{1} [647]. ^{2} [866, 3565]. ^{3} [3333]. ^{∘} [3129]. ^{(}R) [84]. H_{∞} [623]. k [639, 1168, 1450, 1750, 2013]. M [860]. N [104, 392, 1727, 3076]. Q [3002, 3305]. r [2496]. t [392]. x [1760].

* [2943].

-based [1945]. -centric [3166]. -class [994]. -connected [2013]. -coverage [1168, 1750]. -covering [639]. -Dimension [860]. -Dimensional [104, 2496]. -disjoint [275]. -Folded [1760]. -gram [1727]. -grams [3076]. -learning [3002]. -learning-based [3305]. -out-of- [392]. -threshold [3532].

1 [74, 2593, 2616, 2640, 2654, 2670, 2685, 2700, 2715, 2737, 2756, 2779, 2804, 2828, 2853, 2889, 2912, 2936, 2954, 2969, 2992, 3003, 3023, 3046, 3064]. 127 [2661]. 132 [2516]. 134 [2591]. 142 [2587]. 15 [2605, 2629, 2647, 2662, 2677, 2693, 2707, 2726, 2746, 2767, 2792, 2816, 2841, 2872, 2900, 2924, 2944, 2962, 2985, 2998, 3012, 3035, 3056, 3072]. 173 [2888]. 184 [2987]. 19 [2886, 2941, 3382].

2 [8, 2007, 2926]. 2015 [1659]. 2016 [2058]. 2019 [2593, 2605]. 2020 [2616, 2629, 2640, 2647, 2654, 2662, 2670, 2677, 2685, 2693, 2700, 2707, 2715, 2726, 2737, 2746, 2756, 2767, 2779, 2792, 2804, 2816, 2828, 2841]. 2021 [2853, 2872, 2889, 2900, 2912, 2924, 2936, 2944, 2954, 2962, 2969, 2985, 2992, 2998, 3003, 3012, 3023, 3035, 3046, 3056, 3064, 3072]. 2022 [3081, 3089, 3100, 3110, 3122, 3135, 3163, 3176, 3199, 3211, 3229]. 2023 [3239, 3251, 3256, 3265, 3281, 3300, 3311, 3325, 3340, 3365, 3379, 3392]. 2024 [3405, 3431, 3446, 3460, 3479, 3490, 3506, 3515, 3520, 3541]. 2P [1742]. 2P-Lookup [1742].

35 [868]. 3rd [1088].

4 [1525]. 4.0 [2994, 3242]. 46 [1412, 1765]. 4G [2114].

5.0 [3450]. 5G [1739, 1828, 1941, 1942, 2044, 2063, 2072, 2114, 2120, 2122, 2167, 2213, 2216, 2219, 2220, 2223, 2235, 2243, 2250, 2252, 2256, 2267, 2667, 2709, 2710, 2719, 2751, 2769, 2798, 2893, 2897, 2935, 3008, 3031, 3055, 3086, 3095, 3137, 3150, 3157, 3175, 3201, 3304, 3316, 3322, 3375, 3388, 3409, 3417, 3458, 3513]. 5G-based [2072, 2751]. 5G-Enabled [2216]. 5G/6G [3201, 3458]. 5G/6G-enabled [3458].

6G [3201, 3288, 3324, 3347, 3357, 3523, 3544, 3569]. 6G-enabled [3458]. 6G-Underwater [3288]. 6LoWPAN [1379, 1581, 3408]. 6LoWPAN-based [1581].

802.11 [809, 1005, 1156, 1187, 1400, 1443, 1454, 1481, 1773, 1783, 2027]. 802.11-based [809, 1156, 1187, 1454, 1481]. 802.11ah [2964]. 802.11e [700, 1525]. 802.11n [1042, 1417, 1841]. 802.11p [810]. 802.15.4 [1320]. 802.15.4-based [719]. 802.16 [375, 528, 930]. 802.16e [899, 1149, 1614]. 802.16e/m [899]. 802.16j [939]. 802.16m [1409].

900 [1417]. 94 [2339].

A-MSDU [1042]. A1 [797]. AAA [290]. ABC [2799]. ABC-RuleMiner [2799]. abnormal [714, 1596]. abnormality [2754]. ABR [134]. abstraction [458]. abstractions [144]. AC [3048]. Academic [2150, 2454]. ACAI [200]. ACCC [1709]. accelerate [3207]. Accelerated [1768, 2358, 2503, 3283]. accelerating [2957]. acceleration [907, 1963, 3333, 3378]. accelerometers [2801]. acceptable [265]. Access [59, 83, 188, 290, 311, 428, 450, 492, 725, 813, 909, 973, 978, 1024, 1027, 1035, 1064, 1188, 1221, 1238, 1244, 1246, 1400, 1498, 1773, 1775, 1786, 1828, 1872, 1918, 1953, 1965, 1972, 2112, 2137, 2149, 2196, 2213, 2233, 2235, 2243, 2246, 2267, 2273, 2332, 2334, 2349, 2377, 2433, 2460, 2486, 2502, 2513, 2521, 2523, 2531, 2563, 2604, 2622, 2639, 2714, 2786, 2795, 2820, 2864, 2930, 2974, 2987, 3031, 3038, 3050, 3107, 3134, 3151, 3157, 3227, 3250, 3274, 3384, 3409, 3413, 3513, 3522]. accessibility [33, 348]. accessing [140, 202, 257, 910, 2849]. Accident [913]. Accomplishing [2301]. according [2470]. accountability [1087, 2233]. accounts [1854, 2232, 3401]. ACCPndn [1469]. Accreditation [3190]. Accumulating [809]. Accuracy [509, 1604, 2533]. Accuracy-aware [1604]. Accurate [1219, 1274, 1357, 2134, 2564, 3338, 3477]. accurately [2134]. achievable [1500]. achieve [574, 598]. Achievements [3262]. Achieving [613, 833, 1291, 2188, 2776]. ACK [1841]. Acknowledgement [1783]. ACO [777, 1310, 2550]. Acoustic [1274, 1764, 1805, 2017–2019, 2022, 2023, 2864, 3351]. acquisition [1333, 3071]. across [2604, 2879, 3154]. Action [3563]. Action-branching [3563]. activated [478]. activation [2810]. Active [103, 155, 197, 322, 381, 623, 633, 1058, 1063, 1084, 1122, 1213, 1219, 1251, 1357, 1529, 1622, 2035, 2976, 3018, 3419, 3444, 3477]. ActiveGuardian [3477]. activeness [2681]. activities [1272]. Activity [17, 351, 362, 667, 1268, 2084, 2200, 2407, 2533, 2785, 2801, 3172, 3496]. Actor [669, 776, 818, 1794, 3017, 3319]. actor-based [3017]. actuator [916, 1096, 1611, 1868, 2368]. Ad [155, 258, 273, 279, 292, 320, 357, 415, 442, 472, 474, 477, 529, 536, 541, 546, 572, 585, 597, 662, 679, 685, 691, 694, 721, 724, 736, 764, 767, 781, 790, 828, 835, 845, 880, 888, 898, 915, 926, 956, 983, 1009, 1019, 1026, 1048, 1124, 1127, 1155, 1157, 1173, 1190, 1205, 1276, 1315, 1326, 1329, 1351, 1354, 1399, 1419, 1475, 1481, 1501, 1521, 1526, 1575, 1603, 1606, 1647, 1696, 1709, 1747, 1751, 1754, 1761, 1819, 1900, 1903, 2078, 2086, 2164, 2220, 2312, 2336, 2353, 2377, 2405, 2468, 2858, 2961, 3040, 3105, 3161]. Ad-Hoc [477, 529, 536, 597, 691, 926, 1315, 1326, 1475, 1575, 1900, 2336, 2858, 2961, 3040]. adaptability [1295]. Adaptation [154, 293– 295, 297, 488, 524, 690, 710, 776, 1097, 1156, 1221, 1835, 2256, 2559, 2974, 2987]. adapted [2099]. Adapting [87, 872, 3394]. Adaptive [4, 46, 63, 94, 245, 312, 367, 369, 418, 440, 466, 538, 545, 558, 606, 626, 679, 703, 709, 736, 770, 819, 838, 928, 979, 1035, 1045, 1051, 1052, 1121, 1146, 1152, 1216, 1225, 1247, 1300, 1311, 1343, 1347, 1425, 1439, 1469, 1470, 1484, 1491, 1518, 1593, 1648, 1696, 1738, 1765, 1779, 1845, 1861, 1867, 1884, 2003, 2034, 2048, 2074, 2096, 2129, 2169, 2201, 2305, 2313, 2319, 2347, 2355, 2393, 2424, 2450, 2485, 2515, 2610, 2675, 2691, 2761, 2764, 2795, 2862, 2864, 2972, 2990, 3157, 3192, 3295, 3338, 3348, 3386, 3425, 3459, 3488, 3499]. adaptive-learning [2347]. adaptive-predictive [709]. adaptively [515]. adaptivity [296]. added [346]. adders [2823]. Adding [1550]. additions [920]. address [321, 1581, 1768, 2813]. addresses [379]. addressing [1222, 3124]. adjacency [2364]. adjustable [1318]. adjusting [447, 3561]. Adjustment [1038, 1328, 1680, 2060, 2201, 2542]. administration [90]. Administrative [1709]. administrators [1430]. admission [555, 612, 724, 872, 1291, 1343, 1765, 1985, 3108]. adopted [3272]. Adopting [2931, 2994]. adoption [1494, 2385]. ads [3261]. Advance [2179]. Advanced [408, 537, 548, 650, 727, 1126, 1465, 1561, 1598, 1630, 1969, 2000, 2291, 2789, 2961, 3550]. Advancements [2188, 3355, 3484]. Advances [503, 544, 1094, 1821, 2244, 2506, 2742, 2956, 3025, 3091, 3113, 3219, 3273, 3384, 3452]. advantage [161, 761]. Adversarial [2717, 2733, 2982, 3126, 3141, 3309, 3397]. advertised [379]. advertising [2236, 3261]. advice [805]. advising [1213]. advisory [62]. AEP [2472]. AEP-PPA [2472]. aerial [1209, 1386, 2061, 2072, 2141, 2353, 2613, 3220, 3262, 3294, 3321, 3324]. AFCD [1728]. affect [318, 1717]. affecting [99]. affective [2626]. Affinity [2056]. against [288, 353, 356, 592, 657, 751, 845, 983, 1147, 1754, 1825, 1840, 1960, 2090, 2204, 2215, 2319, 2449, 2494, 2581, 2735, 2761, 2919, 3014, 3444, 3448]. Age [3126, 3330]. agenda [2592]. Agent [34, 138, 184, 187, 191, 197, 200, 212, 225, 227, 250, 265, 299, 300, 302, 304, 305, 319, 326, 355, 446, 544– 546, 548, 549, 584, 716, 748, 847, 967, 1088, 1099–1102, 1104–1106, 1157, 1232, 1466, 1751, 1928, 2355, 2430, 2975, 3226, 3422, 3453, 3462, 3475, 3526, 3545]. agent-assisted [748]. Agent-based [138, 187, 197, 200, 212, 265, 305, 326, 544, 545, 1088, 1104, 1106, 1466, 2430]. Agent-oriented [1099]. agents [63, 226, 228, 230, 252, 254, 303, 307, 365, 403, 407, 448, 551, 723, 1738, 1923, 2269, 3088]. aggregate [919, 2349]. aggregated [1339]. Aggregating [2760]. Aggregation [350, 438, 734, 775, 793, 882, 895, 1008, 1018, 1042, 1112, 1207, 1214, 1232, 1562, 1591, 1734, 2039, 2068, 2180, 2323, 2373, 2379, 2390, 2528, 2546, 2790, 2813, 2830, 3014, 3215, 3258, 3395, 3469]. agile [1334, 1464, 1501, 2576]. Agile-SD [1464]. agnostic [3552]. agreement [264, 585, 886, 960, 1012, 1135, 1544, 1595, 1994, 2341, 2395, 3408]. Agreements [217, 2000]. AGRIB [1008]. AGRIB-BHF [1008]. agriculture [2613, 2782, 2994, 3519]. ahead [1732, 1925]. AI [716, 2852, 2983, 3268, 3329, 3399, 3484, 3529, 3533]. AI-based [3533]. AI-driven [2852]. AI-enhanced [3484]. AI-powered [3399]. AID [1526]. Aided [26, 124, 966, 1209, 2176, 2219, 3330, 3331, 3453, 3563]. AIoT [3519, 3525]. AIPISteg [1622]. air [3179, 3469]. airborne [1366, 3221]. aircraft [1543]. AKA3 [3131]. Alert [470, 1505, 2067]. alerting [325]. alerts [3476]. algebra [374, 1092, 1706]. algebra-based [1092]. Algebraic [165]. Algorithm [11, 256, 262, 277, 323, 350, 422, 434, 437, 521, 528, 572, 573, 581, 583, 672, 691, 694, 721, 730, 748, 854, 866, 882, 884, 889, 917, 930, 947, 979, 993, 995, 1002, 1039, 1091, 1140, 1141, 1168, 1216, 1233, 1243, 1247, 1250, 1263, 1295, 1443, 1447, 1464, 1475, 1500, 1527, 1566, 1614, 1645, 1699, 1710, 1721, 1756, 1762, 1833, 1938, 1946, 1996, 2029, 2031, 2032, 2083, 2132, 2134, 2140, 2141, 2239, 2272, 2298, 2318, 2364, 2380, 2388, 2395, 2405, 2463, 2529, 2609, 2646, 2651, 2796, 2838, 2862, 2884, 2949, 3060, 3063, 3116, 3162, 3210, 3249, 3278, 3351, 3359, 3438, 3456, 3485, 3497, 3513]. algorithm [30, 189, 319, 394, 476, 619, 770, 773, 777, 797, 801, 914, 922, 924, 1090, 1310, 1341, 1436, 1503, 1520, 1535, 1537, 1560, 1610, 1692, 1741, 1871, 1988, 1991, 2117, 2305, 2368, 2782, 2860, 2927, 2931, 3145, 3186, 3284, 3305, 3344, 3412, 3429, 3433, 3477, 3531]. algorithmic [677]. Algorithmically [3020]. Algorithms [3, 13, 70, 269, 326, 327, 532, 614, 739, 818, 850, 872, 918, 928, 977, 1107, 1120, 1303, 1305, 1353, 1371, 1373, 1382, 1390, 1517, 1629, 1723, 1761, 1784, 1939, 1980, 2087, 2167, 2392, 2417, 2667, 2966, 3137, 3220, 3233, 3443, 3467]. alias [1358]. alignment [504, 2261]. alignment-free [504, 2261]. all-packets-based [2981]. all-pairs [2318]. alleviation [1804, 3411]. allocate [1711]. Allocation [359, 475, 528, 573, 581, 670, 679, 727, 923, 980, 1103, 1114, 1146, 1149, 1154, 1269, 1338, 1341, 1343, 1344, 1366, 1470, 1491, 1700, 1762, 1765, 1847, 1858, 1946, 2074, 2098, 2129, 2155, 2223, 2230, 2243, 2313, 2400, 2579, 2583, 2674, 2706, 2710, 2735, 2813, 2874, 2894, 2959, 3051, 3096, 3169, 3243, 3283, 3357, 3414, 3430, 3482, 3508, 3511, 3522, 3527, 3539, 3560]. allow [776, 1213]. ALOHA [192]. alone [943]. alpha [218]. ALSketch [3338]. alternates [2643]. alternating [1458]. alternative [2529]. alternatives [702]. ALVS [3192]. Ambient [617, 1247]. AME [1465]. AME-WPC [1465]. America [37]. AMI [1457, 1562, 2299]. amid [2886]. AmoebaNet [2302]. among [168, 1853, 1964, 3026]. amplification [2540]. Anagram [1223]. analyser [401]. analyses [1803, 2744]. analysing [270]. Analysis [110, 111, 152, 192, 233, 235, 258, 322, 346, 365, 401, 422, 432, 444, 445, 507, 584, 621, 642, 667, 712, 722, 765, 815, 835, 885, 897, 971, 1004, 1041, 1044, 1069, 1070, 1104, 1128, 1131, 1164, 1192, 1196, 1314, 1323, 1340, 1345, 1369, 1483, 1512, 1542, 1563, 1578, 1580, 1616, 1662, 1665, 1669, 1698, 1726, 1727, 1785, 1792, 1843, 1848, 1859, 1904, 1910, 1918, 1968, 1972, 1974, 2107, 2109, 2144, 2168, 2190, 2208, 2309, 2337, 2354, 2357, 2431, 2434, 2454, 2497, 2498, 2522, 2652, 2797, 2803, 2914, 2915, 2938, 2997, 3016, 3030, 3052, 3068, 3219, 3277, 3448, 3483, 3562]. analysis [210, 414, 514, 591, 822, 1022, 1026, 1355, 1490, 1607, 1627, 1899, 2067, 2219, 2226, 2368, 2590, 3367, 3399, 3482, 3522, 3531]. analytic [1675]. Analytical [261, 368, 435, 1053, 1235, 2062, 2362, 2491, 2544, 2625, 2720, 3030, 3417]. Analytics [1564, 1901, 2121, 2221, 2234, 2418, 2626, 2753, 3099, 3488]. analytics-based [2234]. analyze [188, 1568]. Analyzing [1081, 1440, 1813, 1957, 2926]. Anatomizing [2322]. anchor [1362, 1486, 1939]. anchoring [1431]. anchorless [2762]. Andean [3116]. Android [1041, 2112, 2182, 2407, 2408, 2689, 3559]. Android-based [1041]. ANFIS [1157]. angle [2029]. animations [225]. ANN [378]. ANN-based [378]. annotation [101]. Announcement [203, 204]. anomalies [1703, 2588]. anomalous [1773]. Anomaly [245, 479, 620, 676, 904, 1201, 1245, 1572, 1627, 1671, 2067, 2175, 2222, 2357, 2414, 2611, 2736, 3009, 3236, 3309, 3434]. anomaly-based [2067, 3009, 3309]. anonymity [276, 1595, 2176]. anonymization [1138, 1480, 3397]. anonymous [387, 425, 582, 661, 1294, 2139, 2157, 2183, 2373, 2472, 2910, 3375]. Ant [350, 394, 581, 777, 918, 1171, 1182, 1520, 1527, 1627]. ant-based [581, 1527].

ANTALG [1310]. AntCom [3368]. antenna [3354]. antennas [273]. Anti [631, 712, 1169, 1378, 1693, 2734, 2736, 3130, 3363, 3368, 3416]. anti-collision [631, 1378]. anti-concept [2736]. anti-disturbance [2734]. anti-forensics [3416]. Anti-repackaging [3363]. anti-rumor [3130]. anti-spam [712]. Anti-swindle [1169]. anti-tracking [3368]. anticipative [447]. antithetic [3026]. antithetic-parties [3026]. AntNet [1295]. Ants [1295]. anycast [1157]. AODV [250, 353, 857]. AP [611, 2093, 2750, 3565]. AP-dense [2093]. API [3370]. API-MalDetect [3370]. app [2133, 2163, 2381, 2880]. APPA [2373]. Appl [868, 1412, 1765, 1766, 2339, 2516, 2587, 2661, 2888, 2987]. appliance [268, 655]. Applicability [1017]. Application [6, 75, 122, 186, 208, 213, 218, 239, 267, 402, 405, 463, 478, 527, 711, 738, 840, 1105, 1166, 1212, 1265, 1282, 1373, 1418, 1426, 1475, 1506, 1512, 1551, 1577, 1683, 1845, 1923, 1935, 1972, 2160, 2170, 2233, 2234, 2304, 2390, 2445, 2500, 2566, 2711, 2766, 2788, 2824, 2861, 2886, 2907, 2959, 2965, 2983, 3049, 3116, 3242, 3275, 3290, 3390, 3473, 3559]. Application-aware [1282, 2959]. application-based [3242]. application-layer [186]. Application-level [239]. Applications [5, 8, 15, 60, 163, 178, 200, 202, 237, 255, 341, 355, 448, 468, 469, 480, 501, 534, 536, 563, 569, 574, 630, 644, 688, 696, 708, 771, 780, 804, 932, 940, 953, 982, 1068, 1078, 1117, 1390, 1403, 1425, 1471, 1564, 1604, 1689, 1759, 1771, 1772, 1778, 1824, 1897, 1901, 1913, 1930, 1996, 2037, 2052, 2070, 2080, 2083, 2105, 2175, 2205, 2255, 2267, 2321, 2337, 2359, 2365, 2378, 2388, 2436, 2454, 2483, 2486, 2512, 2514, 2544, 2560, 2576, 2684, 2689, 2702, 2742, 2770, 2771, 2806, 2845, 2932, 3062, 3109, 3179, 3208, 3225, 3330, 3349, 3383, 3458, 3502, 3512, 3567]. applications [65, 100, 179, 241, 359, 376, 408, 426, 501, 514, 531, 555, 558, 578, 626, 628, 643, 729, 733, 745, 971, 1047, 1053, 1061, 1077, 1121, 1204, 1290, 1292, 1320, 1388, 1436, 1442, 1483, 1533, 1597, 1641, 1658, 1733, 1802, 1809, 1852, 1915, 1988, 1998, 2069, 2144, 2179, 2181, 2290, 2383, 2513, 2763, 2821, 2906, 2916, 2951, 3025, 3044, 3058, 3138, 3149, 3235, 3285, 3433, 3440, 3454, 3498, 3557]. Applied [430, 805, 977, 2298, 3040, 3273]. Apply [184]. Applying [678, 1213, 2958]. appointment [901]. Approach [6, 12, 46, 60, 95, 103, 177, 187, 191, 229, 230, 241, 258, 290, 366, 384, 497, 507, 545, 568, 596, 642, 674, 679, 725, 736, 910, 961, 1010, 1011, 1032, 1038, 1040, 1067, 1104, 1118, 1129, 1152, 1168, 1201, 1207, 1296, 1329, 1346, 1362, 1387, 1392, 1492, 1504, 1628, 1634, 1679, 1710, 1711, 1714, 1794, 1830, 1861, 1864, 1867, 1870, 1909, 1968, 2015, 2052, 2061, 2129, 2144, 2205, 2222, 2236, 2253, 2281, 2319, 2348, 2370, 2371, 2471, 2542, 2632, 2699, 2704, 2740, 2774, 2848, 2917, 2946, 2957, 2981, 2983, 3027, 3069, 3125, 3128, 3228, 3390, 3420, 3470, 3558, 3565]. approach [147, 188, 192, 212, 529, 547, 658, 677, 714, 742, 908, 929, 989, 1074, 1083, 1119, 1153, 1162, 1281, 1331, 1338, 1365, 1412, 1487, 1608, 1656, 1718, 1781, 1941, 2066, 2089, 2153, 2347, 2400, 2440, 2476, 2585, 2591, 2645, 2688, 2929, 2935, 3007, 3048, 3250, 3290, 3344, 3382, 3388]. approaches [193, 895, 1234, 1473, 1515, 1753, 1775, 1807, 1844, 2133, 2232, 2311, 2398, 2692, 2942, 3260, 3272, 3287, 3318, 3434]. Approaching [1668]. appropriateness [2280]. approximation [783]. apps [2652, 3374]. April [2670, 2677, 2936, 2944, 3110, 3281, 3460]. apropos [1300].

APSP [2318]. AQM [1231, 1490, 3535]. aquifer [13]. Arabic [146]. arbitrary [3415]. arbitrated [39]. ARC [824]. Archimedes [1363]. architectonics [2797]. architectural [3481]. Architecture [18, 52, 57, 110, 123, 130, 290, 313, 335, 413, 492, 519, 540, 550, 565, 574, 609, 644, 705, 709, 778, 811, 814, 831, 871, 873, 962, 970, 1081, 1096, 1122, 1312, 1390– 1393, 1584, 1624, 1625, 1675, 1682, 1745, 1788, 1796, 1811, 1829, 2080, 2096, 2128, 2129, 2143, 2229, 2287, 2288, 2316, 2343, 2492, 2504, 2519, 2556, 2653, 2836, 2886, 3010, 3097, 3112, 3167, 3168, 3224, 3232, 3315, 3356, 3375, 3384, 3435]. Architectures [46, 178, 253, 409, 753, 897, 940, 999, 1164, 1325, 1431, 1597, 1749, 1809, 1929, 2015, 2193, 2498, 2742, 3161, 3450, 3498, 3519]. archives [164]. Arduino [1320]. Area [111, 497, 696, 741, 797, 890, 1365, 1441, 1457, 1557, 2010, 2176, 2298, 2351, 2464, 2557, 2595, 2620, 2740, 2744, 2859, 2867, 2892, 3104, 3487]. areas [1926, 2231, 2308]. AREP [2022]. ARF [989]. ARGUS [2465]. ARIMA [2032]. ARKHAM [2000]. Armed [2402, 3558]. ARQ [1021]. Arrangement [1242]. arrays [16]. arrival [1251]. Arriving [800]. arrowhead [1886]. Art [1123, 1309, 1453, 1460, 1506, 1574, 1684, 1740, 1763, 2069, 2212, 2232, 2303, 2623, 2745, 2769, 2771, 2808, 2866, 2983, 3029, 3352, 3503, 3508]. articles [1128]. Artificial [14, 228, 545, 549, 1259, 1263, 1439, 1645, 2692, 2709, 2870, 3134, 3159, 3519]. ASAP [1300]. Asian [2848]. aside [2247]. ASPA [2862]. aspects [120, 302, 1369]. assembly [229, 1072]. assess [17, 805]. Assessing [220, 377, 1494]. assessment [385, 1053, 1107, 1234, 1775, 1799, 2009, 2353, 2492, 2608]. asset [646, 3552]. Assignment [132, 328, 474, 481, 556, 559, 611, 660, 728, 955, 1005, 1141, 1429, 1654, 1753, 1866, 2371, 2417, 2471, 2741, 3377, 3394, 3468]. assignments [805]. Assist [1219, 2291]. assistance [2771, 3295]. assistants [1743]. Assisted [337, 638, 748, 945, 1098, 1576, 2347, 2351, 2411, 2501, 2568, 2676, 2959, 3008, 3189, 3248, 3376, 3412, 3475, 3526, 3539]. assisting [353, 2806]. assistive [178, 571]. associated [860, 1929, 2345, 3018]. association [306, 463, 1004, 3005, 3512]. associations [1085]. assumptions [3277]. assurance [2465]. assuring [2148, 3416]. asteroid [2028]. asthma [2938]. astronomical [396, 2028]. asymmetric [2022, 2341, 3368]. asymmetrical [2231]. asynchronous [46, 371, 451, 621, 2691, 3188, 3497]. ATM [38, 60, 65, 75, 201, 283]. atrial [3268]. attachable [151]. Attached [467]. Attack [237, 402, 498, 592, 822, 1086, 1136, 1185, 1561, 1754, 2007, 2090, 2113, 2164, 2204, 2230, 2312, 2407, 2567, 2717, 2735, 2796, 2877, 2981, 2995, 2996, 3048, 3087, 3155, 3191, 3275, 3352, 3389, 3454, 3486, 3514, 3551, 3553, 3554, 3564, 3565]. attacker [3093]. Attacks [258, 286, 288, 356, 403, 435, 453, 657, 675, 837, 983, 996, 1004, 1139, 1147, 1180, 1202, 1369, 1671, 1702, 1776, 1825, 1840, 1870, 1923, 1960, 2002, 2008, 2042, 2222, 2234, 2294, 2319, 2345, 2429, 2449, 2494, 2540, 2576, 2581, 2628, 2664, 2733, 2739, 2819, 2844, 2870, 2919, 2943, 3001, 3014, 3066, 3093, 3194, 3293, 3397, 3427, 3448, 3474, 3540, 3562]. attempts [511]. attention [2725, 3403, 3543]. attention-based [2725]. attentional [2921].

Attenuating [3564]. attestation [832, 1180, 1872]. Attribute [1956, 2190, 2195, 2332, 2341, 2423]. attribute-based [2195, 2332]. attributes [290, 3242, 3274, 3559]. auctions [2105]. audio [367, 2626, 2766]. audit [2570]. auditing [1277, 1662, 1894, 1922, 2404, 2426, 2470, 2522]. Augmented [1225, 1997, 3329]. Augmenting [144, 1048]. August [2737, 2746, 2992, 2998, 3163, 3340, 3515]. Authenticated [713, 976, 1012, 1544, 2185, 2341, 2351]. Authentication [349, 354, 483, 489, 509, 560, 609, 661, 664, 699, 811, 817, 856, 886, 892, 909, 960, 1055, 1056, 1210, 1392, 1438, 1563, 1583, 1595, 1810, 1908, 1959, 1994, 2011, 2071, 2099, 2114, 2120, 2139, 2157, 2176, 2200, 2273, 2278, 2279, 2316, 2349, 2443, 2472, 2638, 2719, 2843, 2909, 3001, 3029, 3038, 3050, 3055, 3095, 3131, 3171, 3183, 3227, 3361, 3376, 3408, 3464, 3475, 3532, 3536]. authoring [125]. authoritative [2902]. authority [2233, 2332, 2423, 2714]. authorization [290, 406, 1134, 1392, 2182, 3038]. authorized [3375]. authorship [1727]. Auto [1658, 2052, 2359, 2711, 2803, 3155]. Auto-Regressive [2803]. Auto-scaling [1658, 2052, 2359, 2711]. autoencoder [2869, 3048, 3198, 3486]. autoencoder-random [3486]. automata [529, 1296, 1337, 1864]. automata-based [1296, 1337]. automate [667]. Automated [302, 308, 1062, 2089, 2144, 2676, 2787, 2995, 3327, 3370, 3478]. Automatic [146, 890, 1028, 1145, 1702, 2725, 2749, 2948, 3021, 3053, 3134, 3143, 3548]. automating [412, 1904, 3204]. automation [2116]. automaton [2338, 3351]. automaton-based [3351]. AutoML [3267]. automotive [366]. Autonomic [428, 481, 547, 922, 1062, 1383, 2581]. Autonomous [410, 551, 613, 617, 905, 1589, 1738, 2545, 2734, 2810, 2891, 2929, 2968, 3385]. autonomy [305]. autoscaler [2621]. AUV [2377, 2480]. AV1 [2836]. Availability [548, 993, 1242, 1361, 1550, 1574, 2212, 2240, 2578, 3187]. Available [219, 808, 873, 963, 1187, 2159]. avatar [1179]. AVC [615]. average [973]. avionic [2887]. avionics [1366]. Avoidance [1111, 1647, 1944, 2032, 2646]. avoiding [785, 1680, 1787, 3555]. Awards [1659, 2058, 2538, 3364]. Aware [199, 200, 264, 279, 325, 341, 358, 452, 476, 535, 556, 625, 643, 716, 735, 775, 794, 807, 897, 923, 974, 1034, 1111, 1157, 1199, 1232, 1282, 1289, 1333, 1347, 1382, 1395, 1405, 1429, 1480, 1489, 1521, 1522, 1536, 1541, 1604, 1630, 1683, 1694, 1751, 1784, 1831, 1891, 1931, 1985, 1996, 2049, 2052, 2055, 2083, 2084, 2140, 2162, 2211, 2282, 2292, 2313, 2314, 2337, 2361, 2369, 2377, 2458, 2471, 2486, 2541, 2565, 2603, 2649, 2666, 2759, 2786, 2795, 2903, 2947, 2948, 2959, 2973, 3026, 3049, 3051, 3054, 3087, 3109, 3125, 3169, 3190, 3193, 3216, 3244, 3320, 3349, 3425, 3472, 3545, 3563]. aware [391, 400, 474, 520, 619, 1037, 1270, 1321, 1442, 1533, 1548, 1619, 1733, 1770, 1805, 1834, 1847, 1855, 1892, 1915, 1970, 1981, 2056, 2057, 2146, 2149, 2204, 2250, 2317, 2349, 2375, 2396, 2464, 2485, 2508, 2562, 2585, 2619, 2681, 2683, 2697, 2725, 2799, 2867, 2884, 2907, 2967, 3149, 3285, 3362, 3387, 3428]. Awareness [296, 550, 762, 766, 940, 1396, 2289, 2393, 2520, 3289]. awareness-control [2520]. AWG [2364]. AXML [1122].

B [740]. B5G [3005, 3175, 3180, 3283, 3544]. B5G/6G [3544]. Ba [161]. Babies [314]. BAC [3376]. BAC-CRL [3376]. back [992, 2778]. back-haul [992]. back-projection [2778]. backbone [859, 1090, 1387, 1808, 2013, 2531]. backdoor [3454]. backend [202]. Background [2960]. backhaul [2883, 3010]. backhauls [591]. backoff [691]. Backscatter [3235]. backup [671, 1027]. bad [3224]. Baguette [2928]. balance [1881, 2651]. Balanced [336, 786, 977, 1112, 1586, 1750, 1833, 2180, 2198, 2569, 3278, 3485]. Balancer [1484]. Balancing [327, 383, 575, 630, 770, 779, 966, 1033, 1138, 1221, 1443, 1653, 1705, 1736, 1761, 1784, 1915, 1920, 1943, 1977, 1980, 1982, 2081, 2255, 2272, 2416, 2876, 3005, 3358, 3410, 3425]. band [2047, 2851]. Bandit [3558]. Bandits [2402]. Bandwidth [438, 528, 619, 808, 866, 873, 895, 919, 963, 1079, 1121, 1146, 1149, 1154, 1157, 1187, 1252, 1591, 1650, 1711, 1937, 2159, 2243, 2735, 3074, 3160, 3530]. Bandwidth-adaptive [1121]. bandwidth-efficient [3530]. BankSearch [209]. banyan [768]. Banzhaf [1632]. bar [125]. BARC [476]. bargain [3248]. bargaining [1114, 1167, 2223, 2820]. barrier [1496]. barriers [1896]. BART [808]. base [102, 522, 2228, 2276, 2320, 2372, 2509, 2561, 2794, 2881, 3055, 3421]. Based [6, 12, 90, 217, 244, 262, 287, 308, 341, 446, 478, 524, 542, 644, 666, 740, 769, 853, 872, 886, 943, 962, 964, 973, 980, 991, 1019, 1021, 1044, 1092, 1145, 1156, 1167, 1170, 1190, 1214, 1224, 1242, 1354, 1464, 1622, 1710, 1722, 1790, 1799, 1807, 1846, 1870, 1885, 1901, 1964, 1979, 1985, 1996, 1997, 2026, 2032, 2042, 2070, 2159, 2162, 2199, 2218, 2222, 2264, 2271, 2281, 2314, 2380, 2395, 2416, 2430, 2503, 2566, 2576, 2599, 2604, 2632, 2732, 2791, 2818, 2821, 2866, 2918, 3038, 3060, 3097, 3112, 3115, 3134, 3157, 3190, 3263, 3283, 3294, 3335, 3390, 3462, 3561]. based [94, 107, 113, 155, 178, 251, 268, 271, 276, 295, 326, 343, 378, 385, 392, 406, 415, 494, 545, 667, 705, 724, 776, 781, 854, 860, 892, 916, 956, 979, 1017, 1040, 1058, 1104, 1110, 1118, 1129, 1142, 1157, 1200, 1285, 1307, 1312, 1358, 1482, 1484, 1519, 1536, 1545, 1552, 1623, 1661, 1676, 1679, 1700, 1709, 1824, 1922, 2107, 2120, 2132, 2142, 2154, 2184, 2192, 2200, 2229, 2253, 2255, 2297, 2364, 2401, 2446, 2515, 2533, 2547, 2550, 2639, 2781, 2783, 2796, 2803, 2840, 2957, 3010, 3033, 3041, 3045, 3069, 3078, 3124, 3210, 3231, 3274, 3319, 3338, 3351, 3512, 3523, 3537]. based [44, 80, 177, 200, 229, 250, 294, 315, 353, 381, 382, 401, 411, 427, 443, 448, 479, 511, 544, 581, 595, 663, 664, 693, 713, 818, 825, 832, 921, 925, 987, 1087, 1201, 1256, 1293, 1297, 1301, 1349, 1363, 1420, 1425, 1476, 1512, 1525, 1527, 1532, 1575, 1616, 1631, 1648, 1668, 1762, 1810, 1820, 1929, 1931, 1938, 1939, 1945, 1971, 2020, 2022, 2099, 2175, 2209, 2236, 2252, 2292, 2332, 2338, 2348, 2369, 2370, 2384, 2450, 2452, 2540, 2600, 2644, 2733, 2751, 2778, 2836, 2843, 2916, 2975, 3028, 3050, 3054, 3087, 3202, 3216, 3267, 3284, 3295, 3336, 3357, 3369, 3396, 3403, 3425, 3453, 3459, 3465, 3492, 3495]. based [138, 187, 196, 266, 290, 293, 320, 349, 354, 361, 366, 393, 403, 424, 460, 483, 504– 506, 519, 554, 560, 585, 639, 668, 672, 692, 701, 703, 715, 730, 748, 750, 772, 811, 828, 856, 997, 1002, 1024, 1027, 1028, 1041, 1056, 1120, 1231, 1337, 1371, 1438, 1454, 1546, 1639, 1642, 1697, 1717, 1796, 1812, 1825, 1861, 1906, 1966, 1993, 2031, 2053, 2062, 2085, 2141, 2171, 2195, 2220, 2273, 2282, 2328, 2415, 2559, 2567, 2622, 2651, 2713, 2723, 2764, 2765, 2775, 2786, 2838, 2860, 2877, 2884, 2910, 2922, 2941, 2956, 2981, 3062, 3063, 3129, 3130, 3168, 3196, 3243, 3245, 3258, 3296, 3306, 3350, 3354, 3370, 3371, 3470, 3525, 3533, 3540]. based [29, 93, 128, 156, 273, 279, 307, 364, 507, 555, 583, 626, 645, 707, 714, 716, 760, 767, 774, 824, 898, 938, 971, 984, 1010, 1051, 1059, 1079, 1080, 1106, 1115, 1150, 1152, 1155, 1253, 1296, 1336, 1351, 1375, 1387, 1466, 1494, 1640, 1727, 1808, 1840, 1864, 1865, 1908, 1927, 1928, 1989, 2021, 2048, 2148, 2156, 2190, 2213, 2224, 2240, 2278, 2317, 2440, 2474, 2523, 2556, 2611, 2650, 2659, 2668, 2683, 2725, 2741, 2809, 2849, 2899, 2927, 2931, 2934, 2953, 3017, 3071, 3119, 3167, 3187, 3194, 3228, 3234, 3242, 3268, 3302, 3304, 3332, 3337, 3386, 3437, 3496, 3497, 3524, 3550, 3564]. based [20, 70, 121, 174, 180, 208, 212, 259, 265, 288, 311, 348, 493, 517, 529, 609, 620, 646, 685, 756, 773, 777, 806, 901, 926, 986, 1022, 1148, 1153, 1162, 1180, 1266, 1286, 1310, 1313, 1341, 1421, 1501, 1625, 1638, 1643, 1651, 1742, 1743, 1802, 1804, 1815, 1828, 1943, 1970, 1981, 2002, 2009, 2067, 2072, 2091, 2112, 2115, 2180, 2181, 2230, 2234, 2275, 2387, 2396, 2398, 2423, 2461, 2475, 2543, 2585, 2587, 2688, 2722, 2784, 2799, 2810, 2820, 2825, 2911, 2929, 2960, 2984, 3093, 3096, 3120, 3138, 3172, 3222, 3235, 3261, 3269, 3271, 3279, 3308, 3314, 3361, 3382, 3399, 3404, 3434, 3440, 3510]. based [69, 164, 171, 179, 189, 197, 242, 305, 367, 369, 516, 525, 623, 710, 719, 739, 755, 795, 812, 906, 960, 999, 1000, 1012, 1020, 1063, 1064, 1074, 1088, 1098, 1116, 1149, 1187, 1211, 1217, 1235, 1251, 1274, 1379, 1380, 1433, 1436, 1439, 1456, 1462, 1481, 1534, 1535, 1538, 1543, 1544, 1581, 1585, 1602, 1620, 1641, 1712, 1733, 1871, 1898, 1953–1956, 1963, 1977, 2097, 2138, 2185, 2210, 2279, 2284, 2305, 2330, 2341, 2407, 2413, 2468, 2479, 2638, 2653, 2665, 2674, 2717, 2755, 2777, 2824, 2908, 2940, 3007, 3029, 3067, 3145, 3191, 3198, 3215, 3227, 3253, 3286, 3303, 3305, 3309, 3374, 3377, 3389, 3397, 3426, 3469, 3509, 3543]. based [97, 338, 809, 881, 914, 954, 1327, 1487, 1540, 1748, 1775, 2057, 2198, 2634, 2827, 3002, 3009, 3233, 3391, 3451]. bases [10]. basic [2758]. Basics [1427]. batch [494, 957]. batch-based [494]. Batches [3132]. batching [139]. bats [2848]. BatTalk [2848]. Battery [476]. Bayesian [436, 1656, 1840, 1893, 2263]. BBWS [3235]. BCGs [1100]. BcN [384]. BCON [2604]. BDI [226, 1923]. BDI2DoS [1923]. BDP [1273]. beacon [1274, 2450, 2750, 3561]. Beacon-based [2450]. beaconing [1354]. beacons [1283]. bearer [3011]. beat [3268]. Bedside [22]. Bee [694, 1019, 1263, 1645]. Behavior [682, 1193, 1463, 1955, 2115, 2234, 2459, 2461, 2865, 3271, 3298]. Behavior-based [2115]. behavioral [2094, 2278, 2799]. behaviors [1596, 3085]. behaviour [2645]. behaviours [652]. Belief [1266, 2724]. Benchmarking [793, 2578]. benchmarks [53]. benefit [1195]. Benefits [958, 1809]. Benford [1193]. benign [2511]. Best [444, 930, 1659, 2058, 2538, 3364]. Beta [1540]. Beta-based [1540]. Better [526, 1525, 2920]. between [283, 608, 1073, 1078, 1085, 1523, 1965, 2441, 2578, 2763, 3374]. betweenness [854, 1849]. Beyond [771, 1887, 2487, 2710, 2769, 3008, 3150, 3201]. BGP [1367]. BHF [1008]. Bi [316, 2827, 3438]. Bi-level [2827]. Bi-modal [316]. bi-objective [3438]. bias [3440, 3564]. bibliographic [32]. bibliography [941]. Bibliometric [1785, 2763, 3219]. bidding [672]. Bidirectional [1049, 1150, 1306, 3441, 3511]. BIER [2809]. BIER-FRR [2809]. bifurcation [742]. Big [1564, 1565, 1569, 1675, 1721, 1740, 1937, 2069, 2302, 2327, 2365, 2396, 2440, 2544, 2626, 2772, 2840, 3097, 3099]. BilMesh [1166]. binary [1243, 2230]. binding [737, 957]. bio [532, 614, 922, 1107, 1130, 1993, 2261, 2321]. bio-cryptosystems [2261]. bio-inspired [532, 614, 922, 1107, 1130, 2321]. bio-signal [1993]. Biocode [505]. biodiversity [37]. Biofeedback [567]. biofilm [17]. biological [4]. biomedical [1199, 1985]. biometric [509, 517, 1286, 1438]. biometric-based [517]. biometrics [483, 503, 507, 560]. biometrics-based [483, 560]. birds [2094]. bistatic [1496]. Bit [506, 947, 1159, 1727, 2400, 3359]. Bit-level [1727]. bit-shuffling [3359]. bit-strings [506]. bitcoin [2915, 2950]. BitDew [458]. BitTorrent [584, 1159, 2885]. black [2205, 2717]. black-box [2205, 2717]. blacklist [1152]. blacklist-based [1152]. BLE [2750]. blind [59, 564, 714, 3359, 3449]. Blink [153]. bloatware [2689]. block [930, 1289, 1371, 1508, 1783, 1841]. block-aware [1289]. BlockAudit [2570]. Blockchain [2138, 2273, 2385, 2391, 2483, 2556, 2560, 2562, 2604, 2622, 2623, 2628, 2732, 2745, 2769, 2775, 2783, 2818, 2839, 2845, 2856, 2857, 2897, 2898, 2910, 2911, 2915, 2916, 2930, 2943, 2952, 2956, 2960, 2966, 2967, 2984, 3028, 3068, 3078, 3079, 3097, 3115, 3120, 3151, 3154, 3190, 3204, 3225, 3227, 3242, 3245, 3247, 3253, 3261, 3274, 3296, 3315, 3323, 3328, 3335, 3347, 3355, 3374, 3376, 3383, 3396, 3401, 3413, 3435, 3441, 3451, 3462, 3465, 3475, 3484, 3494, 3495, 3498, 3567]. Blockchain-Assisted [3376, 3475]. Blockchain-based [2138, 2273, 2556, 2622, 2732, 2775, 2818, 2984, 3028, 3078, 3097, 3190, 3227, 3253, 3261, 3296, 3335, 3396, 3462]. Blockchain-driven [3465]. blockchain-enabled [3154, 3413]. blockchain-IoT [3115]. blockchains [2403, 2899, 3352]. blocking [994, 1523]. blocks [809, 1132, 2441]. BlockSecIoTNet [2556]. blog [257, 2827]. bloom [988, 1120, 1148, 1649, 1963, 2004, 2496, 2989]. Bloom-filter [2989]. bluetooth [280, 340, 357, 386, 395, 621, 671, 801, 1651, 1787, 3223, 3292]. Bluetree [497]. blurred [2778]. Board [223, 232, 240, 263, 284, 309, 324, 330, 339, 347, 380, 398, 416, 441, 455, 471, 482, 490, 502, 518, 533, 543, 552, 589, 627, 649, 687, 717, 743, 788, 820, 849, 869, 893, 931, 981, 1015, 1030, 1046, 1060, 1093, 1125, 1144, 1176, 1206, 1249, 1267, 1279, 1298, 1316, 1350, 1364, 1374, 1384, 1402, 1413, 1428, 1444, 1452, 1468, 1485, 1502, 1530, 1570, 1594, 1609, 1626, 1646, 1660, 1677, 1688, 1704, 1716, 1729, 1744, 1757, 2147, 2158, 2166, 2177, 2187, 2197, 2206, 2217, 2227, 2237, 2248, 2259]. board [1215, 1767, 1780, 1806, 1816, 1827, 1851, 1857, 1877, 1889, 1902, 1921, 1932, 1947, 1958, 1976, 1986, 1999, 2006, 2016, 2025, 2043, 2050, 2059, 2065, 2077, 2088, 2101, 2110, 2118, 2119, 2125, 2268, 2277, 2286, 2296, 2306, 2315, 2326, 2340, 2352, 2366, 2386, 2399, 2410, 2420, 2428, 2437, 2447, 2457, 2467, 2477, 2488, 2499, 2510, 2517, 2524, 2530, 2537, 2548, 2558, 2573, 2574, 2594, 2606, 2617, 2630, 2641, 2648, 2655, 2663, 2671, 2678, 2686, 2694, 2701, 2708, 2716, 2727, 2738, 2747, 2757, 2768, 2780, 2793, 2805, 2817, 3271]. Board [2829, 2842, 2854, 2873, 2890, 2901, 2913, 2925, 2937, 2945, 2955, 2963, 2970, 2979, 2986, 2993, 2999, 3004, 3013, 3024, 3036, 3047, 3057, 3065, 3073, 3082, 3090, 3101, 3111, 3123, 3136, 3148, 3164, 3177, 3200, 3212, 3230, 3240, 3252, 3257, 3266, 3282, 3301, 3312, 3326, 3341, 3366, 3380, 3393, 3406, 3432, 3447, 3461, 3480, 3491, 3507, 3516, 3521, 3542]. boards [1334]. Body [316, 567, 2010, 2064, 2095, 2176, 2351, 2557, 2744]. BOFUSS [2758]. Boltzmann [2484]. Book [3291]. boosting [247]. Booting [1041]. bootstrap [442]. border [172, 1132]. Borel [1100]. Bot [2688]. BotFlex [1513]. both [2094]. botnet [1129, 1513, 1807, 2632]. botnet-based [1129]. botnets [3126]. Bottleneck [381, 998, 1041, 1631, 2008]. bound [360, 394, 2062, 3463]. Boundary [1368, 2782, 3400, 3428]. bounded [464, 1003, 2368]. bounding [636]. Bounds [352, 971, 1226]. box [2205, 2717]. BPF [3094]. braided [525]. BRAINS [257]. branch [833]. branching [3563]. bribery [1189]. bridge [3415]. BRIP [3494]. BRNN [2634]. broadband [147, 314, 444, 725, 870, 871, 1036, 1244]. Broadcast [602, 848, 946, 1055, 1356, 1505, 1641, 2446, 3105, 3544]. Broadcast/multicast [3544]. Broadcasting [721, 898, 1039, 1326, 1699]. broader [1613]. brokering [216, 1167, 1784, 2055]. brokers [202, 3153, 3306]. bronchial [2938]. browser [1398, 1685]. browsing [40, 1041, 2091, 2221, 2278]. BSeIn [2273]. BSNs [2072]. BTRES [1540]. budget [1650, 2146, 2555]. Buffer [233, 2491, 3424]. buffered [188, 192, 994, 1287]. buffering [2014]. buffers [2433]. build [184]. Building [190, 199, 231, 283, 662, 916, 2283, 2295, 2827]. buildings [2504, 3302]. built [2092]. Burr [3367]. burst [930]. burstiness [3039]. bursty [963]. bus [16, 1643]. Bush [314]. Business [642, 643, 646, 1067, 1105, 1547]. Business-aware [643]. Buyer [1801]. bypassing [1191]. Byzantine [1372]. Byzantine-resilient [1372].

C [1378, 2320, 2521, 2728, 3482]. C-RAN [2320, 2521, 2728, 3482]. Cache [499, 950, 1116, 1395, 1489, 1685, 1690, 1979, 2004, 2042, 2156, 2219, 2317, 2764, 2902, 3077]. cache-aided [2219]. cache-aware [2317]. Cache-based [1979, 2042]. cache-enabled [3077]. caches [21, 1106]. Caching [261, 604, 781, 795, 815, 1118, 1184, 1250, 1473, 1709, 1938, 1967, 1983, 2285, 2431, 2495, 2619, 2633, 2696, 2795, 2850, 2920, 2922, 2958, 3032, 3051, 3077, 3129, 3209, 3332, 3376, 3501, 3531]. CADeComp [341]. calculating [21]. calculation [1208]. Calculus [915, 2062, 2066]. Caledonian [3145]. Calendar [175]. Calibrating [1288]. calibration [2489, 3348]. call [86, 1229, 1648, 2791]. Callimachus [174]. calling [191, 1542]. calls [3370]. CALM [2371]. cameras [1334, 2577]. campus [760, 929, 1272]. CAN [201, 375, 478, 670, 1037, 3015]. CAN-based [478]. canals [2965]. Cancelable [506, 737, 2261]. candidate [722]. capabilities [328, 1539, 2773]. Capability [311]. Capability-based [311]. capacities [1469, 3533]. Capacity [556, 559, 1349, 1395, 1631, 2014, 2371, 2761]. capacity-aware [1395]. capital [1404]. capsule [381]. capsule-based [381]. capture [1063, 2642]. Capturing [3292, 3403]. CAQ [2240]. Car [1051, 1285, 3115]. Car-sharing [3115]. car-to-car [1051]. card [1056]. Cardiac [2178]. Cardinality [1911, 3421]. Cards [219, 225, 483, 560, 811, 856]. Cards-to-presentation [225]. Care [314, 362, 645, 2264, 2266, 2281, 2309, 2476, 2591]. CARED [819]. Carlo [326, 1878, 1968]. carrier [624, 1360, 1630, 1651, 2666]. CARSA [716]. cascade [1950]. cascading [1185]. case [37, 251, 759, 857, 873, 1391, 1476, 1477, 1726, 2525, 3066, 3115, 3322]. cases [805, 2498, 2504, 3384]. Cataclysm [376]. catalogues [138]. categorization [2785]. Categorizing [64]. category [209, 1029]. CATS [1051]. Cautious [819]. Cayley [1100]. CBDI [1632]. CDA [2830]. CDMA [529]. CDMA/TDMA [529]. CDN [2720]. CDN-P2P [2720]. CDS [1833]. CDT [3530]. CE [2850, 3412]. CE-D2D [2850]. CE-VNE [3412]. Cell [229, 2375, 2508, 2810, 2862, 3183]. cells [2935]. Cellular [740, 773, 1215, 1258, 1407, 1424, 1507, 1737, 1766, 1941, 2098, 2114, 2167, 2228, 2256, 2276, 2411, 2521, 2561, 2676, 3112, 3208, 3224, 3331, 3569]. cellular-assisted [2676]. cellular/802.11 [1407]. Center [1538, 1593, 1619, 1694, 1698, 1834, 1907, 2045, 2057, 2246, 2251, 2444, 2572, 2649, 2674, 2800, 2894, 2917, 3039, 3193, 3333, 3350, 3499, 3504]. centered [76, 2941]. Centers [654, 1400, 1415, 1567, 1831, 1832, 1915, 1916, 1934, 1968, 2083, 2239, 3254]. Central [3409]. centrality [854, 3492]. Centralized [1912]. centrally [2411]. centrally-coordinated [2411]. centre [641]. Centric [171, 187, 404, 537, 547, 850, 855, 1116, 1186, 1227, 1358, 1390, 1391, 1394, 1395, 1397, 1414, 1473, 1488, 1489, 1522, 1602, 1625, 1637, 1639, 1649, 1876, 1899, 2109, 2219, 2224, 2271, 2317, 2413, 2466, 2622, 2691, 2696, 2753, 2989, 3166, 3209, 3248, 3360]. centroid [750]. Certificate [1054, 1606, 1865, 3279, 3376]. certificateless [2176]. Certification [767]. Certification-based [767]. certified [2443, 2719]. CGSNet [3556]. chain [351, 642, 835, 1070, 1841, 1865, 2021, 2038, 2431, 2843, 2911, 3361, 3469]. chain-based [1865]. chaining [1791, 2198, 2588, 2609, 3086, 3549]. chains [1606, 2704, 2743]. Challenge [769, 2481]. challenged [1882, 3058]. Challenges [648, 669, 846, 911, 932, 1175, 1290, 1385, 1386, 1391, 1414, 1418, 1426, 1427, 1578, 1597, 1599, 1611, 1674, 1682, 1684, 1697, 1701, 1703, 1732, 1739, 1798, 1801, 1852, 1856, 1913, 1925, 1961, 2010, 2054, 2063, 2075, 2109, 2186, 2208, 2244, 2265, 2385, 2397, 2483, 2504, 2532, 2551, 2560, 2571, 2657, 2695, 2712, 2745, 2771, 2798, 2875, 2891, 2893, 2898, 2916, 2923, 2930, 2947, 2956, 2996, 3010, 3029, 3079, 3107, 3113, 3152, 3170, 3242, 3262, 3310, 3315, 3323, 3324, 3329, 3331, 3337, 3346, 3347, 3384, 3395, 3441, 3457, 3458, 3567]. chameleon [554]. change [1029, 1080]. changes [698, 926, 1029, 1478, 3224]. Channel [450, 556, 564, 592, 611, 636, 728, 787, 825, 864, 876, 885, 888, 1005, 1023, 1048, 1096, 1136, 1141, 1149, 1150, 1166, 1181, 1339, 1361, 1470, 1635, 1654, 1733, 1753, 1773, 1866, 2042, 2108, 2127, 2151, 2226, 2240, 2295, 2371, 2389, 2392, 2450, 2590, 2639, 2703, 2950, 3059, 3253, 3368]. channel-availability [2240]. Channel-aware [1733]. channel-quality [2240]. Channel-reserved [2639]. channels [2060, 2115, 2926, 3511]. Chaos [5]. chaotic [1117, 2463, 2611, 3425]. Characteristic [517]. Characteristics [12, 334, 765, 860, 1254, 1357, 1726]. characterization [1717, 3298]. Characterizing [435, 1685, 2586]. charger [2258]. chargers [3296]. Charging [2646, 2679, 2882, 3495, 3546]. charts [346]. chassis [366]. check [548, 2192]. check-in [2192]. checkability [2293]. checking [686, 1054, 2474, 2588, 3043]. checkpointing [3231]. Chief [106]. Chimera [131]. Chinese [2421]. chip [2162, 3116, 3390]. choice [304, 598]. choices [97]. Chord [423]. Chordal [70]. chronic [2264]. chronometry [2345]. chunks [3501]. Churn [372, 575]. Churn-resilient [372]. Cinema [397]. CIoT [2638]. ciphers [1508]. circuits [2188, 3018]. CIT [1936]. cities [1404, 2316, 2406, 2472, 2484, 2897, 2956, 3173, 3260, 3523]. city [2479, 2486, 2500, 2576, 3382]. CL [3565]. CL-AP [3565]. claimer [1616]. class [9, 74, 411, 994, 2184, 3564]. class-based [411]. classic [860]. Classical [881, 2321]. Classification [318, 393, 401, 463, 708, 812, 834, 987, 1590, 1663, 1812, 1929, 2084, 2117, 2133, 2174, 2234, 2381, 2415, 2551, 2572, 2599, 2657, 2669, 2702, 2777, 2800, 2869, 2971, 3143, 3263, 3348, 3547]. classified [941]. classifier [44, 247, 401, 3267]. classifiers [620]. classifying [127, 2133]. classroom [101, 540]. clear [1607]. click [2865]. click-requests [2865]. clicks [1000]. Client [235, 251, 313, 492, 597, 680, 890, 1227, 1278, 2730, 3459]. Client-server [313]. client-side [1278, 2730]. Client/server [492, 680]. clinical [2280]. clique [522, 724]. clique-based [724]. clock [1147, 1284, 2478]. clocks [2165]. clone [1608, 1616, 3087]. Cloning [967]. close [1965]. closed [133]. closed-system [133]. closer [1636]. Cloud [657, 780, 910, 970, 1091, 1197, 1229, 1275, 1306, 1307, 1309, 1355, 1359, 1415, 1418, 1426, 1465, 1536, 1550, 1583, 1584, 1621, 1623, 1648, 1662, 1676, 1684, 1715, 1722, 1731, 1749, 1776, 1821, 1831, 1842, 1856, 1875, 1897, 1922, 1980, 1996, 2012, 2049, 2055, 2069, 2071, 2079, 2083, 2099, 2109, 2154, 2160, 2172, 2176, 2196, 2211, 2215, 2239, 2283, 2293, 2303, 2313, 2314, 2345, 2416, 2419, 2430, 2458, 2463, 2470, 2503, 2521, 2545, 2603, 2635, 2644, 2774, 2788, 2836, 2863, 2868, 2946, 3019, 3038, 3078, 3162, 3166, 3193, 3216, 3233, 3313, 3372, 3425, 3430, 3467, 3489, 3492, 3505, 3508, 3536]. cloud [553, 650, 655, 659, 827, 908, 934, 952, 1003, 1089, 1180, 1203, 1241, 1277, 1322, 1341, 1373, 1382, 1416, 1460, 1494, 1515, 1548, 1549, 1568, 1574, 1578, 1628, 1634, 1644, 1656, 1665, 1667, 1687, 1701, 1736, 1772, 1774, 1784, 1795, 1796, 1818, 1822, 1826, 1828, 1832, 1834, 1861, 1880, 1894, 1925, 1934, 1935, 1949, 2005, 2037, 2081, 2121, 2129, 2190, 2207, 2212, 2214, 2233, 2265, 2351, 2396, 2404, 2423, 2522, 2549, 2555, 2583, 2621, 2638, 2645, 2651, 2711, 2712, 2714, 2787, 2822, 2826, 2838, 2849, 2888, 2948, 2951, 3032, 3037, 3096, 3153, 3182, 3196, 3237, 3306, 3339, 3359, 3374, 3378, 3410, 3416, 3438, 3499, 3540]. cloud [653, 867, 935, 1033, 1083, 1194, 1236, 1278, 1302, 1388, 1533, 1538, 1547, 1629, 1775, 1871, 1936, 1953, 2137, 2138, 2143, 2173, 2305, 2327, 2363, 2426, 2627, 2706, 2752, 2810, 2907, 2909, 3097, 3145, 3334, 3522, 3531]. Cloud-aided [2176]. cloud-assisted [2351]. Cloud-based [1796, 1996, 2314, 2396, 2638, 2836, 3038]. cloud-edge [3313, 3378]. cloud-edge-device [3334]. Cloud-RANs [2810]. CloudFile [1953]. Cloudlet [1225, 1536, 1597, 2419, 2513]. cloudlet-based [1536]. clouds [462, 1194, 1491, 1608, 1650, 1936, 2052, 2056, 2539, 3140, 3471]. Cluster [235, 269, 320, 529, 585, 883, 921, 1257, 1401, 1709, 1710, 2056, 2292, 2547, 2585, 2975]. Cluster-based [320, 529, 585, 921, 1709, 1710, 2547, 2585]. Cluster-Tap [2292]. clustered [1313, 3119]. Clustering [209, 476, 572, 620, 903, 928, 986, 1044, 1084, 1177, 1207, 1214, 1258, 1305, 1332, 1337, 1423, 1566, 1906, 1939, 1977, 2091, 2253, 2320, 2348, 2508, 2599, 2667, 2860, 3208, 3228, 3284, 3512]. Clustering-based [2599, 3512]. Clustering-oriented [2860]. clusters [327, 1216, 1920, 2697]. CMC [98, 99]. CMI [2621]. CNN [2810, 3033, 3545]. CNN-based [2810]. Co [52, 844, 868, 1259, 1557, 1608, 2150, 2827, 3278, 3321]. co-evolutionary [3278]. Co-FAIS [1259]. co-location [1608]. Co-Occurrence [2827]. co-occurring [844, 868]. co-processor [52]. co-publication [2150]. Co-scheduling [3321]. co-simulation [1557]. COACH [1054]. coal [2170]. coalescing [2773]. coalition [2223]. coalitional [3216, 3425]. COalitions [975]. CoAP [1552, 2861]. CoAP-based [1552]. code [548, 1101, 1802, 2725]. codebook [242]. Coded [1226, 1407, 3376]. codes [2871, 3510]. Coding [525, 601, 704, 735, 787, 958, 984, 1021, 1095, 1115, 1330, 1639, 1641, 1642, 3187, 3457, 3509]. coding-aware [735]. Coding-based [1021, 1639, 3187]. CoDRA [2112]. coercion [1189]. coexistence [3030, 3197]. CoFIND [104]. coflow [2482, 3075]. cognition [823]. Cognitive [527, 664, 790, 896, 1001, 1035, 1114, 1130, 1190, 1221, 1238, 1268, 1303, 1305, 1330, 1361, 1386, 1427, 1590, 1657, 1718, 1747, 1751, 1754, 1759, 1900, 1928, 1941, 1989, 2047, 2060, 2075, 2086, 2108, 2124, 2151, 2171, 2209, 2220, 2240, 2245, 2345, 2511, 2551, 2637, 2696, 2731, 2751, 2798, 2846, 2851, 2858, 3052, 3194, 3354, 3483]. COIN [1253, 2409, 2661]. COINS [975]. cold [3440]. collaborating [758, 1923]. Collaboration [116, 299, 301, 362, 839, 1061, 1065, 1077, 1078, 1534, 2896, 3334, 3459]. Collaborative [104, 107, 231, 238, 306, 313, 356, 470, 536, 667, 683, 705, 712, 729, 745, 753, 754, 756, 757, 759, 842, 1054, 1062–1064, 1066, 1070–1072, 1074– 1076, 1079, 1083, 1084, 1184, 1228, 1282, 1315, 1500, 1579, 1661, 1825, 1838, 2283, 2395, 2495, 2696, 2706, 2723, 2820, 2850, 3378]. collecting [1497, 3426]. Collection [112, 883, 1057, 1140, 1519, 2087, 2109, 2164, 2270, 2292, 2393, 2424, 2480, 2675, 2679, 2880, 2975, 3053, 3061, 3062, 3330, 3337, 3518]. collections [111]. Collective [761, 1163, 1463, 1911, 3357]. collector [2990]. Collision [631, 1378, 1647, 1859, 2040, 2353, 3555]. collision-avoiding [3555]. collision-free [2040]. collisions [1072]. collusion [1672, 2163]. collusion-resistant [1672]. collusive [435]. colonization [1520]. Colony [350, 394, 777, 918, 1171, 1182, 1263, 1627, 1645]. colored [2848]. coloring [2582]. colour [974]. combat [1923]. combating [2232]. Combination [526, 993]. combine [1745]. combined [22, 782, 908, 1632, 1711]. combiner [962]. Combining [508, 1396]. come [2188]. command [1129]. Comments [2522]. commerce [186, 187, 400, 411, 2469]. commercial [237, 633]. commercially [219]. committee [3451]. committee-based [3451]. commodity [911, 2210, 2552, 2564, 3207]. common [864]. communicating [386]. Communication [58, 71, 143, 155, 252, 302, 303, 343, 395, 422, 493, 565, 569, 594, 662, 729, 757, 823, 946, 977, 1051, 1100, 1107, 1205, 1260, 1294, 1296, 1320, 1362, 1385, 1417, 1440, 1545, 1558, 1562, 1579, 1587, 1598, 1691, 1737, 1739, 1758, 1778, 1829, 1843, 1878, 1891, 1929, 1942, 1966, 1989, 2024, 2030, 2063, 2072, 2116, 2186, 2210, 2229, 2274, 2331, 2335, 2387, 2406, 2427, 2441, 2615, 2636, 2748, 2802, 2832, 2846, 2850, 2857, 2949, 3070, 3112, 3113, 3119, 3167, 3208, 3235, 3270, 3289, 3331, 3368, 3375, 3423, 3436, 3452, 3563]. communication-based [3235]. communication-efficient [946]. communication/computation [1878]. Communications [100, 198, 293, 320, 501, 688, 821, 879, 1047, 1050, 1209, 1315, 1410, 1457, 1507, 1552, 1554, 1556, 1560, 1666, 1786, 1876, 1956, 2044, 2098, 2165, 2223, 2339, 2358, 2397, 3059, 3098, 3288, 3346, 3483, 3530]. communities [85, 107, 1183, 2560]. Community [387, 1183, 1513, 1721, 2194, 2203, 2827, 3011, 3210]. community-driven [1513]. Comparative [465, 1231, 1273, 1325, 1508, 1580, 1662, 2208]. Comparison [54, 70, 191, 193, 484, 798, 843, 881, 894, 1009, 1320, 1353, 1575, 2019, 2247, 2806, 2815, 2914]. Comparison-based [1575]. comparisons [327]. Compatibility [1927]. Compatibility-based [1927]. competitive [63, 161, 3132]. complementary [2320, 2595]. Complete [1504, 3107, 3263]. Completion [2462, 2482, 3130]. complex [629, 1198, 1537, 1849, 2646, 3007, 3040, 3140, 3210]. complexity [84]. Component [128, 174, 177, 341, 832, 1627]. component-based [128, 174, 177, 341, 832]. components [157, 1886, 3352]. composed [270]. Composing [2743]. composite [862, 1302, 2293, 2544, 3565]. composites [78]. Composition [162, 301, 358, 580, 617, 622, 716, 814, 970, 979, 1032, 1534, 1573, 1880, 1996, 2211, 2311, 2430, 2553, 2675, 2772]. Comprehensive [933, 1013, 1124, 1317, 1360, 1376, 1579, 1590, 1644, 1665, 1747, 1880, 2044, 2051, 2095, 2128, 2152, 2241, 2339, 2414, 2442, 2549, 2557, 2626, 2628, 2680, 2690, 2705, 2710, 2720, 2846, 2876, 2934, 2958, 2997, 3014, 3091, 3096, 3171, 3219, 3242, 3299, 3317, 3318, 3336, 3402, 3411, 3448, 3503, 3568]. Compressed [152, 578, 2323, 2328, 2409, 2661, 3314]. compression [558, 749, 775, 1214, 1353, 2424, 3198]. Compressive [1586, 1824, 2180, 2284, 2388, 2718, 3196, 3463]. compromise [403]. compromised [489, 2232]. Comput [868, 1412, 1765, 1766, 2339, 2516, 2587, 2661, 2888, 2987]. Computation [227, 1388, 1471, 1794, 1835, 1875, 1878, 1966, 2438, 2464, 2469, 2643, 2811, 2929, 3127, 3142, 3336, 3502]. Computational [241, 342, 715, 1355, 1664, 2049, 3117, 3487, 3489]. Computer [26, 61, 76, 100, 103, 124, 241, 321, 348, 385, 501, 760, 803, 1671, 2436, 2445, 2763, 3236]. computer-aided [124]. computer-based [760]. computers [178, 3271]. Computing [83, 169, 170, 172, 191, 246, 331, 374, 378, 440, 448, 496, 574, 650, 657, 659, 779, 780, 970, 1081, 1099, 1175, 1241, 1275, 1373, 1418, 1426, 1453, 1536, 1564, 1583, 1584, 1623, 1648, 1665, 1667, 1684, 1701, 1731, 1826, 1875, 1920, 1922, 1925, 1980, 1982, 2003, 2055, 2071, 2080, 2140, 2214, 2267, 2303, 2313, 2318, 2365, 2419, 2438, 2463, 2471, 2549, 2578, 2639, 2702, 2764, 2795, 2811, 2838, 2863, 2868, 2892, 2906, 2922, 2938, 2946, 2947, 3019, 3051, 3078, 3107, 3112, 3125, 3127, 3129, 3142, 3156, 3166, 3205, 3243, 3245, 3272, 3295, 3321, 3336, 3349, 3372, 3384, 3413, 3425, 3456, 3501, 3513, 3526]. computing [129, 141, 167, 168, 173, 408, 538, 553, 745, 821, 908, 909, 934, 1077, 1082, 1130, 1203, 1355, 1515, 1531, 1551, 1578, 1628, 1629, 1774, 1795, 2005, 2126, 2137, 2143, 2173, 2188, 2190, 2207, 2265, 2305, 2404, 2425, 2442, 2464, 2479, 2525, 2645, 2712, 2759, 2788, 2808, 2849, 2888, 2929, 2931, 2974, 2983, 2984, 2987, 3002, 3096, 3165, 3170, 3181, 3182, 3203, 3233, 3237, 3260, 3305, 3329, 3387, 3433, 3438, 3441, 3498, 3539]. computing-assisted [3539]. computing-based [3112]. computing-current [3272]. computing-enabled [3245]. computing-inspired [2938]. Concast [2379]. conceal [3307]. Concealed [2115]. concept [610, 642, 2736, 3417]. Concepts [96, 607, 663, 1271]. Conceptual [1611, 1687]. concurrent [800, 1281, 1352, 2883, 3105]. Conditional [909, 1313]. condor [3116]. conferencing [566, 962, 1292]. Confidence [2668, 2736]. confidentiality [1389]. confidentiality-preserving [1389]. Configurable [426, 1547]. Configuration [635, 697, 741, 1225, 1481, 1581, 1914]. Configurations [2642]. configure [3533]. configured [487, 3322]. confinement [634]. conflict [1862, 2604]. conflicting [439]. conflicts [1233, 2024, 3106]. Congestion [340, 422, 671, 684, 725, 941, 1050, 1111, 1113, 1237, 1280, 1304, 1360, 1422, 1436, 1441, 1464, 1469, 1511, 1692, 1793, 1878, 1907, 1944, 2032, 2040, 2057, 2159, 2444, 2471, 2650, 3319, 3332, 3343, 3350, 3411, 3538]. congestion-aware [2057, 2471]. congestion-less [2040]. conjunctions [3171]. conjunctive [576, 2412, 2776]. conjunctive-subset [576]. Connected [739, 797, 889, 1240, 1331, 1366, 1412, 1723, 1916, 2013, 2029, 3327]. connecting [898]. Connection [149, 574, 621, 872, 1208, 2334]. connection-less [621]. connectionist [10, 245]. connectionless [3223]. connections [391, 2939]. Connectivity [520, 799, 818, 926, 964, 1034, 1048, 1123, 1188, 1215, 1275, 1499, 1670, 1912, 2336, 2585, 3280]. Consecutive [2982, 3231]. consensus [858, 1243, 2882, 2966, 3245]. conservation [2224, 2446, 2650]. consideration [1850, 2894]. considerations [1909, 2325, 2592]. considering [652, 662, 1977, 2180, 2824, 3186]. consistency [1230, 2301, 2578, 3556]. consistent [2333, 3069]. consolidation [1415, 1546, 1828, 1831, 2140, 2635, 2822, 3499]. Constrained [132, 738, 778, 1040, 1221, 1278, 1287, 1291, 1296, 1376, 1628, 1670, 2031, 2146, 2290, 2523, 2861, 2904, 3549]. Constraint [363, 759, 1135, 1676, 2361, 2364, 2784, 3361, 3412]. Constraint-based [2364]. constraints [278, 328, 405, 783, 1019, 1504, 1779, 2555, 3246, 3280]. construct [173, 1401]. Constructing [10, 185, 196, 1057, 1253, 1496, 1896]. construction [280, 532, 681, 739, 930, 1756, 1787]. constructions [1972]. consumer [1278]. consumers [2055]. consumption [294, 334, 971, 1288, 1520, 1630, 2242, 2282, 2448]. Contact [1170, 2878, 3144, 3339]. container [633, 2503, 2806, 3231, 3306, 3540]. container-based [2503, 3231, 3306, 3540]. Containerized [3349, 3409, 3466]. containers [2304, 2711]. contaminant [17, 2146]. contamination [1449]. Contemporary [1454, 3042, 3113, 3126, 3384, 3398]. Content [40, 271, 295, 388, 424, 432, 452, 456, 457, 459– 462, 493, 565, 810, 923, 950, 1104, 1116, 1118, 1153, 1178, 1194, 1197, 1236, 1299, 1358, 1395, 1478, 1522, 1639, 1649, 1672, 1853, 2161, 2413, 2422, 2446, 2448, 2485, 2508, 2543, 2587, 2633, 2696, 2786, 2920, 3032, 3169, 3457, 3557, 3566]. Content-aware [923]. content-based [295, 1153, 2413, 2786]. content-centric [1358, 1522, 2413]. content-oriented [2161]. content-to-content [2448]. contention [1079, 1142, 2201, 2559]. contention-based [1079, 1142]. Contents [51, 72, 91, 117, 142, 225, 744, 812, 1085, 1097, 1403, 1437]. Context [199, 200, 296, 341, 358, 361, 535, 580, 625, 669, 716, 766, 877, 940, 1037, 1062, 1271, 1282, 1396, 1457, 1663, 1676, 1839, 1985, 2112, 2289, 2316, 2393, 2435, 2486, 2523, 2799, 2947, 3026, 3087, 3137, 3149, 3185, 3218, 3524]. Context- [535]. Context-aware [199, 200, 341, 716, 1037, 1985, 2486, 2799, 2947, 3026, 3087, 3149]. context-awareness [296]. Context-based [2112]. Context-oriented [3137]. contexts [1565]. Contextual [3558]. continued [744]. continuity [1229]. Continuous [840, 1433, 2200, 2880, 3029, 3437, 3456]. continuum [1499]. contour [900]. Contra [2943]. Contra-* [2943]. contract [2665, 2930, 3215, 3503]. contract-based [3215]. Contracts [2818, 2908, 2916, 2977, 3335]. Control [4, 191, 222, 229, 235, 279, 287, 290, 384, 388, 419, 422, 478, 492, 555, 612, 663, 677, 682, 693, 718, 723, 724, 726, 742, 791, 797, 807, 813, 854, 872, 917, 941, 978, 997, 1039, 1111, 1129, 1151, 1165, 1238, 1246, 1280, 1291, 1304, 1343, 1408, 1422, 1435, 1458, 1464, 1469, 1511, 1548, 1555, 1559, 1765, 1820, 1872, 1985, 2159, 2196, 2225, 2233, 2273, 2276, 2332, 2338, 2394, 2444, 2446, 2455, 2460, 2466, 2520, 2523, 2535, 2563, 2622, 2639, 2650, 2714, 2749, 2786, 2883, 2930, 2965, 2988, 3074, 3108, 3151, 3161, 3169, 3174, 3220, 3274, 3319, 3332, 3343, 3350, 3382, 3402, 3411, 3413, 3424]. control [311, 340, 618, 623, 671, 678, 684, 689, 909, 1064, 1113, 1217, 1237, 1284, 1352, 1419, 1436, 1692, 1775, 1786, 1953, 2087, 2112, 2137, 2149, 2305, 2349, 2468, 2604, 3077, 3226, 3227, 3538]. controllable [1365]. controlled [425, 697, 2412]. Controller [42, 66, 623, 1962, 2027, 2928, 2972, 3278, 3302, 3371, 3485, 3487, 3532, 3535]. controllers [2131, 2333]. controlling [80]. CONUGA [132]. convenience [3394]. convergecast [1336]. converged [586]. convergence [542, 853, 2898, 2923, 3347, 3481]. converging [3343]. conversations [144]. conversion [20, 1097]. convertible [713]. convex [1387]. conveying [3423]. convolution [3007, 3286]. Convolutional [2869, 3092, 3389, 3403, 3412]. Cooperation [173, 299, 305, 307, 557, 658, 806, 1451, 1685, 2209, 2219, 2863, 2888, 3181]. Cooperation-based [2209]. Cooperative [121, 197, 604, 672, 751, 925, 1021, 1080, 1082, 1167, 1259, 1263, 1340, 1348, 1349, 1407, 1518, 1521, 1633, 1709, 1941, 2009, 2020, 2047, 2097, 2285, 2387, 2460, 2493, 2528, 2619, 2904, 3005, 3075, 3165, 3302, 3327, 3344, 3377, 3492]. Cooperative-centrality [3492]. cooperative-communication-enabled [2387]. coordinate [1313]. Coordinated [555, 611, 668, 1258, 1706, 2167, 2203, 2361, 2411, 3423]. coordinating [1088, 2830]. Coordination [123, 806, 915, 1501, 2040, 2305, 2728, 3420]. copula [2401]. copula-based [2401]. Copy [47, 1509, 1798]. Core [738, 1195, 1429, 1788, 1886, 2162, 2358, 2417, 2531, 3316, 3381]. core-edge [1195]. CORM [803]. Corners [2249]. Corona [1640]. correcting [599]. correction [1113, 2462]. correlated [874, 1177]. correlation [470, 507, 904, 1585, 2067]. correlations [3403]. Corrigendum [868, 1412, 1765, 1766, 2587, 2661, 2888, 2987]. cosine [973]. Cost [509, 546, 598, 683, 1132, 1442, 1459, 1494, 1548, 1560, 1601, 1616, 1658, 1782, 1784, 1796, 1832, 1881, 1901, 2052, 2250, 2272, 2425, 2441, 2513, 2529, 2718, 2784, 2806, 2928, 2948, 2972, 3109, 3116, 3124, 3254, 3313, 3410, 3467, 3531]. Cost-aware [1442, 1784, 2052, 2948]. Cost-effective [2425, 2928, 2972, 3254]. cost-efficient [1658, 1782, 1901, 3124]. Cost-optimal [2513]. Cost EACP [2972]. costs [2689]. COTS [2781]. counter [2567]. Counteracting [1180]. Countering [2144, 2449, 2943]. countermeasures [822, 1697, 2113, 2887, 3014]. countermining [435]. counters [1378]. counting [2076]. coupled [570, 606]. Coupling [2479]. course [97, 979]. course-support [97]. courses [94]. Coverage [449, 520, 715, 791, 797, 799, 955, 1002, 1162, 1168, 1328, 1331, 1365, 1368, 1412, 1496, 1499, 1504, 1585, 1750, 1781, 1782, 1864, 1942, 2356, 2585, 2602, 2949, 3280]. Coverage-aware [520]. covering [639, 2031]. covert [2115, 2226, 2295, 2389, 3253, 3454]. COVID [2886, 2941, 3382]. COVID-19 [2886, 2941, 3382]. CPL [282]. CPS [236, 3362]. CPU [656]. CR [2075]. CR-IWSN [2075]. CRAMP [3512]. CRAN [2710]. crawler [207]. CRBAC [405]. CRC [1210, 3291]. Creating [36, 163, 754]. creative [47]. Credential [960, 1253, 1810, 3190]. credential-based [1253]. Credibility [1341]. Credibility-based [1341]. CREME [3053]. Cress [270]. CRIMP [2584]. crisis [2397, 2584]. criteria [247]. criterion [2406, 2822]. critic [3319]. Critical [99, 530, 578, 626, 669, 847, 942, 1123, 1355, 1499, 1624, 1632, 1918, 2582, 2988, 3027, 3208]. criticality [1799, 3167]. CRL [3376]. CRNs [2851]. CROP [2151]. Cross [342, 415, 445, 600, 624, 727, 733, 871, 1280, 1343, 1348, 1521, 1630, 1765, 2018, 2042, 2086, 2117, 2372, 2550, 2569, 2624, 2683, 2843, 3027, 3243, 3362, 3415, 3530, 3556]. cross-carrier-aware [1630]. Cross-consistency [3556]. cross-domain [2117, 2843]. Cross-interface [3362, 3530]. Cross-Layer [342, 415, 445, 600, 624, 727, 733, 1280, 1343, 1348, 1521, 1765, 2018, 2372, 2569, 3027, 3243]. cross-layered [2683]. cross-rollup [3415]. Cross-VM [2042]. Crossed [2013]. crosses [272]. crossroads [647]. Crow [3145, 3513]. Crowd [413, 1266, 1420, 1541, 1748, 1824, 2105, 2432, 2475, 2821, 3063, 3174]. Crowd-based [2821]. crowd-sensing [1824, 2105]. crowd-sourcing [1420]. crowds [1395, 1804]. crowdsensing [1420, 1901, 2335, 2490, 2918, 3114, 3215, 3217, 3320, 3394, 3449, 3468]. crowdsourced [2887]. crowdsourcing [1631, 1642, 1669, 2350, 2911, 3028, 3294, 3377, 3435, 3495]. crowdsourcing-based [1642]. CRT [392]. Crypt [3237]. Crypt-OR [3237]. Cryptanalysis [560]. crypto [2966]. crypto-currencies [2966]. Cryptocurrencies [2745]. cryptocurrency [3016]. Cryptographic [1224, 1376, 2403]. cryptography [243, 825, 2443, 2719]. cryptosystem [49, 504, 737]. cryptosystems [196, 2261]. CSO [1830]. CTC [2830]. CTF [2642]. CTRA [2646]. Cube [2013]. cubes [451]. cubic [3333]. CUR [3463]. currencies [2966]. currency [3259]. Current [93, 895, 2017, 2041, 2773, 2898, 2933, 3272]. curriculum [3424]. cursor [222]. Curve [698]. customer [1031]. Customizable [2431, 2536]. customization [780, 2034, 2687]. customizing [2376]. cuts [2405]. cutting [1074]. CVE [160]. Cyber [1624, 1870, 2144, 2254, 2449, 2456, 2465, 2516, 3099, 3143, 3246, 3290]. Cyber-ARGUS [2465]. cyber-attacks [1870]. cyber-physical [3246]. cyber-resilient [1624]. cyber-security [3246]. cybersecure [2897]. Cybersecurity [2984, 3052, 3184]. cycle [670, 2439, 3497]. cycled [1532, 2231, 3027]. cyclic [3497].

D [40, 246, 655, 695, 1071, 1085, 1250, 1585, 2142, 2222, 3555]. D-FACE [2222]. D-Miner [1085]. D-SCIDS [246]. D2D [1507, 1739, 1829, 1956, 2030, 2223, 2229, 2274, 2633, 2850, 2857, 3169, 3208, 3270, 3331]. D3M [1431]. D3QN [3391]. DA-DRLS [2515]. DAC [158]. DACS [373]. DAG [3433]. Damage [545, 1868, 1917]. Dark [1854, 2249]. Data [110, 191, 249, 350, 458, 651, 836, 850, 852, 1109, 1112, 1138, 1232, 1286, 1333, 1339, 1346, 1358, 1377, 1381, 1399, 1400, 1415, 1469, 1519, 1562, 1566, 1567, 1569, 1605, 1662, 1679, 1754, 1755, 1763, 1824, 1831, 1850, 1901, 1905, 1916, 1922, 2045, 2068, 2083, 2100, 2107, 2113, 2123, 2156, 2164, 2169, 2196, 2225, 2275, 2323, 2327, 2373, 2383, 2424, 2474, 2511, 2539, 2572, 2603, 2607, 2649, 2650, 2664, 2675, 2730, 2752, 2783, 2788, 2819, 2839, 2840, 2879, 2880, 2918, 2957, 2959, 2975, 3031, 3038, 3048, 3097, 3099, 3179, 3185, 3187, 3209, 3218, 3258, 3289, 3320, 3330, 3396, 3419, 3530]. data [124, 158, 172, 237, 247, 567, 654, 734, 815, 816, 839, 844, 855, 859, 861, 868, 883, 993, 1057, 1122, 1140, 1160, 1192, 1289, 1306, 1320, 1396, 1398, 1497, 1586, 1628, 1637, 1644, 1675, 1694, 1698, 1699, 1719, 1772, 1805, 1826, 1834, 1858, 1859, 1907, 1910, 1934, 1937, 1949, 1967, 1968, 2004, 2028, 2153, 2174, 2239, 2246, 2292, 2358, 2365, 2379, 2393, 2418, 2440, 2444, 2470, 2473, 2480, 2565, 2567, 2580, 2676, 2786, 2800, 2830, 2847, 2849, 2875, 2910, 2917, 2951, 2953, 2982, 2990, 3039, 3062, 3071, 3143, 3154, 3167, 3193, 3273, 3333, 3337, 3350, 3359, 3362, 3463, 3470, 3499, 3504, 3517, 3525, 3555]. data [69, 431, 447, 451, 605, 749, 793, 867, 908, 952, 993, 1277, 1380, 1388, 1401, 1533, 1564, 1565, 1593, 1602, 1610, 1619, 1686, 1740, 1832, 1852, 1915, 1953, 1990, 2046, 2057, 2069, 2126, 2138, 2180, 2185, 2204, 2212, 2251, 2270, 2302, 2363, 2396, 2404, 2423, 2485, 2490, 2544, 2546, 2626, 2627, 2674, 2679, 2772, 2807, 2824, 2852, 2894, 2908, 2940, 3014, 3043, 3061, 3138, 3195, 3198, 3215, 3222, 3223, 3254, 3314, 3416, 3426, 3442, 3471, 3531]. data-aware [2083]. data-based [2918]. data-center [2917]. data-centric [850, 855, 1637]. data-collecting [3426]. data-driven [2153]. Data-intensive [1320, 1533, 2539, 2788, 2951]. data-plane [2546]. database [37, 92, 237, 386, 403, 1161, 2260, 3264]. databases [32, 45, 915, 1549, 2787]. Datacast [692]. datacenter [1307, 1436, 1568, 1713, 1717, 2741, 2884, 3160, 3343, 3358]. datacenters [2635, 3074]. dataset [209, 1971, 3053]. datasets [110, 2871]. date [2964, 3091]. DBMS [1931]. DBT [233]. DC [1436, 3361]. DC-PUF [3361]. DC-Vegas [1436]. DCN [3222]. DCNs [2972]. DCS [631]. DCT [3333]. DDC [2439]. DDoS [286, 402, 1687, 1693, 1923, 2222, 2429, 2497, 2576, 2581, 2735, 2796, 2981, 2995, 2996, 3022, 3191, 3514, 3540, 3543, 3554]. DDSE [2132]. de-anonymization [3397]. Dead [1295]. Deadline [2555, 3125, 3238, 3424]. Deadline-aware [3125]. deadlines [3162]. deal [1719, 2345]. dealing [2656]. debris [325]. debris-flow [325]. December [2593, 2605, 2828, 2841, 3064, 3072, 3229]. Decentralized [470, 492, 1110, 1255, 1329, 1589, 1853, 1955, 2218, 2402, 2422, 2423, 2556, 2691, 2732, 2908, 3227, 3302, 3451]. decision [77, 304, 488, 672, 1494, 1764, 2280, 2430, 3482]. decision-making [1494]. decision-taking [488]. decoder [1790]. decoding [2284]. decomposition [16, 3463]. Deconstructing [467]. deconvolution [396]. decoupled [332]. decryption [2195]. Dedup [2730]. deduplication [1389, 2470, 2730]. Deep [911, 1450, 1808, 1955, 2143, 2269, 2320, 2421, 2481, 2515, 2687, 2697, 2725, 2733, 2739, 2755, 2800, 2803, 2810, 2815, 2859, 2860, 2869, 2917, 2919, 2929, 2971, 2997, 3041, 3048, 3058, 3093, 3149, 3191, 3192, 3194, 3202, 3228, 3233, 3275, 3339, 3344, 3348, 3370, 3397, 3399, 3402, 3424, 3426, 3429, 3440, 3470, 3502, 3526, 3534, 3539, 3550, 3564]. deep-learning-based [2733]. Deer [3284]. Defeating [3397]. Defence [657, 3102, 3299]. defend [1840]. Defending [2104, 2494]. Defense [356, 1126, 1135, 1147, 1492, 1693, 2215, 2312, 2511, 2825, 2838, 3146, 3290, 3296, 3525]. Defenses [2664, 3001]. deficit [1146]. Defined [426, 865, 1446, 1486, 1591, 1592, 1619, 1678, 1693, 1697, 1711, 1725, 1837, 1855, 1870, 1874, 1910, 1940, 1952, 1961, 2000, 2008, 2037, 2045, 2131, 2152, 2229, 2243, 2272, 2285, 2300, 2329, 2333, 2334, 2375, 2382, 2434, 2445, 2491, 2494, 2504, 2532, 2535, 2598, 2599, 2631, 2660, 2672, 2680, 2729, 2800, 2825, 2844, 2867, 2876, 2905, 2917, 2919, 2972, 2995, 2996, 3006, 3022, 3041, 3045, 3049, 3060, 3102, 3161, 3187, 3249, 3278, 3317, 3319, 3344, 3436, 3474, 3485, 3487, 3543]. defining [3450]. definition [962, 1281]. degradation [3466]. Degree [1090, 1208, 2132]. degree-descending [2132]. Delay [328, 394, 402, 422, 464, 608, 747, 778, 802, 825, 860, 902, 949, 1049, 1059, 1142, 1157, 1170, 1196, 1200, 1226, 1240, 1252, 1287, 1349, 1420, 1436, 1442, 1593, 1686, 1832, 1928, 1970, 1973, 2014, 2031, 2062, 2361, 2439, 2508, 2565, 2614, 2884, 3071, 3074, 3344, 3353, 3483, 3512]. Delay-aware [2508]. delay-based [1436, 2884]. delay-constrained [778]. Delay-cost [1832]. delay-guaranteed [747, 3344]. delay-range-dependent [422]. delay-sensitive [1142, 1593, 3512]. Delay-Tolerant [860, 902, 1240]. Delayed [1945]. Delays [422, 874, 2631]. delegatable [2412]. delegated [3097]. delegation [244, 1972]. deletion [2138]. delineation [2782]. Delivering [296, 397]. Delivery [56, 431, 457, 459, 460, 462, 553, 655, 701, 810, 859, 870, 978, 1049, 1088, 1104, 1197, 1287, 1299, 1518, 1525, 1639, 1805, 2266, 2633, 2920, 3294, 3362, 3544, 3557]. demand [58, 139, 183, 213, 220, 279, 314, 480, 519, 829, 864, 888, 1098, 1195, 1283, 1352, 1442, 1514, 1603, 1699, 1722, 2258, 2644, 2885, 3345, 3546]. demand-driven [1283]. demands [2656]. Dempster [1638]. DeMS [1920]. Denial [285, 1515, 1561, 1687, 2007, 2494, 2877, 3146]. denial-of-service [285, 2494]. Denko [789]. denoising [2263]. Dense [851, 1378, 2054, 2093, 2369, 2380, 2579, 3527]. densely [1528, 3183]. density [1499, 1566, 1939, 2091, 3228]. department [1135]. Dependability [610, 663, 692, 2009, 3522]. Dependable [501, 688, 690, 698, 1592, 2442]. dependency [3361, 3472]. dependency-aware [3472]. Dependent [422, 604, 1919, 3334, 3372]. deployable [3381]. deployed [626, 1331, 1412, 1528, 2625, 2965, 3183, 3421]. deploying [329, 794, 1137, 2743]. Deployment [61, 138, 341, 430, 551, 611, 639, 1171, 1225, 1331, 1363, 1382, 1412, 1421, 1597, 1617, 1621, 1638, 1640, 1679, 1758, 1942, 1994, 2053, 2142, 2146, 2304, 2416, 2561, 2585, 2711, 2749, 2928, 2949, 2968, 3204, 3384, 3513]. deployments [964, 3523]. derivation [1158]. derived [298]. descending [2132, 2927]. description [1095, 2012]. Descriptors [85]. Design [8, 9, 19, 38, 82, 89, 96, 101, 111, 113, 114, 139, 161, 183, 238, 249, 250, 257, 265, 302, 322, 329, 384, 417, 429, 438, 623, 636, 665, 677, 722, 754, 771, 1020, 1064, 1065, 1070, 1104, 1112, 1186, 1248, 1301, 1348, 1363, 1399, 1400, 1467, 1491, 1501, 1582, 1599, 1618, 1755, 1836, 1844, 1845, 1859, 1867, 1974, 2023, 2038, 2103, 2172, 2208, 2219, 2322, 2368, 2415, 2424, 2466, 2569, 2592, 2765, 2784, 3061, 3094, 3120, 3238, 3439]. designed [913]. designers [92]. Designing [272, 517, 1169, 1718, 3557]. designs [2856]. desktop [377]. destination [338]. destinations [2094]. desynchronization [2183]. detailed [998, 2839]. detect [956, 1334, 2441, 3566]. Detecting [282, 402, 699, 837, 1087, 1128, 1131, 1259, 1262, 1272, 1348, 2205, 2405, 2451, 2540, 2739, 2784, 2819, 3066, 3259, 3543, 3551]. Detection [245, 246, 258, 262, 315, 382, 446, 479, 527, 530, 545, 666, 715, 730, 858, 933, 935, 961, 1050, 1152, 1183, 1201, 1228, 1245, 1311, 1368, 1513, 1572, 1580, 1596, 1616, 1627, 1632, 1661, 1675, 1695, 1721, 1776, 1818, 1825, 1840, 1860, 1870, 1917, 1971, 2007, 2163, 2164, 2193, 2194, 2203, 2222, 2236, 2291, 2354, 2384, 2414, 2484, 2489, 2494, 2596, 2632, 2657, 2713, 2723, 2791, 2796, 2815, 2870, 2877, 2934, 2942, 2981, 2995, 2997, 3020, 3022, 3033, 3053, 3087, 3141, 3210, 3236, 3255, 3263, 3267, 3342, 3370, 3402, 3417, 3429, 3448, 3466, 3486, 3514, 3528, 3529, 3533, 3553, 3554, 3559]. detection [153, 211, 247, 248, 285, 289, 470, 676, 890, 934, 1086, 1449, 1585, 1746, 1770, 1798, 1838, 1848, 1924, 2002, 2067, 2076, 2144, 2146, 2294, 2459, 2611, 2645, 2736, 2754, 2755, 2782, 2902, 2996, 3009, 3048, 3092– 3094, 3268, 3275, 3290, 3293, 3309, 3318, 3389, 3399, 3400, 3404, 3411, 3434, 3550]. detection/localization [1917]. detector [1671, 2175]. detectors [1655, 2717]. determination [11, 1382]. determine [352]. determining [3228]. deterministic [288, 988]. DetLogic [2205]. Detouring [381]. develop [1213]. developing [5, 107, 531, 1913, 2308]. Development [17, 24, 160, 469, 542, 569, 729, 756, 842, 862, 954, 1069, 1075, 2018, 3080, 3242, 3328, 3395, 3464]. developments [2657]. Device [511, 521, 524, 1518, 1685, 1737, 1739, 1829, 1951, 2072, 2076, 2097, 2098, 2153, 2229, 2264, 2291, 2335, 2427, 2433, 2552, 2636, 2785, 2832, 3029, 3112, 3185, 3308, 3334, 3509]. device-differentiated [1518]. Device-free [2153, 2552, 2785]. Device-to-Device [1685, 1737, 1739, 1829, 2098, 2229, 2335, 2636, 2832, 3509]. Devices [96, 180, 386, 772, 801, 1133, 1186, 1278, 1471, 1508, 1628, 1724, 1872, 1993, 2178, 2182, 2289, 2303, 2324, 2519, 2724, 2784, 2806, 2834, 2946, 3361, 3374]. Dew [2938]. Dewey [1381]. DFA [1450]. DHT [1730]. Diagnosing [3197]. diagnosis [945, 1105, 1575, 2401, 2462, 3095]. diagnostic [2280]. diagram [3103]. diagrams [59]. diameter [616]. dictionary [3314]. dictionary-based [3314]. DIDS [3002]. difference [87, 3497]. different [2179, 2232, 2298, 2470]. Differential [838, 2298, 2673, 2794, 2927, 3468]. Differentially [2982]. differentiated [570, 1518, 3157]. differentiation [260]. Diffie [698]. DiffServ [217, 619, 866, 919]. DiffServ-based [217]. diffusion [1957, 2009, 2039, 2490, 2586, 3130]. Digital [31, 33–35, 37, 136, 164, 298, 317, 367, 397, 424, 443, 494, 531, 1113, 2908, 3117, 3357, 3481, 3505, 3561]. dilemma [1771]. Dimension [860]. Dimensional [104, 470, 514, 855, 860, 930, 1120, 1380, 1499, 2018, 2496, 2577]. DIMY [3144]. DIPOR [2154]. Direct [21, 2041, 2433, 3207]. directed [616]. direction [1029, 1033, 1101, 1458, 1841, 2525, 2762, 3113, 3408]. Directional [568, 1162, 1318, 1499, 1652, 3354]. Directional-antenna-based [3354]. Directions [167, 503, 1309, 1369, 1373, 1427, 1460, 1644, 1755, 1759, 1798, 1821, 1852, 1979, 2126, 2190, 2207, 2265, 2303, 2394, 2419, 2507, 2742, 2875, 2898, 2947, 2996, 3079, 3151, 3166, 3201, 3219, 3272, 3315, 3323, 3346, 3384, 3404, 3457, 3458, 3503]. Directory [3444]. dirty [2005]. disabilities [796]. disaggregating [2760]. disaggregation [2546]. disagreement [2723]. disagreement-based [2723]. Disambiguation [2419]. disaster [325, 1007, 1209, 1926, 2199, 2218, 2397, 2501, 2514, 2832, 3297]. disclosing [3303]. disconnected [430]. discontinuous [1630]. Discover [3104]. Discovering [760, 1004, 1066, 1085, 1919]. Discovery [27, 214, 333, 459, 496, 498, 541, 617, 640, 829, 875, 924, 1106, 1163, 1239, 1380, 1427, 1651, 1803, 1813, 1956, 2003, 2028, 2072, 2379, 3011, 3121, 3217, 3449, 3497]. Discrete [196, 368, 973, 3482]. Discrete-time [368, 3482]. discriminant [514]. discrimination [1861]. discriminative [988]. discussion [1822]. Discussions [172, 2712]. disease [2264, 2866]. dishonest [2345]. Disjoint [275, 280, 525, 681, 777, 1270, 1493, 1937]. dispatch [1560]. Dispatcher [780]. dispensers [2899]. display [695, 1227]. disruption [996]. Dissecting [2915]. disseminating [778]. dissemination [736, 1062, 1289, 1329, 1339, 1686, 2199, 2676, 2807, 2830, 2990, 3105]. dissension [3557]. distance [81, 122, 636, 1200, 1464, 1977, 2090, 2796, 2822, 3428]. distance-bounding [636]. DISTILLER [2971]. distorted [508]. distortion [693]. Distributed [43, 57, 109, 110, 137, 157, 190, 191, 227, 238, 246, 253, 285, 287, 291, 351, 354, 386, 478, 481, 547, 569, 574, 651, 715, 730, 759, 771, 844, 868, 887, 1039, 1099, 1119, 1164, 1179, 1207, 1230, 1243, 1250, 1312, 1314, 1338, 1371, 1431, 1447, 1458, 1479, 1493, 1503, 1559, 1579, 1582, 1661, 1664, 1687, 1755, 1756, 1814, 1830, 1838, 1858, 1895, 1903, 1930, 1966, 1969, 2028, 2098, 2107, 2126, 2142, 2193, 2222, 2269, 2305, 2314, 2348, 2405, 2431, 2539, 2564, 2596, 2651, 2739, 2787, 2836, 2870, 2923, 2949, 3030, 3069, 3353, 3354, 3356, 3369, 3481, 3488, 3545]. distributed [85, 116, 198, 288, 289, 312, 321, 335, 336, 383, 451, 557, 603, 640, 660, 670, 739, 924, 1064, 1121, 1153, 1235, 1257, 1284, 1380, 1515, 1560, 1707, 1935, 1990, 2008, 2047, 2165, 2718, 2782, 2877, 2977, 3203, 3237, 3279, 3471]. Distributed-SOM [2008]. distribution [388, 432, 458, 461, 463, 561, 565, 619, 692, 801, 825, 863, 1118, 1403, 1474, 1548, 1672, 1812, 2092, 2137, 2143, 2146, 2298, 2422, 2508, 2531, 2564, 2672, 3149, 3169, 3338, 3367, 3400, 3457]. distributor [249]. distributors [523]. disturbance [2734]. Div [1084]. Div-clustering [1084]. diverse [2151]. diversification [2511]. Diversity [1340, 1455, 1632]. divisible [151]. Division [973, 1532, 2417, 2827, 2874]. DLM [1945]. DLS [959]. DLT [2923]. DMSP [1110]. DMVL [1154]. DNA [2849]. DNC [2860]. DNN [2543, 2587, 2717, 3378]. DNN-based [2717]. DNN-HMM [2543, 2587]. DNP3 [1563]. DNS [266, 311, 1357, 2221, 2902]. Docker [2304]. Document [19, 125, 209, 221]. documents [427]. DogitaLS1 [35]. Domain [6, 136, 150, 523, 564, 886, 937, 1087, 1102, 1201, 1961, 2117, 2466, 2473, 2729, 2843, 2928, 2967, 2980, 3069, 3373, 3532]. domain-based [1201]. Domain-independent [1102]. Domain-specific [2473, 2980]. Domains [2604, 3020, 3371]. Domatic [1108]. domestic [2867]. dominated [898]. dominating [739, 889]. Doppler [22]. Doppler-echo [22]. Dortmund [35]. double [139, 233, 616, 3009]. double-loop [616]. double-rate [139]. down [1936]. downlink [612, 773, 930, 1405, 1762]. download [1139]. downloads [975, 1033]. DP [2794, 2982]. DP-GAN [2982]. DPDRM [2218]. DPMQ [1059]. DQN [3371]. DQN-based [3371]. drainage [1830]. DRED [368]. Drift [2515, 2736]. drive [1139, 1192]. drive-by [1139]. Driven [19, 160, 753, 842, 901, 903, 1032, 1075, 1283, 1343, 1513, 1528, 1541, 1714, 1765, 1769, 1777, 1896, 1981, 2153, 2172, 2444, 2579, 2600, 2852, 3021, 3080, 3112, 3215, 3359, 3465, 3505]. drivers [1919]. driving [1919, 2676, 3327]. DRL [2917]. DRL-R [2917]. DRLS [2515]. drone [1782, 2561, 2881, 3061, 3420, 3426]. drone-mounted [2881]. DroneCells [2881]. drones [3321, 3323]. drop [768, 1059]. droplets [2283]. dropping [845]. DRX [1630, 2274]. DS [1514]. DSR [983]. DTL [3429]. DTL-IDS [3429]. DTN [1643, 2148, 2199, 2218]. Dual [435, 737, 858, 1033, 1724, 1796, 1867, 2350]. dual-direction [1033]. dual-homed [1724]. Dual-key-binding [737]. dual-purpose [435]. Dual-side [2350]. dueling [2687, 2810]. duplex [1349, 2851, 2858]. duplication [699]. duration [1178, 1235]. during [24, 393, 475, 2489, 3540]. dust [2709]. duty [1532, 2231, 2439, 3027, 3497]. duty-cycle [3497]. duty-cycled [1532, 2231, 3027]. DV [1256]. DV-hop [1256]. DVB [692]. DVB-H [692]. DVEs [562]. DVFS [1831]. DVFS-aware [1831]. DWDM [532]. dyadic [317]. Dynamic [94, 143, 144, 150, 250, 251, 256, 259, 264, 295, 297, 337, 355, 406, 436, 457, 459, 562, 580, 581, 598, 609, 654, 671, 776, 811, 856, 858, 859, 861, 872, 877, 892, 903, 921, 923, 978, 985, 987, 989, 1033, 1038, 1059, 1214, 1221, 1258, 1282, 1291, 1328, 1354, 1441, 1483, 1486, 1536, 1761, 1779, 1847, 1867, 1914, 1949, 2011, 2037, 2154, 2184, 2185, 2196, 2271, 2276, 2313, 2336, 2357, 2359, 2383, 2422, 2438, 2439, 2471, 2474, 2511, 2529, 2559, 2579, 2714, 2782, 2825, 2880, 2922, 2990, 3062, 3118, 3124, 3134, 3264, 3304, 3338, 3382, 3386, 3387, 3422, 3425, 3467, 3499, 3546]. dynamic [549, 698, 741, 929, 952, 959, 1035, 1149, 1154, 1960, 2103, 2627, 2691, 2729, 2736, 2794, 3253, 3290, 3368, 3435]. dynamically [832, 2112]. Dynamics [3, 1438, 1715, 1957, 2297, 2586, 2752].

e-catalogues [138]. E-coin [1253]. E-commerce [400, 411, 2469]. e-course [979]. E-eco [1834]. e-government [335]. e-health [3227]. e-healthcare [2525]. e-learning [118, 119, 137, 469, 1437, 2597]. E-MAC [1647]. e-mail [393, 401, 1066]. e-mails [703]. e-Voting [1253, 3034]. EAF [1347]. earliest [3238]. earliest-deadline-first [3238]. Early [475, 792, 2222, 2291, 2354, 2791, 2938, 3474]. ease [729]. eavesdroppers [2761]. eavesdropping [2104]. EBDL [3253]. eBPF [3442]. EC [2282]. ECC [1810, 3050]. ECC-based [3050]. ECCDH [3416]. ECG [20, 22, 578, 1286, 2263, 3268]. echo [22]. eCloudRFID [644]. ECN [2649]. eco [1834]. Economic [725, 1560, 1655, 2498]. ecosystem [206, 610]. ecosystems [607, 2844]. Edge [345, 854, 1195, 1489, 2126, 2267, 2358, 2419, 2425, 2438, 2464, 2471, 2495, 2500, 2525, 2555, 2612, 2639, 2644, 2676, 2753, 2764, 2788, 2795, 2811, 2826, 2850, 2922, 2929, 2948, 2958, 2974, 2987, 3002, 3007, 3032, 3051, 3096, 3107, 3109, 3118, 3127, 3129, 3140, 3142, 3156, 3165, 3170, 3205, 3245, 3269, 3272, 3295, 3305, 3313, 3329, 3334, 3336, 3348, 3349, 3371, 3372, 3378, 3384, 3410, 3413, 3433, 3459, 3488, 3489, 3492, 3498, 3500–3502, 3505, 3513, 3526, 3536, 3539]. edge-based [3269, 3492]. Edge-centric [2753]. edge-cloud [2555, 2948, 3032, 3372]. edge-enabled [2500]. Edge-limited [345]. edge-private [3505]. edge-to-core [2358]. EdgeDrones [3321]. EDHRP [1528]. editing [1083]. Editor [106]. Editorial [31, 223, 232, 240, 263, 284, 309, 324, 330, 339, 347, 380, 398, 416, 441, 455, 471, 482, 490, 502, 518, 533, 543, 552, 589, 627, 649, 687, 717, 743, 788, 820, 849, 869, 893, 931, 981, 1015, 1030, 1046, 1060, 1093, 1125, 1144, 1176, 1206, 1249, 1267, 1279, 1298, 1316, 1350, 1364, 1374, 1384, 1402, 1413, 1428, 1444, 1452, 1468, 1485, 1502, 1530, 1570, 1594, 1609, 1626, 1646, 1660, 1677, 1688, 1704, 1716, 1729, 1744, 1757, 1767, 1780, 1806, 1816, 1827, 1851, 1857, 1877, 1889, 1902, 1921]. Editorial [1, 2, 8, 224, 310, 1016, 1047, 1411, 1659, 1932, 1947, 1958, 1975, 1976, 1986, 1999, 2006, 2016, 2025, 2043, 2050, 2058, 2059, 2065, 2077, 2088, 2101, 2110, 2118, 2119, 2125, 2147, 2158, 2166, 2177, 2187, 2197, 2206, 2217, 2227, 2237, 2248, 2259, 2268, 2277, 2286, 2296, 2306, 2315, 2326, 2340, 2352, 2366, 2386, 2399, 2410, 2420, 2428, 2437, 2447, 2457, 2467, 2477, 2488, 2499, 2510, 2517, 2518, 2524, 2530, 2537, 2538, 2548, 2558, 2573, 2574, 2594, 2606, 2617, 2630, 2641, 2648, 2655, 2663, 2671, 2678, 2686, 3364]. Editorial [2694, 2701, 2708, 2716, 2727, 2738, 2747, 2757, 2768, 2780, 2793, 2805, 2817, 2829, 2842, 2854, 2873, 2890, 2901, 2913, 2925, 2937, 2945, 2955, 2963, 2970, 2979, 2986, 2993, 2999, 3004, 3013, 3024, 3036, 3047, 3057, 3065, 3073, 3082, 3090, 3101, 3111, 3123, 3136, 3148, 3164, 3177, 3200, 3212, 3230, 3240, 3252, 3257, 3266, 3282, 3301, 3312, 3326, 3341, 3366, 3380, 3393, 3406, 3432, 3447, 3461, 3480, 3491, 3507, 3516, 3521, 3542]. editors [1975, 2518]. education [88, 1977]. educational [433, 1213, 2107]. EEE [2773]. EEHSC [1337]. Effective [75, 434, 975, 1139, 1296, 1419, 1571, 2087, 2192, 2388, 2425, 2489, 2502, 2928, 2972, 3116, 3253, 3254, 3368, 3416]. Effectiveness [1242, 2167]. Effects [17, 1166]. Efficacy [41]. efficiency [220, 815, 853, 888, 969, 1448, 1472, 1523, 1529, 1539, 1548, 1562, 1735, 1756, 1761, 1774, 2128, 2347, 2360, 2398, 2439, 2509, 2696, 2718, 2751, 2881, 3098, 3287, 3425, 3457, 3560]. Efficient [46, 269, 276, 554, 563, 572, 585, 590, 654, 661, 681, 740, 810, 829, 866, 875, 899, 905, 923, 944, 947, 1025, 1107, 1139, 1160, 1250, 1299, 1326, 1333, 1370, 1380, 1389, 1406, 1447, 1465, 1488, 1497, 1518, 1525, 1528, 1595, 1614, 1621, 1633, 1691, 1693, 1710, 1789, 1794, 1831, 1901, 1931, 1949, 1988, 1992, 2084, 2120, 2127, 2140, 2162, 2218, 2258, 2318, 2423, 2424, 2429, 2460, 2462, 2463, 2471, 2474, 2480, 2542, 2627, 2631, 2632, 2635, 2643, 2778, 2783, 2801, 2858, 2868, 3005, 3086, 3124, 3125, 3140, 3162, 3169, 3188, 3213, 3314, 3332, 3335, 3351, 3353, 3390, 3414, 3485, 3499, 3532]. efficient [30, 58, 60, 100, 189, 221, 234, 292, 359, 362, 383, 429, 483, 486, 497, 528– 530, 558, 576, 640, 702, 797, 811, 840, 851, 855, 859, 957, 995, 1005, 1083, 1086, 1140, 1177, 1187, 1191, 1216, 1337, 1371, 1401, 1514, 1522, 1535, 1543, 1544, 1610, 1628, 1634, 1658, 1752, 1760, 1815, 1828, 1833, 1834, 1881, 1906, 1984, 1989, 1994, 2021, 2024, 2037, 2072, 2087, 2122, 2183, 2188, 2238, 2251, 2274, 2287, 2408, 2435, 2466, 2472, 2555, 2565, 2602, 2695, 2798, 2835, 2857, 2885, 2935, 2940, 2965, 2989, 3037, 3043, 3071, 3075, 3116, 3119, 3138, 3198, 3344, 3354, 3368, 3477, 3497, 3500, 3511, 3524, 3530, 3555]. efficient [922, 946, 1308, 1619, 1782, 2713]. effort [1869]. EGHR [2120]. egress [311]. eHealth [1325, 2275]. eigen [2002]. ELA [3485]. ELA-RCP [3485]. Elastic [1472, 1700, 1991, 2417, 2836, 2874]. elasticity [1623, 1936, 2126]. election [425]. elections [151]. electric [1745, 2882, 3296]. electricity [1548]. electrocardiogram [2279]. electrocardiogram-based [2279]. electronic [101, 187, 424, 1189]. elephant [3477]. eliminate [712]. Elimination [1289, 1862]. elitism [2931]. elitism-based [2931]. Elliptic [698]. Elliptic-Curve [698]. EM-B [740]. Email [420, 686, 812, 2415, 3255]. eMBB [3417]. embed [3510]. Embedded [82, 268, 683, 1334, 1543, 1891, 2165, 3185]. Embedding [1439, 1479, 1717, 2361, 2656, 2704, 2729, 2766, 2860, 3186, 3249, 3353, 3412]. emergencies [475]. emergency [761, 1752, 2397, 3108, 3561]. emergent [1926]. Emerging [1094, 1330, 1987, 2017, 2054, 2419, 2436, 2456, 2516, 2554, 2996, 3182, 3201, 3299, 3448]. EMM [2108]. emotion [316, 2597]. emotions [674]. emphasis [2597]. EMPIOT [2324]. Empirical [582, 1507, 2325, 2642, 2915, 3146]. employment [2511]. empowered [3159, 3388, 3462]. Empowering [1725]. Emulation [239, 2036, 2433, 3134]. EmuSocket [239]. Enable [59, 1148, 2091, 3134]. Enabled [397, 406, 540, 642, 643, 1096, 1400, 1860, 2216, 2302, 2387, 2455, 2456, 2500, 2504, 2514, 2516, 2568, 2751, 2880, 2910, 2990, 3077, 3127, 3154, 3244, 3245, 3249, 3259, 3260, 3413, 3440, 3458, 3492, 3546]. enablers [3107]. enables [1348]. Enabling [85, 523, 537, 648, 872, 1552, 1598, 1869, 1885, 1887, 2034, 2039, 2040, 2045, 2896, 2932, 3144, 3158, 3383, 3409]. encoder [693]. encoding [521, 1381, 2256, 2284, 2755, 2836, 3510]. encountered [1200]. encounters [2859]. encouraging [2774]. Encrypted [486, 2099, 2174, 2453, 2603, 2627, 2721, 2730, 2777, 2870, 2971, 3237]. Encryption [243, 576, 595, 713, 946, 1543, 1549, 1550, 1949, 1972, 2137, 2173, 2185, 2190, 2195, 2341, 2412, 2413, 2423, 2500, 2910]. encryption-based [2185]. End [186, 372, 418, 524, 698, 894, 1010, 1031, 1067, 1217, 1226, 1252, 1369, 1552, 1745, 1984, 2500, 2607, 2699, 2810, 2892, 2967, 3391, 3419]. End-to-End [186, 698, 894, 1010, 1217, 1226, 1552, 1745, 1984, 2500, 2699, 2810, 2892, 2967, 3391, 3419]. end-user [1067]. endeavors [3528]. endpoint [962]. energies [568]. Energy [597, 693, 702, 853, 897, 899, 944, 971, 1107, 1112, 1196, 1199, 1221, 1232, 1270, 1288, 1291, 1333, 1337, 1347, 1371, 1448, 1472, 1489, 1506, 1518, 1521, 1522, 1525, 1526, 1528, 1536, 1541, 1548, 1586, 1630, 1633, 1634, 1669, 1683, 1691, 1694, 1711, 1735, 1756, 1758, 1761, 1789, 1828, 1831, 1833, 1906, 1909, 1931, 1989, 1996, 2001, 2021, 2072, 2128, 2140, 2162, 2218, 2224, 2228, 2274, 2282, 2298, 2313, 2324, 2406, 2446, 2455, 2480, 2485, 2509, 2555, 2569, 2635, 2650, 2798, 2801, 2868, 2893, 2935, 2952, 3005, 3091, 3098, 3125, 3162, 3169, 3287, 3344, 3351, 3354, 3414, 3470, 3485, 3511, 3560, 3563, 3569]. energy [279, 558, 797, 851, 1014, 1022, 1177, 1191, 1216, 1274, 1296, 1466, 1514, 1533, 1535, 1538, 1539, 1544, 1553, 1555, 1619, 1651, 1680, 1774, 1834, 1941, 1970, 1977, 1992, 2180, 2242, 2287, 2325, 2347, 2398, 2408, 2439, 2464, 2565, 2602, 2695, 2696, 2773, 2977, 3018, 3071, 3138, 3223, 3285, 3292, 3410, 3497, 3530]. Energy-Aware [279, 897, 1199, 1347, 1489, 1521, 1533, 1536, 1541, 1619, 1683, 1694, 1931, 1970, 1996, 2313, 2464, 2485, 3563]. Energy-balanced [1112, 1586, 2180]. energy-bandwidth [1711]. energy-based [1274]. energy-constrained [1221]. Energy-delay [1196]. Energy-efficiency [1735]. Energy-efficient [558, 1191, 1216, 1333, 1371, 1518, 1522, 1525, 1535, 1633, 1828, 1834, 1931, 2162, 2274, 2408, 2480, 2555, 2801, 2935, 3005, 3125, 3162, 3169, 3344, 3351, 3354, 3414, 3485, 3497, 3511]. energy-harvesting [1022, 2569]. energy-optimization-based [1538]. energy-saving [597]. enforced [691]. enforcement [265, 1559, 3168]. engine [208, 266, 488, 1462, 3159]. Engineering [5, 74, 384, 469, 536, 596, 804, 1028, 1040, 1195, 1714, 1883, 2621, 2837, 3006]. engines [53, 479, 1271, 1718]. English [146]. Enhance [33, 1936, 2807, 3374]. Enhanced [287, 415, 454, 605, 700, 710, 775, 867, 927, 1042, 1056, 1214, 1218, 1278, 1342, 1409, 1603, 2108, 2130, 2184, 2373, 2527, 2681, 2791, 3093, 3141, 3190, 3217, 3372, 3428, 3454, 3484, 3514]. Enhancement [447, 1137, 1215, 1340, 1363, 2468, 2920, 3208, 3235]. Enhancing [71, 610, 757, 877, 879, 1676, 1761, 1825, 2066, 2181, 2433, 2723, 2833, 2885, 3218, 3417, 3523]. Ensemble [211, 1661, 2236, 3052, 3255, 3267, 3487, 3496, 3550]. ensembles [318]. ensure [908]. ensuring [1722]. entanglement [3475]. enterprise [171, 435, 440, 812, 1068]. enterprises [161]. ENTICE [1466]. entity [45, 2730]. entropy [1844, 1954, 2724]. entry [391]. entry-level [391]. envelopment [815]. environment [5, 36, 83, 97, 98, 134, 170, 173, 191, 226, 254, 333, 390, 405, 550, 571, 586, 610, 823, 898, 1017, 1064, 1072, 1135, 1229, 1430, 1519, 1601, 1656, 1764, 1818, 1942, 2049, 2093, 2160, 2211, 2257, 2290, 2327, 2363, 2435, 2463, 2503, 2545, 2555, 2638, 2644, 2711, 2754, 2764, 2849, 2851, 2858, 2909, 2931, 2933, 2946, 2948, 2968, 3002, 3032, 3038, 3096, 3145, 3162, 3184, 3186, 3309, 3459, 3473, 3500, 3517, 3540]. environment-friendly [1764]. environment-oriented [2968]. environmental [1869, 2170]. Environments [3, 88, 93, 108, 115, 116, 137, 163, 180, 238, 378, 430, 439, 466, 481, 531, 577, 603, 680, 711, 746, 779, 810, 853, 909, 956, 1003, 1076, 1179, 1230, 1250, 1280, 1376, 1378, 1385, 1388, 1434, 1453, 1549, 1579, 1621, 1644, 1736, 1743, 1880, 1882, 2036, 2055, 2092, 2139, 2497, 2597, 2642, 2723, 2749, 2868, 2879, 3075, 3205, 3305, 3378, 3425, 3444, 3489, 3508, 3531, 3544]. EON [2529]. EPC [643, 1158]. EPCglobal [640, 642]. ephemeral [573]. Epidemic [432, 1849, 2100]. Epidemiological [1346]. episodes [498]. equations [1898]. equipment [105, 1630]. equipments [925]. equivalent [2660]. era [2078, 3005, 3225, 3360]. ERGID [1752]. error [599, 809, 1003, 1113, 1917, 2251, 3463]. error-free [809]. escalation [3412]. essential [3331]. establishing [1723]. establishment [1158]. estimate [2179, 2625, 3155]. estimating [326, 1631]. Estimation [436, 453, 564, 768, 787, 808, 926, 963, 1150, 1187, 1357, 1387, 1487, 1497, 1505, 1546, 1560, 1642, 1707, 1808, 1817, 1911, 2001, 2032, 2090, 2159, 2204, 2371, 2493, 2552, 2634, 2644, 3367, 3421, 3428]. Ethereum [3120]. Ethernet [521, 741, 1360, 1907, 1984, 2159, 3207]. Euclidean [2029]. EUMD [2713]. evaluate [815]. evaluated [1257]. Evaluating [53, 1007, 1074, 1195, 1231, 1623, 1882, 1984, 2502, 2836]. Evaluation [250, 351, 368, 389, 391, 411, 438, 439, 450, 454, 484, 863, 875, 878, 928, 954, 999, 1009, 1020, 1043, 1053, 1071, 1072, 1122, 1186, 1247, 1399, 1459, 1526, 1540, 1558, 1582, 1607, 1618, 1724, 1862, 1865, 1867, 2038, 2172, 2179, 2202, 2299, 2322, 2344, 2660, 2668, 2689, 2703, 2765, 2787, 3061, 3102, 3137, 3204, 3238, 3417, 3471]. evasion [2144]. EVE [190]. evenly [670]. event [629, 1120, 1136, 1328, 1528, 1738, 1794, 1896, 3518]. event-driven [1896]. events [397, 758, 1926, 2222, 3155]. Eventual [858]. everything [3346]. Evidence [48, 1951]. EViMA [3162]. Evolution [364, 557, 838, 2298, 2927, 3225, 3351]. Evolutionary [230, 353, 721, 1040, 1569, 1647, 2132, 2405, 2609, 3278, 3394, 3443]. evolved [2015]. evolved-packet [2015]. evolving [93, 245, 1566]. Ex [1100]. Ex-BCGs [1100]. exact [1504, 1700]. examination [99, 2623]. examining [812, 998]. example [3397]. Exchange [39, 221, 613, 905, 976, 1122, 1944, 2275]. exchanges [2078]. exclusion [70]. execution [46, 407, 542, 1121, 1426, 2290]. exercising [2489]. exfiltration [2113]. exhaustive [2797]. existing [32, 2114]. exogenous [3155]. Expanded [1100]. expansible [354]. Expectation [901]. Expectation-Maximization [901]. expected [652]. expenditure [2644]. Experience [692, 760, 1213, 1445, 1676, 2283, 2561, 3283]. Experiences [65, 176, 607]. experiment [104, 547]. Experimental [33, 1166, 3409]. experimentation [313, 683, 1725]. experiments [1879]. expert [10, 1265, 3489]. explicit [1793, 2535, 3343]. Exploitation [477, 838, 2664]. Exploiting [140, 338, 876, 1095, 1773, 2150, 3566]. exploration [838, 1613]. exploratory [897]. explore [2752]. Exploring [930, 1073, 1084, 1349, 1664, 2492, 2728, 3007, 3465, 3498]. explosion [1011]. exponentiation [2293]. expression [834, 1074, 1462, 3232, 3314]. expressive [316]. extended [602, 1398, 2837, 3094]. extendible [771]. Extending [312, 2338]. Extensibility [86, 126–128, 130, 131]. Extension [150, 581, 720, 1633, 2301]. extensive [3220]. external [1936, 2113]. extracted [516]. Extracting [208, 784, 2461]. extraction [55, 95, 550, 2381, 2869, 3172]. extracts [1951]. eye [153].

fabric [1968]. face [316, 514, 515, 527, 1074, 2222, 3434]. Facial [222, 527]. facilitating [1177, 3105]. facilities [38]. facing [995]. fact [3214]. factor [1810, 2139, 2719, 2909, 2938, 3038]. factories [3513]. factorization [382, 1313]. factors [99]. fading [1326]. Failure [858, 1185, 1311, 1968, 2136, 3186]. failures [948]. Fair [221, 613, 619, 974, 1146, 1728, 1946, 2245, 2444, 2864]. Fairness [866, 919, 1759, 1965, 2220, 2681, 2724, 3453, 3456]. fairness-aware [2681]. fairness-based [2220]. FAIS [1259]. Fake [3015]. falling [1995]. False [712, 734, 2511, 2819, 3048]. false-data [2511]. FAMD [2884]. familiar [3185]. family [739]. FAN [2976]. Fano [1790]. FAPKC3 [49]. far [2188]. farming [2862]. Fast [382, 410, 429, 437, 605, 706, 833, 988, 1020, 1033, 1207, 1214, 1300, 1342, 1433, 1462, 1664, 1693, 1771, 1875, 1907, 1943, 2053, 2122, 2245, 2372, 2409, 2609, 2661, 2976, 2981, 3183, 3420]. Faster [578]. FastHand [3183]. FASTSCALE [2609]. Fault [365, 407, 446, 618, 681, 779, 858, 875, 1008, 1038, 1105, 1115, 1191, 1296, 1493, 1523, 1575, 1584, 1707, 1848, 2061, 2162, 2246, 2541, 2555, 2578, 2821, 2824, 2988, 3103, 3128, 3302, 3351, 3475, 3522]. fault-aware [2162]. Fault-proof [2821]. fault-tolerance [1296]. Fault-tolerant [407, 618, 1115, 1191, 1523, 2061, 2555, 2988, 3302, 3351, 3475]. faults [2203]. faulty [2401, 2441]. fax [43]. FB [2318]. FB-APSP [2318]. FD [2939]. Feasibility [1604, 2167, 3292]. feature [222, 282, 568, 666, 812, 1245, 1495, 2381, 2827, 3172, 3496]. feature-based [812]. features [508, 516, 527, 766, 987, 1438, 1807, 2459, 2554, 2869, 3550, 3566]. featuring [180, 219, 225, 1096]. February [2640, 2647, 2889, 2900, 3089, 3256, 3431]. FEC [782]. federate [145]. Federated [404, 1477, 1822, 3213, 3308, 3313, 3342, 3395, 3418, 3445, 3451, 3536, 3539]. federation [1749, 2774]. federations [3069]. federator [3536]. feed [370]. feedback [663, 704, 807, 1006, 1301, 2097]. feedback-based [1301]. Feedforward [11, 1469]. feeds [524]. femtocell [936, 1119, 1338, 1762, 2579]. fenced [2231]. FI [686, 997, 1965, 2041, 2552, 2637, 2781, 2964, 3030, 3547].

FIB [1815]. fiber [1027, 2513]. fiber-wireless [1027, 2513]. fibers [1027]. fibrillation [3268]. fictitious [1119]. fidelity [2433, 2871]. Field [1668, 2599, 2776, 2899, 3527]. fields [168, 2289]. Fifth [1941]. Fighting [845, 1189]. file [84, 111, 447, 458, 558, 694, 1033, 1133, 1211, 1474, 1488, 2173, 2307, 2354]. files [793, 910]. filming [1702]. Filter [179, 446, 1062, 1120, 1148, 1152, 1649, 1963, 2496, 2665, 2904, 2989, 3416]. filter-based [446, 1148, 2665]. filtering [104, 286, 751, 906, 961, 1370, 2395]. filters [988, 2004]. find [2029]. Finding [703, 889, 1262, 1565]. Fine [890, 1962, 2192, 2233, 2239, 2273, 2800, 2894, 3133, 3413]. Fine-grained [890, 1962, 2192, 2233, 2239, 2273, 2800, 3133, 3413]. Finger [510]. fingerprint [504–506, 508, 2089, 2261, 2407]. fingerprint-based [505]. fingerprinting [1297, 1672, 2703, 3271, 3277]. finite [188, 498, 994, 2625, 2837]. finite-buffered [994]. Finnish [1477]. FinTech [2145]. fire [2370, 3537]. Firewall [3290]. Firm [3473]. firmware [3363, 3472]. first [664, 1700, 3238]. first-last-exact [1700]. fit [1700, 1741]. fitness [2652]. FiWi [1027]. fixed [1499, 1937]. fixed-orientation [1499]. Flag [2642]. flash [1395, 1620, 1804, 2222]. Flaws [632]. FLBA [1443]. flexgrid [1472]. Flexibilizing [2656]. Flexible [47, 118, 119, 187, 198, 214, 496, 831, 916, 1676, 1712, 3133, 3191, 3254, 3433, 3509]. flexibly [1092]. FlexMon [3133]. flock [674]. flocking [1946, 2061]. flocking-inspired [1946]. flood [2234]. Flooding [356, 602, 996, 1509, 1830, 2008, 2581, 2844, 3474]. Flora [37]. Flow [158, 325, 345, 750, 982, 1235, 1446, 1447, 1619, 1934, 1963, 2210, 2305, 2364, 2381, 2453, 2497, 2697, 2721, 2729, 2754, 2800, 2813, 2884, 3077, 3226, 3293, 3362, 3389, 3399, 3504]. Flow-Aware [2884]. flows [1086, 1166, 1627, 1910, 2189, 2434, 3160, 3477]. FlowVisor [3443]. Fluctuating [452, 2874]. flying [2353, 2881]. focus [2808]. focused [3492]. Fog [1901, 2080, 2373, 2404, 2418, 2419, 2475, 2479, 2485, 2568, 2579, 2659, 2702, 2759, 2806, 2808, 2875, 2892, 2909, 2931, 2946, 2947, 2983, 3031, 3112, 3125, 3149, 3170, 3181, 3190, 3243, 3260, 3309, 3329, 3387, 3423, 3508, 3536]. Fog-assisted [2568]. Fog-based [1901, 2475, 2659]. Fog-Cloud [3508]. fog-driven [2579]. Fog-enhanced [2373, 3190]. fog-robotic [3423]. fog-to-cloud [2404]. fog/edge [3170]. Folded [1760]. footprints [2221, 3416]. force [1074]. forcing [979, 3214]. forecast [2001, 2320]. Forecaster [3134]. Forecasting [1359, 3155]. Forensic [1192, 1727]. forensics [990, 1674, 1951, 3416]. forest [2370, 3172, 3486]. forgery [699, 1798]. form [301]. Formal [24, 54, 411, 1032, 1198, 1563, 3439, 3522]. Formally [2588]. formation [497, 801, 2223, 2392, 2847, 2976, 3109]. formats [3076]. forming [992]. formulas [1288]. formulation [40, 728]. Forward [1113, 1505, 2385, 2776, 2960]. Forwarder [1356]. Forwarding [465, 972, 1009, 1170, 1200, 1399, 1493, 1522, 1602, 1815, 2035, 2169, 2203, 2429, 2501, 2527, 2607, 2660, 2905, 3077, 3195, 3248, 3558]. Fostering [771]. Foundations [1599]. Fountain [1113]. four [1404]. Fourier [136, 2564]. fourth [1300]. fourth-generation [1300]. FPGA [1312, 2358, 2682].

FPGA-Accelerated [2358]. FPGA-based [1312]. FPGAs [46]. FPMIPv6 [2015]. fractal [1207, 1969]. Fractional [136, 218]. fragmentation [271, 753]. fragments [959]. Frame [635, 755, 809, 1042, 1525]. Frame-based [755, 1525]. frame-length [635]. FrameIDTV [531]. Framework [27, 50, 62, 77, 123, 138, 176, 251, 302, 308, 333, 461, 485, 492, 531, 563, 588, 624, 643, 644, 655, 663, 667, 683, 753, 776, 785, 824, 839, 916, 943, 992, 1045, 1085, 1161, 1163, 1198, 1282, 1294, 1333, 1367, 1459, 1494, 1557, 1611, 1616, 1648, 1687, 1814, 1879, 1904, 1905, 1931, 1993, 2040, 2048, 2074, 2148, 2193, 2304, 2307, 2308, 2376, 2384, 2432, 2452, 2458, 2465, 2469, 2498, 2500, 2539, 2545, 2562, 2576, 2622, 2752, 2801, 2877, 2977, 2984, 3037, 3049, 3074, 3093, 3143, 3168, 3202, 3243, 3263, 3294, 3335, 3370, 3437, 3455, 3462, 3492, 3495, 3499, 3505, 3564]. framework [69, 112, 127, 181, 226, 265, 352, 369, 646, 762, 952, 1433, 1547, 1558, 1775, 1886, 2002, 2067, 2181, 2284, 2665, 2682, 2720, 3070, 3165, 3191, 3204, 3237, 3261, 3377, 3429, 3440]. frameworks [1355, 1415, 1802, 2513, 2652, 2772, 3398, 3471]. Fraud [1695, 1860, 2236]. Free [139, 348, 504, 809, 838, 879, 917, 1018, 1112, 1191, 1222, 1334, 1347, 1362, 1838, 1859, 2040, 2076, 2090, 2091, 2153, 2261, 2552, 2643, 2774, 2776, 2785, 2965, 3113, 3527]. Free-Riders [2774]. freezes [2048]. FREME [1462]. frequency [564, 973, 1472, 1588]. Frequent [498, 530, 1963, 2262]. Fresnel [2781]. friend [2921]. friendly [17, 817, 1237, 1378, 1764, 2761, 2961]. FriendShare [2307]. fronthauling [3452]. FRR [2809]. FSFC [2665]. FSQCN [1907]. fuel [2899]. FUIR [1719]. full [1006, 1349, 2851, 2858, 3401]. full-duplex [1349, 2851, 2858]. Fully [1589, 1903, 2185]. fully-decentralized [1589]. function [453, 554, 1791, 1938, 2038, 2282, 2588, 2592, 2666, 2743, 2812, 3469, 3500]. functional [1749, 1998, 3469]. functionalities [2554]. functionality [3307]. Functions [1224, 1308, 1445, 1705, 2053, 2252, 2542, 2598, 2642, 2803, 3086, 3139, 3386]. fundamental [50]. Fundamentals [2571, 3107, 3273, 3384]. Fusing [1719]. Fusion [505, 839, 1377, 1638, 3470]. Fusion-based [1638]. Future [226, 503, 632, 803, 853, 953, 1309, 1369, 1373, 1386, 1427, 1460, 1644, 1736, 1759, 1797, 1798, 1821, 1852, 1883, 1979, 2041, 2044, 2046, 2068, 2126, 2190, 2207, 2212, 2265, 2303, 2339, 2394, 2492, 2507, 2525, 2551, 2692, 2742, 2745, 2798, 2808, 2814, 2875, 2891, 2898, 2947, 2956, 2996, 3000, 3068, 3079, 3151, 3166, 3175, 3201, 3219, 3235, 3272, 3315, 3323, 3337, 3346, 3404, 3408, 3457, 3458, 3569]. FUZZ [3472]. fuzzing [3472]. Fuzzy [44, 247, 262, 479, 504, 598, 654, 805, 1004, 1103, 1146, 1168, 1259, 1354, 1443, 1831, 1971, 1981, 2282, 2369, 2627, 2838]. fuzzy-based [44].

G [74, 2271, 2274, 2287, 2331, 2339, 2358, 2361, 2443, 2461, 2486, 2487, 2497, 2505, 2521, 2581]. gain [161]. gains [1771]. Game [353, 557, 672, 1119, 1167, 1211, 1338, 1430, 1448, 1715, 1745, 1762, 1943, 1944, 2009, 2319, 2348, 2863, 2888, 3130, 3216, 3248, 3294, 3303, 3304, 3425]. game-based [672, 1943, 2009, 3294, 3303, 3304]. game-like [1430]. game-theoretic [2319]. games [272, 765, 2215, 2545]. gaming [1467]. GAN [2982, 3214]. gap [2190]. gaps [3287]. Garbled [1294, 2188]. gas [2244]. gateway [774, 1090, 1181, 1743, 1779, 1994, 2703, 3250]. gateways [819, 1885]. gathering [1586, 1679, 1850, 2107, 3138]. gating [3092]. Gaussian [2777]. gaze [156]. GBC [1938]. GBC-based [1938]. GBF [179]. Gbps [2760]. GCS [303]. GCS-MA [303]. Gen2 [1158]. gene [1074]. General [1324, 2301]. generalised [172]. generalization [3439]. generalized [3534]. Generated [493, 3020]. Generating [225, 834, 1271, 1971]. generation [38, 94, 337, 352, 456, 457, 505, 769, 1062, 1265, 1300, 1342, 1681, 1720, 1728, 1852, 1876, 1898, 1941, 2127, 2343, 2590, 2596, 2725, 2814, 2837, 2926, 2958, 3025, 3202, 3331, 3346, 3529, 3568]. Generation-2 [2926]. generative [2766, 2982, 3141, 3309, 3445, 3529]. generator [2564]. Generic [18, 145, 1294, 1814, 2027, 2583, 2610]. Genesis [1204]. Genetic [3, 12, 13, 163, 242, 247, 521, 979, 1761, 1833, 2272, 2297, 2305, 2330, 2621, 2931, 3249, 3278, 3429]. genetics [262]. genetics-based [262]. genre [765]. geo [1935, 2539]. geo-distributed [1935, 2539]. geodistributed [3531]. Geographic [1532, 1873]. geographical [1915]. geometric [513, 1324, 2298]. geometry [715]. geospatial [915, 1532]. GeoXACML [1872]. gestures [316]. GET [2234]. give [805]. glance [1573]. Global [1123, 1595, 2183, 3190, 3210, 3464]. glucose [2297]. glucose-dynamics [2297]. GMPLS [1040]. goal [436]. goes [2584]. good [2580, 3224]. goodput [2092]. Google [1192, 2752]. GopJam [2346]. GOR [3221]. gorilla [3372]. GOSPF [1789]. gossiping [2346]. government [335]. GöwFed [3342]. GPGPU [2238]. GPS [865, 884, 917]. GPS-free [917]. GPSR [1354]. GPU [1768, 1963, 3207]. GPU-accelerated [1768]. GPU-based [1963]. GPU-NFV [3207]. GR [1294]. gradient [938]. grained [405, 890, 1962, 2192, 2233, 2239, 2273, 2800, 3133, 3413]. gram [1727, 3076]. grammar [179]. grams [3076]. grant [1614, 3322, 3527]. grant-free [3527]. granted [2235]. granularity [580, 1322, 1620, 1726, 2894]. Graph [1242, 1480, 1648, 1919, 2029, 2357, 2582, 2878, 2921, 2981, 3067, 3179, 3264, 3389, 3403, 3412, 3487, 3534, 3549]. Graph-Based [1242]. graph-coloring [2582]. graphical [1438, 2608]. graphics [655, 1790, 2725]. Graphs [1100, 1542, 1721, 2980]. gravitational [1141, 2136]. GRBC [2252]. GRBC-based [2252]. greater [2580]. greedy [722, 1009, 1171, 1365, 2741, 3125]. GreedyZero [1233]. Green [969, 1367, 1429, 1498, 1517, 1524, 1527, 1531, 1536, 1537, 1541, 1544, 1545, 1668, 1745, 1823, 1919, 1931, 2053, 2063, 2140, 2276, 2287, 2695, 2740, 2798, 2862, 3091, 3445, 3527]. Green-IoT [3091]. grey [401, 3390]. Grid [110, 184, 333, 342, 343, 419, 439, 440, 550, 605, 622, 625, 630, 665, 779, 824, 855, 861, 862, 956, 964, 971, 980, 1002, 1103, 1239, 1293, 1296, 1381, 1466, 1554– 1557, 1559, 1560, 1653, 1733, 1745, 1778, 1809, 1982, 2141, 2299, 2331, 2514, 2819, 2977, 3246]. grid-based [964, 1002]. grids [447, 924, 993, 1167, 1472, 1558, 2456, 2481, 2516, 3048]. GROM [3534]. grooming [969, 1234, 2529]. Ground [3294, 3324, 3469]. Ground-vehicles [3294]. Group [132, 252, 303, 304, 388, 442, 460, 585, 594, 673, 1024, 1071, 1385, 1587, 1598, 1748, 1894, 1938, 2120, 2281, 2341, 2636, 2681, 2734, 3092, 3221, 3303]. Group-based [460, 1024, 1748, 2120]. group-choice [304]. group-gating [3092]. Group-oblivious [3221]. group-oriented [388]. grouping [632, 1153]. grouping-proofs [632]. groups [275, 1319]. groupware [123]. growth [444, 1168]. GRS [1748]. GRU [2919]. GSA [1845, 2975]. GSA-RPI [2975]. GSM [140]. guarantee [473, 1146, 1148, 3021, 3074, 3108, 3424, 3456]. guaranteed [747, 1252, 3344]. guarantees [58, 109, 559, 2950, 3196]. Guard [3293]. guarding [3537]. guessing [3286]. Guest [8, 31, 1016, 1047]. guided [2736, 3172, 3424]. guided-deep [3424]. guidelines [632, 1689]. guides [940, 1182]. guiding [3556].

H [480, 692]. H-QoS-demand [480]. H.264 [615]. H.264/AVC [615]. H.265 [2836]. HADES [3414]. Hadoop [1317, 2398]. Hail [3020]. half [2858]. half- [2858]. hallucinations [185]. HaLow [2964]. hand [1038]. hand-off [1038]. handhelds [236]. handler [2008]. Handling [948, 1794, 2000]. handoff [468, 672, 731, 1020, 1590, 1769, 1943, 2075, 2124, 2171, 2368]. handover [523, 813, 937, 1034, 1300, 1342, 1908, 2120, 2122, 2372, 3050, 3055, 3183]. Hands [348, 527]. Hands-free [348]. hard [47, 58, 227, 2003, 3118]. hard-copy [47]. Hardening [3478]. Hardware [19, 2171, 2329, 2398, 2491, 2784, 3167, 3238, 3271, 3472]. hardware-based [2398, 3167]. harmonically [1601]. harmonious [1073]. harmonize [3557]. Harnessing [462, 3445]. harvesting [1022, 1844, 2325, 2380, 2455, 2569, 3410, 3530]. Hash [323, 383, 554, 781, 988, 1224, 1922, 2198]. Hash-based [2198]. hashing [321, 513]. Hassan [3291]. haul [992, 2529]. HCCA [700]. HCube [1567]. HD [1106, 3472]. HD-FUZZ [3472]. HDFS [2130]. heads [883, 928]. healing [561]. Health [362, 852, 1743, 1895, 1917, 1990, 2262, 2266, 2281, 2349, 2476, 2591, 2652, 2866, 2938, 3227, 3359]. health-care [362]. health-meteorological [2938]. Healthcare [314, 480, 587, 1111, 1134, 1995, 2186, 2279, 2308, 2483, 2523, 2525, 2930, 2941, 3042, 3044, 3080, 3242, 3315, 3356, 3462]. heart [2489]. Hellman [698]. help [39]. Help-Exchange [39]. helpers [1340]. HELPHED [3255]. HeporCloud [2868]. Here [2584]. Heterogeneous [134, 328, 369, 372, 422, 427, 449, 731, 895, 992, 1034, 1137, 1229, 1266, 1307, 1321, 1331, 1337, 1409, 1412, 1432, 1491, 1579, 1658, 1675, 1692, 1715, 1741, 1852, 1909, 1933, 1942–1944, 2009, 2032, 2033, 2107, 2157, 2193, 2196, 2356, 2553, 2565, 2583, 2586, 2614, 2635, 2735, 2744, 2830, 2868, 3137, 3297, 3358, 3414, 3421, 3489, 3545]. Heterogeneous-belief [1266]. HetNets [1269, 2375]. Heuristic [147, 606, 1318, 1652, 2053, 2416, 2935, 2946]. heuristic-based [2416]. heuristics [1831]. HFedDI [3308]. hidden [1000, 2777]. Hiding [507, 844, 868, 1227, 3307]. Hierarchical [206, 388, 428, 440, 892, 907, 921, 929, 1177, 1381, 1762, 2011, 2219, 2316, 2714, 2850, 2957, 2990, 3011, 3233, 3453]. Hierarchical-game-based [1762]. hierarchy [291]. High [249, 418, 462, 569, 684, 757, 962, 1036, 1237, 1273, 1281, 1304, 1464, 1664, 1728, 1768, 1790, 1812, 2334, 2533, 2541, 2546, 2615, 2682, 2718, 3049, 3160, 3339, 3525]. high-bandwidth [3160]. high-BDP [1273]. high-definition [962]. high-efficiency [2718]. high-end [418]. high-level [569]. High-performance [1237, 1768, 2682, 3049]. high-precision [3525]. High-speed [249, 684, 1273, 1304, 1464, 2546, 3339]. highly [690, 1873, 3140, 3254]. highway [2695]. highways [1017]. hill [914]. HiMAQ [3453]. history [1200]. history-based [1200]. hit [21]. HMM [2543, 2587]. Hoarder [1905]. Hoc [155, 258, 273, 279, 292, 320, 357, 415, 442, 472, 474, 477, 529, 536, 541, 546, 572, 585, 597, 662, 679, 685, 691, 694, 721, 724, 736, 764, 767, 781, 790, 828, 835, 845, 880, 888, 898, 915, 926, 956, 983, 1009, 1019, 1026, 1048, 1124, 1127, 1155, 1157, 1173, 1190, 1205, 1276, 1315, 1326, 1329, 1351, 1354, 1399, 1419, 1475, 1481, 1501, 1521, 1526, 1575, 1603, 1606, 1647, 1696, 1709, 1747, 1751, 1754, 1761, 1819, 1900, 1903, 2086, 2164, 2220, 2312, 2336, 2353, 2377, 2405, 2468, 2858, 2961, 3040, 3105, 3161]. HOCA [1111]. hole [715, 1014, 1191, 1368, 1680]. holes [2602]. Holistic [624, 2414, 3189, 3276, 3285]. holonic [229]. Home [268, 438, 480, 1855, 2070, 2116, 2281, 2559, 3475]. homed [1724]. homeland [1126, 1130, 1131, 1135]. homes [1813, 2659]. homing [770]. homogeneous [1108, 1216]. Homomorphic [2123, 2185]. Honeybees [1710]. honeynet [3523]. honeynet-based [3523]. honeypot [3307]. honeypots [3398]. hop [464, 480, 570, 606, 845, 939, 1079, 1166, 1256, 1348, 1481, 1500, 1507, 1571, 1833, 2108, 2209, 3530]. HOPASS [3074]. Hopfield [7]. hopping [2392]. horizontal [3308]. hospital [76]. hospitals [629]. Host [741, 2224, 3404]. host-based [3404]. hostile [2968]. hosting [251]. hosts [972]. hot [1401, 1973, 2827]. hot-spot [1973]. hotspot [1289]. hotspots [1853, 2422]. HPC [897]. HPM [907]. HRCal [2489]. HRSPC [1664]. HSecGR [1873]. HSIRD [2586]. HTML [146]. HTTP [454, 657, 785, 1807, 2007, 2048, 2115, 2234, 3295, 3459]. HTTP-based [1807]. HTTP-DoS [657]. HTTP-GET [2234]. HTTP-MPLEX [454]. HTTP/2 [2007]. hull [2029]. Human [362, 860, 1073, 1254, 1369, 2076, 2084, 2153, 2533, 2785, 3172, 3496]. humanitarian [641]. Hybla [2362]. Hybrid [15, 154, 248, 254, 450, 500, 542, 641, 684, 774, 779, 1011, 1040, 1088, 1183, 1289, 1504, 1528, 1664, 1690, 1786, 1819, 1865, 1866, 1875, 1918, 1920, 1934, 2003, 2023, 2096, 2106, 2129, 2226, 2543, 2587, 2614, 2645, 2699, 2720, 2750, 2787, 2868, 2929, 3006, 3033, 3255, 3305, 3314, 3417, 3471, 3472, 3486, 3489, 3513, 3537, 3564]. Hydra [3020]. HYDRANET [79]. Hypergraph [1954]. hyperledger [2941]. hyperledger-based [2941]. Hypermedia [78, 86, 121, 124, 126–128, 131, 143, 171, 174, 332]. Hypermedia-based [121, 171]. hypertension [2421]. hypertext [129, 172, 454].

I/O [327, 1154, 3193, 3207]. I/O-intensive [327]. IaaS [1241, 1701, 2283, 2416]. IBFD [2851]. IC [3209]. IC-IoT [3209]. ICIC [727]. ICMetric [3070]. ICN [1400, 2940]. ICN-based [2940]. ICN-enabled [1400]. icon [125]. iContainer [3231]. ID [856, 1118, 1164]. ID-based [856, 1118]. ID/locator [1164]. IDA [2154]. IDA-based [2154]. identical [3271]. identifiability [1652]. Identification [503, 512, 526, 554, 568, 578, 620, 686, 690, 1000, 1286, 1449, 1512, 1622, 1807, 1917, 1963, 2002, 2085, 2178, 2453, 2721, 2933, 2975, 3045, 3143, 3308, 3386, 3510]. identify [3271]. Identifying [444, 1671, 1727, 1749, 1870, 2865, 3477]. identities [320]. Identity [196, 485, 517, 609, 713, 772, 811, 1012, 2474, 2775, 3119, 3234]. Identity-based [196, 713, 772, 1012, 2474, 3119]. IDEWEP [396]. IDFC [742]. IDS [3429, 3564]. IEEE [375, 528, 700, 809, 810, 899, 930, 939, 1005, 1149, 1156, 1187, 1320, 1400, 1409, 1443, 1454, 1481, 1525, 1614, 1773, 1783, 1841, 2027, 2092, 2744, 2964, 2976]. II [2137]. IIoT [2140, 2628, 3005, 3186, 3505, 3553]. image [396, 507, 699, 1132, 1320, 1621, 2462, 2733, 2871, 3037, 3165, 3348]. images [317, 443, 3237]. imaging [2280]. imbalanced [3141]. immortal [3231]. Immune [14, 228, 545, 723, 1259, 1439, 2692]. immune-inspired [723]. immunization [3444]. immutability [2952]. iMote2 [1320]. Impact [26, 464, 638, 1382, 1455, 1474, 1498, 1679, 1758, 1882, 1917, 2502, 3066, 3348]. impacting [2322]. impacts [1256, 1367]. impaired [571, 884, 1219]. impatience [1031]. imperfect [787]. Implantable [2178]. Implementation [93, 114, 183, 257, 265, 302, 317, 329, 366, 386, 389, 438, 584, 1400, 1555, 1618, 1789, 2368, 2466, 2481, 2787, 2809, 3061, 3094, 3120, 3238, 3322]. implementations [1285]. Implementing [133, 1438, 2494]. implications [2902]. implicit [2201]. Importance [1031, 1602, 1850]. Imposing [405]. impression [2236]. Improve [434, 530, 620, 809, 866, 913, 967, 1242, 1295, 1491, 1655, 1756, 1862, 1964, 2534, 2561, 3443]. Improved [395, 599, 602, 633, 860, 1136, 1141, 1403, 1741, 1906, 1942, 2136, 2175, 2288, 3208, 3210, 3214, 3428]. Improvement [560, 989, 1023, 1162, 1471, 1511, 1824, 2522, 3453, 3547]. improvements [848, 2005]. Improving [323, 497, 509, 573, 593, 888, 993, 1230, 1285, 1365, 1546, 1730, 1893, 1965, 2155, 2381, 2439, 2452, 2881]. Improvised [2336]. IMS [853, 1860]. IMS-based [853]. IMS-enabled [1860]. in-band [2851]. In-cluster [1257]. In-Memory [1705]. In-network [793, 1895, 2039, 2435]. In-order [800]. in/for [2121]. inband [2223]. incast [1698, 2572]. Incentive [424, 830, 1211, 1228, 1266, 1510, 2820, 2847, 3206, 3215]. incentive-based [251, 424]. INcentiveS [975]. incentivized [2977]. incidents [3476]. included [2595]. income [2308]. incoming [2466]. incomplete [1466, 2473]. inconsistency [1087, 1179]. incorporating [2681]. increase [2280]. increasing [1004]. Incremental [514, 900, 1272, 1314, 1560]. independent [526, 579, 1102, 1721, 1950, 2608, 2835]. Index [23, 51, 72, 91, 117, 182, 195, 215, 751, 1177, 1293, 1632, 2117, 3417]. indexed [1922]. indexing [53, 89, 513]. Individual [512, 3303]. individual-group [3303]. individuals [1727]. Indoor [430, 638, 906, 951, 1297, 1312, 1420, 1601, 1604, 1712, 1869, 2085, 2093, 2703, 2762, 2781]. indoor-outdoor [2085]. induction [2297]. Industrial [521, 1165, 1777, 2073, 2075, 2096, 2142, 2144, 2368, 2369, 2392, 2424, 2450, 2481, 2541, 2783, 2939, 3117, 3318, 3402, 3448, 3505]. Industry [2244, 2273, 3242, 3322, 3450, 3567]. inelastic [733]. Inertial [3292]. Inference [479, 904, 1840, 1855, 1870, 1893, 1981, 2089, 2208, 2395, 2407, 2435, 3378]. inference-based [1840, 1870]. inferential [1995, 2753]. Inferring [3076]. Influence [1613, 1849, 1950, 2132, 2168, 2490, 2927]. Influential [2807]. informal [163]. Information [27, 33, 63, 83, 89, 90, 95, 102, 124, 206, 241, 244, 254, 259, 289, 537, 550, 587, 617, 634, 666, 719, 734, 736, 737, 778, 784, 848, 1006, 1018, 1136, 1173, 1245, 1329, 1390–1394, 1397, 1407, 1414, 1466, 1473, 1476, 1488, 1489, 1560, 1565, 1625, 1719, 1876, 1944, 1954, 1957, 1959, 2109, 2186, 2199, 2301, 2317, 2466, 2473, 2490, 2622, 2691, 2703, 2762, 2989, 3050, 3087, 3088, 3209, 3210, 3248, 3330, 3510, 3511]. information-based [666, 3087]. Information-Centric [537, 1390, 1394, 1488, 1489, 1625, 1876, 2466, 2622, 2989, 3209]. informational [433]. infrastructure [200, 234, 445, 530, 571, 669, 729, 828, 855, 1091, 1241, 1323, 1440, 1561, 1701, 2027, 2107, 2250, 2568, 2893, 2897, 3109, 3297]. infrastructure-as-a-service [3109]. infrastructure-based [828]. infrastructure-less [1440]. Infrastructures [847, 2582, 2583, 2621, 2895, 3288, 3349, 3414]. InFreD [1838]. initiated [2658]. initiative [941]. Injection [2331, 2511, 3048, 3128, 3454]. innovation [2045]. Innovations [299, 300]. innovative [401, 633, 997, 1065, 1310]. inpatient [645]. input [96, 947, 2665]. input-queued [947]. inputs [351]. inquiry [36]. inquiry-learning [36]. insider [1825, 1840]. insights [1049, 2219, 3219]. inspection [911, 1450, 2409, 2481, 2661, 3232]. inspired [532, 607, 614, 723, 838, 922, 1107, 1130, 1845, 1946, 2129, 2136, 2321, 2370, 2371, 2938]. instance [648, 2601]. instances [1494, 1658]. instantaneous [848]. instantiation [1145]. INstantly [257]. instruction [28, 87]. instructional [18]. Instructors [97]. insulated [713]. INT [3350]. INT-based [3350]. integer [30]. Integrated [32, 52, 138, 230, 333, 937, 969, 987, 1165, 1388, 1392, 1983, 2015, 2247, 2281, 3182, 3324, 3422, 3469]. Integrating [25, 247, 323, 2196, 2856]. Integration [478, 651, 663, 755, 997, 1068, 1386, 1684, 2034, 2960, 3155, 3172, 3498, 3544]. Integrity [548, 1662, 1826, 2212, 2327, 2474, 3043, 3363]. Intelligence [241, 293, 766, 881, 914, 1129, 1710, 1951, 2676, 2870, 3025, 3117, 3159, 3519]. Intelligence-based [293]. Intelligent [8, 15, 25, 105, 210, 211, 228, 248, 325, 370, 689, 720, 850, 997, 1065, 1143, 1334, 1743, 1838, 1933, 1942, 2143, 2612, 2690, 2799, 2917, 3262, 3318, 3371, 3450, 3476]. Intensive [314, 327, 1066, 1320, 1533, 2539, 2788, 2951, 3193]. intent [1063, 3021]. intent-driven [3021]. intention [151, 703]. Inter [523, 615, 886, 1087, 1786, 1888, 1936, 2466, 3373]. inter-domain [886, 1087, 2466, 3373]. INTER-IoT [1888]. inter-M2M [1786]. Inter-operation [1936]. inter-PMIPv6-domain [523]. inter-prediction [615]. interacting [171]. Interaction [96, 99, 161, 272, 577, 746, 1073, 1106, 2896]. interactions [122, 282, 652].

Interactive [3, 28, 143, 144, 261, 366, 519, 531, 540, 574, 747, 2885]. INtERCEDE [677]. Intercepting [3020]. interchange [332]. interconnect [876]. interconnected [216]. Interconnecting [2191]. Interconnection [100, 283, 375, 994, 1233, 1760]. interdependence [2121]. Interest [493, 1892, 2225, 2604, 3195, 3248]. Interests [2169]. interface [38, 73, 101, 143, 348, 487, 2408, 2680, 3362, 3530]. interfaces [96, 972, 2773]. Interfacing [66, 226]. Interference [556, 638, 936, 1419, 1739, 1787, 1911, 1927, 2375, 2728, 3124, 3511]. interleaved [973]. intermediary [187]. intermediary-centric [187]. intermediate [1095]. intermittent [1188, 2191]. Intermittently [1240]. Internal [2491]. Internet [68, 71, 73, 94, 107, 123, 148, 154, 213, 287, 301, 323, 424, 532, 555, 701, 796, 853, 912, 953, 995, 1001, 1024, 1073, 1105, 1193, 1260, 1383, 1637, 1673, 1684, 1734, 1752, 1763, 1839, 1888, 1904, 1978, 1987, 1989, 1992, 1998, 2068, 2070, 2096, 2139, 2241, 2310, 2328, 2365, 2419, 2429, 2480, 2481, 2487, 2492, 2503, 2507, 2512, 2523, 2527, 2531, 2563, 2571, 2578, 2636, 2672, 2692, 2739, 2742, 2783, 2789, 2806, 2845, 2848, 2851, 2878, 2898, 2994, 3042, 3087, 3126, 3134, 3171, 3206, 3209, 3287, 3303, 3318, 3354, 3383, 3410, 3445, 3448, 3463, 3492, 3505, 3527, 3528, 3537, 3569]. Internet [179, 260, 376, 949, 965, 1088, 1090, 1177, 1185, 1188, 1261, 1292, 1498, 1692, 1924, 1988, 1990, 1991, 1995, 2091, 2181, 2249, 2325, 2343, 2349, 2387, 2412, 2475, 2525, 2562, 2722, 2820, 2856, 2964, 3118, 3250, 3309, 3360, 3418]. Internet-based [94, 701]. Internet-of-Things [2503, 2563, 2789, 3171, 3209, 3354]. Internet-supported [107]. interoperability [1885, 1888, 2492, 2941, 3552]. Interoperable [824, 1559, 1883, 1886, 2196]. interoperation [756]. intersection [1324]. intersections [1944]. intertwining [305]. interval [261, 750, 2668]. intervehicle [343]. intervention [231]. interworking [478, 1275]. intra [615, 1516]. intra- [615]. intra-LMA [1516]. Intrabody [2973]. intradomain [1195]. Introducing [78]. introduction [2254]. Intruder [150, 1675]. Intrusion [211, 246–248, 262, 285, 289, 382, 436, 470, 479, 666, 933– 935, 1152, 1259, 1459, 1596, 1599, 1661, 1746, 1776, 1818, 1825, 1870, 1924, 1971, 2193, 2484, 2494, 2596, 2723, 2755, 2815, 2934, 2997, 3009, 3033, 3053, 3092, 3094, 3267, 3309, 3318, 3342, 3399, 3402, 3404, 3429, 3486, 3528, 3529, 3533]. invalidation [499]. invariants [2611]. inverse [2263]. Inverted [1293]. Investigating [209, 2848]. Investigation [1079, 1345]. Investigations [1409, 3492]. involving [1160, 2730]. IoE [3184]. IoMT [2886]. IoNT [3284]. IoT [1073, 1883, 1885, 1887, 1888, 1901, 1988, 1996, 1997, 2184, 2255, 2267, 2311, 2321, 2323, 2324, 2369, 2373, 2378, 2385, 2388, 2390, 2415, 2419, 2456, 2460, 2506, 2515, 2516, 2566, 2568, 2591, 2595, 2607, 2639, 2652, 2695, 2699, 2722, 2723, 2751, 2797, 2821, 2830, 2843, 2858, 2861, 2862, 2866, 2880, 2899, 2909, 2910, 2918, 2923, 2931, 2932, 2946, 2956, 2957, 2959, 3038, 3044, 3054, 3062, 3091, 3103, 3115, 3121, 3125, 3151, 3159, 3168, 3173, 3179, 3209, 3218, 3219, 3245, 3258, 3285, 3289, 3297, 3330, 3347, 3362, 3383, 3411, 3437, 3453, 3465, 3473, 3498, 3517, 3523, 3524, 3545, 3553, 3560]. IoT [1625, 1743, 1994, 2072, 2202, 2290, 2316, 2347, 2363, 2372, 2435, 2476, 2519, 2554, 2556, 2628, 2638, 2669, 2714, 2724, 2754, 2784, 2814, 2826, 2834, 2897, 2960, 2965, 3080, 3138, 3213, 3260, 3308, 3361, 3363, 3374, 3440]. IoT-based [1743, 2866, 2899, 2956, 3258]. IoT-Enabled [2456, 2516, 2568, 3260, 3440]. IoT-ML [3080]. IoT-Oriented [2363]. IoT/IIoT [3553]. IoT/LTE [3297]. IOTA [3152]. IoUT [3287]. IoV [3084, 3173]. IP [181, 216, 233, 356, 377, 532, 582, 583, 692, 769, 819, 941, 972, 1215, 1387, 1406, 1472, 1487, 1622, 1627, 1793, 1808, 2155, 2362, 2452, 3011, 3084, 3298]. IP-based [692, 769]. IP-forwarding [972]. IP-over-WDM [1387]. IP/DWDM [532]. IP/MPLS [1472]. IP/MPLS-over-flexgrid [1472]. iPersea [1730]. IPFS [3494]. IPTV [782, 1113, 1445]. IPv4 [329, 360]. IPv6 [329, 360, 774, 937, 1078, 1096, 1174, 1255, 1516, 1581, 1726, 1768, 2372]. IPv6-enabled [1096]. IPv6-IoT [2372]. iris [513]. irregularity [1408]. isolated [2191]. isolation [237]. isomorphism [3264]. ISP [1690, 1717]. Issue [31, 81, 86, 118, 119, 141, 241, 333, 408, 501, 503, 590, 821, 870, 1335, 1390, 1531, 1948, 2017, 2214, 2506, 2789]. issued [1135]. Issues [83, 89, 96, 125, 148, 167, 553, 766, 799, 895, 939, 1204, 1248, 1277, 1315, 1325, 1385, 1386, 1506, 1508, 1684, 1731, 1795, 1801, 1829, 1856, 2010, 2075, 2079, 2080, 2128, 2212, 2322, 2507, 2551, 2554, 2628, 2695, 2709, 2808, 2845, 2887, 2930, 3019, 3068, 3079, 3242, 3272, 3299, 3323, 3408, 3441, 3502]. item [653, 1719, 2921]. items [2469]. iteration [947]. iterations [1117]. itineraries [355]. IWSN [2075].

J [868, 1412, 1765, 1766, 2339, 2516, 2587, 2661, 2888, 2987]. J2ME [186]. Jaccard [1200]. Jackson [1031]. JamholeHunter [3551]. jammers [2761]. jamming [592, 1588, 2319, 2346, 2961]. January [2616, 2629, 2853, 2872, 3081, 3239, 3251, 3405]. Janus [435]. Java [711, 2851]. JavaSim [2851]. JavaSim-IBFD-CRNs [2851]. JDER [1200]. jitter [1226]. JNCA [31, 2102]. Job [1416, 2651, 2697, 3425]. Job-progress-aware [2697]. join [2260]. Joint [612, 679, 728, 773, 1005, 1238, 1301, 1338, 1619, 1762, 1866, 1934, 2240, 2482, 2495, 2601, 2609, 2679, 2883, 2921, 2922, 2974, 2987, 3174, 3320, 3424]. Journal [241, 501, 2763]. JPAS [2697]. juice [1702]. JuiceCaster [1702]. July [2715, 2726, 2978, 2985, 3147, 3325, 3506]. June [2700, 2707, 2969, 3135, 3311, 3490]. Jungle [2003]. just [693].

Kalman [3416]. KEDGEN2 [1158]. KEPPAN [477]. Kerberos [354, 1860]. Kernel [678, 906, 2796, 2981, 3094]. Kernel-based [906, 2796]. Key [49, 208, 291, 292, 442, 491, 505, 561, 576, 585, 698, 713, 737, 825, 886, 960, 976, 985, 1012, 1070, 1135, 1158, 1255, 1544, 1595, 1612, 1705, 1844, 1898, 1994, 2010, 2080, 2127, 2137, 2289, 2341, 2346, 2351, 2412, 2443, 2554, 2636, 2719, 3070, 3396, 3408]. key-insulated [713]. Key-less [2346]. Key-Value [1705]. keyed [2173]. keypoints [513]. keys [349]. keystroke [1438]. keyword [827, 2412, 2603, 2627, 2776]. keyword-field-free [2776]. keywords [576]. Khorramshahr [1811]. kidney [2264]. kitchen [3537]. KNN [620].

KNN-based [620]. knocking [2071]. Knowledge [6, 10, 121, 161, 194, 231, 294, 338, 366, 477, 761, 794, 1063, 1066, 2980, 3205, 3274]. Knowledge-Aware [794]. Knowledge-based [6, 294, 366]. knowledge-intensive [1066]. Kullback [2822].

L [2240]. L-CAQ [2240]. LA-EEHSC [1337]. Label [1252, 2462, 3228]. labels [3253]. labor [1165]. laboratory [683]. lack [823]. LaCoDa [1916]. Lagged [1469]. Lagrange [1324]. lake [3556]. LAM [2638]. LAM-CIoT [2638]. LAN [105]. land [2914]. landscape [3218, 3448]. Language [113, 115, 126, 469, 1213]. Languages [45, 86]. LANs [1974, 3108]. LAPGN [2301]. Large [79, 100, 110, 517, 603, 929, 948, 977, 1050, 1082, 1208, 1299, 1421, 1463, 1771, 1808, 1911, 1962, 2008, 2028, 2093, 2165, 2238, 2292, 2512, 2529, 2676, 2786, 2813, 3250, 3302, 3399, 3420, 3421]. large-scale [79, 100, 517, 603, 929, 948, 1050, 1082, 1299, 1421, 1463, 1808, 1911, 1962, 2165, 2238, 2786, 2813, 3250, 3302, 3399, 3420]. large-sized [2008]. Large-volume [2676]. laser [42]. last [1700, 3294]. last-mile [3294]. Latency [466, 944, 1146, 1227, 1236, 1336, 1514, 1728, 1847, 2189, 2250, 2272, 2513, 2578, 3109, 3167, 3518]. Latency-aware [1847, 2250]. latency-sensitive [1236]. lateral [3444]. Latest [1448]. launching [780]. law [1193]. laws [802, 1010]. laws-based [1010]. Layer [186, 190, 342, 393, 415, 419, 445, 600, 624, 693, 727, 733, 768, 871, 1020, 1036, 1280, 1323, 1343, 1348, 1403, 1521, 1765, 1797, 1891, 1900, 2018, 2086, 2234, 2284, 2372, 2550, 2569, 2624, 2861, 3027, 3074, 3201, 3243, 3247]. layer-3 [393]. Layer-two [3247]. Layered [154, 319, 423, 566, 658, 1299, 1792, 1916, 2424, 2683]. LBSNs [1463, 2192]. LCC [915, 2462]. LDoS [3525]. LEACH [986, 1044]. LEACH-Based [1044]. lead [3268]. leakage [1605, 2580]. learn [228]. Learnable [207]. Learned [28, 2989]. learners [26]. Learning [7, 14, 26, 36, 47, 63, 93, 122, 124, 125, 163, 236, 238, 262, 297, 361, 421, 469, 510, 540, 555, 760, 979, 1076, 1213, 1220, 1272, 1296, 1337, 1437, 1469, 1808, 1864, 2048, 2171, 2236, 2269, 2338, 2415, 2462, 2515, 2542, 2632, 2657, 2690, 2702, 2723, 2733, 2739, 2765, 2789, 2790, 2796, 2800, 2815, 2819, 2871, 2917, 2919, 2942, 2971, 2997, 3000, 3006, 3045, 3053, 3058, 3062, 3145, 3166, 3192, 3194, 3202, 3213, 3255, 3270, 3295, 3319, 3336, 3338, 3351, 3357, 3370, 3395, 3402, 3418, 3422, 3424, 3453, 3462, 3470, 3487, 3514, 3525, 3526, 3528, 3533, 3534, 3545, 3550, 3564, 3565]. learning [26, 81, 85, 98, 118, 119, 121, 137, 319, 529, 766, 989, 1116, 1484, 1928, 1955, 2143, 2330, 2347, 2421, 2597, 2645, 2688, 2697, 2722, 2755, 2799, 2859, 2929, 2958, 2965, 2974, 2987, 3002, 3048, 3085, 3088, 3093, 3149, 3182, 3191, 3198, 3226, 3233, 3269, 3275, 3305, 3308, 3339, 3344, 3361, 3382, 3397, 3426, 3429, 3440, 3445, 3451, 3510, 3539, 3547]. learning-based [2048, 2171, 2632, 2929, 3062, 3202, 3233, 3338, 3382, 3440, 3470]. learning-powered [3528]. learning-resistant [3361]. leasing [395, 3115]. Least [2029]. LED [3369]. LED-based [3369]. Ledger [2923, 2977, 3279, 3481]. ledger-based [3279]. Left [2291]. legality [979].

Leibler [2822]. Leisure [1169]. LENA [3322]. length [317, 635, 662, 1308, 2295]. LEO [1503]. less [530, 621, 1440, 2040, 2346]. Lessons [28, 577]. letter [19]. letter-driven [19]. Level [130, 176, 217, 239, 264, 391, 568, 569, 588, 622, 757, 765, 1288, 1370, 1575, 1596, 1727, 2000, 2172, 2827, 2975, 3196]. leveling [1620]. levels [1107, 2470]. Leveraging [629, 1278, 2127, 2490, 2773, 3031, 3559]. lexical [2082]. LFA [2754]. LIANA [1314]. libraries [31, 33, 34]. library [35, 37, 298]. license [315]. licensing [420]. LIDDE [2927]. life [648, 1019, 1105, 3108, 3426]. life-based [1019]. life-saving [3108]. lifecycle [2834]. Lifespan [1476, 1617, 1968, 2585]. Lifetime [250, 720, 863, 968, 1137, 1318, 1363, 1377, 1504, 1633, 2338, 2348]. Light [1603, 1894, 2249, 2323, 2533, 2766, 3059]. Light-weight [1603, 1894, 2766]. lighting [938, 2695]. Lightweight [385, 558, 909, 915, 1255, 1508, 1990, 2176, 2185, 2333, 2441, 2443, 2522, 2607, 2638, 2719, 2731, 3038, 3055, 3119, 3383, 3464, 3553]. like [1430]. likelihood [1671, 2090]. limitation [1356]. limited [345, 920, 1509]. LIMOS [2468]. line [98, 557, 1476, 1558, 1794, 2760]. Linear [47, 514, 594, 708, 1890]. linguistic [1727]. Link [275, 556, 598, 685, 920, 926, 996, 1036, 1092, 1155, 1247, 1351, 1417, 1419, 1505, 1706, 1756, 1820, 2022, 2151, 2240, 2330, 2631, 2770, 2844, 3007, 3270, 3474]. link-based [2022]. link-oriented [2240]. link-scheduling [3270]. link-stability [1351]. link-state [1155]. links [712, 963, 1558, 1615, 1758, 2191, 2231]. Linux [484, 972, 1273, 1464, 3094]. Linux-based [1464]. LISP [2288]. List [401, 744, 3376]. listener [1431]. lists [1441]. literature [1304, 1317, 1332, 1736, 2068, 2211, 2430, 2592, 2692, 2818, 3218, 3427, 3481]. Live [397, 656, 747, 1159, 1314, 2005, 2179, 2207, 2314, 3192, 3409]. liveness [82, 3105]. LiveSD [990]. LMA [1516]. LMAAS [3038]. LMAAS-IoT [3038]. Load [327, 383, 459, 575, 630, 770, 779, 786, 966, 1033, 1443, 1484, 1548, 1653, 1705, 1736, 1761, 1784, 1833, 1881, 1915, 1920, 1943, 1980, 1982, 2081, 2097, 2198, 2255, 2416, 2569, 2644, 2651, 2672, 2864, 2876, 3002, 3005, 3157, 3358, 3367, 3425, 3485]. load-adaptive [2864, 3157]. Load-balanced [786, 2198, 2569]. Load-balancing [630, 1980, 3005]. loading [614]. loads [1882]. Local [741, 959, 1194, 1353, 1597, 2210, 2740, 2927, 3104, 3210, 3381]. local-influence-descending [2927]. localisation [1312]. Locality [515, 1081, 1664, 3198]. Locality-based [3198]. Localization [583, 638, 639, 951, 1191, 1256, 1274, 1283, 1297, 1362, 1495, 1576, 1604, 1633, 1712, 1833, 1917, 2076, 2085, 2090, 2153, 2230, 2577, 2703, 2750, 2762, 2781, 2821, 2859, 2918, 3188, 3428]. Localization-free [1191]. Localized [818]. locally [2790]. locating [489, 2602]. Location [273, 414, 604, 645, 664, 740, 1222, 1379, 1433, 1482, 1608, 1617, 1682, 1770, 1813, 1814, 1872, 1903, 1945, 1952, 1988, 2055, 2134, 2135, 2192, 2216, 2292, 2374, 2387, 2526, 2887, 2953, 3114, 3244, 3321]. Location-aware [2055, 3244]. Location-based [664, 1379, 1433, 2387]. location-free [1222]. location-privacy [2135]. location-sharing [1952]. location-wise [1617]. locations [147, 2134]. locator [1164]. logarithm [196]. logic [805, 1354, 2163, 2205, 2282]. logical [1264]. logistic [2611, 2840]. logistics [641, 1089]. logs [2570, 2736, 3236]. Long [1219, 1550, 1633, 2512, 2529]. long-haul [2529]. Long-term [1550, 2512]. long-thin [1633]. longitudinal [3224]. look [1224]. look-up [1224]. Lookup [321, 336, 907, 988, 1406, 1610, 1742, 1768, 1969, 2989]. loop [616, 2643]. loop-free [2643]. loose [301]. LoRa [3530]. loss [274, 352, 434, 782]. losses [1560]. lossless [1353, 3203]. Lossy [1311, 2748, 3198]. lost [1262]. Lostrego [2107]. Low [683, 772, 944, 1132, 1311, 1336, 1410, 1508, 1514, 1558, 1601, 1651, 1728, 1800, 1884, 2165, 2308, 2441, 2513, 2541, 2607, 2620, 2718, 2748, 2784, 2806, 2892, 3002, 3018, 3167, 3223, 3292, 3514, 3518]. Low-cost [1132, 1601, 2441, 2718, 2784]. Low-end [2607]. Low-latency [2513, 3518]. low-power [1410, 1884]. low-precision [2165]. low-rate [3514]. low-resource [1508]. low-voltage [1558]. LPCN [2029]. LR [2975]. LR-WPAN [2975]. LSTM [3048]. LSTM-Autoencoder [3048]. LTE [727, 1343, 1344, 1405, 1511, 1598, 1630, 1765, 1788, 1836, 1908, 1943, 2074, 2179, 2184, 2226, 2375, 2675, 3030, 3131, 3297]. LTE-A [1511, 1943, 2184, 2226, 2675]. LTE-Advance [2179]. LTE-Advanced [727, 1598, 1630]. LTE-U [3030]. LTE-U/NR-U [3030]. LTE/LTE [1943, 2675]. LTE/LTE-A [1943, 2675].

m [74, 186, 433, 866, 899]. m-commerce [186]. m-informational [433]. M/G/1 [74]. M2M [1282, 1362, 1666, 1691, 1786, 2039, 2559, 3483]. MA [303]. MAC [234, 371, 1020, 1142, 1345, 1481, 1514, 1521, 1647, 1735, 1800, 1859, 1891, 1900, 1974, 2220, 2322, 2590, 2624, 2864, 3018, 3369]. MAC-layer [1900]. Machine [367, 656, 702, 989, 999, 1003, 1307, 1415, 1546, 1666, 1667, 1832, 1871, 1913, 1934, 2048, 2207, 2242, 2313, 2542, 2645, 2657, 2690, 2702, 2722, 2789, 2799, 2819, 2871, 2942, 2951, 2958, 3000, 3045, 3053, 3088, 3162, 3166, 3182, 3193, 3269, 3361, 3482, 3499, 3510, 3514, 3522, 3528, 3547]. machine-based [999]. machine-learning [989]. machine-learning-based [3510]. Machine-to-Machine [702, 1666, 1913]. machines [498, 1180, 2484, 2621, 2837]. macroblock [693]. macroprocessing [67]. MADDPG [3388]. MADDPG-empowered [3388]. Mahalanobis [2796]. Mahbub [3291]. Mahjong [1169]. mail [56, 393, 401, 1066]. mails [703]. main [879, 1325, 1529]. Maintaining [3187]. maintenance [155, 357, 762, 3095, 3096]. Major [1644]. majority [3564]. MAKA [2183]. make [3009]. Makespan [2907, 3216]. Makespan-aware [3216]. Making [304, 1494, 1553, 1886, 2430]. MalDetect [3370]. Malicious [548, 956, 1854, 1960, 2645, 3045, 3141]. MalInsight [2384]. MAlleable [3234]. malware [987, 1131, 1785, 2009, 2067, 2144, 2384, 2459, 2580, 2586, 2657, 3370]. Mama [646]. Man [375, 1369]. Man-At-The-End [1369]. manage [1855]. Managed [3158]. Management [13, 76, 102, 105, 110, 149, 184, 267, 291, 361, 365, 406, 412, 419, 458, 544, 547, 562, 596, 648, 740, 769, 835, 836, 852, 855, 861, 870, 916, 936, 985, 1036, 1070, 1173, 1198, 1220, 1241, 1255, 1282, 1296, 1307, 1397, 1404, 1425, 1460, 1486, 1489, 1541, 1582, 1690, 1713, 1739, 1811, 1822, 1825, 1831, 1878, 2010, 2015, 2030, 2035, 2037, 2109, 2218, 2225, 2228, 2271, 2310, 2378, 2397, 2455, 2512, 2515, 2555, 2634, 2636, 2775, 2782, 2875, 2906, 2914, 2948, 2957, 2959, 3091, 3112, 3124, 3150, 3166, 3170, 3173, 3234, 3269, 3386, 3396, 3419, 3422, 3473]. management [161, 292, 369, 414, 491, 517, 549, 623, 641, 646, 834, 894, 902, 1020, 1022, 1148, 1235, 1251, 1261, 1359, 1360, 1388, 1481, 1619, 1625, 1740, 1871, 1945, 1990, 2103, 2182, 2351, 2398, 2425, 2479, 2658, 3189, 3476]. Managing [84, 158, 335, 547, 1265, 2027]. MANEMO [878]. MANET [359, 539, 757, 889, 2224, 2446, 2650]. MANETs [557, 724, 733, 841, 857, 918, 943, 996, 1156, 1184, 1310, 1321, 1435, 1449, 1571, 1615, 1742, 1865, 2127, 2344, 2885]. manifold [510, 949, 1116]. manipulation [271]. manner [780]. MANs [696]. Manufacturing [1322, 1534]. many [2599]. many-field [2599]. map [3063]. MAPLE [2542]. mapped [21]. mapping [603, 1091, 1092, 1164, 1525, 1706, 1869, 2073, 2162, 2210, 2584, 3116, 3162, 3390, 3469]. MapReduce [1491, 2903]. maps [593]. MAQD [1981]. March [2654, 2662, 2912, 2924, 3100, 3265, 3446]. MARIO [2752]. maritime [790, 2024]. marker [619, 866, 919, 974]. Market [622, 980, 1715, 3243]. Market-based [980, 3243]. marketplaces [3154]. markets [187, 3218]. marking [288, 1593, 2649, 2884]. marking-aware [2649]. Markov [1000, 1841, 1865, 2431, 2777, 3482]. Markovian [1708]. markup [427]. MARS [2880]. mashups [1160]. mask [3434]. Massive [1916, 2496, 3134, 3523]. massively [466, 481, 2545]. Master [84, 2843]. Master-slave [2843]. match [3378]. matching [508, 1450, 1462, 1649, 2350, 3028, 3232, 3294, 3314, 3494]. matchmakers [610]. matchmaking [358, 563]. MATEM [2148]. mathematical [728, 1568]. matheuristic [2740]. matrix [382, 546, 670, 973, 1058, 1313, 1387, 1487, 1808]. matter [1637]. mature [3288]. max [1447]. max-flow [1447]. maximal [777, 3496]. maximization [901, 923, 968, 1613, 1858, 1950, 2132, 2256, 2490, 2927, 3132]. maximize [843, 1318, 2348]. Maximizing [978, 1504, 1771, 2356, 2360]. Maximum [352, 559, 955, 1500, 2090, 2791]. May [2685, 2693, 2954, 2962, 3122, 3300, 3479]. MBone [62]. MC [2334, 3546]. MC-enabled [3546]. MC-VAP [2334]. MCC [1802, 3489]. MCDM [2148]. MCOP [193]. MCS [3437]. MCTE [3524]. MCTE-RPL [3524]. MDiag [945]. MDP [1973]. mean [3527]. mean-field [3527]. means [238, 1553]. measure [1849]. Measurement [402, 465, 873, 1208, 1558, 1759, 2164, 2324, 2631, 2675, 3292, 3338, 3518]. measurements [1755, 2323, 2493, 3426]. Measuring [1003, 3103]. MEC [2959, 3375, 3489, 3512]. MEC-assisted [2959]. MEC-MCC [3489]. mechanism [154, 252, 356, 392, 445, 447, 450, 495, 630, 700, 768, 781, 830, 956, 959, 1054, 1082, 1087, 1115, 1122, 1171, 1189, 1196, 1211, 1228, 1242, 1360, 1409, 1525, 1589, 1591, 1742, 1840, 1943, 1964, 2201, 2396, 2416, 2485, 2486, 2562, 2619, 2638, 2653, 2658, 2674, 2735, 2741, 2765, 2820, 2843, 2882, 2920, 2974, 2987, 2990, 3096, 3108, 3215, 3245, 3334, 3378, 3500, 3543, 3554]. Mechanisms [86, 126, 127, 131, 764, 822, 899, 967, 1239, 1422, 1431, 1451, 1510, 1612, 1736, 1787, 1819, 1880, 2042, 2068, 2105, 2312, 2502, 2943, 3022, 3028, 3102, 3114, 3448]. MedHypChain [2941]. media [58, 164, 253, 255, 581, 747, 1341, 1727, 1867, 3250]. mediated [652, 1397]. Medical [443, 1840, 2178, 2462, 2941, 3097, 3126]. Medicine [1959, 2421]. Medium [813, 978, 1246, 1786, 1894, 2149, 2377, 2460, 2639]. meet [3502]. meeting [759, 876]. meets [2728, 3269]. MEGA [1320]. member [2658]. membership [2496]. Memory [429, 599, 789, 1406, 1529, 1610, 1620, 1705, 1815, 2005, 2943, 3197]. memory-based [1620]. Memory-efficient [1406]. MEMS [1264]. mention [3104]. merge [2024]. merging [1183, 1406]. Merkle [1922, 2185]. Mesh [264, 556, 590–592, 594–598, 601, 689, 728, 790, 792, 864, 1005, 1020, 1090, 1095, 1110, 1115, 1141, 1166, 1181, 1247, 1470, 1588, 1618, 1753, 1779, 1866, 1967, 2026, 2162, 2883, 2988, 3103, 3442]. mesh-based [2162]. mesh/ad [790]. meshed [256]. Message [43, 140, 332, 346, 670, 711, 1018, 1049, 1055, 1512, 1812, 1882, 2501, 3076, 3105, 3415, 3561]. messages [201, 602, 775, 1214, 1505, 2087]. messaging [135, 177, 680]. Meta [130, 877, 2416, 2935]. meta-heuristic [2935]. meta-level [130]. meta-portals [877]. MetaCDN [462]. metadata [138, 312, 427, 458, 952]. metaheuristic [1906]. Metarouting [3548]. Metaverse [3346, 3458]. meteorological [2938]. metering [1323, 1561, 2899]. meters [3061]. method [21, 201, 430, 522, 564, 842, 920, 1071, 1072, 1075, 1154, 1208, 1255, 1366, 1458, 1533, 1534, 1538, 1593, 1596, 1622, 1804, 1833, 1881, 1892, 1898, 1918, 1984, 2097, 2156, 2363, 2409, 2453, 2459, 2461, 2464, 2540, 2550, 2661, 2675, 2679, 2703, 2721, 2755, 2791, 2799, 2953, 3085, 3359, 3389, 3391, 3394, 3534]. methodologies [897, 1213, 3517]. Methodology [444, 536, 551, 2547, 3271]. Methods [54, 333, 583, 721, 1254, 1318, 1492, 1703, 1759, 1813, 1842, 2071, 2294, 2509, 2583, 2597, 2815, 3014, 3156, 3171, 3395, 3404]. meticulous [1246]. Metric [740, 1190, 1655, 1696, 2395]. Metric-based [740, 1190]. Metrics [363, 598, 805, 1623, 2583, 2625, 2824, 2825, 2831]. metropolitan [611, 696, 2464]. MHT [1922]. MHz [1417]. MIB [105]. MicaZ [1320]. micro [92, 937, 1555, 2827]. micro-blog [2827]. Micro-Grid [1555]. Micro2vec [3236]. microblogging [1957]. microcell [481]. Microcomputer [15, 20]. microcomputer-based [20]. microcontroller [29]. microdata [3568]. microeconomics [1623]. microeconomics-based [1623]. microprocessor [22]. microservice [2304, 2566, 2601, 2653]. microservices [2711, 3236]. Microsoft [84]. Middlebox [1847]. Middleware [199, 253, 415, 418, 534– 536, 539, 541, 542, 630, 647, 733, 763, 778, 794, 824, 1276, 1370, 1568, 2202, 2566]. MIDS [27]. Mieso [789]. Migration [98, 656, 1171, 1232, 1415, 1636, 1832, 2005, 2207, 2729, 2922, 2974, 2987, 3043, 3349, 3409, 3443]. MIH [2015, 2372]. MIH-base [2372]. MIH-FPMIPv6 [2015]. MIH-SPFP [2372]. mile [3294]. millimeter [2883]. milling [1074]. Mimic [3290]. MIMO [787, 1348, 2213, 3188]. mindsets [47]. Miner [1085]. mines [2170]. Minimalist [674]. minimise [1233]. minimization [1349, 1769, 1928, 3330, 3353]. minimize [1011, 1520, 2482]. minimized [3415]. Minimizing [274, 1927, 2086, 2242, 2931, 2991]. Minimum [739, 1973, 2405, 2529, 2791]. Mining [69, 210, 247, 306, 498, 1066, 1286, 1703, 1852, 1892, 2160, 2262, 2454, 2473, 2752, 3236, 3298, 3517]. minislot [1110]. Minutiae [505, 506]. minutiae-based [506]. MIRACLE [1721]. misbehaving [134, 734]. misbehavior [845]. misinformation [2991]. miss [833]. missing [1262, 2441, 2451, 2713]. mission [2465, 2988, 3208]. mission-critical [2988, 3208]. missions [3321]. MIST [1901]. mitigate [1247, 1367]. Mitigating [1608, 1830, 2156, 2540, 2580, 2754, 3355]. Mitigation [466, 1383, 1441, 1515, 1545, 1687, 2346, 2382, 2429, 2961, 2996, 3093, 3146, 3159, 3191, 3293, 3474, 3486]. mixed [363, 791, 3167, 3462]. mixed-criticality [3167]. mixture [2777, 3367]. ML [423, 3080, 3126, 3166]. ML-Chord [423]. MMIO [3472]. MMR [939]. mMTC [2120]. mmWave [3563]. Mobile [12, 191, 254, 326, 355, 365, 386, 387, 409, 415, 442, 541, 550, 587, 625, 644, 661, 692, 694, 721, 724, 764, 769, 835, 846, 872, 887, 940, 948, 973, 990, 1001, 1133, 1146, 1157, 1172, 1188, 1196, 1204, 1212, 1232, 1275, 1285, 1335, 1354, 1361, 1385, 1418, 1424, 1426, 1434, 1506, 1536, 1583, 1595, 1633, 1669, 1676, 1738, 1748, 1751, 1766, 1872, 1875, 1885, 1895, 1939, 1946, 2105, 2133, 2164, 2200, 2265, 2288, 2289, 2303, 2320, 2342, 2372, 2381, 2419, 2432, 2471, 2486, 2495, 2880, 2914, 2968, 3010, 3019, 3112, 3129, 3142, 3178, 3217, 3272, 3291, 3427, 3449, 3551]. mobile [83, 252, 279, 307, 403, 407, 448, 461, 473, 475, 480, 499, 511, 536, 538, 558, 565, 579, 622, 673, 682, 711, 723, 729, 757, 758, 760, 767, 851, 871, 876, 880, 917, 926, 937, 944, 1101, 1121, 1131, 1186, 1264, 1266, 1286, 1340, 1351, 1362, 1373, 1407, 1408, 1455, 1471, 1475, 1477, 1510, 1541, 1575, 1606, 1614, 1628, 1648, 1743, 1761, 1826, 1882, 1896, 1925, 1951, 1981, 1993, 2033, 2049, 2071, 2122, 2179, 2236, 2240, 2258, 2295, 2316, 2335, 2378, 2423, 2438, 2472, 2527, 2565, 2581, 2600, 2922, 2961, 2990, 3062, 3063, 3105, 3114, 3127, 3156, 3383, 3387, 3421, 3468, 3502, 3526]. mobile [135, 292, 303, 472, 474, 520, 529, 548, 549, 551, 591, 662, 665, 685, 770, 802, 891, 939, 1009, 1029, 1098, 1155, 1174, 1225, 1256, 1276, 1278, 1355, 1365, 1377, 1481, 1526, 1603, 1607, 1608, 1696, 1709, 1712, 1819, 1852, 1952, 1953, 1977, 2091, 2135, 2173, 2330, 2355, 2459, 2461, 2619, 2653, 2788, 2807, 2867, 2929, 2975, 3029, 3215, 3261, 3305, 3320, 3374, 3394, 3539]. mobile-edge [3156]. mobile-health [1895]. mobile-to-mobile [558]. MObility [233, 337, 357, 369, 468, 673, 769, 891, 894, 937, 945, 957, 1024, 1026, 1029, 1036, 1235, 1254, 1256, 1275, 1342, 1344, 1379, 1404, 1431, 1440, 1486, 1516, 1576, 1582, 1595, 1636, 1751, 1846, 1945, 2014, 2015, 2049, 2094, 2149, 2183, 2271, 2321, 2411, 2752, 2803, 2914, 3031, 3112, 3115, 3214, 3244, 3289, 3349, 3387, 3464, 3475]. Mobility-assisted [945, 1576, 2411]. Mobility-Aware [2049, 2149, 3349, 3387]. mobility-enabled [3244]. modal [316]. modality [1097]. mode [695, 858, 1534, 1917, 2271, 3289]. mode-shape [1917]. Model [10, 45, 75, 116, 124, 233, 261, 319, 349, 353, 354, 360, 368, 378, 382, 406, 407, 423, 439, 442, 488, 557, 570, 580, 584, 667, 672, 678, 686, 716, 738, 746, 753, 756, 803, 832, 842, 907, 980, 1001, 1024, 1032, 1049, 1070, 1075, 1134, 1138, 1214, 1280, 1419, 1465, 1516, 1536, 1549, 1556, 1568, 1600, 1635, 1648, 1679, 1696, 1714, 1722, 1745, 1825, 1841, 1950, 2002, 2032, 2062, 2082, 2092, 2160, 2172, 2297, 2309, 2336, 2344, 2345, 2355, 2362, 2365, 2370, 2418, 2431, 2543, 2586, 2587, 2625, 2668, 2766, 2781, 2786, 2863, 2888, 2965, 3030, 3033, 3052, 3096, 3130, 3267, 3296, 3487, 3525]. model [197, 218, 405, 435, 719, 949, 1000, 1064, 1086, 1235, 1456, 2056, 2103, 2421, 2468, 2736, 3286, 3303, 3557]. model-based [667, 2032, 3525]. Model-driven [753, 842, 1032, 1075, 1714]. model-free [2965]. Modeling [6, 165, 248, 346, 361, 705, 823, 835, 842, 885, 1058, 1067, 1069, 1071, 1302, 1410, 1440, 1478, 1483, 1554, 1683, 1843, 1910, 1957, 1971, 2215, 2226, 2247, 2337, 2370, 2434, 2454, 2753, 2811, 2926, 3040, 3059, 3185, 3222, 3472]. Modelling [24, 238, 929, 994, 1198, 1424, 1563, 1766, 2007, 2329, 2430, 2491, 2531, 2578, 2608]. Models [74, 138, 171, 548, 553, 767, 860, 1029, 1178, 1254, 1256, 1268, 1359, 1408, 1507, 1663, 1759, 1792, 1897, 2014, 2100, 2507, 2509, 2544, 2608, 2634, 2777, 2790, 3019, 3352, 3360, 3381, 3514, 3529, 3533]. modern [194, 1315, 1376, 2608]. modification [2784]. modified [394, 505]. Modular [1968, 2293, 3345]. modularity [1954]. modulation [2011, 2400]. modulator [3350]. module [2895]. monitor [22, 3133]. Monitoring [38, 259, 265, 480, 527, 545, 711, 714, 723, 951, 1109, 1165, 1208, 1445, 1463, 1557, 1917, 1997, 2079, 2170, 2388, 2408, 2440, 2458, 2519, 2612, 2895, 3178, 3440]. Monte [326, 1878, 1968]. Monte-Carlo [1968]. MOO [114]. Moodle [1437]. motes [1320]. motion [818, 1179, 2493, 2734, 2778, 3292]. motion-based [818]. motion-blurred [2778]. Motivation [1369, 1418, 1426, 1597]. mounted [1946, 2881]. move [1796, 1798]. movement [1029, 1283, 3444, 3552]. Moving [539, 846, 1283, 2511, 2825, 3102]. moving-target [2511]. MP3 [183]. MP3-music [183]. MPEG [152, 1525]. MPEG-4 [1525]. mPHASiS [587]. mPillBox [3455]. MPLEX [454]. MPLS [1040, 1252, 2498]. MPLS-over-flexgrid [1472]. MPLS/GMPLS [1040]. MPR [898]. MPTCP [2534, 3039]. MQTT [2879]. MR [808]. MR-BART [808]. MSAR [1696]. MSDP [777]. MSDU [1042]. MsM [2566]. MuDi [1566]. MuDi-Stream [1566]. MULBS [759]. Multi [190, 250, 302, 363, 401, 437, 458, 480, 580, 584, 592, 594, 603, 637, 682, 713, 726, 728, 738, 811, 839, 845, 847, 860, 885, 939, 949, 951, 961, 1019, 1048, 1076, 1100, 1102, 1105, 1120, 1135, 1141, 1179, 1181, 1230, 1322, 1339, 1435, 1493, 1500, 1550, 1566, 1571, 1596, 1618, 1630, 1690, 1724, 1788, 1792, 1831, 1875, 1928, 1961, 2055, 2083, 2084, 2108, 2129, 2133, 2209, 2233, 2257, 2267, 2332, 2364, 2432, 2450, 2463, 2503, 2603, 2704, 2729, 2795, 2928, 3038, 3069, 3071, 3107, 3129, 3196, 3220, 3321, 3369, 3372, 3453, 3459, 3462, 3466, 3467, 3501, 3526, 3532, 3545, 3558, 3563]. multi [163, 258, 300, 304, 319, 405, 423, 435, 459, 464, 470, 555, 609, 695, 722, 750, 800, 808, 855, 997, 1005, 1059, 1096, 1140, 1166, 1263, 1296, 1348, 1439, 1446, 1470, 1481, 1483, 1507, 1591, 1603, 1654, 1733, 1753, 1866, 1981, 1994, 2047, 2210, 2230, 2287, 2334, 2400, 2402, 2415, 2423, 2438, 2443, 2581, 2590, 2621, 2627, 2679, 2714, 2719, 2728, 2827, 2838, 2907–2909, 2931, 2965, 2967, 2974, 2987, 3011, 3132, 3159, 3181, 3226, 3268, 3382, 3384, 3388, 3394, 3422, 3513, 3524, 3530, 3549]. Multi-access [2267, 2795, 2974, 2987, 3107, 3384, 3513]. Multi-agent [300, 302, 304, 319, 584, 847, 1100, 1102, 1105, 1928, 3226, 3422, 3453, 3462, 3526, 3545]. multi-agent-based [250]. multi-application [405, 2907]. Multi-Armed [2402, 3558]. multi-authenticated [713]. Multi-authority [2233, 2332, 2423, 2714]. Multi-Aware [1981]. multi-band [2047]. Multi-Batches [3132]. multi-candidate [722]. multi-channel [592, 728, 885, 1005, 1048, 1096, 1181, 1339, 1470, 1654, 1733, 1753, 1866, 2108, 2450, 2590]. Multi-characteristics [860]. multi-city [3382]. Multi-classification [2133]. multi-classifier [401]. multi-client [3459]. multi-cloud [1550, 2055, 3467]. multi-commodity [2210]. multi-connection [2334]. multi-constrained [1296]. Multi-constraint [363]. multi-context [3524]. multi-controller [3532]. Multi-core [738]. multi-dimensional [470, 855, 1120]. Multi-domain [1961, 2729, 2928, 2967, 3069, 3532]. multi-factor [2909, 3038]. multi-flow [750, 1446, 2364]. multi-gateway [1181, 1994]. multi-grained [405]. Multi-granularity [580, 1322]. Multi-hop [464, 480, 845, 939, 1166, 1348, 1481, 1500, 1507, 1571, 2108, 2209, 3530]. multi-issued [1135]. multi-keyword [2603, 2627]. multi-layered [423, 1792]. MUlti-lead [3268]. Multi-level [1596, 3196]. multi-location [3321]. Multi-manifold [949]. multi-mode [695]. Multi-neural [3129]. Multi-objective [603, 1263, 1439, 2129, 2230, 2257, 2463, 2503, 2621, 2679, 2838, 2931, 2965, 3071, 3159, 3181, 3372]. multi-operator [1788]. multi-partite [3549]. multi-party [2084]. Multi-path [258, 800, 1140, 1493, 1603, 1724]. Multi-point [1435, 2728]. multi-protocol [458]. multi-provider [459]. multi-radio [592, 728, 1005, 1141, 1166, 1181, 1470, 1618, 1654, 1753, 1866]. multi-range [726]. multi-RAT [1591]. Multi-rate [808, 1181]. multi-robots [682]. multi-secret [594]. Multi-sector [726]. multi-sender [437]. multi-sensor [839, 997]. multi-server [609, 811, 1179, 1230, 2443, 2719]. multi-sink [1493]. multi-site [1875]. multi-slot [3369]. multi-stage [435, 2704, 3011]. multi-stream [1630]. multi-subtopic [2827]. multi-swarm [637]. Multi-target [951]. multi-task [2432, 3394]. multi-tenant [1690, 2581, 3466]. multi-tier [555, 961, 1483, 2083, 2287, 3388, 3501]. multi-type [2908]. Multi-UAV [3526, 3563]. multi-unmanned [3220]. multi-user [163, 190, 1076, 2400, 2438, 3372]. multi-view [2415]. Multi-worker [2432]. Multiagent [903, 2519]. Multicast [60, 139, 154, 181, 189, 278, 281, 322, 323, 328, 344, 372, 394, 397, 532, 595, 658, 685, 728, 747, 764, 782, 829, 841, 843, 871, 1019, 1153, 1181, 1182, 1226, 1255, 1269, 1394, 1431, 1615, 1783, 1842, 1866, 2057, 2251, 2658, 2748, 2809, 3051, 3084, 3160, 3169, 3221, 3249, 3544]. Multicast-aware [3051]. multicast-unicast [1255]. Multicasting [1287, 1303, 1837, 1970, 2152]. multichannel [888, 1020, 1912, 2051, 2322, 2683]. multicloud [2257, 2774]. multicriteria [11]. multidata [653]. multidimensional [1969, 2344]. multidisciplinary [1839, 2194]. multihop [1451, 1940, 2625]. multilayer [11, 1472, 3548]. Multimedia [26, 41, 60, 65, 66, 68, 85, 89, 122, 162, 165, 225, 281, 293– 295, 297, 359, 468, 488, 501, 565, 586, 688, 691, 696, 700, 950, 1094, 1097, 1343, 1403, 1518, 1543, 1672, 1708, 1732, 1765, 1802, 1910, 1983, 2155, 2365, 2577, 2634, 2691, 2718, 2846, 2898, 3362]. Multimodal [507, 568, 2626, 2971]. multiobjective [1040, 1234, 2914, 3468]. multipacket [1974]. multiparty [962, 2790]. Multipath [472, 525, 681, 841, 996, 1040, 1281, 1352, 1461, 1532, 1777, 1797, 2189, 2541, 3027, 3276]. multiplayer [2545]. Multiple [188, 449, 479, 487, 511, 581, 611, 758, 770, 783, 944, 963, 972, 973, 1095, 1159, 1226, 1278, 1319, 1438, 1495, 1601, 1614, 1635, 1721, 1773, 1837, 1937, 1949, 1996, 2097, 2131, 2213, 2235, 2243, 2269, 2328, 2333, 2360, 2474, 2508, 2564, 2604, 2656, 2939, 3154, 3186, 3371, 3430, 3546]. multiple-network-interface [487]. Multiple-target [2328]. multiplexer [75]. Multiplexing [2417, 2874]. Multiplexing-Elastic [2874]. multiplication [30]. multipliers [1458]. Multipoint [566, 602, 3423]. multirate [983]. multiservice [559]. multisink [2683]. multisource [1639]. multisourced [2314]. Multistage [994, 1233]. multitask [2971]. multitenant [2396]. multitenant-aware [2396]. multiuser [466]. multivariate [2401]. multiview [2314]. MuSC [2704]. MUSE [3268]. museum [124]. music [183, 970]. MUSiCAL [124]. MUSLI [143]. Mutual [70, 276, 561, 666, 960, 1595, 2273, 2909]. mutual-healing [561]. mutualistic [2921]. MutualRec [2921].

N [1815]. N-FIB [1815]. Name [880, 1610, 1649, 1815, 2156, 2989]. name-based [1815]. Named [1398, 1399, 1469, 1610, 1967, 2004, 2169, 2225, 2383, 2567, 2607, 2650, 2839, 3195, 3289, 3396, 3419, 3555]. Named-Data [3419]. names [1358]. Namespace [2567]. Namespace-based [2567]. naming [1396]. NAND [1425, 1620]. nano [1400, 3284]. nano-sensor [3284]. Nanonetworks [2973]. narrowband [1558, 3383]. NASA [53]. Nash [2223]. NAT [1037]. NationTelescope [1463]. native [65, 1802]. natively [994]. nature [607, 2129, 2370]. nature-inspired [607]. navigate [1601]. Navigating [3560, 3569]. Navigation [125, 884, 1420, 1769, 2867]. Navigation-driven [1769]. NB [2862, 3383]. NB-IoT [2862, 3383]. NBP [2567]. NDN [1964, 2379, 2446, 3120, 3195, 3332, 3558]. NDN-based [3120]. NDNBrowser [1398]. NDNrel [1964]. near [1878]. near-optimal [1878]. nearest [1293]. NectaRSS [370]. needs [2610]. Nefis [3509]. Negative [382, 1367, 1783]. Neglected [3288]. negotiation [304, 461, 980, 1103]. Neighbor [841, 1293, 1378, 1427, 2165, 3497]. neighbor-to-neighbor [2165]. Neighborhood [1457]. neighbors [1080]. neighbour [3055]. neighbourhood [1833]. NeMHIP [1053]. NEMO [957, 1025, 1945]. Neonatal [314]. nested [640, 957, 1025, 1969]. nested-NEMO [1025]. net [1494, 2430, 2668]. net-based [1494]. NetInf [1391]. nets [2982, 3102]. Netw [1412, 1765, 1766, 2339, 2516, 2587, 2661, 2888, 2987]. Network [61, 79, 90, 107, 109, 239, 241, 274, 276, 308, 365, 368, 384, 399, 412, 417, 446, 525, 526, 534, 586, 601, 617, 637, 771, 776, 848, 868, 872, 873, 904, 1021, 1124, 1170, 1202, 1220, 1227, 1269, 1327, 1333, 1392, 1393, 1407, 1409, 1430, 1475, 1492, 1512, 1556, 1557, 1562, 1564, 1572, 1627, 1639, 1642, 1674, 1846, 1895, 2054, 2162, 2225, 2229, 2252, 2380, 2406, 2441, 2520, 2542, 2595, 2611, 2614, 2656, 2749, 2763, 2825, 2850, 2869, 2907, 2976, 3006, 3031, 3041, 3069, 3112, 3128, 3134, 3141, 3178, 3187, 3316, 3317, 3338, 3342, 3381, 3390, 3439, 3442, 3478, 3556]. network [10, 338, 381, 385, 390, 418, 421, 436, 468, 480, 544, 692, 730, 735, 792, 920, 923, 951, 958, 963, 968, 1038, 1039, 1069, 1078, 1087, 1091, 1097, 1109, 1137, 1201, 1260, 1272, 1291, 1302, 1304, 1330, 1340, 1363, 1420, 1423, 1489, 1582, 1596, 1645, 1649, 1671, 1761, 1771, 1844, 1939, 1945, 2039, 2040, 2045, 2062, 2083, 2108, 2191, 2193, 2269, 2307, 2336, 2348, 2361, 2392, 2433, 2458, 2486, 2498, 2511, 2521, 2557, 2568, 2598, 2620, 2666, 2761, 2815, 2865, 2876, 2893, 2950, 3005, 3054, 3066, 3105, 3129, 3137, 3155, 3193, 3210, 3224, 3248, 3267, 3351, 3353, 3403, 3421, 3457, 3523, 3534, 3559, 3563]. network [7, 39, 50, 60, 100, 102, 156, 259, 290, 311, 357, 364, 450, 487, 530, 571, 579, 596, 613, 620, 640, 712, 714, 863, 896, 972, 982, 984, 1002, 1010, 1027, 1028, 1059, 1070, 1085, 1115, 1152, 1168, 1198, 1208, 1263, 1289, 1296, 1313, 1318, 1344, 1366, 1382, 1390, 1432, 1455, 1480, 1511, 1517, 1520, 1614, 1636, 1640, 1717, 1726, 1753, 1770, 1882, 1941, 1968, 2036, 2053, 2066, 2170, 2216, 2285, 2287, 2302, 2309, 2536, 2543, 2556, 2565, 2587, 2625, 2725, 2883, 2915, 2981, 2990, 3011, 3096, 3116, 3132, 3139, 3150, 3304, 3348, 3362, 3373, 3382, 3409, 3428, 3434, 3464, 3470, 3518, 3550]. network [218, 219, 408, 442, 460, 467, 491, 501, 549, 641–643, 662, 704, 741, 793, 826, 834, 922, 929, 942, 957, 1009, 1023, 1026, 1086, 1166, 1188, 1211, 1229, 1233, 1414, 1439, 1469, 1481, 1561, 1585, 1593, 1601, 1619, 1705, 1735, 1914, 1940, 1957, 1960, 1977, 2002, 2010, 2018, 2023, 2219, 2270, 2305, 2408, 2414, 2435, 2459, 2468, 2490, 2501, 2589, 2592, 2682, 2687, 2729, 2744, 2755, 2794, 2810, 2812, 2824, 2827, 2844, 2860, 2897, 2905, 2920, 3016, 3021, 3055, 3092, 3181, 3186, 3254, 3259, 3269, 3275, 3286, 3389, 3401, 3403, 3412, 3450, 3477, 3509, 3522, 3529]. Network-aware [1333, 1382, 2458, 2907]. Network-based [107, 2825, 2860, 3389]. network-Cloud [1302]. Network-coded [1407]. network-density-based [1939]. network-on-chip [3116]. network-side [2865]. network-wide [848, 1201]. network/interface [2408]. Networked [42, 101, 164, 618, 677, 1534, 2436, 2933, 3071, 3088]. Networking [144, 630, 664, 741, 1016, 1043, 1116, 1127, 1335, 1358, 1395, 1397, 1398, 1469, 1473, 1488, 1522, 1531, 1578, 1591, 1610, 1625, 1639, 1725, 1837, 1874, 1876, 1879, 1929, 1948, 1951, 1955, 1961, 2000, 2004, 2045, 2131, 2156, 2224, 2235, 2329, 2333, 2383, 2445, 2466, 2494, 2504, 2532, 2567, 2599, 2607, 2622, 2650, 2660, 2672, 2691, 2802, 2905, 2989, 2995, 3022, 3045, 3102, 3161, 3269, 3289, 3291, 3319, 3396, 3419, 3555]. Networks [9, 233, 721, 732, 769, 778, 845, 866, 875, 887, 915, 948, 964, 1036, 1183, 1190, 1199, 1222, 1243, 1246, 1258, 1273, 1328, 1332, 1333, 1339, 1346, 1399, 1445, 1464, 1479, 1521, 1587, 1592, 1618, 1681, 1689, 1710, 1720, 1737, 1759, 1788, 1842, 1843, 1854, 1867, 1868, 1870, 1890, 1909, 1979, 1985, 2032, 2086, 2152, 2157, 2164, 2203, 2220, 2244, 2253, 2301, 2319, 2323, 2336, 2356, 2395, 2491, 2531, 2541, 2569, 2572, 2614, 2633, 2649, 2696, 2874, 2892, 2919, 3000, 3030, 3041, 3103, 3118, 3157, 3180, 3187, 3195, 3209, 3220, 3283, 3381, 3467, 3474, 3485, 3487, 3543, 3546, 3553, 3560]. networks [111, 155, 216, 277, 328, 343, 415, 521, 535, 542, 545, 572, 588, 601, 616, 617, 669, 681, 691, 694, 700, 706, 718, 724, 740, 764, 776, 781, 804, 813, 835, 836, 850, 854, 858, 870, 892, 928, 977, 991, 1019, 1095, 1110, 1112, 1157, 1200, 1228, 1299, 1300, 1311, 1345, 1361, 1404, 1443, 1493, 1504, 1518, 1519, 1524, 1545, 1557, 1611, 1633, 1690, 1693, 1700, 1703, 1751, 1823, 1878, 1900, 1912, 1944, 2001, 2045, 2060, 2064, 2120, 2132, 2142, 2231, 2245, 2271, 2377, 2382, 2405, 2411, 2424, 2495, 2598, 2732, 2748, 2770, 2851, 2869, 2914, 3058, 3112, 3278, 3417]. networks [256, 264, 268, 275, 322, 388, 437, 438, 456, 541, 565, 570, 575, 581, 594, 595, 598, 606, 679, 731, 752, 803, 860, 916, 932, 944, 969, 985, 1034, 1040, 1078, 1099, 1104, 1107, 1141–1143, 1147, 1165, 1181, 1218, 1232, 1247, 1282, 1326, 1330, 1363, 1422, 1424, 1429, 1503, 1528, 1540, 1576, 1595, 1612, 1613, 1617, 1631, 1711, 1738, 1754, 1756, 1762, 1766, 1820, 1884, 1910, 1939, 2022, 2039, 2063, 2114, 2184, 2209, 2243, 2276, 2321, 2342, 2401, 2452, 2547, 2550, 2551, 2561, 2577, 2579, 2608, 2666, 2675, 2690, 2710, 2778, 3050, 3060, 3137, 3142, 3175, 3179, 3208, 3270, 3280, 3475]. networks [11, 162, 278, 320, 344, 350, 371, 374, 403, 452, 497, 519, 528, 556, 585, 592, 597, 611, 682, 708, 720, 723, 728, 735, 748, 791, 816, 818, 825, 882, 917, 930, 939, 992, 997, 1052, 1062, 1094, 1120, 1127, 1132, 1171, 1173, 1221, 1264, 1337, 1368, 1408, 1427, 1458, 1495, 1499, 1525, 1532, 1575, 1588, 1590, 1642, 1657, 1694, 1773, 1779, 1810, 1825, 1896, 1903, 1918, 2020, 2035, 2040, 2139, 2169, 2194, 2232, 2258, 2272, 2274, 2292, 2338, 2345, 2357, 2369, 2370, 2392, 2434, 2444, 2454, 2559, 2600, 2615, 2637, 2680, 2795, 2961, 3027, 3086, 3105, 3201, 3258, 3354, 3357, 3369, 3402, 3436, 3453, 3465, 3501]. networks [279, 283, 321, 461, 477, 583, 590, 600, 673, 696, 697, 703, 715, 722, 767, 774, 828, 855, 864, 871, 880, 936, 938, 1007, 1031, 1057, 1096, 1150, 1255, 1318, 1336, 1351, 1375, 1387, 1391, 1396, 1453, 1454, 1470, 1497, 1509, 1553, 1580, 1606, 1675, 1678, 1680, 1707, 1717, 1739, 1750, 1760, 1805, 1808, 1840, 1845, 1848, 1858, 1860, 1882, 1906, 1967, 2019, 2031, 2033, 2047, 2075, 2098, 2124, 2167, 2171, 2300, 2353, 2417, 2439, 2460, 2513, 2535, 2581, 2596, 2769, 2800, 2830, 2884, 2917, 2949, 3039, 3119, 3131, 3150, 3160, 3221, 3304, 3333, 3350, 3359, 3387, 3411, 3422, 3497, 3502, 3504, 3548]. networks [147, 258, 273, 337, 356, 377, 428, 449, 459, 499, 547, 559, 567, 619, 626, 707, 726, 773, 797, 798, 801, 851, 852, 895, 902, 905, 971, 978, 983, 1005, 1042, 1079, 1090, 1111, 1115, 1215, 1226, 1244, 1254, 1259, 1268, 1303, 1338, 1348, 1451, 1472, 1501, 1523, 1539, 1597, 1598, 1643, 1651, 1669, 1698, 1718, 1732, 1777, 1803, 1804, 1855, 1859, 1908, 1970, 1981, 2008, 2014, 2021, 2051, 2116, 2122, 2168, 2176, 2179, 2201, 2240, 2246, 2250, 2343, 2358, 2372, 2445, 2683, 2741, 2846, 2887, 2927, 2965, 2988, 3008, 3025, 3040, 3183, 3194, 3324, 3331, 3343, 3423, 3434, 3530]. networks [64, 292, 430, 464, 472, 474, 476, 520, 522, 529, 546, 561, 676, 685, 702, 719, 749, 768, 793, 819, 843, 891, 901, 926, 945, 966, 986, 994, 1014, 1022, 1035, 1148, 1153, 1154, 1162, 1217, 1237, 1238, 1252, 1266, 1274, 1281, 1365, 1379, 1394, 1419, 1433, 1456, 1481, 1498, 1514, 1537, 1603, 1620, 1652, 1741, 1747, 1793, 1819, 1828, 1849, 1866, 1933, 1935, 1943, 1952, 1962, 1994, 2076, 2151, 2223, 2230, 2238, 2251, 2299, 2312, 2351, 2362, 2368, 2449, 2461, 2493, 2505, 2619, 2623, 2631, 2658, 2669, 2718, 2729, 2798, 2807, 2911, 2921, 2958, 2996, 3077, 3133, 3188, 3189, 3204, 3344, 3358, 3399, 3400]. networks [181, 221, 243, 369, 372, 414, 435, 495, 525, 591, 684, 689, 710, 736, 766, 790, 795, 799, 802, 822, 837, 881, 888, 914, 959, 960, 1048, 1098, 1108, 1119, 1187, 1191, 1240, 1256, 1276, 1287, 1305, 1342, 1353, 1377, 1407, 1447, 1448, 1476, 1486, 1507, 1526, 1544, 1602, 1607, 1619, 1647, 1697, 1709, 1728, 1764, 1769, 1781, 1794, 1850, 1898, 1954, 2013, 2026, 2057, 2090, 2095, 2100, 2135, 2149, 2150, 2210, 2256, 2331, 2334, 2360, 2374, 2413, 2443, 2450, 2464, 2478, 2673, 2674, 2719, 2731, 2740, 2812, 2839, 2858, 2864, 2940, 2991, 3007, 3014, 3067, 3117, 3141, 3249, 3309, 3412, 3427, 3469, 3539, 3551]. Networks [1020, 1044, 1155, 1329, 1446, 1457, 1586, 1600, 1686, 1696, 1893, 1928, 2017, 2073, 2183, 2317, 2976, 3502]. Neural [9–11, 64, 156, 243, 549, 730, 776, 1469, 1643, 1689, 2269, 2725, 2794, 2860, 2869, 2897, 2921, 3092, 3118, 3129, 3134, 3348, 3359, 3434, 3502, 3534]. NeuralSens [776]. neuro [1981]. neuro-fuzzy [1981]. neuron [1231]. neuron-based [1231]. news [1085, 3015]. Next [352, 456, 769, 1342, 1681, 1720, 1728, 1852, 1876, 2343, 2590, 2596, 2958, 3025, 3331, 3346]. Next-generation [1342, 1720, 1852, 1876, 2343, 2958, 3025, 3331]. NextG [3117]. nexus [3117]. NFC [1135, 2104]. NFRA [1378]. NFRA-C [1378]. NFV [1705, 2034, 2247, 2749, 2895, 3124, 3207, 3414]. NFV-based [3054, 3124]. NFV/SDN [2895]. Niched [1182]. NIDS [484, 998]. NLP [3404]. NN [409]. NN-queries [409]. No [3146]. nodal [3423]. Node [320, 381, 403, 430, 522, 588, 837, 948, 995, 1002, 1029, 1171, 1228, 1363, 1455, 1493, 1580, 1632, 1633, 1706, 1723, 1758, 1779, 1799, 1804, 1846, 1937, 2026, 2029, 2031, 2288, 2401, 2526, 2782, 3087, 3181, 3186, 3280, 3284, 3400, 3421, 3428]. node-based [1804]. node-disjoint [1493, 1937]. nodes [335, 489, 734, 770, 917, 956, 968, 1001, 1137, 1348, 1362, 1365, 1368, 1589, 1620, 1645, 1939, 2056, 2378, 2449, 2455, 2461, 2612, 2807, 3387]. noise [1247]. NOMA [2235, 3175, 3422]. non [48, 321, 382, 801, 958, 1167, 1224, 1348, 1745, 1749, 1890, 1998, 2009, 2082, 2213, 2235, 2243, 2364, 3075]. non-adjacency [2364]. non-cooperative [1167, 1348, 2009, 3075]. non-cryptographic [1224]. non-functional [1749, 1998]. non-green [1745]. non-lexical [2082]. non-linear [1890]. non-negative [382]. non-orthogonal [2213, 2235, 2243]. non-progressive [958]. non-repudiation [48]. non-uniform [801]. non-uniformly [321]. non-zero-sum [2009]. noncoherent [730]. nonlinear [2611]. nonparametric [2089]. norm [1064]. normalization [92, 3007]. North [37]. note [101, 166, 2575]. notice [2339, 2516, 2591, 2721]. noticeable [693]. notification [1793, 1907]. noughts [272]. Novel [11, 277, 357, 378, 407, 422, 468, 488, 507, 588, 594, 695, 730, 771, 814, 854, 907, 976, 1000, 1014, 1086, 1092, 1100, 1116, 1215, 1281, 1344, 1367, 1470, 1477, 1537, 1581, 1680, 1723, 1786, 1831, 1898, 1977, 2008, 2103, 2132, 2183, 2282, 2323, 2330, 2550, 2668, 2674, 2725, 2851, 3033, 3070, 3128, 3286, 3305, 3308, 3342, 3356, 3369, 3382, 3475]. November [2804, 2816, 3046, 3056, 3211, 3392]. NR [2709, 3322]. NR-U [3030]. ns [3322]. ns-3 [3322]. NT [84]. Null [1588]. Null-frequency [1588]. Number [1080, 2060]. numeric [3236].

O [1154, 3193, 3207]. O-intensive [327]. OASys [1334]. obfuscation [1550, 2449, 3468]. Object [7, 24, 54, 124, 176, 815, 900, 2316, 2717, 3165, 3237]. object-level [176]. object-oriented [54, 124]. object-removal [3237]. objective [603, 1263, 1439, 1831, 2129, 2230, 2257, 2282, 2463, 2503, 2621, 2679, 2815, 2838, 2931, 2965, 3071, 3159, 3181, 3372, 3438]. Objectives [439, 2856, 3161]. Objects [85, 705, 752, 1964, 2487, 2699]. oblivious [392, 1820, 3221]. observation [73]. observations [22]. obstacle [3428]. obstacle-aware [3428]. occupancy [3550]. Occurrence [2827]. occurring [844, 868]. ocean [3178]. OCR [19]. October [2779, 2792, 3023, 3035, 3199, 3379, 3541]. OEM [95]. OFDM [564, 1417]. OFDMA [1149, 1762]. off [1038, 1523, 2375]. offer [1024]. offline [772, 927, 1253, 2584]. offload [3433]. Offloading [1186, 1355, 1382, 1471, 1802, 1835, 1875, 2049, 2229, 2303, 2411, 2438, 2464, 2471, 2811, 2847, 2929, 3127, 3142, 3156, 3165, 3181, 3272, 3305, 3313, 3334, 3336, 3348, 3372, 3410, 3456, 3489, 3502, 3526, 3539]. offs [517, 2578]. oil [2244]. OLSR [1435]. omnidirectional [3313]. on-board [1215]. On-demand [58, 279, 314, 829, 864, 888, 1098, 1352, 1603, 1699, 2258, 2644, 3345, 3546]. On-device [3185]. On-line [98, 557, 1476, 1794]. on-street [1334]. One [402, 675, 860, 976, 1380, 1833, 2094, 2831, 3381]. one-dimensional [1380]. one-hop [1833]. One-shot [3381]. one-stop [2831]. one-time [675, 976]. One-way [402]. Online [481, 494, 516, 620, 690, 765, 772, 823, 1218, 1459, 1952, 2078, 2232, 2361, 2453, 2469, 2545, 2621, 2721, 2796, 2991, 3048, 3132, 3259, 3306, 3438, 3467, 3531]. online/offline [772]. only [1262, 2116]. ontological [755]. ontologically [335]. Ontology [427, 756, 913, 925, 1169]. Ontology-based [427, 756, 925]. Open [81, 112, 128, 131, 174, 766, 823, 1204, 1248, 1277, 1418, 1426, 1506, 1508, 1597, 1674, 1684, 1788, 1883, 2079, 2080, 2244, 2376, 2507, 2554, 2571, 2642, 2798, 2845, 2930, 3034, 3242, 3310, 3322, 3398, 3400, 3408, 3455, 3458, 3502]. Open-Source [2642, 3322, 3398]. OpenFlow [1697, 2062, 2741, 2758, 3493]. OpenFlow-based [1697]. OpenNCP [2275]. OpenNCP-based [2275]. OpenPATH [3049]. OpenStack [2283]. operating [465, 678]. operation [1489, 1936]. operational [2656]. operations [541]. Operator [579, 1788, 2066]. operator-independent [579]. Operators [165]. OPNET [377]. Opportunistic [535, 624, 679, 707, 978, 1007, 1073, 1098, 1127, 1150, 1228, 1254, 1334, 1424, 1455, 1509, 1607, 1613, 1669, 1766, 1882, 1885, 2087, 2135, 2468, 2501, 2831, 3320, 3427, 3551]. Opportunities [932, 1386, 1414, 1701, 1852, 1982, 2481, 2483, 2560, 2893, 2916, 2923, 3029, 3079, 3310, 3346, 3347, 3484, 3567]. Opportunity [1565, 1900, 2559]. Optical [256, 275, 278, 344, 532, 570, 606, 798, 1233, 1234, 1244, 1681, 1700, 2023, 2285, 2417, 2690, 2874, 3113]. optical-acoustic [2023]. OPTICS [3263]. Optimal [68, 193, 321, 521, 634, 635, 639, 723, 1197, 1291, 1301, 1331, 1412, 1446, 1636, 1645, 1708, 1782, 1845, 1878, 2210, 2276, 2293, 2513, 2526, 2598, 2633, 2975, 3243, 3305, 3513]. optimally [1745, 2847]. optimisation [977]. optimistic [3415]. Optimization [11, 279, 342, 608, 622, 651, 679, 718, 719, 738, 759, 777, 835, 866, 878, 957, 992, 1019, 1025, 1040, 1068, 1116, 1119, 1121, 1171, 1182, 1194, 1236, 1257, 1264, 1302, 1323, 1387, 1418, 1423, 1487, 1538, 1545, 1569, 1627, 1634, 1668, 1796, 1799, 1890, 1895, 1918, 1931, 1998, 2030, 2053, 2086, 2129, 2142, 2201, 2207, 2230, 2238, 2246, 2250, 2253, 2298, 2304, 2320, 2463, 2482, 2503, 2509, 2601, 2649, 2679, 2691, 2883, 2903, 2939, 3031, 3041, 3051, 3071, 3159, 3165, 3181, 3189, 3222, 3280, 3313, 3324, 3381, 3385, 3390, 3425, 3467, 3504, 3526, 3531, 3534, 3550]. optimization-based [1387, 2053, 2253, 3550]. Optimization/simulation [1545]. Optimization/simulation-based [1545]. optimize [721, 1529]. Optimized [255, 615, 901, 910, 914, 1111, 1469, 1833, 1909, 1966, 2243, 2330, 2378, 2561, 2612, 2838, 3032, 3267, 3429, 3462, 3470]. optimizer [637, 3372]. optimizers [2142]. Optimizing [656, 1404, 1421, 2066, 2388, 2750, 3017, 3088, 3160, 3457, 3493]. options [846]. Oracle [3355]. ORCEF [1459]. orchestrated [2172]. Orchestration [542, 1539, 1834, 2458, 2806, 3269]. orchestrator [2868]. order [422, 573, 800, 1549, 2002]. organisms [2463]. organization [27, 500, 922]. organizational [6, 129]. organizations [2196]. organized [826, 1338]. organizing [7, 593, 726, 855, 936, 1163, 1375, 1651, 2040, 2061]. orientation [1312, 1499]. Oriented [54, 124, 129, 199, 331, 335, 388, 408, 409, 433, 540, 763, 794, 831, 832, 969, 1075, 1099, 1393, 1437, 1505, 1676, 1781, 1881, 2161, 2240, 2255, 2322, 2363, 2505, 2698, 2860, 2883, 2968, 3077, 3137]. Orthogonal [704, 973, 2213, 2235, 2243]. OSIA [800]. OSPF [1789, 2301]. other [122, 168]. Out-of-order [800]. outage [1758, 2086]. outages [2241]. outbreak [3015]. outcast [1698, 2444]. outdoor [884, 2085]. outlier [3263]. output [1308]. Outsourced [1482, 1549, 1628, 2174, 2327, 2442, 3043]. Outsourcing [867, 2195, 2260, 2293]. over-the-top [1229]. Overcoming [134, 3440]. overflow [3293]. overhead [421, 1546, 3339]. overlaid [802, 1758]. overlapping [395, 1441, 3401]. overlay [372, 412, 452, 617, 747, 921, 966, 1057, 1242, 1281, 1327, 1445, 1453, 1804]. overlay-aware [452]. overload [376, 400, 1217]. OverShard [3401]. overview [796, 912, 1667, 1888, 2394, 2763, 3016, 3018]. ownership [287, 554, 1389, 3154].

P [2377]. P-AUV [2377]. P2P [144, 220, 281, 338, 423, 658, 706, 751, 885, 924, 1058, 1081, 1163, 1211, 1327, 1453, 1510, 1742, 2720, 3457]. P4 [2546, 2682, 2760, 2809, 3238, 3273]. P4-based [2809]. pacemaker [29]. package [236, 640]. Packet [288, 434, 463, 465, 559, 765, 782, 845, 911, 978, 1149, 1152, 1221, 1226, 1251, 1354, 1450, 1524, 2015, 2067, 2295, 2389, 2409, 2481, 2546, 2599, 2661, 2682, 2835, 2837, 3207, 3238, 3248]. packet-forwarding [465]. Packet-level [765]. packet-reordering [2389]. packets [2760, 2981]. page [146, 271, 1620]. page-granularity [1620]. pages [64]. paging [3180]. PAgIoT [1734]. pain [2689]. pairing [2803]. pairs [2318]. palliative [2309]. palmprint [568, 737]. pandemic [2886, 2941]. Pandora [433]. Panoramic [208]. PANs [1787]. Paper [20, 1434, 1659, 2058, 2538, 3364]. Papers [86]. paradigm [152, 227, 647, 2229, 2441, 2554, 2771, 3109, 3175, 3272]. paradigms [4, 54, 211, 2419, 3182, 3269]. paradox [2833]. ParaFlow [1962]. Parallel [30, 396, 447, 1216, 1450, 1721, 1881, 1962, 2238, 2279, 2939, 3378, 3399, 3554]. parallelism [2142]. Parallelization [2337]. parallelized [958, 2743]. parameter [352, 623, 1497, 1918, 2493]. parameters [7, 1288]. parametric [1817]. Pareto [1040, 2253]. Pareto-based [1040]. PARIOT [3363]. parking [1334, 1973, 2891, 3385, 3550]. PARS [1529]. Part [8, 299]. Parti [2848]. Parti-colored [2848]. partial [281, 1382, 1504, 1524, 1864, 1896]. particle [679, 906, 1257, 1423, 2142, 2230, 2503, 2904, 3071]. parties [3026]. partite [3549]. partition [242, 521, 1108, 1462, 1977, 2239, 3067, 3378]. partition-match [3378]. Partitioning [1121, 1373, 1799, 1954, 1960, 2303, 3362]. party [976, 1012, 1088, 1894, 2084, 2724, 2730, 3536]. Parzen [730]. passive [570, 606, 635, 645, 699, 1133, 1481, 1681, 2153, 2200, 2285, 2449, 3018, 3391]. password [675, 856, 892, 1056, 2157, 3286]. password-based [892]. password-guessing [3286]. passwords [675]. past [1797]. patching [431, 887, 2602]. patent [1065]. Path [12, 155, 258, 264, 275, 403, 495, 546, 662, 800, 959, 1087, 1140, 1252, 1493, 1603, 1724, 1805, 1919, 1939, 2141, 2358, 2480, 2535, 2679, 2741, 3526, 3546, 3554]. path-based [403]. path-traceable [495]. Paths [963, 1092, 1270, 1706, 1937, 2318]. pathways [26]. Patient [76, 587, 2941]. Patient-centered [76, 2941]. Patients [2264, 2281, 2291]. pattern [115, 515, 1450, 1462, 1861, 1892, 2160, 2192, 2200, 2473, 3232]. Patterns [332, 530, 604, 844, 868, 1004, 1478, 2262, 2359, 2634, 3514]. PAVE [2150]. pay [42, 661, 2091]. pay-as-you-use [42]. pay-per-use [2091]. pay-TV [661]. payload [2067, 2175, 2760, 3045, 3232]. payload-based [2175]. payment [424, 2950]. payments [1571]. PC [66, 386]. PCA [1487]. PDP [1862]. PEA [2279]. peacekeeping [2254]. peak [973]. peak-to-average [973]. pedagogical [230]. PEDCKS [2412]. pedestrians [1926]. Peer [276, 331, 337, 421, 432, 437, 452, 495, 500, 519, 537, 566, 575, 581, 668, 694, 752, 806, 839, 907, 921, 975, 1082, 1098, 1118, 1250, 1375, 1669, 1672, 1718, 1811, 1956, 2360]. Peer-assisted [1098]. Peer-to-Peer [276, 337, 421, 432, 437, 452, 495, 500, 519, 537, 566, 575, 694, 752, 806, 839, 907, 921, 975, 1082, 1118, 1250, 1375, 1669, 1672, 1718, 2360]. PeerCast [372]. peering [459]. penalty [2668]. Pending [2225]. penetration [3478]. people [59, 571, 796, 1219, 1602]. people-centric [1602]. per-port [2649]. perceived [2762, 3424]. perception [3327]. perceptual [693]. percolation [1499]. Performability [3102]. Performance [26, 80, 111, 191, 192, 233, 250, 327, 351, 368, 391, 411, 418, 432, 434, 445, 450, 454, 460, 462, 484, 494, 570, 583, 621, 738, 768, 775, 833, 848, 860, 863, 875, 894, 897, 914, 928, 972, 998, 1023, 1036, 1053, 1069, 1071, 1164, 1214, 1234, 1285, 1323, 1340, 1345, 1455, 1471, 1483, 1507, 1582, 1618, 1679, 1724, 1758, 1768, 1784, 1812, 1834, 1841, 1862, 1910, 1964, 1968, 1974, 2019, 2179, 2202, 2219, 2247, 2271, 2337, 2371, 2431, 2434, 2452, 2502, 2544, 2590, 2625, 2635, 2660, 2728, 2744, 2787, 2806, 2868, 3030, 3049, 3124, 3137, 3216, 3235, 3493]. performance [70, 638, 954, 994, 1237, 1491, 2008, 2180, 2181, 2299, 2334, 2682, 2903, 3061, 3238, 3482]. Performance-aware [1834, 2903]. Performance-to-Power [3216]. performances [967, 1320]. perimeter [1496, 1671]. period [1300]. periodic [2262, 2388]. periodically [1529, 3426]. permission [2652]. permissions [84, 3559]. permutation [1210]. persistence [2383]. person [578, 1571]. person-to-person [1571]. personal [497, 503, 852, 1650, 1743]. personality [512]. personalization [877, 2833]. personalized [138, 480, 979, 1138, 1212, 1565, 1892, 2150]. Perspective [274, 765, 769, 783, 953, 1185, 1375, 1569, 1888, 2030, 2534, 2702, 2709, 2797, 2846, 2914, 3016, 3189, 3527]. Perspectives [1611, 2017, 2041, 2186, 2343, 2893, 3337]. persuasion [3562]. Pervasive [236, 325, 362, 571, 607, 644, 794, 909, 1134, 1396, 1403, 1453, 1467, 1948, 1951, 1955, 2378, 2859, 2906]. PETAL [1903]. Petri [1494, 2430, 2668]. PGFit [2652]. PGTCN [3286]. PGWL [1620]. PHAODV [1321]. phase [1705, 1930, 2951]. phenomena [3178]. phenotype [453]. phishing [675, 961, 3255, 3562]. phone [1438, 1608, 1712, 2218]. phones [1285]. Physical [683, 907, 1166, 1717, 3201, 3246]. Physics [1623]. physiological [480, 2626]. Piconet [340, 1787]. PID [3535]. PileCast [1159]. pillbox [3455]. PIMA [3234]. pipeline [2835]. pipelined [16]. pitch [55]. pitfalls [2402]. pixel [443]. pixel-based [443]. Placement [269, 1090, 1151, 1381, 1645, 1667, 1723, 1779, 1782, 1927, 1934, 1935, 2014, 2031, 2083, 2180, 2250, 2252, 2313, 2363, 2482, 2508, 2513, 2539, 2542, 2598, 2601, 2609, 2666, 2750, 2931, 2951, 2983, 3086, 3139, 3140, 3193, 3250, 3278, 3280, 3381, 3414, 3467, 3485, 3487, 3500, 3512, 3549]. places [161, 3185]. plan [1102]. Plane [2394, 2546, 2824, 3086, 3167, 3222, 3273, 3442, 3525]. planned [477, 2287]. planner [18]. Planning [12, 17, 44, 637, 716, 1168, 1263, 1437, 1770, 1939, 2092, 2281, 2480, 2536, 2679, 3546]. plant [1168]. plate [315]. plateau [3556]. Platform [190, 427, 865, 1619, 1887, 1966, 1997, 2178, 2267, 2324, 2896, 2910]. platforms [911, 2564, 2806, 2826, 3328]. play [391]. playbacks [388]. Pleasure [2689]. plexi [1884]. PMIPv6 [523, 1516]. PMP [3378]. PnP [1096]. pocket [386, 1720]. PODA [2559]. PODEM [2784]. Point [1435, 1892, 2264, 2334, 2728, 2975, 3071, 3339]. point-of-care [2264]. points [1400, 1486, 1495, 1965]. poisoning [2156]. Poissonity [3197]. polar [2029]. polar-angle [2029]. policies [660, 806, 1708, 2588, 3193]. policing [376]. Policy [139, 265, 342, 406, 419, 440, 588, 697, 779, 1045, 1059, 1559, 1700, 1862, 2195, 2276, 2333, 2523, 2555, 2633, 2951, 3106, 3168, 3413]. Policy-based [406, 2523]. policy-enforcement [3168]. polling [1992]. pollution [751, 3179]. polygon [2029]. polynomial [442]. PON [782, 1113]. pooling [2534]. pools [2943]. popular [465, 1817]. Popularity [1178, 1742, 1755, 2470, 2485]. population [1375]. port [463, 1811, 2071, 2494, 2649]. port-knocking [2071]. port-scanning [2494]. portability [537]. Portable [1043]. portals [877]. portraying [3565]. Position [991, 1017, 1028, 1274, 1354, 1758, 2141, 2377, 2862]. Position-based [991, 1017, 1028, 1354]. positional [3076]. Positioning [1219, 2033, 2093]. positions [1779]. positives [712]. possible [2054]. Post [617, 793, 2199, 2218, 2514, 2832]. post-composition [617]. post-disaster [2218, 2514, 2832]. Post-processing [793]. posture [1995]. Potential [1114, 1231, 1338, 2163, 2402, 3347]. Potentials [3113]. Power [105, 279, 473, 612, 679, 693, 772, 816, 899, 969, 973, 1003, 1010, 1022, 1038, 1119, 1238, 1311, 1321, 1338, 1401, 1405, 1409, 1410, 1458, 1520, 1523, 1529, 1555, 1557, 1558, 1650, 1657, 1685, 1710, 1713, 1762, 1800, 1820, 1836, 1871, 1884, 1891, 1992, 2143, 2243, 2331, 2361, 2408, 2541, 2620, 2748, 2883, 2892, 3169, 3216, 3511, 3527]. power-aware [1891]. power-efficiency [1523]. power-efficient [1401]. power-grid [1557]. Power-saving [473, 816, 899]. powered [859, 1758, 2001, 2228, 2343, 2347, 2438, 2893, 3399, 3494, 3528]. Powerless [2178]. PPA [1892, 2472]. PPR [3216]. PPR-RM [3216]. Practical [221, 749, 965, 1098, 2167, 2195, 2325]. practice [85, 1462]. practices [444, 2456, 2516]. PRADD [1805]. Pre [247, 442, 1649, 1967, 2773, 3523]. Pre-6G [3523]. Pre-caching [1967]. pre-coalescing [2773]. pre-screening [247]. pre-searching [1649]. pre-sharing [442]. preamble [3157]. Precise [1284]. precision [2165, 2613, 2782, 3525]. precoding [973]. predetermined [1617]. Predicting [436, 1074, 1179, 1478, 2448, 2476, 2591]. Prediction [325, 606, 615, 654, 833, 891, 1029, 1059, 1139, 1251, 1254, 1465, 1656, 1751, 1808, 1861, 1868, 1897, 1900, 2005, 2136, 2160, 2192, 2359, 2371, 2455, 2610, 2637, 2764, 2770, 2938, 3007, 3041, 3096, 3214, 3258, 3267, 3403, 3565]. prediction-based [2764]. predictive [709, 950, 2803]. predominant [872]. Preemptive [2797, 3108]. Preface [106–108, 118, 126, 159, 205]. preference [1097, 1892, 2867]. preference-aware [1892, 2867]. preferences [3028]. prefetching [1708, 1983, 2764]. prefetching/caching [1983]. prefix [1649, 1726]. Prepare [3009]. prescriptions [2280]. presence [98, 858, 1344]. present [1797]. presentation [225]. presentations [165]. Preservation [952, 2064, 2078, 2216, 2363, 2475, 2731]. Preserving [306, 515, 827, 909, 1055, 1324, 1389, 1433, 1434, 1549, 1562, 1595, 1628, 1734, 1772, 1863, 1894, 2114, 2123, 2174, 2196, 2293, 2299, 2331, 2350, 2373, 2390, 2404, 2426, 2472, 2528, 2840, 2880, 2882, 3067, 3114, 3144, 3156, 3158, 3202, 3215, 3217, 3234, 3237, 3261, 3264, 3308, 3356, 3360, 3418, 3449, 3464, 3468, 3494, 3529, 3568]. Press [3291]. Prevailing [2456, 2516]. prevent [675, 996, 2774]. preventing [2048]. Prevention [498, 725, 934, 1165, 1605, 1776, 2042, 2251, 2494, 2866, 3474]. Pricing [216, 1634, 1715, 2595, 2847, 3385]. PriFoB [3190]. Primary [1268, 2086]. primitive [276]. primitives [174, 2403]. Principal [1627]. principle [693]. principled [335]. Principles [432, 1467, 2957, 3457, 3498, 3562]. printers [42]. prioritization [1132, 3547]. prioritized [844, 868, 968]. Priority [691, 994, 1184, 1733, 2184, 2369, 2741, 2764, 3157]. priority-aware [1733]. Priority-based [2764, 3157]. PriSIEM [3158]. Privacy [306, 485, 634, 698, 734, 817, 827, 867, 879, 909, 927, 952, 1055, 1138, 1324, 1433, 1434, 1562, 1628, 1734, 1772, 1814, 1863, 1894, 1899, 1925, 1952, 1988, 2064, 2078, 2084, 2105, 2114, 2134, 2135, 2174, 2196, 2204, 2216, 2293, 2299, 2331, 2349, 2350, 2363, 2373, 2374, 2390, 2391, 2404, 2426, 2472, 2475, 2506, 2528, 2567, 2673, 2712, 2718, 2731, 2789, 2794, 2833, 2840, 2852, 2880, 2882, 2887, 2902, 2950, 3026, 3067, 3114, 3126, 3144, 3156, 3158, 3190, 3202, 3215, 3217, 3218, 3234, 3237, 3242, 3261, 3264, 3274, 3299, 3308, 3320, 3356, 3360, 3418, 3441, 3449, 3464, 3468, 3494, 3529, 3545, 3568]. Privacy-aware [2084, 2204, 2349, 3190, 3545]. privacy-enhancing [3218].

Privacy-friendly [817]. privacy-personalization [2833]. Privacy-Preserving [306, 909, 1324, 1433, 1628, 1734, 1772, 1863, 1894, 2114, 2174, 2196, 2293, 2299, 2331, 2350, 2390, 2404, 2426, 2472, 2528, 2880, 2882, 3067, 3114, 3144, 3156, 3158, 3202, 3215, 3217, 3234, 3237, 3261, 3356, 3418, 3464, 3468, 3494, 3529, 3568]. privacy-protection [2718]. Privacy-respecting [2105]. private [1036, 2056, 2950, 2982, 3505, 3513]. privilege [2182]. Proactive [103, 369, 604, 706, 752, 1588, 1967, 2359, 2610, 2906, 2988]. proactively [477]. proactively-planned [477]. Probabilistic [21, 631, 634, 730, 1059, 1222, 1456, 1963, 2401, 2681]. Probability [436, 1523, 2559]. Probability-based [2559]. Probe [1217]. Probe-based [1217]. probing [1357, 2093]. Problem [160, 193, 1011, 1014, 1151, 1324, 1356, 1652, 1781, 1850, 1914, 3086, 3153, 3355, 3487, 3549]. problems [614, 1092, 1890, 2463, 2931]. procedure [191]. Process [4, 6, 121, 129, 171, 304, 358, 406, 642, 646, 754, 929, 995, 1067, 1069, 1145, 1288, 1547, 1905, 2365, 2418, 2957, 3482]. process-centric [171]. Process-context [358]. process-level [1288]. process-oriented [129]. processes [57, 194, 1066, 1070, 2803]. processing [19, 43, 629, 651, 793, 882, 1120, 1184, 1213, 1293, 1370, 1790, 1855, 1895, 1931, 2126, 2379, 2835, 3179, 3471]. processor [16, 52, 2682]. Processors [858, 1771, 2835]. product [648, 756, 1064, 1069, 2834]. product-instance [648]. production [1867]. productive [2262]. productive-periodic [2262]. products [3132]. professional [85, 1036]. PROFIBUS [283]. Profile [361, 524]. profiles [872, 1727, 2094]. profiling [998, 1041, 1288, 1359, 1506, 1746, 2384, 3093]. Profit [400]. Profit-aware [400]. profitable [3557]. profits [1771]. Program [479, 780, 2407, 2664]. program-based [479]. programmable [438, 2358, 2682, 3133, 3525]. Programming [805, 1074, 1914, 2163, 2297, 2652, 2835, 3273]. Programming-based [2297]. programs [46]. progress [1607, 2697, 2798, 3441]. Progression [2551]. Progressive [844, 868, 958, 2797, 3085]. project [1416, 1565]. projection [515, 2778]. projects [1325]. prolonging [1617]. prolongs [1377]. prominent [3328]. promoting [2994]. promotion [2866]. Proof [642, 1389, 2154, 2212, 2821, 3075]. proof-of-concept [642]. proofs [632]. Propagation [653, 904, 915, 1058, 1476, 2370]. properties [379, 1749, 1998]. property [1180, 2784, 3105]. property-based [2784]. PRoPHET [2199]. proportional [974]. proposal [615, 828, 1247, 3481]. proposals [1276]. proposed [2042]. ProSe [3297]. prospect [2044, 2339]. prospects [2745]. prosumer [2952]. prosumer-side [2952]. protect [657]. Protected [2761, 2874]. Protecting [751, 847, 1899, 2576, 3359]. Protection [264, 275, 277, 344, 400, 443, 530, 669, 698, 737, 1027, 1577, 1814, 2086, 2134, 2374, 2391, 2417, 2581, 2712, 2718, 2734, 2794, 3026, 3335, 3523]. Protections [278]. Protocol [155, 250, 273, 307, 322, 371, 388, 395, 454, 458, 538, 584, 597, 631, 661, 664, 695, 715, 757, 774, 807, 811, 813, 829, 851, 860, 875, 886, 942, 976, 1005, 1024, 1043, 1103, 1110, 1111, 1155, 1210, 1232, 1321, 1354, 1372, 1378, 1423, 1435, 1469, 1521, 1526, 1528, 1588, 1606, 1615, 1734, 1752, 1777, 1783, 1789, 1805, 1891, 1906, 1922, 1969, 1974, 1981, 2018, 2021, 2022, 2084, 2157, 2159, 2165, 2209, 2220, 2292, 2316, 2333, 2338, 2450, 2451, 2460, 2472, 2480, 2600, 2602, 2614, 2636, 2683, 2748, 2835, 2837, 2861, 2880, 2973, 3050, 3131, 3183, 3219, 3361, 3369, 3475, 3552]. protocol [198, 234, 304, 474, 609, 685, 741, 747, 888, 986, 1012, 1055, 1158, 1248, 1252, 1258, 1352, 1356, 1394, 1481, 1514, 1544, 1786, 1977, 2057, 2089, 2122, 2151, 2176, 2183, 2330, 2351, 2372, 2443, 2719, 2724, 2864, 3085, 3227, 3493]. Protocol-independent [2835]. protocols [188, 272, 276, 357, 472, 539, 585, 632, 680, 735, 771, 798, 817, 863, 881, 894, 984, 996, 1016, 1028, 1032, 1044, 1059, 1142, 1143, 1174, 1190, 1205, 1240, 1246, 1253, 1303, 1345, 1351, 1390, 1432, 1461, 1517, 1575, 1607, 1662, 1720, 1735, 1747, 1800, 1801, 1823, 1846, 1979, 2082, 2095, 2108, 2120, 2149, 2299, 2322, 2547, 2589, 2624, 2680, 2742, 2797, 2814, 2858, 2961, 3018, 3076, 3105, 3246, 3247]. Prototype [20, 44, 1169, 1556, 1619, 1839, 1997]. prototyping [230]. provable [1012]. provably [2472, 3050, 3464]. provably-secure [2472]. Provenance [2046, 2208]. provide [1212]. provided [565]. Provider [288, 459, 1715, 2672]. Provider-based [288]. providers [1088, 1416, 1548, 2091, 2595, 2959, 3430]. Providing [475, 592, 731]. provision [619, 668, 907, 974, 1105, 1629, 2644]. provision-aware [619]. provisioned [2356]. provisioning [217, 275, 586, 625, 943, 1302, 1676, 1897, 1901, 2038, 2396, 2406, 2545, 2759, 2892, 2972, 3153, 3306, 3438]. Proximity [907, 1742, 2953]. Proxy [251, 255, 337, 486, 595, 937, 1118, 1908, 1972, 2372, 2910]. proxy-based [1118]. pruning [3118]. pseudonym [1080]. PSO [1421, 3428]. PSSB [691]. PSUM [1672]. psychosis [185]. Public [49, 349, 576, 825, 1598, 1945, 2347, 2404, 2412, 2443, 2719, 2778, 2914, 2966, 3070]. publication [2150, 3568]. Publicly [2138, 2390, 3043]. publish [653]. publish/subscribe [653]. Publisher [166, 2575]. publishing [37, 2204, 2982]. PUF [927, 3361]. PUF-based [3361]. PUF-enhanced [927]. pull [1197]. punish [956]. Puppet [1561]. purpose [435]. Push [1063, 1197, 3289].

Q [2687, 2810, 3041, 3280, 3453]. Q-Connectivity [3280]. Q-Coverage [3280]. Q-learning-based [3453]. Q-Network [2687, 2810, 3041]. QCM2R [2683]. QMIX [3563]. QO [1023]. QO-STBC [1023]. QoE [598, 1343, 1405, 1765, 1817, 1909, 2155, 2256, 2314, 2634, 2698, 2846, 3417]. QoE-aware [2314]. QoE-driven [1343, 1765]. QoE-oriented [2698]. QoE/QoS [1405]. QoE/QoS-aware [1405]. QoS [58, 83, 116, 149, 164, 218, 334, 342, 345, 363, 391, 397, 414, 473, 474, 480, 532, 598, 606, 624, 700, 748, 831, 854, 1005, 1148, 1155, 1182, 1220, 1348, 1429, 1454, 1777, 1779, 1874, 1918, 1985, 2148, 2201, 2211, 2317, 2322, 2375, 2562, 2683, 2759, 2846, 2883, 2967, 3060, 3108, 3157, 3362]. QoS-aware [391, 1405, 2211, 2562, 2683, 2759, 3362]. QoS-Based [164, 748, 854, 1985, 3060]. QoS-differentiated [3157]. QoS-driven [1777]. QoS-enabled [397]. QoS-oriented [2883]. QoS/QoE [2846]. QQIGSA [1845]. QR [2871, 3510]. Q’tron [243]. quadrivalent [1845]. qualitative [1009, 1971]. Quality [71, 109, 154, 234, 260, 598, 608, 612, 669, 692, 719, 783, 1149, 1170, 1335, 1445, 1474, 1505, 1522, 1525, 1541, 1717, 1722, 1756, 1763, 1824, 1858, 1874, 2146, 2151, 2179, 2240, 3000, 3088, 3196, 3215, 3320, 3466, 3547]. quality-driven [1541, 3215]. Quality-of-sensing [2146]. Quality-of-service [109, 234, 260, 608, 612]. quality-of-services [783]. Quantifying [964]. Quantitative [414, 999, 2264]. quantized [1907]. Quantum [838, 1845, 2136, 2137, 2330, 2823, 3097]. quantum-inspired [838, 1845, 2136]. quarantining [706]. queries [32, 409, 1177, 1327, 1433, 1869]. Query [40, 45, 421, 495, 651, 852, 882, 901, 915, 1120, 1122, 1293, 1357, 1370, 1931, 1969, 1981, 2496, 2600, 3264]. query-driven [901, 2600]. querying [855]. Queue [352, 623, 1059, 1251, 1668, 2159, 3419]. queue-based [1668]. queued [947]. queueing [368, 738]. queues [874, 1011]. queuing [402, 1022, 1728, 3222]. QUIC [3436, 3493]. Quick [3540]. quicSDN [3436]. QUIET [551]. Quokka [1847]. QWRP [2600].

R [192, 2730, 2917, 3076]. R&D [1565]. R-ALOHA [192]. R-Dedup [2730]. R-gram [3076]. RA [3168]. RAaaS [3430]. RaaS [2376]. RaceCC [3343]. radar [1496]. Radient [1610]. Radio [280, 340, 426, 592, 646, 664, 728, 896, 1005, 1035, 1141, 1166, 1181, 1190, 1221, 1238, 1268, 1303, 1305, 1330, 1361, 1386, 1427, 1470, 1590, 1618, 1635, 1654, 1657, 1747, 1753, 1754, 1759, 1828, 1836, 1866, 1900, 1928, 1989, 2047, 2060, 2075, 2086, 2098, 2108, 2124, 2151, 2171, 2209, 2220, 2223, 2240, 2245, 2486, 2511, 2521, 2551, 2624, 2731, 2751, 2798, 2846, 2851, 2858, 3018, 3031, 3063, 3194, 3409, 3522]. Radio-Mama [646]. radios [1027]. rail [2441]. rails [3339]. rain [2709]. RAN [2320, 2521, 2687, 2728, 2795, 3310, 3482]. RAN-aware [2795]. Random [973, 1218, 1308, 1340, 1721, 1918, 2130, 2761, 2897, 3157, 3172, 3210, 3223, 3486]. random-interleaved [973]. Randomized [575, 1223]. randomly [626, 1331, 1412]. Randomness [1224]. RANet [3092]. Range [422, 726, 1219, 1327, 1328, 1362, 1680, 1969, 2090, 2953]. Range-free [1362]. ranges [1318]. rank [1529]. ranked [2603]. ranking [1160]. RANs [2810, 3512]. Ransomware [2354, 2791, 3448, 3559]. Rapid [569, 3231]. rapidly [3343]. rare [752]. RAT [1591]. Rate [139, 352, 434, 573, 581, 679, 689, 710, 808, 923, 978, 1156, 1159, 1181, 1251, 1500, 1858, 2489, 2760, 3191, 3293, 3514]. rather [1262]. rating [823]. Ratio [973, 1287, 1671, 3104, 3216]. ratio- [1287]. ration [1200]. ratios [21]. Raze [3106]. RBA [2380]. RBAC [405]. RC [2911]. RC-chain [2911]. RCBAC [2786]. RCP [3485]. rDBF [2496]. RDF [338, 1327]. RDF-based [338, 1327]. RDMA [429]. Re [998, 1884, 1972, 2755, 2910, 3131, 3426]. re- [3131]. re-encoding [2755]. Re-Encryption [1972, 2910]. Re-examining [998]. re-scheduling [1884]. re-train [3426]. REACT [2582]. Reactive [133, 956, 1555, 2449, 2582]. read [2130]. read/write [2130]. Reader [370, 631, 1370, 1378, 1421]. reader-to-reader [631]. readiness [377]. Real [24, 55, 82, 122, 153, 156, 226, 237, 259, 527, 528, 635, 645, 689, 733, 804, 807, 808, 836, 840, 884, 915, 1105, 1161, 1218, 1224, 1425, 1555, 1556, 1641, 1699, 1800, 1830, 1860, 1891, 1930, 1966, 1995, 2096, 2107, 2368, 2408, 2441, 2637, 2870, 3020, 3038, 3225, 3232, 3293, 3372, 3426, 3525]. real-life [1105, 3426]. Real-time [24, 55, 82, 122, 153, 156, 226, 237, 259, 527, 528, 645, 689, 733, 804, 807, 808, 836, 840, 884, 1161, 1224, 1425, 1556, 1699, 1800, 1830, 1860, 1891, 1930, 1966, 1995, 2096, 2107, 2368, 2408, 2441, 2637, 2870, 3020, 3038, 3232, 3293, 3372, 3525]. real-world [3225]. realistic [898, 1029, 1971, 2299]. Reality [185, 190, 1225, 1997, 2656, 2932, 3547]. realms [433]. realtime [1139]. reasoning [716]. reasoning-based [716]. recall [1438]. recall-based [1438]. RECAR [3555]. Receiver [865, 2444]. Receiver-driven [2444]. Receivers [2887]. reception [1630, 1974]. Rechargeable [1858, 2258, 3546]. reciprocity [2127]. Recognition [7, 316, 362, 510, 511, 514, 515, 690, 1663, 2084, 2200, 2497, 2533, 2597, 2645, 2675, 2688, 2733, 2785, 2801, 3165, 3172, 3496, 3537]. Recognizing [2543, 2587]. recommendation [610, 1065, 1084, 1456, 1565, 1719, 1821, 1892, 2150, 2387, 2653, 2681, 2880]. recommendations [298, 2068, 2345, 2692, 2742, 2921, 3219, 3246]. recommended [3242]. recommender [1172]. Reconfigurable [832, 2112, 3160, 3232]. reconfiguration [428, 920, 1371, 3021, 3069, 3388]. Reconnection [3213]. Reconnection-Efficient [3213]. reconstruction [497, 952, 1010, 1995, 3179]. record [20]. recover [403]. recovery [389, 833, 843, 1095, 1209, 1407, 1524, 1606, 1664, 2251, 2514, 2832]. Recruitment [1748, 2918]. rectangular [2625]. rectify [959]. recurrent [2794, 2869]. recursive [277, 323, 1372, 1542]. recursively [447]. RED [352, 434, 742, 819, 1006, 3284]. Reddit [1476]. redeployment [720]. redirection [459]. Redo [1071]. RED’s [434]. reduce [782, 1435, 2169, 2907]. reduced [670, 3428]. Reducing [421, 434, 1179, 1630, 2189, 3039]. reduction [17, 280, 973, 1494, 2902, 3258]. Redundancy [775, 1664, 2791]. Redundancy-aware [775]. redundant [447, 1002, 1679, 2169]. reference [803, 2814]. referential [3478]. Refinement [2000, 2791]. refinement-based [2791]. ReFSM [2837]. Region [315, 699, 1177, 2020, 2380, 2625, 2646, 3421, 3425]. region-avoidance [2646]. Region-aware [3425]. Region-based [315]. region-duplication [699]. regional [3096]. registers [1284]. registration [429]. registries [617]. Registry [3349]. regression [1257, 2136, 2611, 2840, 3155]. Regressive [2803]. regular [834, 1462, 3105, 3232, 3314]. regularization [3033]. Regularized [3141, 3172]. Regulating [2879]. regulation [746, 2941]. rehabilitation [527, 932]. REHDFS [2130]. reinforce [830]. Reinforcement [297, 319, 766, 1955, 2515, 2632, 2917, 2965, 2974, 2987, 3006, 3062, 3145, 3149, 3191, 3192, 3194, 3226, 3233, 3295, 3319, 3336, 3344, 3357, 3382, 3422, 3424, 3426, 3462, 3526, 3533, 3534, 3539, 3545]. reinforcement-learning-based [3336]. rejection [393]. related [1004, 1128, 1315, 2005, 2117, 2393]. relation [1551]. relational [92]. relations [1001, 1964]. relationship [45, 374, 3104]. Relationships [168, 1106, 2448]. Relative [1284, 1602, 1922, 2762, 3076, 3463]. relative-error [3463]. Relative-importance [1602]. Relay [280, 430, 480, 671, 774, 787, 939, 984, 1079, 1114, 1349, 1435, 1645, 1723, 2031, 3331]. relay-aided [3331]. relay-based [1349]. relaying [602]. relays [1946, 2939]. relevance [2136, 2791]. Reliability [278, 326, 599, 662, 1366, 1664, 1735, 1774, 1805, 1843, 1984, 2069, 2148, 2541, 2615, 2666, 2939, 3216, 3217]. Reliability-aware [1805, 2666]. Reliability-enhanced [3217]. Reliable [437, 474, 476, 525, 586, 826, 833, 841, 882, 1167, 1292, 1339, 1394, 1606, 1658, 1783, 2072, 2116, 2199, 2266, 2292, 2307, 2469, 2527, 2565, 2940, 3105, 3194, 3418, 3485, 3518, 3561]. reliant [1101]. relying [1129]. remnants [1192]. remote [28, 62, 66, 191, 313, 314, 362, 483, 509, 560, 683, 856, 1056, 1277, 2443, 2474, 3095, 3268, 3556]. remotely [2173]. removable [2952]. removal [1247, 3237]. Rendering [257]. rendezvous [2245, 2975, 3071]. Renewable [1915, 2228, 2893]. Renewable-aware [1915]. renting [1424, 1766]. reordering [2389]. repackaging [3363]. repair [1102]. repairing [2033]. repeatability [1879]. replacement [815, 950, 2764]. Replica [861, 966, 1242, 1375, 1580, 1804, 2555, 2948]. Replica-aided [966]. replicas [1474, 2474]. replicating [111, 1401]. replication [269, 604, 605, 752, 837, 993, 1147, 1358, 1474, 1644, 2425, 2485, 2838, 3467]. report [1300]. reporter [1616]. reports [830]. representation [1122, 1495, 3130]. representations [3236]. Representing [270]. reprogrammable [66]. REPS [3131]. REPS-AKA3 [3131]. repudiation [48]. Reputation [453, 593, 601, 823, 826, 828, 830, 921, 1540, 1714, 1863, 2355, 2469, 2571, 2818, 2911, 3096, 3245]. Reputation-based [2911]. request [459, 604, 947, 1079, 2035, 2601, 3223]. request-redirection [459]. requests [1184, 2865]. requirement [2364]. Requirements [669, 1175, 2244, 2656, 2956]. rerouting [2754]. Research [5, 167, 577, 1001, 1052, 1063, 1122, 1248, 1309, 1317, 1325, 1386, 1414, 1427, 1506, 1725, 1759, 1821, 1839, 1852, 1961, 2041, 2068, 2079, 2249, 2265, 2343, 2394, 2419, 2483, 2521, 2551, 2571, 2592, 2657, 2692, 2808, 2898, 2964, 3068, 3113, 3151, 3273, 3287, 3323, 3347, 3458, 3503]. researches [2797]. reservation [798, 864, 1867, 3385]. reservations [2242]. reserved [2639]. residents’ [2473]. residual [1221]. resilience [453, 575, 1493, 1687, 1730, 1896, 2582, 2808, 3231]. Resilient [372, 383, 1372, 1545, 1624, 1711, 1935, 2210, 2809, 2821, 3187]. resistant [1672, 3361]. resolution [686, 768, 2778]. Resource [147, 359, 419, 423, 439, 459, 475, 530, 549, 550, 727, 924, 980, 1089, 1103, 1114, 1127, 1220, 1239, 1241, 1269, 1282, 1322, 1338, 1341, 1343, 1344, 1376, 1382, 1508, 1628, 1629, 1636, 1656, 1670, 1701, 1751, 1765, 1822, 1831, 1836, 1847, 1878, 1895, 1901, 1946, 2003, 2030, 2037, 2074, 2098, 2109, 2129, 2140, 2155, 2218, 2223, 2228, 2290, 2310, 2313, 2396, 2435, 2471, 2479, 2515, 2534, 2545, 2579, 2583, 2610, 2635, 2644, 2687, 2710, 2774, 2784, 2868, 2874, 2892, 2959, 2984, 3017, 3051, 3054, 3124, 3153, 3166, 3169, 3189, 3243, 3283, 3304, 3324, 3357, 3361, 3414, 3422, 3425, 3508, 3511, 3560]. resource [1278, 1359, 1543, 2126, 2674, 2706, 2894, 3140, 3233, 3430, 3539]. resource-aware [3054]. resource-awareness [550]. resource-constrained [1278, 2290]. resource-constraint [2784, 3361]. resource-efficient [1543]. Resource-utilization-aware [2140]. resources [25, 1167, 1897, 2523, 2698]. respecting [2105]. response [761, 1459, 1599, 1669, 1752, 2397, 3367]. RESTful [1573]. restorable [843]. Restoration [520, 818, 1842, 1912]. Restoring [1670]. RESTRAIN [3124]. Restricted [2484]. restructuring [340, 1787]. results [2100, 2325]. retention [956, 3283]. Retracted [2044, 2456, 2476]. Retraction [2339, 2516, 2591, 2721]. retransmission [1042, 1859]. Retrievability [2154, 2212]. Retrieval [89, 157, 164, 254, 2473, 3037]. retrieving [206, 2940]. reusable [2188]. reuse [124, 220, 611, 2402, 3205]. Revealing [3218]. Revenue [3132]. reverse [1028, 1841, 2022, 2837]. reversible [2540, 2823]. Review [764, 766, 769, 856, 918, 933, 934, 942, 986, 1045, 1123, 1142, 1173, 1248, 1275, 1277, 1304, 1317, 1373, 1432, 1434, 1454, 1506, 1578, 1644, 1674, 1681, 1687, 1736, 1746, 1747, 1753, 1822, 1880, 2068, 2070, 2076, 2079, 2113, 2121, 2170, 2194, 2211, 2224, 2232, 2311, 2397, 2414, 2446, 2553, 2557, 2592, 2623, 2650, 2659, 2692, 2770, 2775, 2785, 2818, 2823, 2845, 2856, 2997, 3000, 3010, 3014, 3019, 3044, 3068, 3126, 3166, 3170, 3178, 3218, 3242, 3246, 3275, 3287, 3291, 3318, 3330, 3395, 3402, 3404, 3411, 3427, 3452, 3481, 3484, 3519, 3567]. Reviewer [3241, 3407]. Reviewers [2111, 2367, 2618, 2855, 3083]. Reviewing [1821]. revised [2345]. revisited [1223]. Revocable [2182, 2332]. Revocation [1612, 2195, 3279, 3376]. revolution [169]. reward [2668, 3102, 3320]. rewarding [2105]. rewards [1863]. RF [1758, 2076, 2343]. RF-powered [2343]. RFID [628–639, 641–647, 927, 1158, 1168, 1169, 1210, 1219, 1262, 1263, 1370, 1420, 1421, 1770, 1911, 2153, 2181, 2451, 2926]. RFID-assisted [638]. RFID-based [639, 644, 1420]. RFID-enabled [643]. RFLS [2821]. rich [493, 1204, 3075]. ride [3494]. ride-sharing [3494]. Rider [1838]. Riders [2774]. ridgelet [317]. Rigorous [862]. RIM [1602]. ring [614, 2013]. rings [70, 1057, 1969]. Ripple [653]. Rise [1785, 2232, 2657, 2669, 2942, 3277]. Risk [275, 626, 1165, 1545, 1775, 2353, 2786, 2933, 3155, 3386]. risk-aware [2786]. Risk-based [626, 1775]. risks [1608]. RITS [1922]. RL [806, 3459]. RL-based [3459]. RLS [3535]. RMTS [2478]. RNN [2794]. RNN-DP [2794]. road [1433, 2758, 3550]. roadmap [3347]. roads [913]. Roadside [1048]. roaming [445, 2183, 3464, 3475]. robin [1146]. Robot [180, 219, 948, 1417, 1611, 2577, 2867, 2968]. robotic [229, 2493, 2949, 3423]. Robotics [3079]. robots [12, 682, 876, 3566]. Robust [392, 513, 590, 618, 623, 690, 810, 1057, 1151, 1560, 1960, 2053, 2057, 2395, 2478, 2585, 2634, 2740, 2748, 3095, 3278, 3382, 3535, 3555]. Robustness [1185, 1295, 2322, 2381]. role [90, 244, 291, 1064, 2186, 2531, 2994]. role-based [90, 244, 1064]. roll [2492]. rollup [3345, 3415]. root [2182, 2292]. rootkits [3307]. roots [276]. rotation [699]. round [1146, 2631, 3353]. round-trip [2631, 3353]. route [155, 357, 421, 662, 829, 878, 957, 1025, 2066, 2449, 2529]. route-maintenance [155]. routed [258]. router [1779]. routers [249, 323, 352, 1406, 1768]. routes [280]. ROuting [250, 279, 328, 351, 353, 363, 539, 681, 722, 728, 732, 735, 748, 764, 781, 783, 829, 841, 851, 855, 860, 938, 944, 991, 996, 1005, 1017–1019, 1087, 1115, 1140, 1143, 1155, 1157, 1181, 1190, 1199, 1248, 1295, 1351, 1354, 1423, 1429, 1457, 1500, 1503, 1527, 1528, 1532, 1567, 1588, 1615, 1673, 1710, 1720, 1732, 1747, 1789, 1823, 1846, 2020, 2022, 2026, 2034, 2082, 2086, 2095, 2108, 2141, 2148, 2151, 2209, 2253, 2342, 2376, 2377, 2417, 2446, 2482, 2547, 2550, 2565, 2600, 2614, 2624, 2831, 2858, 2878, 2917, 2973, 3041, 3060, 3062, 3174, 3187, 3219, 3221, 3534, 3555, 3561]. routing [189, 218, 273, 394, 472, 474, 495, 525, 546, 685, 702, 707, 777, 863, 881, 888, 901, 914, 922, 986, 994, 1035, 1059, 1153, 1182, 1205, 1252, 1270, 1294, 1310, 1321, 1432, 1461, 1520, 1603, 1607, 1619, 1643, 1654, 1686, 1696, 1733, 1735, 1752, 1777, 1819, 1866, 1873, 1928, 1940, 1977, 1981, 2021, 2051, 2087, 2094, 2238, 2317, 2330, 2589, 2601, 2683, 2812, 2920, 2967, 3018, 3285, 3344, 3549]. routing-scheduling [1619]. RPI [2975]. RPIFL [3418]. RPL [2282, 3219, 3244, 3524]. RPTD [3217]. RSCAT [3319]. RSS [370, 524, 954, 1160, 2032, 2493]. RSS PROYECT [954]. RSSI [2096, 2859]. RTT [2032]. rule [247, 306, 1004, 2239, 2799]. rule-based [2799]. ruled [2485]. RuleMiner [2799]. rules [1855]. rumor [2370, 3130]. Run [2552]. running [1471]. runoff [151]. Runtime [3335]. rural [314, 2308]. RVC [3245].

S [1516]. S-PMIPv6 [1516]. SA [1421, 3417]. SA-PSO [1421]. SaaS [1416]. safe [2290]. SafeCoder [3510]. safely [2134]. Safety [82, 913, 1598, 1997, 2347, 3510, 3537]. SAML [2099]. Sampling [530, 1031, 1218, 1537, 1858, 3195]. SAODV [857]. SAR [2733]. SARAH [731]. Satellite [1503, 1793, 2362, 2589, 2940, 3422, 3560]. saturation [3389]. save [1553]. saving [473, 597, 816, 899, 1405, 1409, 1685, 1871, 1992, 2773, 3108]. savings [1941]. SBAC [2622]. Scalability [323, 1397, 1722, 2038, 2167, 3068, 3099]. Scalable [176, 252, 253, 345, 349, 376, 418, 655, 693, 855, 1160, 1293, 1358, 1406, 1462, 1479, 1488, 1610, 1792, 1811, 1815, 1837, 1950, 2335, 2458, 2519, 2535, 2539, 2564, 2568, 2609, 2809, 2826, 2950, 3254, 3309, 3327, 3373, 3435, 3454, 3470]. scale [79, 100, 385, 517, 603, 611, 929, 948, 1050, 1082, 1208, 1299, 1421, 1463, 1808, 1911, 1962, 2039, 2165, 2238, 2292, 2496, 2512, 2786, 2813, 3250, 3302, 3399, 3420]. scaleable [111]. Scaling [79, 802, 1658, 2052, 2359, 2711, 2948, 3401]. scanner [385]. scanning [2494]. scans [3066]. scattered [2813]. scatternet [280, 497, 671, 801]. scenario [397, 1053, 1758, 2400]. scenarios [312, 640, 846, 1007, 1050, 1876, 2215, 2501, 2514, 3161, 3213, 3322]. Scheduled [1615]. Scheduled-links [1615]. scheduler [1013, 1405, 1491, 3145]. schedules [1230]. Scheduling [342, 439, 440, 452, 473, 522, 556, 624, 626, 654, 656, 670, 758, 759, 773, 782, 800, 947, 968, 993, 1082, 1110, 1140, 1149, 1150, 1216, 1281, 1291, 1336, 1339, 1349, 1352, 1388, 1416, 1442, 1446, 1529, 1533, 1535, 1538, 1614, 1619, 1641, 1665, 1699, 1701, 1733, 1820, 1823, 1836, 1847, 1864, 1881, 1884, 1910, 1920, 1930, 1937, 1966, 2047, 2056, 2097, 2129, 2167, 2184, 2193, 2242, 2257, 2258, 2274, 2364, 2369, 2392, 2425, 2463, 2503, 2515, 2541, 2549, 2651, 2697, 2698, 2788, 2883, 2903, 2935, 2946, 2947, 3002, 3075, 3125, 3140, 3203, 3216, 3233, 3238, 3270, 3305, 3321, 3322, 3362, 3391, 3425, 3438, 3500, 3546]. schema [3561]. schematic [59]. Scheme [58, 151, 162, 250, 251, 259, 329, 344, 393, 411, 424, 425, 443, 466, 468, 479, 483, 504, 513, 554, 558, 560, 562, 594, 606, 611, 699, 713, 722, 724, 727, 740, 750, 816, 825, 864, 892, 938, 1021, 1025, 1056, 1080, 1150, 1156, 1167, 1200, 1299, 1300, 1344, 1363, 1470, 1482, 1522, 1595, 1651, 1664, 1700, 1772, 1810, 1826, 1894, 1901, 1908, 1920, 1967, 1983, 1992, 2001, 2024, 2139, 2154, 2218, 2252, 2258, 2263, 2271, 2285, 2317, 2373, 2422, 2522, 2527, 2550, 2596, 2603, 2612, 2635, 2783, 2794, 2859, 3038, 3087, 3095, 3109, 3119, 3124, 3223, 3295, 3308, 3332, 3428, 3464, 3468, 3470, 3530, 3546]. scheme [242, 359, 373, 486, 489, 529, 561, 704, 782, 909, 957, 960, 1020, 1037, 1116, 1292, 1543, 1793, 1994, 2138, 2173, 2182, 2251, 2426, 2478, 2679, 2718, 2835, 2882, 2908, 3021, 3043, 3055, 3067, 3238, 3279, 3469, 3509]. schemes [494, 509, 701, 843, 856, 878, 1117, 1231, 1235, 1266, 1383, 1580, 1582, 1616, 1667, 1959, 2038, 2046, 2075, 2114, 2124, 3482]. SCIDS [246]. science [163, 2302]. scientific [110, 2257, 2621, 3198, 3216]. SCoT [2161]. scrambling [443]. screening [247]. Scripting [86, 126]. SCTP [468, 1511]. SD [1464, 1693]. SD-Anti-DDoS [1693]. SDN [1535, 1874, 1962, 2034, 2057, 2062, 2106, 2229, 2238, 2247, 2255, 2302, 2396, 2440, 2452, 2498, 2509, 2514, 2534, 2536, 2576, 2634, 2674, 2705, 2728, 2824, 2895, 2922, 2928, 2957, 2967, 3010, 3054, 3106, 3168, 3191, 3222, 3443, 3465, 3486, 3525, 3561]. SDN-Based [2057, 2396, 2452, 2634, 2674, 3010, 3168, 3191, 3222, 3465, 3561]. SDN-enabled [2302, 2514]. SDN-SFC-based [2255]. SDN/NFV [3054]. SDN/NFV-based [3054]. SDNs [2656, 2813, 3033]. SDRS [2918]. SDS [2252]. SDWAN [3532]. sea [3324]. SEAL [2576]. Seamless [663, 813, 937, 1229, 1275, 1342, 1426, 1636, 2085, 3552]. Search [53, 87, 189, 208, 266, 495, 576, 795, 838, 1141, 1194, 1271, 1567, 2132, 2134, 2136, 2412, 2463, 2603, 2627, 2776, 2927, 3425, 3428, 3513]. searchable [1949]. Searching [169, 694, 827, 1649]. second [422, 432, 573]. second-order [422, 573]. secondary [2731]. Secrecy [2761]. secret [594, 1086]. sector [726]. Secure [244, 292, 320, 373, 425, 486, 496, 507, 595, 597, 609, 636, 659, 665, 681, 724, 734, 827, 946, 957, 1044, 1083, 1174, 1248, 1308, 1385, 1391, 1461, 1559, 1563, 1587, 1592, 1606, 1625, 1628, 1673, 1754, 1844, 1873, 1894, 1972, 1990, 1993, 2046, 2122, 2161, 2188, 2273, 2275, 2307, 2349, 2372, 2380, 2404, 2472, 2511, 2512, 2562, 2570, 2576, 2589, 2622, 2636, 2665, 2692, 2719, 2724, 2730, 2783, 2790, 2857, 2903, 2910, 2928, 2960, 2997, 3014, 3034, 3050, 3055, 3070, 3119, 3131, 3138, 3309, 3356, 3383, 3435, 3451, 3464, 3532, 3552]. SecureBallot [3034]. Secured [3037, 3244, 3416, 3505]. Securing [355, 983, 2090, 2413, 2445, 3448, 3486]. Security [50, 84, 103, 150, 158, 177, 186, 212, 241, 250, 287, 307, 385, 442, 538, 588, 590, 591, 593, 596, 608, 633, 634, 657, 663, 731, 734, 737, 784, 811, 832, 857, 896, 927, 1117, 1130, 1131, 1315, 1425, 1624, 1697, 1731, 1776, 1792, 1829, 1904, 1925, 1978, 2054, 2063, 2114, 2128, 2148, 2164, 2172, 2178, 2214, 2215, 2252, 2393, 2397, 2456, 2470, 2475, 2506, 2516, 2520, 2522, 2556, 2608, 2628, 2675, 2699, 2705, 2709, 2712, 2745, 2778, 2789, 2797, 2825, 2834, 2886, 2898, 3019, 3097, 3099, 3126, 3152, 3158, 3187, 3196, 3201, 3242, 3246, 3299, 3310, 3315, 3375, 3384, 3427, 3458, 3465, 3498]. security [103, 197, 399, 548, 553, 908, 1012, 1101, 1126, 1180, 1608, 1795, 1856, 2722, 2814, 2852, 3503]. security-related [1315, 2393]. sedentary [307]. SEEDGT [3138]. seeking [1262]. segment [277, 2754]. segmentation [3339, 3428, 3439, 3478, 3556]. segmentations [2357]. segments [375, 2191]. SEL [3553]. Selection [334, 378, 493, 538, 666, 722, 831, 883, 955, 995, 1089, 1269, 1309, 1319, 1356, 1435, 1486, 1613, 1635, 1715, 1938, 2002, 2211, 2240, 2381, 2406, 2408, 2432, 2486, 2807, 2907, 3071, 3121, 3134, 3137, 3172, 3496, 3554]. Selective [615, 928, 1023, 1042, 1337, 2093, 2550, 2857, 3052]. selectors [3233]. Self [7, 487, 500, 593, 697, 726, 826, 855, 872, 922, 936, 1035, 1101, 1163, 1338, 1481, 1484, 1651, 1696, 2033, 2040, 2061, 2201, 2313, 2443, 2610, 2719, 2838, 3063, 3197, 3228, 3274, 3381]. self-adapting [872]. Self-adaptive [1696, 2201, 2313, 2610]. self-certified [2443, 2719]. self-configuration [697, 1481]. Self-configured [487]. self-defense [2838]. self-determining [3228]. Self-Learning [1484]. self-organization [500, 922]. self-organized [826, 1338]. self-organizing [7, 593, 726, 855, 936, 1163, 1651, 2040, 2061]. Self-reliant [1101]. self-repairing [2033]. self-similarity [3197]. self-updating [3063]. selfish [1228]. selfishness [353]. seller [1801]. SEM [1892]. Semantic [40, 610, 755, 795, 925, 1062, 1064, 1134, 1271, 1290, 1380, 1547, 1551, 1589, 1855, 1869, 1888, 2096, 2379, 2603, 2755, 2827, 3298, 3494, 3556, 3566]. Semantic-aware [2603]. semantic-based [795]. semantical [1892]. Semi [768, 1000, 1503, 2235, 2415, 2723, 3125, 3354, 3434, 3556]. Semi-Distributed [1503, 3354]. Semi-granted [2235]. semi-greedy [3125]. semi-Markov [1000]. Semi-supervised [2415, 2723, 3434, 3556]. semipublic [243]. send [3424]. sender [437, 1505, 2658]. sender-initiated [2658]. sender-oriented [1505]. SENSE [1228]. Sensing [449, 792, 955, 1266, 1318, 1328, 1377, 1541, 1543, 1651, 1738, 1748, 1754, 1824, 1869, 1900, 2047, 2105, 2146, 2178, 2180, 2200, 2284, 2328, 2388, 2455, 2475, 2528, 2551, 2613, 2718, 2821, 2859, 3063, 3194, 3196, 3235, 3354, 3399, 3463, 3556]. sensing-based [2180]. sensitive [844, 868, 1142, 1236, 1593, 2231, 3109, 3512]. sensitivity [1004, 1825]. sensitivity-based [1825]. Sensor [308, 588, 669, 681, 732, 776, 778, 813, 836, 850, 854, 887, 892, 916, 928, 948, 951, 964, 1038, 1044, 1112, 1142, 1143, 1199, 1232, 1240, 1243, 1246, 1256, 1265, 1311, 1319, 1328, 1332, 1333, 1339, 1346, 1353, 1422, 1423, 1447, 1448, 1493, 1504, 1519, 1524, 1528, 1540, 1586, 1587, 1600, 1611, 1612, 1617, 1679, 1710, 1794, 1844, 1846, 1850, 1868, 1890, 1893, 1912, 1979, 1985, 2001, 2017, 2018, 2064, 2073, 2090, 2100, 2142, 2157, 2231, 2244, 2253, 2298, 2356, 2374, 2401, 2405, 2424, 2450, 2547, 2569, 2577, 2673, 3000, 3178, 3179, 3185, 3284, 3317, 3351, 3546]. sensor [350, 371, 403, 571, 587, 600, 626, 682, 697, 708, 715, 720, 722, 748, 791, 816, 818, 839, 851, 855, 882, 917, 932, 938, 944, 955, 968, 971, 985, 997, 1002, 1039, 1052, 1057, 1096, 1120, 1132, 1137, 1140, 1147, 1171, 1221, 1283, 1289, 1297, 1318, 1320, 1336, 1337, 1363, 1368, 1408, 1495, 1497, 1499, 1532, 1541, 1576, 1580, 1645, 1668, 1675, 1680, 1707, 1750, 1756, 1758, 1805, 1810, 1845, 1848, 1858, 1896, 1906, 1918, 1939, 2019, 2020, 2022, 2031, 2033, 2040, 2075, 2139, 2170, 2258, 2292, 2300, 2321, 2338, 2342, 2345, 2439, 2565, 2600, 2615, 2990, 3027, 3071, 3119, 3196, 3258, 3280, 3470, 3497]. sensor [430, 449, 450, 476, 489, 491, 520, 522, 525, 561, 567, 641, 676, 719, 749, 793, 795, 797, 799, 822, 837, 852, 881, 901, 914, 922, 942, 945, 959, 960, 986, 1014, 1022, 1108, 1111, 1148, 1162, 1191, 1259, 1274, 1365, 1379, 1419, 1432, 1514, 1585, 1589, 1601, 1620, 1640, 1652, 1732, 1735, 1781, 1804, 1859, 1898, 1914, 1981, 1994, 2013, 2021, 2023, 2036, 2051, 2066, 2076, 2095, 2146, 2149, 2201, 2230, 2351, 2368, 2449, 2478, 2493, 2683, 2718, 2864, 3014, 3344, 3391, 3428, 3440]. sensor-based [308]. sensor-cloud [3196]. sensor/robot [948]. sensors [730, 1496, 1663, 2348, 2487, 3302]. sensory [2435]. sentences [208]. sentient [1784]. sentiment [2117]. SentiRelated [2117]. separation [714, 1164, 1166, 1195, 3157]. September [2756, 2767, 3003, 3012, 3176, 3365, 3520]. Sequence [346, 1301, 1812, 3270]. Sequence-to-sequence [3270]. Sequential [844, 868, 1086, 2160, 2820]. Seraph [3532]. series [874]. Server [268, 311, 313, 484, 492, 609, 660, 680, 683, 741, 811, 999, 1179, 1227, 1230, 1277, 1285, 1415, 2140, 2293, 2443, 2719, 3146, 3367]. server-based [268]. server-centric [1227]. serverless [680, 962]. servers [145, 202, 261, 269, 411, 709, 1217, 1983]. Service [109, 162, 184, 199, 202, 214, 217, 264, 287, 331, 358, 378, 396, 458, 496, 540, 541, 570, 580, 586, 608, 610, 621, 625, 640, 668, 669, 692, 716, 763, 784, 814, 870, 875, 925, 943, 1031, 1032, 1085, 1089, 1105, 1163, 1302, 1335, 1382, 1416, 1477, 1482, 1522, 1573, 1614, 1676, 1701, 1715, 1722, 1776, 1784, 1791, 1884, 1887, 1903, 1996, 2000, 2007, 2012, 2034, 2134, 2174, 2211, 2255, 2311, 2376, 2379, 2430, 2494, 2595, 2644, 2672, 2704, 2743, 2759, 2863, 2922, 2931, 2959, 2974, 2987, 3000, 3069, 3076, 3109, 3121, 3128, 3150, 3306, 3357, 3417, 3466, 3500, 3533, 3540]. service [62, 140, 149, 212, 234, 260, 285, 301, 333– 336, 408, 409, 547, 553, 579, 607, 612, 652, 653, 711, 794, 831, 874, 1063, 1088, 1194, 1229, 1236, 1241, 1309, 1380, 1393, 1460, 1515, 1534, 1547, 1548, 1561, 1634, 1687, 1874, 1880, 1892, 1956, 2038, 2143, 2183, 2198, 2250, 2266, 2302, 2505, 2588, 2601, 2772, 2877, 2888, 3146, 3430, 3438, 3442, 3469, 3547]. service-centric [547]. service-level-agreement [264]. Service-oriented [199, 335, 540, 763, 794, 831, 2255, 2505]. Services [34, 58, 61, 79, 90, 107, 134, 135, 173, 177, 199, 216, 253, 270, 333, 346, 387, 391, 418, 493, 501, 519, 537, 542, 590, 622, 688, 692, 705, 783, 794, 827, 862, 1094, 1145, 1196, 1214, 1278, 1290, 1312, 1315, 1343, 1494, 1624, 1683, 1714, 1765, 1817, 1821, 1987, 1988, 2007, 2099, 2172, 2216, 2376, 2387, 2472, 2539, 2553, 2595, 2609, 2720, 2749, 2799, 2826, 3026, 3107, 3157, 3158, 3196, 3231, 3450, 3464, 3466, 3476]. servlets [943]. servlets-based [943]. Session [278, 555, 2383]. session-based [555]. sessions [393, 581]. set [322, 345, 739, 2031, 3497]. set-based [739]. set-covering-based [2031]. set-up [322, 345]. sets [889, 898, 2660]. settings [613]. setup [2529]. several [690]. several-speaker [690]. SFC [2255]. SFTO [3293]. SFTO-Guard [3293]. SGMA [2235]. Shafer [1638]. shape [1917]. shapes [2298]. SHARC [655]. sharding [3352, 3401, 3435]. Shared [264, 275, 277, 754, 1382, 1837, 3115]. shared-path [264]. shared-risk [275]. sharing [244, 289, 423, 442, 459, 558, 594, 974, 1119, 1211, 1322, 1488, 1772, 1853, 1952, 2004, 2011, 2128, 2307, 2354, 2422, 2774, 2783, 2879, 2910, 2984, 3115, 3316, 3494]. Shedding [2249]. shock [2656]. short [140, 1464, 2117, 2189]. short-distance [1464]. shortcut [3442]. shortest [546, 1820, 1919, 2141, 2318]. shortest-paths [2318]. shot [3381]. should [161]. show [674]. shuffling [3359]. shut [1936]. side [1136, 1278, 1460, 1719, 2042, 2350, 2497, 2730, 2865, 2952]. Sidechain [2623]. sided [3028]. SIFT [513]. sightseeing [1212]. sign [404, 824]. sign-on [404, 824]. signal [155, 1051, 1300, 1408, 1993, 2953, 3179]. signal-based [155]. signal-range [2953]. signaling [1441]. signalling [733]. signals [2263, 2626]. signature [486, 494, 516, 579, 905, 1477, 1908, 3234, 3532]. signature-based [1908]. signatures [834, 2123]. signcryption [772]. signed [2395]. significant [87]. Similarity [209, 795, 1567, 2002, 2395, 3197]. Simple [276, 739, 1178, 1251, 1907]. simplification [884]. simulated [102, 206]. Simulation [76, 230, 238, 254, 326, 413, 607, 642, 1127, 1161, 1168, 1231, 1557, 1558, 2036]. simulation-based [326, 1545]. simulations [1878]. simulator [5, 2851]. simulators [771]. SINET [1938]. single [363, 404, 570, 606, 624, 824, 947, 1277, 2276, 2293, 2590, 2703, 3271]. single-bit [947]. single-board [3271]. single-hop [570, 606]. single-iteration [947]. Singular [16]. sink [1151, 1493, 1620, 2180, 2342, 2449, 2565, 2600, 3062]. sinkhole [1449]. sinks [944]. SINR [1820]. SIONA [1393]. SIP [369, 468, 485, 943, 1217, 1544]. SIP-based [369, 1544]. site [67, 1875]. sites [400, 1085, 2091]. situation [325, 2520]. situation-aware [325]. situational [2199]. situations [2397, 3297]. size [463, 1116, 1221, 1512, 1812, 1882]. sized [2008]. sizes [2760]. sketch [2540, 3338]. sketch-based [2540]. skew [1147]. Skewly [1401]. skyline [563]. Skype [2453, 2721]. SLA [1307, 1831]. SLA-based [1307]. slave [2843]. sleep [1514, 1535, 1823, 1864]. sleep-scheduling [1823]. sleeping [2276]. slice [2331, 3388]. slicing [1245, 2505, 2687, 3021, 3304, 3316, 3523]. sliding [1484]. slot [395, 474, 2713, 3369]. slots [1773]. slotted [431, 2392, 3223]. slow [691, 1326, 3191, 3293]. slow-rate [3191, 3293]. slow-start [691]. SlowDoS [2870]. SMAC [1022]. SMAC-based [1022]. Small [385, 910, 2862, 2935, 3183]. small-cell [3183]. small-scale [385]. SMART [159, 273, 483, 560, 629, 811, 853, 856, 1018, 1056, 1260, 1296, 1323, 1356, 1438, 1466, 1554, 1556, 1557, 1559, 1560, 1733, 1745, 1778, 1809, 1813, 1985, 2070, 2155, 2218, 2279, 2299, 2303, 2316, 2331, 2349, 2406, 2421, 2456, 2460, 2462, 2464, 2472, 2479, 2481, 2484, 2486, 2487, 2500, 2504, 2507, 2514, 2516, 2560, 2566, 2576, 2646, 2659, 2665, 2684, 2696, 2699, 2766, 2818, 2819, 2891, 2897, 2908, 2916, 2930, 2956, 2977, 3048, 3061, 3096, 3173, 3215, 3246, 3260, 3296, 3302, 3335, 3503, 3513, 3519, 3523, 3550]. smart-wireless [3302]. smartcard [52]. SmartDR [2832]. SmartEdge [2500]. smarter [3455]. SmartITS [2085]. Smartphone [1519, 1631, 1663, 1840, 1885, 2085, 2199, 2200, 2264, 2533, 3172, 3185, 3496].

Smartphone-based [1519, 1631, 1885, 2085, 2264]. smartphone-embedded [3185]. smartphones [1288, 1702, 2278, 2408, 2487, 2689, 2765]. SmartSteganogaphy [2766]. SmartValidator [3143]. smileys [674]. SMS [433]. SMS-oriented [433]. SMTP [393]. SNMP [259]. Snort [484]. SOA [647, 663]. SOAP [775, 1207, 1214]. SOARware [647]. Social [98, 122, 148, 364, 374, 453, 535, 703, 712, 746, 812, 823, 1001, 1034, 1084, 1085, 1183, 1211, 1218, 1266, 1403, 1456, 1476, 1480, 1553, 1703, 1727, 1730, 1772, 1803, 1854, 1948, 1949, 1951–1957, 1970, 2132, 2168, 2216, 2232, 2309, 2317, 2370, 2395, 2454, 2461, 2619, 2681, 2770, 2807, 2847, 2896, 2920, 2927, 2991, 3067, 3130, 3132, 3185, 3259, 3303, 3467, 3473]. Social-aware [535, 2619]. Social-based [1970, 2317]. Social-connectivity-aware [1034]. SocialHide [1814]. Socially [3169]. Socially-aware [3169]. socio [2371]. socio-inspired [2371]. Sociopsychological [1600]. socket [487]. soft [227, 246, 559, 733, 2505, 3267]. SoftIoT [3054]. Softswitches [2198]. Software [17, 33, 157, 178, 426, 540, 551, 644, 805, 865, 962, 1213, 1446, 1486, 1531, 1591, 1592, 1619, 1678, 1693, 1697, 1711, 1714, 1725, 1768, 1837, 1870, 1874, 1910, 1940, 1961, 2000, 2008, 2045, 2131, 2152, 2229, 2243, 2272, 2285, 2300, 2329, 2333, 2334, 2375, 2382, 2398, 2434, 2445, 2491, 2494, 2504, 2509, 2532, 2535, 2598, 2599, 2631, 2660, 2672, 2680, 2729, 2758, 2800, 2825, 2844, 2876, 2905, 2917, 2919, 2942, 2972, 2995, 2996, 3006, 3022, 3041, 3045, 3049, 3060, 3066, 3102, 3161, 3187, 3249, 3278, 3317, 3319, 3344, 3386, 3436, 3474, 3485, 3487, 3543]. Software-base [2509]. Software-based [962, 2398]. Software-Defined [426, 865, 1446, 1591, 1693, 1725, 2000, 2008, 2272, 2300, 2329, 2333, 2334, 2382, 2491, 2494, 2535, 2599, 2631, 2672, 2729, 2825, 2876, 2905, 2917, 2996, 3022, 3045, 3049, 3161, 3187, 3249, 3278, 3317, 3319, 3344, 3436, 3485, 3487, 3543]. softwarized [3450]. solar [859, 2001]. solar-powered [859]. SOLS [1482]. Solution [622, 634, 810, 853, 872, 937, 1011, 1014, 1151, 1229, 1324, 1518, 1647, 2047, 2223, 2275, 2610, 2952, 3242, 3414, 3536, 3538]. Solutions [600, 648, 857, 879, 1175, 1290, 1376, 1385, 1391, 1408, 1652, 1795, 2054, 2212, 2532, 2572, 2613, 2628, 2722, 2923, 2960, 3068, 3078, 3086, 3152, 3188, 3337, 3439, 3519]. solve [1092, 1914, 3549]. Solving [614, 2024, 2833, 3487]. SOM [2008]. SONR [826]. sound [632]. Source [714, 781, 1788, 1949, 2497, 2642, 2658, 3034, 3322, 3398]. source-based [781]. source-multicast [2658]. source-side [2497]. sources [134, 1160, 1745, 2228]. sourcing [1420, 2432]. southbound [3436]. sovereign [3274]. Space [159, 539, 949, 1011, 1245, 1375, 1380, 1585, 2103, 2337, 2417, 2813, 2874, 2878, 3050, 3113, 3324, 3428, 3469]. space-air-ground [3469]. space-aware [2337]. Space4time [1236]. spam [379, 393, 712, 1655, 2232, 2943]. spam-advertised [379]. spamming [1272]. sparrow [3425]. sparse [670, 2204, 2323, 3130]. sparsity [1719]. Spatial [863, 882, 1585, 2350, 2400, 2402, 2869, 2904, 3354, 3403]. Spatial-temporal [863, 2869, 2904, 3403]. spatially [1340, 3030]. Spatio [904, 1153, 2485, 2735, 2752, 3389]. Spatio-temporal [904, 1153, 2485, 2735, 2752, 3389]. Spatiotemporal [2357, 2511]. speaker [690]. SpeCH [2539]. Special [31, 81, 86, 118, 119, 141, 241, 408, 501, 503, 541, 590, 821, 870, 1335, 1390, 1531, 1948, 2017, 2214, 2506, 2789]. specific [207, 831, 2473, 2980]. specification [24, 54, 333, 411, 1145]. specifications [133]. Specifying [129, 305, 2588]. spectral [1799, 2881]. Spectrum [750, 1035, 1119, 1221, 1238, 1424, 1470, 1501, 1590, 1700, 1754, 1764, 1766, 1836, 2011, 2047, 2103, 2128, 2171, 2368, 2417, 2505, 2528, 2551, 2751, 3194, 3354]. spectrum-agile [1501]. spectrum-based [1501]. spectrum-sharing [1119]. Speculative [1227, 1450]. speech [55, 59, 478, 540, 690]. speech-enabled [540]. speed [249, 684, 1029, 1273, 1304, 1464, 1728, 2546, 3339]. SPFP [2372]. spiral [1363]. Split [3180]. SplitStream [3488]. splitting [1909]. SPO [2903]. spoken [318]. spontaneous [200, 597]. spot [1494, 1658, 1973]. Spread [605, 750, 1501, 2991, 3015, 3303]. square [1002]. SRA [1042]. SRA-MSDU [1042]. SRLG [275, 777]. SRP [3335]. SRv6 [2658]. SSDF [3194]. SSDH [1754]. SSO [445]. stability [422, 685, 926, 1155, 1351, 2026]. stability-based [685, 2026]. stabilize [1006]. stable [218, 322, 2918]. stack [2018]. Stackelberg [2215, 3304]. stacking [3009]. stage [435, 786, 1011, 1092, 1301, 2223, 2704, 2847, 3011, 3145]. stamped [1922]. stand [943]. stand-alone [943]. standalone [2895]. standard [615, 664]. standardization [67, 1809]. Standards [824, 1290, 3246]. standards-based [824]. star [570, 606]. Starfish [2342]. start [691, 3440]. starvation [139]. starvation-free [139]. State [323, 498, 996, 1011, 1123, 1155, 1196, 1309, 1453, 1460, 1506, 1574, 1684, 1740, 1763, 2069, 2212, 2232, 2303, 2330, 2623, 2703, 2745, 2769, 2771, 2808, 2837, 2866, 2974, 2983, 2987, 3029, 3203, 3352, 3435, 3503, 3508]. State-of-art [2745]. State-of-the-Art [1309, 1453, 1460, 1506, 1684, 1740, 1763, 2212, 2232, 2623, 2771, 2866, 3029, 3352]. Stateful [286, 2035, 2429]. stateless [2426]. static [987, 1131, 1510, 1927, 2652, 2781]. station [2320, 2561, 3421]. stationary [2946]. stations [1614, 2228, 2276, 2881, 3055]. statistic [1152]. statistic-based [1152]. statistical [258, 555, 1226, 1245]. statistics [710]. statistics-based [710]. Status [1054, 1467, 1865, 1961, 2041, 2476, 2591]. STBC [1023]. Stealthy [2007]. Steering [2198]. steganographic [1622]. steganography [785, 2766, 3359]. step [1636, 2552, 3385]. stimulation [1451]. stitching [959, 1475]. Stochastic [804, 1879, 2812, 3102, 3140, 3487]. stochastical [2803]. stone [2094]. stop [2831, 3015, 3424]. Storage [107, 109, 139, 462, 467, 509, 816, 827, 910, 1133, 1197, 1224, 1278, 1306, 1371, 1401, 1425, 1550, 1555, 1664, 1796, 2037, 2069, 2138, 2154, 2196, 2212, 2246, 2404, 2423, 2426, 2470, 2485, 2714, 2732, 3253]. Storages [1705]. store [2688]. storing [910, 1949, 3037]. storm [817, 1356, 1966]. story [41]. stove [3537]. strategies [575, 815, 915, 1009, 1322, 1399, 1435, 1471, 1474, 1640, 1848, 2014, 2169, 2224, 2382, 2595, 2650, 2914, 2996, 3444, 3504, 3522]. Strategizing [3037]. Strategy [499, 605, 618, 861, 902, 950, 996, 1036, 1092, 1149, 1170, 1283, 1375, 1388, 1617, 1620, 1643, 1680, 1706, 1709, 1764, 2132, 2330, 2345, 2375, 2463, 2607, 2927, 2929, 3032, 3075, 3145, 3184, 3195, 3203, 3248, 3306, 3332, 3385, 3399]. Strategy-proof [3075]. stream [220, 249, 281, 528, 1352, 1566, 1630, 1852, 2107, 2126, 2683, 3032, 3203]. stream-based [2107]. streamed [295]. Streaming [181, 183, 255, 267, 269, 452, 581, 695, 709, 747, 923, 1041, 1098, 1159, 1250, 1299, 2048, 2074, 2179, 2185, 2314, 2360, 2396, 2691, 2885, 3084, 3192, 3262, 3276, 3295, 3459, 3509]. streams [17, 844, 868, 1905, 2360]. street [938, 1334]. strength [3104]. strict [3108]. stride [1450]. stride- [1450]. strings [506]. strong [858, 1477, 1595]. Structural [141, 167–170, 172, 173, 929, 1917]. Structure [11, 1018, 1112, 1371]. Structure-free [1018, 1112]. structured [32, 537]. Structures [1140, 2123, 2185, 3254]. Structuring [174]. STSIR [3303]. Student [101, 1437]. Student-oriented [1437]. students [1213]. studies [33]. Study [37, 235, 391, 453, 460, 494, 583, 759, 857, 897, 994, 1127, 1166, 1231, 1235, 1246, 1273, 1325, 1355, 1391, 1476, 1498, 1508, 1717, 1726, 1785, 1841, 1955, 2310, 2525, 3030, 3066, 3115, 3146, 3322, 3409]. SU-MIMO [2213]. Sub [515, 651, 1135, 1183, 1878, 3268]. Sub-beat [3268]. sub-communities [1183]. sub-department [1135]. sub-optimal [1878]. sub-pattern [515]. sub-query [651]. subband [568]. subchannel [1762]. subgraph [3264]. subject [64]. subjective [2179]. subscribe [653]. subscriber [1614]. subscription [1062]. subset [576, 3496]. subspace [564, 900]. Subsystems [2155]. subtopic [2827]. subtraction [1136]. successful [161]. suggesting [1821]. Suitability [1117, 2933]. sum [2009]. summarization [1109]. Super [668, 2778]. Super-peer-based [668]. super-resolution [2778]. superposition [787]. supervised [1825, 2199, 2415, 2462, 2723, 2765, 3009, 3434, 3556]. supplementary [2526]. supply [351, 642, 1070]. Support [62, 77, 79, 81, 97, 118, 119, 121, 122, 125, 233, 325, 334, 337, 367, 377, 541, 569, 624, 669, 673, 700, 757, 831, 1096, 1220, 1312, 1342, 1379, 1403, 1431, 1664, 2184, 2280, 2308]. supported [107, 388, 695, 841, 1803, 3290]. Supporting [47, 120, 158, 231, 314, 362, 418, 444, 451, 643, 689, 992, 994, 1032, 1327, 1464, 1481, 1754, 1974, 2049, 2245, 2260, 2470, 3043, 3167, 3274, 3473]. supports [813]. surfaces [3352, 3400]. surfing [2115]. Surgical [3444]. Surveillance [198, 308, 626, 720, 840, 1106, 1132, 1320, 1334, 1538, 1638, 2231, 2452, 2526, 2568, 2899, 3126, 3174, 3434, 3545]. Survey [417, 472, 669, 732, 735, 763, 836, 847, 991, 1122, 1124, 1325, 1330, 1332, 1415, 1461, 1473, 1572, 1583, 1584, 1587, 1605, 1612, 1653, 1673, 1682, 1691, 1695, 1737, 1749, 1755, 1774, 1778, 1791, 1798, 1800, 1807, 1813, 1823, 1835, 1846, 1890, 1897, 1913, 1925, 1978, 1979, 2042, 2045, 2054, 2061, 2073, 2080, 2152, 2241, 2249, 2289, 2310, 2312, 2391, 2430, 2507, 2509, 2547, 2572, 2598, 2615, 2620, 2624, 2628, 2642, 2664, 2667, 2705, 2710, 2732, 2819, 2831, 2839, 2916, 2933, 2966, 2980, 3058, 3059, 3078, 3079, 3091, 3121, 3127, 3142, 3152, 3209, 3219, 3220, 3224, 3276, 3317, 3328, 3408]. survey [148, 599, 600, 701, 767, 880, 935, 936, 982, 984, 985, 1013, 1143, 1174, 1239– 1241, 1351, 1360, 1376, 1422, 1453, 1471, 1508, 1510, 1573, 1575, 1576, 1579, 1590, 1661, 1678, 1697, 1731, 1739, 1763, 1801, 1809, 1818, 1829, 1844, 1874, 1959, 1980, 2012, 2044, 2063, 2069, 2114, 2124, 2128, 2145, 2164, 2207, 2300, 2339, 2374, 2392, 2414, 2456, 2516, 2521, 2549, 2551, 2597, 2613, 2680, 2690, 2728, 2742, 2769, 2811, 2815, 2833, 2876, 2891, 2932, 2934, 2947, 2983, 3042, 3098, 3107, 3113, 3114, 3150, 3153, 3156, 3171, 3206, 3247, 3273, 3299, 3336, 3337, 3352, 3384, 3387, 3396, 3398, 3504, 3517, 3568]. survey [457, 491, 553, 660, 676, 733, 749, 799, 837, 879, 881, 895, 896, 1203, 1205, 1244, 1261, 1268, 1305, 1315, 1359, 1451, 1457, 1629, 1637, 1640, 1654, 1666, 1686, 1713, 1732, 1735, 1740, 1795, 1802, 1817, 1819, 1856, 1924, 2014, 2036, 2046, 2051, 2075, 2081, 2095, 2106, 2126, 2131, 2135, 2149, 2168, 2190, 2212, 2270, 2287, 2294, 2327, 2427, 2442, 2445, 2504, 2525, 2532, 2563, 2589, 2626, 2669, 2698, 2709, 2722, 2802, 2807, 2814, 2834, 2844, 2846, 2905, 2930, 2958, 2964, 2991, 2994, 3001, 3008, 3080, 3117, 3139, 3182, 3285, 3331, 3450, 3476, 3503]. survivability [592, 999, 1346, 2100, 2468]. Survivable [256, 277, 344, 969, 1027, 3186]. suspicious [3259]. sustainable [1008, 1915, 2893, 3445]. SVM [2934]. SVM-based [2934]. Swarm [637, 679, 682, 806, 881, 914, 977, 1129, 1257, 1423, 1710, 2142, 2230, 2503, 3025, 3071, 3174, 3420]. swarm-based [806]. swarming [1650]. swindle [1169]. swipe [2765]. swipe-based [2765]. switch [521, 786, 1301, 2375, 2758, 3102, 3443]. switch-device [521]. switch-off [2375]. switch-over [3102]. switched [521, 1252, 1720, 1984]. Switches [947, 2329, 2364, 2491, 2546, 2760, 2894, 3238]. switching [60, 872, 1546, 2271, 3049]. sybil [1730, 2312]. Sybil-resilience [1730]. SyD [536]. symbiotic [2463]. symbol [426]. symbolic [143]. symmetric [1376, 1949]. SYN [356, 3367]. synchronization [1284, 1314, 1893, 1917, 2165, 2478]. synchronized [817, 1884, 2896]. synchronizers [426]. Synchronizing [3525]. Synchronous [1142]. synergy [1887, 3560]. syntax [2725]. syntax-aware [2725]. Synthetic [3202, 3529, 3568]. System [10, 14, 16, 19, 24, 39, 43, 44, 82, 84, 131, 133, 157, 158, 183, 194, 213, 230, 246, 248, 257, 325, 410, 413, 432, 433, 448, 473, 478, 480, 500, 545, 587, 601, 610, 624, 663, 703, 784, 833, 840, 933, 934, 951, 973, 979, 997, 1018, 1051, 1122, 1133, 1164, 1165, 1186, 1265, 1334, 1362, 1371, 1438, 1527, 1559, 1575, 1695, 1803, 1811, 1860, 1863, 1971, 1997, 2015, 2028, 2037, 2273, 2280, 2391, 2406, 2460, 2489, 2494, 2526, 2533, 2552, 2634, 2791, 2880, 2918, 2919, 2941, 3034, 3112, 3190, 3196, 3234, 3246, 3342, 3356, 3396, 3423, 3449, 3473, 3489, 3494]. system [25, 35, 69, 139, 146, 174, 198, 208, 280, 303, 312, 336, 340, 372, 389, 399, 404, 438, 633, 677, 678, 806, 826, 954, 965, 1065, 1098, 1149, 1169, 1209, 1259, 1313, 1439, 1459, 1490, 1540, 1543, 1555, 1560, 1565, 1643, 1658, 1712, 1748, 1871, 1886, 1898, 1952, 1953, 1990, 2112, 2146, 2601, 2821, 2857, 3061, 3094, 3293, 3368, 3388, 3510, 3535, 3557]. system-level [1575]. Systematic [665, 883, 920, 934, 986, 1045, 1304, 1317, 1432, 1736, 1880, 2068, 2073, 2170, 2211, 2311, 2384, 2553, 2592, 2659, 2692, 2818, 2856, 3000, 3019, 3044, 3068, 3218, 3287, 3318, 3330, 3395, 3404, 3427, 3481]. Systems [8, 15, 71, 77, 89, 103, 228, 245, 291, 302, 326, 332, 361, 440, 467, 501, 593, 634, 635, 645, 651, 666, 667, 688, 725, 753, 832, 847, 880, 885, 921, 1041, 1071, 1075, 1099, 1100, 1105, 1172, 1198, 1202, 1231, 1239, 1253, 1285, 1291, 1294, 1314, 1425, 1510, 1550, 1559, 1599, 1605, 1630, 1661, 1664, 1719, 1755, 1830, 1836, 1838, 1883, 1891, 1895, 1929, 1930, 1945, 1969, 2141, 2154, 2162, 2165, 2193, 2262, 2332, 2415, 2431, 2451, 2521, 2610, 2684, 2692, 2695, 2775, 2808, 2818, 2934, 2966, 2997, 3017, 3033, 3124, 3125, 3236, 3245, 3313, 3315, 3320, 3513, 3528, 3545]. systems [127, 128, 167, 192, 197, 220, 248, 289, 300, 305, 465, 517, 537, 564, 618, 628, 660, 665, 712, 751, 761, 830, 890, 946, 1082, 1130, 1158, 1227, 1421, 1531, 1682, 1774, 1886, 1911, 1933, 1959, 2266, 2279, 2355, 2414, 2479, 2706, 2720, 2856, 2859, 2988, 3001, 3080, 3099, 3203, 3227, 3318, 3336, 3339, 3377, 3404, 3435, 3529].

T [1905]. T-Hoarder [1905]. table [988, 1224, 2225, 2849, 3293, 3359]. table-driven [3359]. tables [383]. Tabu [189]. tactical [2449, 3221]. tag [475, 639, 1135, 1262, 2713]. tagged [192]. tags [2153, 2451, 2713]. Taiwan [325]. Taking [488, 761]. tandem [738, 1011]. TAO [165]. Tap [2292]. Target [449, 951, 955, 1052, 1495, 1497, 1652, 1770, 1781, 1782, 2328, 2451, 2511, 2733, 2825, 2904, 3102, 3104, 3188]. targeted [286, 1185, 3261]. targeting [1363]. targets [1319]. Task [206, 366, 654, 660, 955, 1075, 1388, 1404, 1491, 1533, 1538, 1920, 2129, 2230, 2350, 2416, 2432, 2463, 2907, 3002, 3028, 3125, 3165, 3272, 3334, 3377, 3394, 3410, 3468, 3526, 3539]. task-oriented [1075]. tasks [238, 1216, 1881, 2097, 2482, 2906, 2918, 3162, 3372, 3533]. taxi [2752]. taxonomic [2081]. taxonomies [1418, 2186]. Taxonomy [260, 491, 663, 767, 880, 940, 1142, 1190, 1202, 1204, 1277, 1369, 1373, 1426, 1427, 1506, 1515, 1577, 1607, 1674, 1740, 1774, 1776, 1813, 1959, 1982, 2046, 2079, 2109, 2208, 2244, 2365, 2398, 2418, 2571, 2624, 2667, 2798, 2833, 2934, 2942, 2947, 2983, 2996, 3121, 3182, 3242, 3458, 3502]. TBDClust [2091]. Tc [2412]. Tc-PEDCKS [2412]. TCP [360, 742, 819, 894, 941, 1215, 1231, 1237, 1273, 1436, 1464, 1490, 1593, 1692, 1698, 1724, 1741, 2027, 2189, 2362, 2434, 2444, 2572, 2884, 3333, 3350, 3367, 3436, 3535]. TCP-FIT [1741]. TCP/AQM [1231, 1490, 3535]. TCP/IP [819, 941]. TCP/RED [742]. TD [3168]. TD-RA [3168]. TDD [2505]. TDMA [529, 1110, 2019, 3369]. TDMA-based [1110, 3369]. teachers [231]. teaching [47, 120]. team [229]. team-based [229]. teaming [226, 299]. Teams [228]. teamwork [301]. technical [57, 939, 3113]. technique [233, 271, 274, 366, 1003, 1033, 1160, 1668, 1989, 2129, 2137, 2369, 2401, 2579, 2672, 3033, 3112, 3138, 3208]. Techniques [220, 334, 457, 469, 526, 599, 732, 877, 911, 935, 936, 986, 1191, 1213, 1275, 1278, 1303, 1367, 1492, 1572, 1576, 1577, 1644, 1653, 1689, 1691, 1736, 1818, 1835, 2109, 2143, 2179, 2208, 2440, 2549, 2582, 2664, 2785, 2789, 2819, 2876, 2933, 2958, 2997, 3044, 3059, 3146, 3170, 3182, 3206, 3258, 3277, 3285, 3307, 3336, 3370, 3547]. Technologies [939, 940, 992, 1061, 1077, 1126, 1244, 1778, 1913, 2054, 2080, 2186, 2620, 2623, 2684, 2802, 2814, 2833, 2923, 2932, 3201, 3218, 3299, 3384, 3452, 3458, 3481, 3519]. Technology [62, 148, 160, 183, 343, 628, 846, 997, 1219, 1386, 2033, 2076, 2099, 2486, 2745, 2839, 2886, 3018, 3484]. Teeth [511]. tele [120, 2262, 2266]. tele-health [2262, 2266]. Tele-Teaching [120]. telecardiology [578]. Telecare [1959]. Telecom [542]. telecommunication [105]. Telehealth [2281]. Telemedicine [846, 2297, 2436, 2866]. telemonitoring [1325, 2291]. teleoperation [180, 219]. telephony [71]. telepresentations [96]. teleprotection [2514]. telescope [3066]. television [531, 1178]. TelosB [1320]. TELPAC [2602]. Temperature [2973]. Temperature-aware [2973]. template [900, 2261]. templates [170, 506]. Temporal [10, 165, 863, 904, 960, 1153, 1271, 1478, 1810, 2064, 2485, 2673, 2735, 2752, 2869, 2904, 3040, 3286, 3389, 3403]. temporal-credential-based [960, 1810]. Temporary [2952]. tenant [1690, 2037, 2581, 3466]. Tenant-Defined [2037]. tenants [2508]. Tensor [2877, 3130, 3463]. tensor-CUR [3463]. term [1550, 2512]. Termination [1542]. Termite [914]. Termite-hill [914]. terms [3242]. ternary [1058]. Terra [3518]. terrain [2646]. terrestrial [3422]. terrestrial-satellite [3422]. terrorism [1128]. terrorism-related [1128]. test [112, 448, 805, 2837]. testbed [791, 1166, 2171]. tester [29]. testing [390, 1224, 3478]. tests [527, 582]. tethering [1258]. Text [53, 113, 144, 1128, 2626, 2786]. text-based [113]. texts [2117]. Thank [2111, 2367, 2618, 2855, 3083]. Thankyou [3241, 3407]. Theatres [746]. Themis [170]. theoretic [557, 1245, 2319]. theory [98, 623, 1169, 1211, 1448, 1551, 1638, 2348, 3222, 3248]. there [3104]. thin [890, 1227, 1633]. Things [796, 912, 995, 1001, 1073, 1088, 1177, 1260, 1261, 1383, 1552, 1637, 1673, 1684, 1734, 1752, 1763, 1839, 1888, 1904, 1924, 1978, 1987– 1991, 1995, 1998, 2068, 2070, 2096, 2139, 2181, 2310, 2325, 2328, 2343, 2349, 2365, 2412, 2419, 2429, 2480, 2481, 2503, 2507, 2512, 2523, 2525, 2527, 2562, 2563, 2571, 2636, 2692, 2706, 2722, 2739, 2742, 2783, 2789, 2806, 2845, 2848, 2851, 2856, 2898, 2964, 2994, 3042, 3087, 3126, 3134, 3171, 3206, 3209, 3250, 3287, 3303, 3309, 3318, 3354, 3360, 3383, 3410, 3418, 3445, 3448, 3492, 3505, 3519, 3527, 3528, 3537, 3569]. thinking [47, 754]. third [337, 1894, 2730, 3536]. third-party [2730, 3536]. Thompson [3195]. ThorFI [3128]. threat [1383, 3143, 3159, 3448]. threats [857, 1131, 1795, 2144, 2215, 2456, 2516]. Three [419, 465, 786, 974, 976, 999, 1012, 1079, 2018, 2139, 2577, 2847, 3026]. Three-dimensional [2018, 2577]. three-factor [2139]. Three-layer [419]. three-party [976, 1012]. three-stage [786, 2847]. threshold [1593, 2649, 3532]. thrombosis [2291]. Throughput [360, 464, 802, 809, 843, 989, 1558, 1771, 1790, 2637, 3453, 3483]. Throwboxes [2014]. ticket [151]. tier [555, 961, 1416, 1423, 1483, 2083, 2287, 3313, 3377, 3388, 3467, 3501]. tiered [2031]. tight [963]. tiled [3129]. Time [24, 55, 82, 122, 153, 156, 226, 237, 259, 368, 395, 527, 528, 559, 578, 645, 670, 675, 689, 733, 804, 807, 808, 836, 840, 884, 915, 976, 1023, 1041, 1135, 1161, 1177, 1218, 1224, 1287, 1314, 1375, 1425, 1469, 1556, 1641, 1699, 1800, 1830, 1860, 1884, 1891, 1893, 1917–1919, 1922, 1930, 1966, 1984, 1995, 2091, 2096, 2107, 2124, 2165, 2368, 2378, 2392, 2408, 2412, 2441, 2482, 2602, 2608, 2612, 2637, 2870, 3020, 3038, 3232, 3293, 3353, 3367, 3372, 3456, 3482, 3525]. Time-based [2091]. time-bounded [2368]. time-constrained [1287]. Time-continuous [3456]. time-correlated [1177]. time-critical [1918]. time-dependent [1919]. Time-efficient [3353]. Time-Lagged [1469]. Time-optimized [2378, 2612]. time-selective [1023]. time-slot [395]. time-triggered [670]. Timeline [3225]. timeliness [1613]. timely [1339, 2831]. timeouts [2902]. timer [1991]. times [874]. timescale [1634]. timing [2567]. timing-based [2567]. TM [1760]. ToF [2096]. together [760]. token [70, 2099]. token- [70]. tolerance [1038, 1296, 1584, 2541, 2824, 3103]. Tolerant [407, 618, 681, 779, 825, 860, 875, 902, 1008, 1059, 1115, 1170, 1191, 1200, 1240, 1287, 1420, 1442, 1523, 1686, 1707, 1970, 2014, 2061, 2246, 2555, 2578, 2614, 2988, 3302, 3351, 3475, 3522]. tolerate [1936]. tomography [1475]. tomorrow [1941]. tool [5, 17, 84, 92, 145, 954, 1513, 1618, 2094, 2408]. toolchain [3053]. toolkit [239, 1936, 2000]. tools [33, 842, 879, 911, 1202, 2953, 3395, 3398]. top [1229]. topic [207, 2827]. topic-specific [207]. topics [650]. Topological [417]. topologies [1717, 1723, 2319]. Topology [206, 739, 797, 807, 854, 917, 920, 992, 1039, 1100, 1151, 1264, 1397, 1408, 1419, 1435, 1756, 1760, 1916, 1966, 2338, 2468, 3109, 3174, 3220, 3351]. Tor [2249, 3277]. TOSSIM [1243]. total [2242]. Touch [59, 2278, 3029, 3150, 3319]. Touch-based [3029]. TouchWB [2278]. tourism [338, 1172]. tourist [2422]. tourists [1853]. Tournament [2774]. tours [1212]. TPA [2426]. trace [793, 3437, 3529]. Traceability [3154]. traceable [495]. traceback [410, 750, 1086]. traceroute [2754]. traces [1314, 2837]. Tracing [150, 2953, 3144, 3416]. track [884, 2441]. tracker [156]. Tracking [153, 222, 527, 840, 900, 906, 951, 1052, 1265, 1319, 1441, 1712, 1926, 2085, 2096, 2153, 2328, 2904, 3368, 3391, 3437]. trade [517, 1523, 2578]. trade-off [1523]. trade-offs [517, 2578]. Tradeoff [1196, 1832, 1836]. TRADER [1292]. trading [1466, 1657, 2952, 2977, 3304]. traditional [1330, 1427]. Traffic [38, 74, 210, 218, 267, 287, 523, 526, 530, 573, 596, 620, 689, 714, 765, 804, 834, 864, 902, 963, 994, 1010, 1018, 1040, 1049–1051, 1109, 1193, 1195, 1201, 1207, 1234, 1387, 1409, 1435, 1487, 1512, 1527, 1751, 1807, 1808, 1812, 1866, 1909, 1919, 1960, 1963, 1966, 1967, 2133, 2184, 2198, 2251, 2305, 2320, 2354, 2381, 2409, 2440, 2459, 2466, 2497, 2529, 2531, 2543, 2564, 2587, 2634, 2661, 2669, 2744, 2777, 2812, 2865, 2869, 2870, 2874, 2971, 3006, 3027, 3031, 3039, 3041, 3108, 3133, 3141, 3226, 3228, 3289, 3298, 3338, 3358, 3381, 3399, 3403, 3463, 3477, 3529, 3538, 3547, 3550, 3559]. Traffic-flow [2497]. traffics [700, 2295, 2389]. TrailChain [3154]. trailing [564]. train [2441, 3426]. trained [2790]. Training [41, 80, 1430, 3529]. traits [512]. trajectories [2752]. trajectory [44, 1138, 1642, 1899, 2216, 2794, 2975, 3202, 3568]. Transaction [886, 2562, 2908]. Transaction-based [886]. transactions [424, 3016, 3259]. transceiver [2590]. transcoding [322, 427, 2495, 2850]. Transfer [191, 194, 281, 392, 436, 447, 454, 458, 800, 1352, 1937, 3198, 3402, 3429, 3511, 3530]. transferring [201]. transfers [2027]. transform [317, 973, 1201, 3155]. Transformation [136, 852, 3173]. transformations [170]. transformed [2564]. Transformer [3437]. Transformer-based [3437]. transforming [47]. transient [461, 3209]. transit [2473]. Transition [1196, 1878]. Transitioning [3436]. translation [146]. transmission [528, 691, 807, 871, 937, 1114, 1281, 1292, 1347, 1352, 1525, 1680, 1754, 1984, 2219, 2565, 2798, 2861, 3223, 3258, 3382, 3416, 3424, 3561]. transmissions [3563]. transmit [2116]. transmit-only [2116]. Transmitter [2011]. transmitting [959]. transparent [2570, 3279, 3494]. Transport [68, 942, 1323, 1394, 1797, 1979, 2161, 2189, 2502]. transport-layer [1797]. transportation [1643, 1945, 2024]. transports [633]. trapdoor [554]. TRAQR [2967]. travel [3394]. traversal [1037]. Treading [647]. treatment [2421]. Tree [70, 519, 532, 921, 944, 1177, 1336, 1922, 2108, 2185, 3249]. tree-based [70, 1336]. trees [322, 1837]. Trend [3000]. Trends [93, 331, 534, 911, 1448, 1467, 1644, 1736, 1982, 2017, 2046, 2212, 2657, 2798, 2808, 2898, 2960, 3044, 3058, 3151, 3170, 3315, 3323, 3329]. triage [475]. trickle [1991]. TridentShell [3454]. trie [1406]. triggered [670, 2124]. trip [828, 2631, 3353]. tripartite [3557]. triple [391, 554]. triple-trapdoor [554]. Trojans [2784]. troops [3372]. trouble [3009]. Troubleshooting [366, 3538]. TRP [2108]. Trust [353, 557, 767, 784, 822, 828, 835, 956, 1232, 1261, 1456, 1540, 1600, 1603, 1714, 1825, 1953, 2103, 2148, 2156, 2344, 2355, 2395, 2507, 2571, 2668, 2863, 2888, 2967, 3170, 3296, 3355, 3374, 3415, 3524]. Trust-based [353, 956, 1456, 2156, 3524]. trust-enhanced [1603]. Trust-minimized [3415]. trusted [492, 662, 821, 834, 1133, 2843]. trustworthiness [2527]. trustworthiness-enhanced [2527]. Trustworthy [1571, 2046, 2387, 2442, 2977, 3492, 3494]. truth [3217, 3355, 3449]. truthful [830]. TSC [1284]. TSCH [2976]. TSDLMRA [189]. TSOIA [995]. tswTCM [866]. TTAF [3377]. TTL [707]. tunable [623]. tunnel [329]. tunneling [3548]. turbines [1868]. tutorial [984, 2442, 2831, 2878]. Tutoring [25]. TV [661, 692, 2103]. tweet [3104]. tweets [3104]. twin [3357, 3505]. twin-driven [3505]. twins [3117, 3481, 3561]. Twitter [1905, 3104]. Two [193, 319, 351, 514, 597, 622, 787, 825, 930, 1011, 1270, 1301, 1416, 1423, 1499, 1634, 1705, 1810, 1930, 2031, 2094, 2223, 2284, 2719, 2724, 2951, 2975, 3028, 3074, 3145, 3247, 3313, 3377, 3385, 3439, 3467]. two-channel [825]. two-dimensional [514, 930, 1499]. two-factor [1810, 2719]. Two-layer [2284, 3074]. two-layered [319]. Two-level [622, 2975]. two-party [2724]. Two-phase [1705, 1930, 2951]. Two-sided [3028]. Two-stage [1301, 2223, 3145]. two-step [3385]. Two-tier [1416, 1423, 3313, 3377, 3467]. two-tiered [2031]. two-timescale [1634]. two-way [787]. type [272, 2908]. types [2543, 2587, 3225].

U [3030]. U/NR [3030]. UAV [1929, 2061, 2347, 2526, 2771, 3008, 3127, 3174, 3189, 3330, 3337, 3453, 3526, 3563, 3567]. UAV-Aided [3330, 3453]. UAV-assisted [2347, 3189]. UAV-based [1929, 3337]. UAV-Enabled [3127]. UAVs [1127, 1942, 2487, 2802, 3299]. Ubiquitous [140, 162, 293, 294, 378, 496, 565, 613, 853, 905, 1220, 2036, 2406]. UD-WCMA [2001]. UDDI [333]. UDP [360, 2434, 2581]. ugly [3224]. UHF [2153]. Ultra [730, 780, 2054, 2380, 2579, 3167, 3527]. ultra-dense [2054, 3527]. Ultra-Dispatcher [780]. ultra-low [3167]. ultra-wideband [730]. ultralightweight [817, 1210]. ultrasound [22]. unattended [697, 1346, 2100]. Uncertain [995, 1195, 3389]. uncertainties [3184]. Uncertainty [1274, 1770, 2585, 3221]. Uncertainty-aware [1770, 2585]. Uncovering [2975, 3562]. underground [2170]. underlaid [2223]. underlaying [2098]. underprovisioned [2452]. understanding [435]. Underwater [732, 912, 917, 1408, 1532, 1735, 1764, 1805, 1846, 1850, 2017– 2023, 2480, 2615, 2864, 3098, 3178, 3287, 3288, 3323, 3351, 3391, 3555]. Undo [1071]. Undo/Redo [1071]. unequal [1977]. uneven [3400]. Unfolding [3225]. unforgeability [2776]. unicast [323, 702, 871, 1255]. unidirectional [992]. unification [1294]. Unified [188, 873, 1380, 1682, 2148, 2895]. Uniform [801, 3105]. uniformly [321]. Unit [249, 478, 1790, 3377, 3409]. unit-based [3377]. units [314, 449, 2039, 3292]. Universal [66, 574, 3536]. Unknown [2713, 2968, 3085, 3228]. unlinkable [3234]. unlinkably [151]. unlocking [2765]. Unmanned [1209, 1386, 1543, 2072, 2141, 2613, 3220, 3294, 3323]. unmanned-aerial [3294]. unreliable [2621, 2926]. unsolicited [1614]. unstructured [421, 495, 706, 752, 1375]. unsupervised [209, 2688, 3009]. until [1031]. untraceable [1810, 2157]. untrusted [2293]. UOWC [3369]. up-to-date [2964, 3091]. Update [451, 957, 1230, 1446, 2333, 2536, 3232]. updates [653]. updating [1227, 2195, 3063]. upgradation [3011]. uplink [773, 973, 1079, 1146, 1836, 2184, 2213]. uplink/downlink [773]. upon [307, 986]. upper [360]. urban [267, 314, 925, 1017, 1280, 1519, 1643, 1830, 2039]. urgent [758]. urinalysis [2264]. URL [379]. URLLC [3322]. URLs [1224]. Usability [125, 1068]. Usage [265, 298, 678, 785, 872, 1669, 3017, 3389]. Use [42, 67, 77, 888, 1006, 1278, 1494, 2091, 2178, 2504, 2772]. User [17, 73, 96, 132, 138, 163, 190, 192, 274, 361, 404, 483, 493, 524, 539, 560, 565, 831, 856, 877, 892, 952, 1000, 1053, 1056, 1067, 1076, 1097, 1196, 1268, 1460, 1595, 1630, 1663, 1676, 1719, 1746, 1813, 1855, 1863, 1899, 1952, 2086, 2192, 2219, 2234, 2271, 2278, 2400, 2426, 2438, 2443, 2461, 2561, 2799, 2867, 2902, 3001, 3005, 3031, 3086, 3185, 3289, 3360, 3372, 3424, 3456]. User-centric [404, 1899, 2219, 2271, 3360]. User-defined [1952]. user-friendly [17]. User-Group [132]. user-perceived [3424]. User-provided [565]. User-side [1460]. user-specific [831]. users [34, 135, 188, 667, 1084, 1849, 1894, 2200, 2731, 2803, 3104, 3473]. userspace [2758]. Using [36, 59, 62, 157, 183, 201, 202, 244, 249, 262, 272, 395, 403, 431, 436, 512, 598, 601, 614, 620, 637, 675, 694, 721, 738, 805, 815, 859, 877, 900, 904, 980, 1029, 1050, 1107, 1117, 1141, 1146, 1183, 1200, 1219, 1272, 1295, 1354, 1397, 1445, 1507, 1601, 1627, 1635, 1751, 1823, 1849, 1868, 1872, 1914, 2032, 2094, 2163, 2200, 2218, 2229, 2231, 2309, 2323, 2431, 2484, 2494, 2577, 2580, 2608, 2634, 2739, 2762, 2781, 2796, 2801, 2819, 2869, 2881, 2968, 3084, 3103, 3179, 3195, 3210, 3307, 3426, 3443, 3478, 3513, 3514, 3529, 3561, 3562]. using [4, 14, 63, 74, 144, 211, 228, 320, 346, 350, 363, 366, 427, 463, 468, 479, 483, 506, 513, 536, 551, 560, 568, 569, 571, 573, 602, 683, 791, 811, 823, 856, 917, 958, 976, 988, 1038, 1138, 1168, 1197, 1334, 1362, 1406, 1420, 1450, 1469, 1475, 1499, 1527, 1645, 1648, 1649, 1663, 1668, 1706, 1719, 1761, 1792, 1810, 1825, 1833, 1841, 1844, 1923, 1937, 1939, 2004, 2011, 2033, 2062, 2085, 2090, 2272, 2550, 2582, 2624, 2650, 2668, 2761, 2800, 2838, 2848, 2870, 2899, 2977, 2982, 2997, 3062, 3102, 3104, 3116, 3192, 3258, 3267, 3359, 3383, 3421, 3422, 3429, 3437, 3463, 3470, 3564]. using [22, 139, 192, 235, 243, 248, 270, 311, 377, 394, 407, 516, 578, 629, 633, 746, 795, 806, 857, 952, 1022, 1113, 1136, 1169, 1180, 1271, 1284, 1352, 1381, 1417, 1638, 1658, 1672, 1705, 1893, 1911, 2072, 2143, 2210, 2238, 2304, 2305, 2443, 2449, 2459, 2493, 2562, 2626, 2682, 2711, 2719, 2830, 2862, 2967, 3000, 3048, 3055, 3094, 3149, 3167, 3191, 3226, 3309, 3361, 3416, 3423, 3547]. USRP [865]. USRP-platform [865]. utilities [2689]. Utility [279, 773, 923, 1138, 1445, 1480, 2243, 3410]. Utility-aware [1480]. utility-based [279, 773]. Utility-optimized [2243]. Utilization [1106, 1127, 1568, 2140, 2435, 2713, 2751, 3193, 3441]. utilizations [1836]. Utilizing [46, 2871]. UWSN [1846]. UWSNs [3391].

V2N [3095]. vaccination [3382]. valet [2891, 3385]. validation [1554, 2082, 3322]. validations [1865]. Value [16, 235, 346, 889, 1705]. VANET [1017, 1026, 3561]. VANETs [734, 903, 957, 1021, 1024, 1049, 1054, 1055, 1080, 1505, 2336, 3279, 3376]. VAP [2334]. variable [1308]. variable-length [1308]. variables [3155]. variant [49]. variants [1273, 2019, 3517]. variation [394]. various [1246, 2760, 3014]. vault [504]. VBR [75, 528]. VCR [61]. VDFChain [3451]. VE [2861]. VE-CoAP [2861]. VEACON [913]. vEB [519]. vEB-tree-based [519]. vector [367, 1257, 2136, 3155]. vectors [2113]. Vegas [1436]. vehicle [554, 846, 1048, 1050, 1052, 1209, 1265, 1527, 1946, 2033, 2216, 2331, 2882, 3220, 3296, 3495]. vehicle-aided [1209]. vehicle-mounted [1946]. Vehicle-to-Roadside [1048]. vehicle-to-vehicle [1050, 3495]. vehicles [903, 1386, 1745, 2072, 2387, 2475, 2487, 2734, 3294]. Vehicular [572, 736, 810, 828, 875, 913, 991, 1016, 1019, 1020, 1026, 1047, 1048, 1124, 1153, 1173, 1203, 1205, 1280, 1329, 1642, 1944, 2129, 2272, 2336, 2550, 2911, 2920, 3040, 3539]. vein [510]. velocity [1455]. Velox [2290]. Ventricular [2291]. Venue [2150]. verifiability [1879, 2078]. Verifiable [2138, 2260, 2390, 3043, 3213, 3451]. verifiably [486]. Verification [516, 1032, 1087, 1306, 1826, 2327, 3190]. verify [2871]. Verifying [652]. versatile [2636]. versioning [1133]. versus [604]. vertex [2405]. vertical [672, 731, 1034]. VGITS [343]. via [47, 130, 674, 698, 923, 955, 989, 997, 1357, 1365, 1366, 1458, 1664, 1878, 1894, 2034, 2346, 2415, 2508, 2649, 2723, 2773, 2867, 2971, 3006, 3134, 3181, 3208, 3298, 3424, 3451, 3472, 3526, 3565]. VIBE [851]. Video [56, 73, 75, 112, 139, 181, 220, 261, 267, 269, 298, 431, 516, 519, 566, 626, 658, 693, 701, 709, 871, 923, 937, 951, 962, 1098, 1106, 1118, 1159, 1281, 1292, 1299, 1347, 1525, 1538, 1543, 1722, 1909, 1967, 2048, 2074, 2179, 2256, 2284, 2285, 2389, 2452, 2696, 2718, 2778, 2795, 2836, 2861, 2885, 3032, 3084, 3129, 3192, 3196, 3262, 3276, 3488, 3501, 3509]. video-on-demand [139, 220, 1722, 2885]. videoconferencing [148, 377]. videos [152, 2314]. view [208, 755, 1179, 1317, 2415, 2724]. viewer [236]. viewing [2896]. viewpoint [1347]. viewpoints [208]. viral [3015]. Virtual [88, 108, 113, 115, 116, 137, 163, 185, 190, 314, 366, 390, 466, 481, 577, 603, 655, 656, 705, 746, 999, 1002, 1003, 1072, 1076, 1078, 1091, 1154, 1179, 1180, 1230, 1250, 1307, 1406, 1415, 1439, 1479, 1546, 1636, 1656, 1667, 1711, 1717, 1804, 1832, 1842, 1871, 1934, 1997, 2013, 2053, 2207, 2210, 2242, 2313, 2334, 2361, 2542, 2592, 2598, 2600, 2621, 2656, 2666, 2729, 2867, 2894, 2932, 2951, 3139, 3162, 3186, 3193, 3249, 3250, 3259, 3297, 3346, 3353, 3401, 3412, 3482, 3499, 3522, 3547]. virtual-currency-enabled [3259]. virtual-machine [3162]. Virtualization [659, 1089, 1322, 1403, 1705, 2812, 3373, 3386]. Virtualized [654, 1483, 1690, 1713, 2267, 2609, 2610, 3316]. Virtualizing [3207]. virus [3303]. Visible [3059]. Vision [348, 2875, 3000, 3450]. vision-based [348]. visiting [1853]. Visual [45, 243, 308, 516, 968, 1495, 1601, 1617, 1750, 1781, 2356, 2626, 2871, 3346]. Visual-based [516]. visual/virtual [3346]. visualization [40, 57, 120, 567, 2357]. Visualizing [171, 364, 1463, 2554]. visually [571, 884, 1219]. Vital [2461]. Vitruvius [1265]. Vivaldi [583]. VM [1927, 2042, 2290]. VNE [1092, 1706, 3412]. VNET6 [1078]. VNF [2592, 3069, 3316, 3549]. VNF-based [3069]. VoD [885, 2720]. voice [71, 144, 540, 582, 1229, 2389]. voiced [55]. VoIP [698, 886, 1013, 1123, 1405, 1544, 1860, 2295]. volatility [1548]. voltage [1558]. VoLTE [2389]. Volume [51, 72, 91, 117, 142, 182, 215, 2676]. Volumes [2102]. voluntary [251]. voluntary/incentive [251]. voluntary/incentive-based [251]. Volunteer [1175]. Voronoi [1162, 3103]. Voronoi-based [1162]. voter [425]. voter-controlled [425]. Voting [318, 965, 1189, 1253, 3034, 3245, 3267]. VP9 [2836]. VPN [2699]. VQ [242]. VQ-based [242]. VREFL [3213]. vulnerabilities [2205, 2382, 2797, 2942]. vulnerability [385, 2942, 3066, 3386]. vulnerability-based [3386]. VWCA [572].

W [855]. W-Grid [855]. waiting [559]. wake [2624, 3018]. wake-up [2624, 3018]. wakeup [3061]. walks [1721]. walkthrough [2554]. WAN [145]. WAP [180]. WAP-based [180]. warehouse [1811]. warning [792]. WaspMote [1320]. waste [3173]. Watchdog [2199]. water [2146]. watermark [1086]. watermarking [136, 242, 367, 750, 1801]. watermarks [2908]. wave [2883]. wavelength [328, 798, 1429]. wavelet [3155]. way [402, 787, 1925, 2385, 2960]. WBAN [1345, 1581]. WBLE [93]. WBNs [2092]. WBT [194]. WCMA [2001]. WDM [256, 264, 275, 277, 278, 328, 344, 570, 606, 798, 1387, 1429, 1523]. wear [1620]. wear-leveling [1620]. wearable [2801]. Web [36, 47, 56, 64, 67, 69, 77, 80, 93, 95, 124, 125, 149, 157, 176, 202, 207, 209, 210, 225, 266, 268, 270, 271, 333, 334, 336, 358, 361, 364, 385, 396, 400, 411, 457, 542, 580, 683, 712, 716, 746, 814, 815, 879, 929, 954, 1032, 1128, 1186, 1196, 1214, 1271, 1290, 1398, 1478, 1552, 1577, 1658, 1683, 1884, 1915, 2052, 2099, 2115, 2175, 2205, 2221, 2278, 2359, 2383, 2502, 2553, 2688, 2896, 3093, 3146, 3290, 3454, 3566]. web-based [69, 80, 93, 361, 364, 385, 954]. web-centric [1186]. Web-flexible [47]. Web-OEM [95]. web-service [1884]. Web-Telecom [542]. weblogs [3566]. website [3277]. Weibull [2531]. Weight [889, 1603, 1894, 2323, 2533, 2766]. weight-value [889]. weighted [515, 782, 1503, 2221, 2395, 3550]. WeLearn.Mobile [236]. Wersync [2896]. wheel [2600]. wheelchairs [1601]. which [228]. while [1348]. white [2103]. Who [3104]. Wi [997, 1965, 2041, 2552, 2637, 2781, 2964, 3030, 3547]. Wi-Fi [997, 1965, 2041, 2552, 2637, 2781, 2964, 3030, 3547]. Wi-Run [2552]. Wide [36, 56, 69, 77, 111, 124, 616, 848, 1201, 1557, 2595, 2620, 2892, 3487]. wide-area [111, 1557]. wideband [730]. Wiener [446]. WiFi [1020, 1098, 2229, 2750]. WiFi-based [1020, 1098]. wildcard [1327]. wildfire [2568]. WiMAX [473, 624, 673, 872, 973, 1079, 1146]. WiMesh [1618]. wind [1868]. window [882, 1484, 2201, 2559, 3350]. Windows [84, 178, 484, 990, 3370]. Windows-based [178]. Winnowing [751]. wired [1107, 1304, 1498, 1724, 1823, 2032]. wired-wireless [2032]. wired/wireless [1724]. Wireless [135, 233, 478, 491, 588, 596, 601, 669, 681, 691, 718, 725, 732, 740, 776, 778, 781, 804, 813, 836, 845, 846, 854, 887, 892, 928, 932, 942, 951, 964, 1043, 1095, 1110, 1112, 1141, 1167, 1199, 1222, 1240, 1243, 1246, 1256, 1300, 1311, 1328, 1333, 1335, 1339, 1346, 1385, 1399, 1410, 1518, 1519, 1521, 1524, 1587, 1611, 1618, 1689, 1710, 1724, 1800, 1846, 1850, 1868, 1890, 1909, 1912, 1974, 1979, 1985, 2001, 2032, 2100, 2120, 2142, 2157, 2178, 2203, 2231, 2244, 2253, 2298, 2319, 2324, 2325, 2347, 2405, 2424, 2450, 2460, 2633, 2744, 3000, 3010, 3276, 3317, 3546]. wireless [201, 350, 371, 403, 480, 528, 556, 585, 592, 594, 595, 597, 598, 611, 679, 682, 708, 720, 723, 728, 735, 791, 792, 816, 818, 916, 944, 985, 992, 1034, 1038, 1039, 1094, 1097, 1120, 1132, 1137, 1142, 1143, 1147, 1165, 1181, 1232, 1247, 1282, 1326, 1330, 1337, 1340, 1363, 1368, 1422, 1423, 1427, 1458, 1489, 1495, 1528, 1532, 1540, 1576, 1588, 1612, 1631, 1645, 1699, 1756, 1758, 1779, 1810, 1820, 1844, 1884, 1891, 1896, 1903, 1918, 1939, 2020, 2040, 2063, 2139, 2258, 2292, 2338, 2345, 2348, 2401, 2438, 2446, 2547, 2557, 2577, 2600, 2615, 2761, 3027, 3098, 3161, 3175, 3258, 3280, 3421, 3501, 3511]. wireless [180, 258, 375, 428, 449, 450, 571, 578, 590, 600, 612, 697, 715, 722, 726, 748, 749, 773, 797, 855, 864, 882, 895, 901, 917, 938, 971, 984, 1002, 1005, 1042, 1057, 1090, 1096, 1111, 1115, 1150, 1171, 1215, 1244, 1255, 1259, 1289, 1315, 1320, 1336, 1347, 1348, 1408, 1451, 1454, 1470, 1497, 1501, 1580, 1589, 1597, 1604, 1640, 1675, 1680, 1707, 1732, 1753, 1777, 1845, 1848, 1859, 1906, 1908, 1981, 1989, 2010, 2021, 2031, 2051, 2075, 2170, 2176, 2201, 2300, 2353, 2394, 2397, 2439, 2513, 2565, 2683, 2698, 2949, 2988, 2990, 3008, 3071, 3108, 3119, 3302, 3423, 3428, 3470, 3497, 3530]. wireless [219, 414, 430, 522, 529, 561, 591, 676, 685, 702, 710, 719, 741, 766, 790, 793, 799, 802, 822, 837, 881, 891, 914, 922, 945, 959, 960, 986, 1022, 1023, 1027, 1044, 1108, 1148, 1162, 1166, 1187, 1237, 1274, 1342, 1353, 1365, 1379, 1419, 1432, 1447, 1448, 1514, 1585, 1586, 1600, 1601, 1620, 1641, 1647, 1735, 1769, 1794, 1866, 1893, 1898, 1914, 1940, 1994, 2013, 2023, 2026, 2066, 2073, 2090, 2095, 2127, 2149, 2230, 2312, 2334, 2368, 2374, 2464, 2468, 2478, 2673, 2718, 2740, 3014, 3117, 3189, 3235, 3291, 3344, 3400, 3450]. wireless-networking [1043]. wise [1617, 2199]. Wise-PRoPHET [2199]. within [206, 1671, 3005]. within-perimeter [1671]. without [741, 812, 1859, 2730, 3475]. witness [1616]. WLAN [689, 809, 989, 1013, 2590]. WLANs [234, 906, 1349, 2027, 2402]. WMN [2371].

WMNs [593, 1654]. WMSN [2284]. wolf [3390]. words [703]. work [757]. worker [2432]. Workflow [406, 753, 1148, 1382, 1388, 1665, 1803, 2257, 2651, 3162, 3216]. workflow-enabled [406]. Workflow-supported [1803]. workflows [389, 2621]. workforce [758]. workload [929, 1465, 1828, 1861, 2160, 2359, 2948, 3181, 3488]. workload-adaptive [3488]. workload-aware [2948]. workloads [327, 1307, 1483, 3197]. works [2891]. World [36, 56, 69, 77, 124, 1404, 2308, 2460, 2646, 2684, 3225]. World-Wide [124]. worlds [113]. Worm [2984]. wormhole [258, 983, 2090, 3551]. worms [706, 1058]. WPAN [2975]. WPANs [395]. WPC [1465]. WRANs [2103]. write [2130]. WRSNs [2679]. WSN [863, 1248, 1638, 1670, 1733, 1864, 2501, 2566, 2968, 3330]. WSN-Assisted [2501]. WSN-based [1733, 2566]. WSNs [489, 859, 898, 955, 1151, 1161, 1257, 1331, 1412, 1461, 1535, 1616, 1833, 1917, 2009, 2180, 2181, 2322, 2585, 2586, 2602, 2939, 3062, 3138, 3263]. WWW’ [25, 40, 78, 81, 97, 118, 119]. WWW-based [97].

xDSL [391]. Xhaul [2122]. XML [138, 177, 657, 1004, 1122, 1699]. XML-DoS [657]. XML-related [1004]. XOR [321]. XORshift [1117]. XSS [2294].

YJNCA [2591, 2721].

Z [2742]. Zero [564, 1090, 1869, 2009, 2159, 3150, 3274, 3296, 3319, 3569]. Zero-Degree [1090]. zero-effort [1869]. zero-energy [3569]. Zero-queue [2159]. ZigBee [641, 997, 1336]. zk [3345]. zk-rollup [3345]. zone [603, 2761, 2782]. zones [1449, 2781]. ZSM [3150].

References

Maurer:1996:E

[1] Hermann Maurer. Editorial. Journal of Network and Computer Applications, 19(1):1, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

      science/article/pii/S1084804596900016.

Jain:1996:E

[2] L. C. Jain. Editorial. Journal of Network and Computer Applications, 19(1):3, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

      science/article/pii/S1084804596900028.

Dawid:1996:DGA

[3] Herbert Dawid and Kurt Hornik. The dynamics of genetic algorithms in interactive environments. Journal of Network and Computer Applications, 19(1):5–19, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480459690003X.

Karr:1996:APC

[4] Charles L. Karr. Adaptive process control using biological paradigms. Journal of Network and Computer Applications, 19(1):21–44, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804596900041.

Kuwata:1996:CSD

[5] Kaihei Kuwata, Yuji Kajitani, Masahide Watanabe, and Ryu Katayama. Chaos simulator as a developing tool and a research environment for applications of chaos engineering. Journal of Network and Computer Applications, 19(1):45–66, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804596900053.

Bose:1996:KBA

[6] Ranjit Bose and Vijayan Sugumaran. Knowledge-based approach to domain modeling: organizational process modeling application. Journal of Network and Computer Applications, 19(1):67–89, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804596900065.

Mital:1996:LPO

[7] Dinesh P. Mital, Eam Khwang Teoh, and P. N. Suganthan. Learning parameters for object recognition by the self-organizing Hopfield network. Journal of Network and Computer Applications, 19(1):91– 108, January 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900077.

Jain:1996:GEI

[8] L. C. Jain. Guest editorial: Intelligent systems: design and applications. Part 2. Journal of Network and Computer Applications, 19(2):

109, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900089.

A:1996:DCN

[9] David V. Sánchez A. On the design of a class of neural networks. Journal of Network and Computer Applications, 19(2):111– 118, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900090.

Civelek-Alpaslan:1996:TNN

[10] Ferda N. Civelek-Alpaslan and Kathleen M. Swigger. A temporal neural network model for constructing connectionist expert system knowledge bases. Journal of Network and Computer Applications, 19(2):119– 133, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900107.

Kottathra:1996:NMO

[11] Krishna Kottathra and Yianni Attikiouzel. A novel multicriteria optimization algorithm for the structure determination of multilayer feedforward neural networks. Journal of Network and Computer Applications, 19(2):135–147, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900119.

Ashiru:1996:CGB

[12] Ibikunle Ashiru, Chris Czarnecki, and Tom Routen. Characteristics of a genetic based approach to path planning for mobile robots. Journal of Network and Computer Applications, 19(2):149–169, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900120.

Cedeno:1996:GAA

[13] Walter Cedeño and Rao V. Vemuri. Genetic algorithms in aquifer management. Journal of Network and Computer Applications, 19(2):171– 187, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900132.

Hunt:1996:LUA

[14] John E. Hunt and Denise E. Cooke. Learning using an artificial immune system. Journal of Network and Computer Applications, 19(2): 189–212, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900144.

Medsker:1996:MAH

[15] Larry R. Medsker. Microcomputer applications of hybrid intelligent systems. Journal of Network and Computer Applications, 19(2):213– 234, April 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900156.

Pan:1996:SVD

[16] Yi Pan and Mounir Hamdi. Singular value decomposition on processor arrays with a pipelined bus system. Journal of Network and Computer Applications, 19(3):235–248, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804596900168.

Kuhlman:1996:DUF

[17] J. M. Kuhlman and M. J. Lyell. Development of a user-friendly software planning tool to assess effects of biofilm activity on contaminant reduction in streams. Journal of Network and Computer Applications, 19(3):249–264, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480459690017X.

Tan:1996:AGI

[18] Sun-Teck Tan. Architecture of a generic instructional planner. Journal of Network and Computer Applications, 19(3):265–274, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900181.

Bourbakis:1996:HDL

[19] N. Bourbakis, N. Pereira, and S. Mertoguno. Hardware design of a letter-driven OCR and document processing system. Journal of Network and Computer Applications, 19(3):275–294, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900193.

Wang:1996:MBP

[20] Jian Tao Wang and Dinesh P. Mital. A microcomputer-based prototype for ECG paper record conversion. Journal of Network and Computer Applications, 19(3):295–307, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480459690020X.

Rajaram:1996:PMC

[21] Gokul Rajaram and V. Rajaraman. A probabilistic method for calculating hit ratios in direct mapped caches. Journal of Network and Computer Applications, 19(3):309–319, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804596900211.

Padmanabhan:1996:BMC

[22] K. Padmanabhan, D. Nedumaran, and S. Ananthi. Bedside monitor for combined ECG and ultrasound Doppler-echo observations using a microprocessor. Journal of Network and Computer Applications, 19(3):321–334, July 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900223.

Anonymous:1996:I

[23] Anonymous. Index. Journal of Network and Computer Applications, 19 (4):i–vi, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900302.

ODonoghue:1996:OMF

[24] P. G. O’Donoghue and M. H. Murphy. Object modelling and formal specification during real-time system development. Journal of Network and Computer Applications, 19(4):335–352, October 1996. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900235.

Nkambou:1996:IWR

[25] Roger Nkambou and Gilles Gauthier. Integrating WWW resources in an Intelligent Tutoring System. Journal of Network and Computer Applications, 19(4):353–366, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900247.

Jih:1996:ILP

[26] Hueyching Janice Jih. The impact of learners’ pathways on learning performance in multimedia Computer Aided Learning. Journal of Network and Computer Applications, 19(4):367–380, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900259.

Helm:1996:MFI

[27] Daniel J. Helm, Raymond J. D’Amore, and Puck-Fai Yan. MIDS: a framework for information organization and discovery. Journal of Network and Computer Applications, 19(4):381–394, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900260.

Maly:1996:IRI

[28] K. Maly, C. Wild, C. M. Overstreet, H. Abdel-Wahab, A. Gupta, A. Youssef, E. Stoica, R. Talla, and A. Prabhu. Interactive remote instruction: lessons learned. Journal of Network and Computer Applications, 19(4):395–404, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900272.

Radhakrishnan:1996:MBP

[29] D. Radhakrishnan. A microcontroller based pacemaker tester. Journal of Network and Computer Applications, 19(4):405–413, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804596900284.

Singer:1996:EAP

[30] Benjamin Singer and George Saon. An efficient algorithm for parallel integer multiplication. Journal of Network and Computer Applications, 19(4):415–418, October 1996. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900296.

Anonymous:1997:GES

[31] Anonymous. Guest editorial: Special issue of JNCA on the digital libraries. Journal of Network and Computer Applications, 20(1):1– 2, January 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900351.

Lim:1997:IQE

[32] Ee-Peng Lim and Ying Lu. Integrated queries to existing bibliographic and structured databases. Journal of Network and Computer Applications, 20(1):3–24, January 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900363.

Sugimoto:1997:ESS

[33] Shigeo Sugimoto, Akira Maeda, Tetsuo Sakaguchi, Koichi Tabata, and Takehisa Fujita. Experimental studies on software tools to enhance accessibility to information in digital libraries. Journal of Network and Computer Applications, 20(1):25–44, January 1997. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804596900375.

Sanchez:1997:ASU

[34] Alfredo J. Sánchez and John J. Leggett. Agent services for users of digital libraries. Journal of Network and Computer Applications, 20(1): 45–58, January 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900387.

Tochtermann:1997:DDD

[35] Klaus Tochtermann, Thomas Alders, and Andreas Seifert. DogitaLS1: the Dortmund digital library system. Journal of Network and Computer Applications, 20(1):59–73, January 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804596900399.

Alloway:1997:CIL

[36] Gene Alloway, Nathan Bos, Kathleen Hamel, Tracy Hammerman, Elisabeth Klann, Joseph Krajcik, David Lyons, Terry Madden, Jon Margerum-Leys, James Reed, Nancy Scala, Elliot Soloway, Ioanna Vekiri, and Raven Wallace. Creating an inquiry-learning environment using the World Wide Web. Journal of Network and Computer Applications, 20(1):75–85, January 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804596900405.

Schnase:1997:FND

[37] John L. Schnase, Deborah L. Kama, Kay L. Tomlinson, José Alfredo Sánchez, Edward L. Cunnius, and Nancy R. Morin. The Flora of North America digital library: a case study in biodiversity database publishing. Journal of Network and Computer Applications, 20(1):87– 103, January 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804596900417.

Tripp:1997:DAI

[38] Gerald Tripp. On the design of an ATM interface with facilities for traffic monitoring and generation. Journal of Network and Computer Applications, 20(2):105–121, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804596900338.

Losleben:1997:HEA

[39] Paul Losleben, Alice S. Tull, and Erming Zhou. Help-Exchange: an arbitrated system for a help network. Journal of Network and Computer Applications, 20(2):123–134, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480459690034X.

Fowler:1997:VWS

[40] Richard H. Fowler, Wendy A. L. Fowler, and Jorge L. Williams. 3D visualization of WWW semantic content for browsing and query formulation. Journal of Network and Computer Applications, 20(2):135– 149, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804597900424.

Bielenberg:1997:ESM

[41] Daniel R. Bielenberg and Ted Carpenter-Smith. Efficacy of story in multimedia training. Journal of Network and Computer Applications, 20(2):151–159, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804597900436.

Lee:1997:PYU

[42] B. S. Lee, T. K. Y. Chan, and P. C. Leong. A ‘pay-as-you-use’ controller for networked laser printers. Journal of Network and Computer Applications, 20(2):161–169, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804597900473.

Hui:1997:DFM

[43] S. C. Hui, K. Y. Chan, and G. Y. Qian. Distributed fax message processing system. Journal of Network and Computer Applications, 20(2):171–190, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480459790045X.

Khoo:1997:PFB

[44] L. P. Khoo and M. Y. Teo. A prototype fuzzy-based classifier system for trajectory planning. Journal of Network and Computer Applications, 20(2):191–202, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804597900461.

Chan:1997:VQL

[45] Hock C. Chan. Visual query languages for entity relationship model databases. Journal of Network and Computer Applications, 20(2):203– 221, April 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804597900448.

Ghosh:1997:AAE

[46] Sumit Ghosh. An asynchronous approach to efficient execution of programs on adaptive architectures utilizing FPGAs. Journal of Network and Computer Applications, 20(3):223–252, July 1997. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804597900527.

Borkowski:1997:STL

[47] Ellen Yu Borkowski, David Henry, Lida L. Larsen, and Deborah Mateik. Supporting teaching and learning via the Web: transforming hard-copy linear mindsets into Web-flexible creative thinking. Journal of Network and Computer Applications, 20(3):253–265, July 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804597900552.

Zhou:1997:ENR

[48] Jianying Zhou and Dieter Gollmann. Evidence and non-repudiation. Journal of Network and Computer Applications, 20(3):267–281, July 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804597900564.

Tao:1997:VPK

[49] Renji Tao and Shihua Chen. A variant of the public key cryptosystem FAPKC3. Journal of Network and Computer Applications, 20(3): 283–303, July 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804597900576.

Schumacher:1997:FFN

[50] H. J. Schumacher and Sumit Ghosh. A fundamental framework for network security. Journal of Network and Computer Applications, 20(3): 305–322, July 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804597900588.

Anonymous:1997:ICV

[51] Anonymous. Index and Contents to Volume 20. Journal of Network and Computer Applications, 20(4):i–x, October 1997. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804597900643.

Bock:1997:ICP

[52] Holger Bock, Wolfgang Mayerwieser, Karl C. Posch, Reinhard Posch, and Volker Schindler. An integrated co-processor architecture for a smartcard. Journal of Network and Computer Applications, 20(4):323– 337, October 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804597900485.

Esler:1997:NIB

[53] Sandra L. Esler and Michael L. Nelson. NASA indexing benchmarks: evaluating text search engines. Journal of Network and Computer Applications, 20(4):339–353, October 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804597900497.

Iglewski:1997:CFS

[54] Michal Iglewski and Tomasz Müldner. Comparison of formal specification methods and object-oriented paradigms. Journal of Network and Computer Applications, 20(4):355–377, October 1997. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804597900503.

George:1997:RTP

[55] Don E. George and E. Salari. Real-time pitch extraction of voiced speech. Journal of Network and Computer Applications, 20(4):379– 387, October 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804597900515.

Foo:1997:DVM

[56] Schubert Foo and Siu Cheung Hui. Delivery of video mail on the World Wide Web. Journal of Network and Computer Applications, 20(4):389–403, October 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804597900539.

Leisch:1997:ADV

[57] Bernhard Leisch and Peter Schuhmayer. An architecture for distributed visualization of technical processes. Journal of Network and Computer Applications, 20(4):405–420, October 1997. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804597900540.

Lu:1998:ECS

[58] Guojun Lu and Chester Kang. An efficient communication scheme for media on-demand services with hard QoS guarantees. Journal of Network and Computer Applications, 21(1):1–15, January 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804598900595.

Blenkhorn:1998:UST

[59] P. Blenkhorn and D. G. Evans. Using speech and touch to enable blind people to access schematic diagrams. Journal of Network and Computer Applications, 21(1):17–29, January 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804598900601.

Mir-Fakhraei:1998:EMA

[60] Nader Mir-Fakhraei. An efficient multicast approach in an ATM switching network for multimedia applications. Journal of Network and Computer Applications, 21(1):31–39, January 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804598900613.

Ang:1998:DVS

[61] Ee-Luang Ang, Syin Chan, and Bu-Sung Lee. Deployment of VCR services on a computer network. Journal of Network and Computer Applications, 21(1):41–52, January 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804598900625.

Ratcliffe:1998:UMT

[62] Mark Ratcliffe and Tim Paul Davies. Using MBone technology to support a framework for a remote advisory service. Journal of Network and Computer Applications, 21(1):53–68, January 1998. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804598900637.

Khan:1998:AIA

[63] Imran Khan and Howard C. Card. Adaptive information agents using competitive learning. Journal of Network and Computer Applications, 21(2):69–89, April 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804598900650.

Vlajic:1998:CWP

[64] N. Vlajic and H. C. Card. Categorizing Web pages on the subject of neural networks. Journal of Network and Computer Applications, 21(2):91–105, April 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804598900662.

Zeadally:1998:EMA

[65] Sherali Zeadally and Weiyou Cui. Experiences with multimedia applications over native ATM. Journal of Network and Computer Applications, 21(2):107–123, April 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804598900674.

Tan:1998:IRU

[66] E. C. Tan, A. Wahab, and H. W. Yong. Interfacing reprogrammable universal remote controller with multimedia PC. Journal of Network and Computer Applications, 21(2):125–134, April 1998. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804598900686.

Charlton:1998:UMW

[67] C. C. Charlton, P. H. Leng, R. K. Lloyd, and R. Turnbull. The use of macroprocessing for Web site standardization. Journal of Network and Computer Applications, 21(2):135–148, April 1998. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804598900698.

Parr:1998:OMT

[68] Gerard Parr and Kevin Curran. Optimal multimedia transport on the Internet. Journal of Network and Computer Applications, 21(2):149– 161, April 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804598900704.

Wong:1998:FWW

[69] Johnny S. K. Wong, Rishi Nayar, and Armin R. Mikler. A framework for a World Wide Web-based Data Mining system. Journal of Network and Computer Applications, 21(3):163–185, July 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804598900716.

Saxena:1998:PCT

[70] P. C. Saxena and Sangita Gupta. A performance comparison of tokenand tree-based mutual exclusion algorithms on Chordal Rings. Journal of Network and Computer Applications, 21(3):187–201, July 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804598900728.

Chin:1998:EQI

[71] K. V. Chin, S. C. Hui, and S. Foo. Enhancing the quality of Internet voice communication for Internet telephony systems. Journal of Network and Computer Applications, 21(3):203–218, July 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480459890073X.

Anonymous:1998:CIV

[72] Anonymous. Contents and Index to Volume 21. Journal of Network and Computer Applications, 21(4):i–viii, October 1998. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599900826.

Prihavec:1998:UIV

[73] Bor Prihavec and Franc Solina. User interface for video observation over the Internet. Journal of Network and Computer Applications, 21(4):219–237, October 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900747.

Mitrou:1998:TEU

[74] Nikolas M. Mitrou and Konstantine Kavidopoulos. Traffic engineering using a class of M/G/1 models. Journal of Network and Computer Applications, 21(4):239–271, October 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900759.

Rikli:1998:EMV

[75] Nasser-Eddine Rikli. An effective model for VBR video with application to an ATM multiplexer. Journal of Network and Computer Applications, 21(4):273–285, October 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900760.

Moreno:1998:PCC

[76] L. Moreno, R. M. Aguilar, C. A. Martin, J. D. Piñeiro, J. I. Estévez, J. L. Sánchez, J. F. Sigut, and V. I. Jiménez. Patient-centered computer simulation in hospital management. Journal of Network and Computer Applications, 21(4):287–310, October 1998. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804599900772.

Chen:1999:FDS

[77] Wei-Chou Chen, Tzung-Pei Hong, and Rong Jeng. A framework of decision support systems for use on the World Wide Web. Journal of Network and Computer Applications, 22(1):1–17, January 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599900784.

Helic:1999:IHC

[78] Denis Helic, Hermann Maurer, and Nick Scherbakov. Introducing hypermedia composites to WWW. Journal of Network and Computer Applications, 22(1):19–32, January 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900796.

Chawla:1999:YNS

[79] Hamesh Chawla, Geoff Dillon, and Riccardo Bettati. HYDRANET: Network support for scaling of large-scale services. Journal of Network and Computer Applications, 22(1):33–50, January 1999. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599900802.

Hasebrook:1999:WBT

[80] Joachim Hasebrook. Web-based training, performance and controlling. Journal of Network and Computer Applications, 22(1):51–64, January 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599900814.

Ghaoui:1999:SIS

[81] Claude Ghaoui and M. J. Taylor. Special issue on: ‘support for Open and Distance Learning on the WWW’. Journal of Network and Computer Applications, 22(1):65–68, January 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480459990084X.

Klapuri:1999:SLR

[82] Harri Klapuri, Jarmo Takala, and Jukka Saarinen. Safety, liveness and real-time in embedded system design. Journal of Network and Computer Applications, 22(2):69–89, April 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804599900838.

Kumar:1999:IAQ

[83] M. Kumar, S. Venkatesh, K.-Y. Lim, and H. Santoso. Information access and QoS issues in a mobile computing environment. Journal of Network and Computer Applications, 22(2):91–118, April 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599900851.

Hanner:1999:MWN

[84] K. Hanner and R. Hörmanseder. Managing Windows NT^{(}R) file system permissions — a security tool to master the complexity of Microsoft Windows NT^{(}R) file system permissions. Journal of Network and Computer Applications, 22(2):119–131, April 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804599900863.

Steeples:1999:EPL

[85] Christine Steeples and Peter Goodyear. Enabling professional learning in distributed communities of practice: Descriptors for multimedia objects. Journal of Network and Computer Applications, 22(2):133– 145, April 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599900875.

Wiil:1999:CPS

[86] Uffe Kock Wiil, Kenneth M. Anderson, and Michael Bieber. Call for Papers: Special issue on: Hypermedia extensibility mechanisms and scripting languages. Journal of Network and Computer Applications, 22(2):147–148, April 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900929.

Gilbert:1999:AIS

[87] J. E. Gilbert and C. Y. Han. Adapting instruction in search of ‘a significant difference’. Journal of Network and Computer Applications, 22(3):149–160, July 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900887.

Slator:1999:VEE

[88] Brian M. Slator, Paul Juell, Philip E. McClean, Bernhardt Saini-Eidukat, Donald P. Schwert, Alan R. White, and Curt Hill. Virtual environments for education. Journal of Network and Computer Applications, 22(3):161–174, July 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900899.

Lu:1999:DIM

[89] Guojun Lu. Design issues of multimedia information indexing and retrieval systems. Journal of Network and Computer Applications, 22(3): 175–198, July 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599900905.

Ahn:1999:TRB

[90] Gail J. Ahn and Ravi Sandhu. Towards role-based administration in network information services. Journal of Network and Computer Applications, 22(3):199–213, July 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900917.

Anonymous:1999:ICV

[91] Anonymous. Index and Contents to Volume 22. Journal of Network and Computer Applications, 22(4):i–viii, October 1999. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599901017.

Du:1999:MNT

[92] Hongbo Du and Laurent Wery. Micro: a normalization tool for relational database designers. Journal of Network and Computer Applications, 22(4):215–232, October 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804599900966.

Mioduser:1999:WBL

[93] David Mioduser, Rafi Nachmias, Avigail Oren, and Orly Lahav. Webbased learning environments (WBLE): Current implementation and evolving trends. Journal of Network and Computer Applications, 22(4): 233–247, October 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599900930.

Carro:1999:DGA

[94] Rosa María Carro, Estrella Pulido, and Pilar Rodríguez. Dynamic generation of adaptive Internet-based courses. Journal of Network and Computer Applications, 22(4):249–257, October 1999. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599900942.

Iocchi:1999:WOA

[95] Luca Iocchi. The Web-OEM approach to Web information extraction. Journal of Network and Computer Applications, 22(4):259–269, October 1999. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599900954.

Hurst:2000:UIT

[96] Wolfgang Hürst. User interfaces for telepresentations — Input devices, interaction concepts and design issues. Journal of Network and Computer Applications, 23(1):1–15, January 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804599900978.

deBoer:2000:ICW

[97] Wim de Boer and Betty Collis. Instructors’ choices for a WWW-based course-support environment. Journal of Network and Computer Applications, 23(1):17–26, January 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480459990098X.

Tu:2000:LLM

[98] Chih-Hsiung Tu. On-line learning migration: from social learning theory to social presence theory in a CMC environment. Journal of Network and Computer Applications, 23(1):27–37, January 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599900991.

Tu:2000:CEF

[99] Chih-Hsiung Tu. Critical examination of factors affecting interaction on CMC. Journal of Network and Computer Applications, 23(1):39– 58, January 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599901005.

Mok:2000:EIN

[100] Chwei-King Mok and Nader F. Mir. An efficient interconnection network for large-scale computer communications applications. Journal of Network and Computer Applications, 23(2):59–75, April 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804599901029.

Berque:2000:DIS

[101] Dave Berque, Andrew Hutcheson, David K. Johnson, Larry Jovanovic, Kyle Moore, Carl Singer, and Kevin M. Slattery. The design of an interface for student note annotation in a networked electronic classroom. Journal of Network and Computer Applications, 23(2):77– 91, April 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804599901066.

Pattinson:2000:SNM

[102] Colin Pattinson. A simulated network management information base. Journal of Network and Computer Applications, 23(2):93–107, April 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804500901048.

Eschelbeck:2000:ASP

[103] Gerhard Eschelbeck. Active Security — a proactive approach for computer security systems. Journal of Network and Computer Applications, 23(2):109–130, April 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804500901036.

Dron:2000:CED

[104] Jon Dron, Richard Mitchell, Phil Siviter, and Chris Boyne. CoFIND — an experiment in N-dimensional collaborative filtering. Journal of Network and Computer Applications, 23(2):131–142, April 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450090105X.

Lee:2000:PLM

[105] Bu-Sung Lee, Chiew-Tong Lau, and Nicholas C. H. Vun. Power LAN MIB for management of intelligent telecommunication equipment. Journal of Network and Computer Applications, 23(2):143–154, April 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500901073.

Anonymous:2000:PEC

[106] Anonymous. Preface by the Editor in Chief. Journal of Network and Computer Applications, 23(3):155, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804500901085.

Dempsey:2000:PDN

[107] Bert J. Dempsey and Micah Beck. Preface to ‘Developing New Networkbased Storage Services for Internet-supported Collaborative Communities’. Journal of Network and Computer Applications, 23(3):157– 159, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500901176.

Tomek:2000:PVE

[108] Ivan Tomek. Preface to ‘Virtual Environments’. Journal of Network and Computer Applications, 23(3):161–162, July 2000. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500901188.

Chuang:2000:DNS

[109] John Chung-I Chuang and Marvin A. Sirbu. Distributed network storage service with quality-of-service guarantees. Journal of Network and Computer Applications, 23(3):163–185, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804500901097.

Chervenak:2000:DGT

[110] Ann Chervenak, Ian Foster, Carl Kesselman, Charles Salisbury, and Steven Tuecke. The data grid: Towards an architecture for the distributed management and analysis of large scientific datasets. Journal of Network and Computer Applications, 23(3):187–200, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500901103.

Dempsey:2000:PAS

[111] Bert J. Dempsey. Performance analysis of a scaleable design for replicating file collections in wide-area networks. Journal of Network and Computer Applications, 23(3):201–218, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804500901115.

Slaughter:2000:OVF

[112] Laura Slaughter, Gary Marchionini, and Gary Geisler. Open video: a framework for a test collection. Journal of Network and Computer Applications, 23(3):219–245, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804500901127.

Cicognani:2000:LDT

[113] Anna Cicognani. Language and design in text-based virtual worlds. Journal of Network and Computer Applications, 23(3):247–274, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804500901139.

Tomek:2000:DIM

[114] Ivan Tomek. The design and implementation of a MOO. Journal of Network and Computer Applications, 23(3):275–289, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500901140.

Diaz:2000:PLV

[115] Alicia Díaz and Alejandro Fernández. A pattern language for virtual environments. Journal of Network and Computer Applications, 23(3): 291–309, July 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500901152.

Choukair:2000:QMC

[116] Zièd Choukair and Damien Retailleau. A QoS model for collaboration through Distributed Virtual Environments. Journal of Network and Computer Applications, 23(3):311–334, July 2000. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500901164.

Anonymous:2000:ICV

[117] Anonymous. Index and contents to volume 23. Journal of Network and Computer Applications, 23(4):iii–viii, October 2000. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500900171.

Maurer:2000:PSI

[118] H. Maurer. Preface for the special issue ‘support for flexible e-Learning on the WWW’. Journal of Network and Computer Applications, 23(4): 335, October 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500900092.

Ghaoui:2000:SFL

[119] Claude Ghaoui and M. J. Taylor. Support for flexible e-Learning on the WWW: a special issue. Journal of Network and Computer Applications, 23(4):337–338, October 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804500900109.

Brand:2000:STT

[120] O. Brand, L. Petrak, D. Sturzebecher, and M. Zitterbart. Supporting Tele-Teaching: visualization aspects. Journal of Network and Computer Applications, 23(4):339–355, October 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804500900110.

Wang:2000:HBS

[121] Weigang Wang, Jörg M. Haake, Jessica Rubart, and Daniel A. Tietze. Hypermedia-based support for cooperative learning of process knowledge. Journal of Network and Computer Applications, 23(4):357– 379, October 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500900122.

Pekkola:2000:MAS

[122] Samuli Pekkola, Mike Robinson, Jonni Korhonen, Saku Hujala, and Tero Toivonen. Multimedia application to support distance learning and other social interactions in real-time. Journal of Network and Computer Applications, 23(4):381–399, October 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804500900134.

Dommel:2000:CFA

[123] Hans-Peter Dommel and J. J. Garcia-Luna-Aceves. A coordination framework and architecture for Internet groupware. Journal of Network and Computer Applications, 23(4):401–427, October 2000. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500900146.

MacGlashan:2000:MOO

[124] Stuart K. MacGlashan and Malcolm J. Taylor. MUSiCAL: an objectoriented hypermedia data model for the reuse of museum information for computer-aided learning and the World-Wide Web. Journal of Network and Computer Applications, 23(4):429–453, October 2000. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500900158.

Ghaoui:2000:DIB

[125] Claude Ghaoui. Document icon bar for the support of authoring, learning and navigation on the Web: Usability issues. Journal of Network and Computer Applications, 23(4):455–475, October 2000. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450090016X.

Wiil:2001:PHE

[126] Uffe Kock Wiil and Kenneth M. Anderson. Preface to ‘hypermedia extensibility mechanisms and scripting languages’. Journal of Network and Computer Applications, 24(1):1–5, January 2001. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450090119X.

Wiil:2001:FCE

[127] Uffe Kock Wiil. A framework for classifying extensibility mechanisms in hypermedia systems. Journal of Network and Computer Applications, 24(1):7–18, January 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804500901206.

Nurnberg:2001:ECB

[128] Peter J. Nürnberg. Extensibility in component-based open hypermedia systems. Journal of Network and Computer Applications, 24(1):19– 38, January 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500901218.

Noll:2001:SPO

[129] John Noll and Walt Scacchi. Specifying process-oriented hypertext for organizational computing. Journal of Network and Computer Applications, 24(1):39–61, January 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450090122X.

Demeyer:2001:EML

[130] Serge Demeyer. Extensibility via a meta-level architecture. Journal of Network and Computer Applications, 24(1):63–74, January 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804500901231.

Anderson:2001:EMC

[131] Kenneth M. Anderson. The extensibility mechanisms of the Chimera open hypermedia system. Journal of Network and Computer Applications, 24(1):75–86, January 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804500901243.

Ahn:2001:CCU

[132] Gail-Joon Ahn and Kwangjo Kim. CONUGA: Constrained user-group assignment. Journal of Network and Computer Applications, 24(2):87– 100, April 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500901255.

Klapuri:2001:IRC

[133] Harri Klapuri, Jarmo Takala, and Jukka Saarinen. Implementing reactive closed-system specifications. Journal of Network and Computer Applications, 24(2):101–123, April 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804500901279.

Li:2001:ASO

[134] Guoyi Li, Ian Li-Jin Thng, and Lawrence Wai Choong Wong. ABR services for overcoming misbehaving sources in a heterogeneous environment. Journal of Network and Computer Applications, 24(2):125– 149, April 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804500901292.

Tan:2001:WMS

[135] David Hua Min Tan, Siu Cheung Hui, and Chiew Tong Lau. Wireless messaging services for mobile users. Journal of Network and Computer Applications, 24(2):151–166, April 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804500901267.

Djurovic:2001:DWF

[136] Igor Djurovic, Srdjan Stankovic, and Ioannis Pitas. Digital watermarking in the fractional Fourier transformation domain. Journal of Network and Computer Applications, 24(2):167–173, April 2001. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804500901280.

Bouras:2001:LTD

[137] C. Bouras, A. Philopoulos, and Th. Tsiatsos. e-Learning through distributed virtual environments. Journal of Network and Computer Applications, 24(3):175–199, July 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804501901316.

Liu:2001:DPC

[138] Duen-Ren Liu, Yuh-Jaan Lin, Chung-Min Chen, and Ya-Wen Huang. Deployment of personalized e-catalogues: an agent-based framework integrated with XML metadata and user models. Journal of Network and Computer Applications, 24(3):201–228, July 2001. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804501901328.

Poon:2001:DSF

[139] W. F. Poon and K. T. Lo. Design of a starvation-free storage system for a multicast video-on-demand system using a double-rate batching policy. Journal of Network and Computer Applications, 24(3):229– 247, July 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804501901304.

Tang:2001:EGS

[140] Ming-Chung Tang, Chun-Nun Chou, Ching-Hui Tang, D. C. Pan, and Wei-Kuan Shih. Exploiting GSM short message service for ubiquitous accessing. Journal of Network and Computer Applications, 24(3):249– 267, July 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480450190133X.

Wiil:2001:SIS

[141] Uffe Kock Wiil, David L. Hicks, and Peter J. Nürnberg. Special Issue on ‘structural Computing’. Journal of Network and Computer Applications, 24(3):269–271, July 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804501901341.

Anonymous:2001:VC

[142] Anonymous. Volume Contents. Journal of Network and Computer Applications, 24(4):i–viii, October 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804501901390.

Lennon:2001:MHI

[143] J. A. Lennon and H. Maurer. MUSLI: a hypermedia interface for dynamic, interactive, and symbolic communication. Journal of Network and Computer Applications, 24(4):273–291, October 2001. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804501901377.

Lennon:2001:ATV

[144] J. A. Lennon and H. Maurer. Augmenting text and voice conversations with dynamic, interactive abstractions using P2P networking. Journal of Network and Computer Applications, 24(4):293–306, October 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804501901389.

Sabri:2001:GTF

[145] Nawel Sabri and Chantal Taconet. A generic tool to federate WAN servers. Journal of Network and Computer Applications, 24(4):307– 331, October 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804501901365.

Zantout:2001:AEA

[146] Rached N. Zantout and Ahmed A. Guessoum. An automatic English– Arabic HTML page translation system. Journal of Network and Computer Applications, 24(4):333–357, October 2001. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804501901353.

Saxena:2002:HAR

[147] P. C. Saxena, D. R. Choudhury, G. Gabrani, S. Gupta, M. Bhardwaj, and M. Chopra. A heuristic approach to resource locations in broadband networks. Journal of Network and Computer Applications, 25(1):1–35, January 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502901268.

Kouadio:2002:TSI

[148] M. Kouadio and U. Pooch. Technology on social issues of videoconferencing on the Internet: a survey. Journal of Network and Computer Applications, 25(1):37–56, January 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502901256.

Rhee:2002:CMQ

[149] Yoon-Jung Rhee, Eun-Sil Hyun, and Tai-Yun Kim. Connection management for QoS service on the Web. Journal of Network and Computer Applications, 25(1):57–68, January 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804502901281.

Jang:2002:ITT

[150] Heejin Jang and Sangwook Kim. Intruder tracing through dynamic extension of a security domain. Journal of Network and Computer Applications, 25(2):69–91, April 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502901335.

Fan:2002:UDI

[151] Chun-I Fan and Chin-Laung Lei. An unlinkably divisible and intention attachable ticket scheme for runoff elections. Journal of Network and Computer Applications, 25(2):93–107, April 2002. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502901323.

Farag:2002:NPA

[152] Waleed E. Farag and Hussein Abdel-Wahab. A new paradigm for analysis of MPEG compressed videos. Journal of Network and Computer Applications, 25(2):109–127, April 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502901293.

Morris:2002:BDR

[153] T. Morris, P. Blenkhorn, and Farhan Zaidi. Blink detection for real-time eye tracking. Journal of Network and Computer Applications, 25(2): 129–143, April 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480450290130X.

Lee:2002:HQA

[154] Bu-Sung Lee, Chai-Kiat Yeo, and Ruijin Fu. Hybrid quality adaptation mechanism for layered multicast over the Internet. Journal of Network and Computer Applications, 25(2):145–159, April 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450290127X.

Chang:2002:ARM

[155] Chih-Yung Chang and Shin-Chih Tu. Active route-maintenance protocol for signal-based communication path in ad hoc networks. Journal of Network and Computer Applications, 25(3):161–177, July 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502901438.

Piratla:2002:NNB

[156] Nischal M. Piratla and Anura P. Jayasumana. A neural network based real-time gaze tracker. Journal of Network and Computer Applications, 25(3):179–196, July 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502901426.

Behle:2002:RSC

[157] Anita Behle, Michael Kirchhof, Manfred Nagl, and Ralf Welter. Retrieval of software components using a distributed web system. Journal of Network and Computer Applications, 25(3):197–222, July 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502901347.

Hormanseder:2002:MDF

[158] Rudolf Hörmanseder, Johann Murauer, and Jörg R. Mühlbacher. Managing data flow in a DAC supporting security system. Journal of Network and Computer Applications, 25(3):223–238, July 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502901311.

Daoud:2002:PSS

[159] Fawzi Daoud and Takahiko Nomura. Preface to ‘smart spaces’. Journal of Network and Computer Applications, 25(4):239–242, October 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804502901384.

Economou:2002:PDC

[160] Daphne Economou, William L. Mitchell, and Steve R. Pettifer. Problem driven CVE technology development. Journal of Network and Computer Applications, 25(4):243–262, October 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804502901414.

Nomura:2002:DBS

[161] Takahiko Nomura. Design of ‘Ba’ for successful Knowledge Management — how enterprises should design the places of interaction to gain competitive advantage. Journal of Network and Computer Applications, 25(4):263–278, October 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502901396.

Kosuga:2002:MSC

[162] Masakatsu Kosuga, Naoki Kirimoto, Tatsuya Yamazaki, Tomonori Nakanishi, Masakazu Masuzaki, and Kazuo Hasuike. A multimedia service composition scheme for ubiquitous networks. Journal of Network and Computer Applications, 25(4):279–293, October 2002. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502901402.

Maher:2002:CGA

[163] Christopher Maher and Margaret Corbit. Creating genetic applications for informal science learning in multi-user virtual environments. Journal of Network and Computer Applications, 25(4):295– 308, October 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804502901360.

Yamazaki:2002:QBM

[164] Tatsuya Yamazaki. QoS-based media retrieval for networked digital archives. Journal of Network and Computer Applications, 25(4):309– 317, October 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804502901372.

LoPresti:2002:TTA

[165] Stéphane Lo Presti, Didier Bert, and Andrzej Duda. TAO: Temporal Algebraic Operators for modeling multimedia presentations. Journal of Network and Computer Applications, 25(4):319–342, October 2002. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804502901359.

vanderTogt:2003:PN

[166] Hilde van der Togt. Publisher’s Note. Journal of Network and Computer Applications, 26(1):1–2, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804502000619.

Wiil:2003:SCR

[167] Uffe K. Wiil, Peter J. Nürnberg, and David L. Hicks. Structural computing: research directions, systems and issues. Journal of Network and Computer Applications, 26(1):3–9, January 2003. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502000565.

Nurnberg:2003:RAS

[168] Peter J. Nürnberg and Monica M. C. Schraefel. Relationships among structural computing and other fields. Journal of Network and Computer Applications, 26(1):11–26, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804502000577.

Hicks:2003:SRS

[169] David L. Hicks and Uffe K. Wiil. Searching for revolution in structural computing. Journal of Network and Computer Applications, 26(1):27– 45, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804502000589.

Anderson:2003:STT

[170] Kenneth M. Anderson, Susanne A. Sherba, and William Van Lepthien. Structural templates and transformations: the Themis structural computing environment. Journal of Network and Computer Applications, 26(1):47–71, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502000590.

Wang:2003:VIH

[171] Weigang Wang. Visualizing and interacting with hypermedia-based process-centric enterprise models. Journal of Network and Computer Applications, 26(1):73–93, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804502000620.

Millard:2003:DDB

[172] David E. Millard. Discussions at the data border: from generalised hypertext to structural computing. Journal of Network and Computer Applications, 26(1):95–114, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804502000632.

Wiil:2003:CSC

[173] Uffe K. Wiil, Samir Tata, and David L. Hicks. Cooperation services in the Construct structural computing environment. Journal of Network and Computer Applications, 26(1):115–137, January 2003. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804502000656.

Tzagarakis:2003:SPC

[174] Manolis Tzagarakis, Dimitris Avramidis, Maria Kyriakopoulou, Monica M. C. Schraefel, Michalis Vaitis, and Dimitris Christodoulakis. Structuring primitives in the Callimachus component-based open hypermedia system. Journal of Network and Computer Applications, 26(1):139– 162, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804502000644.

Anonymous:2003:C

[175] Anonymous. Calendar. Journal of Network and Computer Applications, 26(1):ifc, January 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000109.

Prabhakaran:2003:EOL

[176] B. Prabhakaran, Yuguang Tu, and Yin Wu. Experiences with an objectlevel scalable web framework. Journal of Network and Computer Applications, 26(2):163–196, April 1, 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450300016X.

Komathy:2003:SXM

[177] K. Komathy, V. Ramachandran, and P. Vivekanandan. Security for XML messaging services — a component-based approach. Journal of Network and Computer Applications, 26(2):197–211, April 1, 2003. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804503000031.

Evans:2003:AAS

[178] Gareth Evans and Paul Blenkhorn. Architectures of assistive software applications for Windows-based computers. Journal of Network and Computer Applications, 26(2):213–228, April 1, 2003. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450300002X.

Zaki:2003:GGB

[179] M. Zaki, M. G. Darwish, and G. Osman. GBF: a grammar based filter for Internet applications. Journal of Network and Computer Applications, 26(3):229–257, ???? 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000195.

Sgouros:2003:RTE

[180] Nikitas M. Sgouros and Stelios Gerogiannakis. Robot teleoperation environments featuring WAP-based wireless devices. Journal of Network and Computer Applications, 26(3):259–271, ???? 2003. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804503000171.

Yeo:2003:FMV

[181] C. K. Yeo, B. S. Lee, and M. H. Er. A framework for multicast video streaming over IP networks. Journal of Network and Computer Applications, 26(3):273–289, ???? 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000043.

Anonymous:2003:IV

[182] Anonymous. Index to volume 26. Journal of Network and Computer Applications, 26(4):III–IV, November 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804503000481.

Chang:2003:DIM

[183] Ching-Lung Chang and Kuo-Sheng Hsu. Design and implementation of MP3-music on demand system using streaming technology. Journal of Network and Computer Applications, 26(4):291–321, November 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804503000286.

Chunlin:2003:AAB

[184] Li Chunlin and Li Layuan. Apply agent to build grid service management. Journal of Network and Computer Applications, 26(4):323–340, November 2003. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804503000183.

Banks:2004:CHP

[185] Jasmine Banks, Geoffery Ericksson, Kevin Burrage, Peter Yellowlees, Sean Ivermee, and Jennifer Tichon. Constructing the hallucinations of psychosis in Virtual Reality. Journal of Network and Computer Applications, 27(1):1–11, January 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000316.

Itani:2004:JAL

[186] Wassim Itani and Ayman Kayssi. J2ME application-layer end-to-end security for m-commerce. Journal of Network and Computer Applications, 27(1):13–32, January 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000304.

Liu:2004:ABA

[187] Duen-Ren Liu and Tzyy-Feng Hwang. An agent-based approach to flexible commerce in intermediary-centric electronic markets. Journal of Network and Computer Applications, 27(1):33–48, January 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804503000390.

Sheikh:2004:UAA

[188] Asrar Sheikh, Tao Wan, and Zaki Alakhdhar. A unified approach to analyze multiple access protocols for buffered finite users. Journal of Network and Computer Applications, 27(1):49–76, January 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804503000298.

Wang:2004:TEM

[189] Heng Wang, Jie Fang, Hua Wang, and Ya-Min Sun. TSDLMRA: an efficient multicast routing algorithm based on Tabu search. Journal of Network and Computer Applications, 27(2):77–90, April 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804503000535.

Bouras:2004:DVR

[190] C. Bouras and Th. Tsiatsos. Distributed virtual reality: building a multiuser layer for the EVE Platform. Journal of Network and Computer Applications, 27(2):91–111, April 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804503000523.

Aderounmu:2004:PCR

[191] G. A. Aderounmu. Performance comparison of remote procedure calling and mobile agent approach to control and data transfer in distributed computing environment. Journal of Network and Computer Applications, 27(2):113–129, April 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000547.

Sheikh:2004:PAB

[192] A. U. H. Sheikh, T. Wan, S. Z. Alakhdhar, I. Lambadaris, and M. Devetsikiotis. Performance analysis of buffered R-ALOHA systems using tagged user approach. Journal of Network and Computer Applications, 27(3):131–150, August 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804503000663.

Yang:2004:CTO

[193] Wen-Lin Yang. A comparison of two optimal approaches for the MCOP problem. Journal of Network and Computer Applications, 27(3):151– 162, August 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804503000560.

Helic:2004:KTP

[194] Denis Helic, Hermann Maurer, and Nick Scerbakov. Knowledge transfer processes in a modern WBT system. Journal of Network and Computer Applications, 27(3):163–190, August 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804503000559.

Anonymous:2004:I

[195] Anonymous. Index. Journal of Network and Computer Applications, 27(4):III–IV, November 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000384.

Lee:2004:CIB

[196] Wei-Bin Lee and Kuan-Chieh Liao. Constructing identity-based cryptosystems for discrete logarithm based cryptosystems. Journal of Network and Computer Applications, 27(4):191–199, November 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804504000268.

Zaki:2004:CAB

[197] M. Zaki and Tarek S. Sobh. A cooperative agent-based model for active security systems. Journal of Network and Computer Applications, 27(4):201–220, November 2004. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450400027X.

Velastin:2004:FCP

[198] Sergio A. Velastin, Benny Lo, and Jie Sun. A flexible communications protocol for a distributed surveillance system. Journal of Network and Computer Applications, 27(4):221–253, November 2004. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804503000651.

Gu:2005:SOM

[199] Tao Gu, Hung Keng Pung, and Da Qing Zhang. A service-oriented middleware for building context-aware services. Journal of Network and Computer Applications, 28(1):1–18, January 2005. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804504000451.

Khedr:2005:AAB

[200] M. Khedr and A. Karmouch. ACAI: agent-based context-aware infrastructure for spontaneous applications. Journal of Network and Computer Applications, 28(1):19–44, January 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804504000293.

Erturk:2005:NMT

[201] Ismail Erturk. A new method for transferring CAN messages using wireless ATM. Journal of Network and Computer Applications, 28(1):45– 56, January 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000281.

Chen:2005:USB

[202] Huamin Chen and Prasant Mohapatra. Using service brokers for accessing backend servers for web applications. Journal of Network and Computer Applications, 28(1):57–74, January 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804504000256.

Anonymous:2005:Aa

[203] Anonymous. Announcement. Journal of Network and Computer Applications, 28(2):I, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000761.

Anonymous:2005:Ab

[204] Anonymous. Announcement. Journal of Network and Computer Applications, 28(2):II, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000773.

Anonymous:2005:P

[205] Anonymous. Preface. Journal of Network and Computer Applications, 28(2):75–76, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000025.

Walker:2005:HTT

[206] Reginald L. Walker. Hierarchical task topology for retrieving information from within a simulated information ecosystem. Journal of Network and Computer Applications, 28(2):77–96, April 2005. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804504000037.

Rungsawang:2005:LTS

[207] A. Rungsawang and N. Angkawattanawit. Learnable topic-specific web crawler. Journal of Network and Computer Applications, 28(2):97– 114, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000086.

Sunayama:2005:PVS

[208] Wataru Sunayama and Masahiko Yachida. Panoramic View System for extracting key sentences based on viewpoints and application to a search engine. Journal of Network and Computer Applications, 28(2): 115–127, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000074.

Sinka:2005:BWD

[209] Mark P. Sinka and David W. Corne. The BankSearch web document dataset: investigating unsupervised clustering and category similarity. Journal of Network and Computer Applications, 28(2):129– 146, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000062.

Wang:2005:IWT

[210] Xiaozhe Wang, Ajith Abraham, and Kate A. Smith. Intelligent web traffic mining and analysis. Journal of Network and Computer Applications, 28(2):147–165, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000050.

Mukkamala:2005:IDU

[211] Srinivas Mukkamala, Andrew H. Sung, and Ajith Abraham. Intrusion detection using an ensemble of intelligent paradigms. Journal of Network and Computer Applications, 28(2):167–182, April 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804504000049.

Shakshuki:2005:ABA

[212] Elhadi Shakshuki, Zhonghai Luo, and Jing Gong. An agent-based approach to security service. Journal of Network and Computer Applications, 28(3):183–208, August 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000487.

Bouras:2005:ADS

[213] Ch. Bouras, A. Gkamas, I. Nave, D. Primpas, A. Shani, O. Sheory, K. Stamos, and Y. Tzruya. Application on demand system over the Internet. Journal of Network and Computer Applications, 28(3):209– 232, August 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000475.

Gu:2005:TFS

[214] Tao Gu, Hung Keng Pung, and Jian Kang Yao. Towards a flexible service discovery. Journal of Network and Computer Applications, 28(3):233–248, August 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000463.

Anonymous:2005:IV

[215] Anonymous. Index to Volume 28. Journal of Network and Computer Applications, 28(4):III–IV, November 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804505000251.

DiSorte:2005:PBS

[216] D. Di Sorte and G. Reali. Pricing and brokering services over interconnected IP networks. Journal of Network and Computer Applications, 28(4):249–283, November 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000505.

Bouras:2005:SLA

[217] Christos Bouras and Afrodite Sevasti. Service level agreements for DiffServ-based services’ provisioning. Journal of Network and Computer Applications, 28(4):285–302, November 2005. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804504000499.

Zhang:2006:FAS

[218] Bing-Yi Zhang and Ya-Min Sun. Fractional alpha stable network traffic model and its application in QoS routing. Journal of Network and Computer Applications, 29(1):1–10, January 2006. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000032.

Thompson:2006:RTF

[219] Elizabeth A. Thompson, Eric Harmison, Robert Carper, Robert Martin, and Jim Isaacs. Robot teleoperation featuring commercially available wireless network cards. Journal of Network and Computer Applications, 29(1):11–24, January 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000682.

dePinho:2006:AES

[220] Leonardo Bidese de Pinho and Claudio Luis de Amorim. Assessing the efficiency of stream reuse techniques in P2P video-on-demand systems. Journal of Network and Computer Applications, 29(1):25– 45, January 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000670.

Zhang:2006:PEF

[221] N. Zhang, Q. Shi, M. Merabti, and R. Askwith. Practical and efficient fair document exchange over networks. Journal of Network and Computer Applications, 29(1):46–61, January 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804504000669.

Morris:2006:FFT

[222] T. Morris and V. Chauhan. Facial feature tracking for cursor control. Journal of Network and Computer Applications, 29(1):62–80, January 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804504000517.

Anonymous:2006:EBa

[223] Anonymous. Editorial board. Journal of Network and Computer Applications, 29(1):CO2, January 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000640.

Ichalkaranje:2005:E

[224] Nikhil Ichalkaranje. Editorial. Journal of Network and Computer Applications, 29(2–3):81–82, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804504000591.

Nakano:2005:CPW

[225] Yukiko I. Nakano, Toshihiro Murayama, Masashi Okamoto, Daisuke Kawahara, Qing Li, Sadao Kurohashi, and Toyoaki Nishida. Cards-to-presentation on the web: generating multimedia contents featuring agent animations. Journal of Network and Computer Applications, 29(2–3): 83–104, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804504000608.

Urlings:2005:FFI

[226] Pierre Urlings, Christos Sioutis, Jeff Tweedale, Nikhil Ichalkaranje, and Lakhmi Jain. A future framework for interfacing BDI agents in a realtime teaming environment. Journal of Network and Computer Applications, 29(2–3):105–123, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480450400061X.

Krishnamurthy:2005:DAP

[227] E. V. Krishnamurthy and V. K. Murthy. Distributed agent paradigm for soft and hard computation. Journal of Network and Computer Applications, 29(2–3):124–146, August/April 2005/2006. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804504000621.

Fyfe:2005:TIA

[228] Colin Fyfe and Lakhmi Jain. Teams of intelligent agents which learn using artificial immune systems. Journal of Network and Computer Applications, 29(2–3):147–159, August/April 2005/2006. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804504000633.

Jarvis:2005:TBH

[229] J. Jarvis, R. Rönnquist, D. McFarlane, and L. Jain. A team-based holonic approach to robotic assembly cell control. Journal of Network and Computer Applications, 29(2–3):160–176, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804504000645.

Morch:2005:EAP

[230] Anders I. Mørch, Jan A. Dolonen, and Jan Eirik B. Nævdal. An evolutionary approach to prototyping pedagogical agents: from simulation to integrated system. Journal of Network and Computer Applications, 29 (2–3):177–199, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804504000657.

Chen:2005:STI

[231] Weiqin Chen. Supporting teachers’ intervention in collaborative knowledge building. Journal of Network and Computer Applications, 29(2–3): 200–215, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000020.

Anonymous:2005:EBa

[232] Anonymous. Editorial board. Journal of Network and Computer Applications, 29(2–3):CO2, August/April 2005/2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804505000743.

Akintola:2006:PAD

[233] A. A. Akintola, G. A. Aderounmu, and M. O. Adigun. Performance analysis of double buffer technique (DBT) model for mobility support in wireless IP networks. Journal of Network and Computer Applications, 29(4):217–234, November 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000299.

Wu:2006:EQS

[234] Shih-Lin Wu, Sui Fan-Jiang, and Zi-Tsan Chou. An efficient quality-of-service MAC protocol for infrastructure WLANs. Journal of Network and Computer Applications, 29(4):235–261, November 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000159.

Park:2006:SCC

[235] Hea-Sook Park and Doo-Kwon Baik. A study for control of client value using cluster analysis. Journal of Network and Computer Applications, 29(4):262–276, November 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000160.

Loidl:2006:TPL

[236] Susanne Loidl. Towards pervasive learning: WeLearn.Mobile. A CPS package viewer for handhelds. Journal of Network and Computer Applications, 29(4):277–293, November 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804505000147.

Liu:2006:RTD

[237] Peng Liu, Hai Wang, and Lunquan Li. Real-time data attack isolation for commercial database applications. Journal of Network and Computer Applications, 29(4):294–320, November 2006. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000068.

Bravo:2006:CDE

[238] Crescencio Bravo, Miguel Á. Redondo, Manuel Ortega, and M. Felisa Verdejo. Collaborative distributed environments for learning design tasks by means of modelling and simulation. Journal of Network and Computer Applications, 29(4):321–342, November 2006. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000056.

Avvenuti:2006:ALN

[239] M. Avvenuti and A. Vecchio. Application-level network emulation: the EmuSocket toolkit. Journal of Network and Computer Applications, 29(4):343–360, November 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000044.

Anonymous:2006:EBb

[240] Anonymous. Editorial board. Journal of Network and Computer Applications, 29(4):CO2, November 2006. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000154.

Abraham:2007:NIS

[241] Ajith Abraham, Kate Smith, Ravi Jain, and Lakhmi Jain. Network and information security: a computational intelligence approach: Special issue of Journal of Network and Computer Applications. Journal of Network and Computer Applications, 30(1):1–3, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000378.

Wang:2007:VBW

[242] Feng-Hsing Wang, Lakhmi C. Jain, and Jeng-Shyang Pan. VQ-based watermarking scheme with genetic codebook partition. Journal of Network and Computer Applications, 30(1):4–23, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450500038X.

Yue:2007:SEV

[243] Tai-Wen Yue and Suchen Chiang. The semipublic encryption for visual cryptography using Q’tron neural networks. Journal of Network and Computer Applications, 30(1):24–41, January 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000391.

Ahn:2007:TSI

[244] Gail-Joon Ahn, Badrinath Mohan, and Seng-Phil Hong. Towards secure information sharing using role-based delegation. Journal of Network and Computer Applications, 30(1):42–59, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000408.

Liao:2007:AAD

[245] Yihua Liao, V. Rao Vemuri, and Alejandro Pasos. Adaptive anomaly detection with evolving connectionist systems. Journal of Network and Computer Applications, 30(1):60–80, January 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450500041X.

Abraham:2007:DSD

[246] Ajith Abraham, Ravi Jain, Johnson Thomas, and Sang Yong Han. D-SCIDS: Distributed soft computing intrusion detection system. Journal of Network and Computer Applications, 30(1):81–98, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804505000421.

Ozyer:2007:IDI

[247] Tansel Özyer, Reda Alhajj, and Ken Barker. Intrusion detection by integrating boosting genetic fuzzy classifier and data mining criteria for rule pre-screening. Journal of Network and Computer Applications, 30(1):99–113, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000433.

Peddabachigari:2007:MID

[248] Sandhya Peddabachigari, Ajith Abraham, Crina Grosan, and Johnson Thomas. Modeling intrusion detection system using hybrid intelligent systems. Journal of Network and Computer Applications, 30(1):114– 132, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000445.

ElKateeb:2007:HSR

[249] Ali El Kateeb. High-speed routers design using data stream distributor unit. Journal of Network and Computer Applications, 30(1):133– 144, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000603.

Hongsong:2007:DPE

[250] Chen Hongsong, Ji Zhenzhou, Hu Mingzeng, Fu Zhongchuan, and Jiang Ruixiang. Design and performance evaluation of a multi-agent-based dynamic lifetime security scheme for AODV routing protocol. Journal of Network and Computer Applications, 30(1):145–166, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000597.

Choong:2007:CDP

[251] Khong Neng Choong, B. M. Ali, V. Prakash, and Yoke Chek Yee. The case for a dynamic proxy framework with voluntary/incentivebased client hosting scheme. Journal of Network and Computer Applications, 30(1):167–185, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000573.

Jafarpour:2007:SGC

[252] Hojjat Jafarpour, Nasser Yazdani, and Navid Bazzaz-zadeh. A scalable group communication mechanism for mobile agents. Journal of Network and Computer Applications, 30(1):186–208, January 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000561.

Kalogeraki:2007:DMA

[253] V. Kalogeraki, D. Zeinalipour-Yazti, D. Gunopulos, and A. Delis. Distributed middleware architectures for scalable media services. Journal of Network and Computer Applications, 30(1):209–243, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480450500055X.

Lu:2007:MAI

[254] Tainchi Lu and Chinghao Hsu. Mobile agents for information retrieval in hybrid simulation environment. Journal of Network and Computer Applications, 30(1):244–264, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804505000548.

Guo:2007:OSM

[255] Hui Guo, Guobin Shen, Zhiguang Wang, and Shipeng Li. Optimized streaming media proxy and its applications. Journal of Network and Computer Applications, 30(1):265–281, January 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000536.

Guo:2007:DSA

[256] Lei Guo, Hongfang Yu, and Lemin Li. Dynamic survivable algorithm for meshed WDM optical networks. Journal of Network and Computer Applications, 30(1):282–295, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804505000354.

Leu:2007:DIB

[257] Jenq-Shiou Leu, Yuan-Po Chi, and Wei-Kuan Shih. Design and implementation of Blog Rendering and Accessing INstantly system (BRAINS). Journal of Network and Computer Applications, 30(1):296– 307, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000366.

Qian:2007:DWA

[258] Lijun Qian, Ning Song, and Xiangfang Li. Detection of wormhole attacks in multi-path routed wireless ad hoc networks: a statistical analysis approach. Journal of Network and Computer Applications, 30(1):308–330, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000342.

Shin:2007:RTN

[259] Kwang Sik Shin, Jin Ha Jung, Jin Young Cheon, and Sang Bang Choi. Real-time network monitoring scheme based on SNMP for dynamic information. Journal of Network and Computer Applications, 30(1):331–353, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000330.

Zhou:2007:QSD

[260] Xiaobo Zhou, Jianbin Wei, and Cheng-Zhong Xu. Quality-of-service differentiation on the Internet: a taxonomy. Journal of Network and Computer Applications, 30(1):354–383, January 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000329.

Kim:2007:AMI

[261] Suneuy Kim and Chita R. Das. An analytical model for interval caching in interactive video servers. Journal of Network and Computer Applications, 30(1):384–413, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000317.

Abadeh:2007:IDU

[262] M. Saniee Abadeh, J. Habibi, and C. Lucas. Intrusion detection using a fuzzy genetics-based learning algorithm. Journal of Network and Computer Applications, 30(1):414–428, January 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000305.

Anonymous:2007:EBa

[263] Anonymous. Editorial board. Journal of Network and Computer Applications, 30(1):CO2, January 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000579.

He:2007:DSL

[264] Rongxi He, Bin Lin, and Lemin Li. Dynamic service-level-agreement aware shared-path protection in WDM mesh networks. Journal of Network and Computer Applications, 30(2):429–444, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000476.

Stephen:2007:DIA

[265] B. Stephen and L. Petropoulakis. The design and implementation of an agent-based framework for acceptable usage policy monitoring and enforcement. Journal of Network and Computer Applications, 30(2): 445–465, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000464.

Liang:2007:WSE

[266] Wang Liang, Guo Yi-Ping, and Fang Ming. Web search engine based on DNS. Journal of Network and Computer Applications, 30(2):466– 478, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000452.

Esteve:2007:VSA

[267] M. Esteve, C. E. Palau, J. Martínez-Nohales, and B. Molina. A video streaming application for urban traffic management. Journal of Network and Computer Applications, 30(2):479–498, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000440.

Filibeli:2007:EWS

[268] M. Can Filibeli, Oznur Ozkasap, and M. Reha Civanlar. Embedded web server-based home appliance networks. Journal of Network and Computer Applications, 30(2):499–514, April 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000257.

Zhou:2007:EAV

[269] Xiaobo Zhou and Cheng-Zhong Xu. Efficient algorithms of video replication and placement on a cluster of streaming servers. Journal of Network and Computer Applications, 30(2):515–540, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000130.

Turner:2007:RAC

[270] Kenneth J. Turner. Representing and analysing composed web services using Cress. Journal of Network and Computer Applications, 30(2): 541–562, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000129.

Christos:2007:WCM

[271] Bouras Christos, Kounenis Giorgos, and Misedakis Ioannis. A web content manipulation technique based on page Fragmentation. Journal of Network and Computer Applications, 30(2):563–585, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000117.

Abidin:2007:DIP

[272] Siti Z. Z. Abidin, Min Chen, and Phil W. Grant. Designing interaction protocols using noughts and crosses type games. Journal of Network and Computer Applications, 30(2):586–613, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000105.

Quintero:2007:LRP

[273] Alejandro Quintero, Da Yu Li, and Harold Castro. A location routing protocol based on smart antennas for ad hoc networks. Journal of Network and Computer Applications, 30(2):614–636, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000099.

Bai:2007:NTM

[274] Y. Bai and M. R. Ito. A new technique for minimizing network loss from users’ perspective. Journal of Network and Computer Applications, 30(2):637–649, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000087.

Guo:2007:NSR

[275] Lei Guo, Jin Cao, Hongfang Yu, and Lemin Li. A new shared-risk link groups (SRLG)-disjoint path provisioning with shared protection in WDM optical networks. Journal of Network and Computer Applications, 30(2):650–661, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000883.

Chang:2007:SEM

[276] Chin-Chen Chang, Chih-Yang Lin, and Keng-Chu Lin. Simple efficient mutual anonymity protocols for peer-to-peer network based on primitive roots. Journal of Network and Computer Applications, 30(2):662– 676, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000871.

Cao:2007:NRS

[277] Jin Cao, Lei Guo, Hongfang Yu, and Lemin Li. A novel recursive shared segment protection algorithm in survivable WDM networks. Journal of Network and Computer Applications, 30(2):677–694, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450500086X.

Lin:2007:PMS

[278] Rongping Lin, Sheng Wang, and Lemin Li. Protections for multicast session in WDM optical networks under reliability constraints. Journal of Network and Computer Applications, 30(2):695–705, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000858.

Min:2007:DUB

[279] Chan-Ho Min and Sehun Kim. On-demand utility-based power control routing for energy-aware optimization in mobile ad hoc networks. Journal of Network and Computer Applications, 30(2):706– 727, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000846.

Yu:2007:RRD

[280] Gwo-Jong Yu, Chih-Yung Chang, Kuei-Ping Shih, and Shih-Chieh Lee. Relay reduction and disjoint routes construction for scatternet over Bluetooth radio system. Journal of Network and Computer Applications, 30(2):728–749, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450500072X.

Shin:2007:PTP

[281] Kwang-Sik Shin, Jin-Ha Jung, Wan-Oh Yoon, and Sang-Bang Choi. P2P transfer of partial stream in multimedia multicast. Journal of Network and Computer Applications, 30(2):750–774, April 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000627.

Xu:2007:DFI

[282] Yiqun Xu, Luigi Logrippo, and Jacques Sincennes. Detecting feature interactions in CPL. Journal of Network and Computer Applications, 30(2):775–799, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000615.

Ozcelik:2007:BIB

[283] Ibrahim Özçelik and Hüseyin Ekiz. Building an interconnection between PROFIBUS and ATM networks. Journal of Network and Computer Applications, 30(2):800–817, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000585.

Anonymous:2007:EBb

[284] Anonymous. Editorial board. Journal of Network and Computer Applications, 30(2):CO2, April 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000033.

Zhou:2007:DDS

[285] Xiaobo Zhou and Cheng-Zhong Xu. Distributed denial-of-service and intrusion detection. Journal of Network and Computer Applications, 30(3):819–822, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000457.

Chen:2007:SDA

[286] Shigang Chen, Yong Tang, and Wenliang Du. Stateful DDoS attacks and targeted filtering. Journal of Network and Computer Applications, 30(3):823–840, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000469.

Bossardt:2007:EIS

[287] Matthias Bossardt, Thomas Dübendorfer, and Bernhard Plattner. Enhanced Internet security by a distributed traffic control service based on traffic ownership. Journal of Network and Computer Applications, 30(3):841–857, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804505000470.

Siris:2007:PBD

[288] Vasilios A. Siris and Ilias Stavrakis. Provider-based deterministic packet marking against distributed DoS attacks. Journal of Network and Computer Applications, 30(3):858–876, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804505000482.

Peng:2007:ISD

[289] Tao Peng, Christopher Leckie, and Kotagiri Ramamohanarao. Information sharing for distributed intrusion detection systems. Journal of Network and Computer Applications, 30(3):877–899, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000494.

Lopez:2007:NAC

[290] Gabriel López, Oscar Cánovas, Antonio F. Gómez, Jesús D. Jiménez, and Rafael Marín. A network access control approach based on the AAA architecture and authorization attributes. Journal of Network and Computer Applications, 30(3):900–919, August 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804505000500.

Li:2007:KMR

[291] Celia Li, Cungang Yang, and Richard Cheung. Key management for role hierarchy in distributed systems. Journal of Network and Computer Applications, 30(3):920–936, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804505000512.

Wu:2007:SEK

[292] Bing Wu, Jie Wu, Eduardo B. Fernandez, Mohammad Ilyas, and Spyros Magliveras. Secure and efficient key management in mobile ad hoc networks. Journal of Network and Computer Applications, 30(3):937– 954, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804505000524.

Magoulas:2007:IBA

[293] George D. Magoulas and George Ghinea. Intelligence-based adaptation for ubiquitous multimedia communications. Journal of Network and Computer Applications, 30(3):955–957, August 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000026.

Jannach:2007:KBM

[294] Dietmar Jannach and Klaus Leopold. Knowledge-based multimedia adaptation for ubiquitous multimedia consumption. Journal of Network and Computer Applications, 30(3):958–982, August 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000063.

Cranley:2007:DCB

[295] Nicola Cranley, Philip Perry, and Liam Murphy. Dynamic contentbased adaptation of streamed multimedia. Journal of Network and Computer Applications, 30(3):983–1006, August 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000051.

OGrady:2007:DAT

[296] M. J. O’Grady, G. M. P. O’Hare, and C. Donaghey. Delivering adaptivity through context-awareness. Journal of Network and Computer Applications, 30(3):1007–1033, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000075.

Charvillat:2007:RLD

[297] Vincent Charvillat and Romulus Grigoras. Reinforcement learning for dynamic multimedia adaptation. Journal of Network and Computer Applications, 30(3):1034–1058, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480450600004X.

Bollen:2007:UDR

[298] Johan Bollen, Michael L. Nelson, Gary Geisler, and Raquel Araujo. Usage derived recommendations for a video digital library. Journal of Network and Computer Applications, 30(3):1059–1083, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804506000038.

Phillips-Wren:2007:IAC

[299] Gloria Phillips-Wren, Anders Morch, Jeff Tweedale, and Nikhil Ichalkaranje. Innovations in agent collaboration, cooperation and Teaming, Part 2. Journal of Network and Computer Applications, 30(3):1085– 1088, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000348.

Tweedale:2007:IMA

[300] J. Tweedale, N. Ichalkaranje, C. Sioutis, B. Jarvis, A. Consoli, and G. Phillips-Wren. Innovations in multi-agent systems. Journal of Network and Computer Applications, 30(3):1089–1115, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480450600035X.

Padgham:2007:ICS

[301] Lin Padgham and Wei Liu. Internet collaboration and service composition as a loose form of teamwork. Journal of Network and Computer Applications, 30(3):1116–1135, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000361.

DiStefano:2007:FDA

[302] Antonella Di Stefano, Giuseppe Pappalardo, Corrado Santoro, and Emiliano Tramontana. A framework for the design and automated implementation of communication aspects in multi-agent systems. Journal of Network and Computer Applications, 30(3):1136–1152, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804506000373.

Xu:2007:GMG

[303] Wei Xu, Jiannong Cao, Beihong Jin, Jing Li, and Liang Zhang. GCS-MA: a group communication system for mobile agents. Journal of Network and Computer Applications, 30(3):1153–1172, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000385.

Wanyama:2007:PMA

[304] Tom Wanyama and Behrouz Homayoun Far. A protocol for multiagent negotiation in a group-choice decision making process. Journal of Network and Computer Applications, 30(3):1173–1195, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804506000397.

Weiss:2007:SIC

[305] Gerhard Weiß, Matthias Nickles, Michael Rovatsos, and Felix Fischer. Specifying the intertwining of cooperation and autonomy in agent-based systems. Journal of Network and Computer Applications, 30(3):1196– 1215, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000403.

Zhan:2007:PPC

[306] Justin Zhan, Stan Matwin, and LiWu Chang. Privacy-preserving collaborative association rule mining. Journal of Network and Computer Applications, 30(3):1216–1227, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000415.

Ouardani:2007:SPM

[307] Abdelhamid Ouardani, Samuel Pierre, and Hanifa Boucheneb. A security protocol for mobile agents based upon the cooperation of sedentary agents. Journal of Network and Computer Applications, 30(3):1228– 1243, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000427.

Aguilar-Ponce:2007:NSB

[308] Ruth Aguilar-Ponce, Ashok Kumar, J. Luis Tecpanecatl-Xihuitl, and Magdy Bayoumi. A network of sensor-based framework for automated visual surveillance. Journal of Network and Computer Applications, 30(3):1244–1271, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000439.

Anonymous:2007:EBc

[309] Anonymous. Editorial board. Journal of Network and Computer Applications, 30(3):ifc, August 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000161.

He:2007:E

[310] Xiangjian He. Editorial. Journal of Network and Computer Applications, 30(4):1273–1274, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000683.

Suzuki:2007:CBE

[311] Shinichi Suzuki, Yasushi Shinjo, Toshio Hirotsu, Kozo Itano, and Kazuhiko Kato. Capability-based egress network access control by using DNS server. Journal of Network and Computer Applications, 30 (4):1275–1282, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000725.

Wu:2007:EMS

[312] Yanni Wu, Kuo Zhang, Xiaoge Wang, and Jinlan Tian. Extending metadata with scenarios in adaptive distributed system. Journal of Network and Computer Applications, 30(4):1283–1294, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000701.

Callaghan:2007:CSA

[313] M. J. Callaghan, J. Harkin, E. McColgan, T. M. McGinnity, and L. P. Maguire. Client-server architecture for collaborative remote experimentation. Journal of Network and Computer Applications, 30 (4):1295–1308, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000749.

McGregor:2007:DVN

[314] Carolyn McGregor, Bruce Kneale, and Mark Tracy. On-demand Virtual Neonatal Intensive Care units supporting rural, remote and urban healthcare with Bush Babies Broadband. Journal of Network and Computer Applications, 30(4):1309–1323, November 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000750.

Jia:2007:RBL

[315] Wenjing Jia, Huaifeng Zhang, and Xiangjian He. Region-based license plate detection. Journal of Network and Computer Applications, 30 (4):1324–1333, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000762.

Gunes:2007:BME

[316] Hatice Gunes and Massimo Piccardi. Bi-modal emotion recognition from expressive face and body gestures. Journal of Network and Computer Applications, 30(4):1334–1345, November 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000774.

Xia:2007:NDI

[317] Junjun Xia, Lin Ni, and Y. Miao. A new digital implementation of ridgelet transform for images of dyadic length. Journal of Network and Computer Applications, 30(4):1346–1355, November 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000695.

Morrison:2007:VES

[318] Donn Morrison and Liyanage C. De Silva. Voting ensembles for spoken affect classification. Journal of Network and Computer Applications, 30(4):1356–1365, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000737.

Wang:2007:TLM

[319] Ben-Nian Wang, Yang Gao, Zhao-Qian Chen, Jun-Yuan Xie, and Shi-Fu Chen. A two-layered multi-agent reinforcement learning model and algorithm. Journal of Network and Computer Applications, 30 (4):1366–1376, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000713.

Lee:2007:SCC

[320] Jung-San Lee and Chin-Chen Chang. Secure communications for clusterbased ad hoc networks using node identities. Journal of Network and Computer Applications, 30(4):1377–1396, November 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000798.

Martinez:2007:OXH

[321] Christopher J. Martinez, Wei-Ming Lin, and Parimal Patel. Optimal XOR hashing for non-uniformly distributed address lookup in computer networks. Journal of Network and Computer Applications, 30 (4):1397–1427, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000786.

Duysburgh:2007:DAS

[322] Bart Duysburgh, Thijs Lambrecht, Filip De Turck, Bart Dhoedt, and Piet Demeester. Design and analysis of a stable set-up protocol for transcoding multicast trees in active networks. Journal of Network and Computer Applications, 30(4):1428–1444, November 2007. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450600049X.

Al-Talib:2007:IMS

[323] S. A. Al-Talib, B. M. Ali, S. Khatun, and S. Subramaniam. Improving the multicast state scalability in Internet routers by integrating hash algorithm with recursive unicast. Journal of Network and Computer Applications, 30(4):1445–1454, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000488.

Anonymous:2007:EBd

[324] Anonymous. Editorial board. Journal of Network and Computer Applications, 30(4):ifc, November 2007. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000367.

Kung:2008:ISA

[325] Hsu-Yang Kung, Hao-Hsiang Ku, Che-I Wu, and Ching-Yu Lin. Intelligent and situation-aware pervasive system to support debris-flow disaster prediction and alerting in Taiwan. Journal of Network and Computer Applications, 31(1):1–18, January 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000506.

Daoud:2008:MCS

[326] Mosaab Daoud and Qusay H. Mahmoud. Monte Carlo simulationbased algorithms for estimating the reliability of mobile agent-based systems. Journal of Network and Computer Applications, 31(1):19– 31, January 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000518.

Qin:2008:PCL

[327] Xiao Qin. Performance comparisons of load balancing algorithms for I/O-intensive workloads on clusters. Journal of Network and Computer Applications, 31(1):32–46, January 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000531.

Chen:2008:MRW

[328] Ming-Tsung Chen, B. M. T. Lin, and Shian-Shyong Tseng. Multicast routing and wavelength assignment with delay constraints in WDM networks with heterogeneous capabilities. Journal of Network and Computer Applications, 31(1):47–65, January 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000543.

Raste:2008:DIS

[329] Tushar M. Raste and D. B. Kulkarni. Design and implementation scheme for deploying IPv4 over IPv6 tunnel. Journal of Network and Computer Applications, 31(1):66–72, January 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480450600052X.

Anonymous:2008:EBa

[330] Anonymous. Editorial board. Journal of Network and Computer Applications, 31(1):CO2, January 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000549.

Makris:2008:TPP

[331] Christos Makris and Athanasios Tsakalidis. Trends in peer to peer and service oriented computing. Journal of Network and Computer Applications, 31(2):73–74, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000336.

Arroyo:2008:PMI

[332] Sinuhe Arroyo, Miguel-Angel Sicilia, and José-Manuel López-Cobo. Patterns of message interchange in decoupled hypermedia systems.

Journal of Network and Computer Applications, 31(2):75–92, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804506000269.

Pastore:2008:SDM

[333] Serena Pastore. The service discovery methods issue: a web services UDDI specification framework integrated in a grid environment. Journal of Network and Computer Applications, 31(2):93– 107, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000270.

Diamadopoulou:2008:TSW

[334] Vassiliki Diamadopoulou, Christos Makris, Yannis Panagis, and Evangelos Sakkopoulos. Techniques to support Web Service selection and consumption with QoS characteristics. Journal of Network and Computer Applications, 31(2):108–130, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000282.

Votis:2008:OPS

[335] K. Votis, C. Alexakos, B. Vassiliadis, and S. Likothanassis. An ontologically principled service-oriented architecture for managing distributed e-government nodes. Journal of Network and Computer Applications, 31(2):131–148, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804506000294.

Sioutas:2008:BDW

[336] S. Sioutas, E. Sakkopoulos, L. Drossos, and S. Sirmakessis. Balanced distributed web service lookup system. Journal of Network and Computer Applications, 31(2):149–162, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000300.

Ranasinghe:2008:DPA

[337] Sampath N. Ranasinghe, Vasilis Friderikos, and Hamid Aghvami. Dynamic proxy assisted mobility support for third generation peer-to-peer networks. Journal of Network and Computer Applications, 31(2):163– 178, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000312.

Kanellopoulos:2008:ETD

[338] Dimitris N. Kanellopoulos and Alkiviadis A. Panagopoulos. Exploiting tourism destinations’ knowledge in an RDF-based P2P network. Journal of Network and Computer Applications, 31(2):179– 200, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000324.

Anonymous:2008:EBb

[339] Anonymous. Editorial board. Journal of Network and Computer Applications, 31(2):ifc, April 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450800012X.

Yu:2008:CCB

[340] Gwo-Jong Yu and Chih-Yung Chang. Congestion control of Bluetooth radio system by piconet restructuring. Journal of Network and Computer Applications, 31(3):201–223, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804507000112.

Ayed:2008:CCA

[341] Dhouha Ayed, Chantal Taconet, Guy Bernard, and Yolande Berbers. CADeComp: Context-aware deployment of component-based applications. Journal of Network and Computer Applications, 31(3):224– 257, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000100.

Chunlin:2008:CLO

[342] Li Chunlin and Li Layuan. Cross-layer optimization policy for QoS scheduling in computational grid. Journal of Network and Computer Applications, 31(3):258–284, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804506000828.

Chen:2008:VBI

[343] Lin Chen, Changjun Jiang, and Junjie Li. VGITS: ITS based on intervehicle communication networks and grid technology. Journal of Network and Computer Applications, 31(3):285–302, August 2008. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804506000816.

Liao:2008:MPS

[344] Luhua Liao, Lemin Li, and Sheng Wang. Multicast protection scheme in survivable WDM optical networks. Journal of Network and Computer Applications, 31(3):303–316, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804506000804.

Lembke:2008:ELS

[345] James Lembke and Byung Kyu Choi. Edge-limited scalable QoS flow set-up. Journal of Network and Computer Applications, 31(3):317– 337, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804506000671.

Khendek:2008:MAV

[346] Ferhat Khendek and Tong Zheng. Modeling and analysis of value added services using message sequence charts. Journal of Network and Computer Applications, 31(3):338–356, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804506000555.

Anonymous:2008:EBc

[347] Anonymous. Editorial board. Journal of Network and Computer Applications, 31(3):ifc, August 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000301.

Varona:2008:HFV

[348] Javier Varona, Cristina Manresa-Yee, and Francisco J. Perales. Handsfree vision-based interface for computer accessibility. Journal of Network and Computer Applications, 31(4):357–374, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000210.

Liu:2008:SAM

[349] Hongjun Liu, Ping Luo, and Daoshun Wang. A scalable authentication model based on public keys. Journal of Network and Computer Applications, 31(4):375–386, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804508000106.

Liao:2008:DAW

[350] Wen-Hwa Liao, Yucheng Kao, and Chien-Ming Fan. Data aggregation in wireless sensor networks using ant colony algorithm. Journal of Network and Computer Applications, 31(4):387–401, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450800009X.

Bhaskar:2008:ARD

[351] Vidhyacharan Bhaskar and Patrick Lallement. Activity routing in a distributed supply chain: Performance evaluation with two inputs. Journal of Network and Computer Applications, 31(4):402–428, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000052.

Zheng:2008:FDB

[352] Bing Zheng and Mohammed Atiquzzaman. A framework to determine bounds of maximum loss rate parameter of RED queue for next generation routers. Journal of Network and Computer Applications, 31(4):429–445, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000040.

Komathy:2008:TBE

[353] K. Komathy and P. Narayanasamy. Trust-based evolutionary game model assisting AODV routing against selfishness. Journal of Network and Computer Applications, 31(4):446–471, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000039.

Liu:2008:DEA

[354] Hongjun Liu, Ping Luo, and Daoshun Wang. A distributed expansible authentication model based on Kerberos. Journal of Network and Computer Applications, 31(4):472–486, November 2008. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000027.

Garrigues:2008:SDI

[355] Carles Garrigues, Sergi Robles, and Joan Borrell. Securing dynamic itineraries for mobile agent applications. Journal of Network and Computer Applications, 31(4):487–508, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804507000781.

Safa:2008:CDM

[356] Haidar Safa, Mohamad Chouman, Hassan Artail, and Marcel Karam. A collaborative defense mechanism against SYN flooding attacks in IP networks. Journal of Network and Computer Applications, 31(4):509– 534, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480450700077X.

Sahoo:2008:NRM

[357] Prasan Kumar Sahoo, Chih-Yung Chang, and Sheng-Wen Chang. Novel route maintenance protocols for the Bluetooth ad hoc network with mobility. Journal of Network and Computer Applications, 31(4):535–558, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000768.

Han:2008:PCA

[358] Weili Han, Xingdong Shi, and Ronghua Chen. Process-context aware matchmaking for web service composition. Journal of Network and Computer Applications, 31(4):559–576, November 2008. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804507000756.

Varaprasad:2008:ERA

[359] G. Varaprasad, R. S. D. Wahidabanu, and P. Venkataram. An efficient resource allocation scheme for multimedia applications in MANET. Journal of Network and Computer Applications, 31(4):577– 584, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000744.

Gamess:2008:UBM

[360] Eric Gamess and Rina Surós. An upper bound model for TCP and UDP throughput in IPv4 and IPv6. Journal of Network and Computer Applications, 31(4):585–602, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804507000732.

Kritikou:2008:UPM

[361] Yiouli Kritikou, Panagiotis Demestichas, Evgenia Adamopoulou, Konstantinos Demestichas, Michael Theologou, and Maria Paradia. User Profile Modeling in the context of web-based learning management systems. Journal of Network and Computer Applications, 31(4):603–627, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000720.

Osmani:2008:HAR

[362] Venet Osmani, Sasitharan Balasubramaniam, and Dmitri Botvich. Human activity recognition in pervasive health-care: Supporting efficient remote collaboration. Journal of Network and Computer Applications, 31(4):628–655, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000719.

Khadivi:2008:MCQ

[363] P. Khadivi, S. Samavi, and T. D. Todd. Multi-constraint QoS routing using a new single mixed metrics. Journal of Network and Computer Applications, 31(4):656–676, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804507000707.

Prieto:2008:VEW

[364] Beatriz Prieto, Fernando Tricas, Juan J. Merelo, Antonio Mora, and Alberto Prieto. Visualizing the evolution of a web-based social network. Journal of Network and Computer Applications, 31(4):677–698, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000690.

Al-Kasassbeh:2008:AMA

[365] Mouhammd Al-Kasassbeh and Mo Adda. Analysis of mobile agents in network fault management. Journal of Network and Computer Applications, 31(4):699–711, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804507000689.

Liang:2008:TTI

[366] Janus S. Liang. The troubleshooting task implementation in automotive chassis using virtual interactive technique and knowledge-based approach. Journal of Network and Computer Applications, 31(4):712–734, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000677.

Wang:2008:NAD

[367] Xiang-Yang Wang, Pan-Pan Niu, and Wei Qi. A new adaptive digital audio watermarking based on support vector machine. Journal of Network and Computer Applications, 31(4):735–749, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804507000525.

Abdel-Jaber:2008:PED

[368] Hussein Abdel-Jaber, Mike Woodward, Fadi Thabtah, and Amer Abu-Ali. Performance evaluation for DRED discrete-time queueing network analytical model. Journal of Network and Computer Applications, 31(4):750–770, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000513.

Yee:2008:SBP

[369] Y. C. Yee, K. N. Choong, Andy L. Y. Low, and S. W. Tan. SIPbased proactive and adaptive mobility management framework for heterogeneous networks. Journal of Network and Computer Applications, 31(4):771–792, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000501.

Samper:2008:NIR

[370] Juan J. Samper, Pedro A. Castillo, Lourdes Araujo, J. J. Merelo, Óscar Cordón, and Fernando Tricas. NectaRSS, an intelligent RSS feed reader. Journal of Network and Computer Applications, 31(4):793–806, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000495.

Liao:2008:AMP

[371] Wen-Hwa Liao and Hsiao-Hsien Wang. An asynchronous MAC protocol for wireless sensor networks. Journal of Network and Computer Applications, 31(4):807–820, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804507000458.

Zhang:2008:PCR

[372] Jianjun Zhang, Ling Liu, Lakshmish Ramaswamy, and Calton Pu. Peer-Cast: Churn-resilient end system multicast on heterogeneous overlay networks. Journal of Network and Computer Applications, 31(4):821– 850, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000343.

Odagiri:2008:SDS

[373] Kazuya Odagiri, Rihito Yaegashi, Masaharu Tadauchi, and Naohiro Ishii. Secure DACS Scheme. Journal of Network and Computer Applications, 31(4):851–861, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000331.

Khan:2008:CSN

[374] Javed I. Khan and Sajid S. Shaikh. Computing in social networks with relationship algebra. Journal of Network and Computer Applications, 31(4):862–878, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450700032X.

Ozcelik:2008:ICS

[375] Ibrahim Ozcelik. Interconnection of CAN segments through IEEE 802.16 wireless MAN. Journal of Network and Computer Applications, 31(4):879–890, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804507000318.

Urgaonkar:2008:CSO

[376] Bhuvan Urgaonkar and Prashant Shenoy. Cataclysm: Scalable overload policing for Internet applications. Journal of Network and Computer Applications, 31(4):891–920, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804507000306.

Salah:2008:ARI

[377] K. Salah, P. Calyam, and M. I. Buhari. Assessing readiness of IP networks to support desktop videoconferencing using OPNET. Journal of Network and Computer Applications, 31(4):921–943, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804507000148.

Cai:2008:NAB

[378] Haibin Cai, Fang Pu, Runcai Huang, and Qiying Cao. A novel ANN-based service selection model for ubiquitous computing environments. Journal of Network and Computer Applications, 31(4):944–965, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804507000136.

Georgiou:2008:PSA

[379] Eleni Georgiou, Marios D. Dikaiakos, and Athena Stassopoulou. On the properties of spam-advertised URL addresses. Journal of Network and Computer Applications, 31(4):966–985, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804507000124.

Anonymous:2008:EBd

[380] Anonymous. Editorial board. Journal of Network and Computer Applications, 31(4):ifc, November 2008. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000751.

Huang:2009:BAN

[381] Tzu-Chi Huang, Ce-Kuen Shieh, and Yu-Ben Miao. Bottleneck Active Node Detouring for capsule-based active network. Journal of Network and Computer Applications, 32(1):1–30, January 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000428.

Guan:2009:FID

[382] Xiaohong Guan, Wei Wang, and Xiangliang Zhang. Fast intrusion detection based on a non-negative matrix factorization model. Journal of Network and Computer Applications, 32(1):31–44, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000404.

Wu:2009:REL

[383] Di Wu, Ye Tian, and Kam-Wing Ng. Resilient and efficient load balancing in distributed hash tables. Journal of Network and Computer Applications, 32(1):45–60, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000398.

Choi:2009:EAD

[384] JungYul Choi, Jaebong Lee, ShinJung Woo, and ByoungKwon Shim. An engineering approach to the design of a BcN control network. Journal of Network and Computer Applications, 32(1):61–77, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480450800043X.

Davies:2009:LWB

[385] Pete Davies and Theodore Tryfonas. A lightweight web-based vulnerability scanner for small-scale computer network security assessment. Journal of Network and Computer Applications, 32(1):78–95, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000416.

Artail:2009:DMD

[386] Hassan Artail, Manal Shihab, and Haidar Safa. A distributed mobile database implementation on Pocket PC mobile devices communicating over Bluetooth. Journal of Network and Computer Applications, 32(1):96–115, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000386.

Demestichas:2009:TAM

[387] Konstantinos P. Demestichas, Evgenia F. Adamopoulou, John G. Markoulidakis, and Michael E. Theologou. Towards Anonymous Mobile Community services. Journal of Network and Computer Applications, 32(1):116–134, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000283.

Fortino:2009:HCP

[388] Giancarlo Fortino, Carlo Mastroianni, and Wilma Russo. A hierarchical control protocol for group-oriented playbacks supported by content distribution networks. Journal of Network and Computer Applications, 32(1):135–157, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000271.

Yu:2009:IER

[389] Meng Yu, Peng Liu, and Wanyu Zang. The implementation and evaluation of a recovery system for workflows. Journal of Network and Computer Applications, 32(1):158–183, January 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000246.

Fagen:2009:VEN

[390] Wade A. Fagen, João W. Cangussu, and Ram Dantu. A virtual environment for network testing. Journal of Network and Computer Applications, 32(1):184–214, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450800026X.

Papagianni:2009:PES

[391] Chryssa A. Papagianni, Nikolaos D. Tselikas, Evangelos A. Kosmatos, Stauros Papapanagiotou, and Iakovos S. Venieris. Performance evaluation study for QoS-aware triple play services over entry-level xDSL connections. Journal of Network and Computer Applications, 32(1):215– 225, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000258.

Chang:2009:ROT

[392] Chin-Chen Chang and Jung-San Lee. Robust t-out-of-n oblivious transfer mechanism based on CRT. Journal of Network and Computer Applications, 32(1):226–235, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000234.

Marsono:2009:SRS

[393] Muhammad N. Marsono, M. Watheq El-Kharashi, and Fayez Gebali. A spam rejection scheme during SMTP sessions based on layer-3 email classification. Journal of Network and Computer Applications, 32(1):236–257, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000222.

Wang:2009:MRD

[394] Hua Wang, Zhao Shi, and Shuai Li. Multicast routing for delay variation bound using a modified ant colony algorithm. Journal of Network and Computer Applications, 32(1):258–272, January 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000209.

Chen:2009:OCP

[395] Yuh-Shyan Chen, Yun-Wei Lin, and Chih-Yung Chang. An overlapping communication protocol using improved time-slot leasing for Bluetooth WPANs. Journal of Network and Computer Applications, 32(1):273–292, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000088.

Almeida:2009:IWS

[396] Francisco Almeida, Vicente Blanco, Carlos Delgado, Francisco de Sande, and Adrián Santos. IDEWEP: Web service for astronomical parallel image deconvolution. Journal of Network and Computer Applications, 32(1):293–313, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000076.

DiSorte:2009:QEM

[397] D. Di Sorte, M. Femminella, and G. Reali. QoS-enabled multicast for delivering live events in a Digital Cinema scenario. Journal of Network and Computer Applications, 32(1):314–344, January 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000064.

Anonymous:2009:EBa

[398] Anonymous. Editorial board. Journal of Network and Computer Applications, 32(1):ifc, January 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000891.

Zhou:2009:NSS

[399] Wanlei Zhou and Yang Xiang. Network and system security. Journal of Network and Computer Applications, 32(2):345–346, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804508000465.

Yue:2009:PAO

[400] Chuan Yue and Haining Wang. Profit-aware overload protection in E-commerce Web sites. Journal of Network and Computer Applications, 32(2):347–356, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000477.

Islam:2009:IAM

[401] Md Rafiqul Islam, Wanlei Zhou, Minyi Guo, and Yang Xiang. An innovative analyser for multi-classifier e-mail classification based on grey list analysis. Journal of Network and Computer Applications, 32(2):357– 366, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000489.

Lu:2009:OWQ

[402] Wei-Zhou Lu, Wei-Xuan Gu, and Shun-Zheng Yu. One-way queuing delay measurement and its application on detecting DDoS attack. Journal of Network and Computer Applications, 32(2):367– 376, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000490.

Li:2009:UMA

[403] Bai Li and Lynn Batten. Using mobile agents to recover from node and database compromise in path-based DoS attacks in wireless sensor networks. Journal of Network and Computer Applications, 32(2):377– 387, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000507.

Suriadi:2009:UCF

[404] Suriadi Suriadi, Ernest Foo, and Audun Jøsang. A user-centric federated single sign-on system. Journal of Network and Computer Applications, 32(2):388–401, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000519.

Zou:2009:CIM

[405] Deqing Zou, Ligang He, Hai Jin, and Xueguang Chen. CRBAC: Imposing multi-grained constraints on the RBAC model in the multiapplication environment. Journal of Network and Computer Applications, 32(2):402–411, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000520.

Cao:2009:PBA

[406] Jian Cao, Jinjun Chen, Haiyan Zhao, and Minglu Li. A policybased authorization model for workflow-enabled dynamic process management. Journal of Network and Computer Applications, 32(2):412– 422, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000532.

Qu:2009:NFT

[407] Wenyu Qu, Masaru Kitsuregawa, Hong Shen, and Zhiguang Shan. A novel fault-tolerant execution model by using of mobile agents. Journal of Network and Computer Applications, 32(2):423–432, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804508000544.

Younas:2009:SIS

[408] Muhammad Younas, Irfan Awan, and Laurence T. Yang. Special issue on service oriented computing and advanced network applications. Journal of Network and Computer Applications, 32(2):433, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000556.

Waluyo:2009:MSO

[409] Agustinus Borgy Waluyo, David Taniar, Wenny Rahayu, and Bala Srinivasan. Mobile service oriented architectures for NN-queries. Journal of Network and Computer Applications, 32(2):434–447, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804508000568.

Durresi:2009:FAS

[410] Arjan Durresi, Vamsi Paruchuri, and Leonard Barolli. Fast autonomous system traceback. Journal of Network and Computer Applications, 32(2): 448–454, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480450800057X.

Holton:2009:CBS

[411] D. R. W. Holton, I. Nafea, M. Younas, and I. Awan. A classbased scheme for E-commerce web servers: Formal specification and performance evaluation. Journal of Network and Computer Applications, 32(2):455–460, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000581.

Al-Oqily:2009:TAO

[412] I. Al-Oqily and A. Karmouch. Towards automating overlay network management. Journal of Network and Computer Applications, 32(2): 461–473, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000593.

Lozano:2009:NSA

[413] M. Lozano, P. Morillo, J. M. Orduña, V. Cavero, and G. Vigueras. A new system architecture for crowd simulation. Journal of Network and Computer Applications, 32(2):474–482, March 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450800060X.

Zhang:2009:QAL

[414] Yan Zhang, Laurence T. Yang, Jianhua Ma, and Jun Zheng. Quantitative analysis of location management and QoS in wireless networks. Journal of Network and Computer Applications, 32(2):483– 489, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000611.

Denko:2009:ECL

[415] Mieso K. Denko, Elhadi Shakshuki, and Haroon Malik. Enhanced crosslayer based middleware for mobile ad hoc networks. Journal of Network and Computer Applications, 32(2):490–499, March 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000623.

Anonymous:2009:EBb

[416] Anonymous. Editorial board. Journal of Network and Computer Applications, 32(2):ifc, March 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000083.

Abd-El-Barr:2009:TND

[417] Mostafa Abd-El-Barr. Topological network design: a survey. Journal of Network and Computer Applications, 32(3):501–509, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450800101X.

Lee:2009:AMS

[418] Byoung-Dai Lee, Jon B. Weissman, and Young-Kwang Nam. Adaptive middleware supporting scalable performance for high-end network services. Journal of Network and Computer Applications, 32(3):510– 524, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000994.

Li:2009:TLC

[419] Chunlin Li and Layuan Li. Three-layer control policy for grid resource management. Journal of Network and Computer Applications, 32(3): 525–537, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000982.

Joung:2009:EL

[420] Yuh-Jzer Joung and Chu-Jui Yang. Email licensing. Journal of Network and Computer Applications, 32(3):538–549, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000970.

Ciraci:2009:RQO

[421] Selim Ciraci, Ibrahim Körpeoğu, and Özgür Ulusoy. Reducing query overhead through route learning in unstructured peer-to-peer network. Journal of Network and Computer Applications, 32(3):550– 567, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000878.

Guo:2009:NDR

[422] Songtao Guo, Gang Feng, Xiaofeng Liao, and Qun Liu. Novel delayrange-dependent stability analysis of the second-order congestion control algorithm with heterogeneous communication delays. Journal of Network and Computer Applications, 32(3):568–577, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000866.

Lu:2009:MCM

[423] Eric Jui-Lin Lu, Yung-Fa Huang, and Shu-Chiu Lu. ML-Chord: a multilayered P2P resource sharing model. Journal of Network and Computer Applications, 32(3):578–588, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804508000854.

Lin:2009:IBE

[424] Shi-Jen Lin and Ding-Chyu Liu. An incentive-based electronic payment scheme for digital content transactions over the Internet. Journal of Network and Computer Applications, 32(3):589–598, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000842.

Carroll:2009:SAV

[425] Thomas E. Carroll and Daniel Grosu. A secure and anonymous votercontrolled election scheme. Journal of Network and Computer Applications, 32(3):599–606, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000830.

Tachwali:2009:CSS

[426] Y. Tachwali, W. J. Barnes, and H. Refai. Configurable symbol synchronizers for software-defined radio applications. Journal of Network and Computer Applications, 32(3):607–615, May 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000829.

Hsu:2009:PTH

[427] I-Ching Hsu, Li-Pin Chi, and Sheau-Shong Bor. A platform for transcoding heterogeneous markup documents using ontology-based metadata. Journal of Network and Computer Applications, 32(3):616– 629, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000817.

Samdanis:2009:AHR

[428] Konstantinos Samdanis, Vasilis Friderikos, and A. Hamid Aghvami. Autonomic hierarchical reconfiguration for wireless access networks. Journal of Network and Computer Applications, 32(3):630–641, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804508000738.

Ou:2009:EDF

[429] Li Ou, Xubin He, and Jizhong Han. An efficient design for fast memory registration in RDMA. Journal of Network and Computer Applications, 32(3):642–651, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000726.

Tarng:2009:RND

[430] Jenn-Hwan Tarng, Bing-Wen Chuang, and Pei-Chen Liu. A relay node deployment method for disconnected wireless sensor networks: Applied in indoor environments. Journal of Network and Computer Applications, 32(3):652–659, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000714.

Chand:2009:VDD

[431] Satish Chand, Bijendra Kumar, and Hari Om. Video Data Delivery using Slotted Patching. Journal of Network and Computer Applications, 32(3):660–665, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000702.

Ozkasap:2009:PPA

[432] Oznur Ozkasap, Mine Caglar, and Ali Alagoz. Principles and performance analysis of SeCond: a system for epidemic peer-to-peer content distribution. Journal of Network and Computer Applications, 32(3): 666–683, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000696.

Boukas:2009:PSO

[433] Lambros Boukas, Georgios Kambourakis, and Stefanos Gritzalis. Pandora: an SMS-oriented m-informational system for educational realms. Journal of Network and Computer Applications, 32(3):684–702, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804508000672.

Abbasov:2009:ERA

[434] Babek Abbasov and Serdar Korukoglu. Effective RED: an algorithm to improve RED’s performance by reducing packet loss rate. Journal of Network and Computer Applications, 32(3):703–709, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000684.

Zhang:2009:JDP

[435] Zonghua Zhang and Pin-Han Ho. Janus: a dual-purpose analytical model for understanding, characterizing and countermining multi-stage collusive attacks in enterprise networks. Journal of Network and Computer Applications, 32(3):710–720, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804508000660.

Feng:2009:PIG

[436] Li Feng, Wei Wang, Lina Zhu, and Yi Zhang. Predicting intrusion goal using dynamic Bayesian network with transfer probability estimation. Journal of Network and Computer Applications, 32(3):721– 732, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804508000659.

Firooz:2009:FRM

[437] Mohammad Hamed Firooz, Alireza Nasiri Avanaki, Mohammad Reza Pakravan, and Keivan Ronasi. A fast and reliable multi-sender algorithm for peer-to-peer networks. Journal of Network and Computer Applications, 32(3):733–740, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000647.

Huang:2009:DIE

[438] Tzu-Chi Huang, Sherali Zeadally, Naveen Chilamkurti, and Ce-Kuen Shieh. Design, implementation, and evaluation of a Programmable Bandwidth Aggregation System for home networks. Journal of Network and Computer Applications, 32(3):741–759, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000635.

Chunlin:2009:RSC

[439] Li Chunlin, Zhong Jin Xiu, and Li Layuan. Resource scheduling with conflicting objectives in grid environments: Model and evaluation. Journal of Network and Computer Applications, 32(3):760–769, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508000453.

Abawajy:2009:AHS

[440] J. H. Abawajy. Adaptive hierarchical scheduling policy for enterprise grid computing systems. Journal of Network and Computer Applications, 32(3):770–779, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804508000441.

Anonymous:2009:EBc

[441] Anonymous. Editorial board. Journal of Network and Computer Applications, 32(3):ifc, May 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000265.

Feng:2009:SBM

[442] Li Feng, Zili Li, and Yi Zhang. Security bootstrap model of key presharing by polynomial group in mobile Ad Hoc Network. Journal of Network and Computer Applications, 32(4):781–787, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000447.

Hu:2009:PBS

[443] Jiankun Hu and Fengling Han. A pixel-based scrambling scheme for digital medical images protection. Journal of Network and Computer Applications, 32(4):788–794, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804509000423.

Bouras:2009:IBP

[444] C. Bouras, E. Giannaka, and Thrasyvoulos Tsiatsos. Identifying best practices for supporting broadband growth: Methodology and analysis. Journal of Network and Computer Applications, 32(4):795– 807, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480450900040X.

Sanchez:2009:PAC

[445] Manuel Sánchez, Gabriel López, Óscar Cánovas, and Antonio F. Gómez-Skarmeta. Performance analysis of a cross-layer SSO mechanism for a roaming infrastructure. Journal of Network and Computer Applications, 32(4):808–823, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000411.

Al-Kasassbeh:2009:NFD

[446] Mouhammd Al-Kasassbeh and Mo Adda. Network fault detection with Wiener filter-based agent. Journal of Network and Computer Applications, 32(4):824–833, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000393.

Yang:2009:EAR

[447] Chao-Tung Yang, Ming-Feng Yang, and Wen-Chung Chiang. Enhancement of anticipative recursively adjusting mechanism for redundant parallel file transfer in data grids. Journal of Network and Computer Applications, 32(4):834–845, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000381.

Ilarri:2009:SBM

[448] Sergio Ilarri, Eduardo Mena, and Arantza Illarramendi. A system based on mobile agents to test mobile computing applications. Journal of Network and Computer Applications, 32(4):846–865, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450900037X.

Shih:2009:TCW

[449] Kuei-Ping Shih, Hung-Chang Chen, Chien-Min Chou, and Bo-Jun Liu. On target coverage in wireless heterogeneous sensor networks with multiple sensing units. Journal of Network and Computer Applications, 32(4):866–877, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450900006X.

Sahoo:2009:PEW

[450] Prasan Kumar Sahoo, Jang-Ping Sheu, and Yu-Chia Chang. Performance evaluation of wireless sensor network with hybrid channel access mechanism. Journal of Network and Computer Applications, 32(4): 878–888, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509000034.

Wu:2009:SAU

[451] Dan Wu, Chi Hong Cheong, and Man Hon Wong. Supporting asynchronous update for distributed data cubes. Journal of Network and Computer Applications, 32(4):889–900, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804509000058.

Li:2009:COA

[452] Jiaming Li and Chai Kiat Yeo. Content and overlay-aware scheduling for peer-to-peer streaming in fluctuating networks. Journal of Network and Computer Applications, 32(4):901–912, July 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000022.

Khan:2009:PRE

[453] Javed I. Khan and Sajid S. Shaikh. A phenotype reputation estimation function and its study of resilience to social attacks. Journal of Network and Computer Applications, 32(4):913–924, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804508001008.

Mattson:2009:HME

[454] Robert L. R. Mattson and Somnath Ghosh. HTTP-MPLEX: an enhanced hypertext transfer protocol and its performance evaluation. Journal of Network and Computer Applications, 32(4):925–939, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000046.

Anonymous:2009:EBd

[455] Anonymous. Editorial board. Journal of Network and Computer Applications, 32(4):ifc, July 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000605.

Fortino:2009:NGC

[456] Giancarlo Fortino and Carlo Mastroianni. Next generation content networks. Journal of Network and Computer Applications, 32(5):941–942, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509000757.

Ravi:2009:SDW

[457] Jayashree Ravi, Zhifeng Yu, and Weisong Shi. A survey on dynamic Web content generation and delivery techniques. Journal of Network and Computer Applications, 32(5):943–960, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000526.

Fedak:2009:BDM

[458] Gilles Fedak, Haiwu He, and Franck Cappello. BitDew: a data management and distribution service with multi-protocol file transfer and metadata abstraction. Journal of Network and Computer Applications, 32(5):961–975, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000587.

Pathan:2009:RDR

[459] Mukaddim Pathan and Rajkumar Buyya. Resource discovery and request-redirection for dynamic load sharing in multi-provider peering content delivery networks. Journal of Network and Computer Applications, 32(5):976–990, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804509000502.

Lloret:2009:SPG

[460] Jaime Lloret, Miguel Garcia, Diana Bri, and Juan R. Diaz. Study and performance of a group-based Content Delivery Network. Journal of Network and Computer Applications, 32(5):991–999, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000551.

Molina:2009:NFC

[461] Benjamin Molina, Salvatore F. Pileggi, Manuel Esteve, and Carlos E. Palau. A negotiation framework for content distribution in mobile transient networks. Journal of Network and Computer Applications, 32 (5):1000–1011, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450900054X.

Broberg:2009:MHS

[462] James Broberg, Rajkumar Buyya, and Zahir Tari. MetaCDN: Harnessing ‘Storage Clouds’ for high performance content delivery. Journal of Network and Computer Applications, 32(5):1012–1022, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000514.

Lin:2009:ACU

[463] Ying-Dar Lin, Chun-Nan Lu, Yuan-Cheng Lai, Wei-Hao Peng, and Po-Ching Lin. Application classification using packet size distribution and port association. Journal of Network and Computer Applications, 32 (5):1023–1030, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000484.

Wang:2009:IBD

[464] Ling Wang, Yahia Tachwali, Pramode Verma, and Anjan Ghosh. Impact of bounded delay on throughput in multi-hop networks. Journal of Network and Computer Applications, 32(5):1031–1038, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000563.

Salah:2009:CPF

[465] K. Salah and M. Hamawi. Comparative packet-forwarding measurement of three popular operating systems. Journal of Network and Computer Applications, 32(5):1039–1048, September 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000538.

Hariri:2009:ALM

[466] Behnoosh Hariri, Shervin Shirmohammadi, Mohammad Reza Pakravan, and Mohammad Hossein Alavi. An adaptive latency mitigation scheme for massively multiuser virtual environments. Journal of Network and Computer Applications, 32(5):1049–1063, September 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000496.

Deng:2009:DNA

[467] Yuhui Deng. Deconstructing Network Attached Storage systems. Journal of Network and Computer Applications, 32(5):1064–1072, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000472.

Leu:2009:NNM

[468] Fang-Yie Leu. A novel network mobility handoff scheme using SIP and SCTP for multimedia applications. Journal of Network and Computer Applications, 32(5):1073–1091, September 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000460.

Martinez-Ortiz:2009:LET

[469] Iván Martínez-Ortiz, José-Luis Sierra, Baltasar Fernández-Manjón, and Alfredo Fernández-Valmayor. Language engineering techniques for the development of e-learning applications. Journal of Network and Computer Applications, 32(5):1092–1105, September 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000459.

Zhou:2009:DMD

[470] Chenfeng Vincent Zhou, Christopher Leckie, and Shanika Karunasekera. Decentralized multi-dimensional alert correlation for collaborative intrusion detection. Journal of Network and Computer Applications, 32 (5):1106–1123, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000435.

Anonymous:2009:EBe

[471] Anonymous. Editorial board. Journal of Network and Computer Applications, 32(5):ifc, September 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509000782.

Tarique:2009:SMR

[472] Mohammed Tarique, Kemal E. Tepe, Sasan Adibi, and Shervin Erfani. Survey of multipath routing protocols for mobile ad hoc networks. Journal of Network and Computer Applications, 32(6):1125–1143, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001027.

Liao:2009:PSS

[473] Wen-Hwa Liao and Wen-Ming Yen. Power-saving scheduling with a QoS guarantee in a mobile WiMAX system. Journal of Network and Computer Applications, 32(6):1144–1152, November 2009. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000964.

Wang:2009:RQA

[474] Neng-Chung Wang and Chao-Yang Lee. A reliable QoS aware routing protocol with slot assignment for mobile ad hoc networks. Journal of Network and Computer Applications, 32(6):1153–1166, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000952.

Marti:2009:PER

[475] R. Martí, S. Robles, A. Martín-Campillo, and J. Cucurull. Providing early resource allocation during emergencies: the mobile triage tag. Journal of Network and Computer Applications, 32(6):1167–1182, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509000769.

Watfa:2009:BBA

[476] Mohamed K. Watfa, Omar Mirza, and Jad Kawtharani. BARC: a Battery Aware Reliable Clustering algorithm for sensor networks. Journal of Network and Computer Applications, 32(6):1183–1193, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000745.

Lopez-Nores:2009:KKE

[477] Martín López-Nores, Jorge García-Duque, José J. Pazos-Arias, Yolanda Blanco-Fernández, Manuel Ramos-Cabrer, Alberto Gil-Solla, Rebeca P. Díaz-Redondo, and Ana Fernández-Vilas. KEPPAN: Knowledge exploitation for proactively-planned ad-hoc networks. Journal of Network and Computer Applications, 32(6):1194–1209, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509000721.

Bayilmis:2009:ISA

[478] Cuneyt Bayilmis, Ersoy Kelebekler, Ismail Erturk, Celal Ceken, and Ibrahim Ozcelik. Integration of a speech activated control system and a wireless interworking unit for a CAN-based distributed application. Journal of Network and Computer Applications, 32(6):1210–1218, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509000733.

Hoang:2009:PBA

[479] Xuan Dau Hoang, Jiankun Hu, and Peter Bertok. A program-based anomaly intrusion detection scheme using multiple detection engines and fuzzy inference. Journal of Network and Computer Applications, 32(6):1219–1228, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480450900071X.

Lai:2009:HQD

[480] Chien-Chih Lai, Ren-Guey Lee, Chun-Chieh Hsiao, Hsin-Sheng Liu, and Chun-Chang Chen. A H-QoS-demand personalized home physiological monitoring system over a wireless multi-hop relay network for mobile home healthcare applications. Journal of Network and Computer Applications, 32(6):1229–1241, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804509000708.

Bossche:2009:AMA

[481] Bruno Van Den Bossche, Bart De Vleeschauwer, Tom Verdickt, Filip De Turck, Bart Dhoedt, and Piet Demeester. Autonomic microcell assignment in massively distributed online virtual environments. Journal of Network and Computer Applications, 32(6):1242–1256, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509000575.

Anonymous:2009:EBf

[482] Anonymous. Editorial board. Journal of Network and Computer Applications, 32(6):ifc, November 2009. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001076.

Li:2010:EBB

[483] Chun-Ta Li and Min-Shiang Hwang. An efficient biometrics-based remote user authentication scheme using smart cards. Journal of Network and Computer Applications, 33(1):1–5, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509001192.

Salah:2010:PEC

[484] K. Salah and A. Kahtani. Performance evaluation comparison of Snort NIDS under Linux and Windows Server. Journal of Network and Computer Applications, 33(1):6–15, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804509001040.

Karopoulos:2010:FIP

[485] Giorgos Karopoulos, Georgios Kambourakis, Stefanos Gritzalis, and Elisavet Konstantinou. A framework for identity privacy in SIP. Journal of Network and Computer Applications, 33(1):16–28, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804509001052.

Zhang:2010:ESP

[486] Jianhong Zhang, Chenglian Liu, and Yixian Yang. An efficient secure proxy verifiably encrypted signature scheme. Journal of Network and Computer Applications, 33(1):29–34, January 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509001039.

Mohamed:2010:SCM

[487] Nader Mohamed and Jameela Al-Jaroodi. Self-configured multiplenetwork-interface socket. Journal of Network and Computer Applications, 33(1):35–42, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001015.

Tong:2010:NMA

[488] Ming-Wen Tong, Zong-Kai Yang, and Qing-Tang Liu. A novel model of adaptation decision-taking engine in multimedia adaptation. Journal of Network and Computer Applications, 33(1):43–49, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001003.

Zhang:2010:ASL

[489] Youtao Zhang, Jun Yang, Weijia Li, Linzhang Wang, and Lingling Jin. An authentication scheme for locating compromised sensor nodes in WSNs. Journal of Network and Computer Applications, 33(1):50– 62, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509000976.

Anonymous:2010:EBa

[490] Anonymous. Editorial board. Journal of Network and Computer Applications, 33(1):ifc, January 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001234.

Zhang:2010:WSN

[491] Junqi Zhang and Vijay Varadharajan. Wireless sensor network key management survey and taxonomy. Journal of Network and Computer Applications, 33(2):63–75, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001313.

Han:2010:TDA

[492] Weili Han, Min Xu, Weidong Zhao, and Guofu Li. A trusted decentralized access control framework for the client/server architecture. Journal of Network and Computer Applications, 33(2):76–83, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001489.

Strobbe:2010:IBS

[493] Matthias Strobbe, Olivier Van Laere, Samuel Dauwe, Bart Dhoedt, Filip De Turck, Piet Demeester, Christof van Nimwegen, and Jeroen Vanattenhoven. Interest based selection of user generated content for rich communication services. Journal of Network and Computer Applications, 33(2):84–97, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001441.

Chou:2010:PSO

[494] Cheng-Fu Chou, William C. Cheng, and Leana Golubchik. Performance study of online batch-based digital signature schemes. Journal of Network and Computer Applications, 33(2):98–114, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509001374.

Xu:2010:PTQ

[495] Ming Xu, Shuigeng Zhou, Jihong Guan, and Xiaohua Hu. A pathtraceable query routing mechanism for search in unstructured peer-to-peer networks. Journal of Network and Computer Applications, 33(2): 115–127, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001362.

Moschetta:2010:FSS

[496] Eduardo Moschetta, Rodolfo S. Antunes, and Marinho P. Barcellos. Flexible and secure service discovery in ubiquitous computing. Journal of Network and Computer Applications, 33(2):128–140, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001349.

Lin:2010:ERA

[497] Jenn-Wei Lin and Wei-Shou Wang. An efficient reconstruction approach for improving Bluetree scatternet formation in personal area networks. Journal of Network and Computer Applications, 33(2):141– 155, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001325.

Su:2010:DPA

[498] Ming-Yang Su. Discovery and prevention of attack episodes by frequent episodes mining and finite state machines. Journal of Network and Computer Applications, 33(2):156–167, March 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509001337.

Safa:2010:CIS

[499] Haidar Safa, Hassan Artail, and Mirna Nahhas. A cache invalidation strategy for mobile networks. Journal of Network and Computer Applications, 33(2):168–182, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001210.

Joung:2010:SOH

[500] Yuh-Jzer Joung and Zhang-Wen Lin. On the self-organization of a hybrid peer-to-peer system. Journal of Network and Computer Applications, 33(2):183–202, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001209.

Anonymous:2010:SIJ

[501] Anonymous. Special Issue of Journal of Network and Computer Applications on “Dependable Multimedia Communications: Systems, Services, and Applications”. Journal of Network and Computer Applications, 33(2):203, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001350.

Anonymous:2010:EBb

[502] Anonymous. Editorial board. Journal of Network and Computer Applications, 33(2):ifc, March 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000081.

Khan:2010:SIR

[503] Muhammad Khurram Khan, Mohamed Kamel, and Xudong Jiang. Special issue on: Recent advances and future directions in biometrics personal identification. Journal of Network and Computer Applications, 33(3):205–206, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000548.

Li:2010:AFF

[504] Peng Li, Xin Yang, Kai Cao, Xunqiang Tao, Ruifang Wang, and Jie Tian. An alignment-free fingerprint cryptosystem based on fuzzy vault scheme. Journal of Network and Computer Applications, 33(3):207– 220, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001398.

Liu:2010:MMB

[505] Eryun Liu, Jimin Liang, Liaojun Pang, Min Xie, and Jie Tian. Minutiae and modified Biocode fusion for fingerprint-based key generation. Journal of Network and Computer Applications, 33(3):221– 235, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001386.

Lee:2010:CFT

[506] Chulhan Lee and Jaihie Kim. Cancelable fingerprint templates using minutiae-based bit-strings. Journal of Network and Computer Applications, 33(3):236–246, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001477.

Qi:2010:NIH

[507] Miao Qi, Yinghua Lu, Ning Du, Yinan Zhang, Chengxi Wang, and Jun Kong. A novel image hiding approach based on correlation analysis for secure multimodal biometrics. Journal of Network and Computer Applications, 33(3):247–257, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001404.

Cao:2010:CFD

[508] Kai Cao, Xin Yang, Xunqiang Tao, Peng Li, Yali Zang, and Jie Tian. Combining features for distorted fingerprint matching. Journal of Network and Computer Applications, 33(3):258–267, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480450900143X.

Sarier:2010:IAS

[509] Neyire Deniz Sarier. Improving the accuracy and storage cost in biometric remote authentication schemes. Journal of Network and Computer Applications, 33(3):268–274, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510000068.

Liu:2010:FVR

[510] Zhi Liu, Yilong Yin, Hongjun Wang, Shangling Song, and Qingli Li. Finger vein recognition with manifold learning. Journal of Network and Computer Applications, 33(3):275–282, May 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509001428.

Kim:2010:TRB

[511] Dong-Ju Kim, Jeong-Hoon Shin, and Kwang-Seok Hong. Teeth recognition based on multiple attempts in mobile device. Journal of Network and Computer Applications, 33(3):283–292, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000044.

Delgado-Gomez:2010:IIU

[512] David Delgado-Gómez, Federico Sukno, David Aguado, Carlos Santacruz, and Antonio Artés-Rodriguez. Individual identification using personality traits. Journal of Network and Computer Applications, 33(3):293–299, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001453.

Mehrotra:2010:RII

[513] Hunny Mehrotra, Banshidhar Majhi, and Phalguni Gupta. Robust iris indexing scheme using geometric hashing of SIFT keypoints. Journal of Network and Computer Applications, 33(3):300–313, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804509001416.

Wang:2010:ITD

[514] Jian-Gang Wang, Eric Sung, and Wei-Yun Yau. Incremental twodimensional linear discriminant analysis with applications to face recognition. Journal of Network and Computer Applications, 33(3):314– 322, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000020.

Wang:2010:AWS

[515] Jianzhong Wang, Baoxue Zhang, Shuyan Wang, Miao Qi, and Jun Kong. An adaptively weighted sub-pattern locality preserving projection for face recognition. Journal of Network and Computer Applications, 33(3):323–332, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804509001490.

Yasuda:2010:VBO

[516] Kumiko Yasuda, Daigo Muramatsu, Satoshi Shirato, and Takashi Matsumoto. Visual-based online signature verification using features extracted from video. Journal of Network and Computer Applications, 33 (3):333–341, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804509001465.

Shaikh:2010:CTO

[517] Siraj A. Shaikh and Joseph R. Rabaiotti. Characteristic tradeoffs in designing large-scale biometric-based identity management systems. Journal of Network and Computer Applications, 33(3):342– 351, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000032.

Anonymous:2010:EBc

[518] Anonymous. Editorial board. Journal of Network and Computer Applications, 33(3):ifc, May 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000305.

Lee:2010:VTB

[519] Chung-Nan Lee, Yung-Cheng Kao, and Ming-Te Tsai. A vEB-tree-based architecture for interactive video on demand services in peer-to-peer networks. Journal of Network and Computer Applications, 33(4): 353–362, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000561.

Tamboli:2010:CAC

[520] Neelofer Tamboli and Mohamed Younis. Coverage-aware connectivity restoration in mobile sensor networks. Journal of Network and Computer Applications, 33(4):363–374, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510000433.

Carro-Calvo:2010:GAS

[521] Leo Carro-Calvo, Sancho Salcedo-Sanz, Jose A. Portilla-Figueras, and E. G. Ortiz-García. A genetic algorithm with switch-device encoding for optimal partition of switched industrial Ethernet networks. Journal of Network and Computer Applications, 33(4):375–382, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000287.

Wang:2010:CBN

[522] Lei Wang, Ruizhong Wei, Yaping Lin, and Bo Wang. A clique base node scheduling method for wireless sensor networks. Journal of Network and Computer Applications, 33(4):383–396, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000275.

Zhong:2010:EIP

[523] Feng Zhong, Chai Kiat Yeo, and Bu Sung Lee. Enabling inter-PMIPv6-domain handover with traffic distributors. Journal of Network and Computer Applications, 33(4):397–409, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510000263.

Bouras:2010:ARF

[524] Christos Bouras, Vassilis Poulopoulos, and Vassilis Tsogkas. Adaptation of RSS feeds based on the user profile and on the end device. Journal of Network and Computer Applications, 33(4):410–421, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000251.

Yang:2010:NCB

[525] Yuwang Yang, Chunshan Zhong, Yamin Sun, and Jingyu Yang. Network coding based reliable disjoint and braided multipath routing for sensor networks. Journal of Network and Computer Applications, 33(4): 422–432, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451000024X.

Callado:2010:BNT

[526] Arthur Callado, Judith Kelner, Djamel Sadok, Carlos Alberto Kamienski, and Stênio Fernandes. Better network traffic identification through the independent combination of techniques. Journal of Network and Computer Applications, 33(4):433–446, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510000238.

Gonzalez-Ortega:2010:RTH

[527] D. González-Ortega, F. J. Díaz-Pernas, M. Martínez-Zarzuela, M. Antón-Rodríguez, J. F. Díez-Higuera, and D. Boto-Giralda. Real-time hands, face and facial features detection and tracking: Application to cognitive rehabilitation tests monitoring. Journal of Network and Computer Applications, 33(4):447–466, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000226.

Lin:2010:EBA

[528] Chih-Peng Lin, Jenhui Chen, and Hsing-Lung Chen. An efficient bandwidth allocation algorithm for real-time VBR stream transmission under IEEE 802.16 wireless networks. Journal of Network and Computer Applications, 33(4):467–476, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000214.

Torkestani:2010:ECB

[529] Javad Akbari Torkestani and Mohammad Reza Meybodi. An efficient cluster-based CDMA/TDMA scheme for wireless mobile ad-hoc networks: a learning automata approach. Journal of Network and Computer Applications, 33(4):477–490, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510000202.

Mahmood:2010:CIP

[530] Abdun Naser Mahmood, Jiankun Hu, Zahir Tari, and Christopher Leckie. Critical infrastructure protection: Resource efficient sampling to improve detection of less frequent patterns in network traffic. Journal of Network and Computer Applications, 33(4):491–502, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000196.

Pequeno:2010:FFD

[531] Henrique S. L. Pequeno, George A. M. Gomes, Rossana M. C. Andrade, José N. de Souza, and Miguel F. de Castro. FrameIDTV: a framework for developing interactive applications on digital television environments. Journal of Network and Computer Applications, 33(4): 503–511, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000184.

Cheng:2010:QMT

[532] Hui Cheng, Xingwei Wang, Shengxiang Yang, Min Huang, and Jiannong Cao. QoS multicast tree construction in IP/DWDM optical Internet by bio-inspired algorithms. Journal of Network and Computer Applications, 33(4):512–522, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000056.

Anonymous:2010:EBd

[533] Anonymous. Editorial board. Journal of Network and Computer Applications, 33(4):ifc, July 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451000069X.

Al-Jaroodi:2010:MTN

[534] Jameela Al-Jaroodi and Nader Mohamed. Middleware trends for network applications. Journal of Network and Computer Applications, 33(5):523–524, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000664.

Boldrini:2010:CSA

[535] C. Boldrini, M. Conti, F. Delmastro, and A. Passarella. Context- and social-aware middleware for opportunistic networks. Journal of Network and Computer Applications, 33(5):525–541, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000524.

Madiraju:2010:MEC

[536] Praveen Madiraju, Srilaxmi Malladi, Janaka Balasooriya, Arthi Hariharan, Sushil K. Prasad, and Anu Bourgeois. A methodology for engineering collaborative and ad-hoc mobile applications using SyD middleware. Journal of Network and Computer Applications, 33(5):542–555, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000421.

Pujol-Ahullo:2010:EPA

[537] Jordi Pujol-Ahulló and Pedro García-López. Enabling portability in advanced information-centric services over structured peer-to-peer systems. Journal of Network and Computer Applications, 33(5):556–568, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451000041X.

Rocha:2010:ASP

[538] Bruno P. S. Rocha, Daniel N. O. Costa, Rande A. Moreira, Cristiano G. Rezende, Antonio A. F. Loureiro, and Azzedine Boukerche. Adaptive security protocol selection for mobile computing. Journal of Network and Computer Applications, 33(5):569–587, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000639.

Lopez:2010:MRP

[539] Pedro García López, Raúl Gracia Tinedo, and Josep M. Banús Alsina. Moving routing protocols to the user space in MANET middleware. Journal of Network and Computer Applications, 33(5):588–602, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000536.

Garcia:2010:VIC

[540] Víctor Manuel Álvarez García, María del Puerto Paule Ruiz, and Juan Ramón Pérez Pérez. Voice interactive classroom, a service-oriented software architecture for speech-enabled learning. Journal of Network and Computer Applications, 33(5):603–610, September 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000408.

Gadallah:2010:MSS

[541] Yasser Gadallah, Mohamed Adel Serhani, and Nader Mohamed. Middleware support for service discovery in special operations mobile ad hoc networks. Journal of Network and Computer Applications, 33 (5):611–619, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000445.

Bo:2010:DWT

[542] Cheng Bo, Zhang Yang, Zhou Peng, Duan Hua, Hu Xiaoxiao, Wang Zheng, and Chen Junliang. Development of Web-Telecom based hybrid services orchestration and execution middleware over convergence networks. Journal of Network and Computer Applications, 33(5):620–630, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000603.

Anonymous:2010:EBe

[543] Anonymous. Editorial board. Journal of Network and Computer Applications, 33(5):ifc, September 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001153.

Jung:2010:AAB

[544] Jason J. Jung, Chung-Ming Ou, Ngoc Thanh Nguyen, and Chong Gun Kim. Advances on agent-based network management. Journal of Network and Computer Applications, 33(6):631–632, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001311.

Chen:2010:ABA

[545] Bo Chen. Agent-based artificial immune system approach for adaptive damage detection in monitoring networks. Journal of Network and Computer Applications, 33(6):633–645, November 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000469.

Wu:2010:CMA

[546] Mary Wu and ChongGun Kim. A cost matrix agent for shortest path routing in ad hoc networks. Journal of Network and Computer Applications, 33(6):646–652, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510000482.

Ray:2010:DAM

[547] Pradeep Ray, N. Parameswaran, and Lundy Lewis. Distributed autonomic management: an approach and experiment towards managing service-centric networks. Journal of Network and Computer Applications, 33(6):653–660, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510000512.

Venkatesan:2010:AMA

[548] S. Venkatesan, C. Chellappan, T. Vengattaraman, P. Dhavachelvan, and Anurika Vaish. Advanced mobile agent security models for code integrity and malicious availability check. Journal of Network and Computer Applications, 33(6):661–671, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510000457.

Wang:2010:DRM

[549] Yao-Tien Wang. A dynamic resource management in mobile agent by artificial neural network. Journal of Network and Computer Applications, 33(6):672–681, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510000470.

Chang:2010:RAI

[550] Yue-Shan Chang and Pei-Chun Shih. A resource-awareness information extraction architecture on mobile grid environment. Journal of Network and Computer Applications, 33(6):682–695, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000494.

Manzoor:2010:QMA

[551] Umar Manzoor and Samia Nefti. QUIET: a Methodology for Autonomous Software Deployment using Mobile Agents. Journal of Network and Computer Applications, 33(6):696–706, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000500.

Anonymous:2010:EBf

[552] Anonymous. Editorial board. Journal of Network and Computer Applications, 33(6):ifc, November 2010. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001451.

Subashini:2011:SSI

[553] S. Subashini and V. Kavitha. A survey on security issues in service delivery models of cloud computing. Journal of Network and Computer Applications, 34(1):1–11, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510001281.

Lin:2011:EVO

[554] Dai-Rui Lin, Chih-I Wang, and D. J. Guan. Efficient vehicle ownership identification scheme based on triple-trapdoor chameleon hash function. Journal of Network and Computer Applications, 34(1):12– 19, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001232.

Muppala:2011:CSB

[555] Sireesha Muppala and Xiaobo Zhou. Coordinated session-based admission control with statistical learning for multi-tier Internet applications. Journal of Network and Computer Applications, 34(1):20– 29, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001761.

Kumar:2011:CIA

[556] Neeraj Kumar, Manoj Kumar, and R. B. Patel. Capacity and interference aware link scheduling with channel assignment in wireless mesh networks. Journal of Network and Computer Applications, 34(1):30– 38, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001700.

Mejia:2011:GTT

[557] Marcela Mejia, Néstor Peña, Jose L. Muñoz, Oscar Esparza, and Marco A. Alzate. A game theoretic trust model for on-line distributed evolution of cooperation in MANETs. Journal of Network and Computer Applications, 34(1):39–51, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510001694.

Sharafeddine:2011:LAC

[558] Sanaa Sharafeddine and Rakan Maddah. A lightweight adaptive compression scheme for energy-efficient mobile-to-mobile file sharing applications. Journal of Network and Computer Applications, 34(1):52– 61, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001682.

Sharafeddine:2011:CAM

[559] Sanaa Sharafeddine. Capacity assignment in multiservice packet networks with soft maximum waiting time guarantees. Journal of Network and Computer Applications, 34(1):62–72, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001669.

Li:2011:CIB

[560] Xiong Li, Jian-Wei Niu, Jian Ma, Wen-Dong Wang, and Cheng-Lian Liu. Cryptanalysis and improvement of a biometrics-based remote user authentication scheme using smart cards. Journal of Network and Computer Applications, 34(1):73–79, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001657.

Tian:2011:MHK

[561] Biming Tian, Song Han, Jiankun Hu, and Tharam Dillon. A mutualhealing key distribution scheme in wireless sensor networks. Journal of Network and Computer Applications, 34(1):80–88, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510001645.

Bouras:2011:DMS

[562] Christos Bouras, Eri Giannaka, and Thrasyvoulos Tsiatsos. A dynamic management scheme for DVEs. Journal of Network and Computer Applications, 34(1):89–101, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451000161X.

Han:2011:ESF

[563] Hyuck Han, Hyungsoo Jung, Hyeonsang Eom, and H. Y. Yeom. An efficient skyline framework for matchmaking applications. Journal of Network and Computer Applications, 34(1):102–115, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001591.

Wang:2011:FDS

[564] Song Wang, Jinli Cao, and Jiankun Hu. A frequency domain subspace blind channel estimation method for trailing zero OFDM systems. Journal of Network and Computer Applications, 34(1):116–120, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001578.

Hu:2011:UPM

[565] Chih-Lin Hu and Chien-An Cho. User-provided multimedia content distribution architecture in mobile and ubiquitous communication networks. Journal of Network and Computer Applications, 34(1):121– 136, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451000158X.

Akkus:2011:PPM

[566] Istemi Ekin Akkus, Öznur Özkasap, and M. Reha Civanlar. Peer-to-peer multipoint video conferencing with layered video. Journal of Network and Computer Applications, 34(1):137–150, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001542.

Rodrigues:2011:BDV

[567] Joel J. P. C. Rodrigues, Orlando R. E. Pereira, and Paulo A. C. S. Neves. Biofeedback data visualization for body sensor networks. Journal of Network and Computer Applications, 34(1):151–158, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001530.

Mansoor:2011:FLM

[568] Atif Bin Mansoor, Hassan Masood, Mustafa Mumtaz, and Shoab A. Khan. A feature level multimodal approach for palmprint identification using directional subband energies. Journal of Network and Computer Applications, 34(1):159–171, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510001438.

Lim:2011:RDD

[569] Mingyu Lim, Bart Kevelham, Niels Nijdam, and Nadia Magnenat-Thalmann. Rapid development of distributed applications using highlevel communication support. Journal of Network and Computer Applications, 34(1):172–182, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510001426.

Huang:2011:PMD

[570] Xiaohong Huang and Maode Ma. A performance model for differentiated service over single-hop passive star coupled WDM optical networks. Journal of Network and Computer Applications, 34(1):183– 193, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001402.

Mpitziopoulos:2011:PAE

[571] Aristides Mpitziopoulos, Charalampos Konstantopoulos, Damianos Gavalas, and Grammati Pantziou. A pervasive assistive environment for visually impaired people using wireless sensor network infrastructure. Journal of Network and Computer Applications, 34(1):194–206, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001396.

Daeinabi:2011:VEC

[572] Ameneh Daeinabi, Akbar Ghaffar Pour Rahbar, and Ahmad Khademzadeh. VWCA: an efficient clustering algorithm in vehicular ad hoc networks. Journal of Network and Computer Applications, 34(1):207–222, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510001384.

Goudarzi:2011:IRA

[573] Pejman Goudarzi. Improving rate allocation for ephemeral traffic using a second-order algorithm. Journal of Network and Computer Applications, 34(1):223–231, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001372.

Huang:2011:UCA

[574] Tzu-Chi Huang. Universal connection architecture for interactive applications to achieve distributed computing. Journal of Network and Computer Applications, 34(1):232–251, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001359.

Fu:2011:RLB

[575] Song Fu, Cheng-Zhong Xu, and Haiying Shen. Randomized load balancing strategies with churn resilience in peer-to-peer networks. Journal of Network and Computer Applications, 34(1):252–261, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510001335.

Zhang:2011:EPK

[576] Bo Zhang and Fangguo Zhang. An efficient public key encryption with conjunctive-subset keywords search. Journal of Network and Computer Applications, 34(1):262–267, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510001293.

Ibanez:2011:LRI

[577] Jesús Ibáñez and Carlos Delgado-Mata. Lessons from research on interaction with virtual environments. Journal of Network and Computer Applications, 34(1):268–281, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451000127X.

Sufi:2011:FPI

[578] Fahim Sufi and Ibrahim Khalil. Faster person identification using compressed ECG in time critical wireless telecardiology applications. Journal of Network and Computer Applications, 34(1):282–293, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001268.

Ruiz-Martinez:2011:MNO

[579] Antonio Ruiz-Martínez, Juan Sánchez-Montesinos, and Daniel Sánchez-Martínez. A mobile network operator-independent mobile signature service. Journal of Network and Computer Applications, 34(1):294– 311, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001256.

Niu:2011:MGC

[580] Wenjia Niu, Gang Li, Zhijun Zhao, Hui Tang, and Zhongzhi Shi. Multi-granularity context model for dynamic Web service composition. Journal of Network and Computer Applications, 34(1):312–326, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001360.

Goudarzi:2011:ABR

[581] Hadi Goudarzi, Amir Hesam Salavati, and Mohammad Reza Pakravan. An ant-based rate allocation algorithm for media streaming in peer to peer networks: Extension to multiple sessions and dynamic networks. Journal of Network and Computer Applications, 34(1):327– 340, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451000113X.

Liberatore:2011:ETA

[582] Marc Liberatore, Bikas Gurung, Brian Neil Levine, and Matthew Wright. Empirical tests of anonymous voice over IP. Journal of Network and Computer Applications, 34(1):341–350, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001074.

Moravek:2011:SPL

[583] Patrik Moravek, Dan Komosny, Radim Burget, Jaroslav Sveda, Tomas Handl, and Lucie Jarosova. Study and performance of localization methods in IP based networks: Vivaldi algorithm. Journal of Network and Computer Applications, 34(1):351–367, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001128.

Costa-Montenegro:2011:IAB

[584] E. Costa-Montenegro, J. C. Burguillo-Rial, F. Gil-Castiñeira, and F. J. González-Castaño. Implementation and analysis of the BitTorrent protocol with a multi-agent model. Journal of Network and Computer Applications, 34(1):368–383, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001086.

Konstantinou:2011:ECB

[585] Elisavet Konstantinou. Efficient cluster-based group key agreement protocols for wireless ad hoc networks. Journal of Network and Computer Applications, 34(1):384–393, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510000962.

Choi:2011:RSP

[586] Jung Yul Choi, Shin Jung Woo, and Byoung Kwon Shim. Reliable service provisioning in converged multimedia network environment. Journal of Network and Computer Applications, 34(1):394–401, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000676.

Kulkarni:2011:MMP

[587] Prajakta Kulkarni and Yusuf Ozturk. mPHASiS: Mobile patient healthcare and sensor information system. Journal of Network and Computer Applications, 34(1):402–417, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510000652.

Claycomb:2011:NNL

[588] William R. Claycomb and Dongwan Shin. A novel node level security policy framework for wireless sensor networks. Journal of Network and Computer Applications, 34(1):418–428, January 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000391.

Anonymous:2011:EBa

[589] Anonymous. Editorial board. Journal of Network and Computer Applications, 34(1):ifc, January 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001852.

Loo:2011:SIE

[590] Jonathan Loo. Special issue on: Efficient and robust security and services of wireless mesh networks. Journal of Network and Computer Applications, 34(2):429–431, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001323.

Zdarsky:2011:SAW

[591] Frank A. Zdarsky, Sebastian Robitzsch, and Albert Banchs. Security analysis of wireless mesh backhauls for mobile networks. Journal of Network and Computer Applications, 34(2):432–442, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000640.

Jiang:2011:PSA

[592] Shanshan Jiang and Yuan Xue. Providing survivability against jamming attack for multi-radio multi-channel wireless mesh networks. Journal of Network and Computer Applications, 34(2):443–454, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510000573.

Bankovic:2011:ISW

[593] Zorana Bankovic, David Fraga, José Manuel Moya, Juan Carlos Vallejo, Pedro Malagón, Álvaro Araujo, Juan-Mariano de Goyeneche, Elena Romero, Javier Blesa, Daniel Villanueva, and Octavio Nieto-Taladriz. Improving security in WMNs with reputation systems and self-organizing maps. Journal of Network and Computer Applications, 34(2):455– 463, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000585.

Hsu:2011:NLM

[594] Ching-Fang Hsu, Guo-Hua Cui, Qi Cheng, and Jing Chen. A novel linear multi-secret sharing scheme for group communication in wireless mesh networks. Journal of Network and Computer Applications, 34(2): 464–468, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451000055X.

Han:2011:PEB

[595] Yiliang Han, Xiaolin Gui, Xuguang Wu, and Xiaoyuan Yang. Proxy encryption based secure multicast in wireless mesh networks. Journal of Network and Computer Applications, 34(2):469–477, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510000974.

Muogilim:2011:WMN

[596] Okechukwu E. Muogilim, Kok-Keong Loo, and Richard Comley. Wireless mesh network security: a traffic engineering management approach. Journal of Network and Computer Applications, 34(2):478– 491, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000627.

Lacuesta:2011:TSE

[597] Raquel Lacuesta, Jaime Lloret, Miguel Garcia, and Lourdes Peñalver. Two secure and energy-saving spontaneous ad-hoc protocol for wireless mesh client networks. Journal of Network and Computer Applications, 34(2):492–505, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000597.

Gomes:2011:UFL

[598] Rafael Lopes Gomes, Waldir Moreira Junior, Eduardo Cerqueira, and Antônio Jorge Abelém. Using fuzzy link cost and dynamic choice of link quality metrics to achieve QoS and QoE in wireless mesh networks. Journal of Network and Computer Applications, 34(2):506– 516, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000615.

Mukati:2011:SME

[599] Altaf Mukati. A survey of memory error correcting techniques for improved reliability. Journal of Network and Computer Applications, 34(2): 517–522, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002043.

Mendes:2011:SCL

[600] Lucas D. P. Mendes and Joel J. P. C. Rodrigues. A survey on cross-layer solutions for wireless sensor networks. Journal of Network and Computer Applications, 34(2):523–534, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510002079.

Chen:2011:RSW

[601] Tingting Chen, Ankur Bansal, and Sheng Zhong. A reputation system for wireless mesh networks using network coding. Journal of Network and Computer Applications, 34(2):535–541, March 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510002274.

Montolio-Aranda:2011:IFB

[602] Pere Montolio-Aranda, Joaquin Garcia-Alfaro, and David Megías. Improved flooding of broadcast messages using extended multipoint relaying. Journal of Network and Computer Applications, 34(2):542– 550, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002225.

Ta:2011:MOZ

[603] Duong Nguyen Binh Ta, Suiping Zhou, Wentong Cai, Xueyan Tang, and Rassul Ayani. Multi-objective zone mapping in large-scale distributed virtual environments. Journal of Network and Computer Applications, 34(2):551–561, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510002195.

Sluijs:2011:CCV

[604] Niels Sluijs, Frédéric Iterbeke, Tim Wauters, Filip De Turck, Bart Dhoedt, and Piet Demeester. Cooperative caching versus proactive replication for location dependent request patterns. Journal of Network and Computer Applications, 34(2):562–574, March 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510002183.

Bsoul:2011:EFS

[605] Mohammad Bsoul, Ahmad Al-Khasawneh, Emad Eddien Abdallah, and Yousef Kilani. Enhanced Fast Spread Replication strategy for Data Grid. Journal of Network and Computer Applications, 34(2):575– 580, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002171.

Huang:2011:HAQ

[606] Xiaohong Huang and Maode Ma. A heuristic adaptive QoS prediction scheme in single-hop passive star coupled WDM optical networks. Journal of Network and Computer Applications, 34(2):581– 588, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451000216X.

Villalba:2011:TNI

[607] Cynthia Villalba and Franco Zambonelli. Towards nature-inspired pervasive service ecosystems: Concepts and simulation experiences. Journal of Network and Computer Applications, 34(2):589–602, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510002146.

Chen:2011:OBS

[608] Jianyong Chen, Huawang Zeng, Cunying Hu, and Zhen Ji. Optimization between security and delay of quality-of-service. Journal of Network and Computer Applications, 34(2):603–608, March 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510002109.

Sood:2011:SDI

[609] Sandeep K. Sood, Anil K. Sarje, and Kuldip Singh. A secure dynamic identity based authentication protocol for multi-server architecture. Journal of Network and Computer Applications, 34(2):609– 618, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002092.

Dong:2011:SCR

[610] Hai Dong, Farookh Khadeer Hussain, and Elizabeth Chang. A service concept recommendation system for enhancing the dependability of semantic service matchmakers in the service ecosystem environment. Journal of Network and Computer Applications, 34(2):619– 631, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002080.

Leu:2011:CMC

[611] Fang-Yie Leu, Ilsun You, and Yao-Tien Huang. A coordinated multiple channel assignment scheme and AP deployment for channel reuse in metropolitan scale wireless networks. Journal of Network and Computer Applications, 34(2):632–643, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510002067.

Saad:2011:JAP

[612] Mohamed Saad. Joint admission and power control for quality-of-service in the wireless downlink. Journal of Network and Computer Applications, 34(2):644–652, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001992.

Shi:2011:AAF

[613] Q. Shi, N. Zhang, M. Merabti, and R. Askwith. Achieving autonomous fair exchange in ubiquitous network settings. Journal of Network and Computer Applications, 34(2):653–667, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510002006.

Bernardino:2011:SRL

[614] Anabela Moreira Bernardino, Eugénia Moreira Bernardino, Juan Manuel Sanchez-Perez, Juan Antonio Gomez-Pulido, and Miguel Angel Vega-Rodriguez. Solving ring loading problems using bio-inspired algorithms. Journal of Network and Computer Applications, 34(2):668– 685, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002018.

Milicevic:2011:HAS

[615] Zoran Milicevic and Zoran Bojkovic. H.264/AVC standard: a proposal for selective intra- and optimized inter-prediction. Journal of Network and Computer Applications, 34(2):686–691, March 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451000202X.

Chen:2011:WDD

[616] Yebin Chen, Ying Li, and Jiankun Wang. On the wide diameter of directed double-loop networks. Journal of Network and Computer Applications, 34(2):692–696, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001839.

Belqasmi:2011:ONA

[617] Fatna Belqasmi, Roch Glitho, and Rachida Dssouli. An overlay network for autonomous information discovery in the post-composition registries of ambient networks. Journal of Network and Computer Applications, 34(2):697–707, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001827.

Zhihong:2011:RFT

[618] Huo Zhihong, Zheng Yuan, and Xu Chang. A robust fault-tolerant control strategy for networked control systems. Journal of Network and Computer Applications, 34(2):708–714, March 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001815.

Sani:2011:PAF

[619] Yusuf Sani and Mohamed Othman. A provision-aware fair bandwidth distribution marker algorithm for DiffServ networks. Journal of Network and Computer Applications, 34(2):715–721, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001980.

Su:2011:UCI

[620] Ming-Yang Su. Using clustering to improve the KNN-based classifiers for online anomaly network traffic identification. Journal of Network and Computer Applications, 34(2):722–730, March 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001785.

Movahhedinia:2011:PAB

[621] Naser Movahhedinia and Behrouz Shahgholi Ghahfarokhi. Performance analysis of Bluetooth asynchronous connection-less service. Journal of Network and Computer Applications, 34(2):731–738, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510001773.

Li:2011:TLM

[622] Chunlin Li and Layuan Li. Two-level market solution for services composition optimization in mobile grid. Journal of Network and Computer Applications, 34(2):739–749, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510001566.

Yu:2011:DPT

[623] Li Yu, Maode Ma, Weidong Hu, Zibo Shi, and Yantai Shu. Design of parameter tunable robust controller for active queue management based on H_{∞} control theory. Journal of Network and Computer Applications, 34(2):750–764, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451000175X.

Lu:2011:CLQ

[624] Jinchang Lu and Maode Ma. Cross-layer QoS support framework and holistic opportunistic scheduling for QoS in single carrier WiMAX system. Journal of Network and Computer Applications, 34(2):765– 773, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001748.

Chunlin:2011:CAS

[625] Li Chunlin and Li Layuan. Context aware service provisioning in mobile grid. Journal of Network and Computer Applications, 34(2):774– 782, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001724.

Pham:2011:RBA

[626] Congduc Pham, Abdallah Makhoul, and Rachid Saadi. Risk-based adaptive scheduling in randomly deployed video sensor networks for critical surveillance applications. Journal of Network and Computer Applications, 34(2):783–795, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001712.

Anonymous:2011:EBb

[627] Anonymous. Editorial board. Journal of Network and Computer Applications, 34(2):ifc, March 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000117.

Sheng:2011:RTS

[628] Quan Z. Sheng, Sherali Zeadally, Aikaterini Mitrokotsa, and Zakaria Maamar. RFID technology, systems, and applications. Journal of Network and Computer Applications, 34(3):797–798, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451000130X.

Yao:2011:LCE

[629] Wen Yao, Chao-Hsien Chu, and Zang Li. Leveraging complex event processing for smart hospitals using RFID. Journal of Network and Computer Applications, 34(3):799–810, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510000949.

Ma:2011:LBM

[630] Yi-Wei Ma, Han-Chieh Chao, Jiann-Liang Chen, and Cheng-Yen Wu. Load-balancing mechanism for the RFID middleware applications over grid networking. Journal of Network and Computer Applications, 34(3):

811–820, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000846.

Gandino:2011:PDR

[631] Filippo Gandino, Renato Ferrero, Bartolomeo Montrucchio, and Maurizio Rebaudengo. Probabilistic DCS: an RFID reader-to-reader anticollision protocol. Journal of Network and Computer Applications, 34(3): 821–832, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000810.

Peris-Lopez:2011:FRG

[632] Pedro Peris-Lopez, Agustin Orfila, Julio C. Hernandez-Castro, and Jan C. A. van der Lubbe. Flaws on RFID grouping-proofs. Guidelines for future sound protocols. Journal of Network and Computer Applications, 34(3):833–845, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000822.

Rizzo:2011:ISC

[633] Francesco Rizzo, Marcello Barboni, Lorenzo Faggion, Graziano Azzalin, and Marco Sironi. Improved security for commercial container transports using an innovative active RFID system. Journal of Network and Computer Applications, 34(3):846–852, May 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000809.

DiPietro:2011:OPS

[634] Roberto Di Pietro and Refik Molva. An optimal probabilistic solution for information confinement, privacy, and security in RFID systems. Journal of Network and Computer Applications, 34(3):853– 863, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000895.

Bueno-Delgado:2011:OFL

[635] M. V. Bueno-Delgado and J. Vales-Alonso. On the optimal framelength configuration on real passive RFID systems. Journal of Network and Computer Applications, 34(3):864–876, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000986.

Hancke:2011:DSD

[636] G. P. Hancke. Design of a secure distance-bounding channel for RFID. Journal of Network and Computer Applications, 34(3):877– 887, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000883.

Chen:2011:RNP

[637] Hanning Chen, Yunlong Zhu, Kunyuan Hu, and Tao Ku. RFID network planning using a multi-swarm optimizer. Journal of Network and Computer Applications, 34(3):888–901, May 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000780.

Papapostolou:2011:RAI

[638] Apostolia Papapostolou and Hakima Chaouchi. RFID-assisted indoor localization and the impact of interference on its performance. Journal of Network and Computer Applications, 34(3):902–913, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000834.

Ku:2011:OCT

[639] Wei-Shinn Ku, Kazuya Sakai, and Min-Te Sun. The optimal k-covering tag deployment for RFID-based localization. Journal of Network and Computer Applications, 34(3):914–924, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451000086X.

Manzanares-Lopez:2011:EDD

[640] Pilar Manzanares-Lopez, Juan Pedro Muñoz-Gea, Josemaria Malgosa-Sanahuja, and Juan Carlos Sanchez-Aarnoutse. An efficient distributed discovery service for EPCglobal network in nested package scenarios. Journal of Network and Computer Applications, 34(3):925– 937, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000925.

Yang:2011:HZR

[641] Huanjia Yang, Lili Yang, and Shuang-Hua Yang. Hybrid ZigBee RFID sensor network for humanitarian logistics centre management. Journal of Network and Computer Applications, 34(3):938–948, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000913.

Jakkhupan:2011:BPA

[642] Worapot Jakkhupan, Somjit Arch-int, and Yuefeng Li. Business process analysis and simulation for the RFID and EPCglobal Network enabled supply chain: a proof-of-concept approach. Journal of Network and Computer Applications, 34(3):949–957, May 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000779.

Nam:2011:BAF

[643] Taewoo Nam and Keunhyuk Yeom. Business-aware framework for supporting RFID-enabled applications in EPC Network. Journal of Network and Computer Applications, 34(3):958–971, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000950.

Amaral:2011:EMS

[644] Leonardo A. Amaral, Fabiano P. Hessel, Eduardo A. Bezerra, Jerônimo C. Corrêa, Oliver B. Longhi, and Thiago F. O. Dias. eCloudR-FID — a mobile software framework architecture for pervasive RFIDbased applications. Journal of Network and Computer Applications, 34 (3):972–979, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000792.

Najera:2011:RTL

[645] Pablo Najera, Javier Lopez, and Rodrigo Roman. Real-time location and inpatient care systems based on passive RFID. Journal of Network and Computer Applications, 34(3):980–989, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510000858.

Vo:2011:RMR

[646] Chuong C. Vo, Naveen Chilamkurti, Seng W. Loke, and Torab Torabi. Radio-mama: an RFID based business process framework for asset management. Journal of Network and Computer Applications, 34(3):990– 997, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510000871.

Chaudhry:2011:STT

[647] Maria Chaudhry, Ali Hammad Akbar, Qanita Ahmad, and Imran Sarwar. SOARware^{1}: Treading through the crossroads of RFID middleware and SOA paradigm. Journal of Network and Computer Applications, 34(3):998–1014, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000901.

Ranasinghe:2011:ETL

[648] Damith C. Ranasinghe, Mark Harrison, Kary Främling, and Duncan McFarlane. Enabling through life product-instance management: Solutions and challenges. Journal of Network and Computer Applications, 34(3):1015–1031, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000937.

Anonymous:2011:EBc

[649] Anonymous. Editorial board. Journal of Network and Computer Applications, 34(3):ifc, May 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000348.

Li:2011:ATC

[650] Keqiu Li, Laurence T. Yang, and Xuemin Lin. Advanced topics in cloud computing. Journal of Network and Computer Applications, 34(4):1033– 1034, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001347.

Chen:2011:OSQ

[651] Gang Chen, Yongwei Wu, Jia Liu, Guangwen Yang, and Weimin Zheng. Optimization of sub-query processing in distributed data integration systems. Journal of Network and Computer Applications, 34(4):1035– 1042, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001050.

Zhou:2011:VMS

[652] Z. Zhou, L. T. Yang, S. Bhiri, L. Shu, N. Xiong, and M. Hauswirth. Verifying mediated service interactions considering expected behaviours. Journal of Network and Computer Applications, 34(4):1043–1053, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510001049.

Zhu:2011:RPS

[653] Yuqing Zhu, Jianmin Wang, and Chaokun Wang. Ripple: a publish/subscribe service for multidata item updates propagation in the cloud. Journal of Network and Computer Applications, 34(4):1054– 1067, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001001.

Kong:2011:EDT

[654] Xiangzhen Kong, Chuang Lin, Yixin Jiang, Wei Yan, and Xiaowen Chu. Efficient dynamic task scheduling in virtualized data centers with fuzzy prediction. Journal of Network and Computer Applications, 34(4):1068–1077, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510000998.

Shi:2011:SSG

[655] Weidong Shi, Yang Lu, Zhu Li, and Jonathan Engelsma. SHARC: a scalable 3D graphics virtual appliance delivery framework in cloud. Journal of Network and Computer Applications, 34(4):1078–1087, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510001037.

Jin:2011:OLM

[656] Hai Jin, Wei Gao, Song Wu, Xuanhua Shi, Xiaoxin Wu, and Fan Zhou. Optimizing the live migration of virtual machine by CPU scheduling. Journal of Network and Computer Applications, 34(4):1088– 1096, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001116.

Chonka:2011:CSD

[657] Ashley Chonka, Yang Xiang, Wanlei Zhou, and Alessio Bonti. Cloud security defence to protect cloud computing against HTTP-DoS and XML-DoS attacks. Journal of Network and Computer Applications, 34(4): 1097–1107, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001025.

Shen:2011:LVM

[658] Yanming Shen, Keqiu Li, Junfeng Xu, and Lu Li. Layered video multicast with a P2P cooperation approach. Journal of Network and Computer Applications, 34(4):1108–1112, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510001013.

Lombardi:2011:SVC

[659] Flavio Lombardi and Roberto Di Pietro. Secure virtualization for cloud computing. Journal of Network and Computer Applications, 34(4):1113– 1122, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001062.

Semchedine:2011:TAP

[660] Fouzi Semchedine, Louiza Bouallouche-Medjkoune, and Djamil Aïssani. Task assignment policies in distributed server systems: a survey. Journal of Network and Computer Applications, 34(4):1123–1130, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000282.

Chen:2011:EAA

[661] Tien-Ho Chen, Yen-Chiu Chen, Wei-Kuan Shih, and Hsin-Wen Wei. An efficient anonymous authentication protocol for mobile pay-TV. Journal of Network and Computer Applications, 34(4):1131–1137, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510002031.

Wang:2011:BTR

[662] Jian Wang, Yanheng Liu, and Yu Jiao. Building a trusted route in a mobile ad hoc network considering communication reliability and path length. Journal of Network and Computer Applications, 34(4):1138– 1149, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002055.

Hu:2011:SID

[663] J. Hu, I. Khalil, S. Han, and A. Mahmood. Seamless integration of dependability and security concepts in SOA: a feedback control system based framework and taxonomy. Journal of Network and Computer Applications, 34(4):1150–1159, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510002110.

Kim:2011:LBA

[664] Hyun Sung Kim. Location-based authentication protocol for first cognitive radio networking standard. Journal of Network and Computer Applications, 34(4):1160–1167, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510002286.

Rosado:2011:SDS

[665] David G. Rosado, Eduardo Fernández-Medina, Javier López, and Mario Piattini. Systematic design of secure Mobile Grid systems. Journal of Network and Computer Applications, 34(4):1168–1183, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000026.

Amiri:2011:MIB

[666] Fatemeh Amiri, MohammadMahdi Rezaei Yousefi, Caro Lucas, Azadeh Shakery, and Nasser Yazdani. Mutual information-based feature selection for intrusion detection systems. Journal of Network and Computer Applications, 34(4):1184–1199, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000038.

Duque:2011:MBF

[667] Rafael Duque, Crescencio Bravo, and Manuel Ortega. A model-based framework to automate the analysis of users’ activity in collaborative systems. Journal of Network and Computer Applications, 34(4):1200– 1209, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000221.

Liu:2011:SPB

[668] Meirong Liu, Timo Koskela, Zhonghong Ou, Jiehan Zhou, Jukka Riekki, and Mika Ylianttila. Super-peer-based coordinated service provision. Journal of Network and Computer Applications, 34(4):1210– 1224, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000245.

Chen:2011:SQS

[669] Jaime Chen, Manuel Díaz, Luis Llopis, Bartolomé Rubio, and José M. Troya. A survey on quality of service support in wireless sensor and actor networks: Requirements and challenges in the context of critical infrastructure protection. Journal of Network and Computer Applications, 34(4):1225–1239, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000257.

Tenruh:2011:MSR

[670] Mahmut Tenruh. Message scheduling with reduced matrix cycle and evenly distributed sparse allocation for time-triggered CAN. Journal of Network and Computer Applications, 34(4):1240–1251, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000269.

Bakhsh:2011:DCC

[671] Sheikh Tahir Bakhsh, Halabi Hasbullah, and Sabeen Tahir. Dynamic Congestion Control through backup relay in Bluetooth scatternet. Journal of Network and Computer Applications, 34(4):1252– 1262, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000270.

Liu:2011:BMC

[672] Xingwei Liu, Xuming Fang, Xu Chen, and Xuesong Peng. A bidding model and cooperative game-based vertical handoff decision algorithm. Journal of Network and Computer Applications, 34(4):1263– 1271, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000294.

Lu:2011:GMS

[673] Qi Lu and Maode Ma. Group mobility support in mobile WiMAX networks. Journal of Network and Computer Applications, 34(4):1272– 1282, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000312.

Ibanez:2011:MAS

[674] Jesús Ibáñez. Minimalist approach to show emotions via a flock of smileys. Journal of Network and Computer Applications, 34(4):1283– 1291, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000324.

Huang:2011:UOT

[675] Chun-Ying Huang, Shang-Pin Ma, and Kuan-Ta Chen. Using one-time passwords to prevent password phishing attacks. Journal of Network and Computer Applications, 34(4):1292–1301, July 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000427.

Xie:2011:ADW

[676] Miao Xie, Song Han, Biming Tian, and Sazia Parvin. Anomaly detection in wireless sensor networks: a survey. Journal of Network and Computer Applications, 34(4):1302–1325, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000580.

Senol:2011:IAA

[677] Sinan Senol, Kemal Leblebicioglu, and Ece G. Schmidt. INtERCEDE: an algorithmic approach to networked control system design. Journal of Network and Computer Applications, 34(4):1326–1341, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000725.

Teigao:2011:AUC

[678] Rafael Teigão, Carlos Maziero, and Altair Santin. Applying a usage control model in an operating system kernel. Journal of Network and Computer Applications, 34(4):1342–1352, July 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000737.

Guo:2011:JOP

[679] Songtao Guo, Chuangyin Dang, and Xiaofeng Liao. Joint opportunistic power and rate allocation for wireless ad hoc networks: an adaptive particle swarm optimization approach. Journal of Network and Computer Applications, 34(4):1353–1365, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000749.

Dean:2011:CSM

[680] Justin Dean, Andrew Harrison, Robert N. Lass, Joe Macker, David Millar, and Ian Taylor. Client/server messaging protocols in serverless environments. Journal of Network and Computer Applications, 34(4):1366– 1379, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000750.

Challal:2011:SED

[681] Y. Challal, A. Ouadjaout, N. Lasla, M. Bagaa, and A. Hadjidj. Secure and efficient disjoint multipath construction for fault tolerant routing in wireless sensor networks. Journal of Network and Computer Applications, 34(4):1380–1397, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000762.

Li:2011:SBC

[682] Wenfeng Li and Weiming Shen. Swarm behavior control of mobile multirobots with wireless sensor networks. Journal of Network and Computer Applications, 34(4):1398–1407, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000774.

Ocaya:2011:FCR

[683] R. O. Ocaya. A framework for collaborative remote experimentation for a physical laboratory using a low cost embedded web server. Journal of Network and Computer Applications, 34(4):1408– 1415, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000786.

Xu:2011:HCC

[684] Wenjun Xu, Zude Zhou, D. T. Pham, C. Ji, M. Yang, and Quan Liu. Hybrid congestion control for high-speed networks. Journal of Network and Computer Applications, 34(4):1416–1428, July 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000798.

Torkestani:2011:LSB

[685] Javad Akbari Torkestani and Mohammad Reza Meybodi. A link stability-based multicast routing protocol for wireless mobile ad hoc networks. Journal of Network and Computer Applications, 34(4):1429– 1440, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000804.

Calvinho:2011:EFI

[686] André Vicente Calvinho and Rui Gustavo Crespo. Email FI identification and resolution with model checking. Journal of Network and Computer Applications, 34(4):1441–1446, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000816.

Anonymous:2011:EBd

[687] Anonymous. Editorial board. Journal of Network and Computer Applications, 34(4):ifc, July 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000956.

Chao:2011:DMC

[688] Han-Chieh Chao, Jean-Pierre Seifert, Shiguo Lian, and Liang Zhou. Dependable multimedia communications: Systems, services, and applications. Journal of Network and Computer Applications, 34(5):1447–1448, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001251.

Zhu:2011:IRC

[689] Rongbo Zhu. Intelligent rate control for supporting real-time traffic in WLAN mesh networks. Journal of Network and Computer Applications, 34(5):1449–1458, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001633.

Shih:2011:RSS

[690] Po-Yi Shih, Po-Chuan Lin, Jhing-Fa Wang, and Yuan-Ning Lin. Robust several-speaker speech recognition with highly dependable online speaker adaptation and identification. Journal of Network and Computer Applications, 34(5):1459–1467, September 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001554.

Deng:2011:PPE

[691] Der-Jiunn Deng. PSSB: Priority enforced slow-start backoff algorithm for multimedia transmission in wireless ad-hoc networks. Journal of Network and Computer Applications, 34(5):1468–1473, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001104.

Leroux:2011:MTS

[692] Philip Leroux, Steven Latré, Nicolas Staelens, Piet Demeester, and Filip De Turck. Mobile TV services through IP Datacast over DVB-H: Dependability of the quality of experience on the IP-based distribution network quality of service. Journal of Network and Computer Applications, 34(5):1474–1488, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001736.

Ji:2011:PML

[693] Wen Ji, Min Chen, Xiaohu Ge, Peng Li, and Yiqiang Chen. A perceptual macroblock layer power control for energy scalable video encoder based on just noticeable distortion principle. Journal of Network and Computer Applications, 34(5):1489–1497, September 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001098.

Dhurandher:2011:UBA

[694] Sanjay K. Dhurandher, Sudip Misra, Puneet Pruthi, Shubham Singhal, Saurabh Aggarwal, and Isaac Woungang. Using bee algorithm for peerto-peer file searching in mobile ad hoc networks. Journal of Network and Computer Applications, 34(5):1498–1508, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001797.

Liao:2011:NSP

[695] I.-Ju Liao, Tin-Yu Wu, Jen-Wen Ding, and Wei-Tsong Lee. A novel 3D streaming protocol supported multi-mode display. Journal of Network and Computer Applications, 34(5):1509–1517, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001670.

Mehmood:2011:MAM

[696] Rashid Mehmood, Raad Alturki, and Sherali Zeadally. Multimedia applications over metropolitan area networks (MANs). Journal of Network and Computer Applications, 34(5):1518–1529, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001414.

Misra:2011:PCS

[697] Sudip Misra and Ankur Jain. Policy controlled self-configuration in unattended wireless sensor networks. Journal of Network and Computer Applications, 34(5):1530–1544, September 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510001244.

Wang:2011:DPP

[698] Chia-Hui Wang and Yu-Shun Liu. A dependable privacy protection for end-to-end VoIP via Elliptic-Curve Diffie–Hellman and dynamic key changes. Journal of Network and Computer Applications, 34(5):1545– 1556, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001803.

Liu:2011:PIA

[699] Guangjie Liu, Junwen Wang, Shiguo Lian, and Zhiquan Wang. A passive image authentication scheme for detecting region-duplication forgery with rotation. Journal of Network and Computer Applications, 34(5):1557–1565, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001621.

Chen:2011:EHM

[700] Yeong-Sheng Chen, Yuan-Wei Lee, and Jong Hyuk Park. Enhanced HCCA mechanism for multimedia traffics with QoS support in IEEE 802.11e networks. Journal of Network and Computer Applications, 34 (5):1566–1571, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510001608.

Ma:2011:SSI

[701] Kevin J. Ma, Radim Bartos, and Swapnil Bhatia. A survey of schemes for Internet-based video delivery. Journal of Network and Computer Applications, 34(5):1572–1586, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000300.

Tekbiyik:2011:EEW

[702] Neyre Tekbiyik and Elif Uysal-Biyikoglu. Energy efficient wireless unicast routing alternatives for machine-to-machine networks. Journal of Network and Computer Applications, 34(5):1587–1614, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000439.

Mao:2011:AMI

[703] Ching-Hao Mao, Hahn-Ming Lee, and Che-Fu Yeh. Adaptive e-mails intention finding system based on words social networks. Journal of Network and Computer Applications, 34(5):1615–1622, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000841.

Yang:2011:OFS

[704] Jun Yang, Bin Dai, Benxiong Huang, and Shui Yu. Orthogonal feedback scheme for network coding. Journal of Network and Computer Applications, 34(5):1623–1633, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000890.

Espada:2011:MAC

[705] Jordán Pascual Espada, Oscar Sanjuán Martínez, Juan Manuel Cueva Lovelle, B. Cristina Pelayo G-Bustelo, Manuel Álvarez Álvarez, and Alejandro González García. Modeling architecture for collaborative virtual objects based on services. Journal of Network and Computer Applications, 34(5):1634–1647, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000907.

Chen:2011:FQP

[706] Ting Chen, Xiao song Zhang, Hua Li, Xiong da Li, and Yue Wu. Fast quarantining of proactive worms in unstructured P2P networks. Journal of Network and Computer Applications, 34(5):1648–1659, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000919.

Prodhan:2011:TBR

[707] Anindya Tahsin Prodhan, Rajkumar Das, Humayun Kabir, and Gholamali C. Shoja. TTL based routing in opportunistic networks. Journal of Network and Computer Applications, 34(5):1660–1670, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451100107X.

Jawhar:2011:LWS

[708] Imad Jawhar, Nader Mohamed, and Dharma P. Agrawal. Linear wireless sensor networks: Classification and applications. Journal of Network and Computer Applications, 34(5):1671–1682, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001081.

Fernandes:2011:APA

[709] Stenio Fernandes, Judith Kelner, and Djamel Sadok. An adaptivepredictive architecture for video streaming servers. Journal of Network and Computer Applications, 34(5):1683–1694, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001093.

Zhang:2011:ESB

[710] Liqiang Zhang, Yu-Jen Cheng, and Xiaobo Zhou. Enhanced statisticsbased rate adaptation for 802.11 wireless networks. Journal of Network and Computer Applications, 34(5):1695–1706, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001147.

Kuehnhausen:2011:AJM

[711] Martin Kuehnhausen and Victor S. Frost. Application of the Java Message Service in mobile monitoring environments. Journal of Network and Computer Applications, 34(5):1707–1716, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001159.

Sadan:2011:SNA

[712] Zac Sadan and David G. Schwartz. Social network analysis of web links to eliminate false positives in collaborative anti-spam systems. Journal of Network and Computer Applications, 34(5):1717–1723, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001160.

Hsu:2011:NIB

[713] Chien-Lung Hsu and Han-Yu Lin. New identity-based key-insulated convertible multi-authenticated encryption scheme. Journal of Network and Computer Applications, 34(5):1724–1731, September 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001172.

Qin:2011:MAN

[714] Tao Qin, Xiaohong Guan, Wei Li, Pinghui Wang, and Qiuzhen Huang. Monitoring abnormal network traffic based on blind source separation approach. Journal of Network and Computer Applications, 34 (5):1732–1742, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001184.

Ma:2011:CGB

[715] Hwa-Chun Ma, Prasan Kumar Sahoo, and Yen-Wen Chen. Computational geometry based distributed coverage hole detection protocol for the wireless sensor networks. Journal of Network and Computer Applications, 34(5):1743–1756, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001196.

Niu:2011:CCA

[716] Wenjia Niu, Gang Li, Hui Tang, Xu Zhou, and Zhongzhi Shi. CARSA: a context-aware reasoning-based service agent model for AI planning of web service composition. Journal of Network and Computer Applications, 34(5):1757–1770, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001202.

Anonymous:2011:EBe

[717] Anonymous. Editorial board. Journal of Network and Computer Applications, 34(5):ifc, September 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001354.

Chen:2011:COW

[718] Jiming Chen and David K. Y. Yau. Control and optimization over wireless networks. Journal of Network and Computer Applications, 34 (6):1771–1772, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001263.

Weng:2011:IQM

[719] Ning Weng, I-Hung Li, and Lucas Vespa. Information quality model and optimization for 802.15.4-based wireless sensor networks. Journal of Network and Computer Applications, 34(6):1773–1783, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510002249.

Kosar:2011:LES

[720] Rabun Kosar, Ilir Bojaxhiu, Ertan Onur, and Cem Ersoy. Lifetime extension for surveillance wireless sensor networks with intelligent redeployment. Journal of Network and Computer Applications, 34(6):1784– 1793, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510002213.

Abdou:2011:UEA

[721] Wahabou Abdou, Adrien Henriet, Christelle Bloch, Dominique Dhoutaut, Damien Charlet, and François Spies. Using an evolutionary algorithm to optimize the broadcasting methods in mobile ad hoc networks. Journal of Network and Computer Applications, 34(6):1794–1804, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000075.

Nguyen:2011:DAM

[722] Dang Tu Nguyen, Wook Choi, Minh Thiep Ha, and Hyunseung Choo. Design and analysis of a multi-candidate selection scheme for greedy routing in wireless sensor networks. Journal of Network and Computer Applications, 34(6):1805–1817, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804510002201.

Liu:2011:OCM

[723] Wenjia Liu and Bo Chen. Optimal control of mobile monitoring agents in immune-inspired wireless monitoring networks. Journal of Network and Computer Applications, 34(6):1818–1826, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510002158.

Chen:2011:CBS

[724] Qing Chen, Zubair Md. Fadlullah, Xiaodong Lin, and Nei Kato. A clique-based secure admission control scheme for mobile ad hoc networks (MANETs). Journal of Network and Computer Applications, 34 (6):1827–1835, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510002237.

Al-Manthari:2011:CPB

[725] Bader Al-Manthari, Nidal Nasser, Najah Abu Ali, and Hossam Hassanein. Congestion prevention in broadband wireless access systems: an economic approach. Journal of Network and Computer Applications, 34(6):1836–1847, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804510002122.

Ryoo:2011:MSM

[726] Jihoon Ryoo and Hwangnam Kim. Multi-sector multi-range control for self-organizing wireless networks. Journal of Network and Computer Applications, 34(6):1848–1860, November 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804510002134.

Lu:2011:CLR

[727] Zhaoming Lu, Yan Yang, Xiangming Wen, Ying Ju, and Wei Zheng. A cross-layer resource allocation scheme for ICIC in LTE-Advanced. Journal of Network and Computer Applications, 34(6):1861–1868, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000063.

Jahanshahi:2011:MFJ

[728] M. Jahanshahi, M. Dehghan, and M. R. Meybodi. A mathematical formulation for joint channel assignment and multicast routing in multichannel multi-radio wireless mesh networks. Journal of Network and Computer Applications, 34(6):1869–1882, November 2011. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000051.

Rodriguez-Covili:2011:CIE

[729] Juan Rodríguez-Covili, Sergio F. Ochoa, José A. Pino, Roc Messeguer, Esunly Medina, and Dolors Royo. A communication infrastructure to ease the development of mobile collaborative applications. Journal of Network and Computer Applications, 34(6):1883–1893, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804510002250.

Li:2011:NPP

[730] Bin Li, Zheng Zhou, Dejian Li, and Weixia Zou. A novel Parzen probabilistic neural network based noncoherent detection algorithm for distributed ultra-wideband sensors. Journal of Network and Computer Applications, 34(6):1894–1902, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804510002262.

Han:2011:PSV

[731] Kyusuk Han, Youngjoon Seo, Sungjune Yoon, James J. (Jong Hyuk) Park, and Taeshik Shon. Providing security vertical handoff in SARAH for heterogeneous networks. Journal of Network and Computer Applications, 34(6):1903–1907, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000233.

Ayaz:2011:SRT

[732] Muhammad Ayaz, Imran Baig, Azween Abdullah, and Ibrahima Faye. A survey on routing techniques in underwater wireless sensor networks. Journal of Network and Computer Applications, 34(6):1908–1927, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001214.

Pease:2011:CLS

[733] Sarogini Grace Pease, Lin Guan, Iain Phillips, and Alan Grigg. Crosslayer signalling and middleware: a survey for inelastic soft real-time applications in MANETs. Journal of Network and Computer Applications, 34(6):1928–1941, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001305.

Rivas:2011:SVP

[734] David Antolino Rivas, José M. Barceló-Ordinas, Manel Guerrero Zapata, and Julián D. Morillo-Pozo. Security on VANETs: Privacy, misbehaving nodes, false information and secure data aggregation. Journal of Network and Computer Applications, 34(6):1942–1955, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001317.

Iqbal:2011:SNC

[735] Muhammad Azhar Iqbal, Bin Dai, Benxiong Huang, A. Hassan, and Shui Yu. Survey of network coding-aware routing protocols in wireless networks. Journal of Network and Computer Applications, 34 (6):1956–1970, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001500.

Bakhouya:2011:AAI

[736] M. Bakhouya, J. Gaber, and P. Lorenz. An adaptive approach for information dissemination in Vehicular Ad hoc Networks. Journal of Network and Computer Applications, 34(6):1971–1978, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001226.

Leng:2011:DKB

[737] Lu Leng and Jiashu Zhang. Dual-key-binding cancelable palmprint cryptosystem for palmprint protection and information security. Journal of Network and Computer Applications, 34(6):1979–1989, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001287.

Deng:2011:MCA

[738] Jeremiah D. Deng and Martin K. Purvis. Multi-core application performance optimization using a constrained tandem queueing model. Journal of Network and Computer Applications, 34(6):1990–1996, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001299.

Wightman:2011:FSD

[739] Pedro M. Wightman and Miguel A. Labrador. A family of simple distributed minimum connected dominating set-based topology construction algorithms. Journal of Network and Computer Applications, 34 (6):1997–2010, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001457.

Biswash:2011:EMB

[740] Sanjay Kumar Biswash and Chiranjeev Kumar. An efficient metricbased (EM-B) location management scheme for wireless cellular networks. Journal of Network and Computer Applications, 34(6):2011–2026, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001470.

Huang:2011:NDH

[741] Tzu-Chi Huang and Kuo-Chih Chu. Networking without Dynamic Host Configuration Protocol server in Ethernet and Wireless Local Area Network. Journal of Network and Computer Applications, 34 (6):2027–2041, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001482.

Sharbafi:2011:INA

[742] Maziar Ahmad Sharbafi and Mohammad Javad Yazdanpanah. IDFC: a new approach to control bifurcation in TCP/RED. Journal of Network and Computer Applications, 34(6):2042–2050, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001494.

Anonymous:2011:EBf

[743] Anonymous. Editorial board. Journal of Network and Computer Applications, 34(6):ifc, November 2011. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001603.

Anonymous:2012:CLC

[744] Anonymous. Contents list continued. Journal of Network and Computer Applications, 35(1):II, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002220.

Shen:2012:CCA

[745] Weiming Shen, Yinsheng Li, and Anne James. Collaborative computing and applications. Journal of Network and Computer Applications, 35(1):1–2, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451100124X.

Paredes:2012:SIR

[746] Hugo Paredes and F. Mario Martins. Social interaction regulation in virtual web environments using the Social Theatres model. Journal of Network and Computer Applications, 35(1):3–19, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000476.

Zhang:2012:OMP

[747] Weizhan Zhang, Qinghua Zheng, Haifei Li, and Feng Tian. An overlay multicast protocol for live streaming and delay-guaranteed interactive media. Journal of Network and Computer Applications, 35(1):20– 28, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000518.

Liu:2012:AAQ

[748] Min Liu, Shijun Xu, and Siyi Sun. An agent-assisted QoS-based routing algorithm for wireless sensor networks. Journal of Network and Computer Applications, 35(1):29–36, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000853.

Srisooksai:2012:PDC

[749] Tossaporn Srisooksai, Kamol Keamarungsi, Poonlap Lamsrichan, and Kiyomichi Araki. Practical data compression in wireless sensor networks: a survey. Journal of Network and Computer Applications, 35(1): 37–59, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000555.

Luo:2012:ICB

[750] Junzhou Luo, Xiaogang Wang, and Ming Yang. An interval centroid based spread spectrum watermarking scheme for multi-flow traceback. Journal of Network and Computer Applications, 35(1):60–71, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000579.

Shin:2012:WPP

[751] Kyuyong Shin and Douglas S. Reeves. Winnowing: Protecting P2P systems against pollution through cooperative index filtering. Journal of Network and Computer Applications, 35(1):72–84, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000531.

Gao:2012:PRR

[752] Guoqiang Gao, Ruixuan Li, Kunmei Wen, and Xiwu Gu. Proactive replication for rare objects in unstructured peer-to-peer networks. Journal of Network and Computer Applications, 35(1):85–96, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000452.

Kim:2012:MDW

[753] Kwanghoon Kim. A model-driven workflow fragmentation framework for collaborative workflow architectures and systems. Journal of Network and Computer Applications, 35(1):97–110, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451100083X.

Du:2012:CSD

[754] Junpeng Du, Shikai Jing, and Jihong Liu. Creating shared design thinking process for collaborative design. Journal of Network and Computer Applications, 35(1):111–120, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451100052X.

Xue:2012:FBO

[755] Yunjiao Xue, Hamada H. Ghenniwa, and Weiming Shen. Frame-based ontological view for semantic integration. Journal of Network and Computer Applications, 35(1):121–131, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804511000488.

Sun:2012:OBI

[756] Hongbo Sun, Wenhui Fan, Weiming Shen, and Tianyuan Xiao. Ontologybased interoperation model of collaborative product development. Journal of Network and Computer Applications, 35(1):132–144, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511000506.

Rodriguez-Covili:2012:HLM

[757] Juan Rodríguez-Covili, Sergio F. Ochoa, and José A. Pino. High level MANET protocol: Enhancing the communication support for mobile collaborative work. Journal of Network and Computer Applications, 35(1):145–155, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000440.

Sun:2012:SMC

[758] Yuqing Sun, Dickson K. W. Chiu, Bin Gong, Xiangxu Meng, and Peng Zhang. Scheduling mobile collaborating workforce for multiple urgent events. Journal of Network and Computer Applications, 35(1):156– 163, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000464.

Enembreck:2012:DCO

[759] Fabrício Enembreck and Jean-Paul André Barthès. Distributed constraint optimization with MULBS: a case study on collaborative meeting scheduling. Journal of Network and Computer Applications, 35(1):164– 175, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000543.

Perez-Sanagustin:2012:DCT

[760] Mar Pérez-Sanagustín, Gustavo Ramirez-Gonzalez, Davinia Hernández-Leo, Mario Muñoz-Organero, Patricia Santos, Josep Blat, and Carlos Delgado Kloos. Discovering the campus together: a mobile and computer-based learning experience. Journal of Network and Computer Applications, 35(1):176–188, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451100049X.

Vivacqua:2012:TAC

[761] Adriana S. Vivacqua and Marcos R. S. Borges. Taking advantage of collective knowledge in emergency response systems. Journal of Network and Computer Applications, 35(1):189–198, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000567.

Talaei-Khoei:2012:FAM

[762] Amir Talaei-Khoei, Pradeep Ray, Nandan Parameshwaran, and Lundy Lewis. A framework for awareness maintenance. Journal of Network and Computer Applications, 35(1):199–210, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001238.

Al-Jaroodi:2012:SOM

[763] Jameela Al-Jaroodi and Nader Mohamed. Service-oriented middleware: a survey. Journal of Network and Computer Applications, 35(1):211– 220, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001512.

Biradar:2012:RMR

[764] Rajashekhar C. Biradar and Sunilkumar S. Manvi. Review of multicast routing mechanisms in mobile ad hoc networks. Journal of Network and Computer Applications, 35(1):221–239, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001573.

Che:2012:PLT

[765] Xianhui Che and Barry Ip. Packet-level traffic analysis of online games from the genre characteristics perspective. Journal of Network and Computer Applications, 35(1):240–252, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001731.

Yau:2012:RLC

[766] Kok-Lim Alvin Yau, Peter Komisarczuk, and Paul D. Teal. Reinforcement learning for context awareness and intelligence in wireless networks: Review, new features and open issues. Journal of Network and Computer Applications, 35(1):253–267, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001755.

Omar:2012:CBT

[767] Mawloud Omar, Yacine Challal, and Abdelmadjid Bouabdallah. Certification-based trust models in mobile ad hoc networks: a survey and taxonomy. Journal of Network and Computer Applications, 35(1):268–286, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001767.

Stergiou:2012:PEB

[768] Eleftherios Stergiou and John Garofalakis. Performance estimation of banyan semi layer networks with drop resolution mechanism. Journal of Network and Computer Applications, 35(1):287–294, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001780.

Al-Surmi:2012:MMI

[769] Ibrahim Al-Surmi, Mohamed Othman, and Borhanuddin Mohd Ali. Mobility management for IP-based next generation mobile networks: Review, challenge and perspective. Journal of Network and Computer Applications, 35(1):295–315, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511001792.

Zhong:2012:ALB

[770] Feng Zhong, Chai Kiat Yeo, and Bu Sung Lee. Adaptive load balancing algorithm for multiple homing mobile nodes. Journal of Network and Computer Applications, 35(1):316–327, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001469.

Barcellos:2012:BNS

[771] Marinho P. Barcellos, Rodolfo S. Antunes, Hisham H. Muhammad, and Ruthiano S. Munaretti. Beyond network simulators: Fostering novel distributed applications and protocols through extendible design. Journal of Network and Computer Applications, 35(1):328–339, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001536.

Li:2012:IBO

[772] Fagen Li, Muhammad Khurram Khan, Khaled Alghathbar, and Tsuyoshi Takagi. Identity-based online/offline signcryption for low power devices. Journal of Network and Computer Applications, 35(1):340–347, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451100155X.

Saad:2012:UBA

[773] Walid Saad, Zaher Dawy, and Sanaa Sharafeddine. A utility-based algorithm for joint uplink/downlink scheduling in wireless cellular networks. Journal of Network and Computer Applications, 35(1):348– 356, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001561.

Nie:2012:IBR

[774] Jing Nie. IPv6 based relay gateway protocol in hybrid networks. Journal of Network and Computer Applications, 35(1):357–364, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001585.

Al-Shammary:2012:RAS

[775] Dhiah Al-Shammary and Ibrahim Khalil. Redundancy-aware SOAP messages compression and aggregation for enhanced performance. Journal of Network and Computer Applications, 35(1):365–381, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451100172X.

Canete:2012:NNN

[776] Eduardo Cañete, Jaime Chen, R. Marcos Luque, and Bartolomé Rubio. NeuralSens: a neural network based framework to allow dynamic adaptation in wireless sensor and actor networks. Journal of Network and Computer Applications, 35(1):382–393, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001743.

Rostami:2012:MAM

[777] Mohammad Javad Rostami, Azadeh Alsadat Emrani Zarandi, and Seyed Mohamad Hoseininasab. MSDP with ACO: a maximal SRLG disjoint routing algorithm based on ant colony optimization. Journal of Network and Computer Applications, 35(1):394–402, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001779.

Beaubrun:2012:MAD

[778] Ronald Beaubrun, Jhon-Fredy Llano-Ruiz, Benoit Poirier, and Alejandro Quintero. A middleware architecture for disseminating delayconstrained information in wireless sensor networks. Journal of Network and Computer Applications, 35(1):403–411, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001809.

Balasangameshwara:2012:HPF

[779] Jasma Balasangameshwara and Nedunchezhian Raju. A hybrid policy for fault tolerant load balancing in grid computing environments. Journal of Network and Computer Applications, 35(1):412–422, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001834.

Huang:2012:PUD

[780] Tzu-Chi Huang. Program Ultra-Dispatcher for launching applications in a customization manner on cloud computing. Journal of Network and Computer Applications, 35(1):423–446, January 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001846.

Chang:2012:HCM

[781] Chao-Tsun Chang. Hash caching mechanism in source-based routing for wireless ad hoc networks. Journal of Network and Computer Applications, 35(1):447–458, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001858.

Zare:2012:FSC

[782] Sajjad Zare and Akbar Ghaffarpour Rahbar. An FEC scheme combined with weighted scheduling to reduce multicast packet loss in IPTV over PON. Journal of Network and Computer Applications, 35(1):459– 468, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451100186X.

Huang:2012:RMQ

[783] Jun Huang, Xiaohong Huang, and Yan Ma. Routing with multiple quality-of-services constraints: an approximation perspective. Journal of Network and Computer Applications, 35(1):469–479, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001871.

Bahtiyar:2012:ETI

[784] Serif Bahtiyar and Mehmet Ufuk Çaglayan. Extracting trust information from security system of a service. Journal of Network and Computer Applications, 35(1):480–490, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511001895.

Blasco:2012:FAS

[785] Jorge Blasco, Julio Cesar Hernandez-Castro, José María de Fuentes, and Benjamín Ramos. A framework for avoiding steganography usage over HTTP. Journal of Network and Computer Applications, 35(1):491– 501, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001901.

Hu:2012:LBT

[786] Bing Hu, Kwan L. Yeung, and Zhaoyang Zhang. Load-balanced threestage switch. Journal of Network and Computer Applications, 35(1):502– 509, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001986.

Krikidis:2012:MTW

[787] Ioannis Krikidis and John S. Thompson. MIMO two-way relay channel with superposition coding and imperfect channel estimation. Journal of Network and Computer Applications, 35(1):510–516, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511002001.

Anonymous:2012:EBa

[788] Anonymous. Editorial board. Journal of Network and Computer Applications, 35(1):ifc, January 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451100213X.

Gunes:2012:MMD

[789] Mesut Günes and Yan Zhang. In memory of Mieso Denko. Journal of Network and Computer Applications, 35(2):517, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001810.

Zhou:2012:CMW

[790] Ming-Tuo Zhou and Hiroshi Harada. Cognitive maritime wireless mesh/ad hoc networks. Journal of Network and Computer Applications, 35(2):518–526, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451100004X.

Lambrou:2012:TCC

[791] Theofanis P. Lambrou and Christos G. Panayiotou. A testbed for coverage control using mixed wireless sensor networks. Journal of Network and Computer Applications, 35(2):527–537, March 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001123.

Fischer:2012:WMS

[792] Joachim Fischer, Jens-Peter Redlich, Jochen Zschau, Claus Milkereit, Matteo Picozzi, Kevin Fleming, Mihal Brumbulli, Björn Lichtblau, and Ingmar Eveslage. A wireless mesh sensing network for early warning. Journal of Network and Computer Applications, 35(2):538– 547, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001548.

Theodoridis:2012:PPW

[793] Evangelos Theodoridis, Ioannis Chatzigiannakis, and Stefan Dulman. Post-processing in wireless sensor networks: Benchmarking sensor trace files for in-network data aggregation. Journal of Network and Computer Applications, 35(2):548–561, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511000087.

Corredor:2012:KAS

[794] Iván Corredor, José F. Martínez, Miguel S. Familiar, and Lourdes López. Knowledge-aware and Service-Oriented Middleware for deploying pervasive services. Journal of Network and Computer Applications, 35(2):562–576, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001111.

Yang:2012:SSS

[795] Bo Yang and Manohar Mareboyana. Similarity search in sensor networks using semantic-based caching. Journal of Network and Computer Applications, 35(2):577–583, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451100110X.

Domingo:2012:OIT

[796] Mari Carmen Domingo. An overview of the Internet of Things for people with disabilities. Journal of Network and Computer Applications, 35(2):584–596, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002025.

Rizvi:2012:AEE

[797] Sajjad Rizvi, Hassaan Khaliq Qureshi, Syed Ali Khayam, Veselin Rakocevic, and Muttukrishnan Rajarajan. A1: an energy efficient topology control algorithm for connected area coverage in wireless sensor networks. Journal of Network and Computer Applications, 35(2):597– 605, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002062.

Sengupta:2012:CWR

[798] Malabika Sengupta, Swapan Kumar Mondal, and Debashis Saha. A comparison of wavelength reservation protocols for WDM optical networks. Journal of Network and Computer Applications, 35(2):606– 618, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002293.

Zhu:2012:SCC

[799] Chuan Zhu, Chunlin Zheng, Lei Shu, and Guangjie Han. A survey on coverage and connectivity issues in wireless sensor networks. Journal of Network and Computer Applications, 35(2):619–632, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002323.

Wang:2012:OOS

[800] Jingyu Wang, Jianxin Liao, and Tonghong Li. OSIA: Out-of-order Scheduling for In-order Arriving in concurrent multi-path transfer. Journal of Network and Computer Applications, 35(2):633–643, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001822.

Sharafeddine:2012:SFA

[801] Sanaa Sharafeddine, Ibrahim Al-Kassem, and Zaher Dawy. A scatternet formation algorithm for Bluetooth networks with a non-uniform distribution of devices. Journal of Network and Computer Applications, 35(2):644–656, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001913.

Zhang:2012:TDS

[802] Wenjie Zhang and Chai Kiat Yeo. Throughput and delay scaling laws for mobile overlaid wireless networks. Journal of Network and Computer Applications, 35(2):657–667, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511001998.

Hassan:2012:CRM

[803] Hoda Hassan and Mohamed Eltoweissy. CORM: a reference model for future computer networks. Journal of Network and Computer Applications, 35(2):668–680, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002037.

Cordeschi:2012:STE

[804] Nicola Cordeschi, Tatiana Patriarca, and Enzo Baccarelli. Stochastic traffic engineering for real-time applications over wireless networks. Journal of Network and Computer Applications, 35(2):681– 694, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002049.

Jurado:2012:UFL

[805] Francisco Jurado, Miguel A. Redondo, and Manuel Ortega. Using fuzzy logic applied to software metrics and test cases to assess programming assignments and give advice. Journal of Network and Computer Applications, 35(2):695–712, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002050.

Vakili:2012:CCP

[806] Golnaz Vakili and Siavash Khorsandi. Coordination of cooperation policies in a peer-to-peer system using swarm-based RL. Journal of Network and Computer Applications, 35(2):713–722, March 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002074.

Burget:2012:TAF

[807] Radim Burget, Dan Komosny, and Kathiravelu Ganeshan. Topology aware feedback transmission for real-time control protocol. Journal of Network and Computer Applications, 35(2):723–730, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511002086.

Sedighizad:2012:MBM

[808] Mahboobeh Sedighizad, Babak Seyfe, and Keivan Navaie. MR-BART: Multi-Rate Available Bandwidth Estimation in Real-Time. Journal of Network and Computer Applications, 35(2):731–742, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511002098.

Zhu:2012:AEF

[809] Yi hua Zhu, Hui Xu, Kai kai Chi, and Hua Hu. Accumulating error-free frame blocks to improve throughput for IEEE 802.11-based WLAN. Journal of Network and Computer Applications, 35(2):743– 752, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002104.

Calafate:2012:ERC

[810] Carlos T. Calafate, Giancarlo Fortino, Sascha Fritsch, Janio Monteiro, Juan-Carlos Cano, and Pietro Manzoni. An efficient and robust content delivery solution for IEEE 802.11p vehicular environments. Journal of Network and Computer Applications, 35(2):753– 762, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002116.

Li:2012:ESD

[811] Xiong Li, Yongping Xiong, Jian Ma, and Wendong Wang. An efficient and security dynamic identity based authentication protocol for multiserver architecture using smart cards. Journal of Network and Computer Applications, 35(2):763–769, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511002244.

Wang:2012:SFB

[812] Min-Feng Wang, Meng-Feng Tsai, Sie-Long Jheng, and Cheng-Hsien Tang. Social feature-based enterprise email classification without examining email contents. Journal of Network and Computer Applications, 35(2):770–777, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002256.

Dargie:2012:MAC

[813] Waltenegus Dargie. A medium access control protocol that supports a seamless handover in wireless sensor networks. Journal of Network and Computer Applications, 35(2):778–786, March 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002268.

Karunamurthy:2012:NAW

[814] Rajesh Karunamurthy, Ferhat Khendek, and Roch H. Glitho. A novel architecture for Web service composition. Journal of Network and Computer Applications, 35(2):787–802, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451100227X.

Kastaniotis:2012:UDE

[815] Georgios Kastaniotis, Elias Maragos, Christos Douligeris, and Dimitris K. Despotis. Using data envelopment analysis to evaluate the efficiency of web caching object replacement strategies. Journal of Network and Computer Applications, 35(2):803–817, March 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002281.

Liao:2012:PSD

[816] Wen-Hwa Liao and Hung-Chun Yang. A power-saving data storage scheme for wireless sensor networks. Journal of Network and Computer Applications, 35(2):818–825, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451100230X.

Avoine:2012:PFS

[817] Gildas Avoine, Xavier Carpent, and Benjamin Martin. Privacyfriendly synchronized ultralightweight authentication protocols in the storm. Journal of Network and Computer Applications, 35(2):826– 843, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002335.

Imran:2012:LMB

[818] Muhammad Imran, Mohamed Younis, Abas Md Said, and Halabi Hasbullah. Localized motion-based connectivity restoration algorithms for wireless sensor and actor networks. Journal of Network and Computer Applications, 35(2):844–856, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511002347.

Tahiliani:2012:CCA

[819] Mohit P. Tahiliani, K. C. Shet, and T. G. Basavaraju. CARED: Cautious adaptive RED gateways for TCP/IP networks. Journal of Network and Computer Applications, 35(2):857–864, March 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002359.

Anonymous:2012:EBb

[820] Anonymous. Editorial board. Journal of Network and Computer Applications, 35(2):ifc, March 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000033.

Yang:2012:SIT

[821] Laurence T. Yang and Guojun Wang. Special issue on trusted computing and communications. Journal of Network and Computer Applications, 35(3):865–866, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000586.

Yu:2012:TMW

[822] Yanli Yu, Keqiu Li, Wanlei Zhou, and Ping Li. Trust mechanisms in wireless sensor networks: Attack analysis and countermeasures. Journal of Network and Computer Applications, 35(3):867–880, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000592.

Jiang:2012:RRM

[823] Lei Jiang, Lixin Ding, Jianxun Liu, and Jinjun Chen. Reputation rating modeling for open environment lack of communication by using online social cognition. Journal of Network and Computer Applications, 35(3):881–891, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000610.

Qiang:2012:SBI

[824] Weizhong Qiang, Aleksandr Konstantinov, Deqing Zou, and Laurence T. Yang. A standards-based interoperable single sign-on framework in ARC Grid middleware. Journal of Network and Computer Applications, 35(3):892–904, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000609.

Jia:2012:PKD

[825] Zhongtian Jia, Xiaodong Lin, Seng-Hua Tan, Lixiang Li, and Yixian Yang. Public key distribution scheme for delay tolerant networks based on two-channel cryptography. Journal of Network and Computer Applications, 35(3):905–913, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000634.

Wang:2012:SRR

[826] Fei Wang, Furong Wang, Benxiong Huang, and Laurence T. Yang. SONR: a reliable reputation system of self-organized network. Journal of Network and Computer Applications, 35(3):914–926, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000622.

Liu:2012:SPP

[827] Qin Liu, Guojun Wang, and Jie Wu. Secure and privacy preserving keyword searching for cloud storage services. Journal of Network and Computer Applications, 35(3):927–933, May 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000646.

Marmol:2012:TTR

[828] Félix Gómez Mármol and Gregorio Martínez Pérez. TRIP, a trust and reputation infrastructure-based proposal for vehicular ad hoc networks. Journal of Network and Computer Applications, 35(3):934– 941, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000828.

Kharraz:2012:DMR

[829] Mohammad-Amin Kharraz, Hamid Sarbazi-Azad, and Albert Y. Zomaya. On-demand multicast routing protocol with efficient route discovery. Journal of Network and Computer Applications, 35(3):942– 950, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451100066X.

Zhao:2012:IMR

[830] Huanyu Zhao, Xin Yang, and Xiaolin Li. An incentive mechanism to reinforce truthful reports in reputation systems. Journal of Network and Computer Applications, 35(3):951–961, May 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000658.

Zhao:2012:FSS

[831] Laiping Zhao, Yizhi Ren, Mingchu Li, and Kouichi Sakurai. Flexible service selection with user-specific QoS support in service-oriented architecture. Journal of Network and Computer Applications, 35(3): 962–973, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000671.

Gu:2012:SMO

[832] Liang Gu, Guangdong Bai, Yao Guo, Xiangqun Chen, and Hong Mei. Security model oriented attestation on dynamically reconfigurable component-based systems. Journal of Network and Computer Applications, 35(3):974–981, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000683.

Choi:2012:ARS

[833] Min Choi, Jong Hyuk Park, Seungho Lim, and Young-Sik Jeong. Achieving reliable system performance by fast recovery of branch miss prediction. Journal of Network and Computer Applications, 35(3):982– 991, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000695.

Wang:2012:GRE

[834] Yu Wang, Yang Xiang, Wanlei Zhou, and Shunzheng Yu. Generating regular expression signatures for network traffic classification in trusted network management. Journal of Network and Computer Applications, 35(3):992–1000, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000713.

Cho:2012:MAT

[835] Jin-Hee Cho, Ananthram Swami, and Ing-Ray Chen. Modeling and analysis of trust management with trust chain optimization in mobile ad hoc networks. Journal of Network and Computer Applications, 35(3):1001–1012, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511000701.

Diallo:2012:RTD

[836] Ousmane Diallo, Joel J. P. C. Rodrigues, and Mbaye Sene. Real-time data management on wireless sensor networks: a survey. Journal of Network and Computer Applications, 35(3):1013–1021, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002384.

Zhu:2012:DNR

[837] Wen Tao Zhu, Jianying Zhou, Robert H. Deng, and Feng Bao. Detecting node replication attacks in wireless sensor networks: a survey. Journal of Network and Computer Applications, 35(3):1022– 1034, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000161.

Yin:2012:FSA

[838] Jihao Yin, Yifei Wang, and Jiankun Hu. Free Search with Adaptive Differential Evolution Exploitation and Quantum-Inspired Exploration. Journal of Network and Computer Applications, 35(3):1035– 1051, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002360.

Lee:2012:PPC

[839] Panho Lee, Anura P. Jayasumana, H. M. N. Dilum Bandara, Sanghun Lim, and V. Chandrasekar. A peer-to-peer collaboration framework for multi-sensor data fusion. Journal of Network and Computer Applications, 35(3):1052–1066, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002372.

Lee:2012:ECT

[840] Chao-Yang Lee, Shou-Jen Lin, Chen-Wei Lee, and Chu-Sing Yang. An efficient continuous tracking system in real-time surveillance application. Journal of Network and Computer Applications, 35(3):1067– 1073, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511002396.

Biradar:2012:NSR

[841] Rajashekhar C. Biradar and Sunilkumar S. Manvi. Neighbor supported reliable multipath multicast routing in MANETs. Journal of Network and Computer Applications, 35(3):1074–1085, May 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002402.

Gallardo:2012:MDD

[842] Jesús Gallardo, Crescencio Bravo, and Miguel A. Redondo. A modeldriven development method for collaborative modeling tools. Journal of Network and Computer Applications, 35(3):1086–1105, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511002414.

Suraprasert:2012:CRS

[843] S. Suraprasert and J. Fakcharoenphol. Comparison of recovery schemes to maximize restorable throughput in multicast networks. Journal of Network and Computer Applications, 35(3):1106–1115, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511002426.

Keshavamurthy:2012:HCO

[844] Bettahally N. Keshavamurthy, Durga Toshniwal, and Bhavani K. Eshwar. Hiding co-occurring prioritized sensitive patterns over distributed progressive sequential data streams. Journal of Network and Computer Applications, 35(3):1116–1129, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804511002438. See

corrigendum [868].

Baadache:2012:FAP

[845] Abderrahmane Baadache and Ali Belmehdi. Fighting against packet dropping misbehavior in multi-hop wireless ad hoc networks. Journal of Network and Computer Applications, 35(3):1130–1139, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451100244X.

Batistatos:2012:MTM

[846] M. C. Batistatos, G. V. Tsoulos, and G. E. Athanasiadou. Mobile telemedicine for moving vehicle scenarios: Wireless technology options and challenges. Journal of Network and Computer Applications, 35(3): 1140–1150, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000173.

Baig:2012:MAS

[847] Zubair A. Baig. Multi-agent systems for protecting critical infrastructures: a survey. Journal of Network and Computer Applications, 35(3): 1151–1161, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000203.

Banerjee:2012:PIN

[848] Abhik Banerjee, Chuan Heng Foh, Chai Kiat Yeo, and Bu Sung Lee. Performance improvements for network-wide broadcast with instantaneous network information. Journal of Network and Computer Applications, 35(3):1162–1174, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000227.

Anonymous:2012:EBc

[849] Anonymous. Editorial board. Journal of Network and Computer Applications, 35(3):ifc, May 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000422.

Cuzzocrea:2012:IAD

[850] Alfredo Cuzzocrea. Intelligent algorithms for data-centric sensor networks. Journal of Network and Computer Applications, 35(4):1175– 1176, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000689.

Papadopoulos:2012:VEE

[851] Aris Papadopoulos, Alfredo Navarra, Julie A. McCann, and Cristina M. Pinotti. VIBE: an energy efficient routing protocol for dense and mobile sensor networks. Journal of Network and Computer Applications, 35(4):1177–1190, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001068.

Roantree:2012:DTQ

[852] Mark Roantree, Jie Shi, Paolo Cappellari, Martin F. O’Connor, Michael Whelan, and Niall Moyna. Data transformation and query management in personal health sensor networks. Journal of Network and Computer Applications, 35(4):1191–1202, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804511001032.

Bellavista:2012:FIC

[853] Paolo Bellavista, Giuseppe Cardone, Antonio Corradi, and Luca Foschini. The Future Internet convergence of IMS and ubiquitous smart environments: an IMS-based solution for energy efficiency. Journal of Network and Computer Applications, 35(4):1203–1209, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001056.

Cuzzocrea:2012:EBC

[854] Alfredo Cuzzocrea, Alexis Papadimitriou, Dimitrios Katsaros, and Yannis Manolopoulos. Edge betweenness centrality: a novel algorithm for QoS-based topology control over wireless sensor networks. Journal of Network and Computer Applications, 35(4):1210–1217, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804511001135.

Moro:2012:WGS

[855] Gianluca Moro and Gabriele Monti. W-Grid: a scalable and efficient selforganizing infrastructure for multi-dimensional data management, querying and routing in wireless data-centric sensor networks. Journal of Network and Computer Applications, 35(4):1218–1234, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001044.

Madhusudhan:2012:DIB

[856] R. Madhusudhan and R. C. Mittal. Dynamic ID-based remote user password authentication schemes using smart cards: a review. Journal of Network and Computer Applications, 35(4):1235–1248, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000215.

vonMulert:2012:STS

[857] Jan von Mulert, Ian Welch, and Winston K. G. Seah. Security threats and solutions in MANETs: a case study using AODV and SAODV. Journal of Network and Computer Applications, 35(4):1249– 1259, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000331.

Cheng:2012:ESC

[858] Chien-Fu Cheng and Kuo-Tang Tsai. Eventual strong consensus with fault detection in the presence of dual failure mode on processors under dynamic networks. Journal of Network and Computer Applications, 35(4):1260–1276, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000252.

Noh:2012:UDB

[859] Dong Kun Noh and Junbeom Hur. Using a dynamic backbone for efficient data delivery in solar-powered WSNs. Journal of Network and Computer Applications, 35(4):1277–1284, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512000264.

Gao:2012:DMC

[860] Longxiang Gao, Ming Li, Alessio Bonti, Wanlei Zhou, and Shui Yu. M-dimension: Multi-characteristics based routing protocol in human associated delay-tolerant networks with improved performance over one dimensional classic models. Journal of Network and Computer Applications, 35(4):1285–1296, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000276.

Mansouri:2012:DRM

[861] Najme Mansouri and Gholam Hosein Dastghaibyfard. A dynamic replica management strategy in data grid. Journal of Network and Computer Applications, 35(4):1297–1303, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512000288.

Turner:2012:RDC

[862] Kenneth J. Turner and Koon Leai Larry Tan. Rigorous development of composite grid services. Journal of Network and Computer Applications, 35(4):1304–1316, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451200029X.

Senouci:2012:PEN

[863] Mustapha Reda Senouci, Abdelhamid Mellouk, Hadj Senouci, and Amar Aissani. Performance evaluation of network lifetime spatial-temporal distribution for WSN routing protocols. Journal of Network and Computer Applications, 35(4):1317–1328, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512000306.

Mogaibel:2012:DCR

[864] Hassen A. Mogaibel, Mohamed Othman, Shamala Subramaniam, and Nor Asilah Wati Abdul Hamid. On-demand channel reservation scheme for common traffic in wireless mesh networks. Journal of Network and Computer Applications, 35(4):1329–1351, July 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000318.

Thompson:2012:SDG

[865] Elizabeth A. Thompson, Nathan Clem, Isaac Renninger, and Timothy Loos. Software-defined GPS receiver on USRP-platform. Journal of Network and Computer Applications, 35(4):1352–1360, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000343.

Alkharasani:2012:TNE

[866] Ameen M. Alkharasani and Mohamed Othman. M^{2}I^{2} tswTCM: a new efficient optimization marker algorithm to improve fairness bandwidth in DiffServ networks. Journal of Network and Computer Applications, 35(4):1361–1366, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000355.

Zhou:2012:PED

[867] Miao Zhou, Yi Mu, Willy Susilo, Jun Yan, and Liju Dong. Privacy enhanced data outsourcing in the cloud. Journal of Network and Computer Applications, 35(4):1367–1373, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512000367.

Keshavamurthy:2012:CHC

[868] Bettahally N. Keshavamurthy, Durga Toshniwal, and Bhavani K. Eshwar. Corrigendum to “Hiding co-occurring prioritized sensitive patterns over distributed progressive sequential data streams” [J. Network Comput. Appl. 35 (3) (2012) 1116–1129]. Journal of Network and Computer Applications, 35(4):1374, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000859. See [844].

Anonymous:2012:EBd

[869] Anonymous. Editorial board. Journal of Network and Computer Applications, 35(4):ifc, July 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000781.

Bellavista:2012:SIS

[870] Paolo Bellavista, Chi-Ming Chen, and Hossam Hassanein. Special issue on service delivery management in broadband networks. Journal of Network and Computer Applications, 35(5):1375–1376, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451200118X.

Meddour:2012:CLA

[871] Djamal-Eddine Meddour, Alaeddine Abdallah, Toufik Ahmed, and Raouf Boutaba. A cross layer architecture for multicast and unicast video transmission in mobile broadband networks. Journal of Network and Computer Applications, 35(5):1377–1391, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001950.

Both:2012:SAC

[872] Cristiano Bonato Both, Clarissa Cassales Marquezan, Rafael Kunst, Lisandro Zambenedetti Granville, and Juergen Rochol. A self-adapting connection admission control solution for mobile WiMAX: Enabling dynamic switching of admission control algorithms based on predominant network usage profiles. Journal of Network and Computer Applications, 35(5):1392–1401, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511002013.

Aceto:2012:UAN

[873] Giuseppe Aceto, Alessio Botta, Antonio Pescapé, and Maurizio D’Arienzo. Unified architecture for network measurement: the case of available bandwidth. Journal of Network and Computer Applications, 35(5):1402–1414, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001974.

Sandmann:2012:DSQ

[874] Werner Sandmann. Delays in a series of queues with correlated service times. Journal of Network and Computer Applications, 35(5): 1415–1423, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001962.

Abrougui:2012:PEE

[875] Kaouther Abrougui, Azzedine Boukerche, and Hussam Ramadan. Performance evaluation of an efficient fault tolerant service discovery protocol for vehicular networks. Journal of Network and Computer Applications, 35(5):1424–1435, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804511001949.

Hadid:2012:EMC

[876] Nassima Hadid, Alexandre Guitton, and Michel Misson. Exploiting a meeting channel to interconnect mobile robots. Journal of Network and Computer Applications, 35(5):1436–1445, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001937.

Bouras:2012:EMP

[877] Christos Bouras and Vassilis Poulopoulos. Enhancing meta-portals using dynamic user context personalization techniques. Journal of Network and Computer Applications, 35(5):1446–1453, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001925.

Mosa:2012:EMR

[878] Ahmed A. Mosa, Aisha H. Abdalla, and Rashid A. Saeed. Evaluation of MANEMO route optimization schemes. Journal of Network and Computer Applications, 35(5):1454–1472, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000392.

Ruiz-Martinez:2012:SSM

[879] A. Ruiz-Martínez. A survey on solutions and main free tools for privacy enhancing Web communications. Journal of Network and Computer Applications, 35(5):1473–1492, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000665.

Masdari:2012:STN

[880] Mohammad Masdari, Mehdi Maleknasab, and Moazam Bidaki. A survey and taxonomy of name systems in mobile ad hoc networks. Journal of Network and Computer Applications, 35(5):1493–1507, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512000677.

Zungeru:2012:CSI

[881] Adamu Murtala Zungeru, Li-Minn Ang, and Kah Phooi Seng. Classical and swarm intelligence based routing protocols for wireless sensor networks: a survey and comparison. Journal of Network and Computer Applications, 35(5):1508–1536, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000719.

Liu:2012:RSW

[882] Liang Liu, Xiao-Lin Qin, and Gui-Neng Zheng. Reliable spatial window aggregation query processing algorithm in wireless sensor networks. Journal of Network and Computer Applications, 35(5):1537–1547, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000732.

Mamidisetty:2012:SSC

[883] Kranthi K. Mamidisetty, Michael J. Ferrara, and Shivakumar Sastry. Systematic selection of cluster heads for data collection. Journal of Network and Computer Applications, 35(5):1548–1558, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451200032X.

Ivanov:2012:RTG

[884] Rosen Ivanov. Real-time GPS track simplification algorithm for outdoor navigation of visually impaired. Journal of Network and Computer Applications, 35(5):1559–1567, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000409.

He:2012:MAM

[885] Yun He, Zhongyang Xiong, Yufang Zhang, Xiaolin Tan, and Zhixing Li. Modeling and analysis of multi-channel P2P VoD systems. Journal of Network and Computer Applications, 35(5):1568–1578, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512000598.

Battistello:2012:TBA

[886] Patrick Battistello, Joaquin Garcia-Alfaro, and Cyril Delétré. Transactionbased authentication and key agreement protocol for inter-domain VoIP. Journal of Network and Computer Applications, 35(5):1579–1597, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000653.

Ali:2012:DPM

[887] Samia A. Ali, Khaled M. Shaaban, and Islam M. Alkabbany. Distributed patching for mobile wireless sensor networks. Journal of Network and Computer Applications, 35(5):1598–1605, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000690.

Zhou:2012:DRP

[888] Jipeng Zhou, Liyang Peng, Yuhui Deng, and Jianzhu Lu. An on-demand routing protocol for improving channel use efficiency in multichannel ad hoc networks. Journal of Network and Computer Applications, 35 (5):1606–1614, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000707.

Leu:2012:WVA

[889] Seiven Leu and Ruay-Shiung Chang. A weight-value algorithm for finding connected dominating sets in a MANET. Journal of Network and Computer Applications, 35(5):1615–1619, September 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000720.

Vankeirsbilck:2012:AFG

[890] Bert Vankeirsbilck, Dieter Verslype, Nicolas Staelens, Pieter Simoens, Chris Develder, Bart Dhoedt, Filip De Turck, and Piet Demeester. Automatic fine-grained area detection for thin client systems. Journal of Network and Computer Applications, 35(5):1620–1632, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512000744.

Torkestani:2012:MPM

[891] Javad Akbari Torkestani. Mobility prediction in mobile wireless networks. Journal of Network and Computer Applications, 35(5):1633–1645, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000756.

Das:2012:DPB

[892] Ashok Kumar Das, Pranay Sharma, Santanu Chatterjee, and Jamuna Kanta Sing. A dynamic password-based user authentication scheme for hierarchical wireless sensor networks. Journal of Network and Computer Applications, 35(5):1646–1656, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000926.

Anonymous:2012:EBe

[893] Anonymous. Editorial board. Journal of Network and Computer Applications, 35(5):ifc, September 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001208.

Shah:2012:PCE

[894] Peer Azmat Shah, Muhammad Yousaf, Amir Qayyum, and Halabi B. Hasbullah. Performance comparison of end-to-end mobility management protocols for TCP. Journal of Network and Computer Applications, 35 (6):1657–1673, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001166.

Ramaboli:2012:BAH

[895] Allen L. Ramaboli, Olabisi E. Falowo, and Anthony H. Chan. Bandwidth aggregation in heterogeneous wireless networks: a survey of current approaches and issues. Journal of Network and Computer Applications, 35(6):1674–1690, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001452.

Parvin:2012:CRN

[896] Sazia Parvin, Farookh Khadeer Hussain, Omar Khadeer Hussain, Song Han, Biming Tian, and Elizabeth Chang. Cognitive radio network security: a survey. Journal of Network and Computer Applications, 35 (6):1691–1708, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001518.

Benedict:2012:EAP

[897] Shajulin Benedict. Energy-aware performance analysis methodologies for HPC architectures — an exploratory study. Journal of Network and Computer Applications, 35(6):1709–1719, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001798.

Moulahi:2012:BBD

[898] Tarek Moulahi, Salem Nasri, and Hervé Guyennet. Broadcasting based on dominated connecting sets with MPR in a realistic environment for WSNs & ad hoc. Journal of Network and Computer Applications, 35 (6):1720–1727, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001816.

Choi:2012:PSM

[899] JungYul Choi, Min-Gon Kim, Hongkyu Jeong, and Hong-Shik Park. Power-saving mechanisms for energy efficient IEEE 802.16e/m. Journal of Network and Computer Applications, 35(6):1728–1739, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512001877.

Yin:2012:UIS

[900] Jihao Yin, Chongyang Fu, and Jiankun Hu. Using incremental subspace and contour template for object tracking. Journal of Network and Computer Applications, 35(6):1740–1748, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001506.

Shokrzadeh:2012:OQD

[901] Hamid Shokrzadeh, Siavash Khorsandi, and Abolfazl Toroghi Haghighat. Optimized query-driven appointment routing based on Expectation-Maximization in wireless sensor networks. Journal of Network and Computer Applications, 35(6):1749–1761, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451200152X.

Shin:2012:TMS

[902] Kwangcheol Shin, Kyungjun Kim, and Soontae Kim. Traffic management strategy for delay-tolerant networks. Journal of Network and Computer Applications, 35(6):1762–1770, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001531.

Kakkasageri:2012:MDD

[903] M. S. Kakkasageri and S. S. Manvi. Multiagent driven dynamic clustering of vehicles in VANETs. Journal of Network and Computer Applications, 35(6):1771–1780, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001543.

AguiarAmaral:2012:INA

[904] Alexandre Aguiar Amaral, Bruno Bogaz Zarpelão, Leonardo de Souza Mendes, Joel José Puga Coelho Rodrigues, and Mario Lemes Proença Junior. Inference of network anomaly propagation using spatio-temporal correlation. Journal of Network and Computer Applications, 35(6):1781– 1792, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001555.

Shi:2012:EAS

[905] Q. Shi, N. Zhang, and D. Llewellyn-Jones. Efficient autonomous signature exchange on ubiquitous networks. Journal of Network and Computer Applications, 35(6):1793–1806, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001567.

Zhang:2012:KBP

[906] Victoria Ying Zhang and Albert Kai sun Wong. Kernel-based particle filtering for indoor tracking in WLANs. Journal of Network and Computer Applications, 35(6):1807–1817, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001579.

Amad:2012:HNH

[907] Mourad Amad, Ahmed Meddahi, Djamil Aïssani, and Zonghua Zhang. HPM: a novel hierarchical Peer-to-Peer model for lookup acceleration with provision of physical proximity. Journal of Network and Computer Applications, 35(6):1818–1830, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001580.

Sood:2012:CAE

[908] Sandeep K. Sood. A combined approach to ensure data security in cloud computing. Journal of Network and Computer Applications, 35 (6):1831–1838, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001592.

Tan:2012:LCP

[909] Zuowen Tan. A lightweight conditional privacy-preserving authentication and access control scheme for pervasive computing environments. Journal of Network and Computer Applications, 35(6):1839–1846, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001609.

Dong:2012:OAS

[910] Bo Dong, Qinghua Zheng, Feng Tian, Kuo-Ming Chao, Rui Ma, and Rachid Anane. An optimized approach for storing and accessing small files on cloud storage. Journal of Network and Computer Applications, 35(6):1847–1862, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001610.

Antonello:2012:DPI

[911] Rafael Antonello, Stenio Fernandes, Carlos Kamienski, Djamel Sadok, Judith Kelner, István Gódor, Géza Szabó, and Tord Westholm. Deep packet inspection tools and techniques in commodity platforms: Challenges and trends. Journal of Network and Computer Applications, 35 (6):1863–1878, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001622.

Domingo:2012:OIU

[912] Mari Carmen Domingo. An overview of the Internet of underwater things. Journal of Network and Computer Applications, 35(6):1879– 1890, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001646.

Barrachina:2012:VVA

[913] Javier Barrachina, Piedad Garrido, Manuel Fogue, Francisco J. Martinez, Juan-Carlos Cano, Carlos T. Calafate, and Pietro Manzoni. VEACON: a Vehicular Accident Ontology designed to improve safety on the roads. Journal of Network and Computer Applications, 35 (6):1891–1900, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001658.

Zungeru:2012:THP

[914] Adamu Murtala Zungeru, Li-Minn Ang, and Kah Phooi Seng. Termitehill: Performance optimized swarm intelligence based routing algorithm for wireless sensor networks. Journal of Network and Computer Applications, 35(6):1901–1917, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451200166X.

Bakillah:2012:RTQ

[915] Mohamed Bakillah, Mir Abolfazl Mostafavi, Steve H. L. Liang, and Alexander Zipf. Real time query propagation strategies with Lightweight Coordination Calculus (LCC) for ad hoc networks of geospatial databases. Journal of Network and Computer Applications, 35 (6):1918–1933, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001671.

Fortino:2012:FBM

[916] G. Fortino, A. Guerrieri, G. M. P. O’Hare, and A. Ruzzelli. A flexible building management framework based on wireless sensor and actuator networks. Journal of Network and Computer Applications, 35 (6):1934–1952, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001683.

Liu:2012:TCA

[917] Linfeng Liu, Ruchuan Wang, and Fu Xiao. Topology control algorithm for underwater wireless sensor networks using GPS-free mobile sensor nodes. Journal of Network and Computer Applications, 35 (6):1953–1963, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001695.

Singh:2012:ACA

[918] Gurpreet Singh, Neeraj Kumar, and Anil Kumar Verma. Ant colony algorithms in MANETs: a review. Journal of Network and Computer Applications, 35(6):1964–1972, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512001701.

Sudha:2012:AMB

[919] S. Sudha and N. Ammasaigounden. An aggregate marker for bandwidth fairness in DiffServ. Journal of Network and Computer Applications, 35(6):1973–1978, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001725.

Li:2012:SMN

[920] Li Li, Qing-Shan Jia, Hengtao Wang, Ruixi Yuan, and Xiaohong Guan. A systematic method for network topology reconfiguration with limited link additions. Journal of Network and Computer Applications, 35(6):1979–1989, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001737.

Hu:2012:HOC

[921] Chih-Lin Hu and Tzu-Han Kuo. A hierarchical overlay with cluster-based reputation tree for dynamic peer-to-peer systems. Journal of Network and Computer Applications, 35(6):1990–2002, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001749.

deCastro:2012:ABI

[922] Miguel Franklin de Castro, Levi Bayde Ribeiro, and Camila Helena Souza Oliveira. An autonomic bio-inspired algorithm for wireless sensor network self-organization and efficient routing. Journal of Network and Computer Applications, 35(6):2003–2015, November 2012. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001750.

Hajiesmaili:2012:CAR

[923] Mohammad H. Hajiesmaili, Ahmad Khonsari, Ali Sehati, and Mohammad Sadegh Talebi. Content-aware rate allocation for efficient video streaming via dynamic network utility maximization. Journal of Network and Computer Applications, 35(6):2016–2027, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001762.

Torkestani:2012:DRD

[924] Javad Akbari Torkestani. A distributed resource discovery algorithm for P2P grids. Journal of Network and Computer Applications, 35 (6):2028–2036, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001774.

Gregor:2012:OBS

[925] D. Gregor, S. L. Toral, T. Ariza, and F. Barrero. An ontology-based semantic service for cooperative urban equipments. Journal of Network and Computer Applications, 35(6):2037–2050, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001786.

Song:2012:LSE

[926] Qingyang Song, Zhaolong Ning, Shiqiang Wang, and Abbas Jamalipour. Link stability estimation based on link connectivity changes in mobile ad-hoc networks. Journal of Network and Computer Applications, 35 (6):2051–2058, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001804.

Kardas:2012:PEO

[927] Süleyman Kardas, Serkan Çelik, Muhammet Yildiz, and Albert Levi. PUF-enhanced offline RFID security and privacy. Journal of Network and Computer Applications, 35(6):2059–2067, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001828.

Darabkh:2012:PES

[928] Khalid A. Darabkh, Shereen S. Ismail, Mohammad Al-Shurman, Iyad F. Jafar, Eman Alkhader, and Mamoun F. Al-Mistarihi. Performance evaluation of selective and adaptive heads clustering algorithms over wireless sensor networks. Journal of Network and Computer Applications, 35(6):2068–2080, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001841.

Xie:2012:SAM

[929] Yi Xie, J. Hu, S. Tang, and X. Huang. A structural approach for modelling the hierarchical dynamic process of Web workload in a large-scale campus network. Journal of Network and Computer Applications, 35 (6):2081–2091, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001865.

Lai:2012:BBE

[930] Yuan-Cheng Lai and Yen-Hung Chen. A best block exploring algorithm for two-dimensional downlink burst construction in IEEE 802.16 networks. Journal of Network and Computer Applications, 35(6):2092– 2104, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001889.

Anonymous:2012:EBf

[931] Anonymous. Editorial board. Journal of Network and Computer Applications, 35(6):ifc, November 2012. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002007.

Hadjidj:2013:WSN

[932] Abdelkrim Hadjidj, Marion Souil, Abdelmadjid Bouabdallah, Yacine Challal, and Henry Owen. Wireless sensor networks for rehabilitation applications: Challenges and opportunities. Journal of Network and Computer Applications, 36(1):1–15, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002196.

Liao:2013:IDS

[933] Hung-Jen Liao, Chun-Hung Richard Lin, Ying-Chih Lin, and Kuang-Yuan Tung. Intrusion detection system: a comprehensive review. Journal of Network and Computer Applications, 36(1):16–24, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512001944.

Patel:2013:IDP

[934] Ahmed Patel, Mona Taghavi, Kaveh Bakhtiyari, and Joaquim Celestino Júnior. An intrusion detection and prevention system in cloud computing: a systematic review. Journal of Network and Computer Applications, 36(1):25–41, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451200183X.

Modi:2013:SID

[935] Chirag Modi, Dhiren Patel, Bhavesh Borisaniya, Hiren Patel, Avi Patel, and Muttukrishnan Rajarajan. A survey of intrusion detection techniques in Cloud. Journal of Network and Computer Applications, 36(1):42–57, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001178.

Mhiri:2013:SIM

[936] Fadoua Mhiri, Kaouthar Sethom, and Ridha Bouallegue. A survey on interference management techniques in femtocell self-organizing networks. Journal of Network and Computer Applications, 36(1):58–65, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001142.

Hassan:2013:SHI

[937] Md. Mahedi Hassan and Poo Kuan Hoong. Seamless handover integrated solution for video transmission over proxy mobile IPv6 in a micro mobility domain. Journal of Network and Computer Applications, 36(1):66–76, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001130.

Pantoni:2013:GBR

[938] Rodrigo Palucci Pantoni and Dennis Brandão. A gradient based routing scheme for street lighting wireless sensor networks. Journal of Network and Computer Applications, 36(1):77–90, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001129.

Kumar:2013:RTT

[939] D. Satish Kumar and N. Nagarajan. Relay technologies and technical issues in IEEE 802.16j Mobile Multi-hop Relay (MMR) networks. Journal of Network and Computer Applications, 36(1):91–102, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001026.

Emmanouilidis:2013:MGT

[940] Christos Emmanouilidis, Remous-Aris Koutsiamanis, and Aimilia Tasidou. Mobile guides: Taxonomy of architectures, context awareness, technologies and applications. Journal of Network and Computer Applications, 36(1):103–125, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001002.

Lar:2013:ICB

[941] Saleem ullah Lar and Xiaofeng Liao. An initiative for a classified bibliography on TCP/IP congestion control. Journal of Network and Computer Applications, 36(1):126–133, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804512000963.

Rathnayaka:2013:WSN

[942] A. J. Dinusha Rathnayaka and Vidyasagar M. Potdar. Wireless Sensor Network transport protocol: a critical review. Journal of Network and Computer Applications, 36(1):134–146, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511001883.

Bah:2013:SSB

[943] Slimane Bah, Roch Glitho, and Rachida Dssouli. A SIP servlets-based framework for service provisioning in stand-alone MANETs. Journal of Network and Computer Applications, 36(1):147–155, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512002226.

Han:2013:LLE

[944] Seung-Wan Han, In-Seon Jeong, and Seung-Ho Kang. Low latency and energy efficient routing tree for wireless sensor networks with multiple mobile sinks. Journal of Network and Computer Applications, 36(1):156–166, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002202.

Xiong:2013:MMA

[945] Junjie Xiong, Yangfan Zhou, Michael R. Lyu, and Evan F. Y. Young. MDiag: Mobility-assisted diagnosis for wireless sensor networks. Journal of Network and Computer Applications, 36(1):167– 177, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002160.

Zhu:2013:TSC

[946] Wen Tao Zhu. Towards secure and communication-efficient broadcast encryption systems. Journal of Network and Computer Applications, 36(1):178–186, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002159.

Hu:2013:ESI

[947] Bing Hu, Kwan L. Yeung, and Zhaoyang Zhang. An efficient single-iteration single-bit request scheduling algorithm for input-queued switches. Journal of Network and Computer Applications, 36(1):187– 194, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002147.

Akkaya:2013:HLS

[948] Kemal Akkaya, Izzet F. Senturk, and Shanthi Vemulapalli. Handling large-scale node failures in mobile sensor/robot networks. Journal of Network and Computer Applications, 36(1):195–210, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002135.

Wang:2013:MMM

[949] Zhan feng Wang, Ming Chen, Chang you Xing, Jing Feng, Xiang lin Wei, and Hua li Bai. Multi-manifold model of the Internet delay space. Journal of Network and Computer Applications, 36(1):211– 218, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001932.

Famaey:2013:TPC

[950] Jeroen Famaey, Frédéric Iterbeke, Tim Wauters, and Filip De Turck. Towards a predictive cache replacement strategy for multimedia content. Journal of Network and Computer Applications, 36(1):219–227, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001919.

Dai:2013:MTI

[951] Huan Dai, Zhao min Zhu, and Xiao-Feng Gu. Multi-target indoor localization and tracking on video monitoring system in a wireless sensor network. Journal of Network and Computer Applications, 36(1):228– 234, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001907.

Waqar:2013:FPC

[952] Adeela Waqar, Asad Raza, Haider Abbas, and Muhammad Khurram Khan. A framework for preservation of cloud users’ data privacy using dynamic reconstruction of metadata. Journal of Network and Computer Applications, 36(1):235–248, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804512001890.

Jappinen:2013:APF

[953] Pekka Jäppinen, Renata Guarneri, and Luis M. Correia. An applications perspective into the Future Internet. Journal of Network and Computer Applications, 36(1):249–254, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001853.

delaTorre-Diez:2013:DPE

[954] Isabel de la Torre-Díez, Saúl Álvaro-Muñoz, Miguel López-Coronado, and Joel Rodrigues. Development and performance evaluation of a new RSS tool for a Web-based system: RSS PROYECT. Journal of Network and Computer Applications, 36(1):255–261, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451200149X.

Naderan:2013:STA

[955] Marjan Naderan, Mehdi Dehghan, and Hossein Pedram. Sensing task assignment via sensor selection for maximum target coverage in WSNs. Journal of Network and Computer Applications, 36(1):262– 273, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001488.

Braga:2013:RRT

[956] Reinaldo B. Braga, Igor A. Chaves, Carina T. de Oliveira, Rossana M. C. Andrade, José Neuman de Souza, Hervé Martin, and Bruno Schulze. RETENTION: a reactive trust-based mechanism to detect and punish malicious nodes in ad hoc grid environments. Journal of Network and Computer Applications, 36(1):274–283, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001476.

Yeh:2013:SEB

[957] Lo-Yao Yeh, Chun-Chuan Yang, Jee-Gong Chang, and Yi-Lang Tsai. A secure and efficient batch binding update scheme for route optimization of nested NEtwork MObility (NEMO) in VANETs. Journal of Network and Computer Applications, 36(1):284–292, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001464.

Kim:2013:BUP

[958] Minwoo Kim, Karam Park, and Won W. Ro. Benefits of using parallelized non-progressive network coding. Journal of Network and Computer Applications, 36(1):293–305, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804512001440.

Yang:2013:DDL

[959] Ting Yang, Yugeng Sun, Javid Taheri, and Albert Y. Zomaya. DLS: a dynamic local stitching mechanism to rectify transmitting path fragments in wireless sensor networks. Journal of Network and Computer Applications, 36(1):306–315, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001427.

Xue:2013:TCB

[960] Kaiping Xue, Changsha Ma, Peilin Hong, and Rong Ding. A temporalcredential-based mutual authentication and key agreement scheme for wireless sensor networks. Journal of Network and Computer Applications, 36(1):316–323, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001403.

Islam:2013:MTP

[961] Rafiqul Islam and Jemal Abawajy. A multi-tier phishing detection and filtering approach. Journal of Network and Computer Applications, 36(1):324–335, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001397.

Baskaran:2013:SBS

[962] Vishnu Monn Baskaran, Yoong Choon Chang, Jonathan Loo, and Kok-Sheik Wong. Software-based serverless endpoint video combiner architecture for high-definition multiparty video conferencing. Journal of Network and Computer Applications, 36(1):336–352, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001385.

Li:2013:ABE

[963] Mingfu Li, Yueh-Lin Wu, and Chia-Rong Chang. Available bandwidth estimation for the network paths with multiple tight links and bursty traffic. Journal of Network and Computer Applications, 36(1):353– 367, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001373.

Al-Turjman:2013:QCW

[964] Fadi M. Al-Turjman, Hossam S. Hassanein, and Mohamad Ibnkahla. Quantifying connectivity in wireless sensor networks with grid-based deployments. Journal of Network and Computer Applications, 36(1):368– 377, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001361.

Yi:2013:PIV

[965] Xun Yi and Eiji Okamoto. Practical Internet voting system. Journal of Network and Computer Applications, 36(1):378–387, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451200135X.

Wang:2013:RAL

[966] Yuehua Wang, Zhong Zhou, Ling Liu, and Wei Wu. Replica-aided load balancing in overlay networks. Journal of Network and Computer Applications, 36(1):388–401, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512001154.

Rosaci:2013:CMI

[967] D. Rosaci and G. M. L. Sarné. Cloning mechanisms to improve agent performances. Journal of Network and Computer Applications, 36(1): 402–408, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001117.

Hooshmand:2013:VSN

[968] M. Hooshmand, S. M. R. Soroushmehr, P. Khadivi, S. Samavi, and S. Shirani. Visual sensor network lifetime maximization by prioritized scheduling of nodes. Journal of Network and Computer Applications, 36(1):409–419, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001105.

Hou:2013:SPE

[969] Weigang Hou, Lei Guo, and Xiaoxue Gong. Survivable power efficiency oriented integrated grooming in green networks. Journal of Network and Computer Applications, 36(1):420–428, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001099.

Alvaro:2013:NCC

[970] Jesús L. Alvaro and Beatriz Barros. A new cloud computing architecture for music composition. Journal of Network and Computer Applications, 36(1):429–443, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001087.

Peng:2013:ECB

[971] I-Hsuan Peng and Yen-Wen Chen. Energy consumption bounds analysis and its applications for grid based wireless sensor networks. Journal of Network and Computer Applications, 36(1):444–451, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512001075.

Salah:2013:PIF

[972] K. Salah and M. Hamawi. Performance of IP-forwarding of Linux hosts with multiple network interfaces. Journal of Network and Computer Applications, 36(1):452–465, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512001063.

Baig:2013:PAP

[973] Imran Baig, Muhammad Ayaz, and Varun Jeoti. On the peak-to-average power ratio reduction in mobile WiMAX: a discrete cosine transform matrix precoding based random-interleaved orthogonal frequency division multiple access uplink system. Journal of Network and Computer Applications, 36(1):466–475, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512001051.

Looney:2013:PAP

[974] Mary Looney and Oliver Gough. A provision aware proportional fair sharing three colour marker. Journal of Network and Computer Applications, 36(1):476–483, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451200104X.

Belmonte:2013:CCI

[975] Maria-Victoria Belmonte, Manuel Díaz, Jose-Luis Pérez de-la Cruz, and Ana Reyna. COINS: COalitions and INcentiveS for effective peerto-peer downloads. Journal of Network and Computer Applications, 36(1):484–497, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001038.

Lv:2013:NTP

[976] Chao Lv, Maode Ma, Hui Li, Jianfeng Ma, and Yaoyu Zhang. An novel three-party authenticated key exchange protocol using one-time key. Journal of Network and Computer Applications, 36(1):498–503, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000999.

Bernardino:2013:SOA

[977] Eugénia Moreira Bernardino, Anabela Moreira Bernardino, Juan Manuel Sánchez-Pérez, Juan Antonio Gómez Pulido, and Miguel A. Vega Rodríguez. Swarm optimisation algorithms applied to large balanced communication networks. Journal of Network and Computer Applications, 36(1):504–522, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000987.

Salameh:2013:OMA

[978] Haythem Bany Salameh and Osamah S. Badarneh. Opportunistic medium access control for maximizing packet delivery rate in dynamic access networks. Journal of Network and Computer Applications, 36(1): 523–532, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000975.

Chang:2013:PCC

[979] Ting-Yi Chang and Yan-Ru Ke. A personalized e-course composition based on a genetic algorithm with forcing legality in an adaptive learning system. Journal of Network and Computer Applications, 36(1):533–542, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000951.

Adabi:2013:MGR

[980] Sepideh Adabi, Ali Movaghar, Amir Masoud Rahmani, and Hamid Beigy. Market-based grid resource allocation using new negotiation model. Journal of Network and Computer Applications, 36(1):543– 565, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451200063X.

Anonymous:2013:EBa

[981] Anonymous. Editorial board. Journal of Network and Computer Applications, 36(1):ifc, January 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002354.

Li:2013:SNF

[982] Bingdong Li, Jeff Springer, George Bebis, and Mehmet Hadi Gunes. A survey of network flow applications. Journal of Network and Computer Applications, 36(2):567–581, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002676.

Qazi:2013:SDA

[983] Shams Qazi, Raad Raad, Yi Mu, and Willy Susilo. Securing DSR against wormhole attacks in multirate ad hoc networks. Journal of Network and Computer Applications, 36(2):582–592, March 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002664.

Mohammed:2013:STW

[984] Ahmed Hassan Mohammed, Bin Dai, Benxiong Huang, Muhammad Azhar, Guan Xu, Peng Qin, and Shui Yu. A survey and tutorial of wireless relay network protocols based on network coding. Journal of Network and Computer Applications, 36(2):593–610, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002652.

He:2013:DKM

[985] Xiaobing He, Michael Niedermeier, and Hermann de Meer. Dynamic key management in wireless sensor networks: a survey. Journal of Network and Computer Applications, 36(2):611–622, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002573.

Tyagi:2013:SRC

[986] Sudhanshu Tyagi and Neeraj Kumar. A systematic review on clustering and routing techniques based upon LEACH protocol for wireless sensor networks. Journal of Network and Computer Applications, 36(2):623–645, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002482.

Islam:2013:CMB

[987] Rafiqul Islam, Ronghua Tian, Lynn M. Batten, and Steve Versteeg. Classification of malware based on integrated static and dynamic features. Journal of Network and Computer Applications, 36(2):646– 656, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002214.

Huang:2013:FDH

[988] Kun Huang, Gaogang Xie, Rui Li, and Shuai Xiong. Fast and deterministic hash table lookup using discriminative Bloom filters. Journal of Network and Computer Applications, 36(2):657–666, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000088.

Wang:2013:DAT

[989] Chiapin Wang. Dynamic ARF for throughput improvement in 802.11 WLAN via a machine-learning approach. Journal of Network and Computer Applications, 36(2):667–676, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513000027.

Canlar:2013:WML

[990] Eyüp S. Canlar, Mauro Conti, Bruno Crispo, and Roberto Di Pietro. Windows Mobile LiveSD Forensics. Journal of Network and Computer Applications, 36(2):677–684, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002718.

Bilal:2013:PBR

[991] Sardar Muhammad Bilal, Carlos Jesus Bernardos, and Carmen Guerrero. Position-based routing in vehicular networks: a survey. Journal of Network and Computer Applications, 36(2):685–697, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002706.

Kretschmer:2013:TFO

[992] M. Kretschmer, P. Batroff, and G. Ghinea. Topology forming and optimization framework for heterogeneous wireless back-haul networks supporting unidirectional technologies. Journal of Network and Computer Applications, 36(2):698–710, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451200269X.

Mansouri:2013:CDR

[993] Najme Mansouri, Gholam Hosein Dastghaibyfard, and Ehsan Mansouri. Combination of data replication and scheduling algorithm for improving data availability in Data Grids. Journal of Network and Computer Applications, 36(2):711–722, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002688.

Vasiliadis:2013:MPS

[994] D. C. Vasiliadis, G. E. Rizos, and C. Vassilakis. Modelling and performance study of finite-buffered blocking multistage interconnection networks supporting natively 2-class priority routing traffic. Journal of Network and Computer Applications, 36(2):723–737, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002639.

Luo:2013:TEN

[995] Shiliang Luo, Xu Lu, and Lianglun Cheng. TSOIA: an efficient node selection algorithm facing the uncertain process for Internet of Things. Journal of Network and Computer Applications, 36(2):738– 743, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002627.

Cervera:2013:MRS

[996] Gimer Cervera, Michel Barbeau, Joaquin Garcia-Alfaro, and Evangelos Kranakis. A multipath routing strategy to prevent flooding disruption attacks in link state routing protocols for MANETs. Journal of Network and Computer Applications, 36(2):744–755, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002603.

Lian:2013:IMS

[997] Kuang-Yow Lian, Sung-Jung Hsiao, and Wen-Tsai Sung. Intelligent multi-sensor control system based on innovative technology integration via ZigBee and Wi-Fi networks. Journal of Network and Computer Applications, 36(2):756–767, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002597.

Lin:2013:REP

[998] Po-Ching Lin and Jia-Hau Lee. Re-examining the performance bottleneck in a NIDS with detailed profiling. Journal of Network and Computer Applications, 36(2):768–780, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002561.

Yang:2013:QSE

[999] Yan Yang, Yulong Zhang, Alex Hai Wang, Meng Yu, Wanyu Zang, Peng Liu, and Sushil Jajodia. Quantitative survivability evaluation of three virtual machine-based server architectures. Journal of Network and Computer Applications, 36(2):781–790, March 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002536.

Xu:2013:NMU

[1000] C. Xu, C. Du, G. F. Zhao, and S. Yu. A novel model for user clicks identification based on hidden semi-Markov. Journal of Network and Computer Applications, 36(2):791–798, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002524.

An:2013:RSR

[1001] Jian An, Xiaolin Gui, Wendong Zhang, Jinhua Jiang, and Jianwei Yang. Research on social relations cognitive model of mobile nodes in Internet of Things. Journal of Network and Computer Applications, 36(2):799–810, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002512.

Liu:2013:VSG

[1002] Yanheng Liu, Longxiang Suo, Dayang Sun, and Aimin Wang. A virtual square grid-based coverage algorithm of redundant node for wireless sensor network. Journal of Network and Computer Applications, 36(2):811–817, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002500.

Xiao:2013:VMP

[1003] Peng Xiao, Zhigang Hu, Dongbo Liu, Guofeng Yan, and Xilong Qu. Virtual machine power measuring technique with bounded error in cloud environments. Journal of Network and Computer Applications, 36(2):

818–828, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002494.

Chan:2013:DFA

[1004] Gaik-Yee Chan, Chien-Sing Lee, and Swee-Huay Heng. Discovering fuzzy association rule patterns and increasing sensitivity analysis of XML-related attacks. Journal of Network and Computer Applications, 36(2):829–842, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002470.

Peng:2013:EJC

[1005] Yuhuai Peng, Yao Yu, Lei Guo, Dingde Jiang, and Qiming Gai. An efficient joint channel assignment and QoS routing protocol for IEEE 802.11 multi-radio multi-channel wireless mesh networks. Journal of Network and Computer Applications, 36(2):843–857, March 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002299.

Jamali:2013:UFI

[1006] Shahram Jamali, Seyyed Nasser Seyyed Hashemi, and Amir Masoud Eftekhari Moghadam. On the use of a full information feedback to stabilize RED. Journal of Network and Computer Applications, 36(2):858–869, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002287.

Martin-Campillo:2013:EON

[1007] Abraham Martín-Campillo, Jon Crowcroft, Eiko Yoneki, and Ramon Martí. Evaluating opportunistic networks in disaster scenarios. Journal of Network and Computer Applications, 36(2):870–880, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002275.

Kumar:2013:ABS

[1008] Mahesh Kumar and Shishir Kumar. AGRIB-BHF: a sustainable and fault tolerant aggregation. Journal of Network and Computer Applications, 36(2):881–886, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002263.

Al-shugran:2013:QCE

[1009] Mahmoud Al-shugran, Osman Ghazali, Suhaidi Hassan, Kashif Nisar, and A. Suki M. Arif. A qualitative comparison evaluation of the greedy forwarding strategies in Mobile Ad Hoc Network. Journal of Network and Computer Applications, 36(2):887–897, March 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002251.

Nie:2013:PLB

[1010] Laisen Nie, Dingde Jiang, and Lei Guo. A power laws-based reconstruction approach to end-to-end network traffic. Journal of Network and Computer Applications, 36(2):898–907, March 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451200224X.

Ever:2013:HAM

[1011] Enver Ever, Orhan Gemikonakli, Altan Kocyigit, and Eser Gemikonakli. A hybrid approach to minimize state space explosion problem for the solution of two stage tandem queues. Journal of Network and Computer Applications, 36(2):908–926, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002238.

Xiong:2013:NIB

[1012] Hu Xiong, Zhong Chen, and Fagen Li. New identity-based threeparty authenticated key agreement protocol with provable security. Journal of Network and Computer Applications, 36(2):927–932, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512002184.

Nisar:2013:CSS

[1013] Kashif Nisar, Angela Amphawan, Suhaidi Hassan, and Nurul I. Sarkar. A comprehensive survey on scheduler for VoIP over WLAN. Journal of Network and Computer Applications, 36(2):933–948, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001713.

Watfa:2013:NSE

[1014] Mohamed K. Watfa, Haitham Al-Hassanieh, and Samir Salmen. A novel solution to the energy hole problem in sensor networks. Journal of Network and Computer Applications, 36(2):949–958, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512001634.

Anonymous:2013:EBb

[1015] Anonymous. Editorial board. Journal of Network and Computer Applications, 36(2):ifc, March 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000271.

Rodrigues:2013:GEVa

[1016] Joel J. P. C. Rodrigues, Russel Hsing, Min Chen, Bingli Jiao, and Binod Vaidya. Guest editorial on Vehicular Networking Protocols. Journal of Network and Computer Applications, 36(3):959–960, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000805.

Fonseca:2013:APB

[1017] António Fonseca and Teresa Vazão. Applicability of position-based routing for VANET in highways and urban environment. Journal of Network and Computer Applications, 36(3):961–973, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000768.

Nair:2013:SFM

[1018] Rajashree Nair, Ben Soh, Naveen Chilamkurti, and James J. (Jong Hyuk) Park. Structure-free message aggregation and routing in traffic information system (SMART). Journal of Network and Computer Applications, 36(3):974–980, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512001014.

Bitam:2013:BLB

[1019] Salim Bitam and Abdelhamid Mellouk. Bee life-based multi constraints multicast routing optimization for vehicular ad hoc networks. Journal of Network and Computer Applications, 36(3):981–991, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000379.

Zhang:2013:DEF

[1020] Zhenxia Zhang, Azzedine Boukerche, and Hussam Ramadan. Design and evaluation of a fast MAC layer handoff management scheme for WiFi-based multichannel Vehicular Mesh Networks. Journal of Network and Computer Applications, 36(3):992–1000, May 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000380.

Antonopoulos:2013:NCB

[1021] Angelos Antonopoulos, Charalabos Skianis, and Christos Verikoukis. Network coding-based cooperative ARQ scheme for VANETs. Journal of Network and Computer Applications, 36(3):1001–1007, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512000938.

Tadayon:2013:PMS

[1022] Navid Tadayon, Sasan Khoshroo, Elaheh Askari, Honggang Wang, and Howard Michel. Power management in SMAC-based energy-harvesting wireless sensor networks using queuing analysis. Journal of Network and Computer Applications, 36(3):1008–1017, May 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000239.

Wang:2013:PIQ

[1023] Youxiang Wang, Jianquan Wang, Lu Zhaobiao, and Xingang Liu. Performance improvement of QO-STBC over time-selective channel for wireless network. Journal of Network and Computer Applications, 36(3): 1018–1026, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000197.

Lloret:2013:GBP

[1024] Jaime Lloret, Alejandro Canovas, Angel Catalá, and Miguel Garcia. Group-based protocol and mobility model for VANETs to offer Internet access. Journal of Network and Computer Applications, 36(3):1027– 1038, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512000641.

Kabir:2013:ERO

[1025] Md. Humayun Kabir, Mohammad Mukhtaruzzaman, and Mohammed Atiquzzaman. Efficient route optimization scheme for nested-NEMO. Journal of Network and Computer Applications, 36(3):1039–1049, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000039.

Zhang:2013:MAV

[1026] Liren Zhang, Abderrahmane Lakas, Hesham El-Sayed, and Ezedin Barka. Mobility analysis in vehicular ad hoc network (VANET). Journal of Network and Computer Applications, 36(3):1050–1056, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451200255X.

Liu:2013:PBB

[1027] Yejun Liu, Qingyang Song, Rui Ma, Bing Li, and Bo Gong. Protection based on backup radios and backup fibers for survivable Fiber-Wireless (FiWi) access network. Journal of Network and Computer Applications, 36(3):1057–1069, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000234.

Luo:2013:PBA

[1028] Jian-Zhen Luo and Shun-Zheng Yu. Position-based automatic reverse engineering of network protocols. Journal of Network and Computer Applications, 36(3):1070–1077, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000222.

Zarifneshat:2013:UMN

[1029] Masoud Zarifneshat and Pejman Khadivi. Using mobile node speed changes for movement direction change prediction in a realistic category of mobility models. Journal of Network and Computer Applications, 36(3):1078–1090, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000210.

Anonymous:2013:EBc

[1030] Anonymous. Editorial board. Journal of Network and Computer Applications, 36(3):ifc, May 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000829.

Mahdipour:2013:ISJ

[1031] Ebrahim Mahdipour, Amir Masoud Rahmani, and Saeed Setayeshi. Importance sampling for Jackson networks with customer impatience until the end of service. Journal of Network and Computer Applications, 36(4):1091–1101, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000258.

Dumez:2013:MDA

[1032] C. Dumez, M. Bakhouya, J. Gaber, M. Wack, and P. Lorenz. Modeldriven approach supporting formal verification for web service composition protocols. Journal of Network and Computer Applications, 36(4): 1102–1115, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000180.

Mohamed:2013:DDT

[1033] Nader Mohamed, Jameela Al-Jaroodi, and Abdulla Eid. A dual-direction technique for fast file downloads with dynamic load balancing in the Cloud. Journal of Network and Computer Applications, 36(4):1116– 1130, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000155.

Haider:2013:SCA

[1034] Ammar Haider, Iqbal Gondal, and Joarder Kamruzzaman. Socialconnectivity-aware vertical handover for heterogeneous wireless networks. Journal of Network and Computer Applications, 36(4):1131– 1139, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000192.

Talay:2013:SAR

[1035] A. Cagatay Talay and D. Turgay Altilar. Self adaptive routing for dynamic spectrum access in cognitive radio networks. Journal of Network and Computer Applications, 36(4):1140–1151, July 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000167.

Astorga:2013:HPL

[1036] Jasone Astorga, Marina Aguado, Nerea Toledo, and Marivi Higuero. A high performance link layer mobility management strategy for professional private broadband networks. Journal of Network and Computer Applications, 36(4):1152–1163, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000143.

Tseng:2013:CCA

[1037] Chien-Chao Tseng, Chia-Liang Lin, Li-Hsing Yen, Jyun-Yan Liu, and Cheng-Yuan Ho. Can: a context-aware NAT traversal scheme. Journal of Network and Computer Applications, 36(4):1164–1173, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000398.

Geeta:2013:FTW

[1038] D. D. Geeta, N. Nalini, and Rajashekhar C. Biradar. Fault tolerance in wireless sensor network using hand-off and dynamic power adjustment approach. Journal of Network and Computer Applications, 36(4): 1174–1185, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451300043X.

Lee:2013:DTC

[1039] Chao-Yang Lee, Liang-Cheng Shiu, Fu-Tian Lin, and Chu-Sing Yang. Distributed topology control algorithm on broadcasting in wireless sensor network. Journal of Network and Computer Applications, 36(4): 1186–1195, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000489.

El-Alfy:2013:PBH

[1040] El-Sayed M. El-Alfy, Syed N. Mujahid, and Shokri Z. Selim. A Pareto-based hybrid multiobjective evolutionary approach for constrained multipath traffic engineering optimization in MPLS/GMPLS networks. Journal of Network and Computer Applications, 36(4):1196– 1207, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000465.

Lin:2013:BBS

[1041] Ying-Dar Lin, Cheng-Yuan Ho, Yuan-Cheng Lai, Tzu-Hsiung Du, and Shun-Lee Chang. Booting, browsing and streaming time profiling, and bottleneck analysis on Android-based systems. Journal of Network and Computer Applications, 36(4):1208–1218, July 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000623.

Saif:2013:SME

[1042] Anwar Saif and Mohamed Othman. SRA-MSDU: Enhanced A-MSDU frame aggregation with selective retransmission in 802.11n wireless networks. Journal of Network and Computer Applications, 36(4):1219– 1229, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000611.

Alizai:2013:PWN

[1043] Muhammad Hamad Alizai, Hanno Wirtz, Bernhard Kirchen, and Klaus Wehrle. Portable wireless-networking protocol evaluation. Journal of Network and Computer Applications, 36(4):1230–1242, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451300060X.

Masdari:2013:ASL

[1044] Mohammad Masdari, Sadegh Mohammadzadeh Bazarchi, and Moazam Bidaki. Analysis of Secure LEACH-Based Clustering Protocols in Wireless Sensor Networks. Journal of Network and Computer Applications, 36(4):1243–1260, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512002640.

Bakshi:2013:APF

[1045] Aishwarya Bakshi, Amir Talaei-Khoei, and Pradeep Ray. Adaptive policy framework: a systematic review. Journal of Network and Computer Applications, 36(4):1261–1271, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002548.

Anonymous:2013:EBd

[1046] Anonymous. Editorial board. Journal of Network and Computer Applications, 36(4):ifc, July 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001288.

Rodrigues:2013:GEVb

[1047] Joel J. P. C. Rodrigues, Russel Hsing, Min Chen, Bingli Jiao, and Binod Vaidya. Guest editorial on vehicular communications and applications. Journal of Network and Computer Applications, 36(5):1273–1274, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001513.

Campolo:2013:AVR

[1048] C. Campolo, H. A. Cozzetti, A. Molinaro, and R. Scopigno. Augmenting Vehicle-to-Roadside connectivity in multi-channel vehicular Ad Hoc Networks. Journal of Network and Computer Applications, 36 (5):1275–1286, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451200094X.

Liu:2013:IMD

[1049] Yazhi Liu, Jianwei Niu, Jian Ma, Lei Shu, Takahiro Hara, and Wendong Wang. The insights of message delivery delay in VANETs with a bidirectional traffic model. Journal of Network and Computer Applications, 36(5):1287–1294, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000240.

Bauza:2013:TCD

[1050] R. Bauza and J. Gozalvez. Traffic congestion detection in large-scale scenarios using vehicle-to-vehicle communications. Journal of Network and Computer Applications, 36(5):1295–1307, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000628.

Maslekar:2013:CAT

[1051] Nitin Maslekar, Joseph Mouzna, Mounir Boussedjra, and Houda Labiod. CATS: an adaptive traffic signal system based on car-to-car communication. Journal of Network and Computer Applications, 36(5):1308–1315, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512001415.

Lin:2013:RAT

[1052] Kai Lin. Research on adaptive target tracking in vehicle sensor networks. Journal of Network and Computer Applications, 36(5):1316–1323, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451200015X.

Toledo:2013:PEU

[1053] Nerea Toledo, Jean Marie Bonnin, and Marivi Higuero. Performance evaluation of user applications in the ITS scenario: an analytical assessment of the NeMHIP. Journal of Network and Computer Applications, 36(5):1324–1336, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804512000604.

Ganan:2013:CCC

[1054] Carlos Gañán, Jose L. Muñoz, Oscar Esparza, Jorge Mata-Díaz, Juan Hernández-Serrano, and Juanjo Alins. COACH: COllaborative certificate stAtus CHecking mechanism for VANETs. Journal of Network and Computer Applications, 36(5):1337–1351, September 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512000616.

Ying:2013:PPB

[1055] Bidi Ying, Dimitrios Makrakis, and Hussein T. Mouftah. Privacy preserving broadcast message authentication protocol for VANETs. Journal of Network and Computer Applications, 36(5):1352–1364, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804512001439.

Li:2013:ESC

[1056] Xiong Li, Jianwei Niu, Muhammad Khurram Khan, and Junguo Liao. An enhanced smart card based remote user password authentication scheme. Journal of Network and Computer Applications, 36(5):1365– 1371, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000726.

Luu:2013:CRO

[1057] Hai Van Luu and Xueyan Tang. Constructing rings overlay for robust data collection in wireless sensor networks. Journal of Network and Computer Applications, 36(5):1372–1386, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000714.

Chen:2013:PMA

[1058] Ting Chen, Xiao song Zhang, Hong yuan Li, Dong Wang, and Yue Wu. Propagation modeling of active P2P worms based on ternary matrix. Journal of Network and Computer Applications, 36(5):1387–1394, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000702.

Rashid:2013:DPB

[1059] Sulma Rashid, Abdul Hanan Abdullah, Qaisar Ayub, and M. Soperi Mohd Zahid. Dynamic Prediction based Multi Queue (DPMQ) drop policy for probabilistic routing protocols of delay tolerant network. Journal of Network and Computer Applications, 36(5):1395–1402, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000684.

Anonymous:2013:EBe

[1060] Anonymous. Editorial board. Journal of Network and Computer Applications, 36(5):ifc, September 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001537.

Shen:2013:CTA

[1061] Weiming Shen and Weidong Li. Collaboration technologies and applications. Journal of Network and Computer Applications, 36(6):1403– 1404, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002178.

Latre:2013:ACD

[1062] Steven Latré, Jeroen Famaey, John Strassner, and Filip De Turck. Automated context dissemination for autonomic collaborative networks through semantic subscription filter generation. Journal of Network and Computer Applications, 36(6):1405–1417, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000209.

Xu:2013:RAK

[1063] Yuanhao Xu, Guofu Yin, Yong Nie, Zhengwei He, and Xiaolin Dai. Research on an active knowledge push service based on collaborative intent capture. Journal of Network and Computer Applications, 36 (6):1418–1430, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001136.

Wu:2013:DCP

[1064] Tao Wu, Xiaobing Pei, Yongzhong Lu, Changqing Chen, and Liang Gao. A distributed collaborative product design environment based on semantic norm model and role-based access control. Journal of Network and Computer Applications, 36(6):1431–1440, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000568.

Trappey:2013:IPR

[1065] Amy J. C. Trappey, Charles V. Trappey, Chun-Yi Wu, Chin Yuan Fan, and Yi-Liang Lin. Intelligent patent recommendation system for innovative design collaboration. Journal of Network and Computer Applications, 36(6):1441–1450, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000738.

Soares:2013:DCK

[1066] Diego Carvalho Soares, Flávia Maria Santoro, and Fernanda Araujo Baião. Discovering collaborative knowledge-intensive processes through e-mail mining. Journal of Network and Computer Applications, 36 (6):1451–1465, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000453.

Antunes:2013:EUA

[1067] Pedro Antunes, David Simões, Luis Carriço, and José A. Pino. An enduser approach to business process modeling. Journal of Network and Computer Applications, 36(6):1466–1479, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451300101X.

Iqbal:2013:IOU

[1068] Rahat Iqbal, Nazaraf Shah, Anne James, and Tomasz Cichowicz. Integration, optimization and usability of enterprise applications. Journal of Network and Computer Applications, 36(6):1480–1488, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000477.

Fu:2013:MPA

[1069] Leijie Fu, Pingyu Jiang, and Wei Cao. Modeling and performance analysis of product development process network. Journal of Network and Computer Applications, 36(6):1489–1502, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000441.

Lu:2013:CDA

[1070] Ta-Ping Lu, Amy J. C. Trappey, Yi-Kuang Chen, and Yu-Da Chang. Collaborative design and analysis of supply chain network management key processes model. Journal of Network and Computer Applications, 36(6):1503–1511, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001021.

Cheng:2013:GUR

[1071] Yuan Cheng, Fazhi He, Xiantao Cai, and Dejun Zhang. A group Undo/Redo method in 3D collaborative modeling systems with performance evaluation. Journal of Network and Computer Applications, 36(6):1512–1522, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001239.

Shi:2013:CEM

[1072] Jiancheng Shi, Jianhua Liu, Ruxin Ning, and Weiwei Hou. A collisions evaluation method in virtual environment for collaborative assembly. Journal of Network and Computer Applications, 36(6):1523–1530, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451300012X.

Guo:2013:OIE

[1073] Bin Guo, Daqing Zhang, Zhu Wang, Zhiwen Yu, and Xingshe Zhou. Opportunistic IoT: Exploring the harmonious interaction between human and the Internet of Things. Journal of Network and Computer Applications, 36(6):1531–1539, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000052.

Yang:2013:NAP

[1074] Yang Yang, Xinyu Li, Liang Gao, and Xinyu Shao. A new approach for predicting and collaborative evaluating the cutting force in face milling based on gene expression programming. Journal of Network and Computer Applications, 36(6):1540–1550, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000428.

Gallardo:2013:MDT

[1075] Jesús Gallardo, Ana I. Molina, Crescencio Bravo, and Miguel A. Redondo. A model-driven and task-oriented method for the development of collaborative systems. Journal of Network and Computer Applications, 36(6):1551–1565, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001033.

Ibanez:2013:CLM

[1076] María Blanca Ibáñez, José J. García Rueda, David Maroto, and Carlos Delgado Kloos. Collaborative learning in multi-user virtual environments. Journal of Network and Computer Applications, 36(6):1566– 1576, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000040.

Shen:2013:CCT

[1077] Weiming Shen and Weidong Li. Collaboration computing technologies and applications. Journal of Network and Computer Applications, 36 (6):1577–1578, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451300218X.

Gu:2013:VIV

[1078] Dujuan Gu, Xiaohan Liu, Gang Qin, Shuangjian Yan, Ze luo, and Baoping Yan. VNET6: IPv6 virtual network for the collaboration between applications and networks. Journal of Network and Computer Applications, 36(6):1579–1588, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000490.

Rajesh:2013:IUC

[1079] Anbazhagan Rajesh and Rangaswamy Nakkeeran. Investigation on uplink collaborative contention-based bandwidth request for WiMAX three hop relay networks. Journal of Network and Computer Applications, 36(6):1589–1598, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000106.

Pan:2013:CPC

[1080] Yuanyuan Pan and Jianqing Li. Cooperative pseudonym change scheme based on the number of neighbors in VANETs. Journal of Network and Computer Applications, 36(6):1599–1609, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000416.

Castella:2013:ALP

[1081] Damià Castellà, Francesc Giné, Francesc Solsona, and Josep Lluis Lérida. Analyzing locality over a P2P computing architecture. Journal of Network and Computer Applications, 36(6):1610–1619, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451300009X.

Rius:2013:CSM

[1082] Josep Rius, Fernando Cores, and Francesc Solsona. Cooperative scheduling mechanism for large-scale peer-to-peer computing systems. Journal of Network and Computer Applications, 36(6):1620–1631, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000118.

Yeh:2013:ESA

[1083] Sheng-Cheng Yeh, Ming-Yang Su, Hui-Hui Chen, and Chun-Yuen Lin. An efficient and secure approach for a cloud collaborative editing. Journal of Network and Computer Applications, 36(6):1632–1641, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001409.

Wu:2013:DCE

[1084] Hongchen Wu, Xinjun Wang, Zhaohui Peng, and Qingzhong Li. Divclustering: Exploring active users for social collaborative recommendation. Journal of Network and Computer Applications, 36(6):1642–1650, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000544.

Li:2013:DAB

[1085] Ho Leung Li and Vincent T. Y. Ng. Discovering associations between news and contents in social network sites with the D-Miner service framework. Journal of Network and Computer Applications, 36 (6):1651–1659, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001240.

Wang:2013:NSW

[1086] Xiaogang Wang, Ming Yang, and Junzhou Luo. A novel sequential watermark detection model for efficient traceback of secret network attack flows. Journal of Network and Computer Applications, 36 (6):1660–1670, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000246.

Jiang:2013:NAB

[1087] Jian Jiang, Wei Li, Junzhou Luo, and Jing Tan. A network accountability based verification mechanism for detecting inter-domain routing path inconsistency. Journal of Network and Computer Applications, 36 (6):1671–1683, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000064.

Wang:2013:ABH

[1088] JunPing Wang, Qiliang Zhu, and Yan Ma. An agent-based hybrid service delivery for coordinating Internet of Things and 3rd party service providers. Journal of Network and Computer Applications, 36 (6):1684–1695, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001252.

Li:2013:RVS

[1089] Wenfeng Li, Ye Zhong, Xun Wang, and Yulian Cao. Resource virtualization and service selection in cloud logistics. Journal of Network and Computer Applications, 36(6):1696–1704, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451300057X.

Seyedzadegan:2013:ZDA

[1090] Mojtaba Seyedzadegan, Mohamed Othman, Borhanuddin Mohd Ali, and Shamala Subramaniam. Zero-degree algorithm for Internet GateWay placement in backbone wireless mesh networks. Journal of Network and Computer Applications, 36(6):1705–1723, November 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000696.

Hsu:2013:VNM

[1091] Wu-Hsiao Hsu and Yuh-Pyng Shieh. Virtual network mapping algorithm in the cloud infrastructure. Journal of Network and Computer Applications, 36(6):1724–1734, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000660.

Botero:2013:NPA

[1092] Juan Felipe Botero, Miguel Molina, Xavier Hesselbach-Serra, and José Roberto Amazonas. A novel paths algebra-based strategy to flexibly solve the link mapping stage of VNE problems. Journal of Network and Computer Applications, 36(6):1735–1752, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000672.

Anonymous:2013:EBf

[1093] Anonymous. Editorial board. Journal of Network and Computer Applications, 36(6):ifc, November 2013. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002208.

Lim:2014:AMS

[1094] Hyuk Lim, James Chang Wu Yu, and Binod Vaidya. Advances in multimedia services over emerging wireless networks. Journal of Network and Computer Applications, 37(??):1, January 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002415.

De:2014:EMD

[1095] Pradipta De, Nilanjan Banerjee, and Swades De. Exploiting multiple description coding for intermediate recovery in wireless mesh networks. Journal of Network and Computer Applications, 37(??):2–11, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000865.

Neves:2014:MCA

[1096] Paulo A. Neves, Joel J. P. C. Rodrigues, Min Chen, and Athanasios V. Vasilakos. A multi-channel architecture for IPv6-enabled wireless sensor and actuator networks featuring PnP support. Journal of Network and Computer Applications, 37(??):12–24, January 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804511000877.

Lee:2014:MCA

[1097] Soocheol Lee, Seungmin Rho, and Jong Hyuk Park. Multimedia contents adaptation by modality conversion with user preference in wireless network. Journal of Network and Computer Applications, 37(??): 25–32, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000889.

Yoon:2014:PAV

[1098] Hayoung Yoon, JongWon Kim, and Robert Hsieh. Peer-assisted video on-demand streaming system in practical WiFi-based mobile opportunistic networks. Journal of Network and Computer Applications, 37(??): 33–44, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511001524.

Ganzha:2014:AOC

[1099] Maria Ganzha and Marcin Paprzycki. Agent-oriented computing for distributed systems and networks. Journal of Network and Computer Applications, 37(??):45–46, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000982.

Kim:2014:EBC

[1100] Dongsoo Kim, Eric Noel, and K. Wendy Tang. Expanded Borel Cayley Graphs (Ex-BCGs): a novel communication topology for multi-agent systems. Journal of Network and Computer Applications, 37(??):47– 61, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002615.

Srivastava:2014:SRM

[1101] Shashank Srivastava and G. C. Nandi. Self-reliant mobile code: a new direction of agent security. Journal of Network and Computer Applications, 37(??):62–75, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000131.

Komenda:2014:DIM

[1102] Antonín Komenda, Peter Novák, and Michal Pechoucek. Domainindependent multi-agent plan repair. Journal of Network and Computer Applications, 37(??):76–88, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804512002585.

Adabi:2014:NFN

[1103] Sepideh Adabi, Ali Movaghar, Amir Masoud Rahmani, Hamid Beigy, and Hengameh Dastmalchy-Tabrizi. A new fuzzy negotiation protocol for grid resource allocation. Journal of Network and Computer Applications, 37(??):89–126, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000076.

Fortino:2014:ABA

[1104] Giancarlo Fortino, Wilma Russo, and Marino Vaccaro. An agentbased approach for the design and analysis of content delivery networks. Journal of Network and Computer Applications, 37(??):127– 145, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804512002469.

Carrera:2014:RLA

[1105] Álvaro Carrera, Carlos A. Iglesias, Javier García-Algarra, and Dusan Kolarík. A real-life application of multi-agent systems for fault diagnosis in the provision of an Internet business service. Journal of Network and Computer Applications, 37(??):146–154, January 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804512002457.

Niu:2014:IRC

[1106] Wenjia Niu, Gang Li, Endong Tong, Xinghua Yang, Liang Chang, Zhongzhi Shi, and Song Ci. Interaction relationships of caches in agent-based HD video surveillance: Discovery and utilization. Journal of Network and Computer Applications, 37(??):155–169, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000179.

Galan-Jimenez:2014:UBI

[1107] Jaime Galán-Jiménez and Alfonso Gazo-Cervero. Using bio-inspired algorithms for energy levels assessment in energy efficient wired communication networks. Journal of Network and Computer Applications, 37(??):171–185, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000659.

Yu:2014:DPH

[1108] Jiguo Yu, Qingbo Zhang, Dongxiao Yu, Congcong Chen, and Guanghui Wang. Domatic partition in homogeneous wireless sensor networks. Journal of Network and Computer Applications, 37(??):186–193, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000635.

Hoplaros:2014:DSN

[1109] Demetris Hoplaros, Zahir Tari, and Ibrahim Khalil. Data summarization for network traffic monitoring. Journal of Network and Computer Applications, 37(??):194–205, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000593.

Chiang:2014:DMS

[1110] Chi-Tao Chiang, Hung-Chang Chen, Wen-Hwa Liao, and Kuei-Ping Shih. A decentralized minislot scheduling protocol (DMSP) in TDMAbased wireless mesh networks. Journal of Network and Computer Applications, 37(??):206–215, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000556.

Rezaee:2014:HHA

[1111] Abbas Ali Rezaee, Mohammad Hossein Yaghmaee, Amir Masoud Rahmani, and Amir Hossein Mohajerzadeh. HOCA: Healthcare aware optimized congestion avoidance and control protocol for wireless sensor networks. Journal of Network and Computer Applications, 37(??):216– 228, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000520.

Chao:2014:DSF

[1112] Chih-Min Chao and Tzu-Ying Hsiao. Design of structure-free and energy-balanced data aggregation in wireless sensor networks. Journal of Network and Computer Applications, 37(??):229–239, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000519.

Zare:2014:CCI

[1113] Sajjad Zare and Akbar Ghaffarpour Rahbar. Congestion control in IPTV over PON using Digital Fountain forward error correction. Journal of Network and Computer Applications, 37(??):240–252, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000507.

Li:2014:PBR

[1114] Feng Li, Li Wang, and Weidang Lu. Potential bargaining for resource allocation in cognitive relay transmission. Journal of Network and Computer Applications, 37(??):253–258, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513000581.

Peng:2014:FTR

[1115] Yuhuai Peng, Qingyang Song, Yao Yu, and Fei Wang. Fault-tolerant routing mechanism based on network coding in wireless mesh networks. Journal of Network and Computer Applications, 37(??):259– 272, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000532.

Xu:2014:NCS

[1116] Yuemei Xu, Yang Li, Tao Lin, Zihou Wang, Wenjia Niu, Hui Tang, and Song Ci. A novel cache size optimization scheme based on manifold learning in Content Centric Networking. Journal of Network and Computer Applications, 37(??):273–281, January 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000763.

Bahi:2014:SCI

[1117] Jacques M. Bahi, Xiaole Fang, Christophe Guyeux, and Qianxue Wang. Suitability of chaotic iterations schemes using XORshift for security applications. Journal of Network and Computer Applications, 37(??):282– 292, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000751.

Cameron:2014:IBA

[1118] Conor Cameron, Ibrahim Khalil, and Zahir Tari. An ID-based approach to the caching and distribution of peer-to-peer, proxy-based video content. Journal of Network and Computer Applications, 37(??):293– 314, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000799.

Zheng:2014:DPO

[1119] Wei Zheng, Tao Su, Haijun Zhang, Wei Li, Xiaoli Chu, and Xiangming Wen. Distributed power optimization for spectrum-sharing femtocell networks: a fictitious game approach. Journal of Network and Computer Applications, 37(??):315–322, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513000787.

Li:2014:BFB

[1120] Guilin Li, Longjiang Guo, Xing Gao, and Minghong Liao. Bloom filter based processing algorithms for the multi-dimensional event query in wireless sensor networks. Journal of Network and Computer Applications, 37(??):323–333, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000775.

Niu:2014:BAP

[1121] Jianwei Niu, Wenfang Song, and Mohammed Atiquzzaman. Bandwidthadaptive partitioning for distributed execution optimization of mobile applications. Journal of Network and Computer Applications, 37(??): 334–347, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000945.

Phan:2014:AXA

[1122] Binh Viet Phan and Eric Pardede. Active XML (AXML) research: Survey on the representation, system architecture, data exchange mechanism and query evaluation. Journal of Network and Computer Applications, 37(??):348–364, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513000404.

Singh:2014:VSA

[1123] Harjit Pal Singh, Sarabjeet Singh, J. Singh, and S. A. Khan. VoIP: State of art for global connectivity — a critical review. Journal of Network and Computer Applications, 37(??):365–379, January 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000647.

Al-Sultan:2014:CSV

[1124] Saif Al-Sultan, Moath M. Al-Doori, Ali H. Al-Bayatti, and Hussien Zedan. A comprehensive survey on vehicular Ad Hoc network. Journal of Network and Computer Applications, 37(??):380–392, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451300074X.

Anonymous:2014:EBa

[1125] Anonymous. Editorial board. Journal of Network and Computer Applications, 37(??):ifc, January 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002439.

You:2014:ATH

[1126] Ilsun You, Marek R. Ogiela, A. Min Tjoa, and Dongwan Shin. Advanced technologies for homeland defense and security. Journal of Network and Computer Applications, 38(??):1–2, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002646.

Lilien:2014:SSA

[1127] Leszek T. Lilien, Lotfi ben Othmane, Pelin Angin, Andrew DeCarlo, Raed M. Salih, and Bharat Bhargava. A simulation study of ad hoc networking of UAVs with opportunistic resource utilization networks. Journal of Network and Computer Applications, 38(??):3–15, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001173.

Choi:2014:TAD

[1128] Dongjin Choi, Byeongkyu Ko, Heesun Kim, and Pankoo Kim. Text analysis for detecting terrorism-related articles on the web. Journal of Network and Computer Applications, 38(??):16–21, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001215.

Castiglione:2014:BBC

[1129] Aniello Castiglione, Roberto De Prisco, Alfredo De Santis, Ugo Fiore, and Francesco Palmieri. A botnet-based command and control approach relying on swarm intelligence. Journal of Network and Computer Applications, 38(??):22–33, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001161.

Ogiela:2014:CSB

[1130] Lidia Ogiela and Marek R. Ogiela. Cognitive systems and bio-inspired computing in homeland security. Journal of Network and Computer Applications, 38(??):34–42, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513001197.

Seo:2014:DMM

[1131] Seung-Hyun Seo, Aditi Gupta, Asmaa Mohamed Sallam, Elisa Bertino, and Kangbin Yim. Detecting mobile malware threats to homeland security through static analysis. Journal of Network and Computer Applications, 38(??):43–53, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001227.

Irgan:2014:LCP

[1132] Kerem Irgan, Cem Ünsalan, and Sebnem Baydere. Low-cost prioritization of image blocks in wireless sensor networks for border surveillance. Journal of Network and Computer Applications, 38(??):54–64, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451300163X.

Catuogno:2014:TVF

[1133] Luigi Catuogno, Hans Löhr, Marcel Winandy, and Ahmad-Reza Sadeghi. A trusted versioning file system for passive mobile storage devices. Journal of Network and Computer Applications, 38(??):65–75, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513001203.

Li:2014:SAM

[1134] Zang Li, Chao-Hsien Chu, and Wen Yao. A semantic authorization model for pervasive healthcare. Journal of Network and Computer Applications, 38(??):76–87, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001653.

Chen:2014:MIT

[1135] Hsing-Chung Chen. A multi-issued tag key agreement with time constraint for homeland defense sub-department in NFC environment. Journal of Network and Computer Applications, 38(??):88–98, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001185.

Park:2014:ISC

[1136] Jong-Yeon Park, Dong-Guk Han, Okyeon Yi, and JeongNyeo Kim. An improved side channel attack using event information of subtraction. Journal of Network and Computer Applications, 38(??):99–105, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451300115X.

Halder:2014:EWS

[1137] Subir Halder and Sipra Das Bit. Enhancement of wireless sensor network lifetime by deploying heterogeneous nodes. Journal of Network and Computer Applications, 38(??):106–124, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513000957.

Gao:2014:BTP

[1138] Sheng Gao, Jianfeng Ma, Cong Sun, and Xinghua Li. Balancing trajectory privacy and data utility using a personalized anonymization model. Journal of Network and Computer Applications, 38(??):125– 134, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000970.

Jayasinghe:2014:EER

[1139] Gaya K. Jayasinghe, J. Shane Culpepper, and Peter Bertok. Efficient and effective realtime prediction of drive-by download attacks. Journal of Network and Computer Applications, 38(??):135–149, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513000969.

Luu:2014:EAS

[1140] Hai Van Luu and Xueyan Tang. An efficient algorithm for scheduling sensor data collection through multi-path routing structures. Journal of Network and Computer Applications, 38(??):150–162, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513001008.

Doraghinejad:2014:CAM

[1141] Mohammad Doraghinejad, Hossein Nezamabadi-pour, and Ali Mahani. Channel assignment in multi-radio wireless mesh networks using an improved gravitational search algorithm. Journal of Network and Computer Applications, 38(??):163–171, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513001100.

Doudou:2014:SCB

[1142] Messaoud Doudou, Djamel Djenouri, Nadjib Badache, and Abdelmadjid Bouabdallah. Synchronous contention-based MAC protocols for delay-sensitive wireless sensor networks: a review and taxonomy. Journal of Network and Computer Applications, 38(??):172–184, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513000994.

Guo:2014:SIR

[1143] Wenjing Guo and Wei Zhang. A survey on intelligent routing protocols in wireless sensor networks. Journal of Network and Computer Applications, 38(??):185–201, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513001045.

Anonymous:2014:EBb

[1144] Anonymous. Editorial board. Journal of Network and Computer Applications, 38(??):ifc, February 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451300266X.

Borges:2014:ASI

[1145] Hélder Pereira Borges, José Neuman de Souza, Bruno Schulze, and Antonio Roberto Mury. Automatic services instantiation based on a process specification. Journal of Network and Computer Applications, 39(??):1–16, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001082.

Alsahag:2014:FUB

[1146] Ali Mohammed Alsahag, Borhanuddin Mohd Ali, Nor Kamariah Noordin, and Hafizal Mohamad. Fair uplink bandwidth allocation and latency guarantee for mobile WiMAX using fuzzy adaptive deficit round robin. Journal of Network and Computer Applications, 39(??):17– 25, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001070.

Huang:2014:DAC

[1147] Ding-Jie Huang and Wei-Chung Teng. A defense against clock skew replication attacks in wireless sensor networks. Journal of Network and Computer Applications, 39(??):26–37, March 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001069.

Tong:2014:BFB

[1148] Endong Tong, Wenjia Niu, Gang Li, Ding Tang, Liang Chang, Zhongzhi Shi, and Song Ci. Bloom filter-based workflow management to enable QoS guarantee in wireless sensor networks. Journal of Network and Computer Applications, 39(??):38–51, March 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001057.

Yang:2014:DSP

[1149] Feng-Ming Yang, Wei-Mei Chen, and Jean-Lien C. Wu. A dynamic strategy for packet scheduling and bandwidth allocation based on channel quality in IEEE 802.16e OFDMA system. Journal of Network and Computer Applications, 39(??):52–60, March 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001112.

Ning:2014:CEB

[1150] Zhaolong Ning, Qingyang Song, Yang Huang, and Lei Guo. A channel estimation based opportunistic scheduling scheme in wireless bidirectional networks. Journal of Network and Computer Applications, 39(??): 61–69, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001148.

Safa:2014:RTC

[1151] Haidar Safa, Wassim El-Hajj, and Hanan Zoubian. A robust topology control solution for the sink placement problem in WSNs. Journal of Network and Computer Applications, 39(??):70–82, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001124.

Meng:2014:ABB

[1152] Yuxin Meng and Lam-For Kwok. Adaptive blacklist-based packet filter with a statistic-based approach in network intrusion detection. Journal of Network and Computer Applications, 39(??):83–92, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001264.

Shivshankar:2014:STM

[1153] Smitha Shivshankar and Abbas Jamalipour. Spatio-temporal multicast grouping for content-based routing in vehicular networks: a distributed approach. Journal of Network and Computer Applications, 39(??):93– 103, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001392.

Tan:2014:DBD

[1154] Huailiang Tan, Lianjun Huang, Zaihong He, Youyou Lu, and Xubin He. DMVL: an I/O bandwidth dynamic allocation method for virtual networks. Journal of Network and Computer Applications, 39(??):104– 116, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001380.

Moussaoui:2014:LSQ

[1155] Ali Moussaoui, Fouzi Semchedine, and Abdallah Boukerram. A linkstate QoS routing protocol based on link stability for Mobile Ad hoc Networks. Journal of Network and Computer Applications, 39(??):117– 125, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001422.

Benslimane:2014:RAS

[1156] Abderrahim Benslimane and Abderrezak Rachedi. Rate adaptation scheme for IEEE 802.11-based MANETs. Journal of Network and Computer Applications, 39(??):126–139, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513001410.

Budyal:2014:AAB

[1157] V. R. Budyal and S. S. Manvi. ANFIS and agent based bandwidth and delay aware anycast routing in mobile ad hoc networks. Journal of Network and Computer Applications, 39(??):140–151, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451300146X.

Tounsi:2014:KKE

[1158] Wiem Tounsi, Nora Cuppens-Boulahia, Joaquin Garcia-Alfaro, Yannick Chevalier, and Frédéric Cuppens. KEDGEN2: a key establishment and derivation protocol for EPC Gen2 RFID systems. Journal of Network and Computer Applications, 39(??):152–166, March 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001458.

Chadagorn:2014:PMB

[1159] Aukrit Chadagorn, Ibrahim Khalil, Conor Cameron, and Zahir Tari. PileCast: Multiple bit rate live video streaming over BitTorrent. Journal of Network and Computer Applications, 39(??):167–178, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513001446.

Hassan:2014:ESR

[1160] Osama Al-Haj Hassan, Thamer Al-Rousan, Anas Abu Taleb, and Adi Maaita. An efficient and scalable ranking technique for mashups involving RSS data sources. Journal of Network and Computer Applications, 39(??):179–190, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001483.

Diallo:2014:SFR

[1161] Ousmane Diallo, Joel J. P. C. Rodrigues, Mbaye Sene, and Jaime Lloret. Simulation framework for real-time database on WSNs. Journal of Network and Computer Applications, 39(??):191–201, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001471.

Sung:2014:VBC

[1162] Tien-Wen Sung and Chu-Sing Yang. Voronoi-based coverage improvement approach for wireless directional sensor networks. Journal of Network and Computer Applications, 39(??):202–213, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001501.

Mastroianni:2014:SOP

[1163] Carlo Mastroianni and Giuseppe Papuzzo. A self-organizing P2P framework for collective service discovery. Journal of Network and Computer Applications, 39(??):214–222, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513001495.

Kim:2014:PAD

[1164] Younghyun Kim, Haneul Ko, Sangheon Pack, Jong-Hyouk Lee, Seok-Joo Koh, and Heeyoung Jung. Performance analysis of distributed mapping system in ID/locator separation architectures. Journal of Network and Computer Applications, 39(??):223–232, March 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001641.

Gisbert:2014:ISC

[1165] J. R. Gisbert, C. Palau, M. Uriarte, G. Prieto, J. A. Palazón, M. Esteve, O. López, J. Correas, M. C. Lucas-Estañ, P. Giménez, A. Moyano, L. Collantes, J. Gozálvez, B. Molina, O. Lázaro, and A. González. Integrated system for control and monitoring industrial wireless networks for labor risk prevention. Journal of Network and Computer Applications, 39(??):233–252, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001744.

Ulucinar:2014:EPC

[1166] Alper Rifat Ulucinar, Ibrahim Korpeoglu, and Ezhan Karasan. Effects of physical channel separation on application flows in a multi-radio multi-hop wireless mesh network: an experimental study on BilMesh testbed. Journal of Network and Computer Applications, 39(??):253– 265, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001689.

Birje:2014:RRB

[1167] Mahantesh N. Birje, Sunilkumar S. Manvi, and Sajal K. Das. Reliable resources brokering scheme in wireless grids based on non-cooperative bargaining game. Journal of Network and Computer Applications, 39(??): 266–279, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001677.

Lu:2014:FCA

[1168] Shilei Lu and Shunzheng Yu. A fuzzy k-coverage approach for RFID network planning using plant growth simulation algorithm. Journal of Network and Computer Applications, 39(??):280–291, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451300177X.

Tang:2014:DAS

[1169] Jenn Tang. Designing an Anti-swindle Mahjong Leisure Prototype System using RFID and ontology theory. Journal of Network and Computer Applications, 39(??):292–301, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513001720.

Ayub:2014:CQB

[1170] Qaisar Ayub, Sulma Rashid, M. Soperi Mohd Zahid, and Abdul Hanan Abdullah. Contact quality based forwarding strategy for delay tolerant network. Journal of Network and Computer Applications, 39(??): 302–309, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001719.

Liu:2014:ACO

[1171] Xuxun Liu and Desi He. Ant colony optimization with greedy migration mechanism for node deployment in wireless sensor networks. Journal of Network and Computer Applications, 39(??):310– 318, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001707.

Gavalas:2014:MRS

[1172] Damianos Gavalas, Charalampos Konstantopoulos, Konstantinos Mastakas, and Grammati Pantziou. Mobile recommender systems in tourism. Journal of Network and Computer Applications, 39(??):319– 333, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001094.

Kakkasageri:2014:IMV

[1173] M. S. Kakkasageri and S. S. Manvi. Information management in vehicular ad hoc networks: a review. Journal of Network and Computer Applications, 39(??):334–350, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001434.

Modares:2014:SSP

[1174] Hero Modares, Amirhossein Moravejosharieh, Jaime Lloret, and Rosli Salleh. A survey of secure protocols in Mobile IPv6. Journal of Network and Computer Applications, 39(??):351–368, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001732.

Durrani:2014:VCR

[1175] Muhammad Nouman Durrani and Jawwad A. Shamsi. Volunteer computing: requirements, challenges, and solutions. Journal of Network and Computer Applications, 39(??):369–380, March 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001665.

Anonymous:2014:EBc

[1176] Anonymous. Editorial board. Journal of Network and Computer Applications, 39(??):ifc, March 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000137.

Tang:2014:EEH

[1177] Jine Tang, ZhangBing Zhou, Jianwei Niu, and Qun Wang. An energy efficient hierarchical clustering index tree for facilitating time-correlated region queries in the Internet of Things. Journal of Network and Computer Applications, 40(??):1–11, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513001690.

Fraile:2014:SMC

[1178] Francisco Fraile and Juan Carlos Guerri. Simple models of the content duration and the popularity of television content. Journal of Network and Computer Applications, 40(??):12–20, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001859.

Hu:2014:RVI

[1179] Haiyang Hu, Yizhi Ren, Xu Xu, Liguo Huang, and Hua Hu. Reducing view inconsistency by predicting avatars’ motion in multi-server distributed virtual environments. Journal of Network and Computer Applications, 40(??):21–30, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001860.

Varadharajan:2014:CSA

[1180] Vijay Varadharajan and Udaya Tupakula. Counteracting security attacks in virtual machines in the cloud using property based attestation. Journal of Network and Computer Applications, 40(??):31– 45, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001768.

Farzinvash:2014:MRM

[1181] Leili Farzinvash and Mehdi Dehghan. Multi-rate multicast routing in multi-gateway multi-channel multi-radio wireless mesh networks. Journal of Network and Computer Applications, 40(??):46– 60, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001847.

Yin:2014:NAC

[1182] Peng-Yeng Yin, Ray-I. Chang, Chih-Chiang Chao, and Yen-Ting Chu. Niched ant colony optimization with colony guides for QoS multicast routing. Journal of Network and Computer Applications, 40(??):61– 72, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001781.

Arab:2014:CDS

[1183] Mohsen Arab and Mohsen Afsharchi. Community detection in social networks using hybrid merging of sub-communities. Journal of Network and Computer Applications, 40(??):73–84, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001835.

Elfaki:2014:CCP

[1184] Mohamed Ahmed Elfaki, Hamidah Ibrahim, Ali Mamat, Mohamed Othman, and Haidar Safa. Collaborative caching priority for processing requests in MANETs. Journal of Network and Computer Applications, 40(??):85–96, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001884.

Wang:2014:RIU

[1185] Jianwei Wang, Chen Jiang, and Jianfei Qian. Robustness of Internet under targeted attack: a cascading failure perspective. Journal of Network and Computer Applications, 40(??):97–104, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001823.

Park:2014:DEM

[1186] Sehoon Park, Qichen Chen, Hyuck Han, and Heon Y. Yeom. Design and evaluation of mobile offloading system for web-centric devices. Journal of Network and Computer Applications, 40(??):105– 115, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001811.

Zhao:2014:TEE

[1187] Peng Zhao, Xinyu Yang, Wei Yu, Chiyong Dong, Shusen Yang, and Sulabh Bhattarai. Toward efficient estimation of available bandwidth for IEEE 802.11-based wireless networks. Journal of Network and Computer Applications, 40(??):116–125, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451300180X.

Xia:2014:MIA

[1188] Yang Xia and Chai Kiat Yeo. Mobile Internet access over intermittent network connectivity. Journal of Network and Computer Applications, 40(??):126–138, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513001872.

Wu:2014:EVM

[1189] Zhen-Yu Wu, Ju-Chuan Wu, Sung-Chiang Lin, and Charlotte Wang. An electronic voting mechanism for fighting bribery and coercion. Journal of Network and Computer Applications, 40(??):139–150, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804513001999.

Abdelaziz:2014:MBT

[1190] Samar Abdelaziz and Mustafa ElNainay. Metric-based taxonomy of routing protocols for cognitive radio ad hoc networks. Journal of Network and Computer Applications, 40(??):151–163, April 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001896.

Yilmaz:2014:LFE

[1191] Onur Yilmaz, Orhan Dagdeviren, and Kayhan Erciyes. Localizationfree and energy-efficient hole bypassing techniques for fault-tolerant sensor networks. Journal of Network and Computer Applications, 40(??):

164–178, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001902.

Quick:2014:GDF

[1192] Darren Quick and Kim-Kwang Raymond Choo. Google Drive: Forensic analysis of data remnants. Journal of Network and Computer Applications, 40(??):179–193, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002051.

Arshadi:2014:BLB

[1193] Laleh Arshadi and Amir Hossein Jahangir. Benford’s law behavior of Internet traffic. Journal of Network and Computer Applications, 40(0): 194–205, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001951.

Zeng:2014:OCS

[1194] Lingfang Zeng and Yang Wang. Optimization on content service with local search in cloud of clouds. Journal of Network and Computer Applications, 40(??):206–215, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451300194X.

Li:2014:EBC

[1195] Ke Li, Sheng Wang, Shizhong Xu, Xiong Wang, Haojun Huang, Jin Ren, and Bo Zhai. Evaluating the benefit of the core-edge separation on intradomain traffic engineering under uncertain traffic demand. Journal of Network and Computer Applications, 40(??):216– 226, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001938.

Choi:2014:EDT

[1196] Hyun-Ho Choi, Ki-Ho Lee, and Jung-Ryun Lee. Energy-delay tradeoff analysis of user state transition mechanism for mobile web services. Journal of Network and Computer Applications, 40(??):227– 233, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001926.

Guan:2014:PPT

[1197] Xinjie Guan and Baek-Young Choi. Push or pull? Toward optimal content delivery using cloud storage. Journal of Network and Computer Applications, 40(??):234–243, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513001914.

Macia-Perez:2014:FFM

[1198] Francisco Macia-Perez, Iren Lorenzo-Fonseca, and Jose Vicente Berna-Martinez. A formal framework for modelling complex network management systems. Journal of Network and Computer Applications, 40(??): 255–269, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002063.

Abreu:2014:EAR

[1199] Carlos Abreu, Manuel Ricardo, and P. M. Mendes. Energy-aware routing for biomedical wireless sensor networks. Journal of Network and Computer Applications, 40(??):270–278, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451300204X.

Ciobanu:2014:JHB

[1200] R. I. Ciobanu, D. G. Reina, C. Dobre, S. L. Toral, and P. Johnson. JDER: a history-based forwarding scheme for delay tolerant networks using Jaccard distance and encountered ration. Journal of Network and Computer Applications, 40(??):279–291, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513002002.

Jiang:2014:TDB

[1201] Dingde Jiang, Zhengzheng Xu, Peng Zhang, and Ting Zhu. A transform domain-based anomaly detection approach to network-wide traffic. Journal of Network and Computer Applications, 40(??):292– 306, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002038.

Hoque:2014:NAT

[1202] N. Hoque, Monowar H. Bhuyan, R. C. Baishya, D. K. Bhattacharyya, and J. K. Kalita. Network attacks: Taxonomy, tools and systems. Journal of Network and Computer Applications, 40(??):307– 324, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001756.

Whaiduzzaman:2014:SVC

[1203] Md Whaiduzzaman, Mehdi Sookhak, Abdullah Gani, and Rajkumar Buyya. A survey on vehicular cloud computing. Journal of Network and Computer Applications, 40(??):325–344, April 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001793.

Abolfazli:2014:RMA

[1204] Saeid Abolfazli, Zohreh Sanaei, Abdullah Gani, Feng Xia, and Laurence T. Yang. Rich Mobile Applications: Genesis, taxonomy, and open issues. Journal of Network and Computer Applications, 40(??): 345–362, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513001975.

Sharef:2014:VCA

[1205] Baraa T. Sharef, Raed A. Alsaqour, and Mahamod Ismail. Vehicular communication ad hoc routing protocols: a survey. Journal of Network and Computer Applications, 40(??):363–396, April 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513001963.

Anonymous:2014:EBd

[1206] Anonymous. Editorial board. Journal of Network and Computer Applications, 40(??):ifc, April 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000344.

Al-Shammary:2014:DAF

[1207] Dhiah Al-Shammary, Ibrahim Khalil, and Zahir Tari. A distributed aggregation and fast fractal clustering approach for SOAP traffic. Journal of Network and Computer Applications, 41(??):1–14, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002026.

Qin:2014:NCD

[1208] Tao Qin, Xiaohong Guan, Wei Li, Pinghui Wang, and Min Zhu. A new connection degree calculation and measurement method for large scale network monitoring. Journal of Network and Computer Applications, 41(??):15–26, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002130.

Tuna:2014:UAV

[1209] Gurkan Tuna, Bilel Nefzi, and Gianpaolo Conte. Unmanned aerial vehicle-aided communications system for disaster recovery. Journal of Network and Computer Applications, 41(??):27–36, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002075.

Gao:2014:URA

[1210] Lijun Gao, Maode Ma, Yantai Shu, and Yuhua Wei. An ultralightweight RFID authentication protocol with CRC and permutation. Journal of Network and Computer Applications, 41(??):37–46, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002269.

Wu:2014:IMP

[1211] Tin-Yu Wu, Wei-Tsong Lee, Nadra Guizani, and Tzu-Ming Wang. Incentive mechanism for P2P file sharing based on social network and game theory. Journal of Network and Computer Applications, 41(??): 47–55, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002117.

Anacleto:2014:MAP

[1212] Ricardo Anacleto, Lino Figueiredo, Ana Almeida, and Paulo Novais. Mobile application to provide personalized sightseeing tours. Journal of Network and Computer Applications, 41(??):56–64, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002105.

Castro-Schez:2014:EAL

[1213] J. J. Castro-Schez, M. A. Redondo, F. Jurado, and J. Albusac. Experience applying language processing techniques to develop educational software that allow active learning methodologies by advising students. Journal of Network and Computer Applications, 41(??):65– 79, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002166.

Abbas:2014:FDC

[1214] Ahmed Mohammed Abbas, Azuraliza Abu Bakar, and Mohd Zakree Ahmad. Fast dynamic clustering SOAP messages based compression and aggregation model for enhanced performance of Web services. Journal of Network and Computer Applications, 41(??):80–88, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002154.

Sardar:2014:NET

[1215] Bhaskar Sardar, Debashis Saha, and Mahbub Hassan. A novel enhancement of TCP for on-board IP networks with wireless cellular connectivity. Journal of Network and Computer Applications, 41(??):89– 100, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002129.

Liu:2014:AEE

[1216] Wei Liu, Wei Du, Jing Chen, Wei Wang, and GuoSun Zeng. Adaptive energy-efficient scheduling algorithm for parallel tasks on homogeneous clusters. Journal of Network and Computer Applications, 41(??): 101–113, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002142.

Wang:2014:PBE

[1217] Jinzhu Wang, Jianxin Liao, Tonghong Li, Jing Wang, Jingyu Wang, and Qi Qi. Probe-based end-to-end overload control for networks of SIP servers. Journal of Network and Computer Applications, 41(??): 114–125, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002609.

Haralabopoulos:2014:RTE

[1218] Giannis Haralabopoulos and Ioannis Anagnostopoulos. Real time enhanced random sampling of online social networks. Journal of Network and Computer Applications, 41(??):126–134, May 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002592.

Alghamdi:2014:APU

[1219] Saleh Alghamdi, Ron van Schyndel, and Ibrahim Khalil. Accurate positioning using long range active RFID technology to assist visually impaired people. Journal of Network and Computer Applications, 41(??):135–147, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002580.

Atif:2014:NRM

[1220] Yacine Atif and Liren Zhang. Network resource management in support of QoS in ubiquitous learning. Journal of Network and Computer Applications, 41(??):148–156, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002610.

Li:2014:DSA

[1221] Xiaoyuan Li, Dexiang Wang, Janise McNair, and Jianmin Chen. Dynamic spectrum access with packet size adaptation and residual energy balancing for energy-constrained cognitive radio sensor networks. Journal of Network and Computer Applications, 41(??):157– 166, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002403.

Alizai:2014:PLF

[1222] Muhammad Hamad Alizai and Klaus Wehrle. Probabilistic locationfree addressing in wireless networks. Journal of Network and Computer Applications, 41(??):167–181, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513002622.

Pastrana:2014:RAR

[1223] Sergio Pastrana, Agustin Orfila, Juan E. Tapiador, and Pedro Peris-Lopez. Randomized Anagram revisited. Journal of Network and Computer Applications, 41(??):182–196, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513002841.

Ahmad:2014:RTN

[1224] Tahir Ahmad and Usman Younis. Randomness testing of noncryptographic hash functions for real-time hash table based storage and look-up of URLs. Journal of Network and Computer Applications, 41(??):197–205, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002853.

Verbelen:2014:ADC

[1225] Tim Verbelen, Pieter Simoens, Filip De Turck, and Bart Dhoedt. Adaptive deployment and configuration for mobile augmented reality in the cloudlet. Journal of Network and Computer Applications, 41(??):206– 216, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002816.

Raayatpanah:2014:BEE

[1226] M. A. Raayatpanah, H. Salehi Fathabadi, B. H. Khalaj, S. Khodayifar, and P. M. Pardalos. Bounds on end-to-end statistical delay and jitter in multiple multicast coded packet networks. Journal of Network and Computer Applications, 41(??):217–227, May 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451300283X.

Vankeirsbilck:2014:NLH

[1227] Bert Vankeirsbilck, Pieter Simoens, Filip De Turck, Piet Demeester, and Bart Dhoedt. Network latency hiding in thin client systems through server-centric speculative display updating. Journal of Network and Computer Applications, 41(??):228–239, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804513002828.

Ciobanu:2014:SCS

[1228] Radu-Ioan Ciobanu, Ciprian Dobre, Mihai Dascalu, Stefan Trausan-Matu, and Valentin Cristea. SENSE: a collaborative selfish node detection and incentive mechanism for opportunistic networks. Journal of Network and Computer Applications, 41(??):240–249, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000101.

Tran:2014:CVS

[1229] Thang Tran, Maike Kuhnert, and Christian Wietfeld. Cloud voice service as over-the-top solution for seamless voice call continuity in a heterogeneous network environment. Journal of Network and Computer Applications, 41(??):250–262, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000095.

Li:2014:USI

[1230] Yusen Li and Wentong Cai. Update schedules for improving consistency in multi-server distributed virtual environments. Journal of Network and Computer Applications, 41(??):263–273, May 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000083.

Li:2014:CSS

[1231] Fan Li, Jinsheng Sun, Moshe Zukerman, Zhengfei Liu, Qin Xu, Sammy Chan, Guanrong Chen, and King-Tim Ko. A comparative simulation study of TCP/AQM systems for evaluating the potential of neuronbased AQM schemes. Journal of Network and Computer Applications, 41(??):274–299, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451400006X.

Gupta:2014:ETA

[1232] Govind P. Gupta, Manoj Misra, and Kumkum Garg. Energy and trust aware mobile agent migration protocol for data aggregation in wireless sensor networks. Journal of Network and Computer Applications, 41(??):300–311, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000046.

Moudi:2014:GAM

[1233] Mehrnaz Moudi and Mohamed Othman. A GreedyZero algorithm to minimise the conflicts in an Optical Multistage Interconnection Network. Journal of Network and Computer Applications, 41(??):312– 318, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000022.

Rubio-Largo:2014:PAM

[1234] Álvaro Rubio-Largo, Miguel A. Vega-Rodríguez, and David L. González-Álvarez. Performance assessment of multiobjective approaches in optical Traffic Grooming. Journal of Network and Computer Applications, 41(??):319–350, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000253.

Yi:2014:ASD

[1235] Li Yi, Huachun Zhou, Daochao Huang, and Hongke Zhang. An analytical study of distributed mobility management schemes with a flow duration based model. Journal of Network and Computer Applications, 41(??):351–357, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000265.

Zeng:2014:SOL

[1236] Lingfang Zeng, Bharadwaj Veeravalli, and Qingsong Wei. Space4time: Optimization latency-sensitive content service in cloud. Journal of Network and Computer Applications, 41(??):358–368, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000289.

Utsumi:2014:NHP

[1237] Satoshi Utsumi and Salahuddin Muhammad Salim Zabir. A new high-performance TCP friendly congestion control over wireless networks. Journal of Network and Computer Applications, 41(??):369– 378, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000290.

Song:2014:JPC

[1238] Qingyang Song, Zhaolong Ning, Yang Huang, Lei Guo, and Xiaobing Lu. Joint power control and spectrum access in cognitive radio networks. Journal of Network and Computer Applications, 41(??):379– 388, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000319.

Navimipour:2014:RDM

[1239] Nima Jafari Navimipour, Amir Masoud Rahmani, Ahmad Habibizad Navin, and Mehdi Hosseinzadeh. Resource discovery mechanisms in grid systems: a survey. Journal of Network and Computer Applications, 41(??):389–410, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804513002014.

Li:2014:SPI

[1240] Ying Li and Radim Bartos. A survey of protocols for Intermittently Connected Delay-Tolerant Wireless Sensor Networks. Journal of Network and Computer Applications, 41(??):411–423, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804513002087.

Manvi:2014:RMI

[1241] Sunilkumar S. Manvi and Gopal Krishna Shyam. Resource management for Infrastructure as a Service (IaaS) in cloud computing: a survey. Journal of Network and Computer Applications, 41(??):424– 440, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002099.

Lu:2014:ERM

[1242] Ssu-Hsuan Lu, Kuan-Ching Li, Kuan-Chou Lai, and Yeh-Ching Chung. Effectiveness of a replica mechanism to improve availability with Arrangement Graph-Based Overlay. Journal of Network and Computer Applications, 41(??):441–450, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804513002634.

Abdaoui:2014:TDB

[1243] Abderrazak Abdaoui and Tarek M. El-Fouly. TOSSIM and distributed binary consensus algorithm in wireless sensor networks. Journal of Network and Computer Applications, 41(??):451–458, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000113.

Shaddad:2014:SAT

[1244] R. Q. Shaddad, A. B. Mohammad, S. A. Al-Gailani, A. M. Al-hetar, and M. A. Elmagzoub. A survey on access technologies for broadband optical and wireless networks. Journal of Network and Computer Applications, 41(??):459–472, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000058.

Ashfaq:2014:ITF

[1245] Ayesha Binte Ashfaq, Sajjad Rizvi, Mobin Javed, Syed Ali Khayam, Muhammad Qasim Ali, and Ehab Al-Shaer. Information theoretic feature space slicing for statistical anomaly detection. Journal of Network and Computer Applications, 41(??):473–487, May 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000034.

Adhikari:2014:MSV

[1246] Ratnadip Adhikari. A meticulous study of various medium access control protocols for wireless sensor networks. Journal of Network and Computer Applications, 41(??):488–504, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514000241.

Friginal:2014:ANW

[1247] Jesús Friginal, Juan-Carlos Ruiz, David de Andrés, and Antonio Bustos. Ambient noise in wireless mesh networks: Evaluation and proposal of an adaptive algorithm to mitigate link removal. Journal of Network and Computer Applications, 41(??):505–516, May 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000307.

Zin:2014:RPD

[1248] Shazana Md Zin, Nor Badrul Anuar, Miss Laiha Mat Kiah, and Al-Sakib Khan Pathan. Routing protocol design for secure WSN: Review and open research issues. Journal of Network and Computer Applications, 41(??):517–530, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000460.

Anonymous:2014:EBe

[1249] Anonymous. Editorial board. Journal of Network and Computer Applications, 41(??):ifc, May 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451400068X.

Jia:2014:ECA

[1250] Jinyuan Jia, Wei Wang, and Xiaojun Hei. An efficient caching algorithm for peer-to-peer 3D streaming in distributed virtual environments. Journal of Network and Computer Applications, 42(??):1– 11, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451400054X.

Xu:2014:SAQ

[1251] Qin Xu and Jinsheng Sun. A simple active queue management based on the prediction of the packet arrival rate. Journal of Network and Computer Applications, 42(??):12–20, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514000563.

Soorki:2014:LSP

[1252] Mehdi Naderi Soorki and Habib Rostami. Label switched protocol routing with guaranteed bandwidth and end to end path delay in MPLS networks. Journal of Network and Computer Applications, 42(??):21– 38, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000617.

Mateu:2014:CCB

[1253] Víctor Mateu, Francesc Sebé, and Magda Valls. Constructing credentialbased E-voting systems from offline E-coin protocols. Journal of Network and Computer Applications, 42(??):39–44, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000629.

Pirozmand:2014:HMO

[1254] Poria Pirozmand, Guowei Wu, Behrouz Jedari, and Feng Xia. Human mobility in opportunistic networks: Characteristics, models and prediction methods. Journal of Network and Computer Applications, 42(??): 45–58, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000587.

Mehdizadeh:2014:LDM

[1255] Abbas Mehdizadeh, Fazirulhisyam Hashim, and Mohamed Othman. Lightweight decentralized multicast-unicast key management method in wireless IPv6 networks. Journal of Network and Computer Applications, 42(??):59–69, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000812.

Han:2014:IMM

[1256] Guangjie Han, Jia Chao, Chenyu Zhang, Lei Shu, and Qingwu Li. The impacts of mobility models on DV-hop based localization in Mobile Wireless Sensor Networks. Journal of Network and Computer Applications, 42(??):70–79, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000824.

Shakibian:2014:CVE

[1257] Hadi Shakibian and Nasrollah Moghadam Charkari. In-cluster vector evaluated particle swarm optimization for distributed regression in WSNs. Journal of Network and Computer Applications, 42(??):80– 91, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000526.

Akkari:2014:DCP

[1258] Nadine Akkari, Ghadah Aldabbagh, Michel Nahas, and John Cioffi. Dynamic Clustering Protocol for coordinated tethering over cellular networks. Journal of Network and Computer Applications, 42(??):92– 101, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000654.

Shamshirband:2014:CFC

[1259] Shahaboddin Shamshirband, Nor Badrul Anuar, Miss Laiha Mat Kiah, Vala Ali Rohani, Dalibor Petković, Sanjay Misra, and Abdul Nasir Khan. Co-FAIS: Cooperative fuzzy artificial immune system for detecting intrusion in wireless sensor networks. Journal of Network and Computer Applications, 42(??):102–117, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514000666.

Espada:2014:ITS

[1260] Jordán Pascual Espada, Ronald R. Yager, and Bin Guo. Internet of Things: Smart things network and communication. Journal of Network and Computer Applications, 42(??):118–119, June 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000514.

Yan:2014:STM

[1261] Zheng Yan, Peng Zhang, and Athanasios V. Vasilakos. A survey on trust management for Internet of Things. Journal of Network and Computer Applications, 42(??):120–134, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514000575.

Xie:2014:RSF

[1262] Wei Xie, Lei Xie, Chen Zhang, Qiang Wang, Jian Xu, Quan Zhang, and Chaojing Tang. RFID seeking: Finding a lost tag rather than only detecting its missing. Journal of Network and Computer Applications, 42(??):135–142, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000071.

Ma:2014:CAB

[1263] Lianbo Ma, Kunyuan Hu, Yunlong Zhu, and Hanning Chen. Cooperative artificial bee colony algorithm for multi-objective RFID network planning. Journal of Network and Computer Applications, 42(??):143– 162, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000551.

Lakhlef:2014:OLT

[1264] Hicham Lakhlef, Hakim Mabed, and Julien Bourgeois. Optimization of the logical topology for mobile MEMS networks. Journal of Network and Computer Applications, 42(??):163–177, June 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000642.

Cueva-Fernandez:2014:VES

[1265] Guillermo Cueva-Fernandez, Jordán Pascual Espada, Vicente García-Díaz, Cristian González García, and Nestor Garcia-Fernandez. Vitruvius: an expert system for vehicle sensor tracking and managing application generation. Journal of Network and Computer Applications, 42(??):178–188, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000605.

Sun:2014:HBB

[1266] Jiajun Sun and Huadong Ma. Heterogeneous-belief based incentive schemes for crowd sensing in mobile social networks. Journal of Network and Computer Applications, 42(??):189–196, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000538.

Anonymous:2014:EBf

[1267] Anonymous. Editorial board. Journal of Network and Computer Applications, 42(??):ifc, June 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000988.

Saleem:2014:PRU

[1268] Yasir Saleem and Mubashir Husain Rehmani. Primary radio user activity models for cognitive radio networks: a survey. Journal of Network and Computer Applications, 43(??):1–16, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000848.

Ma:2014:NSR

[1269] Dong Ma and Maode Ma. Network selection and resource allocation for multicast in HetNets. Journal of Network and Computer Applications, 43(??):17–26, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000836.

Lin:2014:EAT

[1270] Gongqi Lin, Sieteng Soh, Kwan-Wu Chin, and Mihai Lazarescu. Energy Aware Two Disjoint Paths Routing. Journal of Network and Computer Applications, 43(??):27–41, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514000873.

Xu:2014:GTS

[1271] Zheng Xu, Yunhuai Liu, Lin Mei, Chuanping Hu, and Lan Chen. Generating temporal semantic context of concepts using web search engines. Journal of Network and Computer Applications, 43(??):42– 55, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451400085X.

Fu:2014:DSA

[1272] JuiHsi Fu, PoChing Lin, and SingLing Lee. Detecting spamming activities in a campus network using incremental learning. Journal of Network and Computer Applications, 43(??):56–65, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000630.

Alrshah:2014:CSH

[1273] Mohamed A. Alrshah, Mohamed Othman, Borhanuddin Ali, and Zurina Mohd Hanapi. Comparative study of high-speed Linux TCP variants over high-BDP networks. Journal of Network and Computer Applications, 43(??):66–75, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000903.

Wu:2014:AAE

[1274] Xiaoping Wu, GuoYing Wang, Dan Dai, and Mengjun Tong. Accurate acoustic energy-based localization with beacon position uncertainty in wireless sensor networks. Journal of Network and Computer Applications, 43(??):76–83, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000861.

Gani:2014:RIM

[1275] Abdullah Gani, Golam Mokatder Nayeem, Muhammad Shiraz, Mehdi Sookhak, Md Whaiduzzaman, and Suleman Khan. A review on interworking and mobility techniques for seamless connectivity in mobile cloud computing. Journal of Network and Computer Applications, 43(??):84–102, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000927.

daSilva:2014:MPM

[1276] Eduardo da Silva and Luiz Carlos P. Albini. Middleware proposals for mobile ad hoc networks. Journal of Network and Computer Applications, 43(??):103–120, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000939.

Sookhak:2014:RRD

[1277] Mehdi Sookhak, Hamid Talebian, Ejaz Ahmed, Abdullah Gani, and Muhammad Khurram Khan. A review on remote data auditing in single cloud server: Taxonomy and open issues. Journal of Network and Computer Applications, 43(??):121–141, August 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000940.

Yeo:2014:LCS

[1278] Hui-Shyong Yeo, Xiao-Shen Phang, Hoon-Jae Lee, and Hyotaek Lim. Leveraging client-side storage techniques for enhanced use of multiple consumer cloud storage services on resource-constrained mobile devices. Journal of Network and Computer Applications, 43(??):142– 156, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000897.

Anonymous:2014:EBg

[1279] Anonymous. Editorial board. Journal of Network and Computer Applications, 43(??):ifc, August 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001167.

Jabbarpour:2014:CLC

[1280] Mohammad Reza Jabbarpour, Rafidah Md Noor, Rashid Hafeez Khokhar, and Chih-Heng Ke. Cross-layer congestion control model for urban vehicular environments. Journal of Network and Computer Applications, 44(??):1–16, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514001106.

Wu:2014:NSA

[1281] Jiyan Wu, Bo Cheng, Yanlei Shang, Jun Huang, and Junliang Chen. A novel scheduling approach to concurrent multipath transmission of high definition video in overlay networks. Journal of Network and Computer Applications, 44(??):17–29, September 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451400112X.

ElMougy:2014:CAA

[1282] Amr El Mougy, Aymen Kamoun, Mohamed Ibnkahla, Saïd Tazi, and Khalil Drira. A context and application-aware framework for resource management in dynamic collaborative wireless M2M networks. Journal of Network and Computer Applications, 44(??):30–45, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000964.

Iqbal:2014:DDM

[1283] Anindya Iqbal and Manzur Murshed. On demand-driven movement strategy for moving beacons in sensor localization. Journal of Network and Computer Applications, 44(??):46–62, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000885.

Tian:2014:PRC

[1284] Guo-Song Tian, Yu-Chu Tian, and Colin Fidge. Precise relative clock synchronization for distributed control using TSC registers. Journal of Network and Computer Applications, 44(??):63–71, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451400109X.

Fernandez:2014:SII

[1285] Víctor Fernández, Juan Manuel Orduña, and Pedro Morillo. Server implementations for improving the performance of CAR systems based on mobile phones. Journal of Network and Computer Applications, 44(??):72–82, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514000952.

Sidek:2014:DMM

[1286] Khairul Azami Sidek, Vu Mai, and Ibrahim Khalil. Data mining in mobile ECG based biometric identification. Journal of Network and Computer Applications, 44(??):83–91, September 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000915.

Jiang:2014:DRB

[1287] Guoxing Jiang, Jun Chen, and Yanqing Shen. Delivery ratio- and buffered time-constrained: Multicasting for Delay Tolerant Networks. Journal of Network and Computer Applications, 44(??):92–105, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514001131.

Lin:2014:CPF

[1288] Ying-Dar Lin, Ekarat Rattagan, Yuan-Cheng Lai, Li-Pin Chang, Yun-Chien Yo, Cheng-Yuan Ho, and Shun-Lee Chang. Calibrating parameters and formulas for process-level energy consumption profiling in smartphones. Journal of Network and Computer Applications, 44 (??):106–119, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001118.

Niu:2014:BAH

[1289] Wenjia Niu, Gang Li, Endong Tong, Quan Z. Sheng, Qian Li, Yue Hu, Athanasios V. Vasilakos, and Li Guo. A block-aware hybrid data dissemination with hotspot elimination in wireless sensor network. Journal of Network and Computer Applications, 44(??):120–133, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001258.

Nacer:2014:SWS

[1290] Hassina Nacer and Djamil Aissani. Semantic web services: Standards, applications, challenges and solutions. Journal of Network and Computer Applications, 44(??):134–151, September 2014. CODEN JNCAF3.

ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514001143.

Fang:2014:AOA

[1291] Weiwei Fang, Zhulin An, Lei Shu, Qingyu Liu, Yongjun Xu, and Yuan An. Achieving optimal admission control with dynamic scheduling in energy constrained network systems. Journal of Network and Computer Applications, 44(??):152–160, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514001283.

Wu:2014:TRT

[1292] Jiyan Wu, Yanlei Shang, Chau Yuen, Bo Cheng, and Junliang Chen. TRADER: a reliable transmission scheme to video conferencing applications over the Internet. Journal of Network and Computer Applications, 44(??):161–171, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451400126X.

Ji:2014:SNN

[1293] Changqing Ji, Zhiyang Li, Wenyu Qu, Yujie Xu, and Yuanyuan Li. Scalable nearest neighbor query processing based on Inverted Grid Index. Journal of Network and Computer Applications, 44(??):172–182, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001295.

Madani:2014:GRG

[1294] Shaahin Madani and Ibrahim Khalil. Garbled Routing (GR): a generic framework towards unification of anonymous communication systems. Journal of Network and Computer Applications, 44(??):183–195, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001246.

Lalbakhsh:2014:UDA

[1295] Pooia Lalbakhsh, Bahram Zaeri, and Yi-Ping Phoebe Chen. Using Dead Ants to improve the robustness and adaptability of AntNet routing algorithm. Journal of Network and Computer Applications, 44 (??):196–211, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001271.

Misra:2014:LAB

[1296] Sudip Misra, P. Venkata Krishna, V. Saritha, Harshit Agarwal, and Aditya Ahuja. Learning automata-based multi-constrained faulttolerance approach for effective energy management in smart grid communication network. Journal of Network and Computer Applications, 44(??):212–219, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001301.

Haque:2014:SBI

[1297] Israat Tanzeena Haque. A sensor based indoor localization through fingerprinting. Journal of Network and Computer Applications, 44 (??):220–229, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001313.

Anonymous:2014:EBh

[1298] Anonymous. Editorial board. Journal of Network and Computer Applications, 44(??):ifc, September 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001696.

Bradai:2014:ECD

[1299] Abbas Bradai, Toufik Ahmed, Raouf Boutaba, and Reaz Ahmed. Efficient content delivery scheme for layered video streaming in large-scale networks. Journal of Network and Computer Applications, 45(??):1– 14, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001441.

Chen:2014:ASR

[1300] Jenhui Chen, Zhuxiu Yuan, and Lei Wang. An apropos signal report and adaptive period (ASAP) scheme for fast handover in the fourthgeneration wireless networks. Journal of Network and Computer Applications, 45(??):15–26, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001325.

Huang:2014:OJS

[1301] An Huang and Bing Hu. The optimal joint sequence design in the feedback-based two-stage switch. Journal of Network and Computer Applications, 45(??):27–34, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514001477.

Huang:2014:MOC

[1302] Jun Huang, Guoquan Liu, and Qiang Duan. On modeling and optimization for composite network-Cloud service provisioning. Journal of Network and Computer Applications, 45(??):35–43, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001507.

Qadir:2014:MCR

[1303] Junaid Qadir, Adeel Baig, Asad Ali, and Quratulain Shafi. Multicasting in cognitive radio networks: Algorithms, techniques and protocols. Journal of Network and Computer Applications, 45(??):44– 61, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001659.

Kushwaha:2014:CCH

[1304] Vandana Kushwaha and Ratneshwer Gupta. Congestion control for highspeed wired network: a systematic literature review. Journal of Network and Computer Applications, 45(??):62–78, October 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001453.

Yau:2014:CAC

[1305] Kok-Lim Alvin Yau, Nordin Ramli, Wahidah Hashim, and Hafizal Mohamad. Clustering algorithms for Cognitive Radio networks: a survey. Journal of Network and Computer Applications, 45(??):79–95, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001611.

Husain:2014:BDV

[1306] Mohammad Iftekhar Husain, Steven Y. Ko, Steve Uurtamo, Atri Rudra, and Ramalingam Sridhar. Bidirectional data verification for cloud storage. Journal of Network and Computer Applications, 45(??):96–107, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001386.

Garg:2014:SBV

[1307] Saurabh Kumar Garg, Adel Nadjaran Toosi, Srinivasa K. Gopalaiyengar, and Rajkumar Buyya. SLA-based virtual machine management for heterogeneous workloads in a cloud datacenter. Journal of Network and Computer Applications, 45(??):108–120, October 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001787.

Zhu:2014:SER

[1308] Yan Zhu, Di Ma, Changjun Hu, Gail-Joon Ahn, and Hongxin Hu. Secure and efficient random functions with variable-length output. Journal of Network and Computer Applications, 45(??):121–133, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001817.

Sun:2014:CSS

[1309] Le Sun, Hai Dong, Farookh Khadeer Hussain, Omar Khadeer Hussain, and Elizabeth Chang. Cloud service selection: State-of-the-art and future research directions. Journal of Network and Computer Applications, 45(??):134–150, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451400160X.

Singh:2014:AIA

[1310] Gurpreet Singh, Neeraj Kumar, and Anil Kumar Verma. ANTALG: an Innovative ACO based Routing Algorithm for MANETs. Journal of Network and Computer Applications, 45(??):151–167, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514001465.

Benhamida:2014:AFD

[1311] Fatima Zohra Benhamida, Yacine Challal, and Mouloud Koudil. Adaptive failure detection in low power lossy wireless sensor networks. Journal of Network and Computer Applications, 45(??):168–180, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001763.

Dondo:2014:DFB

[1312] Julio Dondo, Felix Villanueva, David Garcia, David Vallejo, Carlos Glez-Morcillo, and Juan Carlos Lopez. Distributed FPGA-based architecture to support indoor localisation and orientation services. Journal of Network and Computer Applications, 45(??):181–190, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001775.

Rizk:2014:CCM

[1313] Rawya Rizk, Walaa Saber, and Hany Harb. Conditional clustered matrix factorization based network coordinate system. Journal of Network and Computer Applications, 45(??):191–202, October 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001751.

Jabbarifar:2014:LLI

[1314] Masoume Jabbarifar and Michel Dagenais. LIANA: Live incremental time synchronization of traces for distributed systems analysis. Journal of Network and Computer Applications, 45(??):203–214, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001623.

Saied:2014:SCS

[1315] Yosra Ben Saied, Alexis Olivereau, Djamal Zeghlache, and Maryline Laurent. A survey of collaborative services and security-related issues in modern wireless Ad-Hoc communications. Journal of Network and Computer Applications, 45(??):215–227, October 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001842.

Anonymous:2014:EBi

[1316] Anonymous. Editorial board. Journal of Network and Computer Applications, 45(??):ifc, October 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001957.

Polato:2014:CVH

[1317] Ivanilton Polato, Reginaldo Ré, Alfredo Goldman, and Fabio Kon. A comprehensive view of Hadoop research — a systematic literature review. Journal of Network and Computer Applications, 46(??):1–25, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001635.

Mohamadi:2014:HMM

[1318] Hosein Mohamadi, Shaharuddin Salleh, and Mohd Norsyarizad Razali. Heuristic methods to maximize network lifetime in directional sensor networks with adjustable sensing ranges. Journal of Network and Computer Applications, 46(??):26–35, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514001866.

Armaghani:2014:SST

[1319] Farzaneh R. Armaghani, Iqbal Gondal, Joarder Kamruzzaman, and David G. Green. Sensor selection for tracking multiple groups of targets. Journal of Network and Computer Applications, 46(??):36–47, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001647.

Pham:2014:CPI

[1320] Congduc Pham. Communication performances of IEEE 802.15.4 wireless sensor motes for data-intensive applications: a comparison of Wasp-Mote, Arduino MEGA, TelosB, MicaZ and iMote2 for image surveillance. Journal of Network and Computer Applications, 46(??):48–59, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451400188X.

Safa:2014:PPA

[1321] Haidar Safa, Marcel Karam, and Bassam Moussa. PHAODV: Power aware heterogeneous routing protocol for MANETs. Journal of Network and Computer Applications, 46(??):60–71, November 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001830.

Liu:2014:MGR

[1322] Ning Liu, Xiaoping Li, and Weiming Shen. Multi-granularity resource virtualization and sharing strategies in cloud manufacturing. Journal of Network and Computer Applications, 46(??):72–82, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514001933.

Khalifa:2014:TLP

[1323] Tarek Khalifa, Atef Abdrabou, Khaled Bashir Shaban, Maazen Alsabaan, and Kshirasagar Naik. Transport layer performance analysis and optimization for smart metering infrastructure. Journal of Network and Computer Applications, 46(??):83–93, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001878.

Qin:2014:NLS

[1324] Jing Qin, Hongwei Duan, Huawei Zhao, and Jiankun Hu. A new Lagrange solution to the privacy-preserving general geometric intersection problem. Journal of Network and Computer Applications, 46(??):94– 99, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001908.

Hamdi:2014:ESR

[1325] Omessaad Hamdi, Mohamed Aymen Chalouf, Dramane Ouattara, and Francine Krief. eHealth: Survey on research projects, comparative study of telemonitoring architectures and main issues. Journal of Network and Computer Applications, 46(??):100–112, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001672.

Hakami:2014:EBS

[1326] Vesal Hakami and Mehdi Dehghan. Efficient broadcasting in slow fading wireless ad-hoc networks. Journal of Network and Computer Applications, 46(??):113–123, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514001891.

Chang:2014:RBP

[1327] Ing-Chau Chang, Eric Jui-Lin Lu, Shiuan-Yin Huang, and Yi-Hui Chen. An RDF-based P2P overlay network supporting range and wildcard queries. Journal of Network and Computer Applications, 46(??):124– 138, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451400191X.

Alam:2014:DAS

[1328] Kh Mahmudul Alam, Joarder Kamruzzaman, Gour Karmakar, and Manzur Murshed. Dynamic adjustment of sensing range for event coverage in wireless sensor networks. Journal of Network and Computer Applications, 46(??):139–153, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514001660.

Medetov:2014:DAI

[1329] Seytkamal Medetov, Mohamed Bakhouya, Jaafar Gaber, Khalid Zinedine, Maxime Wack, and Pascal Lorenz. A decentralized approach for information dissemination in Vehicular Ad hoc Networks. Journal of Network and Computer Applications, 46(??):154–165, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514001854.

Farooqi:2014:SNC

[1330] Muhammad Zubair Farooqi, Salma Malik Tabassum, Mubashir Husain Rehmani, and Yasir Saleem. A survey on network coding: From traditional wireless networks to emerging cognitive radio networks. Journal of Network and Computer Applications, 46(??):166–181, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002069.

Sevgi:2014:ODR

[1331] Cüneyt Sevgi and Altan Koçyigit. Optimal deployment in randomly deployed heterogeneous WSNs: a connected coverage approach. Journal of Network and Computer Applications, 46(??):182–197, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S1084804514002082. See corrigendum [1412].

Afsar:2014:CSN

[1332] M. Mehdi Afsar and Mohammad-H. Tayarani-N. Clustering in sensor networks: a literature survey. Journal of Network and Computer Applications, 46(??):198–226, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514002124.

Andreou:2014:NAF

[1333] Panayiotis G. Andreou, Demetrios Zeinalipour-Yazti, George S. Samaras, and Panos K. Chrysanthis. A network-aware framework for energyefficient data acquisition in wireless sensor networks. Journal of Network and Computer Applications, 46(??):227–240, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002094.

Lima:2014:OOA

[1334] David H. S. Lima, Andre L. L. Aquino, Heitor S. Ramos, Eliana S. Almeida, and Joel J. P. C. Rodrigues. OASys: an opportunistic and agile system to detect free on-street parking using intelligent boards embedded in surveillance cameras. Journal of Network and Computer Applications, 46(??):241–249, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002057.

Benslimane:2014:SIQ

[1335] Abderrahim Benslimane, Jiankun Hu, and Xuemin (Sherman) Shen. Special issue on “Quality of service in wireless and mobile networking”. Journal of Network and Computer Applications, 46(??):250–251, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002422.

Pan:2014:LLS

[1336] Meng-Shiuan Pan and Ping-Lin Liu. Low latency scheduling for convergecast in ZigBee tree-based wireless sensor networks. Journal of Network and Computer Applications, 46(??):252–263, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001489.

Kumar:2014:ELA

[1337] Neeraj Kumar, Sudhanshu Tyagi, and Der-Jiunn Deng. LA-EEHSC: Learning automata-based energy efficient heterogeneous selective clustering for wireless sensor networks. Journal of Network and Computer Applications, 46(??):264–279, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514001568.

Shahid:2014:DJR

[1338] Adnan Shahid, Saleem Aslam, Hyung Seok Kim, and Kyung-Geun Lee. Distributed joint resource and power allocation in self-organized femtocell networks: a potential game approach. Journal of Network and Computer Applications, 46(??):280–292, November 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001490.

Bagaa:2014:RMC

[1339] Miloud Bagaa, Mohamed Younis, Adlen Ksentini, and Nadjib Badache. Reliable multi-channel scheduling for timely dissemination of aggregated data in wireless sensor networks. Journal of Network and Computer Applications, 46(??):293–304, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514001556.

Ju:2014:PAE

[1340] Peijian Ju, Wei Song, A-Long Jin, and Dizhi Zhou. Performance analysis and enhancement for a cooperative wireless diversity network with spatially random mobile helpers. Journal of Network and Computer Applications, 46(??):305–314, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514001532.

Tang:2014:CBC

[1341] Ruichun Tang, Yuanzhen Yue, Xiangqian Ding, and Yue Qiu. Credibilitybased cloud media resource allocation algorithm. Journal of Network and Computer Applications, 46(??):315–321, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001593.

Zhang:2014:EFH

[1342] Li Jun Zhang and Samuel Pierre. An enhanced fast handover with seamless mobility support for next-generation wireless networks. Journal of Network and Computer Applications, 46(??):322–335, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451400157X.

Ivesic:2014:CLQ

[1343] K. Ivesic, L. Skorin-Kapov, and M. Matijasevic. Cross-layer QoEdriven admission control and resource allocation for adaptive multimedia services in LTE. Journal of Network and Computer Applications, 46(??):336–351, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002173. See corrigendum [1765].

Rejeb:2014:NRA

[1344] Sonia Ben Rejeb, Nidal Nasser, and Sami Tabbane. A novel resource allocation scheme for LTE network in the presence of mobility. Journal of Network and Computer Applications, 46(??):352–361, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001581.

Bradai:2014:IPA

[1345] Nourchène Bradai, Lamia Chaari Fourati, and Lotfi Kamoun. Investigation and performance analysis of MAC protocols for WBAN networks. Journal of Network and Computer Applications, 46(??):362–373, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002197.

Bahi:2014:EAD

[1346] Jacques M. Bahi, Christophe Guyeux, Mourad Hakem, and Abdallah Makhoul. Epidemiological approach for data survivability in unattended wireless sensor networks. Journal of Network and Computer Applications, 46(??):374–383, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002185.

Liu:2014:EEA

[1347] Zhi Liu, Jingyun Feng, Yusheng Ji, and Yongbing Zhang. EAF: Energy-aware adaptive free viewpoint video wireless transmission. Journal of Network and Computer Applications, 46(??):384–394, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514001519.

Rachedi:2014:HMC

[1348] Abderrezak Rachedi, Hakim Badis, and Abderrahim Benslimane. How MIMO cross-layer design enables QoS while detecting non-cooperative nodes in wireless multi-hop networks. Journal of Network and Computer Applications, 46(??):395–406, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514001520.

Huang:2014:DME

[1349] Tao Huang, Baoliu Ye, Song Guo, and Sanglu Lu. Delay minimization by exploring full-duplex capacity and relay-based cooperative scheduling in WLANs. Journal of Network and Computer Applications, 46 (??):407–417, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001544.

Anonymous:2014:EBj

[1350] Anonymous. Editorial board. Journal of Network and Computer Applications, 46(??):ifc, November 2014. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002446.

Moussaoui:2015:SRP

[1351] Ali Moussaoui and Abdallah Boukeream. A survey of routing protocols based on link-stability in mobile ad hoc networks. Journal of Network and Computer Applications, 47(??):1–10, January 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002148.

Wallace:2015:DSC

[1352] T. Daniel Wallace, Khalim Amjad Meerja, and Abdallah Shami. Ondemand scheduling for concurrent multipath transfer using the stream control transmission protocol. Journal of Network and Computer Applications, 47(??):11–22, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451400215X.

Campobello:2015:CLL

[1353] G. Campobello, O. Giordano, A. Segreto, and S. Serrano. Comparison of local lossless compression algorithms for Wireless Sensor Networks. Journal of Network and Computer Applications, 47(??):23– 31, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002203.

Alsaqour:2015:DPB

[1354] Raed Alsaqour, Maha Abdelhaq, Rashid Saeed, Mueen Uddin, Ola Alsukour, Mohammed Al-Hubaishi, and Tariq Alahdal. Dynamic packet beaconing for GPSR mobile ad hoc position-based routing protocol using fuzzy logic. Journal of Network and Computer Applications, 47(??): 32–46, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002033.

Shiraz:2015:SCA

[1355] Muhammad Shiraz, Mehdi Sookhak, Abdullah Gani, and Syed Adeel Ali Shah. A Study on the Critical Analysis of Computational Offloading Frameworks for Mobile Cloud Computing. Journal of Network and Computer Applications, 47(??):47–60, January 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002100.

Vegni:2015:FSS

[1356] Anna Maria Vegni and Enrico Natalizio. Forwarder smart selection protocol for limitation of broadcast storm problem. Journal of Network and Computer Applications, 47(??):61–71, January 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002045.

Ma:2015:ADQ

[1357] Xiaobo Ma, Junjie Zhang, Zhenhua Li, Jianfeng Li, Jing Tao, Xiaohong Guan, John C. S. Lui, and Don Towsley. Accurate DNS query characteristics estimation via active probing. Journal of Network and Computer Applications, 47(??):72–84, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514002239.

Garcia-Reinoso:2015:SDR

[1358] Jaime Garcia-Reinoso, Norberto Fernández, Ivan Vidal, and Jesús Arias Fisteus. Scalable data replication in content-centric networking based on alias names. Journal of Network and Computer Applications, 47(??): 85–98, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002318.

Weingartner:2015:CRM

[1359] Rafael Weingärtner, Gabriel Beims Bräscher, and Carlos Becker Westphall. Cloud resource management: a survey on forecasting and profiling models. Journal of Network and Computer Applications, 47(??): 99–106, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002252.

Mliki:2015:CSC

[1360] Hela Mliki, Lamia Chaari, and Lotfi Kamoun. A comprehensive survey on Carrier Ethernet Congestion Management mechanism. Journal of Network and Computer Applications, 47(??):107–130, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514002070.

Cacciapuoti:2015:CAM

[1361] Angela Sara Cacciapuoti, Marcello Caleffi, Luigi Paura, and Md Arafatur Rahman. Channel availability for mobile cognitive radio networks. Journal of Network and Computer Applications, 47(??):131–136, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514002306.

Karim:2015:RFL

[1362] Lutful Karim, N. Nasser, Q. H. Mahmoud, Alagan Anpalagan, and T. E. Salti. Range-free localization approach for M2M communication system using mobile anchor nodes. Journal of Network and Computer Applications, 47(??):137–146, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001829.

Halder:2015:DAS

[1363] Subir Halder and Sipra Das Bit. Design of an Archimedes’ spiral based node deployment scheme targeting enhancement of network lifetime in wireless sensor networks. Journal of Network and Computer Applications, 47(??):147–167, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002215.

Anonymous:2015:EBa

[1364] Anonymous. Editorial board. Journal of Network and Computer Applications, 47(??):ifc, January 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002537.

Vecchio:2015:IAC

[1365] Massimo Vecchio and Roberto López-Valcarce. Improving area coverage of wireless sensor networks via controllable mobile nodes: a greedy approach. Journal of Network and Computer Applications, 48(??):1– 13, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002355.

Li:2015:NMR

[1366] Ruiying Li, Jingfu Wang, Haitao Liao, and Ning Huang. A new method for reliability allocation of avionics connected via an airborne network. Journal of Network and Computer Applications, 48(??):14– 21, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002331.

Ruiz-Rivera:2015:NFM

[1367] Alejandro Ruiz-Rivera, Kwan-Wu Chin, and Sieteng Soh. A novel framework to mitigate the negative impacts of green techniques on BGP. Journal of Network and Computer Applications, 48(??):22–34, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002409.

Li:2015:CHB

[1368] Wei Li and Wei Zhang. Coverage hole and boundary nodes detection in wireless sensor networks. Journal of Network and Computer Applications, 48(??):35–43, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002392.

Akhunzada:2015:MEA

[1369] Adnan Akhunzada, Mehdi Sookhak, Nor Badrul Anuar, Abdullah Gani, Ejaz Ahmed, Muhammad Shiraz, Steven Furnell, Amir Hayat, and Muhammad Khurram Khan. Man-at-the-end attacks: Analysis, taxonomy, human aspects, motivation and future directions. Journal of Network and Computer Applications, 48(??):44–57, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002379.

Kabir:2015:RLF

[1370] Muhammad Ashad Kabir, Jun Han, and Bonghee Hong. Reader level filtering for efficient query processing in RFID middleware. Journal of Network and Computer Applications, 48(??):58–70, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002343.

Liao:2015:EEA

[1371] Bin Liao, Jiong Yu, Tao Zhang, Guo Binglei, Sun Hua, and Changtian Ying. Energy-efficient algorithms for distributed storage system based on block storage structure reconfiguration. Journal of Network and Computer Applications, 48(??):71–86, February 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002367.

Cheng:2015:RBR

[1372] Chien-Fu Cheng and Kuo-Tang Tsai. A recursive Byzantine-resilient protocol. Journal of Network and Computer Applications, 48(??):87– 98, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002380.

Liu:2015:APA

[1373] Jieyao Liu, Ejaz Ahmed, Muhammad Shiraz, Abdullah Gani, Rajkumar Buyya, and Ahsan Qureshi. Application partitioning algorithms in mobile cloud computing: Taxonomy, review and future directions. Journal of Network and Computer Applications, 48(??):99– 117, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002161.

Anonymous:2015:EBb

[1374] Anonymous. Editorial board. Journal of Network and Computer Applications, 48(??):ifc, February 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002756.

Meng:2015:PST

[1375] Xianfu Meng, Yanli Wang, and Yue Gong. Perspective of space and time based replica population organizing strategy in unstructured peer-to-peer networks. Journal of Network and Computer Applications, 49(??): 1–14, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002628.

Kong:2015:CSM

[1376] Jia Hao Kong, Li-Minn Ang, and Kah Phooi Seng. A comprehensive survey of modern symmetric cryptographic solutions for resource constrained environments. Journal of Network and Computer Applications, 49(??):15–50, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002136.

Yuan:2015:DFP

[1377] Peiyan Yuan and Ping Liu. Data fusion prolongs the lifetime of mobile sensing networks. Journal of Network and Computer Applications, 49(??):51–59, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002641.

Nawaz:2015:NCN

[1378] Faiza Nawaz and Varun Jeoti. NFRA-C, neighbor friendly reader to reader anti-collision protocol with counters for dense reader environments. Journal of Network and Computer Applications, 49(??):60– 67, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002653.

Wang:2015:LBM

[1379] Xiaonan Wang, Deguang Le, Yufeng Yao, and Conghua Xie. Locationbased mobility support for 6LoWPAN wireless sensor networks. Journal of Network and Computer Applications, 49(??):68–77, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514002719.

Zhang:2015:EDS

[1380] Ying Zhang, Houkuan Huang, Hui He, Jing Teng, and Zhuxiao Wang. Efficient distributed semantic based data and service unified discovery with one-dimensional semantic space. Journal of Network and Computer Applications, 49(??):78–87, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514002690.

Rahmani:2015:DPU

[1381] Amir Masoud Rahmani, Zeinab Fadaie, and Anthony T. Chronopoulos. Data placement using Dewey Encoding in a hierarchical data grid. Journal of Network and Computer Applications, 49(??):88–98, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002707.

Moens:2015:SRN

[1382] Hendrik Moens and Filip De Turck. Shared resource network-aware impact determination algorithms for service workflow deployment with partial cloud offloading. Journal of Network and Computer Applications, 49(??):99–111, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451400263X.

Ashraf:2015:AST

[1383] Qazi Mamoon Ashraf and Mohamed Hadi Habaebi. Autonomic schemes for threat mitigation in Internet of Things. Journal of Network and Computer Applications, 49(??):112–127, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804514002732.

Anonymous:2015:EBc

[1384] Anonymous. Editorial board. Journal of Network and Computer Applications, 49(??):ifc, March 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451500003X.

Daghighi:2015:TSG

[1385] Babak Daghighi, Miss Laiha Mat Kiah, Shahaboddin Shamshirband, and Muhammad Habib Ur Rehman. Toward secure group communication in wireless mobile environments: Issues, solutions, and challenges. Journal of Network and Computer Applications, 50(??):1– 14, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002410.

Saleem:2015:ICR

[1386] Yasir Saleem, Mubashir Husain Rehmani, and Sherali Zeadally. Integration of Cognitive Radio Technology with unmanned aerial vehicles: Issues, opportunities, and future research challenges. Journal of Network and Computer Applications, 50(??):15–31, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002811.

Nie:2015:COB

[1387] Laisen Nie, Dingde Jiang, and Lei Guo. A convex optimizationbased traffic matrix estimation approach in IP-over-WDM backbone networks. Journal of Network and Computer Applications, 50(??):32– 38, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002720.

Zeng:2015:ITC

[1388] Lingfang Zeng, Bharadwaj Veeravalli, and Albert Y. Zomaya. An integrated task computation and data management scheduling strategy for workflow applications in cloud environments. Journal of Network and Computer Applications, 50(??):39–48, April 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000119.

Gonzalez-Manzano:2015:ECP

[1389] Lorena González-Manzano and Agustin Orfila. An efficient confidentialitypreserving Proof of Ownership for deduplication. Journal of Network and Computer Applications, 50(??):49–59, April 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000132.

Loo:2015:SII

[1390] Jonathan Loo and Mahdi Aiash. Special issue on information-centric network architecture, protocols, algorithms and applications. Journal of Network and Computer Applications, 50(??):60–63, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002616.

Loo:2015:CSS

[1391] Jonathan Loo and Mahdi Aiash. Challenges and solutions for secure information centric networks: a case study of the NetInf architecture. Journal of Network and Computer Applications, 50(??):64– 72, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001337.

Aiash:2015:IAA

[1392] Mahdi Aiash and Jonathan Loo. An integrated authentication and authorization approach for the network of information architecture. Journal of Network and Computer Applications, 50(??):73–79, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001349.

Xu:2015:SSI

[1393] Mingwei Xu, Zhongxing Ming, Chunmei Xia, Jia Ji, Dan Li, and Dan Wang. SIONA: a Service and Information Oriented Network Architecture. Journal of Network and Computer Applications, 50(??):80– 91, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451400143X.

Stais:2015:RMT

[1394] Charilaos Stais, George Xylomenos, and Alexios Voulimeneas. A reliable multicast transport protocol for information-centric networks. Journal of Network and Computer Applications, 50(??):92–100, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001398.

Kim:2015:CCA

[1395] Dabin Kim, Sung-Won Lee, Young-Bae Ko, and Jae-Hoon Kim. Cache capacity-aware content centric networking under flash crowds. Journal of Network and Computer Applications, 50(??):101–113, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514001416.

Mendes:2015:CDN

[1396] Paulo Mendes. Combining data naming and context awareness for pervasive networks. Journal of Network and Computer Applications, 50(??): 114–125, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002227.

Alzahrani:2015:SIC

[1397] Bander A. Alzahrani, Martin J. Reed, Janne Riihijärvi, and Vassilios G. Vassilakis. Scalability of information centric networking using mediated topology management. Journal of Network and Computer Applications, 50(??):126–133, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001374.

Qiao:2015:NEW

[1398] Xiuquan Qiao, Guoshun Nan, Yue Peng, Lei Guo, Jingwen Chen, Yunlei Sun, and Junliang Chen. NDNBrowser: an extended web browser for named data networking. Journal of Network and Computer Applications, 50(??):134–147, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001428.

Amadeo:2015:FSN

[1399] Marica Amadeo, Claudia Campolo, and Antonella Molinaro. Forwarding strategies in named data wireless ad hoc networks: Design and evaluation. Journal of Network and Computer Applications, 50(??):148– 158, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001404.

Eum:2015:DII

[1400] S. Eum, Y. Shoji, M. Murata, and N. Nishinaga. Design and implementation of ICN-enabled IEEE 802.11 access points as nano data centers. Journal of Network and Computer Applications, 50(??):159– 167, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001799.

Zhang:2015:SRH

[1401] Lingwei Zhang, Yuhui Deng, Weiheng Zhu, Jipeng Zhou, and Frank Wang. Skewly replicating hot data to construct a power-efficient storage cluster. Journal of Network and Computer Applications, 50(??): 168–179, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514001362.

Anonymous:2015:EBd

[1402] Anonymous. Editorial board. Journal of Network and Computer Applications, 50(??):ifc, April 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000168.

Bravo-Torres:2015:IVL

[1403] Jack Fernando Bravo-Torres, Martín López-Nores, Yolanda Blanco-Fernández, José Juan Pazos-Arias, Manuel Ramos-Cabrer, and Alberto Gil-Solla. An improved virtualization layer to support distribution of multimedia contents in pervasive social applications. Journal of Network and Computer Applications, 51(??):1–17, May 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000120.

Berrocal-Plaza:2015:OMM

[1404] Víctor Berrocal-Plaza, Miguel A. Vega-Rodríguez, and Juan M. Sánchez-Pérez. Optimizing the mobility management task in networks of four world capital cities. Journal of Network and Computer Applications, 51 (??):18–28, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000144.

Mushtaq:2015:QQA

[1405] M. Sajid Mushtaq, Scott Fowler, Abdelhamid Mellouk, and Brice Augustin. QoE/QoS-aware LTE downlink scheduler for VoIP with power saving. Journal of Network and Computer Applications, 51(??):29– 46, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000277.

Huang:2015:MEI

[1406] Kun Huang, Gaogang Xie, Yanbiao Li, and Dafang Zhang. Memoryefficient IP lookup using trie merging for scalable virtual routers. Journal of Network and Computer Applications, 51(??):47–58, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514000459.

Liu:2015:NCC

[1407] Yu Liu, Bin Guo, Chi Zhou, and Yu Cheng. Network-coded cooperative information recovery in cellular/802.11 mobile Networks. Journal of Network and Computer Applications, 51(??):59–67, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514000320.

Liu:2015:TCM

[1408] Linfeng Liu, Ningshen Zhang, and Ye Liu. Topology control models and solutions for signal irregularity in mobile underwater wireless sensor networks. Journal of Network and Computer Applications, 51(??): 68–90, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000496.

Anbazhagan:2015:IEP

[1409] Rajesh Anbazhagan and Nakkeeran Rangaswamy. Investigations on enhanced power saving mechanism for IEEE 802.16m network with heterogeneous traffic. Journal of Network and Computer Applications, 51(??): 91–101, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000484.

Dezfouli:2015:MLP

[1410] Behnam Dezfouli, Marjan Radi, Shukor Abd Razak, Tan Hwee-Pink, and Kamalrulnizam Abu Bakar. Modeling low-power wireless communications. Journal of Network and Computer Applications, 51(??):102– 126, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514000472.

Atiquzzaman:2015:E

[1411] Mohammed Atiquzzaman. Editorial. Journal of Network and Computer Applications, 51(??):127–129, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804514000502.

Sevgi:2015:COD

[1412] Cüneyt Sevgi and Altan Koçyigit. Corrigendum to “Optimal deployment in randomly deployed heterogeneous WSNs: a connected coverage approach” [J. Netw. Comput. Appl. 46, (2014), 182–197]. Journal of Network and Computer Applications, 51(??):130, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451400232X. See [1331].

Anonymous:2015:EBe

[1413] Anonymous. Editorial board. Journal of Network and Computer Applications, 51(??):ifc, May 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000302.

Vasilakos:2015:ICN

[1414] Athanasios V. Vasilakos, Zhe Li, Gwendal Simon, and Wei You. Information centric network: Research challenges and opportunities. Journal of Network and Computer Applications, 52(??):1–10, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000272.

Ahmad:2015:SVM

[1415] Raja Wasim Ahmad, Abdullah Gani, Siti Hafizah Ab. Hamid, Muhammad Shiraz, Abdullah Yousafzai, and Feng Xia. A survey on virtual machine migration and server consolidation frameworks for cloud data centers. Journal of Network and Computer Applications, 52(??): 11–25, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000284.

Lin:2015:TTP

[1416] Ying-Dar Lin, Minh-Tuan Thai, Chih-Chiang Wang, and Yuan-Cheng Lai. Two-tier project and job scheduling for SaaS cloud service providers. Journal of Network and Computer Applications, 52(??):26– 36, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000478.

Thompson:2015:RCL

[1417] Elizabeth A. Thompson, Charles McIntosh, James Isaacs, Eric Harmison, and Ross Sneary. Robot communication link using 802.11n or 900 MHz OFDM. Journal of Network and Computer Applications, 52(??): 37–51, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000430.

Ahmed:2015:AOM

[1418] Ejaz Ahmed, Abdullah Gani, Mehdi Sookhak, Siti Hafizah Ab Hamid, and Feng Xia. Application optimization in mobile cloud computing: Motivation, taxonomies, and open challenges. Journal of Network and Computer Applications, 52(??):52–68, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515000417.

Sun:2015:ELI

[1419] Guodong Sun, Lin Zhao, Zhibo Chen, and Guofu Qiao. Effective link interference model in topology control of wireless ad hoc and sensor networks. Journal of Network and Computer Applications, 52(??):69– 78, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000533.

Ji:2015:CCS

[1420] Hao Ji, Lei Xie, Chuyu Wang, Yafeng Yin, and Sanglu Lu. Crowd-Sensing: a crowd-sourcing based indoor navigation using RFID-based delay tolerant network. Journal of Network and Computer Applications, 52(??):79–89, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000508.

Tao:2015:SPB

[1421] Ming Tao, Shuqiang Huang, Yang Li, Min Yan, and Yuyu Zhou. SA-PSO based optimizing reader deployment in large-scale RFID Systems. Journal of Network and Computer Applications, 52(??):90– 100, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500051X.

Ghaffari:2015:CCM

[1422] Ali Ghaffari. Congestion control mechanisms in wireless sensor networks: a survey. Journal of Network and Computer Applications, 52(??):

101–115, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000557.

Elhabyan:2015:TTP

[1423] Riham S. Y. Elhabyan and Mustapha C. E. Yagoub. Two-tier particle swarm optimization protocol for clustering and routing in wireless sensor network. Journal of Network and Computer Applications, 52(??): 116–128, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000429.

Do:2015:MOS

[1424] Tien Van Do, Nam H. Do, Ádám Horváth, and Jinting Wang. Modelling opportunistic spectrum renting in mobile cellular networks. Journal of Network and Computer Applications, 52(??):129–138, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515000454.

Jiang:2015:ASM

[1425] Wei Jiang, Yue Ma, Xia Zhang, Xupeng Wang, and Zili Shao. Adaptive security management of real-time storage applications over NAND based storage systems. Journal of Network and Computer Applications, 52(??):139–153, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000491.

Ahmed:2015:SAE

[1426] Ejaz Ahmed, Abdullah Gani, Muhammad Khurram Khan, Rajkumar Buyya, and Samee U. Khan. Seamless application execution in mobile cloud computing: Motivation, taxonomy, and open challenges. Journal of Network and Computer Applications, 52(??):154– 172, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000545.

Khan:2015:NDT

[1427] Athar Ali Khan, Mubashir Husain Rehmani, and Yasir Saleem. Neighbor discovery in traditional wireless networks and cognitive radio networks: Basics, taxonomy, challenges and future research directions. Journal of Network and Computer Applications, 52(??):173– 190, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000569.

Anonymous:2015:EBf

[1428] Anonymous. Editorial board. Journal of Network and Computer Applications, 52(??):ifc, June 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000685.

Ebrahimzadeh:2015:QAG

[1429] Amin Ebrahimzadeh, Akbar Ghaffarpour Rahbar, and Behrooz Alizadeh. QoS aware green routing and wavelength assignment in core WDM networks. Journal of Network and Computer Applications, 53 (??):1–13, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000570.

Arslan:2015:TNA

[1430] Engin Arslan, Murat Yuksel, and Mehmet Hadi Gunes. Training network administrators in a game-like environment. Journal of Network and Computer Applications, 53(??):14–23, July 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000600.

Figueiredo:2015:DML

[1431] Sérgio Figueiredo, Seil Jeon, Diogo Gomes, and Rui L. Aguiar. D3M: Multicast listener mobility support mechanisms over distributed mobility anchoring architectures. Journal of Network and Computer Applications, 53(??):24–38, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000442.

Tanwar:2015:SRH

[1432] Sudeep Tanwar, Neeraj Kumar, and Joel J. P. C. Rodrigues. A systematic review on heterogeneous routing protocols for wireless sensor network. Journal of Network and Computer Applications, 53(??):39– 56, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000594.

Wang:2015:FPPa

[1433] Yong Wang, Yun Xia, Jie Hou, Shi meng Gao, Xiao Nie, and Qi Wang. A fast privacy-preserving framework for continuous location-based queries in road networks. Journal of Network and Computer Applications, 53 (??):57–73, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500048X.

Arunkumar:2015:RPP

[1434] S. Arunkumar, M. Srivatsa, and M. Rajarajan. A review paper on preserving privacy in mobile environments. Journal of Network and Computer Applications, 53(??):74–90, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515000582.

Boushaba:2015:MPR

[1435] Abdelali Boushaba, Adil Benabbou, Rachid Benabbou, Azeddine Zahi, and Mohammed Oumsis. Multi-point relay selection strategies to reduce topology control traffic for OLSR protocol in MANETs. Journal of Network and Computer Applications, 53(??):91–102, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000636.

Wang:2015:DVD

[1436] Jingyuan Wang, Jiangtao Wen, Chao Li, Zhang Xiong, and Yuxing Han. DC-Vegas: a delay-based TCP congestion control algorithm for datacenter applications. Journal of Network and Computer Applications, 53(??):103–114, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000661.

Caputi:2015:SOP

[1437] Valentina Caputi and Antonio Garrido. Student-oriented planning of elearning contents for Moodle. Journal of Network and Computer Applications, 53(??):115–127, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451500065X.

Liu:2015:IMB

[1438] Chao-Liang Liu, Cheng-Jung Tsai, Ting-Yi Chang, Wang-Jui Tsai, and Po-Kai Zhong. Implementing multiple biometric features for a recallbased graphical keystroke dynamics authentication system on a smart phone. Journal of Network and Computer Applications, 53(??):128– 139, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000612.

Zhang:2015:AMO

[1439] Zhongbao Zhang, Sen Su, Yikai Lin, Xiang Cheng, Kai Shuang, and Peng Xu. Adaptive multi-objective artificial immune system based virtual network embedding. Journal of Network and Computer Applications, 53(??):140–155, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000624.

Lin:2015:AMM

[1440] Zhiting Lin and XiuLong Wu. Analyzing and modeling mobility for infrastructure-less communication. Journal of Network and Computer Applications, 53(??):156–163, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515000752.

Razavi:2015:DSC

[1441] Sara Modarres Razavi, Di Yuan, Fredrik Gunnarsson, and Johan Moe. On dynamic signaling congestion mitigation by overlapping tracking area lists. Journal of Network and Computer Applications, 53(??):164– 172, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000648.

Wang:2015:CAD

[1442] Xiumin Wang, Chau Yuen, Xiaoming Chen, Naveed Ul Hassan, and Yiming Ouyang. Cost-aware demand scheduling for delay tolerant applications. Journal of Network and Computer Applications, 53(??):173– 182, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000740.

Collotta:2015:FFA

[1443] Mario Collotta. FLBA: a fuzzy algorithm for load balancing in IEEE 802.11 networks. Journal of Network and Computer Applications, 53(??): 183–192, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000776.

Anonymous:2015:EBg

[1444] Anonymous. Editorial board. Journal of Network and Computer Applications, 53(??):ifc, July 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000971.

Abdeljaouad:2015:MIQ

[1445] Imad Abdeljaouad and Ahmed Karmouch. Monitoring IPTV quality of experience in overlay networks using utility functions. Journal of Network and Computer Applications, 54(??):1–10, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000788.

Liu:2015:OSM

[1446] Yujie Liu, Yong Li, Yue Wang, and Jian Yuan. Optimal scheduling for multi-flow update in Software-Defined Networks. Journal of Network and Computer Applications, 54(??):11–19, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451500082X.

Homayounnejad:2015:EDM

[1447] Saman Homayounnejad and Alireza Bagheri. An efficient distributed max-flow algorithm for Wireless Sensor Networks. Journal of Network and Computer Applications, 54(??):20–32, August 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000764.

AlSkaif:2015:GTE

[1448] Tarek AlSkaif, Manel Guerrero Zapata, and Boris Bellalta. Game theory for energy efficiency in Wireless Sensor Networks: Latest trends. Journal of Network and Computer Applications, 54(??):33– 61, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000806.

Sanchez-Casado:2015:ICZ

[1449] Leovigildo Sánchez-Casado, Gabriel Maciá-Fernández, Pedro García-Teodoro, and Nils Aschenbruck. Identification of contamination zones for sinkhole detection in MANETs. Journal of Network and Computer Applications, 54(??):62–77, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515000818.

Najam:2015:SPP

[1450] Maleeha Najam, Usman Younis, and Raihan ur Rasool. Speculative parallel pattern matching using stride-k DFA for deep packet inspection. Journal of Network and Computer Applications, 54(??):78– 87, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000867.

Samian:2015:CSM

[1451] Normalia Samian, Zuriati Ahmad Zukarnain, Winston K. G. Seah, Azizol Abdullah, and Zurina Mohd Hanapi. Cooperation stimulation mechanisms for wireless multihop networks: a survey. Journal of Network and Computer Applications, 54(??):88–106, August 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000855.

Anonymous:2015:EBh

[1452] Anonymous. Editorial board. Journal of Network and Computer Applications, 54(??):ifc, August 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001125.

Malatras:2015:SAS

[1453] Apostolos Malatras. State-of-the-art survey on P2P overlay networks in pervasive computing environments. Journal of Network and Computer Applications, 55(??):1–23, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515000879.

Malik:2015:QIB

[1454] Aqsa Malik, Junaid Qadir, Basharat Ahmad, Kok-Lim Alvin Yau, and Ubaid Ullah. QoS in IEEE 802.11-based wireless networks: a contemporary review. Journal of Network and Computer Applications, 55(??):24–46, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000892.

Lin:2015:INV

[1455] Yaguang Lin, Xiaoming Wang, Lichen Zhang, Peng Li, Dan Zhang, and Sen Liu. The impact of node velocity diversity on mobile opportunistic network performance. Journal of Network and Computer Applications, 55(??):47–58, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000831.

Wang:2015:TBP

[1456] Yingjie Wang, Guisheng Yin, Zhipeng Cai, Yuxin Dong, and Hongbin Dong. A trust-based probabilistic recommendation model for social networks. Journal of Network and Computer Applications, 55(??):59–67, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500079X.

Ramirez:2015:RNA

[1457] Diego F. Ramírez and Sandra Céspedes. Routing in Neighborhood Area Networks: a survey in the context of AMI communications. Journal of Network and Computer Applications, 55(??):68–80, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000843.

Liao:2015:DPC

[1458] S. Liao, J. Sun, Y. Chen, Y. Wang, and P. Zhang. Distributed power control for wireless networks via the alternating direction method of multipliers. Journal of Network and Computer Applications, 55(??):81–88, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001058.

Shameli-Sendi:2015:OOR

[1459] Alireza Shameli-Sendi and Michel Dagenais. ORCEF: Online response cost evaluation framework for intrusion response system. Journal of Network and Computer Applications, 55(??):89–107, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515000958.

Rehman:2015:USC

[1460] Zia ur Rehman, Omar Khadeer Hussain, and Farookh Khadeer Hussain. User-side cloud service management: State-of-the-art and future directions. Journal of Network and Computer Applications, 55 (??):108–122, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001071.

Zin:2015:SSM

[1461] Shazana Md Zin, Nor Badrul Anuar, Miss Laiha Mat Kiah, and Ismail Ahmedy. Survey of secure multipath routing protocols for WSNs. Journal of Network and Computer Applications, 55(??):123–153, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000946.

Wang:2015:FPPb

[1462] Kai Wang and Jun Li. FREME: a pattern partition based engine for fast and scalable regular expression matching in practice. Journal of Network and Computer Applications, 55(??):154–169, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001198.

Yang:2015:NMV

[1463] Dingqi Yang, Daqing Zhang, Longbiao Chen, and Bingqing Qu. NationTelescope: Monitoring and visualizing large-scale collective behavior in LBSNs. Journal of Network and Computer Applications, 55 (??):170–180, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001101.

Alrshah:2015:ASL

[1464] Mohamed A. Alrshah, Mohamed Othman, Borhanuddin Ali, and Zurina Mohd Hanapi. Agile-SD: a Linux-based TCP congestion control algorithm for supporting high-speed and short-distance networks. Journal of Network and Computer Applications, 55(??):181–190, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515001186.

Cetinski:2015:AWA

[1465] Katja Cetinski and Matjaz B. Juric. AME-WPC: Advanced model for efficient workload prediction in the cloud. Journal of Network and Computer Applications, 55(??):191–201, September 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001241.

Misra:2015:EAB

[1466] Sudip Misra, Samaresh Bera, Tamoghna Ojha, and Liang Zhou. EN-TICE: Agent-based energy trading with incomplete information in the smart grid. Journal of Network and Computer Applications, 55(??):202– 212, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001083.

Kasapakis:2015:PGS

[1467] Vlasios Kasapakis and Damianos Gavalas. Pervasive gaming: Status, trends and design principles. Journal of Network and Computer Applications, 55(??):213–236, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515001095.

Anonymous:2015:EBi

[1468] Anonymous. Editorial board. Journal of Network and Computer Applications, 55(??):ifc, September 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001472.

Karami:2015:AAC

[1469] Amin Karami. ACCPndn: Adaptive Congestion Control Protocol in Named Data Networking by learning capacities using optimized Time-Lagged Feedforward Neural Network. Journal of Network and Computer Applications, 56(??):1–18, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515001265.

Ning:2015:NAS

[1470] Zhaolong Ning, Qingyang Song, Lei Guo, and Xiangjie Kong. A novel adaptive spectrum allocation scheme for multi-channel multi-radio wireless mesh networks. Journal of Network and Computer Applications, 56(??):19–27, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001289.

Khan:2015:SCO

[1471] Minhaj Ahmad Khan. A survey of computation offloading strategies for performance improvement of applications running on mobile devices. Journal of Network and Computer Applications, 56(??):28– 40, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001320.

Ricciardi:2015:EEE

[1472] Sergio Ricciardi, Francesco Palmieri, Aniello Castiglione, and Davide Careglio. Energy efficiency of elastic frequency grids in multilayer IP/ MPLS-over-flexgrid networks. Journal of Network and Computer Applications, 56(??):41–47, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001423.

Abdullahi:2015:SCA

[1473] Ibrahim Abdullahi, Suki Arif, and Suhaidi Hassan. Survey on caching approaches in Information Centric Networking. Journal of Network and Computer Applications, 56(??):48–59, October 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001381.

Hamrouni:2015:IDQ

[1474] T. Hamrouni, C. Hamdeni, and F. Ben Charrada. Impact of the distribution quality of file replicas on replication strategies. Journal of Network and Computer Applications, 56(??):60–76, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451500137X.

Khan:2015:NTA

[1475] M. S. S. Khan, Anup Kumar, Bin Xie, and Prasanna K. Sahoo. Network tomography application in mobile ad-hoc network using stitching algorithm. Journal of Network and Computer Applications, 56(??):77– 87, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001587.

Haralabopoulos:2015:LPI

[1476] Giannis Haralabopoulos, Ioannis Anagnostopoulos, and Sherali Zeadally. Lifespan and propagation of information in On-line Social Networks: a case study based on Reddit. Journal of Network and Computer Applications, 56(??):88–100, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001307.

Kerttula:2015:NFS

[1477] Esa Kerttula. A novel federated strong mobile signature service — the Finnish case. Journal of Network and Computer Applications, 56(??):101–114, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001319.

Calzarossa:2015:MPT

[1478] Maria Carla Calzarossa and Daniele Tessera. Modeling and predicting temporal patterns of web content changes. Journal of Network and Computer Applications, 56(??):115–123, October 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001332.

Beck:2015:DSE

[1479] Michael Till Beck, Andreas Fischer, Juan Felipe Botero, Claudia Linnhoff-Popien, and Hermann de Meer. Distributed and scalable embedding of virtual networks. Journal of Network and Computer Applications, 56(??):124–136, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001393.

Ninggal:2015:UAS

[1480] Mohd Izuan Hafez Ninggal and Jemal H. Abawajy. Utility-aware social network graph anonymization. Journal of Network and Computer Applications, 56(??):137–148, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515001204.

Wang:2015:PSC

[1481] Sheng-Shih Wang, Chun-Chih Li, Hsieh-Wei Lin, and Kuei-Ping Shih. A passive self-configuration MAC protocol for supporting network management in IEEE 802.11-based multi-hop mobile ad hoc networks. Journal of Network and Computer Applications, 56(??):149–157, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001368.

Chen:2015:SSO

[1482] Chi Chen, Xiaojie Zhu, Peisong Shen, Jing Yu, Hong Zou, and Jiankun Hu. SOLS: a scheme for outsourced location based service. Journal of Network and Computer Applications, 56(??):158–165, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001277.

RahimiZadeh:2015:PMA

[1483] Keyvan RahimiZadeh, Morteza AnaLoui, Peyman Kabiri, and Bahman Javadi. Performance modeling and analysis of virtualized multitier applications under dynamic workloads. Journal of Network and Computer Applications, 56(??):166–187, October 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001642.

Chandakanna:2015:SWB

[1484] Veerabhadra R. Chandakanna and Valli K. Vatsavayi. A sliding window based Self-Learning and Adaptive Load Balancer. Journal of Network and Computer Applications, 56(??):188–205, October 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001538.

Anonymous:2015:EBj

[1485] Anonymous. Editorial board. Journal of Network and Computer Applications, 56(??):ifc, October 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001800.

Bradai:2015:DAP

[1486] Abbas Bradai, Abderrahim Benslimane, and Kamal D. Singh. Dynamic anchor points selection for mobility management in Software Defined Networks. Journal of Network and Computer Applications, 57(??):1–11, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001575.

Zhao:2015:PBO

[1487] Erdun Zhao and Liansheng Tan. A PCA based optimization approach for IP traffic matrix estimation. Journal of Network and Computer Applications, 57(??):12–20, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001691.

Kim:2015:SEF

[1488] Younghoon Kim, Ikjun Yeom, Jun Bi, and Yusung Kim. Scalable and efficient file sharing in information-centric networking. Journal of Network and Computer Applications, 57(??):21–32, November 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001733.

Gur:2015:EAC

[1489] Gürkan Gür. Energy-aware cache management at the wireless network edge for information-centric operation. Journal of Network and Computer Applications, 57(??):33–42, November 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001344.

Xu:2015:NTA

[1490] Qin Xu, Fan Li, Jinsheng Sun, and Moshe Zukerman. A new TCP/ AQM system analysis. Journal of Network and Computer Applications, 57(??):43–60, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001599.

Yang:2015:DAT

[1491] Shin-Jer Yang and Yi-Ru Chen. Design adaptive task allocation scheduler to improve MapReduce performance in heterogeneous clouds. Journal of Network and Computer Applications, 57(??):61–70, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515001757.

Deka:2015:NDA

[1492] Rup Kumar Deka, Kausthav Pratim Kalita, D. K. Bhattacharya, and Jugal K. Kalita. Network defense: Approach’s, methods and techniques. Journal of Network and Computer Applications, 57(??):71–84, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001745.

Chakraborty:2015:FRS

[1493] Suchetana Chakraborty, Sandip Chakraborty, Sukumar Nandi, and Sushanta Karmakar. Fault resilience in sensor networks: Distributed node-disjoint multi-path multi-sink forwarding. Journal of Network and Computer Applications, 57(??):85–101, November 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001770.

Ribas:2015:PNB

[1494] Maristella Ribas, C. G. Furtado, José Neuman de Souza, Giovanni Cordeiro Barroso, Antão Moura, Alberto S. Lima, and Flávio R. C. Sousa. A Petri net-based decision-making framework for assessing cloud services adoption: the use of spot instances for cost reduction. Journal of Network and Computer Applications, 57(??):102–118, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001563.

Li:2015:MFP

[1495] Wei Li, Jorge Portilla, Félix Moreno, Guixuan Liang, and Teresa Riesgo. Multiple feature points representation in target localization of wireless visual sensor networks. Journal of Network and Computer Applications, 57(??):119–128, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001769.

Chen:2015:CPB

[1496] Jiaoyan Chen, Bang Wang, and Wenyu Liu. Constructing perimeter barrier coverage with bistatic radar sensors. Journal of Network and Computer Applications, 57(??):129–141, November 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001782.

Mohajerzadeh:2015:EDC

[1497] Amir Hossein Mohajerzadeh, Mohammad Hossein Yaghmaee, and Afsane Zahmatkesh. Efficient data collecting and target parameter estimation in wireless sensor networks. Journal of Network and Computer Applications, 57(??):142–155, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001290.

Suh:2015:SIW

[1498] Yuhwa Suh, Kiyoung Kim, Aran Kim, and Yongtae Shin. A study on impact of wired access networks for green Internet. Journal of Network and Computer Applications, 57(??):156–168, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001940.

Khanjary:2015:CDC

[1499] Mohammad Khanjary, Masoud Sabaei, and Mohammad Reza Meybodi. Critical density for coverage and connectivity in two-dimensional fixed-orientation directional sensor networks using continuum percolation. Journal of Network and Computer Applications, 57(??):169–181, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001964.

Jiang:2015:CMH

[1500] Dingde Jiang, Zhengzheng Xu, Wenqin Wang, Yuanting Wang, and Yang Han. A collaborative multi-hop routing algorithm for maximum achievable rate. Journal of Network and Computer Applications, 57 (??):182–191, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001356.

Salameh:2015:SSB

[1501] Haythem Bany Salameh. Spread spectrum-based coordination design for spectrum-agile wireless ad hoc networks. Journal of Network and Computer Applications, 57(??):192–201, November 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002027.

Anonymous:2015:EBk

[1502] Anonymous. Editorial board. Journal of Network and Computer Applications, 57(??):ifc, November 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002180.

Guo:2015:WSD

[1503] Zhe Guo and Zheng Yan. A Weighted Semi-Distributed Routing Algorithm for LEO satellite networks. Journal of Network and Computer Applications, 58(??):1–11, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515002015.

Carrabs:2015:HEA

[1504] Francesco Carrabs, Raffaele Cerulli, Ciriaco D’Ambrosio, and Andrea Raiconi. A hybrid exact approach for maximizing lifetime in sensor networks with complete and partial coverage constraints. Journal of Network and Computer Applications, 58(??):12–22, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002040.

Rehman:2015:FLQ

[1505] Osama M. Hussain Rehman, Hadj Bourdoucen, and Mohamed Ould-Khaoua. Forward link quality estimation in VANETs for sender-oriented alert messages broadcast. Journal of Network and Computer Applications, 58(??):23–41, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002064.

Ahmad:2015:RMA

[1506] Raja Wasim Ahmad, Abdullah Gani, Siti Hafizah Ab. Hamid, Feng Xia, and Muhammad Shiraz. A review on mobile application energy profiling: Taxonomy, state-of-the-art, and open research issues. Journal of Network and Computer Applications, 58(??):42–59, December 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002088.

Coll-Perales:2015:EMC

[1507] B. Coll-Perales, J. Gozalvez, and M. Sepulcre. Empirical models of the communications performance of Multi-hop Cellular Networks using D2D. Journal of Network and Computer Applications, 58(??):60– 72, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002039.

Mohd:2015:SLB

[1508] Bassam J. Mohd, Thaier Hayajneh, and Athanasios V. Vasilakos. A survey on lightweight block ciphers for low-resource devices: Comparative study and open issues. Journal of Network and Computer Applications, 58(??):73–93, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002076.

Niu:2015:CLF

[1509] Jianwei Niu, Danning Wang, and Mohammed Atiquzzaman. Copy limited flooding over opportunistic networks. Journal of Network and Computer Applications, 58(??):94–107, December 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002052.

Haddi:2015:SIM

[1510] Fatima Lamia Haddi and Mahfoud Benchaïba. A survey of incentive mechanisms in static and mobile P2P systems. Journal of Network and Computer Applications, 58(??):108–118, December 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002106.

Najm:2015:ISC

[1511] Ihab Ahmed Najm, Mahamod Ismail, Jaime Lloret, Kayhan Zrar Ghafoor, B. B. Zaidan, and Abd Al razak Tareq Rahem. Improvement of SCTP congestion control in the LTE-A network. Journal of Network and Computer Applications, 58(??):119–129, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451500209X.

Hajjar:2015:NTA

[1512] Amjad Hajjar, Jawad Khalife, and Jesús Díaz-Verdejo. Network traffic application identification based on message size analysis. Journal of Network and Computer Applications, 58(??):130–143, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515002167.

Khattak:2015:BCD

[1513] Sheharbano Khattak, Zaafar Ahmed, Affan A. Syed, and Syed Ali Khayam. BotFlex: a community-driven tool for botnet detection. Journal of Network and Computer Applications, 58(??):144–154, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002155.

Wang:2015:DME

[1514] Guijuan Wang, Jiguo Yu, Dongxiao Yu, Haitao Yu, Li Feng, and Pan liu. DS–MAC: an energy efficient demand sleep MAC protocol with low latency for wireless sensor networks. Journal of Network and Computer Applications, 58(??):155–164, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515002131.

Shameli-Sendi:2015:TDD

[1515] Alireza Shameli-Sendi, Makan Pourzandi, Mohamed Fekih-Ahmed, and Mohamed Cheriet. Taxonomy of Distributed Denial of Service mitigation approaches for cloud computing. Journal of Network and Computer Applications, 58(??):165–179, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515002118.

Leu:2015:PIL

[1516] Fang-Yie Leu, Chin-Yu Liu, Jung-Chun Liu, Fuu-Cheng Jiang, and Heru Susanto. S-PMIPv6: an intra-LMA model for IPv6 mobility. Journal of Network and Computer Applications, 58(??):180–191, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002003.

Lloret:2015:GNP

[1517] Jaime Lloret, Kayhan Zrar Ghafoor, Danda B. Rawat, and Youssef Nasser. Green network protocols and algorithms. Journal of Network and Computer Applications, 58(??):192–193, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002441.

Ding:2015:EED

[1518] Ruiqi Ding, Cristina Hava Muntean, and Gabriel-Miro Muntean. Energyefficient device-differentiated cooperative adaptive multimedia delivery solution in wireless networks. Journal of Network and Computer Applications, 58(??):194–207, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451500212X.

Can:2015:SBD

[1519] Zuhal Can and Murat Demirbas. Smartphone-based data collection from wireless sensor networks in an urban environment. Journal of Network and Computer Applications, 58(??):208–216, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451500199X.

Rodriguez-Perez:2015:ACR

[1520] Miguel Rodríguez-Pérez, Sergio Herrería-Alonso, Manuel Fernández-Veiga, and Cándido López-García. An ant colonization routing algorithm to minimize network power consumption. Journal of Network and Computer Applications, 58(??):217–226, December 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001976.

Sami:2015:EAC

[1521] Mahmoud Sami, Nor Kamariah Noordin, Fazirulhysiam Hashim, Shamala Subramaniam, and Ayyoub Akbari-Moghanjoughi. An energyaware cross-layer cooperative MAC protocol for wireless ad hoc networks. Journal of Network and Computer Applications, 58(??):227–240, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001952.

Li:2015:EEQ

[1522] Chengming Li, Wenjing Liu, Lei Wang, Mingchu Li, and Koji Okamura. Energy-efficient quality of service aware forwarding scheme for Content-Centric Networking. Journal of Network and Computer Applications, 58(??):241–254, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001939.

PourEmami:2015:TBP

[1523] Seyed Iman PourEmami and Bahador Bakhshi. On the trade-off between power-efficiency and blocking probability in fault-tolerant WDM networks. Journal of Network and Computer Applications, 58(??):255– 266, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001927.

Daghistani:2015:GPP

[1524] Anas Daghistani, Abderrahman Ben Khalifa, Ahmad Showail, and Basem Shihada. Green partial packet recovery in wireless sensor networks. Journal of Network and Computer Applications, 58(??):267–279, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001915.

Ke:2015:FBM

[1525] Chih-Heng Ke, Chong-Yi Yang, Jiann-Liang Chen, Kayhan Zrar Ghafoor, and Jaime Lloret. Frame-based mapping mechanism for energyefficient MPEG-4 video transmission over IEEE 802.11e networks with better quality of delivery. Journal of Network and Computer Applications, 58(??):280–286, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515001903.

Bakhouya:2015:EEA

[1526] M. Bakhouya, J. Gaber, and P. Lorenz. Energy evaluation of AID protocol in Mobile Ad Hoc Networks. Journal of Network and Computer Applications, 58(??):287–293, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515001897.

Jabbarpour:2015:GVT

[1527] Mohammad Reza Jabbarpour, Rafidah Md Noor, and Rashid Hafeez Khokhar. Green vehicle traffic routing system using ant-based algorithm. Journal of Network and Computer Applications, 58(??):294–308, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001885.

Faheem:2015:EEE

[1528] Muhammad Faheem, Muhammad Zahid Abbas, Gurkan Tuna, and Vehbi Cagri Gungor. EDHRP: Energy efficient event driven hybrid routing protocol for densely deployed wireless sensor networks. Journal of Network and Computer Applications, 58(??):309–326, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001873.

Jia:2015:PSP

[1529] Gangyong Jia, Guangjie Han, Jinfang Jiang, and Joel J. P. C. Rodrigues. PARS: a scheduling of periodically active rank to optimize power efficiency for main memory. Journal of Network and Computer Applications, 58(??):327–336, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001861.

Anonymous:2015:EBl

[1530] Anonymous. Editorial board. Journal of Network and Computer Applications, 58(??):ifc, December 2015. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002398.

Zhou:2016:SIG

[1531] Zhang Bing Zhou, Wendong Xiao, Jucheng Yang, and Faming Gong. Special issue on green networking, computing, and software systems. Journal of Network and Computer Applications, 59(??):1–3, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002647.

Jiang:2016:GMR

[1532] Jinfang Jiang, Guangjie Han, Hui Guo, Lei Shu, and Joel J. P. C. Rodrigues. Geographic multipath routing based on geospatial division in duty-cycled underwater wireless sensor networks. Journal of Network and Computer Applications, 59(??):4–13, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000521.

Zhao:2016:NEA

[1533] Qing Zhao, Congcong Xiong, Ce Yu, Chuanlei Zhang, and Xi Zhao. A new energy-aware task scheduling method for data-intensive applications in the cloud. Journal of Network and Computer Applications, 59(??):14–27, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000909.

Xue:2016:MSC

[1534] Xiao Xue, Shufang Wang, and Baoyun Lu. Manufacturing service composition method based on networked collaboration mode. Journal of Network and Computer Applications, 59(??):28–38, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000934.

Wang:2016:EES

[1535] Yanwen Wang, Hainan Chen, Xiaoling Wu, and Lei Shu. An energyefficient SDN based sleep scheduling algorithm for WSNs. Journal of Network and Computer Applications, 59(??):39–45, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515000910.

Gai:2016:DEA

[1536] Keke Gai, Meikang Qiu, Hui Zhao, Lixin Tao, and Ziliang Zong. Dynamic energy-aware cloudlet-based mobile cloud computing model for green computing. Journal of Network and Computer Applications, 59(??):46– 54, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500123X.

Tong:2016:NGA

[1537] Chao Tong, Yu Lian, Jianwei Niu, Zhongyu Xie, and Yang Zhang. A novel green algorithm for sampling complex networks. Journal of Network and Computer Applications, 59(??):55–62, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001411.

Xiong:2016:EOB

[1538] Yonghua Xiong, Shaoyun Wan, Jinhua She, Min Wu, Yong He, and Keyuan Jiang. An energy-optimization-based method of task scheduling for a cloud video surveillance center. Journal of Network and Computer Applications, 59(??):63–73, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515001459.

Riekstin:2016:OEE

[1539] Ana C. Riekstin, Guilherme C. Januário, Bruno B. Rodrigues, Viviane T. Nascimento, Tereza C. M. B. Carvalho, and Catalin Meirosu. Orchestration of energy efficiency capabilities in networks. Journal of Network and Computer Applications, 59(??):74–87, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001435.

Fang:2016:BBB

[1540] Weidong Fang, Chuanlei Zhang, Zhidong Shi, Qing Zhao, and Lianhai Shan. BTRES: Beta-based Trust and Reputation Evaluation System for wireless sensor networks. Journal of Network and Computer Applications, 59(??):88–94, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451500140X.

Marjanovic:2016:EAQ

[1541] Martina Marjanović, Lea Skorin-Kapov, Kresimir Pripuzić, Aleksandar Antonić, and Ivana Podnar Zarko. Energy-aware and quality-driven sensor management for green mobile crowd sensing. Journal of Network and Computer Applications, 59(??):95–108, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001678.

Long:2016:TAR

[1542] Teng Long and Wenhui Zhang. Termination analysis with recursive calling graphs. Journal of Network and Computer Applications, 59(??):109– 116, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001605.

Xiao:2016:REM

[1543] Chen Xiao, Lifeng Wang, Mengjiao Zhu, and Wendong Wang. A resource-efficient multimedia encryption scheme for embedded video sensing system based on unmanned aircraft. Journal of Network and Computer Applications, 59(??):117–125, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001654.

Zhang:2016:EEA

[1544] Liping Zhang, Shanyu Tang, and Shaohui Zhu. An energy efficient authenticated key agreement protocol for SIP-based green VoIP networks. Journal of Network and Computer Applications, 59(??):126– 133, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001666.

Cholda:2016:OSB

[1545] Piotr Cholda and Piotr Jaglarz. Optimization/simulation-based risk mitigation in resilient green communication networks. Journal of Network and Computer Applications, 59(??):134–157, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001721.

Li:2016:ICV

[1546] Mingfu Li, Jingping Bi, and Zhongcheng Li. Improving consolidation of virtual machine based on virtual switching overhead estimation. Journal of Network and Computer Applications, 59(??):158–167, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500171X.

Yongsiriwit:2016:SFC

[1547] Karn Yongsiriwit, Nour Assy, and Walid Gaaloul. A semantic framework for configurable business process as a service in the cloud. Journal of Network and Computer Applications, 59(??):168–184, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515001708.

Paul:2016:EEA

[1548] Debdeep Paul, Wen-De Zhong, and Sanjay K. Bose. Energy efficiency aware load distribution and electricity cost volatility control for cloud service providers. Journal of Network and Computer Applications, 59(??):185–197, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515001988.

Liu:2016:NOP

[1549] Zheli Liu, Xiaofeng Chen, Jun Yang, Chunfu Jia, and Ilsun You. New order preserving encryption model for outsourced databases in cloud environments. Journal of Network and Computer Applications, 59(??):198–207, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001350.

Celesti:2016:ALT

[1550] Antonio Celesti, Maria Fazio, Massimo Villari, and Antonio Puliafito. Adding long-term availability, obfuscation, and encryption to multicloud storage systems. Journal of Network and Computer Applications, 59(??):208–218, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514002288.

Sun:2016:SRC

[1551] Yunchuan Sun, Cheng Lu, Rongfang Bie, and Junsheng Zhang. Semantic relation computing theory and its application. Journal of Network and Computer Applications, 59(??):219–229, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002240.

Castro:2016:EEE

[1552] Miguel Castro, Antonio J. Jara, and Antonio F. Skarmeta. Enabling end-to-end CoAP-based communications for the Web of Things. Journal of Network and Computer Applications, 59(??):230–236, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514002264.

Lopez-de-Armentia:2016:MSN

[1553] Juan López de Armentia, Diego Casado-Mansilla, and Diego López de Ipiña. Making social networks a means to save energy. Journal of Network and Computer Applications, 59(??):237–246, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804514002276.

Yu:2016:SGC

[1554] Ming Yu and Nirwan Ansari. Smart grid communications: Modeling and validation. Journal of Network and Computer Applications, 59(??):247–249, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002623.

Sbordone:2016:RPC

[1555] D. A. Sbordone, L. Martirano, M. C. Falvo, L. Chiavaroli, B. Di Pietra, I. Bertini, and A. Genovese. Reactive power control for an energy storage system: a real implementation in a Micro-Grid. Journal of Network and Computer Applications, 59(??):250–263, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451500106X.

Cai:2016:NMR

[1556] Ziyuan Cai, Ming Yu, Michael Steurer, Hui Li, and Yizhou Dong. A network model for the real-time communications of a smart grid prototype. Journal of Network and Computer Applications, 59(??):264– 273, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001630.

Bhor:2016:NPG

[1557] Dhananjay Bhor, Kavinkadhirselvan Angappan, and Krishna M. Sivalingam. Network and power-grid co-simulation framework for Smart Grid wide-area monitoring networks. Journal of Network and Computer Applications, 59(??):274–284, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515001447.

Wolkerstorfer:2016:MSF

[1558] Martin Wolkerstorfer, Bernhard Schweighofer, Hannes Wegleiter, Driton Statovci, Herbert Schwaiger, and Walter Lackner. Measurement and simulation framework for throughput evaluation of narrowband power line communication links in low-voltage grids. Journal of Network and Computer Applications, 59(??):285–300, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001617.

Alcaraz:2016:PES

[1559] Cristina Alcaraz, Javier Lopez, and Stephen Wolthusen. Policy enforcement system for secure interoperable control in distributed Smart Grid systems. Journal of Network and Computer Applications, 59(??):301– 314, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515001629.

Zhang:2016:RDS

[1560] Yuan Zhang, Navid Rahbari-Asr, and Mo-Yuen Chow. A robust distributed system incremental cost estimation algorithm for smart grid economic dispatch with communications information losses. Journal of Network and Computer Applications, 59(??):315–324, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515001216.

Yi:2016:PAD

[1561] Ping Yi, Ting Zhu, Qingquan Zhang, Yue Wu, and Li Pan. Puppet attack: a denial of service attack in advanced metering infrastructure network. Journal of Network and Computer Applications, 59(??):325–332, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000880.

Bae:2016:PPE

[1562] Mungyu Bae, Kangho Kim, and Hwangnam Kim. Preserving privacy and efficiency in data communication and aggregation for AMI network. Journal of Network and Computer Applications, 59(??):333– 344, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451500168X.

Amoah:2016:FMA

[1563] Raphael Amoah, Seyit Camtepe, and Ernest Foo. Formal modelling and analysis of DNP3 secure authentication. Journal of Network and Computer Applications, 59(??):345–360, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001228.

Abawajy:2016:NCA

[1564] Jemal H. Abawajy, Albert Y. Zomaya, and Ivan Stojmenovic. Network computing and applications for Big Data analytics. Journal of Network and Computer Applications, 59(??):361, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002635.

Xu:2016:PIR

[1565] Wei Xu, Jianshan Sun, Jian Ma, and Wei Du. A personalized information recommendation system for R&D project opportunity finding in big data contexts. Journal of Network and Computer Applications, 59(??):362–369, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515000466.

Amini:2016:MSM

[1566] Amineh Amini, Hadi Saboohi, Tutut Herawan, and Teh Ying Wah. MuDi-Stream: a multi density clustering algorithm for evolving data stream. Journal of Network and Computer Applications, 59(??):370– 385, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002665.

Villaca:2016:HRS

[1567] R. S. Villaça, R. Pasquini, L. B. de Paula, and M. F. Magalhães. HCube: Routing and similarity search in data centers. Journal of Network and Computer Applications, 59(??):386–398, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514002112.

Mershad:2016:MMA

[1568] Khaleel Mershad, Hassan Artail, Mazen Saghir, Hazem Hajj, and Mariette Awad. A mathematical model to analyze the utilization of a cloud datacenter middleware. Journal of Network and Computer Applications, 59(??):399–415, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804514001921.

Bhattacharya:2016:EOB

[1569] Maumita Bhattacharya, Rafiqul Islam, and Jemal Abawajy. Evolutionary optimization: a big data perspective. Journal of Network and Computer Applications, 59(??):416–426, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804514001805.

Anonymous:2016:EBa

[1570] Anonymous. Editorial board. Journal of Network and Computer Applications, 59(??):ifc, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002544.

Carminati:2016:TEP

[1571] Barbara Carminati, Elena Ferrari, and Ngoc Hong Tran. Trustworthy and effective person-to-person payments over multi-hop MANETs. Journal of Network and Computer Applications, 60(??):1–18, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804515002830.

Ahmed:2016:SNA

[1572] Mohiuddin Ahmed, Abdun Naser Mahmood, and Jiankun Hu. A survey of network anomaly detection techniques. Journal of Network and Computer Applications, 60(??):19–31, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804515002891.

Garriga:2016:RSC

[1573] Martin Garriga, Cristian Mateos, Andres Flores, Alejandra Cechich, and Alejandro Zunino. RESTful service composition at a glance: a survey. Journal of Network and Computer Applications, 60(??):32– 53, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002933.

Nabi:2016:ACS

[1574] Mina Nabi, Maria Toeroe, and Ferhat Khendek. Availability in the cloud: State of the art. Journal of Network and Computer Applications, 60(??):54–67, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002878.

Jarrah:2016:CBS

[1575] Hazim Jarrah, Nurul I. Sarkar, and Jairo Gutierrez. Comparison-based system-level fault diagnosis protocols for mobile ad-hoc networks: a survey. Journal of Network and Computer Applications, 60(??):68– 81, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002957.

Halder:2016:SMA

[1576] Subir Halder and Amrita Ghosal. A survey on mobility-assisted localization techniques in wireless sensor networks. Journal of Network and Computer Applications, 60(??):82–94, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002921.

Prokhorenko:2016:WAP

[1577] Victor Prokhorenko, Kim-Kwang Raymond Choo, and Helen Ashman. Web application protection techniques: a taxonomy. Journal of Network and Computer Applications, 60(??):95–112, January 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002908.

Moura:2016:RAN

[1578] Jose Moura and David Hutchison. Review and analysis of networking challenges in cloud computing. Journal of Network and Computer Applications, 60(??):113–129, January 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451500288X.

Ghate:2016:CDC

[1579] Pritish Vijayanand Ghate and Hemanta Kumar Pati. Collaborative distributed communication in heterogeneous environments: a comprehensive survey. Journal of Network and Computer Applications, 61(??):1– 20, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002271.

Mishra:2016:CAN

[1580] Alekha Kumar Mishra and Ashok Kumar Turuk. A comparative analysis of node replica detection schemes in wireless sensor networks. Journal of Network and Computer Applications, 61(??):21–32, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002945.

Wang:2016:NIA

[1581] Xiaonan Wang, Hongbin Chen, and Deguang Le. A novel IPv6 address configuration for a 6LoWPAN-based WBAN. Journal of Network and Computer Applications, 61(??):33–45, February 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002763.

Ernest:2016:DPE

[1582] Petro P. Ernest, Olabisi E. Falowo, and H. Anthony Chan. Design and performance evaluation of distributed mobility management schemes for network mobility. Journal of Network and Computer Applications, 61(??):46–58, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451500274X.

Alizadeh:2016:AMC

[1583] Mojtaba Alizadeh, Saeid Abolfazli, Mazdak Zamani, Sabariah Baharun, and Kouichi Sakurai. Authentication in mobile cloud computing: a survey. Journal of Network and Computer Applications, 61(??):59– 80, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002258.

Cheraghlou:2016:SFT

[1584] Mehdi Nazari Cheraghlou, Ahmad Khadem-Zadeh, and Majid Haghparast. A survey of fault tolerance architecture in cloud computing. Journal of Network and Computer Applications, 61(??):81–92, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002246.

Xiang:2016:SDC

[1585] Yu Xiang, Zhaoguang Xuan, Meng Tang, Jun Zhang, and Ming Sun. 3D space detection and coverage of wireless sensor network based on spatial correlation. Journal of Network and Computer Applications, 61(??):93–101, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002507.

Lv:2016:EBC

[1586] Cuicui Lv, Qiang Wang, Wenjie Yan, and Yi Shen. Energy-balanced compressive data gathering in Wireless Sensor Networks. Journal of Network and Computer Applications, 61(??):102–114, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002362.

Cheikhrouhou:2016:SGC

[1587] Omar Cheikhrouhou. Secure group communication in wireless sensor networks: a survey. Journal of Network and Computer Applications, 61(??):115–132, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002349.

Lall:2016:NFJ

[1588] Shruti Lall, B. T. J. Maharaj, and P. A. Jansen van Vuuren. Nullfrequency jamming of a proactive routing protocol in wireless mesh networks. Journal of Network and Computer Applications, 61(??):133– 141, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002325.

Rocha:2016:FDS

[1589] Atslands R. Rocha, Flávia C. Delicato, Luci Pirmez, Danielo G. Gomes, and José Neuman de Souza. A fully-decentralized semantic mechanism for autonomous wireless sensor nodes. Journal of Network and Computer Applications, 61(??):142–160, February 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002350.

Kumar:2016:SHC

[1590] Krishan Kumar, Arun Prakash, and Rajeev Tripathi. Spectrum handoff in cognitive radio networks: a classification and comprehensive survey. Journal of Network and Computer Applications, 61(??):161–188, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002295.

Yang:2016:MRB

[1591] Shun-Neng Yang, Shu-Wei Ho, Yi-Bing Lin, and Chai-Hien Gan. A multi-RAT bandwidth aggregation mechanism with software-defined networking. Journal of Network and Computer Applications, 61(??):189– 198, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002374.

Akhunzada:2016:SDS

[1592] Adnan Akhunzada, Abdullah Gani, Nor Badrul Anuar, Ahmed Abdelaziz, Muhammad Khurram Khan, Amir Hayat, and Samee U. Khan. Secure and dependable software defined networks. Journal of Network and Computer Applications, 61(??):199–221, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002842.

Zhang:2016:AMT

[1593] Tao Zhang, Jianxing Wang, Jiawei Huang, Yi Huang, Jianer Chen, and Yi Pan. Adaptive marking threshold method for delay-sensitive TCP in data center network. Journal of Network and Computer Applications, 61(??):222–234, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002751.

Anonymous:2016:EBb

[1594] Anonymous. Editorial board. Journal of Network and Computer Applications, 61(??):ifc, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000229.

Gope:2016:EMA

[1595] Prosanta Gope and Tzonelih Hwang. An efficient mutual authentication and key agreement scheme preserving strong anonymity of the mobile user in global mobility networks. Journal of Network and Computer Applications, 62(??):1–8, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804515002969.

Ji:2016:MLI

[1596] Soo-Yeon Ji, Bong-Keun Jeong, Seonho Choi, and Dong Hyun Jeong. A multi-level intrusion detection method for abnormal network behaviors. Journal of Network and Computer Applications, 62(??):9–17, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515002970.

Shaukat:2016:CDL

[1597] Usman Shaukat, Ejaz Ahmed, Zahid Anwar, and Feng Xia. Cloudlet deployment in local wireless networks: Motivation, architectures, applications, and open challenges. Journal of Network and Computer Applications, 62(??):18–40, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002817.

Prasad:2016:EGC

[1598] Athul Prasad, Andreas Maeder, Konstantinos Samdanis, Andreas Kunz, and Genadi Velev. Enabling group communication for public safety in LTE-Advanced networks. Journal of Network and Computer Applications, 62(??):41–52, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002866.

Inayat:2016:IRS

[1599] Zakira Inayat, Abdullah Gani, Nor Badrul Anuar, Muhammad Khurram Khan, and Shahid Anwar. Intrusion response systems: Foundations, design, and challenges. Journal of Network and Computer Applications, 62(??):53–74, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002994.

Rathore:2016:STM

[1600] Heena Rathore, Venkataramana Badarla, and George K. J. Sociopsychological trust model for Wireless Sensor Networks. Journal of Network and Computer Applications, 62(??):75–87, February 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002714.

Tian:2016:UWV

[1601] Feng Tian, Kuo-Ming Chao, Zuren Feng, Keyi Xing, and Nazaraf Shah. Using a wireless visual sensor network to harmonically navigate multiple low-cost wheelchairs in an indoor environment. Journal of Network and Computer Applications, 62(??):88–99, February 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451500291X.

Yuan:2016:RRI

[1602] Peiyan Yuan, Ping Liu, and Shaojie Tang. RIM: Relative-importance based data forwarding in people-centric networks. Journal of Network and Computer Applications, 62(??):100–111, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515003008.

Xia:2016:LWT

[1603] Hui Xia, Jia Yu, Cheng liang Tian, Zhen kuan Pan, and Edwin Sha. Light-weight trust-enhanced on-demand multi-path routing in mobile ad hoc networks. Journal of Network and Computer Applications, 62(??):112–127, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804515002982.

Luo:2016:AAW

[1604] Chengwen Luo, Hande Hong, Long Cheng, Mun Choon Chan, Jianqiang Li, and Zhong Ming. Accuracy-aware wireless indoor localization: Feasibility and applications. Journal of Network and Computer Applications, 62(??):128–136, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451500301X.

Alneyadi:2016:SDL

[1605] Sultan Alneyadi, Elankayer Sithirasenan, and Vallipuram Muthukkumarasamy. A survey on data leakage prevention systems. Journal of Network and Computer Applications, 62(??):137–152, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516000102.

Omar:2016:SRC

[1606] Mawloud Omar, Hamida Boufaghes, Lydia Mammeri, Amel Taalba, and Abdelkamel Tari. Secure and reliable certificate chains recovery protocol for mobile ad hoc networks. Journal of Network and Computer Applications, 62(??):153–162, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000096.

Yuan:2016:RPR

[1607] Peiyan Yuan, Lilin Fan, Ping Liu, and Shaojie Tang. Recent progress in routing protocols of mobile opportunistic networks: a clear taxonomy, analysis and evaluation. Journal of Network and Computer Applications, 62(??):163–170, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000084.

Vaezpour:2016:NAM

[1608] Seyed Yahya Vaezpour, Rui Zhang, Kui Wu, Jianping Wang, and Gholamali C. Shoja. A new approach to mitigating security risks of phone clone co-location over mobile clouds. Journal of Network and Computer Applications, 62(??):171–184, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000072.

Anonymous:2016:EBc

[1609] Anonymous. Editorial board. Journal of Network and Computer Applications, 62(??):ifc, February 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000679.

Saxena:2016:RSM

[1610] Divya Saxena and Vaskar Raychoudhury. Radient: Scalable, memory efficient name lookup algorithm for named data networking. Journal of Network and Computer Applications, 63(??):1–13, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000060.

Curiac:2016:TWS

[1611] Daniel-Ioan Curiac. Towards wireless sensor, actuator and robot networks: Conceptual framework, challenges and perspectives. Journal of Network and Computer Applications, 63(??):14–23, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000345.

Ge:2016:SKR

[1612] Mengmeng Ge, Kim-Kwang Raymond Choo, Huai Wu, and Yong Yu. Survey on key revocation mechanisms in wireless sensor networks. Journal of Network and Computer Applications, 63(??):24– 38, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000333.

Han:2016:EBI

[1613] Meng Han, Mingyuan Yan, Zhipeng Cai, and Yingshu Li. An exploration of broader influence maximization in timeliness networks with opportunistic selection. Journal of Network and Computer Applications, 63(??):39–49, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000059.

Liao:2016:ESA

[1614] Wen-Hwa Liao, Ssu-Chi Kuai, and Chen Liu. Efficient scheduling algorithm for multiple mobile subscriber stations with unsolicited grant service in an IEEE 802.16e network. Journal of Network and Computer Applications, 63(??):50–55, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000655.

Abdulwahid:2016:SLM

[1615] Hasan Abdulwahid, Bin Dai, Benxiong Huang, and Zijing Chen. Scheduled-links multicast routing protocol in MANETs. Journal of Network and Computer Applications, 63(??):56–67, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000552.

Khan:2016:CAF

[1616] Wazir Zada Khan, Md. Shohrab Hossain, Mohammed Y. Aalsalem, N. M. Saad, and Mohammed Atiquzzaman. A cost analysis framework for claimer reporter witness based clone detection schemes in WSNs. Journal of Network and Computer Applications, 63(??):68– 85, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000357.

Ghosal:2016:LPL

[1617] Amrita Ghosal and Subir Halder. Lifespan prolonging location-wise predetermined deployment strategy for visual sensor networks. Journal of Network and Computer Applications, 63(??):86–97, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000631.

Avallone:2016:DIW

[1618] S. Avallone and G. Di Stasi. Design and implementation of WiMesh: a tool for the performance evaluation of multi-radio wireless mesh networks. Journal of Network and Computer Applications, 63(??):98– 109, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000436.

Zhu:2016:JFR

[1619] Hao Zhu, Xiangke Liao, Cees de Laat, and Paola Grosso. Joint flow routing-scheduling for energy efficient software defined data center networks: a prototype of energy-aware network management platform. Journal of Network and Computer Applications, 63(??):110– 124, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000473.

Wei:2016:PGW

[1620] Debao Wei, Libao Deng, Peng Zhang, Liyan Qiao, and Xiyuan Peng. A page-granularity wear-leveling (PGWL) strategy for NAND flash memory-based sink nodes in wireless sensor networks. Journal of Network and Computer Applications, 63(??):125–139, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000424.

Garcia:2016:EID

[1621] Álvaro López García and Enol Fernández del Castillo. Efficient image deployment in cloud environments. Journal of Network and Computer Applications, 63(??):140–149, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000382.

Abdullaziz:2016:AAI

[1622] Osamah Ibrahiem Abdullaziz, Vik Tor Goh, Huo-Chong Ling, and Kok-Sheik Wong. AIPISteg: an active IP identification based steganographic method. Journal of Network and Computer Applications, 63(??):150– 158, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451600059X.

Coutinho:2016:PMB

[1623] Emanuel F. Coutinho, Paulo A. L. Rego, Danielo G. Gomes, and José N. de Souza. Physics and microeconomics-based metrics for evaluating cloud computing elasticity. Journal of Network and Computer Applications, 63(??):159–172, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000369.

Kreutz:2016:CRA

[1624] Diego Kreutz, Oleksandr Malichevskyy, Eduardo Feitosa, Hugo Cunha, Rodrigo da Rosa Righi, and Douglas D. J. de Macedo. A cyberresilient architecture for critical security services. Journal of Network and Computer Applications, 63(??):173–189, March 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000539.

Suarez:2016:SIM

[1625] Javier Suarez, Jose Quevedo, Ivan Vidal, Daniel Corujo, Jaime Garcia-Reinoso, and Rui L. Aguiar. A secure IoT management architecture based on Information-Centric Networking. Journal of Network and Computer Applications, 63(??):190–204, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516000370.

Anonymous:2016:EBd

[1626] Anonymous. Editorial board. Journal of Network and Computer Applications, 63(??):ifc, March 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000837.

Fernandes:2016:NAD

[1627] Gilberto Fernandes, Jr., Luiz F. Carvalho, Joel J. P. C. Rodrigues, and Mario Lemes Proença, Jr. Network anomaly detection using IP flows with Principal Component Analysis and Ant Colony Optimization. Journal of Network and Computer Applications, 64(??):1–11, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000618.

Pasupuleti:2016:ESP

[1628] Syam Kumar Pasupuleti, Subramanian Ramalingam, and Rajkumar Buyya. An efficient and secure privacy-preserving approach for outsourced data of resource constrained mobile devices in cloud computing. Journal of Network and Computer Applications, 64(??):12– 22, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000540.

Zhang:2016:RPA

[1629] Jiangtao Zhang, Hejiao Huang, and Xuan Wang. Resource provision algorithms in cloud computing: a survey. Journal of Network and Computer Applications, 64(??):23–42, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516000643.

Chavarria-Reyes:2016:REC

[1630] Elias Chavarria-Reyes, Etimad Fadel, Suleiman Almasri, and M. G. Abbas Malik. Reducing the energy consumption at the user equipment through multi-stream cross-carrier-aware discontinuous reception (DRX) in LTE-advanced systems. Journal of Network and Computer Applications, 64(??):43–61, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000412.

Gregori:2016:SBC

[1631] Enrico Gregori, Alessandro Improta, Luciano Lenzini, Valerio Luconi, Nilo Redini, and Alessio Vecchio. Smartphone-based crowdsourcing for estimating the bottleneck capacity in wireless networks. Journal of Network and Computer Applications, 64(??):62–75, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000485.

Asif:2016:CBD

[1632] Waqar Asif, Hassaan Khaliq Qureshi, Muttukrishnan Rajarajan, and Marios Lestas. Combined Banzhaf & Diversity Index (CBDI) for critical node detection. Journal of Network and Computer Applications, 64(??):76–88, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451600062X.

Chen:2016:CEE

[1633] Lien-Wu Chen. Cooperative energy-efficient localization with node lifetime extension in mobile long-thin networks. Journal of Network and Computer Applications, 64(??):89–97, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516000461.

Paul:2016:EEC

[1634] Debdeep Paul, Wen-De Zhong, and Sanjay K. Bose. Energy efficient cloud service pricing: a two-timescale optimization approach. Journal of Network and Computer Applications, 64(??):98–112, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000400.

Barbeau:2016:CSU

[1635] Michel Barbeau, Gimer Cervera, Joaquin Garcia-Alfaro, and Evangelos Kranakis. Channel selection using a multiple radio model. Journal of Network and Computer Applications, 64(??):113–123, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516000515.

Melo:2016:OVN

[1636] Márcio Melo, Susana Sargento, and Jorge Carapinha. Optimal virtual network migration: a step closer for seamless resource mobility. Journal of Network and Computer Applications, 64(??):124– 136, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000588.

Qin:2016:WTM

[1637] Yongrui Qin, Quan Z. Sheng, Nickolas J. G. Falkner, Schahram Dustdar, Hua Wang, and Athanasios V. Vasilakos. When things matter: a survey on data-centric Internet of Things. Journal of Network and Computer Applications, 64(??):137–153, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000606.

Senouci:2016:FBS

[1638] Mustapha Reda Senouci, Abdelhamid Mellouk, Nadjib Aitsaadi, and Latifa Oukhellou. Fusion-based surveillance WSN deployment using Dempster–Shafer theory. Journal of Network and Computer Applications, 64(??):154–166, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000576.

Liu:2016:NCB

[1639] Yan Liu and Shun-Zheng Yu. Network coding-based multisource content delivery in Content Centric Networking. Journal of Network and Computer Applications, 64(??):167–175, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516000813.

Rahman:2016:CBD

[1640] Atiq Ur Rahman, Abdulrahman Alharby, Halabi Hasbullah, and Khalid Almuzaini. Corona based deployment strategies in wireless sensor network: a survey. Journal of Network and Computer Applications, 64(??): 176–193, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000771.

Zhan:2016:CBW

[1641] Cheng Zhan and Fuyuan Xiao. Coding based wireless broadcast scheduling in real time applications. Journal of Network and Computer Applications, 64(??):194–203, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000783.

Li:2016:NCC

[1642] Lingzhi Li, Zhe Yang, Jin Wang, Shukui Zhang, and Yanqin Zhu. Network coding with crowdsourcing-based trajectory estimation for vehicular networks. Journal of Network and Computer Applications, 64(??): 204–215, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000758.

Segundo:2016:DRS

[1643] Fabio Rafael Segundo, Eraldo Silveira e Silva, and Jean-Marie Farines. A DTN routing strategy based on neural networks for urban bus transportation system. Journal of Network and Computer Applications, 64(??):216–228, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451600076X.

Milani:2016:CRD

[1644] Bahareh Alami Milani and Nima Jafari Navimipour. A comprehensive review of the data replication techniques in the cloud environments: Major trends and future directions. Journal of Network and Computer Applications, 64(??):229–238, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000795.

Hashim:2016:OPR

[1645] Hashim A. Hashim, B. O. Ayinde, and M. A. Abido. Optimal placement of relay nodes in wireless sensor network using artificial bee colony algorithm. Journal of Network and Computer Applications, 64(??):239– 248, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000527.

Anonymous:2016:EBe

[1646] Anonymous. Editorial board. Journal of Network and Computer Applications, 64(??):ifc, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516001016.

Zhao:2016:MES

[1647] Haitao Zhao, Jibo Wei, Nurul I. Sarkar, and Shengchun Huang. E-MAC: an evolutionary solution for collision avoidance in wireless ad hoc networks. Journal of Network and Computer Applications, 65(??): 1–11, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451600093X.

Kaya:2016:AMC

[1648] Mahir Kaya, Altan Koçyigit, and P. Erhan Eren. An adaptive mobile cloud computing framework using a call graph based model. Journal of Network and Computer Applications, 65(??):12–35, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000965.

Lee:2016:NPM

[1649] Jungwon Lee, Miran Shim, and Hyesook Lim. Name prefix matching using bloom filter pre-searching for content centric network. Journal of Network and Computer Applications, 65(??):36–47, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000916.

Chaabouni:2016:PSP

[1650] Rahma Chaabouni, Marc Sánchez-Artigas, Pedro García-López, and Lluís Pàmies-Juàrez. The power of swarming in personal clouds under bandwidth budget. Journal of Network and Computer Applications, 65(??):48–71, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000801.

Seo:2016:DSB

[1651] Jihun Seo, Keuchul Cho, Wooseong Cho, Gisu Park, and Kijun Han. A discovery scheme based on carrier sensing in self-organizing Bluetooth Low Energy networks. Journal of Network and Computer Applications, 65(??):72–83, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000564.

Tan:2016:HST

[1652] Wilson M. Tan and Stephen A. Jarvis. Heuristic solutions to the target identifiability problem in directional sensor networks. Journal of Network and Computer Applications, 65(??):84–102, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000941.

Patel:2016:SLB

[1653] Deepak Kumar Patel, Devashree Tripathy, and C. R. Tripathy. Survey of load balancing techniques for Grid. Journal of Network and Computer Applications, 65(??):103–119, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000953.

Qu:2016:SRC

[1654] Ying Qu, Bryan Ng, and Winston Seah. A survey of routing and channel assignment in multi-channel multi-radio WMNs. Journal of Network and Computer Applications, 65(??):120–130, April 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300054.

Gillani:2016:EMI

[1655] Fida Gillani, Ehab Al-Shaer, and Basil AsSadhan. Economic metric to improve spam detectors. Journal of Network and Computer Applications, 65(??):131–143, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451630008X.

Shyam:2016:VRP

[1656] Gopal Kirshna Shyam and Sunilkumar S. Manvi. Virtual resource prediction in cloud environment: a Bayesian approach. Journal of Network and Computer Applications, 65(??):144–154, April 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300091.

Khasawneh:2016:PTC

[1657] Mahmoud Khasawneh, Saed Alrabaee, Anjali Agarwal, Nishith Goel, and Marzia Zaman. Power trading in cognitive radio networks. Journal of Network and Computer Applications, 65(??):155–166, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516000977.

Qu:2016:RCE

[1658] Chenhao Qu, Rodrigo N. Calheiros, and Rajkumar Buyya. A reliable and cost-efficient auto-scaling system for web applications using heterogeneous spot instances. Journal of Network and Computer Applications, 65(??):167–180, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300078.

Atiquzzaman:2016:EBP

[1659] Mohammed Atiquzzaman. Editorial: 2015 Best Paper Awards. Journal of Network and Computer Applications, 65(??):A1–A3, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451600117X.

Anonymous:2016:EBf

[1660] Anonymous. Editorial board. Journal of Network and Computer Applications, 65(??):ifc, April 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516001119.

Folino:2016:EBC

[1661] Gianluigi Folino and Pietro Sabatino. Ensemble based collaborative and distributed intrusion detection systems: a survey. Journal of Network and Computer Applications, 66(??):1–16, May 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300248.

Garg:2016:CAC

[1662] Neenu Garg and Seema Bawa. Comparative analysis of cloud data integrity auditing protocols. Journal of Network and Computer Applications, 66(??):17–32, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300224.

Otebolaku:2016:UCR

[1663] Abayomi Moradeyo Otebolaku and Maria Teresa Andrade. User context recognition using smartphone sensors and classification models. Journal of Network and Computer Applications, 66(??):33–51, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516300261.

Li:2016:HHR

[1664] Shiyi Li, Qiang Cao, Shenggang Wan, Lu Qian, and Changsheng Xie. HRSPC: a hybrid redundancy scheme via exploring computational locality to support fast recovery and high reliability in distributed storage systems. Journal of Network and Computer Applications, 66(??): 52–63, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000503.

Masdari:2016:TWS

[1665] Mohammad Masdari, Sima ValiKardan, Zahra Shahi, and Sonay Imani Azar. Towards workflow scheduling in cloud computing: a comprehensive analysis. Journal of Network and Computer Applications, 66(??): 64–82, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451600045X.

Verma:2016:MMM

[1666] Pawan Kumar Verma, Rajesh Verma, Arun Prakash, Ashish Agrawal, Kshirasagar Naik, Rajeev Tripathi, Maazen Alsabaan, Tarek Khalifa, Tamer Abdelkader, and Abdulhakim Abogharaf. Machine-to-machine (M2M) communications: a survey. Journal of Network and Computer Applications, 66(??):83–105, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000990.

Masdari:2016:OVM

[1667] Mohammad Masdari, Sayyid Shahab Nabavi, and Vafa Ahmadi. An overview of virtual machine placement schemes in cloud computing. Journal of Network and Computer Applications, 66(??):106– 127, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000291.

Jiang:2016:AGS

[1668] Fuu-Cheng Jiang, Ilsun You, and Fang-Yi Leu. Approaching green sensor field using queue-based optimization technique. Journal of Network and Computer Applications, 66(??):128–136, May 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515002829.

Phuttharak:2016:MCP

[1669] Jurairat Phuttharak and Seng W. Loke. Mobile crowdsourcing in peer-to-peer opportunistic networks: Energy usage and response analysis. Journal of Network and Computer Applications, 66(??):137– 150, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300285.

Joshi:2016:RCR

[1670] Yatish K. Joshi and Mohamed Younis. Restoring connectivity in a resource constrained WSN. Journal of Network and Computer Applications, 66(??):151–165, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300212.

Grana:2016:LRA

[1671] Justin Grana, David Wolpert, Joshua Neil, Dongping Xie, Tanmoy Bhattacharya, and Russell Bent. A likelihood ratio anomaly detector for identifying within-perimeter computer network attacks. Journal of Network and Computer Applications, 66(??):166–179, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300157.

Qureshi:2016:PPP

[1672] Amna Qureshi, David Megías, and Helena Rifà-Pous. PSUM: Peerto-peer multimedia content distribution using collusion-resistant fingerprinting. Journal of Network and Computer Applications, 66(??):180– 197, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300145.

Airehrour:2016:SRI

[1673] David Airehrour, Jairo Gutierrez, and Sayan Kumar Ray. Secure routing for Internet of Things: a survey. Journal of Network and Computer Applications, 66(??):198–213, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516300133.

Khan:2016:NFR

[1674] Suleman Khan, Abdullah Gani, Ainuddin Wahid Abdul Wahab, Muhammad Shiraz, and Iftikhar Ahmad. Network forensics: Review, taxonomy, and open challenges. Journal of Network and Computer Applications, 66(??):214–235, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300121.

Mohapatra:2016:BDA

[1675] Suvendu Kumar Mohapatra, Prasan Kumar Sahoo, and Shih-Lin Wu. Big data analytic architecture for intruder detection in heterogeneous wireless sensor networks. Journal of Network and Computer Applications, 66(??):236–249, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451630011X.

Chunlin:2016:FSP

[1676] Li Chunlin, Yan Xin, and Li LaYuan. Flexible service provisioning based on context constraint for enhancing user experience in service oriented mobile cloud. Journal of Network and Computer Applications, 66(??):250–261, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300108.

Anonymous:2016:EBg

[1677] Anonymous. Editorial board. Journal of Network and Computer Applications, 66(??):ifc, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300376.

Masoudi:2016:SDN

[1678] Rahim Masoudi and Ali Ghaffari. Software defined networks: a survey. Journal of Network and Computer Applications, 67(??):1– 25, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300297.

Chakraborty:2016:IRS

[1679] Suchetana Chakraborty, Sandip Chakraborty, Sukumar Nandi, and Sushanta Karmakar. Impact of redundant sensor deployment over data gathering performance: a model based approach. Journal of Network and Computer Applications, 67(??):26–42, May 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451630025X.

Liu:2016:NTR

[1680] Xuxun Liu. A novel transmission range adjustment strategy for energy hole avoiding in wireless sensor networks. Journal of Network and Computer Applications, 67(??):43–52, May 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300066.

Abbas:2016:NGP

[1681] Huda Saleh Abbas and Mark A. Gregory. The next generation of passive optical networks: a review. Journal of Network and Computer Applications, 67(??):53–74, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000989.

Pelka:2016:SCT

[1682] Mathias Pelka and Horst Hellbrück. Survey of challenges and towards a unified architecture for location systems. Journal of Network and Computer Applications, 67(??):75–85, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516000928.

Bartalos:2016:MEA

[1683] Peter Bartalos, Yi Wei, M. Brian Blake, Haluk Damgacioglu, Iman Saleh, and Nurcin Celik. Modeling energy-aware web services and application. Journal of Network and Computer Applications, 67(??):

86–98, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000448.

Diaz:2016:SAC

[1684] Manuel Díaz, Cristian Martín, and Bartolomé Rubio. State-of-the-art, challenges, and open issues in the integration of Internet of Things and cloud computing. Journal of Network and Computer Applications, 67(??):99–117, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451600028X.

Zarepour:2016:CPS

[1685] Eisa Zarepour, Abdul Alim Abd Karim, Ajay Sharma, Mahbub Hassan, and Aruna Seneviratne. Characterizing power saving for device-to-device browser cache cooperation. Journal of Network and Computer Applications, 67(??):118–127, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000047.

Sobin:2016:SRD

[1686] Sobin CC, Vaskar Raychoudhury, Gustavo Marfia, and Ankita Singla. A survey of routing and data dissemination in Delay Tolerant Networks. Journal of Network and Computer Applications, 67(??):128– 146, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516000035.

Osanaiye:2016:DDS

[1687] Opeyemi Osanaiye, Kim-Kwang Raymond Choo, and Mqhele Dlodlo. Distributed denial of service (DDoS) resilience in cloud: Review and conceptual cloud DDoS mitigation framework. Journal of Network and Computer Applications, 67(??):147–165, May 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516000023.

Anonymous:2016:EBh

[1688] Anonymous. Editorial board. Journal of Network and Computer Applications, 67(??):ifc, May 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300510.

Ahad:2016:NNW

[1689] Nauman Ahad, Junaid Qadir, and Nasir Ahsan. Neural networks in wireless networks: Techniques, applications and guidelines. Journal of Network and Computer Applications, 68(??):1–27, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300492.

Claeys:2016:HMT

[1690] Maxim Claeys, Daphné Tuncer, Jeroen Famaey, Marinos Charalambides, Steven Latré, George Pavlou, and Filip De Turck. Hybrid multi-tenant cache management for virtualized ISP networks. Journal of Network and Computer Applications, 68(??):28–41, June 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300479.

Ali:2016:EET

[1691] Anum Ali, Ghalib A. Shah, and Junaid Arshad. Energy efficient techniques for M2M communication: a survey. Journal of Network and Computer Applications, 68(??):42–55, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516300467.

Wang:2016:TCC

[1692] Zhiming Wang, Xiaoping Zeng, Xue Liu, Man Xu, Ya Wen, and Li Chen. TCP congestion control algorithm for heterogeneous Internet. Journal of Network and Computer Applications, 68(??):56–64, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300327.

Cui:2016:SAD

[1693] Yunhe Cui, Lianshan Yan, Saifei Li, Huanlai Xing, Wei Pan, Jian Zhu, and Xiaoyang Zheng. SD-Anti-DDoS: Fast and efficient DDoS defense in software-defined networks. Journal of Network and Computer Applications, 68(??):65–79, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300480.

Jiang:2016:EAD

[1694] Han-Peng Jiang, David Chuck, and Wei-Mei Chen. Energy-aware data center networks. Journal of Network and Computer Applications, 68(??):80–89, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300455.

Abdallah:2016:FDS

[1695] Aisha Abdallah, Mohd Aizaini Maarof, and Anazida Zainal. Fraud detection system: a survey. Journal of Network and Computer Applications, 68(??):90–113, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300571.

Son:2016:MMS

[1696] Tran The Son, Hoa Le-Minh, and Nauman Aslam. MSAR: a metric selfadaptive routing model for Mobile Ad Hoc Networks. Journal of Network and Computer Applications, 68(??):114–125, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300601.

Li:2016:SOB

[1697] Wenjuan Li, Weizhi Meng, and Lam For Kwok. A survey on OpenFlowbased Software Defined Networks: Security challenges and countermeasures. Journal of Network and Computer Applications, 68(??):126– 139, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300613.

Qin:2016:ATD

[1698] Yang Qin, Weihong Yang, Yibin Ye, and Yao Shi. Analysis for TCP in data center networks: Outcast and Incast. Journal of Network and Computer Applications, 68(??):140–150, June 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300649.

Hu:2016:RTS

[1699] Wenbin Hu, Zhenyu Qiu, Huan Wang, and Liping Yan. A realtime scheduling algorithm for on-demand wireless XML data broadcasting. Journal of Network and Computer Applications, 68(??):151– 163, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300595.

Chatterjee:2016:SAS

[1700] Bijoy Chand Chatterjee, Waya Fadini, and Eiji Oki. A spectrum allocation scheme based on first-last-exact fit policy for elastic optical networks. Journal of Network and Computer Applications, 68(??):164– 172, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300017.

Madni:2016:RSI

[1701] Syed Hamid Hussain Madni, Muhammad Shafie Abd Latiff, Yahaya Coulibaly, and Shafi’i Muhammad Abdulhamid. Resource scheduling for infrastructure as a service (IaaS) in cloud computing: Challenges and opportunities. Journal of Network and Computer Applications, 68(??):173–200, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300674.

Meng:2016:JTA

[1702] Weizhi Meng, Wang Hao Lee, S. R. Murali, and S. P. T. Krishnan. JuiceCaster: Towards automatic juice filming attacks on smartphones. Journal of Network and Computer Applications, 68(??):201– 212, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300650.

Bindu:2016:MSN

[1703] P. V. Bindu and P. Santhi Thilagam. Mining social networks for anomalies: Methods and challenges. Journal of Network and Computer Applications, 68(??):213–229, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300029.

Anonymous:2016:EBi

[1704] Anonymous. Editorial board. Journal of Network and Computer Applications, 68(??):ifc, June 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300881.

Trajano:2016:TPL

[1705] Alex F. R. Trajano and Marcial P. Fernandez. Two-phase load balancing of In-Memory Key-Value Storages using Network Functions Virtualization (NFV). Journal of Network and Computer Applications, 69(??): 1–13, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300789.

Hesselbach:2016:CNL

[1706] Xavier Hesselbach, Josè Roberto Amazonas, Santi Villanueva, and Juan Felipe Botero. Coordinated node and link mapping VNE using a new paths algebra strategy. Journal of Network and Computer Applications, 69(??):14–26, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451630073X.

Panigrahi:2016:FTD

[1707] T. Panigrahi, Meenakshi Panda, and G. Panda. Fault tolerant distributed estimation in wireless sensor networks. Journal of Network and Computer Applications, 69(??):27–39, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516300777.

Plesca:2016:MPO

[1708] Cezar Plesca, Vincent Charvillat, and Wei Tsang Ooi. Multimedia prefetching with optimal Markovian policies. Journal of Network and Computer Applications, 69(??):40–53, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516300807.

ElKhawaga:2016:ACB

[1709] Sally E. El Khawaga, Ahmed I. Saleh, and Hesham A. Ali. An Administrative Cluster-based Cooperative Caching (ACCC) strategy for Mobile Ad Hoc Networks. Journal of Network and Computer Applications, 69(??):54–76, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300820.

Ari:2016:PEC

[1710] Ado Adamou Abba Ari, Blaise Omer Yenke, Nabila Labraoui, Irepran Damakoa, and Abdelhak Gueroui. A power efficient cluster-based routing algorithm for wireless sensor networks: Honeybees swarm intelligence based approach. Journal of Network and Computer Applications, 69(??):77–97, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300741.

Gomes:2016:CEB

[1711] Rafael L. Gomes, Luiz F. Bittencourt, Edmundo R. M. Madeira, Eduardo Cerqueira, and Mario Gerla. A combined energy-bandwidth approach to allocate resilient virtual software defined networks. Journal of Network and Computer Applications, 69(??):98–106, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300728.

Yuanfeng:2016:FIL

[1712] Du Yuanfeng, Yang Dongkai, Yang Huilin, and Xiu Chundi. Flexible indoor localization and tracking system based on mobile phone. Journal of Network and Computer Applications, 69(??):107–116, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516300716.

Raj:2016:PMV

[1713] V. K. Mohan Raj and R. Shriram. Power management in virtualized datacenter — a survey. Journal of Network and Computer Applications, 69(??):117–133, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300704.

Moyano:2016:MDA

[1714] Francisco Moyano, Carmen Fernandez-Gago, and Javier Lopez. A model-driven approach for engineering trust and reputation into software services. Journal of Network and Computer Applications, 69(??): 134–151, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300698.

Do:2016:DSS

[1715] Cuong T. Do, Nguyen H. Tran, Eui-Nam Huh, Choong Seon Hong, Dusit Niyato, and Zhu Han. Dynamics of service selection and provider pricing game in heterogeneous cloud market. Journal of Network and Computer Applications, 69(??):152–165, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516300625.

Anonymous:2016:EBj

[1716] Anonymous. Editorial board. Journal of Network and Computer Applications, 69(??):ifc, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300972.

Luizelli:2016:HPN

[1717] Marcelo Caggiani Luizelli, Leonardo Richter Bays, Luciana Salete Buriol, Marinho Pilla Barcellos, and Luciano Paschoal Gaspary. How physical network topologies affect virtual network embedding quality: a characterization study based on ISP and datacenter networks. Journal of Network and Computer Applications, 70(??):1–16, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300959.

Saghiri:2016:ADC

[1718] Ali Mohammad Saghiri and Mohammad Reza Meybodi. An approach for designing cognitive engines in cognitive peer-to-peer networks. Journal of Network and Computer Applications, 70(??):17–40, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301084.

Niu:2016:FFU

[1719] Jianwei Niu, Lei Wang, Xiting Liu, and Shui Yu. FUIR: Fusing user and item information to deal with data sparsity by using side information in recommendation systems. Journal of Network and Computer Applications, 70(??):41–50, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300856.

Amah:2016:TNG

[1720] Tekenate E. Amah, Maznah Kamat, Waldir Moreira, Kamalrulnizam Abu Bakar, Satria Mandala, and Marcos A. Batista. Towards next-generation routing protocols for pocket switched networks. Journal of Network and Computer Applications, 70(??):51–88, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516301072.

Liu:2016:MMI

[1721] Xiaoming Liu, Yadong Zhou, Chengchen Hu, and Xiaohong Guan. MIR-ACLE: a multiple independent random walks community parallel detection algorithm for big graphs. Journal of Network and Computer Applications, 70(??):89–101, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300947.

Barba-Jimenez:2016:CBV

[1722] Carlos Barba-Jimenez, Raul Ramirez-Velarde, Andrei Tchernykh, Ramón Rodríguez-Dagnino, Juan Nolazco-Flores, and Raul Perez-Cazares. Cloud based Video-on-Demand service model ensuring quality of service and scalability. Journal of Network and Computer Applications, 70(??):102–113, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300868.

Senel:2016:NRN

[1723] Fatih Senel and Mohamed Younis. Novel relay node placement algorithms for establishing connected topologies. Journal of Network and Computer Applications, 70(??):114–130, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516300832.

Cordero:2016:MPT

[1724] Juan Antonio Cordero. Multi-path TCP performance evaluation in dualhomed (wired/wireless) devices. Journal of Network and Computer Applications, 70(??):131–139, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300819.

Anan:2016:ENR

[1725] Muhammad Anan, Ala Al-Fuqaha, Nidal Nasser, Ting-Yu Mu, and Husnain Bustam. Empowering networking research and experimentation through Software-Defined Networking. Journal of Network and Computer Applications, 70(??):140–155, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516300790.

Li:2016:CAP

[1726] Fuliang Li, Jiahai Yang, Xingwei Wang, Tian Pan, Changqing An, and Jianping Wu. Characteristics analysis at prefix granularity: a case study in an IPv6 network. Journal of Network and Computer Applications, 70(??):156–170, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300030.

Peng:2016:BLG

[1727] Jian Peng, Kim-Kwang Raymond Choo, and Helen Ashman. Bit-level n-gram based forensic authorship analysis on social media: Identifying individuals from linguistic profiles. Journal of Network and Computer Applications, 70(??):171–182, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516300443.

Xue:2016:TFL

[1728] Lin Xue, Suman Kumar, Cheng Cui, Praveenkumar Kondikoppa, Chui-Hui Chiu, and Seung-Jong Park. Towards fair and low latency next generation high speed networks: Afcd queuing. Journal of Network and Computer Applications, 70(??):183–193, July 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300352.

Anonymous:2016:EBk

[1729] Anonymous. Editorial board. Journal of Network and Computer Applications, 70(??):ifc, July 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301138.

Al-Ameen:2016:ITI

[1730] Mahdi Nasrullah Al-Ameen and Matthew Wright. iPersea: Towards improving the Sybil-resilience of social DHT. Journal of Network and Computer Applications, 71(??):1–10, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516301096.

Khan:2016:SSI

[1731] Minhaj Ahmad Khan. A survey of security issues for cloud computing. Journal of Network and Computer Applications, 71(??):11– 29, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301060.

Shen:2016:RWM

[1732] Hang Shen and Guangwei Bai. Routing in wireless multimedia sensor networks: a survey and challenges ahead. Journal of Network and Computer Applications, 71(??):30–49, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516301102.

Yigit:2016:CAR

[1733] Melike Yigit, V. Cagri Gungor, Etimad Fadel, Laila Nassef, Nadine Akkari, and Ian F. Akyildiz. Channel-aware routing and priorityaware multi-channel scheduling for WSN-based smart grid applications. Journal of Network and Computer Applications, 71(??):50– 58, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301114.

Gonzalez-Manzano:2016:PPP

[1734] L. González-Manzano, José M. de Fuentes, Sergio Pastrana, Pedro Peris-Lopez, and Luis Hernández-Encinas. PAgIoT — Privacy-preserving Aggregation protocol for Internet of Things. Journal of Network and Computer Applications, 71(??):59–71, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516301199.

Zenia:2016:EER

[1735] Nusrat Zerin Zenia, Mohammed Aseeri, Muhammad R. Ahmed, Zamshed I. Chowdhury, and M. Shamim Kaiser. Energy-efficiency and reliability in MAC and routing protocols for underwater wireless sensor network: a survey. Journal of Network and Computer Applications, 71(??):72–85, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301230.

Milani:2016:LBM

[1736] Alireza Sadeghi Milani and Nima Jafari Navimipour. Load balancing mechanisms and techniques in the cloud environments: Systematic literature review and future trends. Journal of Network and Computer Applications, 71(??):86–98, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301217.

Gandotra:2016:DDC

[1737] Pimmy Gandotra and Rakesh Kumar Jha. Device-to-device communication in cellular networks: a survey. Journal of Network and Computer Applications, 71(??):99–117, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516301229.

Esch:2016:AES

[1738] Rodrigo R. Esch, Fábio Protti, and Valmir C. Barbosa. Adaptive event sensing in networks of autonomous mobile agents. Journal of Network and Computer Applications, 71(??):118–129, August 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300765.

Noura:2016:SIM

[1739] Mahda Noura and Rosdiadee Nordin. A survey on interference management for Device-to-Device (D2D) communication and its challenges in 5G networks. Journal of Network and Computer Applications, 71(??): 130–150, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300753.

Siddiqa:2016:SBD

[1740] Aisha Siddiqa, Ibrahim Abaker Targio Hashem, Ibrar Yaqoob, Mohsen Marjani, Shahabuddin Shamshirband, Abdullah Gani, and Fariza Nasaruddin. A survey of big data management: Taxonomy and stateof-the-art. Journal of Network and Computer Applications, 71(??):151– 166, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300583.

Wang:2016:TFI

[1741] Jingyuan Wang, Jiangtao Wen, Jun Zhang, Zhang Xiong, and Yuxing Han. TCP-FIT: an improved TCP algorithm for heterogeneous networks. Journal of Network and Computer Applications, 71(??):167– 180, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300340.

Seddiki:2016:LPP

[1742] Manel Seddiki and Mahfoud Benchaïba. 2P-Lookup: Popularity and proximity based P2P lookup mechanism over MANETs. Journal of Network and Computer Applications, 71(??):181–193, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300303.

Santos:2016:IBM

[1743] João Santos, Joel J. P. C. Rodrigues, Bruno M. C. Silva, João Casal, Kashif Saleem, and Victor Denisov. An IoT-based mobile gateway for intelligent personal assistants on mobile health environments. Journal of Network and Computer Applications, 71(??):194– 204, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300273.

Anonymous:2016:EBl

[1744] Anonymous. Editorial board. Journal of Network and Computer Applications, 71(??):ifc, August 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301333.

Attia:2016:GMO

[1745] Mohamed Attia, Hichem Sedjelmaci, Sidi Mohammed Senouci, and El-Hassane Aglzim. Game model to optimally combine electric vehicles with green and non-green sources into an end-to-end smart grid architecture. Journal of Network and Computer Applications, 72(??):1–13, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301254.

Peng:2016:UPI

[1746] Jian Peng, Kim-Kwang Raymond Choo, and Helen Ashman. User profiling in intrusion detection: a review. Journal of Network and Computer Applications, 72(??):14–27, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516301412.

Singh:2016:RPC

[1747] Kishor Singh and Sangman Moh. Routing protocols in cognitive radio ad hoc networks: a comprehensive review. Journal of Network and Computer Applications, 72(??):28–37, September 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301503.

Azzam:2016:GGB

[1748] Rana Azzam, Rabeb Mizouni, Hadi Otrok, Anis Ouali, and Shakti Singh. GRS: a Group-Based Recruitment System for mobile crowd sensing. Journal of Network and Computer Applications, 72(??):38–50, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300170.

Assis:2016:SCF

[1749] M. R. M. Assis and L. F. Bittencourt. A survey on cloud federation architectures: Identifying functional and non-functional properties. Journal of Network and Computer Applications, 72(??):51–71, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301436.

Malek:2016:BCV

[1750] Sakib Md. Bin Malek, Md. Muntakim Sadik, and Ashikur Rahman. On balanced k-coverage in visual sensor networks. Journal of Network and Computer Applications, 72(??):72–86, September 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301400.

Chaudhari:2016:TMA

[1751] Shilpa Shashikant Chaudhari and Rajashekhar C. Biradar. Traffic and mobility aware resource prediction using cognitive agent in mobile ad hoc networks. Journal of Network and Computer Applications, 72(??):87–103, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451630131X.

Qiu:2016:EER

[1752] Tie Qiu, Yuan Lv, Feng Xia, Ning Chen, Jiafu Wan, and Amr Tolba. ERGID: an efficient routing protocol for emergency response Internet of Things. Journal of Network and Computer Applications, 72 (??):104–112, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301308.

Mogaibel:2016:RCA

[1753] Hassen A. Mogaibel, Mohamed Othman, Shamala Subramaniam, and Nor Asilah Wati Abdul Hamid. Review of channel assignment approaches in multi-radio multi-channel wireless mesh network. Journal of Network and Computer Applications, 72(??):113–139, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301291.

Feng:2016:SSS

[1754] Jingyu Feng, Guangyue Lu, Honggang Wang, and Xuanhong Wang. Supporting secure spectrum sensing data transmission against SSDH attack in cognitive radio ad hoc networks. Journal of Network and Computer Applications, 72(??):140–149, September 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301266.

Hamdeni:2016:DPM

[1755] C. Hamdeni, T. Hamrouni, and F. Ben Charrada. Data popularity measurements in distributed systems: Survey and design directions. Journal of Network and Computer Applications, 72(??):150–161, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301205.

Hao:2016:DTC

[1756] Xiaochen Hao, Weijing Liu, Ning Yao, Dehua Geng, and Xida Li. Distributed topology construction algorithm to improve link quality and energy efficiency for wireless sensor networks. Journal of Network and Computer Applications, 72(??):162–170, September 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300686.

Anonymous:2016:EBm

[1757] Anonymous. Editorial board. Journal of Network and Computer Applications, 72(??):ifc, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301643.

Li:2016:INP

[1758] Xian Li, Qiuling Tang, and Changyin Sun. The impact of node position on outage performance of RF energy powered wireless sensor communication links in overlaid deployment scenario. Journal of Network and Computer Applications, 73(??):1–11, September 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301461.

Khan:2016:FCR

[1759] Ubaid Ullah Khan, Naqqash Dilshad, Mubashir Husain Rehmani, and Tariq Umer. Fairness in cognitive radio networks: Models, measurement methods, applications, and future research directions. Journal of Network and Computer Applications, 73(??):12–26, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516301527.

Moudi:2016:FTE

[1760] Mehrnaz Moudi, Mohamed Othman, Kweh Yeah Lun, and Amir Rizaan Abdul Rahiman. x-folded TM: an efficient topology for interconnection networks. Journal of Network and Computer Applications, 73(??):27–34, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301539.

Kaliappan:2016:EEE

[1761] M. Kaliappan, Susan Augustine, and B. Paramasivan. Enhancing energy efficiency and load balancing in mobile ad hoc network using dynamic genetic algorithms. Journal of Network and Computer Applications, 73(??):35–43, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301473.

Han:2016:HGB

[1762] Song Han, Xinbin Li, Zhixin Liu, and Xinping Guan. Hierarchical-game-based algorithm for downlink joint subchannel and power allocation in OFDMA femtocell networks. Journal of Network and Computer Applications, 73(??):44–56, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301515.

Karkouch:2016:DQI

[1763] Aimad Karkouch, Hajar Mousannif, Hassan Al Moatassime, and Thomas Noel. Data quality in Internet of Things: a state-of-the-art survey. Journal of Network and Computer Applications, 73(??):57–81, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301564.

Yao:2016:EFS

[1764] Guidan Yao, Zhigang Jin, and Yishan Su. An environment-friendly spectrum decision strategy for underwater acoustic networks. Journal of Network and Computer Applications, 73(??):82–93, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516301485.

Ivesic:2016:CCL

[1765] K. Ivesic, L. Skorin-Kapov, and M. Matijasevic. Corrigendum to: “Crosslayer QoE-driven admission control and resource allocation for adaptive multimedia services in LTE” [J. Netw. Comput. Appl. 46 (2014) 336–351]. Journal of Network and Computer Applications, 73(??):94– 97, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516300339. See [1343].

Do:2016:CMO

[1766] Tien Van Do, Nam H. Do, Ádám Horváth, and Jinting Wang. Corrigendum to “Modelling opportunistic spectrum renting in mobile cellular networks” [J. Netw. Comput. Appl. 52 (2015) 129–138]. Journal of Network and Computer Applications, 73(??):98, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804515001253.

Anonymous:2016:EBn

[1767] Anonymous. Editorial board. Journal of Network and Computer Applications, 73(??):ifc, September 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301801.

Lin:2016:HPI

[1768] Feng Lin, Gang Wang, Junhai Zhou, Shiwen Zhang, and Xin Yao. High-performance IPv6 address lookup in GPU-accelerated software routers. Journal of Network and Computer Applications, 74(??):1– 10, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301588.

Zhao:2016:NDH

[1769] Yanchao Zhao, Wenzhong Li, and Sanglu Lu. Navigation-driven handoff minimization in wireless networks. Journal of Network and Computer Applications, 74(??):11–20, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451630159X.

Tang:2016:UAR

[1770] Lin Tang, Hui Cao, Li Zheng, and Ningjian Huang. Uncertaintyaware RFID network planning for target detection and target location. Journal of Network and Computer Applications, 74(??):21–30, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301576.

Kogan:2016:LPF

[1771] Kirill Kogan, Alejandro López-Ortiz, Sergey I. Nikolenko, Gabriel Scalosub, and Michael Segal. Large profits or fast gains: a dilemma in maximizing throughput with applications to network processors. Journal of Network and Computer Applications, 74(??):31–43, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301540.

Lyu:2016:PPD

[1772] Chen Lyu, Shi-Feng Sun, Yuanyuan Zhang, Amit Pande, Haining Lu, and Dawu Gu. Privacy-preserving data sharing scheme over cloud for social applications. Journal of Network and Computer Applications, 74(??):44–55, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301606.

Khatua:2016:EAS

[1773] Manas Khatua and Sudip Misra. Exploiting anomalous slots for multiple channel access in IEEE 802.11 networks. Journal of Network and Computer Applications, 74(??):56–65, October 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301552.

Sharma:2016:REE

[1774] Yogesh Sharma, Bahman Javadi, Weisheng Si, and Daniel Sun. Reliability and energy efficiency in cloud computing systems: Survey and taxonomy. Journal of Network and Computer Applications, 74(??):66– 85, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301746.

dosSantos:2016:FRA

[1775] Daniel Ricardo dos Santos, Roberto Marinho, Gustavo Roecker Schmitt, Carla Merkle Westphall, and Carlos Becker Westphall. A framework and risk assessment approaches for risk-based access control in the cloud. Journal of Network and Computer Applications, 74(??):86–97, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301710.

Iqbal:2016:CSA

[1776] Salman Iqbal, Miss Laiha Mat Kiah, Babak Dhaghighi, Muzammil Hussain, Suleman Khan, Muhammad Khurram Khan, and Kim-Kwang Raymond Choo. On cloud security attacks: a taxonomy and intrusion detection and prevention as a service. Journal of Network and Computer Applications, 74(??):98–120, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516301771.

Sepulcre:2016:MQD

[1777] Miguel Sepulcre, Javier Gozalvez, and Baldomero Coll-Perales. Multipath QoS-driven routing protocol for industrial wireless networks. Journal of Network and Computer Applications, 74(??):121–132, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451630162X.

Emmanuel:2016:CTS

[1778] Michael Emmanuel and Ramesh Rayudu. Communication technologies for smart grid applications: a survey. Journal of Network and Computer Applications, 74(??):133–148, October 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301734.

Lin:2016:ARN

[1779] Chun-Cheng Lin, Teng-Huei Chen, and Hui-Hsin Chin. Adaptive router node placement with gateway positions and QoS constraints in dynamic wireless mesh networks. Journal of Network and Computer Applications, 74(??):149–164, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516300844.

Anonymous:2016:EBo

[1780] Anonymous. Editorial board. Journal of Network and Computer Applications, 74(??):ifc, October 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302041.

Zannat:2016:CPV

[1781] Hafsa Zannat, Taslima Akter, Mashrura Tasnim, and Ashikur Rahman. The coverage problem in visual sensor networks: a target oriented approach. Journal of Network and Computer Applications, 75(??):1–15, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301783.

Zorbas:2016:ODP

[1782] Dimitrios Zorbas, Luigi Di Puglia Pugliese, Tahiry Razafindralambo, and Francesca Guerriero. Optimal drone placement and cost-efficient target coverage. Journal of Network and Computer Applications, 75(??):16– 31, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301709.

Daldoul:2016:BNA

[1783] Yousri Daldoul, Djamal-Eddine Meddour, Toufik Ahmed, and Raouf Boutaba. Block Negative Acknowledgement protocol for reliable multicast in IEEE 802.11. Journal of Network and Computer Applications, 75(??):32–46, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301497.

Naha:2016:CAS

[1784] Ranesh Kumar Naha and Mohamed Othman. Cost-aware service brokering and performance sentient load balancing algorithms in the cloud. Journal of Network and Computer Applications, 75(??):47–57, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301874.

Razak:2016:RMB

[1785] Mohd Faizal Ab Razak, Nor Badrul Anuar, Rosli Salleh, and Ahmad Firdaus. The rise of “malware”: Bibliometric analysis of malware study. Journal of Network and Computer Applications, 75(??):58–76, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301904.

Verma:2016:NHM

[1786] Pawan Kumar Verma, Rajesh Verma, Arun Prakash, Rajeev Tripathi, and Kshirasagar Naik. A novel hybrid medium access control protocol for inter-M2M communications. Journal of Network and Computer Applications, 75(??):77–88, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301722.

Zhang:2016:PCR

[1787] Qiaoyun Zhang, Guilin Chen, Liang Zhao, and Chih-Yung Chang. Piconet construction and restructuring mechanisms for interference avoiding in Bluetooth PANs. Journal of Network and Computer Applications, 75(??):89–100, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301862.

Alaez:2016:TOS

[1788] Ricardo Marco Alaez, Jose M. Alcaraz Calero, Fatna Belqasmi, May El-Barachi, Mohamad Badra, and Omar Alfandi. Towards an open source architecture for multi-operator LTE core networks. Journal of Network and Computer Applications, 75(??):101–109, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301916.

DArienzo:2016:GEE

[1789] M. D’Arienzo and S. P. Romano. GOSPF: an energy efficient implementation of the OSPF routing protocol. Journal of Network and Computer Applications, 75(??):110–127, November 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300662.

Ates:2016:HTG

[1790] Ozgur Ates, Selcuk Keskin, and Taskin Kocak. High throughput graphics processing unit based Fano decoder. Journal of Network and Computer Applications, 75(??):128–137, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516301886.

Bhamare:2016:SSF

[1791] Deval Bhamare, Raj Jain, Mohammed Samaka, and Aiman Erbad. A survey on service function chaining. Journal of Network and Computer Applications, 75(??):138–155, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516301989.

Hong:2016:TSS

[1792] Jin B. Hong and Dong Seong Kim. Towards scalable security analysis using multi-layered security models. Journal of Network and Computer Applications, 75(??):156–168, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516301928.

Utsumi:2016:NEC

[1793] Satoshi Utsumi, Salahuddin Muhammad Salim Zabir, and Sumet Prabhavat. A new explicit congestion notification scheme for satellite IP networks. Journal of Network and Computer Applications, 75(??):169– 180, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301953.

Konstantopoulos:2016:EEH

[1794] Charalampos Konstantopoulos, Grammati Pantziou, Ioannis E. Venetis, and Damianos Gavalas. Efficient event handling in Wireless Sensor and Actor Networks: an on-line computation approach. Journal of Network and Computer Applications, 75(??):181–199, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451630193X.

Singh:2016:SCC

[1795] Saurabh Singh, Young-Sik Jeong, and Jong Hyuk Park. A survey on cloud computing security: Issues, threats, and solutions. Journal of Network and Computer Applications, 75(??):200–222, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516301990.

Mansouri:2016:MMC

[1796] Yaser Mansouri and Rajkumar Buyya. To move or not to move: Cost optimization in a dual cloud-based storage architecture. Journal of Network and Computer Applications, 75(??):223–235, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301977.

Habib:2016:PPF

[1797] Sana Habib, Junaid Qadir, Anwaar Ali, Durdana Habib, Ming Li, and Arjuna Sathiaseelan. The past, present, and future of transportlayer multipath. Journal of Network and Computer Applications, 75 (??):236–258, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302028.

Warif:2016:CMF

[1798] Nor Bakiah Abd Warif, Ainuddin Wahid Abdul Wahab, Mohd Yamani Idna Idris, Roziana Ramli, Rosli Salleh, Shahaboddin Shamshirband, and Kim-Kwang Raymond Choo. Copy–move forgery detection: Survey, challenges and future directions. Journal of Network and Computer Applications, 75(??):259–278, November 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302144.

Asif:2016:OBS

[1799] Waqar Asif, Marios Lestas, Hassaan Khaliq Qureshi, and Muttukrishnan Rajarajan. Optimization based spectral partitioning for node criticality assessment. Journal of Network and Computer Applications, 75(??):279–292, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302004.

Bartolomeu:2016:SLP

[1800] Paulo Bartolomeu, Muhammad Alam, Joaquim Ferreira, and José Fonseca. Survey on low power real-time wireless MAC protocols. Journal of Network and Computer Applications, 75(??):293–316, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516302016.

Khan:2016:BSW

[1801] Abid Khan, Farhana Jabeen, Farah Naz, Sabah Suhail, Mansoor Ahmed, and Sarfraz Nawaz. Buyer seller watermarking protocols issues and challenges — a survey. Journal of Network and Computer Applications, 75(??):317–334, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301941.

Shuja:2016:TNC

[1802] Junaid Shuja, Abdullah Gani, Muhammad Habib ur Rehman, Ejaz Ahmed, Sajjad A. Madani, Muhammad Khurram Khan, and Kwangman Ko. Towards native code offloading based MCC frameworks for multimedia applications: a survey. Journal of Network and Computer Applications, 75(??):335–354, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516301898.

Park:2016:WSS

[1803] Minjae Park, Hyun Ahn, and Kwanghoon Pio Kim. Workflow-supported social networks: Discovery, analyses, and system. Journal of Network and Computer Applications, 75(??):355–373, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516301758.

Shao:2016:VRN

[1804] Xun Shao, Masahiro Jibiki, Yuuichi Teranishi, and Nozomu Nishinaga. A virtual replica node-based flash crowds alleviation method for sensor overlay networks. Journal of Network and Computer Applications, 75(??):374–384, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302120.

Nowsheen:2016:PPR

[1805] Nusrat Nowsheen, Gour Karmakar, and Joarder Kamruzzaman. PRADD: a path reliability-aware data delivery protocol for underwater acoustic sensor networks. Journal of Network and Computer Applications, 75(??):385–397, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516000394.

Anonymous:2016:EBp

[1806] Anonymous. Editorial board. Journal of Network and Computer Applications, 75(??):ifc, November 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302247.

Acarali:2016:SAF

[1807] Dilara Acarali, Muttukrishnan Rajarajan, Nikos Komninos, and Ian Herwono. Survey of approaches and features for the identification of HTTP-based botnet traffic. Journal of Network and Computer Applications, 76(??):1–15, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302363.

Nie:2016:TMP

[1808] Laisen Nie, Dingde Jiang, Lei Guo, and Shui Yu. Traffic matrix prediction and estimation based on deep learning in large-scale IP backbone networks. Journal of Network and Computer Applications, 76(??):16– 22, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302351.

Nafi:2016:SSG

[1809] Nazmus S. Nafi, Khandakar Ahmed, Mark A. Gregory, and Manoj Datta. A survey of smart grid architectures, applications, benefits and standardization. Journal of Network and Computer Applications, 76(??):23–36, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302314.

Jiang:2016:UTC

[1810] Qi Jiang, Jianfeng Ma, Fushan Wei, Youliang Tian, Jian Shen, and Yuanyuan Yang. An untraceable temporal-credential-based two-factor authentication scheme using ECC for wireless sensor networks. Journal of Network and Computer Applications, 76(??):37–48, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516302302.

Goudarzi:2016:KSP

[1811] Parisa Goudarzi, Hadi Tabatabaee Malazi, and Mahmood Ahmadi. Khorramshahr: a scalable peer to peer architecture for port warehouse management system. Journal of Network and Computer Applications, 76(??):49–59, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302223.

Lu:2016:HPT

[1812] Chun-Nan Lu, Chun-Ying Huang, Ying-Dar Lin, and Yuan-Cheng Lai. High performance traffic classification based on message size sequence and distribution. Journal of Network and Computer Applications, 76(??):60–74, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451630220X.

Ahvar:2016:AUL

[1813] Ehsan Ahvar, Nafiseh Daneshgar-Moghaddam, Antonio M. Ortiz, Gyu Myoung Lee, and Noel Crespi. On analyzing user location discovery methods in smart homes: a taxonomy and survey. Journal of Network and Computer Applications, 76(??):75–86, December 2016. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302181.

Hwang:2016:SGD

[1814] Ren-Hung Hwang, Yu-Ling Hsueh, Jang-Jiin Wu, and Fu-Hui Huang. SocialHide: a generic distributed framework for location privacy protection. Journal of Network and Computer Applications, 76(??):87–100, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302156.

Saxena:2016:FSM

[1815] Divya Saxena and Vaskar Raychoudhury. N-FIB: Scalable, memory efficient name-based forwarding. Journal of Network and Computer Applications, 76(??):101–109, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516302132.

Anonymous:2016:EBq

[1816] Anonymous. Editorial board. Journal of Network and Computer Applications, 76(??):ifc, December 2016. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302466.

Tsolkas:2017:SPQ

[1817] Dimitris Tsolkas, Eirini Liotou, Nikos Passas, and Lazaros Merakos. A survey on parametric QoE estimation for popular services. Journal of Network and Computer Applications, 77(??):1–17, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516302533.

Mishra:2017:IDT

[1818] Preeti Mishra, Emmanuel S. Pilli, Vijay Varadharajan, and Udaya Tupakula. Intrusion detection techniques in cloud environment: a survey. Journal of Network and Computer Applications, 77(??):18–47, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302417.

Walikar:2017:SHR

[1819] Gyanappa A. Walikar and Rajashekar C. Biradar. A survey on hybrid routing mechanisms in mobile ad hoc networks. Journal of Network and Computer Applications, 77(??):48–63, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302430.

Huang:2017:SBS

[1820] Baogui Huang, Jiguo Yu, Xiuzhen Cheng, Honglong Chen, and Hang Liu. SINR based shortest link scheduling with oblivious power control in wireless networks. Journal of Network and Computer Applications, 77(??):64–72, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302442.

Aznoli:2017:CSR

[1821] Fariba Aznoli and Nima Jafari Navimipour. Cloud services recommendation: Reviewing the recent advances and suggesting the future research directions. Journal of Network and Computer Applications, 77(??):73–86, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302375.

Liaqat:2017:FCR

[1822] Misbah Liaqat, Victor Chang, Abdullah Gani, Siti Hafizah Ab Hamid, Muhammad Toseef, Umar Shoaib, and Rana Liaqat Ali. Federated cloud resource management: Review and discussion. Journal of Network and Computer Applications, 77(??):87–105, January 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302387.

Dabaghi:2017:SGR

[1823] Fahimeh Dabaghi, Zeinab Movahedi, and Rami Langar. A survey on green routing protocols using sleep-scheduling in wired networks. Journal of Network and Computer Applications, 77(??):106–122, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451630234X.

Cheng:2017:CSB

[1824] Long Cheng, Jianwei Niu, Linghe Kong, Chengwen Luo, Yu Gu, Wenbo He, and Sajal K. Das. Compressive sensing based data quality improvement for crowd-sensing applications. Journal of Network and Computer Applications, 77(??):123–134, January 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302338.

Li:2017:ECI

[1825] Wenjuan Li, Weizhi Meng, Lam-For Kwok, and Horace H. S. Ip. Enhancing collaborative intrusion detection networks against insider attacks using supervised intrusion sensitivity-based trust management model. Journal of Network and Computer Applications, 77(??):135–145, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302211.

Lin:2017:DIV

[1826] Chen Lin, Zhidong Shen, Qian Chen, and Frederick T. Sheldon. A data integrity verification scheme in mobile cloud computing. Journal of Network and Computer Applications, 77(??):146–151, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451630176X.

Anonymous:2017:EBa

[1827] Anonymous. Editorial board. Journal of Network and Computer Applications, 77(??):ifc, January 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302636.

Sigwele:2017:EEC

[1828] Tshiamo Sigwele, Atm S. Alam, Prashant Pillai, and Yim F. Hu. Energy-efficient cloud radio access networks by cloud based workload consolidation for 5G. Journal of Network and Computer Applications, 78(??):1–8, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302740.

Gandotra:2017:SDD

[1829] Pimmy Gandotra, Rakesh Kumar Jha, and Sanjeev Jain. A survey on device-to-device (D2D) communication: Architecture and security issues. Journal of Network and Computer Applications, 78(??):9–29, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302727.

Garofalo:2017:DRT

[1830] Giuseppina Garofalo, Andrea Giordano, Patrizia Piro, Giandomenico Spezzano, and Andrea Vinci. A distributed real-time approach for mitigating CSO and flooding in urban drainage systems. Journal of Network and Computer Applications, 78(??):30–42, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302752.

Arianyan:2017:NFM

[1831] Ehsan Arianyan, Hassan Taheri, and Vahid Khoshdel. Novel fuzzy multi objective DVFS-aware consolidation heuristics for energy and SLA efficient resource management in cloud data centers. Journal of Network and Computer Applications, 78(??):43–61, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300169.

Wang:2017:DCT

[1832] Xiumin Wang, Xiaoming Chen, Chau Yuen, Weiwei Wu, Meng Zhang, and Cheng Zhan. Delay-cost tradeoff for virtual machine migration in cloud data centers. Journal of Network and Computer Applications, 78(??):62–72, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302739.

Kumar:2017:EEO

[1833] Gulshan Kumar and Mritunjay Kumar Rai. An energy efficient and optimized load balanced localization method using CDS with one-hop neighbourhood and genetic algorithm in WSNs. Journal of Network and Computer Applications, 78(??):73–82, January 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451630282X.

Rossi:2017:EPA

[1834] Fábio D. Rossi, Miguel G. Xavier, César A. F. De Rose, Rodrigo N. Calheiros, and Rajkumar Buyya. E-eco: Performance-aware energy-efficient cloud data center orchestration. Journal of Network and Computer Applications, 78(??):83–96, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516302569.

Bhattacharya:2017:SAT

[1835] Arani Bhattacharya and Pradipta De. A survey of adaptation techniques in computation offloading. Journal of Network and Computer Applications, 78(??):97–115, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516302570.

Kong:2017:TDR

[1836] Cuiyu Kong and I.-Hsuan Peng. Tradeoff design of radio resource scheduling for power and spectrum utilizations in LTE uplink systems.

Journal of Network and Computer Applications, 78(??):116–124, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302806.

Lin:2017:SMM

[1837] Ying-Dar Lin, Yuan-Cheng Lai, Hung-Yi Teng, Chun-Chieh Liao, and Yi-Chih Kao. Scalable multicasting with multiple shared trees in software defined networking. Journal of Network and Computer Applications, 78 (??):125–133, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302831.

Ghaderzadeh:2017:IIF

[1838] Abdulbaghi Ghaderzadeh, Mehdi Kargahi, and Midia Reshadi. In-FreD: Intelligent Free Rider Detection in collaborative distributed systems. Journal of Network and Computer Applications, 78(??):134–145, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302788.

Lopez-Benitez:2017:PMR

[1839] Miguel López-Benítez, Timothy D. Drysdale, Simon Hadfield, and Mohamed Ismaeel Maricar. Prototype for multidisciplinary research in the context of the Internet of Things. Journal of Network and Computer Applications, 78(??):146–161, January 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302922.

Meng:2017:BIB

[1840] Weizhi Meng, Wenjuan Li, Yang Xiang, and Kim-Kwang Raymond Choo. A Bayesian inference-based detection mechanism to defend medical smartphone networks against insider attacks. Journal of Network and Computer Applications, 78(??):162–169, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302818.

Hossain:2017:PSB

[1841] Md Akbar Hossain, Nurul I. Sarkar, Jairo Gutierrez, and William Liu. Performance study of block ACK and reverse direction in IEEE 802.11n using a Markov chain model. Journal of Network and Computer Applications, 78(??):170–179, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451630296X.

Ayoubi:2017:RMC

[1842] Sara Ayoubi, Chadi Assi, Yiheng Chen, Tarek Khalifa, and Khaled Bashir Shaban. Restoration methods for cloud multicast virtual networks. Journal of Network and Computer Applications, 78(??):180–190, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302764.

Ahmed:2017:RMA

[1843] Waqar Ahmed, Osman Hasan, Usman Pervez, and Junaid Qadir. Reliability modeling and analysis of communication networks. Journal of Network and Computer Applications, 78(??):191–215, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516302776.

Ghosal:2017:SKD

[1844] Amrita Ghosal, Subir Halder, and Stefano Chessa. Secure key design approaches using entropy harvesting in wireless sensor network: a survey. Journal of Network and Computer Applications, 78(??):216–230, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302880.

Mirhosseini:2017:QQQ

[1845] Mina Mirhosseini, Fatemeh Barani, and Hossein Nezamabadi-pour. QQIGSA: a quadrivalent quantum-inspired GSA and its application in optimal adaptive design of wireless sensor networks. Journal of Network and Computer Applications, 78(??):231–241, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302612.

Ahmed:2017:RPB

[1846] Mukhtiar Ahmed, Mazleena Salleh, and M. Ibrahim Channa. Routing protocols based on node mobility for Underwater Wireless Sensor Network (UWSN): a survey. Journal of Network and Computer Applications, 78(??):242–252, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302594.

Li:2017:QLA

[1847] Qing Li, Yong Jiang, Pengfei Duan, Mingwei Xu, and Xi Xiao. Quokka: Latency-Aware Middlebox Scheduling with dynamic resource allocation. Journal of Network and Computer Applications, 78(??):253–266, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302600.

Muhammed:2017:AFD

[1848] Thaha Muhammed and Riaz Ahmed Shaikh. An analysis of fault detection strategies in wireless sensor networks. Journal of Network and Computer Applications, 78(??):267–287, January 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302545.

Wen:2017:UEB

[1849] Sheng Wen, Jiaojiao Jiang, Bo Liu, Yang Xiang, and Wanlei Zhou. Using epidemic betweenness to measure the influence of users in complex networks. Journal of Network and Computer Applications, 78 (??):288–299, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302557.

Cheng:2017:DGP

[1850] Chien-Fu Cheng and Lung-Hao Li. Data gathering problem with the data importance consideration in Underwater Wireless Sensor Networks. Journal of Network and Computer Applications, 78(??):300–312, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302399.

Anonymous:2017:EBb

[1851] Anonymous. Editorial board. Journal of Network and Computer Applications, 78(??):ifc, January 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303241.

urRehman:2017:TNG

[1852] Muhammad Habib ur Rehman, Chee Sun Liew, Teh Ying Wah, and Muhammad Khurram Khan. Towards next-generation heterogeneous mobile data stream mining applications: Opportunities, challenges, and future research directions. Journal of Network and Computer Applications, 79(??):1–24, February 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302995.

kaisar:2017:DCS

[1853] Shahriar kaisar, Joarder Kamruzzaman, Gour Karmakar, and Iqbal Gondal. Decentralized content sharing among tourists in visiting hotspots. Journal of Network and Computer Applications, 79(??):25–40, February 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451630279X.

Adewole:2017:MAD

[1854] Kayode Sakariyah Adewole, Nor Badrul Anuar, Amirrudin Kamsin, Kasturi Dewi Varathan, and Syed Abdul Razak. Malicious accounts: Dark of the social networks. Journal of Network and Computer Applications, 79(??):41–67, February 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303009.

Rana:2017:SAP

[1855] Annie Ibrahim Rana and Brendan Jennings. Semantic aware processing of user defined inference rules to manage home networks. Journal of Network and Computer Applications, 79(??):68–87, February 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302892.

Singh:2017:CSI

[1856] Ashish Singh and Kakali Chatterjee. Cloud security issues and challenges: a survey. Journal of Network and Computer Applications, 79 (??):88–115, February 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302983.

Anonymous:2017:EBc

[1857] Anonymous. Editorial board. Journal of Network and Computer Applications, 79(??):ifc, February 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303344.

Lu:2017:DSR

[1858] Ting Lu, Guohua Liu, Wei Li, Shan Chang, and Wenjing Guo. Distributed sampling rate allocation for data quality maximization in rechargeable sensor networks. Journal of Network and Computer Applications, 80(??):1–9, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516303216.

Sahoo:2017:DAC

[1859] Prasan Kumar Sahoo and Jang-Ping Sheu. Design and analysis of collision free MAC for wireless sensor networks with or without data retransmission. Journal of Network and Computer Applications, 80 (??):10–21, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303174.

Manunza:2017:KRT

[1860] L. Manunza, S. Marseglia, and S. P. Romano. Kerberos: a real-time fraud detection system for IMS-enabled VoIP networks. Journal of Network and Computer Applications, 80(??):22–34, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303228.

Liu:2017:APA

[1861] Chunhong Liu, Chuanchang Liu, Yanlei Shang, Shiping Chen, Bo Cheng, and Junliang Chen. An adaptive prediction approach based on workload pattern discrimination in the cloud. Journal of Network and Computer Applications, 80(??):35–44, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516303198.

Deng:2017:EPC

[1862] Fan Deng and Li-Yong Zhang. Elimination of policy conflict to improve the PDP evaluation performance. Journal of Network and Computer Applications, 80(??):45–57, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516303010.

Busom:2017:PPR

[1863] N. Busom, R. Petrlic, F. Sebé, C. Sorge, and M. Valls. A privacypreserving reputation system with user rewards. Journal of Network and Computer Applications, 80(??):58–66, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303320.

Mostafaei:2017:SSA

[1864] Habib Mostafaei, Antonio Montieri, Valerio Persico, and Antonio Pescapé. A sleep scheduling approach based on learning automata for WSN partial coverage. Journal of Network and Computer Applications, 80(??):67–78, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303204.

Masdari:2017:MCB

[1865] Mohammad Masdari. Markov chain-based evaluation of the certificate status validations in hybrid MANETs. Journal of Network and Computer Applications, 80(??):79–89, February 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303459.

Wang:2017:JMR

[1866] Jihong Wang and Wenxiao Shi. Joint multicast routing and channel assignment for multi-radio multi-channel wireless mesh networks with hybrid traffic. Journal of Network and Computer Applications, 80 (??):90–108, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302909.

Barshan:2017:DED

[1867] Maryam Barshan, Hendrik Moens, Bruno Volckaert, and Filip De Turck. Design and evaluation of a dual dynamic adaptive reservation approach in media production networks. Journal of Network and Computer Applications, 80(??):109–122, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516303083.

Alves:2017:DPW

[1868] Maicon Melo Alves, Luci Pirmez, Silvana Rossetto, Flavia C. Delicato, Claudio M. de Farias, Paulo F. Pires, Igor L. dos Santos, and Albert Y. Zomaya. Damage prediction for wind turbines using wireless sensor and actuator networks. Journal of Network and Computer Applications, 80 (??):123–140, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303435.

Luo:2017:MIS

[1869] Chengwen Luo, Long Cheng, Hande Hong, Kartik Sankaran, Mun Choon Chan, Jianqiang Li, and Zhong Ming. From mapping to indoor semantic queries: Enabling zero-effort indoor environmental sensing. Journal of Network and Computer Applications, 80(??):141–151, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302910.

AlEroud:2017:ICA

[1870] Ahmed AlEroud and Izzat Alsmadi. Identifying cyber-attacks on software defined networks: an inference-based intrusion detection approach. Journal of Network and Computer Applications, 80(??):152–164, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303447.

Yang:2017:VMM

[1871] Chao-Tung Yang, Jung-Chun Liu, Shuo-Tsung Chen, and Kuan-Lung Huang. Virtual machine management system based on the power saving algorithm in cloud. Journal of Network and Computer Applications, 80 (??):165–180, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302971.

Arunkumar:2017:LAA

[1872] Saritha Arunkumar, Berker Soyluoglu, Murat Sensoy, Mudhakar Srivatsa, and Muttukrishnan Rajarajan. Location attestation and access control for mobile devices using GeoXACML. Journal of Network and Computer Applications, 80(??):181–188, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302958.

Boulaiche:2017:HHS

[1873] Mehdi Boulaiche and Louiza Bouallouche-Medjkoune. HSecGR: Highly Secure Geographic Routing. Journal of Network and Computer Applications, 80(??):189–199, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516303460.

Karakus:2017:QSQ

[1874] Murat Karakus and Arjan Durresi. Quality of Service (QoS) in Software Defined Networking (SDN): a survey. Journal of Network and Computer Applications, 80(??):200–218, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303186.

Goudarzi:2017:FHM

[1875] Mohammad Goudarzi, Mehran Zamani, and Abolfazl Toroghi Haghighat. A fast hybrid multi-site computation offloading for mobile cloud computing. Journal of Network and Computer Applications, 80(??):219–231, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303514.

Morelli:2017:ICN

[1876] Alessandro Morelli, Mauro Tortonesi, Cesare Stefanelli, and Niranjan Suri. Information-centric Networking in next-generation communications scenarios. Journal of Network and Computer Applications, 80 (??):232–250, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303423.

Anonymous:2017:EBd

[1877] Anonymous. Editorial board. Journal of Network and Computer Applications, 80(??):ifc, February 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300085.

Cogoni:2017:TCC

[1878] Marco Cogoni, Giovanni Busonera, Paolo Anedda, and Gianluigi Zanetti. Transition to congestion in communication/computation networks for near-optimal and sub-optimal resource management via Monte Carlo simulations. Journal of Network and Computer Applications, 81(??): 1–11, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303472.

Sharma:2017:TRV

[1879] Swati Sharma, Alefiya Hussain, and Huzur Saran. Towards repeatability & verifiability in networking experiments: a stochastic framework. Journal of Network and Computer Applications, 81(??):12–23, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301448.

Vakili:2017:CSR

[1880] Asrin Vakili and Nima Jafari Navimipour. Comprehensive and systematic review of the service composition mechanisms in the cloud environments. Journal of Network and Computer Applications, 81(??):24– 36, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451730005X.

Xin:2017:LBO

[1881] Yu Xin, Zhi-Qiang Xie, and Jing Yang. A load balance oriented cost efficient scheduling method for parallel tasks. Journal of Network and Computer Applications, 81(??):37–46, March 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303496.

Pal:2017:EIN

[1882] Shantanu Pal. Evaluating the impact of network loads and message size on mobile opportunistic networks in challenged environments. Journal of Network and Computer Applications, 81(??):47–58, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303484.

Fortino:2017:EFI

[1883] Giancarlo Fortino, Giuseppe Di Fatta, Sergio F. Ochoa, and Carlos E. Palau. Engineering future interoperable and open IoT systems. Journal of Network and Computer Applications, 81(??):59–61, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300243.

Exarchakos:2017:PAR

[1884] Georgios Exarchakos, Ilker Oztelcan, Dimitris Sarakiotis, and Antonio Liotta. plexi: Adaptive re-scheduling web-service of time synchronized low-power wireless networks. Journal of Network and Computer Applications, 81(??):62–73, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516301424.

Aloi:2017:EII

[1885] G. Aloi, G. Caliciuri, G. Fortino, R. Gravina, P. Pace, W. Russo, and C. Savaglio. Enabling IoT interoperability through opportunistic smartphone-based mobile gateways. Journal of Network and Computer Applications, 81(??):74–84, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516302405.

Varga:2017:MSS

[1886] Pal Varga, Fredrik Blomstedt, Luis Lino Ferreira, Jens Eliasson, Mats Johansson, Jerker Delsing, and Iker Martínez de Soria. Making system of systems interoperable — the core components of the arrowhead framework. Journal of Network and Computer Applications, 81(??):85– 95, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516301965.

Andersen:2017:ESI

[1887] Michael P. Andersen, Gabe Fierro, and David E. Culler. Enabling synergy in IoT: Platform to service and beyond. Journal of Network and Computer Applications, 81(??):96–110, March 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302521.

Ganzha:2017:SII

[1888] Maria Ganzha, Marcin Paprzycki, Wieslaw Pawlowski, Pawel Szmeja, and Katarzyna Wasielewska. Semantic interoperability in the Internet of Things: an overview from the INTER-IoT perspective. Journal of Network and Computer Applications, 81(??):111–124, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804516301618.

Anonymous:2017:EBe

[1889] Anonymous. Editorial board. Journal of Network and Computer Applications, 81(??):ifc, March 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300267.

Asorey-Cacheda:2017:SNL

[1890] R. Asorey-Cacheda, A.-J. Garcia-Sanchez, F. Garcia-Sanchez, and J. Garcia-Haro. A survey on non-linear optimization problems in wireless sensor networks. Journal of Network and Computer Applications, 82(??):1–20, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300024.

Franchino:2017:PAM

[1891] Gianluca Franchino and Giorgio Buttazzo. A power-aware MAC layer protocol for real-time communication in wireless embedded systems. Journal of Network and Computer Applications, 82(??):21–34, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300061.

Zhu:2017:SPS

[1892] Liang Zhu, Changqiao Xu, Jianfeng Guan, and Hongke Zhang. SEM– PPA: a semantical pattern and preference-aware service mining method for personalized point of interest recommendation. Journal of Network and Computer Applications, 82(??):35–46, March 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303502.

Perez-Solano:2017:ITS

[1893] Juan J. Pérez-Solano and Santiago Felici-Castell. Improving time synchronization in Wireless Sensor Networks using Bayesian Inference. Journal of Network and Computer Applications, 82(??):47–55, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300164.

Shen:2017:LWP

[1894] Wenting Shen, Jia Yu, Hui Xia, Hanlin Zhang, Xiuqing Lu, and Rong Hao. Light-weight and privacy-preserving secure cloud auditing scheme for group users via the third party medium. Journal of Network and Computer Applications, 82(??):56–64, March 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300218.

Awad:2017:DNP

[1895] Alaa Awad, Amr Mohamed, Carla-Fabiana Chiasserini, and Tarek Elfouly. Distributed in-network processing and resource optimization over mobile-health systems. Journal of Network and Computer Applications, 82(??):65–76, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730022X.

Kim:2017:CED

[1896] Hyunbum Kim, Heekuck Oh, Paolo Bellavista, and Jalel Ben-Othman. Constructing event-driven partial barriers with resilience in wireless mobile sensor networks. Journal of Network and Computer Applications, 82(??):77–92, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300152.

Amiri:2017:SPM

[1897] Maryam Amiri and Leyli Mohammad-Khanli. Survey on prediction models of applications for resources provisioning in cloud. Journal of Network and Computer Applications, 82(??):93–113, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300231.

Zhan:2017:NKG

[1898] Furui Zhan, Nianmin Yao, Zhenguo Gao, and Guozhen Tan. A novel key generation method for wireless sensor networks based on system of equations. Journal of Network and Computer Applications, 82 (??):114–127, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300395.

Shen:2017:PTP

[1899] Hang Shen, Guangwei Bai, Mei Yang, and Zhonghui Wang. Protecting trajectory privacy: a user-centric analysis. Journal of Network and Computer Applications, 82(??):128–139, March 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300413.

Deka:2017:OPM

[1900] Sanjib K. Deka and Nityananda Sarma. Opportunity prediction at MAClayer sensing for ad-hoc cognitive radio networks. Journal of Network and Computer Applications, 82(??):140–151, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302946.

Arkian:2017:MFB

[1901] Hamid Reza Arkian, Abolfazl Diyanat, and Atefe Pourkhalili. MIST: Fog-based data analytics scheme with cost-efficient resource provisioning for IoT crowdsensing applications. Journal of Network and Computer Applications, 82(??):152–165, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804517300188.

Anonymous:2017:EBf

[1902] Anonymous. Editorial board. Journal of Network and Computer Applications, 82(??):ifc, March 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300681.

Ilkhechi:2017:PFD

[1903] Amir Rahimzadeh Ilkhechi, Ibrahim Korpeoglu, Ugur Güdükbay, and Özgür Ulusoy. PETAL: a fully distributed location service for wireless ad hoc networks. Journal of Network and Computer Applications, 83(??):1–11, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300437.

Ge:2017:FAS

[1904] Mengmeng Ge, Jin B. Hong, Walter Guttmann, and Dong Seong Kim. A framework for automating security analysis of the Internet of Things. Journal of Network and Computer Applications, 83(??):12– 27, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300541.

Congosto:2017:HFP

[1905] Mariluz Congosto, Pablo Basanta-Val, and Luis Sanchez-Fernandez. T-Hoarder: a framework to process Twitter data streams. Journal of Network and Computer Applications, 83(??):28–39, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300486.

Mann:2017:EEC

[1906] Palvinder Singh Mann and Satvir Singh. Energy efficient clustering protocol based on improved metaheuristic in wireless sensor networks. Journal of Network and Computer Applications, 83(??):40–52, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300516.

Ruan:2017:FFS

[1907] Chang Ruan, Jianxin Wang, Wanchun Jiang, Jiawei Huang, Geyong Min, and Yi Pan. FSQCN: Fast and simple quantized congestion notification in data center Ethernet. Journal of Network and Computer Applications, 83(??):53–62, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300450.

Qiu:2017:PSB

[1908] Yue Qiu, Maode Ma, and Xilei Wang. A proxy signature-based handover authentication scheme for LTE wireless networks. Journal of Network and Computer Applications, 83(??):63–71, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300449.

Abbas:2017:OAV

[1909] Nadine Abbas, Hazem Hajj, Zaher Dawy, Karim Jahed, and Sanaa Sharafeddine. An optimized approach to video traffic splitting in heterogeneous wireless networks with energy and QoE considerations. Journal of Network and Computer Applications, 83(??):72–88, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517300140.

Huang:2017:MPA

[1910] Jun Huang, Liqian Xu, Qiang Duan, Cong cong Xing, Jiangtao Luo, and Shui Yu. Modeling and performance analysis for multimedia data flows scheduling in software defined networks. Journal of Network and Computer Applications, 83(??):89–100, April 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730053X.

Park:2017:CEU

[1911] Jonghoon Park, Cheoleun Moon, Ikjun Yeom, and Yusung Kim. Cardinality estimation using collective interference for large-scale RFID systems. Journal of Network and Computer Applications, 83(??):101– 110, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300565.

Chouikhi:2017:CCR

[1912] Samira Chouikhi, Inès El Korbi, Yacine Ghamri-Doudane, and Leila Azouz Saidane. Centralized connectivity restoration in multichannel wireless sensor networks. Journal of Network and Computer Applications, 83(??):111–123, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300619.

Ali:2017:TCD

[1913] Anum Ali, Ghalib A. Shah, Muhammad Omer Farooq, and Usman Ghani. Technologies and challenges in developing machine-to-machine applications: a survey. Journal of Network and Computer Applications, 83(??):124–139, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300620.

Gogu:2017:UDP

[1914] Ada Gogu, Dritan Nace, Enrico Natalizio, and Yacine Challal. Using dynamic programming to solve the Wireless Sensor Network Configuration Problem. Journal of Network and Computer Applications, 83(??):140–154, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300425.

Toosi:2017:RAG

[1915] Adel Nadjaran Toosi, Chenhao Qu, Marcos Dias de Assunção, and Rajkumar Buyya. Renewable-aware geographical load balancing of web applications for sustainable data centers. Journal of Network and Computer Applications, 83(??):155–168, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804517300590.

Chkirbene:2017:LLC

[1916] Zina Chkirbene, Sebti Foufou, Ridha Hamila, Zahir Tari, and Albert Y. Zomaya. LaCoDa: Layered connected topology for massive data centers. Journal of Network and Computer Applications, 83(??):169–180, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300401.

Abdaoui:2017:ITS

[1917] Abderrazak Abdaoui, Tarek M. El Fouly, and Mohamed H. Ahmed. Impact of time synchronization error on the mode-shape identification and damage detection/localization in WSNs for structural health monitoring. Journal of Network and Computer Applications, 83(??):181– 189, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300012.

Leyva-Mayorga:2017:HMQ

[1918] Israel Leyva-Mayorga, Vicent Pla, Jorge Martinez-Bauset, and Mario E. Rivero-Angeles. A hybrid method for the QoS analysis and parameter optimization in time-critical random access wireless sensor networks. Journal of Network and Computer Applications, 83(??):190– 203, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300504.

Sun:2017:DTD

[1919] Yunchuan Sun, Xinpei Yu, Rongfang Bie, and Houbing Song. Discovering time-dependent shortest path on traffic graph for drivers towards green driving. Journal of Network and Computer Applications, 83(??):204–212, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516000497.

Liu:2017:DHS

[1920] Yu Liu, Changjie Zhang, Bo Li, and Jianwei Niu. DeMS: a hybrid scheme of task scheduling and load balancing in computing clusters. Journal of Network and Computer Applications, 83(??):213–220, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804515000922.

Anonymous:2017:EBg

[1921] Anonymous. Editorial board. Journal of Network and Computer Applications, 83(??):ifc, April 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300875.

Garg:2017:RMR

[1922] Neenu Garg and Seema Bawa. RITS–MHT: Relative indexed and time stamped Merkle hash tree based data auditing protocol for cloud computing. Journal of Network and Computer Applications, 84(??):1–13, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300668.

Nunes:2017:BAU

[1923] Ingrid Nunes, Frederico Schardong, and Alberto Schaeffer-Filho. BDI2DoS: an application using collaborating BDI agents to combat DDoS attacks. Journal of Network and Computer Applications, 84(??): 14–24, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300577.

Zarpelao:2017:SID

[1924] Bruno Bogaz Zarpelão, Rodrigo Sanches Miani, Cláudio Toshio Kawakani, and Sean Carlisto de Alvarenga. A survey of intrusion detection in Internet of Things. Journal of Network and Computer Applications, 84(??):25–37, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300802.

Mollah:2017:SPC

[1925] Muhammad Baqer Mollah, Md. Abul Kalam Azad, and Athanasios Vasilakos. Security and privacy challenges in mobile cloud computing: Survey and way ahead. Journal of Network and Computer Applications, 84(??):38–54, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300632.

Solmaz:2017:TPE

[1926] Gürkan Solmaz and Damla Turgut. Tracking pedestrians and emergent events in disaster areas. Journal of Network and Computer Applications, 84(??):55–67, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300784.

Rahman:2017:CBS

[1927] Mahfuzur Rahman and Peter Graham. Compatibility-based static VM placement minimizing interference. Journal of Network and Computer Applications, 84(??):68–81, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804517300656.

Pourpeighambar:2017:MAL

[1928] Babak Pourpeighambar, Mehdi Dehghan, and Masoud Sabaei. Multiagent learning based routing for delay minimization in Cognitive Radio Networks. Journal of Network and Computer Applications, 84(??):82– 92, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300474.

Jawhar:2017:CNU

[1929] Imad Jawhar, Nader Mohamed, Jameela Al-Jaroodi, Dharma P. Agrawal, and Sheng Zhang. Communication and networking of UAVbased systems: Classification and associated architectures. Journal of Network and Computer Applications, 84(??):93–108, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517300814.

Alghamdi:2017:TTP

[1930] Mohammed I. Alghamdi, Xunfei Jiang, Ji Zhang, Jifu Zhang, Minghua Jiang, and Xiao Qin. Towards two-phase scheduling of real-time applications in distributed systems. Journal of Network and Computer Applications, 84(??):109–117, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300607.

Guo:2017:GFD

[1931] Binglei Guo, Jiong Yu, Bin Liao, Dexian Yang, and Liang Lu. A green framework for DBMS based on energy-aware query optimization and energy-efficient query processing. Journal of Network and Computer Applications, 84(??):118–130, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517300851.

Anonymous:2017:EBh

[1932] Anonymous. Editorial board. Journal of Network and Computer Applications, 84(??):ifc, April 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301017.

Tsai:2017:ISH

[1933] Chun-Wei Tsai, Priyanka Rawat, and Ming-Chao Chiang. Intelligent systems for heterogeneous networks. Journal of Network and Computer Applications, 85(??):1–3, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451730125X.

Roh:2017:JFV

[1934] Heejun Roh, Cheoulhoon Jung, Kyunghwi Kim, Sangheon Pack, and Wonjun Lee. Joint flow and virtual machine placement in hybrid cloud data centers. Journal of Network and Computer Applications, 85(??):

4–13, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303101.

Spinnewyn:2017:RAP

[1935] Bart Spinnewyn, Ruben Mennes, Juan Felipe Botero, and Steven Latré. Resilient application placement for geo-distributed cloud networks. Journal of Network and Computer Applications, 85(??):14– 31, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303149.

Kirthica:2017:CCI

[1936] S. Kirthica and Rajeswari Sridhar. CIT: a Cloud Inter-operation Toolkit to enhance elasticity and tolerate shut down of external clouds. Journal of Network and Computer Applications, 85(??):32– 46, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303137.

Hou:2017:BSB

[1937] Aiqin Hou, Chase Q. Wu, Dingyi Fang, Yongqiang Wang, and Meng Wang. Bandwidth scheduling for big data transfer using multiple fixed node-disjoint paths. Journal of Network and Computer Applications, 85(??):47–55, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303022.

Guan:2017:GBC

[1938] Jianfeng Guan, Zhiwei Yan, Su Yao, Changqiao Xu, and Hongke Zhang. GBC-based caching function group selection algorithm for SINET. Journal of Network and Computer Applications, 85(??):56– 63, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303071.

Han:2017:MAN

[1939] Guangjie Han, Chenyu Zhang, Jinfang Jiang, Xuan Yang, and Mohsen Guizani. Mobile anchor nodes path planning algorithms using networkdensity-based clustering in wireless sensor networks. Journal of Network and Computer Applications, 85(??):64–75, May 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303046.

Wang:2017:SDN

[1940] Junfeng Wang, Yiming Miao, Ping Zhou, M. Shamim Hossain, and Sk Md Mizanur Rahman. A software defined network routing in wireless multihop network. Journal of Network and Computer Applications, 85(??):76–83, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303058.

Soliman:2017:FGC

[1941] Samir S. Soliman and Bongyong Song. Fifth generation (5G) cellular and the network for tomorrow: cognitive and cooperative approach for energy savings. Journal of Network and Computer Applications, 85(??):84–93, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303125.

Sharma:2017:IDU

[1942] Vishal Sharma, Kathiravan Srinivasan, Han-Chieh Chao, Kai-Lung Hua, and Wen-Huang Cheng. Intelligent deployment of UAVs in 5G heterogeneous communication environment for improved coverage. Journal of Network and Computer Applications, 85(??):94–105, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516303034.

Tseng:2017:FGB

[1943] Chih-Cheng Tseng, Hwang-Cheng Wang, Kuo-Chang Ting, Chih-Chieh Wang, and Fang-Chang Kuo. Fast game-based handoff mechanism with load balancing for LTE/LTE-A heterogeneous networks. Journal of Network and Computer Applications, 85(??):106–115, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303113.

Chen:2017:CAG

[1944] Chen Chen, Tie Qiu, Jinna Hu, Zhiyuan Ren, Yang Zhou, and Arun Kumar Sangaiah. A congestion avoidance game for information exchange on intersections in heterogeneous vehicular networks. Journal of Network and Computer Applications, 85(??):116–126, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303095.

Ko:2017:DDL

[1945] Haneul Ko, Sangheon Pack, Jong-Hyouk Lee, and Alexandru Petrescu. DLM: Delayed Location Management in network mobility (NEMO)- based public transportation systems. Journal of Network and Computer Applications, 85(??):127–133, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451630306X.

Choi:2017:FIA

[1946] Hyun-Ho Choi and Jung-Ryun Lee. A flocking-inspired algorithm for fair resource allocation in vehicle-mounted mobile relays. Journal of Network and Computer Applications, 85(??):134–142, May 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516303162.

Anonymous:2017:EBi

[1947] Anonymous. Editorial board. Journal of Network and Computer Applications, 85(??):ifc, May 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301194.

Yan:2017:SIP

[1948] Zheng Yan, Jun Liu, and Laurence T. Yang. Special issue on pervasive social networking. Journal of Network and Computer Applications, 86(??):1–2, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730139X.

Liu:2017:ESS

[1949] Chang Liu, Liehuang Zhu, and Jinjun Chen. Efficient searchable symmetric encryption for storing multiple source dynamic social data on cloud. Journal of Network and Computer Applications, 86(??):3–14, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451630217X.

Lu:2017:SIM

[1950] Feng Lu, Weikang Zhang, Liwen Shao, Xunfei Jiang, Peng Xu, and Hai Jin. Scalable influence maximization under independent cascade model. Journal of Network and Computer Applications, 86(??):15–23, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302582.

Quick:2017:PSN

[1951] Darren Quick and Kim-Kwang Raymond Choo. Pervasive social networking forensics: Intelligence and evidence from mobile device extracts. Journal of Network and Computer Applications, 86(??):24–33, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302843.

Sun:2017:UDP

[1952] Gang Sun, Yuxia Xie, Dan Liao, Hongfang Yu, and Victor Chang. User-defined privacy location-sharing system in mobile online social networks. Journal of Network and Computer Applications, 86(??):34–45, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302934.

Yan:2017:CCD

[1953] Zheng Yan and Wangyang Shi. CloudFile: a cloud data access control system based on mobile social trust. Journal of Network and Computer Applications, 86(??):46–58, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804516302879.

Yang:2017:HPS

[1954] Wenyin Yang, Guojun Wang, Md Zakirul Alam Bhuiyan, and Kim-Kwang Raymond Choo. Hypergraph partitioning for social networks based on information entropy modularity. Journal of Network and Computer Applications, 86(??):59–71, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804516302326.

Zhang:2017:SBS

[1955] Yue Zhang, Bin Song, and Peng Zhang. Social behavior study under pervasive social networking based on decentralized deep reinforcement learning. Journal of Network and Computer Applications, 86(??):72– 81, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302867.

Zhang:2017:PDD

[1956] Zufan Zhang, Lisha Wang, Dan Liu, and Yu Zhang. Peer discovery for D2D communications based on social attribute and service attribute. Journal of Network and Computer Applications, 86(??):82– 91, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302715.

Zhou:2017:AMD

[1957] Yadong Zhou, Beibei Zhang, Xiaoxiao Sun, Qinghua Zheng, and Ting Liu. Analyzing and modeling dynamics of information diffusion in microblogging social network. Journal of Network and Computer Applications, 86(??):92–102, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516302168.

Anonymous:2017:EBj

[1958] Anonymous. Editorial board. Journal of Network and Computer Applications, 86(??):ifc, May 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301339.

Masdari:2017:STA

[1959] Mohammad Masdari and Safiyyeh Ahmadzadeh. A survey and taxonomy of the authentication schemes in Telecare Medicine Information Systems. Journal of Network and Computer Applications, 87(??):1– 19, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300978.

Xiong:2017:RDN

[1960] Bing Xiong, Kun Yang, Jinyuan Zhao, and Keqin Li. Robust dynamic network traffic partitioning against malicious attacks. Journal of Network and Computer Applications, 87(??):20–31, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516300637.

Wibowo:2017:MDS

[1961] Franciscus X. A. Wibowo, Mark A. Gregory, Khandakar Ahmed, and Karina M. Gomez. Multi-domain Software Defined Networking: Research status and challenges. Journal of Network and Computer Applications, 87(??):32–45, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300991.

Song:2017:PFG

[1962] Ping Song, Yi Liu, Chi Liu, and Depei Qian. ParaFlow: Fine-grained parallel SDN controller for large-scale networks. Journal of Network and Computer Applications, 87(??):46–59, June 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730111X.

Xiong:2017:FTF

[1963] Sisi Xiong, Yanjun Yao, Michael Berry, Hairong Qi, and Qing Cao. Frequent traffic flow identification through probabilistic Bloom filter and its GPU-based acceleration. Journal of Network and Computer Applications, 87(??):60–72, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730108X.

Antunes:2017:NMB

[1964] Rodolfo S. Antunes, Matheus B. Lehmann, Rodrigo B. Mansilha, Luciano P. Gaspary, and Marinho P. Barcellos. NDNrel: a mechanism based on relations among objects to improve the performance of NDN. Journal of Network and Computer Applications, 87(??):73– 86, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301170.

Ducourthial:2017:IFB

[1965] B. Ducourthial, S. Mottelet, and A. Busson. Improving fairness between close Wi-Fi access points. Journal of Network and Computer Applications, 87(??):87–99, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300966.

Li:2017:RTS

[1966] Chunlin Li, Jing Zhang, and Youlong Luo. Real-time scheduling based on optimized topology and communication traffic in distributed real-time computation platform of storm. Journal of Network and Computer Applications, 87(??):100–115, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517301091.

Majeed:2017:PCP

[1967] Muhammad Faran Majeed, Matthew N. Dailey, Riaz Khan, and Apinun Tunpan. Pre-caching: a proactive scheme for caching video traffic in named data mesh networks. Journal of Network and Computer Applications, 87(??):116–130, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301145.

Moghaddam:2017:MCA

[1968] Reza Farrahi Moghaddam, Vahid Asghari, Fereydoun Farrahi Moghaddam, Yves Lemieux, and Mohamed Cheriet. A Monte-Carlo approach to lifespan failure performance analysis of the network fabric in modular data centers. Journal of Network and Computer Applications, 87(??):131–146, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301169.

Kargar:2017:FAM

[1969] Saeed Kargar and Leyli Mohammad-Khanli. Fractal: an advanced multidimensional range query lookup protocol on nested rings for distributed systems. Journal of Network and Computer Applications, 87(??):147–168, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301303.

Roy:2017:SBE

[1970] Animesh Roy, Soumit Bose, Tamaghna Acharya, and Sipra Das-Bit. Social-based energy-aware multicasting in delay tolerant networks. Journal of Network and Computer Applications, 87(??):169– 184, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301297.

Haider:2017:GRI

[1971] W. Haider, J. Hu, J. Slay, B. P. Turnbull, and Y. Xie. Generating realistic intrusion detection system dataset based on fuzzy qualitative modeling. Journal of Network and Computer Applications, 87(??):185– 192, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301273.

Nunez:2017:PRE

[1972] David Nuñez, Isaac Agudo, and Javier Lopez. Proxy re-encryption: Analysis of constructions and its application to secure access delegation. Journal of Network and Computer Applications, 87(??):193– 209, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301078.

Liu:2017:MMD

[1973] Peng Liu, Biao Xu, Guojun Dai, Zhen Jiang, and Jie Wu. MDP: Minimum delay hot-spot parking. Journal of Network and Computer Applications, 87(??):210–222, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730098X.

Arun:2017:DPA

[1974] I. B. Arun and T. G. Venkatesh. Design and performance analysis of a MAC protocol for wireless LANs supporting multipacket reception. Journal of Network and Computer Applications, 87(??):223– 236, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301121.

Atiquzzaman:2017:ENE

[1975] Mohammed Atiquzzaman. Editorial: New Editors 2017. Journal of Network and Computer Applications, 87(??):A1–A5, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517301510.

Anonymous:2017:EBk

[1976] Anonymous. Editorial board. Journal of Network and Computer Applications, 87(??):ifc, June 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301534.

Zhang:2017:NUC

[1977] De gan Zhang, Si Liu, Ting Zhang, and Zhao Liang. Novel unequal clustering routing protocol considering energy balancing based on network partition & distance for mobile education. Journal of Network and Computer Applications, 88(??):1–9, June 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517301418.

Alaba:2017:ITS

[1978] Fadele Ayotunde Alaba, Mazliza Othman, Ibrahim Abaker Targio Hashem, and Faiz Alotaibi. Internet of Things security: a survey. Journal of Network and Computer Applications, 88(??):10–28, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301455.

Alipio:2017:CBT

[1979] Melchizedek Alipio, Nestor Michael Tiglao, Antonio Grilo, Fawaz Bokhari, Umair Chaudhry, and Shavez Qureshi. Cache-based transport protocols in wireless sensor networks: a survey and future directions. Journal of Network and Computer Applications, 88(??):29– 49, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301443.

Ghomi:2017:LBA

[1980] Einollah Jafarnejad Ghomi, Amir Masoud Rahmani, and Nooruldeen Nasih Qader. Load-balancing algorithms in cloud computing: a survey. Journal of Network and Computer Applications, 88(??):50–71, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517301480.

Saleh:2017:MAQ

[1981] Ahmed I. Saleh, Khaled M. Abo-Al-Ez, and Ahmed A. Abdullah. A Multi-Aware Query Driven (MAQD) routing protocol for mobile wireless sensor networks based on neuro-fuzzy inference. Journal of Network and Computer Applications, 88(??):72–98, June 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300954.

Khan:2017:LBG

[1982] Sumair Khan, Babar Nazir, Iftikhar Ahmed Khan, Shahaboddin Shamshirband, and Anthony T. Chronopoulos. Load balancing in grid computing: Taxonomy, trends and opportunities. Journal of Network and Computer Applications, 88(??):99–111, June 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300838.

Kim:2017:IPC

[1983] Eunsam Kim and Jonathan C. L. Liu. An integrated prefetching/caching scheme in multimedia servers. Journal of Network and Computer Applications, 88(??):112–123, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517300826.

Li:2017:EME

[1984] Ruiying Li, Meinan Li, Haitao Liao, and Ning Huang. An efficient method for evaluating the end-to-end transmission time reliability of a switched Ethernet. Journal of Network and Computer Applications, 88(??):124–133, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300589.

Abreu:2017:SCA

[1985] Carlos Abreu, Francisco Miranda, and P. M. Mendes. Smart contextaware QoS-based admission control for biomedical wireless sensor networks. Journal of Network and Computer Applications, 88(??):134– 145, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300553.

Anonymous:2017:EBl

[1986] Anonymous. Editorial board. Journal of Network and Computer Applications, 88(??):ifc, June 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301686.

Chang:2017:ESI

[1987] Victor Chang, Verena Kantere, and Muthu Ramanchadran. Emerging services for Internet of Things. Journal of Network and Computer Applications, 89(??):1–2, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301467.

Sun:2017:ELP

[1988] Gang Sun, Victor Chang, Muthu Ramachandran, Zhili Sun, Gangmin Li, Hongfang Yu, and Dan Liao. Efficient location privacy algorithm for Internet of Things (IoT) services and applications. Journal of Network and Computer Applications, 89(??):3–13, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804516302429.

Qureshi:2017:EEW

[1989] Faisal Fayyaz Qureshi, Rahat Iqbal, and Muhammad Nabeel Asghar. Energy efficient wireless communication technique based on Cognitive Radio for Internet of Things. Journal of Network and Computer Applications, 89(??):14–25, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300048.

Yang:2017:LDS

[1990] Yang Yang, Xianghan Zheng, and Chunming Tang. Lightweight distributed secure data management system for health Internet of Things. Journal of Network and Computer Applications, 89(??):26– 37, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804516302855.

Yassein:2017:NET

[1991] Muneer Bani Yassein, Shadi Aljawarneh, and Esra’a Masa’deh. A new elastic trickle timer algorithm for Internet of Things. Journal of Network and Computer Applications, 89(??):38–47, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300462.

Chen:2017:EPS

[1992] Chen Chen, Honghui Zhao, Tie Qiu, Mingcheng Hu, Hui Han, and Zhiyuan Ren. An efficient power saving polling scheme in the Internet of Energy. Journal of Network and Computer Applications, 89(??): 48–61, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300036.

Kumar:2017:BSB

[1993] Pradeep Kumar, Rajkumar Saini, Partha Pratim Roy, and Debi Prosad Dogra. A bio-signal based framework to secure mobile devices. Journal of Network and Computer Applications, 89(??):62–71, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517300772.

Wu:2017:EAK

[1994] Fan Wu, Lili Xu, Saru Kumari, Xiong Li, Jian Shen, Kim-Kwang Raymond Choo, Mohammad Wazid, and Ashok Kumar Das. An efficient authentication and key agreement scheme for multi-gateway wireless sensor networks in IoT deployment. Journal of Network and Computer Applications, 89(??):72–85, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804516303150.

Zhang:2017:IRT

[1995] Cong Zhang, Chin-Feng Lai, Ying-Hsun Lai, Zhen-Wei Wu, and Han-Chieh Chao. An inferential real-time falling posture reconstruction for Internet of Healthcare Things. Journal of Network and Computer Applications, 89(??):86–95, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300760.

Baker:2017:EAS

[1996] Thar Baker, Muhammad Asim, Hissam Tawfik, Bandar Aldawsari, and Rajkumar Buyya. An energy-aware service composition algorithm for multiple cloud-based IoT applications. Journal of Network and Computer Applications, 89(??):96–108, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517301108.

Alam:2017:AVR

[1997] Md Fasiul Alam, Serafeim Katsikas, Olga Beltramello, and Stathes Hadjiefthymiades. Augmented and virtual reality based monitoring and safety system: a prototype IoT platform. Journal of Network and Computer Applications, 89(??):109–119, July 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517301315.

Nguyen:2017:ONF

[1998] Xuan Thang Nguyen, Huu Tam Tran, Harun Baraki, and Kurt Geihs. Optimization of non-functional properties in Internet of Things applications. Journal of Network and Computer Applications, 89(??):120– 129, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301285.

Anonymous:2017:EBm

[1999] Anonymous. Editorial board. Journal of Network and Computer Applications, 89(??):ifc, July 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301753.

Machado:2017:AAR

[2000] Cristian Cleder Machado, Juliano Araujo Wickboldt, Lisandro Zambenedetti Granville, and Alberto Schaeffer-Filho. ARKHAM: an advanced refinement toolkit for handling service level agreements in software-defined networking. Journal of Network and Computer Applications, 90(??):1–16, July 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301509.

Dehwah:2017:UWE

[2001] Ahmad H. Dehwah, Shahrazed Elmetennani, and Christian Claudel. UD-WCMA: an energy estimation and forecast scheme for solar powered wireless sensor networks. Journal of Network and Computer Applications, 90(??):17–25, July 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730142X.

Vieira:2017:MOS

[2002] Thiago P. B. Vieira, Danilo F. Tenório, João Paulo C. L. da Costa, Edison P. de Freitas, Giovanni Del Galdo, and Rafael T. de Sousa Júnior. Model order selection and eigen similarity based framework for detection and identification of network attacks. Journal of Network and Computer Applications, 90(??):26–41, July 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517301601.

Zarrin:2017:HHA

[2003] Javad Zarrin, Rui L. Aguiar, and João Paulo Barraca. HARD: Hybrid adaptive resource discovery for jungle computing. Journal of Network and Computer Applications, 90(??):42–73, July 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517301625.

Mun:2017:CSU

[2004] Ju Hyoung Mun and Hyesook Lim. Cache sharing using bloom filters in named data networking. Journal of Network and Computer Applications, 90(??):74–82, July 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301595.

Wu:2017:LMI

[2005] Tin-Yu Wu, Nadra Guizani, and Jhih-Siang Huang. Live migration improvements by related dirty memory prediction in cloud computing. Journal of Network and Computer Applications, 90(??):83–89, July 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301133.

Anonymous:2017:EBn

[2006] Anonymous. Editorial board. Journal of Network and Computer Applications, 90(??):ifc, July 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730187X.

Adi:2017:SDS

[2007] Erwin Adi, Zubair Baig, and Philip Hingston. Stealthy Denial of Service (DoS) attack modelling and detection for HTTP/2 services. Journal of Network and Computer Applications, 91(??):1–13, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301637.

Phan:2017:DSN

[2008] Trung V. Phan, Nguyen Khac Bao, and Minho Park. Distributed-SOM: a novel performance bottleneck handler for large-sized software-defined networks under flooding attacks. Journal of Network and Computer Applications, 91(??):14–25, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517301649.

Shen:2017:NCN

[2009] Shigen Shen, Haiping Ma, En Fan, Keli Hu, Shui Yu, Jianhua Liu, and Qiying Cao. A non-cooperative non-zero-sum game-based dependability assessment of heterogeneous WSNs with malware diffusion. Journal of Network and Computer Applications, 91(??):26–35, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301832.

Masdari:2017:KMW

[2010] Mohammad Masdari, Safiyyeh Ahmadzadeh, and Moazam Bidaki. Key management in wireless Body Area Network: Challenges and issues. Journal of Network and Computer Applications, 91(??):36–51, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301492.

Kumar:2017:TAU

[2011] Vireshwar Kumar, Jung-Min (Jerry) Park, and Kaigui Bian. Transmitter authentication using hierarchical modulation in dynamic spectrum sharing. Journal of Network and Computer Applications, 91(??):52–60, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301935.

Ghazouani:2017:SCS

[2012] Souad Ghazouani and Yahya Slimani. A survey on cloud service description. Journal of Network and Computer Applications, 91(??):61– 74, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301613.

Zhang:2017:CCR

[2013] Jing Zhang, Li Xu, Shuming Zhou, Geyong Min, Yang Xiang, and Jia Hu. Crossed Cube Ring: a k-connected virtual backbone for wireless sensor networks. Journal of Network and Computer Applications, 91(??):75–88, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301650.

Qirtas:2017:TDT

[2014] Malik Muhammad Qirtas, Yasir Faheem, and Mubashir Husain Rehmani. Throwboxes in delay tolerant networks: a survey of placement strategies, buffering capacity, and mobility models. Journal of Network and Computer Applications, 91(??):89–103, August 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517301431.

Michalas:2017:IMF

[2015] Angelos Michalas, Aggeliki Sgora, and Dimitrios D. Vergados. An integrated MIH-FPMIPv6 mobility management approach for evolvedpacket system architectures. Journal of Network and Computer Applications, 91(??):104–119, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517300644.

Anonymous:2017:EBo

[2016] Anonymous. Editorial board. Journal of Network and Computer Applications, 91(??):ifc, August 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301972.

Thampi:2017:SIU

[2017] Sabu M. Thampi, Jaime Lloret Mauri, and Robin Doss. Special issue on Underwater Acoustic Sensor Networks: Emerging trends and current perspectives. Journal of Network and Computer Applications, 92(??):1–2, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301479.

Dhongdi:2017:CLP

[2018] Sarang C. Dhongdi, Pritish Nahar, Rishabh Sethunathan, Lucy J. Gudino, and K. R. Anupama. Cross-layer protocol stack development for three-dimensional underwater Acoustic Sensor Network. Journal of Network and Computer Applications, 92(??):3–19, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517301406.

Mandal:2017:PCT

[2019] Priyatosh Mandal. Performance comparison of TDMA variants for underwater acoustic sensor networks. Journal of Network and Computer Applications, 92(??):20–30, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804517300528.

Javaid:2017:RBC

[2020] Nadeem Javaid, Sheraz Hussain, Ashfaq Ahmad, Muhammad Imran, Abid Khan, and Mohsen Guizani. Region based cooperative routing in underwater wireless sensor networks. Journal of Network and Computer Applications, 92(??):31–41, August 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300206.

Rani:2017:EEC

[2021] Shalli Rani, Syed Hassan Ahmed, Jyoteesh Malhotra, and Rajneesh Talwar. Energy efficient chain based routing protocol for underwater wireless sensor networks. Journal of Network and Computer Applications, 92(??):42–50, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517300176.

Han:2017:AAL

[2022] Guangjie Han, Li Liu, Na Bao, Jinfang Jiang, Wenbo Zhang, and Joel J. P. C. Rodrigues. AREP: an asymmetric link-based reverse routing protocol for underwater acoustic sensor networks. Journal of Network and Computer Applications, 92(??):51–58, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730019X.

Wang:2017:DOA

[2023] Jingjing Wang, Wei Shi, Lingwei Xu, Liya Zhou, Qiuna Niu, and Ju liu. Design of optical-acoustic hybrid underwater wireless sensor network. Journal of Network and Computer Applications, 92(??):59–67, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451730084X.

Sun:2017:ECS

[2024] Weifeng Sun, Chi Zhou, Tie Qiu, and Yuqing Liu. An efficient communication scheme for solving merge conflicts in maritime transportation. Journal of Network and Computer Applications, 92(??):68–76, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517300498.

Anonymous:2017:EBp

[2025] Anonymous. Editorial board. Journal of Network and Computer Applications, 92(??):ifc, August 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302138.

Boushaba:2017:NSB

[2026] Mustapha Boushaba, Abdelhakim Hafid, and Michel Gendreau. Node stability-based routing in Wireless Mesh Networks. Journal of Network and Computer Applications, 93(??):1–12, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517300796.

Sunny:2017:GCM

[2027] Albert Sunny, Sumankumar Panchal, Nikhil Vidhani, Subhashini Krishnasamy, S. V. R. Anand, Malati Hegde, Joy Kuri, and Anurag Kumar. A generic controller for managing TCP transfers in IEEE 802.11 infrastructure WLANs. Journal of Network and Computer Applications, 93(??):13–26, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301947.

Huang:2017:DAD

[2028] Chi-Sheng Huang, Meng-Feng Tsai, Po-Hsuan Huang, Li-Ding Su, and Kuei-Sheng Lee. Distributed asteroid discovery system for large astronomical data. Journal of Network and Computer Applications, 93 (??):27–37, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301157.

Lalem:2017:LLP

[2029] Farid Lalem, Ahcène Bounceur, Madani Bezoui, Massinissa Saoudi, Reinhardt Euler, Tahar Kechadi, and Marc Sevaux. LPCN: Least polarangle connected node algorithm to find a polygon hull in a connected Euclidean graph. Journal of Network and Computer Applications, 93 (??):38–50, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301856.

Ahmad:2017:RMD

[2030] M. Ahmad, M. Azam, M. Naeem, M. Iqbal, A. Anpalagan, and M. Haneef. Resource management in D2D communication: an optimization perspective. Journal of Network and Computer Applications, 93 (??):51–75, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301261.

Ma:2017:DCR

[2031] Chaofan Ma, Wei Liang, and Meng Zheng. Delay constrained relay node placement in two-tiered wireless sensor networks: a set-covering-based algorithm. Journal of Network and Computer Applications, 93 (??):76–90, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517301844.

A:2017:CAA

[2032] Jeyasekar A., Kasmir Raja S. V., and Annie Uthra R. Congestion avoidance algorithm using ARIMA(2,1,1) model-based RTT estimation and RSS in heterogeneous wired-wireless networks. Journal of Network and Computer Applications, 93(??):91–109, September 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302060.

Luo:2017:PTM

[2033] Chengming Luo, Wei Li, Xinnan Fan, Hai Yang, Jianjun Ni, Xuewu Zhang, Gaifang Xin, and Pengfei Shi. Positioning technology of mobile vehicle using self-repairing heterogeneous sensor networks. Journal of Network and Computer Applications, 93(??):110–122, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302096.

Bu:2017:EAR

[2034] Chao Bu, Xingwei Wang, Hui Cheng, Min Huang, Keqin Li, and Sajal K. Das. Enabling Adaptive Routing Service Customization via the integration of SDN and NFV. Journal of Network and Computer Applications, 93(??):123–136, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302084.

Lee:2017:ARM

[2035] Sugi Lee, Younghoon Kim, Ikjun Yeom, and Yusung Kim. Active request management in stateful forwarding networks. Journal of Network and Computer Applications, 93(??):137–149, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302072.

Sharif:2017:USN

[2036] Mohammad Sharif and Abolghasem Sadeghi-Niaraki. Ubiquitous sensor network simulation and emulation environments: a survey. Journal of Network and Computer Applications, 93(??):150–181, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302059.

Maenhaut:2017:DTD

[2037] Pieter-Jan Maenhaut, Hendrik Moens, Bruno Volckaert, Veerle Ongenae, and Filip De Turck. A dynamic Tenant-Defined Storage system for efficient resource management in cloud applications. Journal of Network and Computer Applications, 93(??):182–196, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302114.

Yi:2017:DES

[2038] Bo Yi, Xingwei Wang, and Min Huang. Design and evaluation of schemes for provisioning service function chain with function scalability. Journal of Network and Computer Applications, 93(??):197–214, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302102.

Fan:2017:ENA

[2039] Yao-Chung Fan, Huan Chen, Fang-Yie Leu, and Ilsun You. Enabling in-network aggregation by diffusion units for urban scale M2M networks. Journal of Network and Computer Applications, 93(??):215–227, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517301662.

Kamimura:2017:SON

[2040] Akiya Kamimura and Kohji Tomita. A self-organizing network coordination framework enabling collision-free and congestion-less wireless sensor networks. Journal of Network and Computer Applications, 93 (??):228–244, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302217.

Khan:2017:WFD

[2041] Muhammad Asif Khan, Wael Cherif, Fethi Filali, and Ridha Hamila. Wi-Fi direct research — current status and future perspectives. Journal of Network and Computer Applications, 93(??):245–258, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302230.

Anwar:2017:CVC

[2042] Shahid Anwar, Zakira Inayat, Mohamad Fadli Zolkipli, Jasni Mohamad Zain, Abdullah Gani, Nor Badrul Anuar, Muhammad Khurram Khan, and Victor Chang. Cross-VM cache-based side channel attacks and proposed prevention mechanisms: a survey. Journal of Network and Computer Applications, 93(??):259–279, September 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302205.

Anonymous:2017:EBq

[2043] Anonymous. Editorial board. Journal of Network and Computer Applications, 93(??):ifc, September 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302369.

Ghebleh:2017:RCS

[2044] Reza Ghebleh and Ali Ghaffari. Retracted: A Comprehensive survey on 5G: the prospect of future communications. Journal of Network and Computer Applications, 94(??):1–32, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S1084804517302291. See retraction notice [2339].

Dai:2017:ENI

[2045] Bin Dai, Guan Xu, Bengxiong Huang, Peng Qin, and Yang Xu. Enabling network innovation in data center networks with software defined networking: a survey. Journal of Network and Computer Applications, 94(??):33–49, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302278.

Zafar:2017:TDS

[2046] Faheem Zafar, Abid Khan, Saba Suhail, Idrees Ahmed, Khizar Hameed, Hayat Mohammad Khan, Farhana Jabeen, and Adeel Anjum. Trustworthy data: a survey, taxonomy and future trends of secure provenance schemes. Journal of Network and Computer Applications, 94 (??):50–68, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302229.

Nath:2017:DSC

[2047] Arijit Nath and Nityananda Sarma. A distributed solution for cooperative spectrum sensing scheduling in multi-band cognitive radio networks. Journal of Network and Computer Applications, 94(??):69–77, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302242.

Petrangeli:2017:MLB

[2048] Stefano Petrangeli, Tingyao Wu, Tim Wauters, Rafael Huysegems, Tom Bostoen, and Filip De Turck. A machine learning-based framework for preventing video freezes in HTTP adaptive streaming. Journal of Network and Computer Applications, 94(??):78–92, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451730231X.

Junior:2017:SMA

[2049] Warley Junior, Adriano França, Kelvin Dias, and José N. de Souza. Supporting mobility-aware computational offloading in mobile cloud environment. Journal of Network and Computer Applications, 94(??): 93–108, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302321.

Anonymous:2017:EBr

[2050] Anonymous. Editorial board. Journal of Network and Computer Applications, 94(??):ifc, September 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302527.

Rehan:2017:CSM

[2051] Waqas Rehan, Stefan Fischer, Maaz Rehan, and Mubashir Husain Rehmani. A comprehensive survey on multichannel routing in wireless sensor networks. Journal of Network and Computer Applications, 95(??):1–25, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730228X.

Aslanpour:2017:ASW

[2052] Mohammad Sadegh Aslanpour, Mostafa Ghobaei-Arani, and Adel Nadjaran Toosi. Auto-scaling web applications in clouds: a cost-aware approach. Journal of Network and Computer Applications, 95(??):26–41, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302448.

Marotta:2017:FRO

[2053] Antonio Marotta, Enrica Zola, Fabio D’Andreagiovanni, and Andreas Kassler. A fast robust optimization-based heuristic for the deployment of green virtual network functions. Journal of Network and Computer Applications, 95(??):42–53, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517302461.

Chopra:2017:SUD

[2054] Garima Chopra, Rakesh Kumar Jha, and Sanjeev Jain. A survey on ultra-dense network and emerging technologies: Security challenges and possible solutions. Journal of Network and Computer Applications, 95(??):54–78, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302308.

Heilig:2017:LAB

[2055] Leonard Heilig, Rajkumar Buyya, and Stefan Voß. Location-aware brokering for consumers in multi-cloud computing environments. Journal of Network and Computer Applications, 95(??):79–93, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517302333.

Yokoyama:2017:AAS

[2056] Daniel Yokoyama, Bruno Schulze, Henrique Kloh, Matheus Bandini, and Vinod Rebello. Affinity aware scheduling model of cluster nodes in private clouds. Journal of Network and Computer Applications, 95 (??):94–104, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302497.

Zhu:2017:CAR

[2057] Tingwei Zhu, Dan Feng, Fang Wang, Yu Hua, Qingyu Shi, Yanwen Xie, and Yong Wan. A congestion-aware and robust multicast protocol in SDN-based data center networks. Journal of Network and Computer Applications, 95(??):105–117, October 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730245X.

Atiquzzaman:2017:EBP

[2058] Mohammed Atiquzzaman. Editorial: 2016 Best Paper Awards. Journal of Network and Computer Applications, 95(??):A1–A4, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517302680.

Anonymous:2017:EBs

[2059] Anonymous. Editorial board. Journal of Network and Computer Applications, 95(??):ifc, October 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730262X.

Chao:2017:NCA

[2060] Chih-Min Chao, Yun-Ting Ling, and Cheng-Hong Jiang. Number of channels adjustment for cognitive radio networks. Journal of Network and Computer Applications, 96(??):1–13, October 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302503.

DeBenedetti:2017:FTS

[2061] M. De Benedetti, F. D’Urso, G. Fortino, F. Messina, G. Pappalardo, and C. Santoro. A fault-tolerant self-organizing flocking approach for UAV aerial survey. Journal of Network and Computer Applications, 96(??):14–30, October 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302606.

Koohanestani:2017:AMD

[2062] Amir Khorsandi Koohanestani, Amin Ghalami Osgouei, Hossein Saidi, and Ali Fanian. An analytical model for delay bound of OpenFlow based SDN using network calculus. Journal of Network and Computer Applications, 96(??):31–38, October 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517302485.

Gandotra:2017:SGC

[2063] Pimmy Gandotra and Rakesh Kumar Jha. A survey on green communication and security challenges in 5G wireless communication networks. Journal of Network and Computer Applications, 96(??):39–61, October 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302266.

Diyanat:2017:PTP

[2064] Abolfazl Diyanat, Ahmad Khonsari, and Hosein Shafiei. Preservation of temporal privacy in body sensor networks. Journal of Network and Computer Applications, 96(??):62–71, October 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302473.

Anonymous:2017:EBt

[2065] Anonymous. Editorial board. Journal of Network and Computer Applications, 96(??):ifc, October 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302849.

Syarif:2017:OCA

[2066] Abdusy Syarif, Abdelhafid Abouaissa, and Pascal Lorenz. Operator calculus approach for route optimizing and enhancing wireless sensor network. Journal of Network and Computer Applications, 97(??):1–10, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302709.

Vidal:2017:ACF

[2067] Jorge Maestre Vidal, Ana Lucila Sandoval Orozco, and Luis Javier García Villalba. Alert correlation framework for malware detection by anomalybased packet payload analysis. Journal of Network and Computer Applications, 97(??):11–22, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517302746.

Pourghebleh:2017:DAM

[2068] Behrouz Pourghebleh and Nima Jafari Navimipour. Data aggregation mechanisms in the Internet of things: a systematic review of the literature and recommendations for future research. Journal of Network and Computer Applications, 97(??):23–34, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302692.

Nachiappan:2017:CSR

[2069] Rekha Nachiappan, Bahman Javadi, Rodrigo N. Calheiros, and Kenan M. Matawie. Cloud storage reliability for Big Data applications: a state of the art survey. Journal of Network and Computer Applications, 97(??):35–47, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302734.

Alaa:2017:RSH

[2070] Mussab Alaa, A. A. Zaidan, B. B. Zaidan, Mohammed Talal, and M. L. M. Kiah. A review of smart home applications based on Internet of Things. Journal of Network and Computer Applications, 97 (??):48–65, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302801.

Khan:2017:TPK

[2071] Suleman Khan, Muhammad Shiraz, Laleh Boroumand, Abdullah Gani, and Muhammad Khurram Khan. Towards port-knocking authentication methods for mobile cloud computing. Journal of Network and Computer Applications, 97(??):66–78, November 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302813.

Sharma:2017:EED

[2072] Vishal Sharma, Fei Song, Ilsun You, and Mohammed Atiquzzaman. Energy efficient device discovery for reliable communication in 5G-based IoT and BSNs using unmanned aerial vehicles. Journal of Network and Computer Applications, 97(??):79–95, November 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302916.

Queiroz:2017:SSM

[2073] Diego V. Queiroz, Marcelo S. Alencar, Ruan D. Gomes, Iguatemi E. Fonseca, and Cesar Benavente-Peces. Survey and systematic mapping of industrial Wireless Sensor Networks. Journal of Network and Computer Applications, 97(??):96–125, November 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302771.

Kumar:2017:RAF

[2074] Satish Kumar, Arnab Sarkar, and Arijit Sur. A resource allocation framework for adaptive video streaming over LTE. Journal of Network and Computer Applications, 97(??):126–139, November 1, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730276X.

Oyewobi:2017:SCR

[2075] Stephen S. Oyewobi and Gerhard P. Hancke. A survey of cognitive radio handoff schemes, challenges and issues for industrial wireless sensor networks (CR-IWSN). Journal of Network and Computer Applications, 97(??):140–156, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302795.

Shukri:2017:DFL

[2076] Shaufikah Shukri and Latifah Munirah Kamarudin. Device free localization technology for human detection and counting with RF sensor networks: a review. Journal of Network and Computer Applications, 97 (??):157–174, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302928.

Anonymous:2017:EBu

[2077] Anonymous. Editorial board. Journal of Network and Computer Applications, 97(??):ifc, November 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303041.

Niu:2017:ETP

[2078] Chaoyue Niu, Minping Zhou, Zhenzhe Zheng, Fan Wu, and Guihai Chen. ERA: Towards privacy preservation and verifiability for online ad exchanges. Journal of Network and Computer Applications, 98(??):1–10, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302758.

Syed:2017:CMR

[2079] Hassan Jamil Syed, Abdullah Gani, Raja Wasim Ahmad, Muhammad Khurram Khan, and Abdelmuttlib Ibrahim Abdalla Ahmed. Cloud monitoring: a review, taxonomy, and open research issues. Journal of Network and Computer Applications, 98(??):11–26, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302783.

Hu:2017:SFC

[2080] Pengfei Hu, Sahraoui Dhelim, Huansheng Ning, and Tie Qiu. Survey on fog computing: architecture, key technologies, applications and open issues. Journal of Network and Computer Applications, 98 (??):27–42, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302953.

Thakur:2017:TSL

[2081] Avnish Thakur and Major Singh Goraya. A taxonomic survey on load balancing in cloud. Journal of Network and Computer Applications, 98 (??):43–57, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302825.

Khayou:2017:VMN

[2082] Hussein Khayou and Bakr Sarakbi. A validation model for non-lexical routing protocols. Journal of Network and Computer Applications, 98 (??):58–64, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303028.

Ferdaus:2017:AND

[2083] Md Hasanul Ferdaus, Manzur Murshed, Rodrigo N. Calheiros, and Rajkumar Buyya. An algorithm for network and data-aware placement of multi-tier applications in cloud data centers. Journal of Network and Computer Applications, 98(??):65–83, November 15, 2017. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302989.

Gheid:2017:EPA

[2084] Zakaria Gheid, Yacine Challal, Xun Yi, and Abdelouahid Derhab. Efficient and privacy-aware multi-party classification protocol for human activity recognition. Journal of Network and Computer Applications, 98(??):84–96, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303016.

Kulshrestha:2017:SSB

[2085] Tarun Kulshrestha, Divya Saxena, Rajdeep Niyogi, Vaskar Raychoudhury, and Manoj Misra. SmartITS: Smartphone-based identification and tracking using seamless indoor-outdoor localization. Journal of Network and Computer Applications, 98(??):97–113, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517302965.

Basak:2017:CLO

[2086] Surajit Basak and Tamaghna Acharya. Cross layer optimization for outage minimizing routing in cognitive radio ad hoc networks with primary users’ outage protection. Journal of Network and Computer Applications, 98(??):114–124, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302977.

Yuan:2017:EEC

[2087] Peiyan Yuan, Mingyang Song, and Xiaoyan Zhao. Effective and efficient collection of control messages for opportunistic routing algorithms. Journal of Network and Computer Applications, 98(??):125–130, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303132.

Anonymous:2017:EBv

[2088] Anonymous. Editorial board. Journal of Network and Computer Applications, 98(??):ifc, November 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303260.

Wang:2017:NAA

[2089] YiPeng Wang, Xiaochun Yun, Yongzheng Zhang, Liwei Chen, and Guangjun Wu. A nonparametric approach to the automated protocol fingerprint inference. Journal of Network and Computer Applications, 99(??):1–9, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303223.

Kumar:2017:SRF

[2090] Gulshan Kumar, Mritunjay Kumar Rai, and Rahul Saha. Securing range free localization against wormhole attack using distance estimation and maximum likelihood estimation in Wireless Sensor Networks. Journal of Network and Computer Applications, 99(??):10–16, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451730320X.

Torres:2017:TTB

[2091] Luis Miguel Torres, Eduardo Magaña, Daniel Morató, Santiago Garcia-Jimenez, and Mikel Izal. TBDClust: Time-based density clustering to enable free browsing of sites in pay-per-use mobile Internet providers. Journal of Network and Computer Applications, 99(??):17–27, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303193.

Qu:2017:GDM

[2092] Ying Qu, Bryan Ng, and Michael Homer. A goodput distribution model for planning IEEE 802.11 WBNs in built environments. Journal of Network and Computer Applications, 99(??):28–46, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303181.

Jeon:2017:SAP

[2093] Seokseong Jeon, Jae-Pil Jeong, Young-Joo Suh, Chansu Yu, and Dongsoo Han. Selective AP probing for indoor positioning in a large and AP-dense environment. Journal of Network and Computer Applications, 99(??):47–57, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303119.

Sanchez-Carmona:2017:TBO

[2094] Adrián Sánchez-Carmona, Carlos Borrego, Sergi Robles, and Gerard Garcia-Vandellós. Two birds, one stone: Using mobility behavioral profiles both as destinations and as a routing tool. Journal of Network and Computer Applications, 99(??):58–72, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303211.

Zuhra:2017:RPW

[2095] Fatima Tul Zuhra, Kamalrulnizam Abu Bakar, Adnan Ahmed, and Mohsin Ali Tunio. Routing protocols in wireless body sensor networks: a comprehensive survey. Journal of Network and Computer Applications, 99(??):73–97, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303144.

Pease:2017:HTR

[2096] Sarogini Grace Pease, Paul P. Conway, and Andrew A. West. Hybrid ToF and RSSI real-time semantic tracking with an adaptive industrial Internet of Things architecture. Journal of Network and Computer Applications, 99(??):98–109, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517303235.

Xin:2017:CSM

[2097] Yu Xin, Ya-Di Wang, Zhi-Qiang Xie, and Jing Yang. A cooperative scheduling method based on the device load feedback for multiple tasks scheduling. Journal of Network and Computer Applications, 99 (??):110–119, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303107.

Lucas-Estan:2017:DRR

[2098] M. C. Lucas-Estañ and J. Gozalvez. Distributed radio resource allocation for device-to-device communications underlaying cellular networks. Journal of Network and Computer Applications, 99(??):120–130, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303120.

I:2017:ETB

[2099] Indu I., Rubesh Anand P. M., and Vidhyacharan Bhaskar. Encrypted token based authentication with adapted SAML technology for cloud web services. Journal of Network and Computer Applications, 99(??): 131–145, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303156.

Aliberti:2017:EDS

[2100] Giulio Aliberti, Roberto Di Pietro, and Stefano Guarino. Epidemic data survivability in Unattended Wireless Sensor Networks: New models and results. Journal of Network and Computer Applications, 99 (??):146–165, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302990.

Anonymous:2017:EBw

[2101] Anonymous. Editorial board. Journal of Network and Computer Applications, 99(??):ifc, December 1, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303430.

Atiquzzaman:2017:VJ

[2102] Mohammed Atiquzzaman. 100 volumes of JNCA. Journal of Network and Computer Applications, 100(??):iii–iv, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303697.

Wang:2017:DND

[2103] Cong Wang, Maode Ma, and Zenghua Zhao. Design of a novel dynamic trust model for spectrum management in WRANs of TV white space. Journal of Network and Computer Applications, 100(??):1–10, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303004.

Islam:2017:TDE

[2104] A. B. M. Alim Al Islam, Tusher Chakraborty, Taslim Arefin Khan, Mahabub Zoraf, and Chowdhury Sayeed Hyder. Towards defending eavesdropping on NFC. Journal of Network and Computer Applications, 100 (??):11–23, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303405.

Dimitriou:2017:PRA

[2105] Tassos Dimitriou and Ioannis Krontiris. Privacy-respecting auctions and rewarding mechanisms in mobile crowd-sensing applications. Journal of Network and Computer Applications, 100(??):24–34, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303326.

Sandhya:2017:SHS

[2106] Sandhya, Yash Sinha, and K. Haribabu. A survey: Hybrid SDN. Journal of Network and Computer Applications, 100(??):35–55, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451730317X.

Estevez-Ayres:2017:LDS

[2107] Iria Estévez-Ayres, Jesús Arias Fisteus, and Carlos Delgado-Kloos. Lostrego: a distributed stream-based infrastructure for the real-time gathering and analysis of heterogeneous educational data. Journal of Network and Computer Applications, 100(??):56–68, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303417.

Hashem:2017:ETR

[2108] Mohammed Hashem, Shrief I. Barakat, and Mahmoud A. AttaAlla. Enhanced tree routing protocols for multi-hop and multi-channel cognitive radio network (EMM–TRP). Journal of Network and Computer Applications, 100(??):69–79, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804517303570.

Aslam:2017:ICC

[2109] Sidra Aslam, Saif ul Islam, Abid Khan, Mansoor Ahmed, Adnan Akhundzada, and Muhammad Khurram Khan. Information collection centric techniques for cloud resource management: Taxonomy, analysis and challenges. Journal of Network and Computer Applications, 100 (??):80–94, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303545.

Anonymous:2017:EBx

[2110] Anonymous. Editorial board. Journal of Network and Computer Applications, 100(??):ifc, December 15, 2017. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303624.

Anonymous:2018:TYR

[2111] Anonymous. Thank you reviewers. Journal of Network and Computer Applications, 101(??):iii–ix, January 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804517303867.

Thanigaivelan:2018:CCB

[2112] Nanda Kumar Thanigaivelan, Ethiopia Nigussie, Antti Hakkala, Seppo Virtanen, and Jouni Isoaho. CoDRA: Context-based dynamically reconfigurable access control system for Android. Journal of Network and Computer Applications, 101(??):1–17, January 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730351X.

Ullah:2018:DER

[2113] Faheem Ullah, Matthew Edwards, Rajiv Ramdhany, Ruzanna Chitchyan, M. Ali Babar, and Awais Rashid. Data exfiltration: a review of external attack vectors and countermeasures. Journal of Network and Computer Applications, 101(??):18–54, January 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303569.

Ferrag:2018:SCN

[2114] Mohamed Amine Ferrag, Leandros Maglaras, Antonios Argyriou, Dimitrios Kosmanos, and Helge Janicke. Security for 4G and 5G cellular networks: a survey of existing authentication and privacy-preserving schemes. Journal of Network and Computer Applications, 101(??):55– 82, January 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303521.

Shen:2018:CWS

[2115] Yao Shen, Wei Yang, and Liusheng Huang. Concealed in web surfing: Behavior-based covert channels in HTTP. Journal of Network and Computer Applications, 101(??):83–95, January 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303508.

Parsch:2018:RCT

[2116] Philip Parsch and Alejandro Masrur. On reliable communication in transmit-only networks for home automation. Journal of Network and Computer Applications, 101(??):96–110, January 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303533.

Wang:2018:SCD

[2117] Lei Wang, Jianwei Niu, Houbing Song, and Mohammed Atiquzzaman. SentiRelated: a cross-domain sentiment classification algorithm for short texts through sentiment related index. Journal of Network and Computer Applications, 101(??):111–119, January 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303582.

Anonymous:2018:EBa

[2118] Anonymous. Editorial Board. Journal of Network and Computer Applications, 101(??):ifc, January 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730379X.

Anonymous:2018:EBb

[2119] Anonymous. Editorial Board. Journal of Network and Computer Applications, 102(??):ii, January 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517304101.

Cao:2018:EEG

[2120] Jin Cao, Maode Ma, Hui Li, Yulong Fu, and Xuefeng Liu. EGHR: Efficient group-based handover authentication protocols for mMTC in 5G wireless networks. Journal of Network and Computer Applications, 102(??):1–16, January 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303776.

Ravi:2018:ACI

[2121] Kumar Ravi, Yogesh Khandelwal, Boora Shiva Krishna, and Vadlamani Ravi. Analytics in/for cloud — an interdependence: a review.

Journal of Network and Computer Applications, 102(??):17–37, January 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303764.

Sharma:2018:SEP

[2122] Vishal Sharma, Ilsun You, Fang-Yie Leu, and Mohammed Atiquzzaman. Secure and efficient protocol for fast handover in 5G mobile Xhaul networks. Journal of Network and Computer Applications, 102 (??):38–57, January 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303727.

Emmanuel:2018:SDP

[2123] Naina Emmanuel, Abid Khan, Masoom Alam, Tanveer Khan, and Muhammad Khurram Khan. Structures and data preserving homomorphic signatures. Journal of Network and Computer Applications, 102(??):58–70, January 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303739.

Mir:2018:TTH

[2124] Usama Mir and Zeeshan Ahmed Bhatti. Time triggered handoff schemes in cognitive radio networks: a survey. Journal of Network and Computer Applications, 102(??):71–85, January 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303752.

Anonymous:2018:EBc

[2125] Anonymous. Editorial Board. Journal of Network and Computer Applications, 103(??):i, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517304307.

deAssuncao:2018:DDS

[2126] Marcos Dias de Assunção, Alexandre da Silva Veith, and Rajkumar Buyya. Distributed data stream processing and edge computing: a survey on resource elasticity and future directions. Journal of Network and Computer Applications, 103(??):1–17, February 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303971.

Zhan:2018:EKG

[2127] Furui Zhan, Nianmin Yao, Zhenguo Gao, and Haitao Yu. Efficient key generation leveraging wireless channel reciprocity for MANETs. Journal of Network and Computer Applications, 103(??):18–28, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303909.

Kour:2018:CSS

[2128] Haneet Kour, Rakesh Kumar Jha, and Sanjeev Jain. A comprehensive survey on spectrum sharing: Architecture, energy efficiency and security issues. Journal of Network and Computer Applications, 103 (??):29–57, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303879.

Midya:2018:MOO

[2129] Sadip Midya, Asmita Roy, Koushik Majumder, and Santanu Phadikar. Multi-objective optimization technique for resource allocation and task scheduling in vehicular cloud architecture: a hybrid adaptive nature inspired approach. Journal of Network and Computer Applications, 103(??):58–84, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303958.

Chandakanna:2018:RRR

[2130] Veerabhadra Rao Chandakanna. REHDFS: a random read/write enhanced HDFS. Journal of Network and Computer Applications, 103 (??):85–100, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730396X.

Zhang:2018:SSD

[2131] Yuan Zhang, Lin Cui, Wei Wang, and Yuxiang Zhang. A survey on software defined networking with multiple controllers. Journal of Network and Computer Applications, 103(??):101–118, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303934.

Cui:2018:DNE

[2132] Laizhong Cui, Huaixiong Hu, Shui Yu, Qiao Yan, Zhong Ming, Zhenkun Wen, and Nan Lu. DDSE: a novel evolutionary algorithm based on degree-descending search strategy for influence maximization in social networks. Journal of Network and Computer Applications, 103 (??):119–130, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303995.

Aceto:2018:MCA

[2133] Giuseppe Aceto, Domenico Ciuonzo, Antonio Montieri, and Antonio Pescapé. Multi-classification approaches for classifying mobile app traffic. Journal of Network and Computer Applications, 103(??):131–145, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303740.

Huang:2018:SLS

[2134] Yan Huang, Zhipeng Cai, and Anu G. Bourgeois. Search locations safely and accurately: a location privacy protection algorithm with accurate service. Journal of Network and Computer Applications, 103 (??):146–156, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303983.

Zakhary:2018:LPO

[2135] Sameh Zakhary and Abderrahim Benslimane. On location-privacy in opportunistic mobile networks, a survey. Journal of Network and Computer Applications, 103(??):157–170, February 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303557.

Lou:2018:FPR

[2136] Jungang Lou, Yunliang Jiang, Qing Shen, and Ruiqin Wang. Failure prediction by relevance vector regression with improved quantum-inspired gravitational search. Journal of Network and Computer Applications, 103(??):171–177, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517303946.

Zhou:2018:QTA

[2137] Lu Zhou, Quanlong Wang, Xin Sun, Piotr Kulicki, and Arcangelo Castiglione. Quantum technique for access control in cloud computing II: Encryption and key distribution. Journal of Network and Computer Applications, 103(??):178–184, February 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303922.

Yang:2018:BBP

[2138] Changsong Yang, Xiaofeng Chen, and Yang Xiang. Blockchain-based publicly verifiable data deletion scheme for cloud storage. Journal of Network and Computer Applications, 103(??):185–193, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303910.

Li:2018:TFA

[2139] Xiong Li, Jianwei Niu, Saru Kumari, Fan Wu, Arun Kumar Sangaiah, and Kim-Kwang Raymond Choo. A three-factor anonymous authentication scheme for wireless sensor networks in Internet of Things environments. Journal of Network and Computer Applications, 103 (??):194–204, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302254.

Han:2018:RUA

[2140] Guangjie Han, Wenhui Que, Gangyong Jia, and Wenbo Zhang. Resourceutilization-aware energy efficient server consolidation algorithm for green computing in IIOT. Journal of Network and Computer Applications, 103(??):205–214, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302345.

Lin:2018:SPR

[2141] Qinying Lin, Houbing Song, Xiaolin Gui, Xiaoping Wang, and Saiyu Su. A shortest path routing algorithm for unmanned aerial systems based on grid position. Journal of Network and Computer Applications, 103 (??):215–224, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517302710.

Cao:2018:DOI

[2142] Bin Cao, Jianwei Zhao, Zhihan Lv, Xin Liu, Xinyuan Kang, and Shan Yang. Deployment optimization for 3D industrial wireless sensor networks based on particle swarm optimizers with distributed parallelism. Journal of Network and Computer Applications, 103(??):225–238, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302722.

Zhang:2018:IPD

[2143] Weishan Zhang, Gaowa Wulan, Jia Zhai, Liang Xu, Dehai Zhao, Xin Liu, Su Yang, and Jiehan Zhou. An intelligent power distribution service architecture using cloud computing and deep learning techniques. Journal of Network and Computer Applications, 103(??):239–248, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517302941.

Noor:2018:CCT

[2144] Muzzamil Noor, Haider Abbas, and Waleed Bin Shahid. Countering cyber threats for industrial applications: an automated approach for malware evasion detection and analysis. Journal of Network and Computer Applications, 103(??):249–261, February 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303168.

Gai:2018:SF

[2145] Keke Gai, Meikang Qiu, and Xiaotong Sun. A survey on FinTech. Journal of Network and Computer Applications, 103(??):262–273, February 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303247.

Zeng:2018:QSA

[2146] Deze Zeng, Shiyan Zhang, Lin Gu, Shui Yu, and Zhangjie Fu. Qualityof-sensing aware budget constrained contaminant detection sensor deployment in water distribution system. Journal of Network and Computer Applications, 103(??):274–279, February 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303491.

Anonymous:2018:EBd

[2147] Anonymous. Editorial board. Journal of Network and Computer Applications, 104(??):ii, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300146.

Paul:2018:MUF

[2148] Amrita Bose Paul, Santosh Biswas, Sukumar Nandi, and Sandip Chakraborty. MATEM: A unified framework based on trust and MCDM for assuring security, reliability and QoS in DTN routing. Journal of Network and Computer Applications, 104(??):1–20, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304010.

Zareei:2018:MAM

[2149] Mahdi Zareei, A. K. M. Muzahidul Islam, Cesar Vargas-Rosales, Nafees Mansoor, Shidrokh Goudarzi, and Mubashir Husain Rehmani. Mobilityaware medium access control protocols for wireless sensor networks: A survey. Journal of Network and Computer Applications, 104(??):21–37, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304216.

Yu:2018:PPA

[2150] Shuo Yu, Jiaying Liu, Zhuo Yang, Zhen Chen, Huizhen Jiang, Amr Tolba, and Feng Xia. PAVE: Personalized Academic Venue recommendation Exploiting co-publication networks. Journal of Network and Computer Applications, 104(??):38–47, February 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517304009.

Singh:2018:CCR

[2151] Jai Sukh Paul Singh and Mritunjay Kumar Rai. CROP: Cognitive radio ROuting Protocol for link quality channel diverse cognitive networks. Journal of Network and Computer Applications, 104(??):48–60, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304186.

AlSaeed:2018:MSD

[2152] Zainab AlSaeed, Imtiaz Ahmad, and Iftekhar Hussain. Multicasting in software defined networks: A comprehensive survey. Journal of Network and Computer Applications, 104(??):61–77, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451730406X.

Ruan:2018:DFH

[2153] Wenjie Ruan, Quan Z. Sheng, Lina Yao, Xue Li, Nickolas J. G. Falkner, and Lei Yang. Device-free human localization and tracking with UHF passive RFID tags: A data-driven approach. Journal of Network and Computer Applications, 104(??):78–96, February 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517304228.

Fu:2018:DIB

[2154] Anmin Fu, Yuhan Li, Shui Yu, Yan Yu, and Gongxuan Zhang. DI-POR: An IDA-based dynamic proof of retrievability scheme for cloud storage systems. Journal of Network and Computer Applications, 104 (??):97–106, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517304034.

Canovas:2018:SRA

[2155] Alejandro Cánovas, Miran Taha, Jaime Lloret, and Jesús Tomás. Smart resource allocation for improving QoE in IP multimedia subsystems. Journal of Network and Computer Applications, 104(??):107–116, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304277.

Rezaeifar:2018:TBM

[2156] Zeinab Rezaeifar, Jian Wang, and Heekuck Oh. A trust-based method for mitigating cache poisoning in Name Data Networking. Journal of Network and Computer Applications, 104(??):117–132, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304046.

Amin:2018:UAP

[2157] Ruhul Amin, S. K. Hafizul Islam, Neeraj Kumar, and Kim-Kwang Raymond Choo. An untraceable and anonymous password authentication protocol for heterogeneous wireless sensor networks. Journal of Network and Computer Applications, 104(??):133–144, February 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304058.

Anonymous:2018:EBe

[2158] Anonymous. Editorial board. Journal of Network and Computer Applications, 105(??):ii, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300353.

Bahnasy:2018:ZQE

[2159] Mahmoud Bahnasy, Halima Elbiaze, and Bochra Boughzala. Zeroqueue Ethernet congestion control protocol based on available bandwidth estimation. Journal of Network and Computer Applications, 105(??):1–20, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451730423X.

Amiri:2018:SPM

[2160] Maryam Amiri, Leyli Mohammad-Khanli, and Raffaela Mirandola. A sequential pattern mining model for application workload prediction in cloud environment. Journal of Network and Computer Applications, 105(??):21–62, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517304198.

Vidal:2018:SSC

[2161] Ivan Vidal, Jaime Garcia-Reinoso, Ignacio Soto, Francisco Valera, and Diego Lopez. SCoT: A secure content-oriented transport. Journal of Network and Computer Applications, 105(??):63–78, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300018.

Beechu:2018:EEF

[2162] Naresh Kumar Reddy Beechu, Vasantha Moodabettu Harishchandra, and Nithin Kumar Yernad Balachandra. An energy-efficient fault-aware core mapping in mesh-based network on chip systems. Journal of Network and Computer Applications, 105(??):79–87, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517304265.

Blasco:2018:DAC

[2163] Jorge Blasco, Thomas M. Chen, Igor Muttik, and Markus Roggenbach. Detection of app collusion potential using logic programming. Journal of Network and Computer Applications, 105(??):88–104, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304204.

Liu:2018:DCA

[2164] Gao Liu, Zheng Yan, and Witold Pedrycz. Data collection for attack detection and security measurement in mobile ad hoc networks: A survey. Journal of Network and Computer Applications, 105(??):105–122, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300043.

Naz:2018:TSP

[2165] André Naz, Benoît Piranda, Julien Bourgeois, and Seth Copen Goldstein. A time synchronization protocol for large-scale distributed embedded systems with low-precision clocks and neighbor-to-neighbor communications. Journal of Network and Computer Applications, 105 (??):123–142, March 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804517304253.

Anonymous:2018:EBf

[2166] Anonymous. Editorial board. Journal of Network and Computer Applications, 106(??):ii, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300493.

Nardini:2018:PFS

[2167] G. Nardini, G. Stea, A. Virdis, A. Frangioni, L. Galli, D. Sabella, and G. M. Dell’Aera. Practical feasibility, scalability and effectiveness of coordinated scheduling algorithms in cellular networks towards 5G. Journal of Network and Computer Applications, 106(??):1–16, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300213.

Peng:2018:IAS

[2168] Sancheng Peng, Yongmei Zhou, Lihong Cao, Shui Yu, Jianwei Niu, and Weijia Jia. Influence analysis in social networks: A survey. Journal of Network and Computer Applications, 106(??):17–32, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804518300195.

Iqbal:2018:AFS

[2169] Shahid Md. Asif Iqbal and Asaduzzaman. Adaptive forwarding strategies to reduce redundant interests and data in named data networks. Journal of Network and Computer Applications, 106(??):33–47, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300274.

Muduli:2018:AWS

[2170] Lalatendu Muduli, Devi Prasad Mishra, and Prasanta K. Jana. Application of wireless sensor network for environmental monitoring in underground coal mines: A systematic review. Journal of Network and Computer Applications, 106(??):48–67, March 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517304368.

M:2018:HTL

[2171] Koushik A. M., Elizabeth Bentley, Fei Hu, and Sunil Kumar. A hardware testbed for learning-based spectrum handoff in cognitive radio networks. Journal of Network and Computer Applications, 106(??):68–77, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517303594.

Amato:2018:MDD

[2172] Flora Amato, Nicola Mazzocca, and Francesco Moscato. Model driven design and evaluation of security level in orchestrated cloud services. Journal of Network and Computer Applications, 106(??):78–89, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804517304022.

Yang:2018:RKF

[2173] Li Yang, Ziyi Han, Zhengan Huang, and Jianfeng Ma. A remotely keyed file encryption scheme under mobile cloud computing. Journal of Network and Computer Applications, 106(??):90–99, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517304241.

Li:2018:OPP

[2174] Tong Li, Zhengan Huang, Ping Li, Zheli Liu, and Chunfu Jia. Outsourced privacy-preserving classification service over encrypted data.

Journal of Network and Computer Applications, 106(??):100–110, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804517304356.

Jin:2018:IPB

[2175] Xiaohui Jin, Baojiang Cui, Dong Li, Zishuai Cheng, and Congxian Yin. An improved payload-based anomaly detector for web applications. Journal of Network and Computer Applications, 106(??):111–116, March 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451830002X.

Shen:2018:CAL

[2176] Jian Shen, Ziyuan Gui, Sai Ji, Jun Shen, Haowen Tan, and Yi Tang. Cloud-aided lightweight certificateless authentication protocol with anonymity for wireless body area networks. Journal of Network and Computer Applications, 106(??):117–123, March 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300031.

Anonymous:2018:EBg

[2177] Anonymous. Editorial board. Journal of Network and Computer Applications, 107(??):ii, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300614.

Ellouze:2018:PSC

[2178] Nourhene Ellouze, Slim Rekhis, Noureddine Boudriga, and Mohamed Allouche. Powerless security for cardiac implantable medical devices: Use of wireless identification and sensing platform. Journal of Network and Computer Applications, 107(??):1–21, April 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300237.

Segura-Garcia:2018:PED

[2179] Jaume Segura-Garcia, Santiago Felici-Castell, and Miguel Garcia-Pineda. Performance evaluation of different techniques to estimate subjective quality in live video streaming applications over LTE-Advance mobile networks. Journal of Network and Computer Applications, 107(??):22–37, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300201.

Tirani:2018:PSP

[2180] Shima Pakdaman Tirani and Avid Avokh. On the performance of sink placement in WSNs considering energy-balanced compressive sensingbased data aggregation. Journal of Network and Computer Applications, 107(??):38–55, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300262.

Sobral:2018:FEP

[2181] José V. V. Sobral, Joel J. P. C. Rodrigues, Ricardo A. L. Rabelo, José C. Lima Filho, Natanael Sousa, Harilton S. Araujo, and Raimir Holanda Filho. A framework for enhancing the performance of Internet of Things applications based on RFID and WSNs. Journal of Network and Computer Applications, 107(??):56–68, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518300298.

Tan:2018:RPM

[2182] Yu an Tan, Yuan Xue, Chen Liang, Jun Zheng, Quanxin Zhang, Jiamin Zheng, and Yuanzhang Li. A root privilege management scheme with revocable authorization for Android devices. Journal of Network and Computer Applications, 107(??):69–82, April 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300250.

Xu:2018:NEM

[2183] Guangquan Xu, Jia Liu, Yanrong Lu, Xianjiao Zeng, Yao Zhang, and Xiaoming Li. A novel efficient MAKA protocol with desynchronization for anonymous roaming service in Global Mobility Networks. Journal of Network and Computer Applications, 107(??):83–92, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300407.

Giluka:2018:ECB

[2184] Mukesh Kumar Giluka, Thomas Valerrian Pasca, Tathagat Priyadarshi, and Bheemarjuna Reddy Tamma. Enhanced class based dynamic priority scheduling to support uplink IoT traffic in LTE-A networks. Journal of Network and Computer Applications, 107(??):93–112, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804518300249.

Xu:2018:DFH

[2185] Jian Xu, Laiwen Wei, Yu Zhang, Andi Wang, Fucai Zhou, and Chong zhi Gao. Dynamic fully homomorphic encryption-based Merkle tree for lightweight streaming authenticated data structures. Journal of Network and Computer Applications, 107(??):113–124, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300286.

Aceto:2018:RIC

[2186] Giuseppe Aceto, Valerio Persico, and Antonio Pescapé. The role of information and communication technologies in healthcare: taxonomies, perspectives, and challenges. Journal of Network and Computer Applications, 107(??):125–154, April 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300456.

Anonymous:2018:EBh

[2187] Anonymous. Editorial board. Journal of Network and Computer Applications, 108(??):ii, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451830078X.

Saleem:2018:RAG

[2188] Ahsan Saleem, Abid Khan, Furqan Shahid, M. Masoom Alam, and Muhammad Khurram Khan. Recent advancements in garbled computing: How far have we come towards achieving secure, efficient and reusable garbled circuits. Journal of Network and Computer Applications, 108(??):1–19, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300432.

Dong:2018:RTL

[2189] Pingping Dong, Wenjun Yang, Wensheng Tang, Jiawei Huang, Haodong Wang, Yi Pan, and Jianxin Wang. Reducing transport latency for short flows with multipath TCP. Journal of Network and Computer Applications, 108(??):20–36, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300420.

P:2018:ABE

[2190] Praveen Kumar P, Syam Kumar P, and Alphonse P. J. A. Attribute based encryption in cloud computing: A survey, gap analysis, and future directions. Journal of Network and Computer Applications, 108(??):37–52, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300547.

Lalouani:2018:IIN

[2191] Wassila Lalouani, Mohamed Younis, and Nadjib Badache. Interconnecting isolated network segments through intermittent links. Journal of Network and Computer Applications, 108(??):53–63, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300559.

Cao:2018:EFG

[2192] Jiuxin Cao, Shuai Xu, Xuelin Zhu, Renjun Lv, and Bo Liu. Effective fine-grained location prediction based on user check-in pattern in LBSNs. Journal of Network and Computer Applications, 108(??):64– 75, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300444.

Colom:2018:SFD

[2193] José Francisco Colom, David Gil, Higinio Mora, Bruno Volckaert, and Antonio Manuel Jimeno. Scheduling framework for distributed intrusion detection systems over heterogeneous network architectures. Journal of Network and Computer Applications, 108(??):76–86, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804518300419.

Javed:2018:CDN

[2194] Muhammad Aqib Javed, Muhammad Shahzad Younis, Siddique Latif, Junaid Qadir, and Adeel Baig. Community detection in networks: A multidisciplinary review. Journal of Network and Computer Applications, 108(??):87–111, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300560.

Liu:2018:PAB

[2195] Zechao Liu, Zoe L. Jiang, Xuan Wang, and S. M. Yiu. Practical attribute-based encryption: Outsourcing decryption, attribute revocation and policy updating. Journal of Network and Computer Applications, 108(??):112–123, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518300304.

Esposito:2018:IDP

[2196] Christian Esposito. Interoperable, dynamic and privacy-preserving access control for cloud data storage when integrating heterogeneous organizations. Journal of Network and Computer Applications, 108 (??):124–136, April 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300316.

Anonymous:2018:EBi

[2197] Anonymous. Editorial board. Journal of Network and Computer Applications, 109(??):ii, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301000.

Thai:2018:TLB

[2198] Minh-Tuan Thai, Ying-Dar Lin, Po-Ching Lin, and Yuan-Cheng Lai. Towards load-balanced service chaining by Hash-based Traffic Steering on Softswitches. Journal of Network and Computer Applications, 109(??): 1–10, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300699.

Basu:2018:WPW

[2199] Souvik Basu, Ayanesh Biswas, Siuli Roy, and Sipra DasBit. Wise-PRoPHET: A watchdog supervised PRoPHET for reliable dissemination of post disaster situational information over smartphone based DTN. Journal of Network and Computer Applications, 109(??):11– 23, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300572.

Ehatisham-ul-Haq:2018:CAS

[2200] Muhammad Ehatisham ul Haq, Muhammad Awais Azam, Usman Naeem, Yasar Amin, and Jonathan Loo. Continuous authentication of smartphone users based on activity pattern recognition using passive mobile sensing. Journal of Network and Computer Applications, 109(??):24–35, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300717.

Rao:2018:SAI

[2201] Yuan Rao, Cheng Deng, Gang Zhao, Yan Qiao, Lei yang Fu, Xing Shao, and Ru chuan Wang. Self-adaptive implicit contention window adjustment mechanism for QoS optimization in wireless sensor networks. Journal of Network and Computer Applications, 109(??):36– 52, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300742.

daCruz:2018:PEI

[2202] Mauro A. A. da Cruz, Joel J. P. C. Rodrigues, Arun Kumar Sangaiah, Jalal Al-Muhtadi, and Valery Korotaev. Performance evaluation of IoT middleware. Journal of Network and Computer Applications, 109(??):53–65, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451830064X.

Lopez:2018:CDF

[2203] Ester López and Leandro Navarro. Coordinated detection of forwarding faults in wireless community networks. Journal of Network and Computer Applications, 109(??):66–77, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518300705.

Wu:2018:PAD

[2204] Xuangou Wu, Panlong Yang, Shaojie Tang, Xiao Zheng, and Xiaolin Wang. Privacy-aware data publishing against sparse estimation attack. Journal of Network and Computer Applications, 109(??):78– 88, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300675.

Deepa:2018:DBB

[2205] G. Deepa, P. Santhi Thilagam, Amit Praseed, and Alwyn R. Pais. DetLogic: A black-box approach for detecting logic vulnerabilities in web applications. Journal of Network and Computer Applications, 109(??):89–109, May 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300225.

Anonymous:2018:EBj

[2206] Anonymous. Editorial board. Journal of Network and Computer Applications, 110(??):ii, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301218.

Noshy:2018:OLV

[2207] Mostafa Noshy, Abdelhameed Ibrahim, and Hesham Arafat Ali. Optimization of live virtual machine migration in cloud computing: A survey and future directions. Journal of Network and Computer Applications, 110(??):1–10, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300833.

Sheikh:2018:PIT

[2208] Umber Sheikh, Abid Khan, Bilal Ahmed, Abdul Waheed, and Abdul Hameed. Provenance inference techniques: Taxonomy, comparative analysis and design challenges. Journal of Network and Computer Applications, 110(??):11–26, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300857.

Guirguis:2018:CBM

[2209] Arsany Guirguis, Mohammed Karmoose, Karim Habak, Mustafa El-Nainay, and Moustafa Youssef. Cooperation-based multi-hop routing protocol for cognitive radio networks. Journal of Network and Computer Applications, 110(??):27–42, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518300869.

Yang:2018:RVC

[2210] Qiang Yang, Wei Li, Jose Neuman de Souza, and Albert Y. Zomaya. Resilient virtual communication networks using multi-commodity flow based local optimal mapping. Journal of Network and Computer Applications, 110(??):43–51, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300730.

Hayyolalam:2018:SLR

[2211] Vahideh Hayyolalam and Ali Asghar Pourhaji Kazem. A systematic literature review on QoS-aware service composition and selection in cloud environment. Journal of Network and Computer Applications, 110(??):52–74, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300845.

Tan:2018:SPR

[2212] Choon Beng Tan, Mohd Hanafi Ahmad Hijazi, Yuto Lim, and Abdullah Gani. A survey on proof of retrievability for cloud data integrity and availability: Cloud storage state-of-the-art, issues, solutions and future trends. Journal of Network and Computer Applications, 110(??):75–86, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301048.

Murti:2018:SMB

[2213] Fahri Wisnu Murti, Rahmat Faddli Siregar, and Soo Young Shin. SU-MIMO based uplink non-orthogonal multiple access for 5G. Journal of Network and Computer Applications, 110(??):87–96, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804518300900.

Li:2018:SIS

[2214] Jin Li, Aniello Castiglione, and Changyu Dong. Special issue on security in cloud computing. Journal of Network and Computer Applications, 110(??):97–98, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301036.

Jakobik:2018:SGM

[2215] Agnieszka Jakóbik, Francesco Palmieri, and Joanna Kolodziej. Stackelberg games for modeling defense scenarios against cloud security threats. Journal of Network and Computer Applications, 110(??):99– 107, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518300663.

Liao:2018:LTP

[2216] Dan Liao, Hui Li, Gang Sun, Ming Zhang, and Victor Chang. Location and trajectory privacy preservation in 5G-enabled vehicle social network services. Journal of Network and Computer Applications, 110(??):108–118, May 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300390.

Anonymous:2018:EBk

[2217] Anonymous. Editorial board. Journal of Network and Computer Applications, 111(??):ii, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301395.

Bhattacharjee:2018:DDP

[2218] Suman Bhattacharjee, Siuli Roy, and Sipra DasBit. DPDRM: A decentralized post-disaster resource management scheme using energy efficient smart phone based DTN. Journal of Network and Computer Applications, 111(??):1–16, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300882.

Zhang:2018:HCA

[2219] Hongtao Zhang, Ying Chen, and Zihua Yang. Hierarchical cache-aided transmission cooperation in 5G user-centric network: Performance analysis and design insights. Journal of Network and Computer Applications, 111(??):17–27, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300924.

Li:2018:FBM

[2220] Aohan Li and Guangjie Han. A fairness-based MAC protocol for 5G cognitive radio ad hoc networks. Journal of Network and Computer Applications, 111(??):28–34, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518301061.

Garcia-Dorado:2018:DWF

[2221] José Luis García-Dorado, Javier Ramos, Miguel Rodríguez, and Javier Aracil. DNS weighted footprints for web browsing analytics. Journal of Network and Computer Applications, 111(??):35–48, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804518300894.

Behal:2018:DFA

[2222] Sunny Behal, Krishan Kumar, and Monika Sachdeva. D-FACE: An anomaly based distributed approach for early detection of DDoS attacks and flash events. Journal of Network and Computer Applications, 111(??):49–63, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301115.

Tseng:2018:TSC

[2223] Chih-Cheng Tseng and Jyun-Yao Shih. Two-stage coalition formation and radio resource allocation with Nash bargaining solution for inband underlaid D2D communications in 5G networks. Journal of Network and Computer Applications, 111(??):64–76, June 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300651.

Muchtar:2018:ECS

[2224] Farkhana Muchtar, Abdul Hanan Abdullah, Suhaidi Hassan, and Farhan Masud. Energy conservation strategies in host centric networking based MANET: A review. Journal of Network and Computer Applications, 111(??):77–98, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301073.

Alubady:2018:PIT

[2225] Raaid Alubady, Suhaidi Hassan, and Adib Habbal. Pending interest table control management in named data network. Journal of Network and Computer Applications, 111(??):99–116, June 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804517303600.

Xu:2018:HCC

[2226] Guangliang Xu, Wei Yang, and Liusheng Huang. Hybrid covert channel in LTE-A: Modeling and analysis. Journal of Network and Computer Applications, 111(??):117–126, June 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518300389.

Anonymous:2018:EBl

[2227] Anonymous. Editorial board. Journal of Network and Computer Applications, 112(??):ii, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301541.

Ahmed:2018:RMC

[2228] Faran Ahmed, Muhammad Naeem, Waleed Ejaz, Muhammad Iqbal, and Alagan Anpalagan. Resource management in cellular base stations powered by renewable energy sources. Journal of Network and Computer Applications, 112(??):1–17, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518301085.

Cheng:2018:WOU

[2229] Rung-Shiang Cheng, Chung-Ming Huang, and Sheng-Yuan Pan. WiFi offloading using the device-to-device (D2D) communication paradigm based on the Software Defined Network (SDN) architecture. Journal of Network and Computer Applications, 112(??):18–28, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451830095X.

Sun:2018:ALT

[2230] Ziwen Sun, Yuhui Liu, and Li Tao. Attack localization task allocation in wireless sensor networks based on multi-objective binary particle swarm optimization. Journal of Network and Computer Applications, 112(??):29–40, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301103.

Benzerbadj:2018:SSF

[2231] Ali Benzerbadj, Bouabdellah Kechar, Ahcène Bounceur, and Mohammad Hammoudeh. Surveillance of sensitive fenced areas using duty-cycled wireless sensor networks with asymmetrical links. Journal of Network and Computer Applications, 112(??):41–52, June 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301140.

Kaur:2018:RSC

[2232] Ravneet Kaur, Sarbjeet Singh, and Harish Kumar. Rise of spam and compromised accounts in online social networks: A state-of-the-art review of different combating approaches. Journal of Network and Computer Applications, 112(??):53–88, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518300961.

Li:2018:MAF

[2233] Jin Li, Xiaofeng Chen, Sherman S. M. Chow, Qiong Huang, Duncan S. Wong, and Zheli Liu. Multi-authority fine-grained access control with accountability and its application in cloud. Journal of Network and Computer Applications, 112(??):89–96, June 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300870.

Singh:2018:UBA

[2234] Karanpreet Singh, Paramvir Singh, and Krishan Kumar. User behavior analytics-based classification of application layer HTTP-GET flood attacks. Journal of Network and Computer Applications, 112(??):97– 114, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301176.

Yang:2018:SSG

[2235] Mao Yang, Bo Li, Zhicheng Bai, and Zhongjiang Yan. SGMA: Semigranted multiple access for non-orthogonal multiple access (NOMA) in 5G networking. Journal of Network and Computer Applications, 112(??):115–125, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300912.

Haider:2018:ELB

[2236] Ch. Md. Rakin Haider, Anindya Iqbal, Atif Hasan Rahman, and M. Sohel Rahman. An ensemble learning based approach for impression fraud detection in mobile advertising. Journal of Network and Computer Applications, 112(??):126–141, June 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518300729.

Anonymous:2018:EBm

[2237] Anonymous. Editorial board. Journal of Network and Computer Applications, 113(??):ii, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301668.

Wang:2018:TEP

[2238] Xiong Wang, Qian Zhang, Jing Ren, Shizhong Xu, Sheng Wang, and Shui Yu. Toward efficient parallel routing optimization for large-scale SDN networks using GPGPU. Journal of Network and Computer Applications, 113(??):1–13, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301243.

Jiang:2018:FGR

[2239] Wei Jiang, Wanchun Jiang, Weiping Wang, Haodong Wang, Yi Pan, and Jianxin Wang. A fine-grained rule partition algorithm in cloud data centers. Journal of Network and Computer Applications, 113(??):14– 25, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301127.

Rahman:2018:CJL

[2240] Md. Arafatur Rahman, A. Taufiq Asyhari, Md Zakirul Alam Bhuiyan, Qusay Medhat Salih, and Kamal Zuhairi Bin Zamli. L-CAQ: Joint linkoriented channel-availability and channel-quality based channel selection for mobile cognitive radio networks. Journal of Network and Computer Applications, 113(??):26–35, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518301097.

Aceto:2018:CSI

[2241] Giuseppe Aceto, Alessio Botta, Pietro Marchetta, Valerio Persico, and Antonio Pescapé. A comprehensive survey on Internet outages. Journal of Network and Computer Applications, 113(??):36– 63, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301139.

Tian:2018:MTE

[2242] Wenhong Tian, Majun He, Wenxia Guo, Wenqiang Huang, Xiaoyu Shi, Mingsheng Shang, Adel Nadjaran Toosi, and Rajkumar Buyya. On minimizing total energy consumption in the scheduling of virtual machine reservations. Journal of Network and Computer Applications, 113(??):64–74, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301267.

Huang:2018:UOB

[2243] Xiaohong Huang, Tingting Yuan, and Yan Zhang. Utility-optimized bandwidth and power allocation for non-orthogonal multiple access in software defined 5G networks. Journal of Network and Computer Applications, 113(??):75–86, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300936.

Aalsalem:2018:WSN

[2244] Mohammed Y. Aalsalem, Wazir Zada Khan, Wajeb Gharibi, Muhammad Khurram Khan, and Quratulain Arshad. Wireless sensor networks in oil and gas industry: Recent advances, taxonomy, requirements, and open challenges. Journal of Network and Computer Applications, 113(??):87–97, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301309.

Chao:2018:SFF

[2245] Chih-Min Chao and Hsiang-Yuan Fu. Supporting fast and fair rendezvous for cognitive radio networks. Journal of Network and Computer Applications, 113(??):98–108, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301255.

Qin:2018:FTS

[2246] Yang Qin, Weihong Yang, Xiao Ai, and Lingjian Chen. Fault tolerant storage and data access optimization in data center networks. Journal of Network and Computer Applications, 113(??):109– 118, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301279.

Fahmin:2018:PMC

[2247] Ahmed Fahmin, Yuan-Cheng Lai, Md. Shohrab Hossain, and Ying-Dar Lin. Performance modeling and comparison of NFV integrated with SDN: Under or aside? Journal of Network and Computer Applications, 113(??):119–129, July 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301292.

Anonymous:2018:EBn

[2248] Anonymous. Editorial board. Journal of Network and Computer Applications, 114(??):ii, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301802.

Saleh:2018:SLD

[2249] Saad Saleh, Junaid Qadir, and Muhammad U. Ilyas. Shedding light on the dark corners of the Internet: A survey of Tor research. Journal of Network and Computer Applications, 114(??):1–28, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301280.

Santoyo-Gonzalez:2018:LAC

[2250] Alejandro Santoyo-González and Cristina Cervelló-Pastor. Latencyaware cost optimization of the service infrastructure placement in 5G networks. Journal of Network and Computer Applications, 114(??):29– 37, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301334.

Tseng:2018:EEP

[2251] Hsueh-Wen Tseng, Ting-Ting Yang, Wan-Chi Chang, and Yu-Xiang Lan. An efficient error prevention and recovery scheme for multicast traffic in data center networks. Journal of Network and Computer Applications, 114(??):38–47, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301164.

Guan:2018:GBN

[2252] Jianfeng Guan, Zhijun Wei, and Ilsun You. GRBC-based network security functions placement scheme in SDS for 5G security. Journal of Network and Computer Applications, 114(??):48–56, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518300948.

Elhabyan:2018:POB

[2253] Riham Elhabyan, Wei Shi, and Marc St-Hilaire. A Pareto optimizationbased approach to clustering and routing in wireless sensor networks. Journal of Network and Computer Applications, 114(??):57– 69, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301310.

Robinson:2018:ICP

[2254] Michael Robinson, Kevin Jones, Helge Janicke, and Leandros Maglaras. An introduction to cyber peacekeeping. Journal of Network and Computer Applications, 114(??):70–87, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518301425.

Chien:2018:SSB

[2255] Wei-Che Chien, Chin-Feng Lai, Hsin-Hung Cho, and Han-Chieh Chao. A SDN-SFC-based service-oriented load balancing for the IoT applications. Journal of Network and Computer Applications, 114(??):88– 97, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301358.

Yu:2018:VEA

[2256] Ya-Ju Yu, Ai-Chun Pang, and Ming-Yu Yeh. Video encoding adaptation for QoE maximization over 5G cellular networks. Journal of Network and Computer Applications, 114(??):98–107, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301346.

Hu:2018:MOS

[2257] Haiyang Hu, Zhongjin Li, Hua Hu, Jie Chen, Jidong Ge, Chuanyi Li, and Victor Chang. Multi-objective scheduling for scientific workflow in multicloud environment. Journal of Network and Computer Applications, 114(??):108–122, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301152.

Kaswan:2018:ESS

[2258] Amar Kaswan, Abhinav Tomar, and Prasanta K. Jana. An efficient scheduling scheme for mobile charger in on-demand wireless rechargeable sensor networks. Journal of Network and Computer Applications, 114(??):123–134, July 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518300687.

Anonymous:2018:EBo

[2259] Anonymous. Editorial board. Journal of Network and Computer Applications, 115(??):ii, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301875.

Etemad:2018:VDO

[2260] Mohammad Etemad and Alptekin Küpçü. Verifiable database outsourcing supporting join. Journal of Network and Computer Applications, 115(??):1–19, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301322.

Alam:2018:AFC

[2261] Badiul Alam, Zhe Jin, Wun-She Yap, and Bok-Min Goi. An alignment-free cancelable fingerprint template for bio-cryptosystems.

Journal of Network and Computer Applications, 115(??):20–32, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301450.

Ismail:2018:MPP

[2262] Walaa N. Ismail, Mohammad Mehedi Hassan, Hessah A. Alsalamah, and Giancarlo Fortino. Mining productive-periodic frequent patterns in tele-health systems. Journal of Network and Computer Applications, 115(??):33–47, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301474.

Cuomo:2018:IBS

[2263] Salvatore Cuomo, Raffaele Farina, and Francesco Piccialli. An inverse Bayesian scheme for the denoising of ECG signals. Journal of Network and Computer Applications, 115(??):48–58, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301486.

Akraa:2018:SBP

[2264] Shaymaa Akraa, Anh Pham Tran Tam, Haifeng Shen, Youhong Tang, Ben Zhong Tang, Jimmy Li, and Sandy Walker. A smartphone-based point-of-care quantitative urinalysis device for chronic kidney disease patients. Journal of Network and Computer Applications, 115(??):59– 69, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301449.

Noor:2018:MCC

[2265] Talal H. Noor, Sherali Zeadally, Abdullah Alfazi, and Quan Z. Sheng. Mobile cloud computing: Challenges and future research directions. Journal of Network and Computer Applications, 115(??):70–85, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301504.

Rawashdeh:2018:RSD

[2266] Majdi Rawashdeh, Mohammed Gh. Al Zamil, M. Shamim Hossain, Samer Samarah, Syed Umar Amin, and Ghulam Muhammad. Reliable service delivery in Tele-health care systems. Journal of Network and Computer Applications, 115(??):86–93, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301462.

Hsieh:2018:VMA

[2267] Han-Chuan Hsieh, Jiann-Liang Chen, and Abderrahim Benslimane. 5G virtualized multi-access edge computing platform for IoT applications. Journal of Network and Computer Applications, 115(??):94–102, August 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301577.

Anonymous:2018:EBp

[2268] Anonymous. Editorial Board. Journal of Network and Computer Applications, 116(??):ii, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451830198X.

Gupta:2018:DLD

[2269] Otkrist Gupta and Ramesh Raskar. Distributed learning of deep neural network over multiple agents. Journal of Network and Computer Applications, 116(??):1–8, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301590.

Zhou:2018:SND

[2270] Donghao Zhou, Zheng Yan, Yulong Fu, and Zhen Yao. A survey on network data collection. Journal of Network and Computer Applications, 116(??):9–23, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301607.

Biswash:2018:PBU

[2271] Sanjay Kumar Biswash and Dushantha Nalin K. Jayakody. Performance based user-centric dynamic mode switching and mobility management scheme for 5G networks. Journal of Network and Computer Applications, 116(??):24–34, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301759.

Lin:2018:BLC

[2272] Chun-Cheng Lin, Hui-Hsin Chin, and Wei-Bo Chen. Balancing latency and cost in software-defined vehicular networks using genetic algorithm. Journal of Network and Computer Applications, 116(??):35–41, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301589.

Lin:2018:BBB

[2273] Chao Lin, Debiao He, Xinyi Huang, Kim-Kwang Raymond Choo, and Athanasios V. Vasilakos. BSeIn: a blockchain-based secure mutual authentication with fine-grained access control system for industry 4.0. Journal of Network and Computer Applications, 116(??):42–52, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301619.

Liang:2018:EED

[2274] Jia-Ming Liang, Po-Yen Chang, Jen-Jee Chen, Chien-Feng Huang, and Yu-Chee Tseng. Energy-efficient DRX scheduling for D2D communication in 5G networks. Journal of Network and Computer Applications, 116(??):53–64, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451830170X.

Staffa:2018:OBS

[2275] Mariacarla Staffa, Luigi Sgaglione, Giovanni Mazzeo, Luigi Coppolino, Salvatore D’Antonio, Luigi Romano, Erol Gelenbe, Oana Stan, Sergiu Carpov, Evangelos Grivas, Paolo Campegiani, Luigi Castaldo, Konstantinos Votis, Vassilis Koutkias, and Ioannis Komnios. An OpenNCP-based solution for secure eHealth data exchange. Journal of Network and Computer Applications, 116(??):65–85, August 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301747.

Herreria-Alonso:2018:ODS

[2276] Sergio Herrería-Alonso, Miguel Rodríguez-Pérez, Manuel Fernández-Veiga, and Cándido López-García. An optimal dynamic sleeping control policy for single base stations in green cellular networks. Journal of Network and Computer Applications, 116(??):86–94, August 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301760.

Anonymous:2018:EBq

[2277] Anonymous. Editorial Board. Journal of Network and Computer Applications, 117(??):ii, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518302091.

Meng:2018:TTB

[2278] Weizhi Meng, Yu Wang, Duncan S. Wong, Sheng Wen, and Yang Xiang. TouchWB: Touch behavioral user authentication based on web browsing on smartphones. Journal of Network and Computer Applications, 117(??):1–9, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301723.

Zhang:2018:PPE

[2279] Yin Zhang, Raffaele Gravina, Huimin Lu, Massimo Villari, and Giancarlo Fortino. PEA: Parallel electrocardiogram-based authentication for smart healthcare systems. Journal of Network and Computer Applications, 117(??):10–16, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518301693.

Calcaterra:2018:CDS

[2280] Domenico Calcaterra, Giuseppe Di Modica, Orazio Tomarchio, and Placido Romeo. A clinical decision support system to increase appropriateness of diagnostic imaging prescriptions. Journal of Network and Computer Applications, 117(??):17–29, September 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301735.

Nasir:2018:IPA

[2281] Jamal Abdul Nasir, Shahid Hussain, and Chuangyin Dang. An integrated planning approach towards home health care, telehealth and patients group based care. Journal of Network and Computer Applications, 117(??):30–41, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301711.

Lamaazi:2018:ENE

[2282] Hanane Lamaazi and Nabil Benamar. OF–EC: a novel energy consumption aware objective function for RPL based on fuzzy logic. Journal of Network and Computer Applications, 117(??):42–58, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301838.

Couto:2018:BIC

[2283] Rodrigo S. Couto, Hugo Sadok, Pedro Cruz, Felipe F. da Silva, Tatiana Sciammarella, Miguel Elias M. Campista, Luís Henrique M. K. Costa, Pedro B. Velloso, and Marcelo G. Rubinstein. Building an IaaS cloud with droplets: a collaborative experience with OpenStack. Journal of Network and Computer Applications, 117(??):59–71, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301929.

Yang:2018:TLC

[2284] Yang Yang, Songtao Guo, Guiyan Liu, and Yuanyuan Yang. Twolayer compressive sensing based video encoding and decoding framework for WMSN. Journal of Network and Computer Applications, 117 (??):72–85, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518301942.

Li:2018:CVC

[2285] Yan Li, Shifang Dai, and Weiwei Wu. Cooperative video caching scheme over software defined passive optical network. Journal of Network and Computer Applications, 117(??):86–100, September 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518301930.

Anonymous:2018:EBr

[2286] Anonymous. Editorial Board. Journal of Network and Computer Applications, 118(??):ii, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518302418.

Sofi:2018:SEE

[2287] Ishfaq Bashir Sofi and Akhil Gupta. A survey on energy efficient 5G green network with a planned multi-tier architecture. Journal of Network and Computer Applications, 118(??):1–28, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302029.

Isah:2018:ILM

[2288] Musab Isah, Steven Simpson, and Chris Edwards. An improved LISP mobile node architecture. Journal of Network and Computer Applications, 118(??):29–43, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451830105X.

Capurso:2018:SKF

[2289] Nicholas Capurso, Bo Mei, Tianyi Song, Xiuzhen Cheng, and Jiguo Yu. A survey on key fields of context awareness for mobile devices. Journal of Network and Computer Applications, 118(??):44–60, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301620.

Tsiftes:2018:VVS

[2290] Nicolas Tsiftes and Thiemo Voigt. Velox VM: a safe execution environment for resource-constrained IoT applications. Journal of Network and Computer Applications, 118(??):61–73, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302017.

Glitza:2018:ATL

[2291] Jenny I. Glitza, Frerk Müller von Aschwege, Marco Eichelberg, Nils Reiss, Thomas Schmidt, Christina Feldmann, Regina Wendl, Jan D. Schmitto, and Andreas Hein. Advanced telemonitoring of Left Ventricular Assist Device patients for the early detection of thrombosis. Journal of Network and Computer Applications, 118(??):74–82, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518301437.

Kalaivanan:2018:RLA

[2292] Karunanithy Kalaivanan and Velusamy Bhanumathi. Reliable location aware and Cluster-Tap Root based data collection protocol for large scale wireless sensor networks. Journal of Network and Computer Applications, 118(??):83–101, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302054.

Fu:2018:PPC

[2293] Anmin Fu, Shuai Li, Shui Yu, Yuqing Zhang, and Yinxia Sun. Privacypreserving composite modular exponentiation outsourcing with optimal checkability in single untrusted cloud server. Journal of Network and Computer Applications, 118(??):102–112, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302030.

Sarmah:2018:SDM

[2294] Upasana Sarmah, D. K. Bhattacharyya, and J. K. Kalita. A survey of detection methods for XSS attacks. Journal of Network and Computer Applications, 118(??):113–143, September 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302042.

Liang:2018:BPL

[2295] Chen Liang, Yu an Tan, Xiaosong Zhang, Xianmin Wang, Jun Zheng, and Quanxin Zhang. Building packet length covert channel over mobile VoIP traffics. Journal of Network and Computer Applications, 118 (??):144–153, September 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302200.

Anonymous:2018:EBs

[2296] Anonymous. Editorial Board. Journal of Network and Computer Applications, 119(??):ii, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302534.

DeFalco:2018:GPB

[2297] Ivanoe De Falco, Antonio Della Cioppa, Tomas Koutny, Michal Krcma, Umberto Scafuri, and Ernesto Tarantino. Genetic programmingbased induction of a glucose-dynamics model for telemedicine. Journal of Network and Computer Applications, 119(??):1–13, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804518302157.

Cespedes-Mota:2018:DEA

[2298] Armando Céspedes-Mota, Gerardo Castañón, Alberto F. Martínez-Herrera, Leopoldo Eduardo Cárdenas-Barrón, and Ana María Sarmiento. Differential evolution algorithm applied to wireless sensor distribution on different geometric shapes with area and energy optimization. Journal of Network and Computer Applications, 119(??):14–23, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302145.

Tonyali:2018:RPE

[2299] Samet Tonyali, Ruben Munoz, Kemal Akkaya, and Utku Ozgur. A realistic performance evaluation of privacy-preserving protocols for smart grid AMI networks. Journal of Network and Computer Applications, 119(??):24–41, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302194.

Mostafaei:2018:SDW

[2300] Habib Mostafaei and Michael Menth. Software-defined wireless sensor networks: a survey. Journal of Network and Computer Applications, 119(??):42–56, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302248.

Lv:2018:LAI

[2301] Jianhui Lv, Xingwei Wang, Qingyi Zhang, and Min Huang. LAPGN: Accomplishing information consistency under OSPF in general networks (an extension). Journal of Network and Computer Applications, 119(??):57–69, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302224.

Shah:2018:ASE

[2302] S. A. R. Shah, Wenji Wu, Qiming Lu, Liang Zhang, Sajith Sasidharan, Phil DeMar, Chin Guok, John Macauley, Eric Pouyoul, Jin Kim, and Seo-Young Noh. AmoebaNet: an SDN-enabled network service for big data science. Journal of Network and Computer Applications, 119(??):70–82, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302236.

Gu:2018:POS

[2303] Fei Gu, Jianwei Niu, Zhiping Qi, and Mohammed Atiquzzaman. Partitioning and offloading in smart mobile devices for mobile cloud computing: State of the art and future directions. Journal of Network and Computer Applications, 119(??):83–96, October 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302170.

Wan:2018:ADU

[2304] Xili Wan, Xinjie Guan, Tianjing Wang, Guangwei Bai, and Baek-Yong Choi. Application deployment using Microservice and Docker containers: Framework and optimization. Journal of Network and Computer Applications, 119(??):97–109, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518302273.

Zhang:2018:DCC

[2305] Yongnan Zhang and Yonghua Zhou. Distributed coordination control of traffic network flow using adaptive genetic algorithm based on cloud computing. Journal of Network and Computer Applications, 119 (??):110–120, October 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451830225X.

Anonymous:2018:EBt

[2306] Anonymous. Editorial Board. Journal of Network and Computer Applications, 120(??):ii, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302777.

Jain:2018:FSR

[2307] Abhinav Jain and Sandeep Kumar. FriendShare: a secure and reliable framework for file sharing on network. Journal of Network and Computer Applications, 120(??):1–16, October 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302297.

Bagula:2018:FHS

[2308] A. Bagula, M. Mandava, and H. Bagula. A framework for healthcare support in the rural and low income areas of the developing world. Journal of Network and Computer Applications, 120(??):17–29, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302182.

Moradianzadeh:2018:USN

[2309] Nima Moradianzadeh, Pooya Moradian Zadeh, Ziad Kobti, Sarah Hansen, and Kathryn Pfaff. Using social network analysis to model palliative care. Journal of Network and Computer Applications, 120 (??):30–41, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302285.

Chowdhury:2018:SSI

[2310] Abishi Chowdhury and Shital A. Raut. A survey study on Internet of Things resource management. Journal of Network and Computer Applications, 120(??):42–60, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518302315.

Asghari:2018:SCA

[2311] Parvaneh Asghari, Amir Masoud Rahmani, and Hamid Haj Seyyed Javadi. Service composition approaches in IoT: a systematic review. Journal of Network and Computer Applications, 120(??):61–77, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302376.

Vasudeva:2018:SSA

[2312] Amol Vasudeva and Manu Sood. Survey on sybil attack defense mechanisms in wireless ad hoc networks. Journal of Network and Computer Applications, 120(??):78–118, October 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302303.

Jiang:2018:SAR

[2313] Han-Peng Jiang and Wei-Mei Chen. Self-adaptive resource allocation for energy-aware virtual machine placement in dynamic computing cloud. Journal of Network and Computer Applications, 120(??):119–129, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302352.

Bilal:2018:QAD

[2314] K. Bilal, A. Erbad, and M. Hefeeda. QoE-aware distributed cloud-based live streaming of multisourced multiview videos. Journal of Network and Computer Applications, 120(??):130–144, October 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302364.

Anonymous:2018:EBu

[2315] Anonymous. Editorial Board. Journal of Network and Computer Applications, 121(??):ii, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302856.

Saadeh:2018:HAP

[2316] Maha Saadeh, Azzam Sleit, Khair Eddin Sabri, and Wesam Almobaideen. Hierarchical architecture and protocol for mobile object authentication in the context of IoT smart cities. Journal of Network and Computer Applications, 121(??):1–19, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302339.

Qu:2018:CAS

[2317] Dapeng Qu, Xingwei Wang, Min Huang, Keqin Li, Sajal K. Das, and Sijin Wu. A cache-aware social-based QoS routing scheme in Information Centric Networks. Journal of Network and Computer Applications, 121(??):20–32, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302261.

Junior:2018:FAN

[2318] Dyson Pereira Junior and Emilio Carlos Gomes Wille. FB-APSP: a new efficient algorithm for computing all-pairs shortest-paths. Journal of Network and Computer Applications, 121(??):33–43, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302443.

Anwar:2018:ATA

[2319] Ahmed H. Anwar, George Atia, and Mina Guirguis. Adaptive topologies against jamming attacks in wireless networks: a game-theoretic approach. Journal of Network and Computer Applications, 121(??):44–58, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302169.

Chen:2018:DMT

[2320] Longbiao Chen, Dingqi Yang, Daqing Zhang, Cheng Wang, Jonathan Li, and Thi-Mai-Trang Nguyen. Deep mobile traffic forecast and complementary base station clustering for C-RAN optimization. Journal of Network and Computer Applications, 121(??):59–69, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302455.

Hamidouche:2018:CBI

[2321] Ranida Hamidouche, Zibouda Aliouat, Abdelhak Mourad Gueroui, Ado Adamou Abba Ari, and Lemia Louail. Classical and bio-inspired mobility in sensor networks for IoT applications. Journal of Network and Computer Applications, 121(??):70–88, November 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302340.

Rehan:2018:ARM

[2322] Waqas Rehan, Stefan Fischer, and Maaz Rehan. Anatomizing the robustness of multichannel MAC protocols for WSNs: an evaluation under MAC oriented design issues impacting QoS. Journal of Network and Computer Applications, 121(??):89–118, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302212.

Amarlingam:2018:NLW

[2323] M. Amarlingam, Pradeep Kumar Mishra, P. Rajalakshmi, Sumohana S. Channappayya, and C. S. Sastry. Novel light weight compressed data aggregation using sparse measurements for IoT networks. Journal of Network and Computer Applications, 121(??):119–134, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302613.

Dezfouli:2018:EEM

[2324] Behnam Dezfouli, Immanuel Amirtharaj, and Chia-Chi (Chelsey) Li. EMPIOT: an energy measurement platform for wireless IoT devices. Journal of Network and Computer Applications, 121(??):135–148, November 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302479.

Sanislav:2018:WEH

[2325] Teodora Sanislav, Sherali Zeadally, George Dan Mois, and Silviu Corneliu Folea. Wireless energy harvesting: Empirical results and practical considerations for Internet of Things. Journal of Network and Computer Applications, 121(??):149–158, November 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302595.

Anonymous:2018:EBv

[2326] Anonymous. Editorial Board. Journal of Network and Computer Applications, 122(??):ii, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303278.

Zhou:2018:DIV

[2327] Lei Zhou, Anmin Fu, Shui Yu, Mang Su, and Boyu Kuang. Data integrity verification of the outsourced big data in the cloud environment: a survey. Journal of Network and Computer Applications, 122 (??):1–15, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302601.

Lu:2018:MTT

[2328] Xu Lu, Jun Liu, Wei Qi, and Qingyun Dai. Multiple-target tracking based on compressed sensing in the Internet of Things. Journal of Network and Computer Applications, 122(??):16–23, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302637.

Singh:2018:MSD

[2329] Deepak Singh, Bryan Ng, Yuan-Cheng Lai, Ying-Dar Lin, and Winston K. G. Seah. Modelling software-defined networking: Software and hardware switches. Journal of Network and Computer Applications, 122 (??):24–36, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302625.

Zhang:2018:NOL

[2330] De gan Zhang, Ting Zhang, Yue Dong, Xiao huan Liu, Yu ya Cui, and De xin Zhao. Novel optimized link state routing protocol based on quantum genetic strategy for mobile learning. Journal of Network and Computer Applications, 122(??):37–49, November 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302492.

Zhang:2018:PPC

[2331] Yinghui Zhang, Jin Li, Dong Zheng, Ping Li, and Yangguang Tian. Privacy-preserving communication and power injection over vehicle networks and 5G smart grid slice. Journal of Network and Computer Applications, 122(??):50–60, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518302480.

Imine:2018:RAB

[2332] Youcef Imine, Ahmed Lounis, and Abdelmadjid Bouabdallah. Revocable attribute-based access control in multi-authority systems. Journal of Network and Computer Applications, 122(??):61–76, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302650.

Mattos:2018:LPC

[2333] Diogo Menezes Ferrazani Mattos, Otto Carlos Muniz Bandeira Duarte, and Guy Pujolle. A lightweight protocol for consistent policy update on software-defined networking with multiple controllers. Journal of Network and Computer Applications, 122(??):77–87, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302649.

Xu:2018:MVM

[2334] Chuan Xu, Wenqiang Jin, Xinheng Wang, Guofeng Zhao, and Shui Yu. MC-VAP: a multi-connection virtual access point for high performance software-defined wireless networks. Journal of Network and Computer Applications, 122(??):88–98, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302662.

Mota:2018:TSM

[2335] Vinícius F. S. Mota, Thiago H. Silva, Daniel F. Macedo, Yacine Ghamri-Doudane, and José M. S. Nogueira. Towards scalable mobile crowdsensing through device-to-device communication. Journal of Network and Computer Applications, 122(??):99–106, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302674.

Das:2018:IDN

[2336] Debasis Das and Rajiv Misra. Improvised dynamic network connectivity model for Vehicular Ad-Hoc Networks (VANETs). Journal of Network and Computer Applications, 122(??):107–114, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302716.

Cicirelli:2018:PSA

[2337] Franco Cicirelli, Agostino Forestiero, Andrea Giordano, and Carlo Mastroianni. Parallelization of space-aware applications: Modeling and performance analysis. Journal of Network and Computer Applications, 122 (??):115–127, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302728.

Javadi:2018:LAB

[2338] Mahmood Javadi, Habib Mostafaei, Morshed U. Chowdhurry, and Jemal H. Abawajy. Learning automaton based topology control protocol for extending wireless sensor networks lifetime. Journal of Network and Computer Applications, 122(??):128–136, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302698.

Ghebleh:2018:RNC

[2339] Reza Ghebleh and Ali Ghaffari. Retraction notice to “A comprehensive survey on 5G: the prospect of future communications” [J. Netw. Comput. Appl. 94 (2017) 1–32]. Journal of Network and Computer Applications, 122(??):137, November 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303175. See [2044].

Anonymous:2018:EBw

[2340] Anonymous. Editorial Board. Journal of Network and Computer Applications, 123(??):ii, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303448.

Zhang:2018:AAG

[2341] Qikun Zhang, Yong Gan, Lu Liu, Xianmin Wang, Xiangyang Luo, and Yuanzhang Li. An authenticated asymmetric group key agreement based on attribute encryption. Journal of Network and Computer Applications, 123(??):1–10, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302704.

Habib:2018:SRS

[2342] Md. Ahsan Habib, Sajeeb Saha, Md. Abdur Razzaque, Md. Mamun or Rashid, Giancarlo Fortino, and Mohammad Mehedi Hassan. Starfish routing for sensor networks with mobile sink. Journal of Network and Computer Applications, 123(??):11–22, December 1, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451830273X.

Shao:2018:NGR

[2343] Chenglong Shao, Heejun Roh, and Wonjun Lee. Next-generation RFpowered networks for Internet of Things: Architecture and research perspectives. Journal of Network and Computer Applications, 123 (??):23–31, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302686.

Shabut:2018:MTE

[2344] Antesar M. Shabut, M. Shamim Kaiser, Keshav P. Dahal, and Wenbing Chen. A multidimensional trust evaluation model for MANETs. Journal of Network and Computer Applications, 123(??):32–41, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302327.

Khedim:2018:CCS

[2345] Farah Khedim, Nabila Labraoui, and Ado Adamou Abba Ari. A cognitive chronometry strategy associated with a revised cloud model to deal with the dishonest recommendations attacks in wireless sensor networks. Journal of Network and Computer Applications, 123 (??):42–56, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302881.

DiPietro:2018:GKL

[2346] Roberto Di Pietro and Gabriele Oligeri. GopJam: Key-less jamming mitigation via gossiping. Journal of Network and Computer Applications, 123(??):57–68, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302583.

Sikeridis:2018:WPP

[2347] Dimitrios Sikeridis, Eirini Eleni Tsiropoulou, Michael Devetsikiotis, and Symeon Papavassiliou. Wireless powered Public Safety IoT: a UAVassisted adaptive-learning approach towards energy efficiency. Journal of Network and Computer Applications, 123(??):69–79, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451830290X.

Kassan:2018:GTB

[2348] Sara Kassan, Jaafar Gaber, and Pascal Lorenz. Game theory based distributed clustering approach to maximize wireless sensors network lifetime. Journal of Network and Computer Applications, 123(??):80–88, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302911.

Zhang:2018:SSH

[2349] Yinghui Zhang, Robert H. Deng, Gang Han, and Dong Zheng. Secure smart health with privacy-aware aggregate authentication and access control in Internet of Things. Journal of Network and Computer Applications, 123(??):89–100, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518302935.

Shu:2018:DSP

[2350] Jiangang Shu, Ximeng Liu, Yinghui Zhang, Xiaohua Jia, and Robert H. Deng. Dual-side privacy-preserving task matching for spatial crowdsourcing. Journal of Network and Computer Applications, 123(??): 101–111, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302959.

Wazid:2018:AKM

[2351] Mohammad Wazid, Ashok Kumar Das, and Athanasios V. Vasilakos. Authenticated key management protocol for cloud-assisted body area sensor networks. Journal of Network and Computer Applications, 123 (??):112–126, December 1, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302960.

Anonymous:2018:EBx

[2352] Anonymous. Editorial Board. Journal of Network and Computer Applications, 124(??):ii, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303576.

Mahjri:2018:CRA

[2353] Imen Mahjri, Amine Dhraief, Abdelfettah Belghith, Sofien Gannouni, Issam Mabrouki, and Maram AlAjlan. Collision risk assessment in Flying Ad Hoc aerial wireless networks. Journal of Network and Computer Applications, 124(??):1–13, December 15, 2018. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302984.

Morato:2018:RED

[2354] Daniel Morato, Eduardo Berrueta, Eduardo Magaña, and Mikel Izal. Ransomware early detection by the analysis of file sharing traffic. Journal of Network and Computer Applications, 124(??):14–32, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451830300X.

Shehada:2018:NAT

[2355] Dina Shehada, Chan Yeob Yeun, M. Jamal Zemerly, Mahmoud Al-Qutayri, Yousof Al-Hammadi, and Jiankun Hu. A new adaptive trust and reputation model for Mobile Agent Systems. Journal of Network and Computer Applications, 124(??):33–43, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302996.

AlZishan:2018:MHC

[2356] Abdullah Al Zishan, Imtiaz Karim, Sudipta Saha Shubha, and Ashikur Rahman. Maximizing heterogeneous coverage in over and under provisioned visual sensor networks. Journal of Network and Computer Applications, 124(??):44–62, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518302972.

Liao:2018:AAV

[2357] Qi Liao, Ting Li, and Benjamin A. Blakely. Anomaly analysis and visualization for dynamic networks through spatiotemporal graph segmentations. Journal of Network and Computer Applications, 124(??):63–79, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303035.

Ricart-Sanchez:2018:TFA

[2358] Ruben Ricart-Sanchez, Pedro Malagon, Pablo Salva-Garcia, Enrique Chirivella Perez, Qi Wang, and Jose M. Alcaraz Calero. Towards an FPGA-Accelerated programmable data path for edge-to-core communications in 5G networks. Journal of Network and Computer Applications, 124(??):80–93, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518302923.

Iqbal:2018:DWP

[2359] Waheed Iqbal, Abdelkarim Erradi, and Arif Mahmood. Dynamic workload patterns prediction for proactive auto-scaling of web applications. Journal of Network and Computer Applications, 124(??):94–107, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303102.

Zhang:2018:MSE

[2360] Jianwei Zhang, Xinchang Zhang, Meng Sun, and Chunling Yang. Maximizing streaming efficiency of multiple streams in peer-to-peer networks. Journal of Network and Computer Applications, 124(??):108–120, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303084.

Hejja:2018:OPA

[2361] Khaled Hejja and Xavier Hesselbach. Online power aware coordinated virtual network embedding with 5G delay constraint. Journal of Network and Computer Applications, 124(??):121–136, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303151.

Utsumi:2018:NAM

[2362] Satoshi Utsumi, Salahuddin Muhammad Salim Zabir, Yuto Usuki, Seisho Takeda, Norio Shiratori, Yasushi Kato, and Jeyeon Kim. A new analytical model of TCP Hybla for satellite IP networks. Journal of Network and Computer Applications, 124(??):137–147, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303023.

Xu:2018:IOD

[2363] Xiaolong Xu, Shucun Fu, Lianyong Qi, Xuyun Zhang, Qingxiang Liu, Qiang He, and Shancang Li. An IoT-Oriented data placement method with privacy preservation in cloud environment. Journal of Network and Computer Applications, 124(??):148–157, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518302947.

Chen:2018:CBS

[2364] Shuyue Chen and Chin-Tau Lea. Constraint-based scheduling algorithm with the non-adjacency requirement for multi-flow AWG switches. Journal of Network and Computer Applications, 124(??):158–168, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303096.

Kumari:2018:MBD

[2365] Aparna Kumari, Sudeep Tanwar, Sudhanshu Tyagi, Neeraj Kumar, Michele Maasberg, and Kim-Kwang Raymond Choo. Multimedia big data computing and Internet of Things applications: a taxonomy and process model. Journal of Network and Computer Applications, 124 (??):169–195, December 15, 2018. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303011.

Anonymous:2019:EBa

[2366] Anonymous. Editorial Board. Journal of Network and Computer Applications, 125(??):ii, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303734.

Anonymous:2019:TYR

[2367] Anonymous. Thank you reviewers. Journal of Network and Computer Applications, 125(??):iii–xii, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518303771.

Val:2019:DAI

[2368] I. Val, A. Etxabe, R. Torrego, P. M. Rodriguez, C. Cruces, V. Diez, M. Mendicute, and A. Arriola. Design, analysis and implementation of a time-bounded spectrum handoff algorithm for real-time industrial wireless sensor and actuator networks. Journal of Network and Computer Applications, 125(??):1–16, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303114.

Kharb:2019:FBP

[2369] Seema Kharb and Anita Singhrova. Fuzzy based priority aware scheduling technique for dense industrial IoT networks. Journal of Network and Computer Applications, 125(??):17–27, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303138.

Indu:2019:NIA

[2370] V. Indu and Sabu M. Thampi. A nature — inspired approach based on Forest Fire model for modeling rumor propagation in social networks. Journal of Network and Computer Applications, 125(??):28–41, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451830314X.

Kala:2019:SIC

[2371] Srikant Manas Kala, Vanlin Sathya, M. Pavan Kumar Reddy, Betty Lala, and Bheemarjuna Reddy Tamma. A socio-inspired CALM approach to channel assignment performance prediction and WMN capacity estimation. Journal of Network and Computer Applications, 125(??):42–66, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303126.

Sharma:2019:MSM

[2372] Vishal Sharma, Jianfeng Guan, Jiyoon Kim, Soonhyun Kwon, Ilsun You, Francesco Palmieri, and Mario Collotta. MIH-SPFP: MIH-base secure cross-layer handover protocol for Fast Proxy Mobile IPv6-IoT networks. Journal of Network and Computer Applications, 125(??):67–81, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518302893.

Guan:2019:AAP

[2373] Zhitao Guan, Yue Zhang, Longfei Wu, Jun Wu, Jing Li, Yinglong Ma, and Jingjing Hu. APPA: an anonymous and privacy preserving data aggregation scheme for fog-enhanced IoT. Journal of Network and Computer Applications, 125(??):82–92, January 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303060.

Jiang:2019:SLP

[2374] Jinfang Jiang, Guangjie Han, Hao Wang, and Mohsen Guizani. A survey on location privacy protection in Wireless Sensor Networks. Journal of Network and Computer Applications, 125(??):93–114, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303308.

Rangisetti:2019:IQA

[2375] Anil Kumar Rangisetti and Bheemarjuna Reddy Tamma. Interference and QoS aware cell switch-off strategy for software defined LTE Het-Nets. Journal of Network and Computer Applications, 125(??):115–129, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303163.

Bu:2019:RSR

[2376] Chao Bu, Xingwei Wang, Hui Cheng, Min Huang, and Keqin Li. Routing as a service (RaaS): an open framework for customizing routing services. Journal of Network and Computer Applications, 125(??):130–145, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303321.

Bereketli:2019:PAP

[2377] Alper Bereketli, Muharrem Tümçakir, and Burcu Yeni. P-AUV: Position aware routing and medium access for ad hoc AUV networks. Journal of Network and Computer Applications, 125(??):146–154, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303369.

Kolomvatsos:2019:TOM

[2378] Kostas Kolomvatsos. Time-optimized management of mobile IoT nodes for pervasive applications. Journal of Network and Computer Applications, 125(??):155–167, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303382.

Liao:2019:SCS

[2379] Zhuhua Liao, Zengde Teng, Jian Zhang, Yizhi Liu, Hao Xiao, and Aiping Yi. A Semantic Concast service for data discovery, aggregation and processing on NDN. Journal of Network and Computer Applications, 125(??):168–178, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303394.

Chopra:2019:RRB

[2380] Garima Chopra, Rakesh Kumar Jha, and Sanjeev Jain. RBA: Region based algorithm for secure harvesting in ultra dense network. Journal of Network and Computer Applications, 125(??):179–189, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303072.

Liu:2019:MAT

[2381] Zhen Liu, Ruoyu Wang, Nathalie Japkowicz, Yongming Cai, Deyu Tang, and Xianfa Cai. Mobile app traffic flow feature extraction and selection for improving classification robustness. Journal of Network and Computer Applications, 125(??):190–208, January 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303400.

Deng:2019:VMS

[2382] Shuhua Deng, Xing Gao, Zebin Lu, Zhengfa Li, and Xieping Gao. DoS vulnerabilities and mitigation strategies in software-defined networks. Journal of Network and Computer Applications, 125(??):209–219, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303333.

Qiao:2019:SPD

[2383] Xiuquan Qiao, Pei Ren, Junliang Chen, Wei Tan, M. Brian Blake, and Wangli Xu. Session persistence for dynamic web applications in Named Data Networking. Journal of Network and Computer Applications, 125 (??):220–235, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303370.

Han:2019:MSP

[2384] Weijie Han, Jingfeng Xue, Yong Wang, Zhenyan Liu, and Zixiao Kong. MalInsight: a systematic profiling based malware detection framework. Journal of Network and Computer Applications, 125(??):236–250, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303503.

Makhdoom:2019:BAI

[2385] Imran Makhdoom, Mehran Abolhasan, Haider Abbas, and Wei Ni. Blockchain’s adoption in IoT: the challenges, and a way forward.

Journal of Network and Computer Applications, 125(??):251–279, January 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303473.

Anonymous:2019:EBb

[2386] Anonymous. Editorial Board. Journal of Network and Computer Applications, 126(??):ii, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303941.

Tao:2019:LBT

[2387] Ming Tao, Wenhong Wei, and Shuqiang Huang. Location-based trustworthy services recommendation in cooperative-communication-enabled Internet of Vehicles. Journal of Network and Computer Applications, 126(??):1–11, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303515.

Aziz:2019:EAO

[2388] Ahmed Aziz, Karan Singh, Walid Osamy, and Ahmed M. Khedr. Effective algorithm for optimizing compressive sensing in IoT and periodic monitoring applications. Journal of Network and Computer Applications, 126(??):12–28, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303357.

Zhang:2019:PRC

[2389] Xiaosong Zhang, Liehuang Zhu, Xianmin Wang, Changyou Zhang, Hongfei Zhu, and Yu an Tan. A packet-reordering covert channel over VoLTE voice and video traffics. Journal of Network and Computer Applications, 126(??):29–38, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303527.

Li:2019:PVP

[2390] Tong Li, Chongzhi Gao, Liaoliang Jiang, Witold Pedrycz, and Jian Shen. Publicly verifiable privacy-preserving aggregation and its application in IoT. Journal of Network and Computer Applications, 126 (??):39–44, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303059.

Feng:2019:SPP

[2391] Qi Feng, Debiao He, Sherali Zeadally, Muhammad Khurram Khan, and Neeraj Kumar. A survey on privacy protection in blockchain system. Journal of Network and Computer Applications, 126(??):45–58, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303485.

Kharb:2019:SNF

[2392] Seema Kharb and Anita Singhrova. A survey on network formation and scheduling algorithms for time slotted channel hopping in industrial networks. Journal of Network and Computer Applications, 126 (??):59–87, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303631.

Lin:2019:ASR

[2393] Huaqing Lin, Zheng Yan, and Yulong Fu. Adaptive security-related data collection with context awareness. Journal of Network and Computer Applications, 126(??):88–103, January 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303539.

Mazied:2019:WCP

[2394] EmadelDin A. Mazied, Mustafa Y. ElNainay, Mohammad J. Abdel-Rahman, Scott F. Midkiff, Mohamed R. M. Rizk, Hesham A. Rakha, and Allen B. MacKenzie. The wireless control plane: an overview and directions for future research. Journal of Network and Computer Applications, 126(??):104–122, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303047.

Akilal:2019:RTI

[2395] Karim Akilal, Hachem Slimani, and Mawloud Omar. A robust trust inference algorithm in weighted signed social networks based on collaborative filtering and agreement as a similarity metric. Journal of Network and Computer Applications, 126(??):123–132, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303679.

Vicentini:2019:SBM

[2396] Cleverton Vicentini, Altair Santin, Eduardo Viegas, and Vilmar Abreu. SDN-based and multitenant-aware resource provisioning mechanism for cloud-based big data streaming. Journal of Network and Computer Applications, 126(??):133–149, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804518303643.

Seba:2019:RSC

[2397] Abderazek Seba, Nadia Nouali-Taboudjemat, Nadjib Badache, and Hamida Seba. A review on security challenges of wireless communications in disaster emergency response and crisis management situations. Journal of Network and Computer Applications, 126(??):150–161, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303692.

Shabestari:2019:TSB

[2398] Fatemeh Shabestari, Amir Masoud Rahmani, Nima Jafari Navimipour, and Sam Jabbehdari. A taxonomy of software-based and hardwarebased approaches for energy efficiency management in the Hadoop. Journal of Network and Computer Applications, 126(??):162–177, January 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303667.

Anonymous:2019:EBc

[2399] Anonymous. Editorial Board. Journal of Network and Computer Applications, 127(??):ii, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518304065.

Siregar:2019:BAA

[2400] Rahmat Faddli Siregar, Fahri Wisnu Murti, and Soo Young Shin. Bit allocation approach of spatial modulation for multi-user scenario. Journal of Network and Computer Applications, 127(??):1–8, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303801.

Ghalem:2019:PMC

[2401] Sanaa Kawther Ghalem, Bouabdellah Kechar, Ahcène Bounceur, and Reinhardt Euler. A probabilistic multivariate copula-based technique for faulty node diagnosis in wireless sensor networks. Journal of Network and Computer Applications, 127(??):9–25, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804518303680.

Wilhelmi:2019:PPM

[2402] Francesc Wilhelmi, Sergio Barrachina-Muñoz, Boris Bellalta, Cristina Cano, Anders Jonsson, and Gergely Neu. Potential and pitfalls of Multi-Armed Bandits for decentralized Spatial Reuse in WLANs. Journal of Network and Computer Applications, 127(??):26–42, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303655.

Wang:2019:CPB

[2403] Licheng Wang, Xiaoying Shen, Jing Li, Jun Shao, and Yixian Yang. Cryptographic primitives in blockchains. Journal of Network and Computer Applications, 127(??):43–58, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451830362X.

Tian:2019:PPP

[2404] Hui Tian, Fulin Nan, Chin-Chen Chang, Yongfeng Huang, Jing Lu, and Yongqian Du. Privacy-preserving public auditing for secure data storage in fog-to-cloud computing. Journal of Network and Computer Applications, 127(??):59–69, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303862.

Dagdeviren:2019:DEA

[2405] Orhan Dagdeviren, Vahid Khalilpour Akram, and Ali Farzan. A distributed evolutionary algorithm for detecting minimum vertex cuts for wireless ad hoc and sensor networks. Journal of Network and Computer Applications, 127(??):70–81, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451830331X.

Ahuja:2019:NSC

[2406] Kiran Ahuja and Arun Khosla. Network selection criterion for ubiquitous communication provisioning in smart cities for smart energy system. Journal of Network and Computer Applications, 127(??):82–91, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303783.

Yang:2019:IAA

[2407] Li Yang, Yifang Zhi, Teng Wei, Shui Yu, and Jianfeng Ma. Inference attack in Android activity based on program fingerprint. Journal of Network and Computer Applications, 127(??):92–106, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303898.

Tuysuz:2019:RTP

[2408] Mehmet Fatih Tuysuz, Murat Ucan, and Ramona Trestian. A real-time power monitoring and energy-efficient network/interface selection tool for Android smartphones. Journal of Network and Computer Applications, 127(??):107–121, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303813.

Sun:2019:CFP

[2409] Xiuwen Sun, Hao Li, Dan Zhao, Xingxing Lu, Kaiyu Hou, and Chengchen Hu. COIN: a fast packet inspection method over compressed traffic. Journal of Network and Computer Applications, 127(??):122–134, February 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303904. See corrigendum [2661].

Anonymous:2019:EBd

[2410] Anonymous. Editorial Board. Journal of Network and Computer Applications, 128(??):ii, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519300128.

Belouanas:2019:MAO

[2411] Salah-Eddine Belouanas, Kim-Loan Thai, Prométhée Spathis, and Marcelo Dias de Amorim. Mobility-assisted offloading in centrallycoordinated cellular networks. Journal of Network and Computer Applications, 128(??):1–10, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303795.

Xu:2019:TPT

[2412] Lingling Xu, Jin Li, Xiaofeng Chen, Wanhua Li, Shaohua Tang, and Hao-Tian Wu. Tc-PEDCKS: Towards time controlled public key encryption with delegatable conjunctive keyword search for Internet of Things. Journal of Network and Computer Applications, 128(??):11–20, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303850.

Yang:2019:SCC

[2413] Haomiao Yang, Xiaofen Wang, Chun Yang, Xin Cong, and You Zhang. Securing content-centric networks with content-based encryption. Journal of Network and Computer Applications, 128(??):21–32, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303874.

Moustafa:2019:HRN

[2414] Nour Moustafa, Jiankun Hu, and Jill Slay. A holistic review of Network Anomaly Detection Systems: a comprehensive survey. Journal of Network and Computer Applications, 128(??):33–55, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303886.

Li:2019:DMV

[2415] Wenjuan Li, Weizhi Meng, Zhiyuan Tan, and Yang Xiang. Design of multi-view based email classification for IoT systems via semi-supervised learning. Journal of Network and Computer Applications, 128(??):56–63, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303849.

Adhikari:2019:MHB

[2416] Mainak Adhikari, Sudarshan Nandy, and Tarachand Amgoth. Meta heuristic-based task deployment mechanism for load balancing in IaaS cloud. Journal of Network and Computer Applications, 128(??):64–77, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303989.

Oliveira:2019:RSC

[2417] Helder M. N. S. Oliveira and Nelson L. S. da Fonseca. Routing, spectrum and core assignment algorithms for protection of space division multiplexing elastic optical networks. Journal of Network and Computer Applications, 128(??):78–89, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303977.

Kumari:2019:FDA

[2418] Aparna Kumari, Sudeep Tanwar, Sudhanshu Tyagi, Neeraj Kumar, Reza M. Parizi, and Kim-Kwang Raymond Choo. Fog data analytics: a taxonomy and process model. Journal of Network and Computer Applications, 128(??):90–104, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518304016.

Elazhary:2019:ITI

[2419] Hanan Elazhary. Internet of Things (IoT), mobile cloud, cloudlet, mobile IoT, IoT cloud, fog, mobile edge, and edge emerging computing paradigms: Disambiguation and research directions. Journal of Network and Computer Applications, 128(??):105–140, February 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303497.

Anonymous:2019:EBe

[2420] Anonymous. Editorial Board. Journal of Network and Computer Applications, 129(??):ii, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300360.

Zhang:2019:SCM

[2421] Qingchen Zhang, Changchuan Bai, Zhikui Chen, Peng Li, Shuo Wang, and He Gao. Smart Chinese medicine for hypertension treatment with a deep learning model. Journal of Network and Computer Applications, 129(??):1–8, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518304004.

Kaisar:2019:DCD

[2422] Shahriar Kaisar, Joarder Kamruzzaman, and Gour Karmakar. A dynamic content distribution scheme for decentralized sharing in tourist hotspots. Journal of Network and Computer Applications, 129(??):9– 24, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303990.

Sandor:2019:EDM

[2423] Voundi Koe Arthur Sandor, Yaping Lin, Xiehua Li, Feng Lin, and Shiwen Zhang. Efficient decentralized multi-authority attribute based encryption for mobile cloud data storage. Journal of Network and Computer Applications, 129(??):25–36, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300037.

Chen:2019:LAC

[2424] Siguang Chen, Shujun Zhang, Xiaoyao Zheng, and Xiukai Ruan. Layered adaptive compression design for efficient data collection in industrial wireless sensor networks. Journal of Network and Computer Applications, 129(??):37–45, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300025.

Shao:2019:CER

[2425] Yanling Shao, Chunlin Li, Zhao Fu, Leyue Jia, and Youlong Luo. Cost-effective replication management and scheduling in edge computing. Journal of Network and Computer Applications, 129(??):46–61, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300013.

Zhao:2019:USP

[2426] Haichun Zhao, Xuanxia Yao, Xuefeng Zheng, Tie Qiu, and Huansheng Ning. User stateless privacy-preserving TPA auditing scheme for cloud storage. Journal of Network and Computer Applications, 129(??):62–70, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300050.

Pedhadiya:2019:DDC

[2427] Mittal K. Pedhadiya, Rakesh Kumar Jha, and Hetal G. Bhatt. Device to device communication: a survey. Journal of Network and Computer Applications, 129(??):71–89, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804518303345.

Anonymous:2019:EBf

[2428] Anonymous. Editorial Board. Journal of Network and Computer Applications, 130(??):ii, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930061X.

Liu:2019:EDA

[2429] Gang Liu, Wei Quan, Nan Cheng, Hongke Zhang, and Shui Yu. Efficient DDoS attacks mitigation for stateful forwarding in Internet of Things. Journal of Network and Computer Applications, 130(??):1–13, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300062.

Ahmed:2019:TAB

[2430] Firas D. Ahmed and Mazlina Abdul Majid. Towards agent-based Petri net decision making modelling for cloud service composition: a literature survey. Journal of Network and Computer Applications, 130 (??):14–38, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518303837.

Ben-Ammar:2019:PAD

[2431] Hamza Ben-Ammar, Yassine Hadjadj-Aoul, Gerardo Rubino, and Soraya Ait-Chellouche. On the performance analysis of distributed caching systems using a customizable Markov chain model. Journal of Network and Computer Applications, 130(??):39–51, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300190.

Abououf:2019:MWM

[2432] Menatalla Abououf, Rabeb Mizouni, Shakti Singh, Hadi Otrok, and Anis Ouali. Multi-worker multi-task selection framework in mobile crowd sourcing. Journal of Network and Computer Applications, 130 (??):52–62, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300086.

Imputato:2019:EFN

[2433] Pasquale Imputato and Stefano Avallone. Enhancing the fidelity of network emulation through direct access to device buffers. Journal of Network and Computer Applications, 130(??):63–75, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519300220.

Lai:2019:PMA

[2434] Yuan-Cheng Lai, Ahsan Ali, Md. Shohrab Hossain, and Ying-Dar Lin. Performance modeling and analysis of TCP and UDP flows over software defined networks. Journal of Network and Computer Applications, 130(??):76–88, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300189.

Verma:2019:NCI

[2435] Rahul Kumar Verma, K. K. Pattanaik, Sourabh Bharti, and Divya Saxena. In-network context inference in IoT sensory environment for efficient network resource utilization. Journal of Network and Computer Applications, 130(??):89–103, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519300219.

Celesti:2019:ENC

[2436] Antonio Celesti, Antoine Bagula, Ivanoe De Falco, Pedro Brandão, and Giovanna Sannino. Emerging networked computer applications for telemedicine. Journal of Network and Computer Applications, 130 (??):104–106, March 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300323.

Anonymous:2019:EBg

[2437] Anonymous. Editorial Board. Journal of Network and Computer Applications, 131(??):ii, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930075X.

Li:2019:DMU

[2438] Chunlin Li, Jianhang Tang, and Youlong Luo. Dynamic multiuser computation offloading for wireless powered mobile edge computing. Journal of Network and Computer Applications, 131(??):1–15, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300281.

Liu:2019:DDD

[2439] Yuxin Liu, Anfeng Liu, Ning Zhang, Xiao Liu, Ming Ma, and Yanling Hu. DDC: Dynamic duty cycle for improving delay and energy efficiency in wireless sensor networks. Journal of Network and Computer Applications, 131(??):16–27, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300293.

Queiroz:2019:AST

[2440] Wander Queiroz, Miriam A. M. Capretz, and Mario Dantas. An approach for SDN traffic monitoring based on big data techniques. Journal of Network and Computer Applications, 131(??):28–39, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300244.

Chakraborty:2019:NNP

[2441] Tusher Chakraborty, Novia Nurain, Suraiya Tairin, Taslim Arefin Khan, Jannatun Noor, Md. Rezwanul Islam, and A. B. M. Alim Al Islam. A new network paradigm for low-cost and lightweight real-time communication between train and rail track to detect missing and faulty rail blocks. Journal of Network and Computer Applications, 131(??):40– 54, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300268.

Zhao:2019:TDT

[2442] Minghao Zhao, Chengyu Hu, Xiangfu Song, and Chuan Zhao. Towards dependable and trustworthy outsourced computing: a comprehensive survey and tutorial. Journal of Network and Computer Applications, 131(??):55–65, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930030X.

Ying:2019:LRU

[2443] Bidi Ying and Amiya Nayak. Lightweight remote user authentication protocol for multi-server 5G networks using self-certified public key cryptography. Journal of Network and Computer Applications, 131(??):66–74, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300256.

Huang:2019:RDF

[2444] Jiawei Huang, Shuping Li, Rui Han, and Jianxin Wang. Receiverdriven fair congestion control for TCP outcast in data center networks. Journal of Network and Computer Applications, 131(??):75– 88, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300311.

Sahay:2019:ASD

[2445] Rishikesh Sahay, Weizhi Meng, and Christian D. Jensen. The application of Software Defined Networking on securing computer networks: a survey. Journal of Network and Computer Applications, 131(??):89–108, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930027X.

Farkhana:2019:ECC

[2446] Muchtar Farkhana, Abdullah Abdul Hanan, Hassan Suhaidi, Khader Ahamad Tajudin, and Zamli Kamal Zuhairi. Energy conservation of content routing through wireless broadcast control in NDN based MANET: a review. Journal of Network and Computer Applications, 131(??):109– 132, April 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300049.

Anonymous:2019:EBh

[2447] Anonymous. Editorial Board. Journal of Network and Computer Applications, 132(??):ii, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300979.

Han:2019:PCC

[2448] Jinyoung Han, Daejin Choi, Taejoong Chung, Chen-Nee Chuah, Hyun chul Kim, and Ted Taekyoung Kwon. Predicting content consumption from content-to-content relationships. Journal of Network and Computer Applications, 132(??):1–9, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451930044X.

Thulasiraman:2019:CPC

[2449] P. Thulasiraman, T. Haakensen, and A. Callanan. Countering passive cyber attacks against sink nodes in tactical sensor networks using reactive route obfuscation. Journal of Network and Computer Applications, 132(??):10–21, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300426.

Gomes:2019:ABB

[2450] Ruan D. Gomes, Cesar Benavente-Peces, Iguatemi E. Fonseca, and Marcelo S. Alencar. Adaptive and Beacon-based multi-channel protocol for Industrial Wireless Sensor Networks. Journal of Network and Computer Applications, 132(??):22–39, April 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300396.

Han:2019:PDM

[2451] Wenyuan Han, Weijiang Liu, Kaiye Zhang, Zhiyang Li, and Zhaobin Liu. A protocol for detecting missing target tags in RFID systems. Journal of Network and Computer Applications, 132(??):40–48, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300414.

Go:2019:SBF

[2452] Sharleen Joy Y. Go, Cedric Angelo M. Festin, and Wilson M. Tan. An SDN-based framework for improving the performance of underprovisioned IP Video Surveillance networks. Journal of Network and Computer Applications, 132(??):49–74, April 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300402.

Dong:2019:FOI

[2453] Shi Dong and Raj Jain. Flow online identification method for the encrypted Skype. Journal of Network and Computer Applications, 132(??):75–85, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300074.

Kong:2019:ASN

[2454] Xiangjie Kong, Yajie Shi, Shuo Yu, Jiaying Liu, and Feng Xia. Academic social networks: Modeling analysis, mining and applications. Journal of Network and Computer Applications, 132(??):86–103, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300438.

Ashraf:2019:EMH

[2455] Nouman Ashraf, Muhammad Faizan, Waqar Asif, Hassaan Khaliq Qureshi, Adnan Iqbal, and Marios Lestas. Energy management in harvesting enabled sensing nodes: Prediction and control. Journal of Network and Computer Applications, 132(??):104–117, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519300232.

Gupta:2019:RPE

[2456] Abhishek Gupta, Alagan Anpalagan, Glaucio H. S. Carvalho, Ahmed S. Khwaja, Ling Guan, and Isaac Woungang. RETRACTED: Prevailing and emerging cyber threats and security practices in IoT-Enabled smart grids: a survey. Journal of Network and Computer Applications, 132(??):118–148, April 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300207. See [2516].

Anonymous:2019:EBi

[2457] Anonymous. Editorial Board. Journal of Network and Computer Applications, 133(??):ii, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301067.

Jabbarifar:2019:SNA

[2458] Masoume Jabbarifar, Alireza Shameli-Sendi, and Bettina Kemme. A scalable network-aware framework for cloud monitoring orchestration. Journal of Network and Computer Applications, 133(??):1–14, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300505.

Wang:2019:MMD

[2459] Shanshan Wang, Zhenxiang Chen, Qiben Yan, Bo Yang, Lizhi Peng, and Zhongtian Jia. A mobile malware detection method using behavior features in network traffic. Journal of Network and Computer Applications, 133(??):15–25, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804518304028.

Mahmud:2019:ECM

[2460] Md. Tareq Mahmud, Md. Obaidur Rahman, Mohammad Mehedi Hassan, Ahmad Almogren, and Mengchu Zhou. An efficient cooperative medium access control protocol for wireless IoT networks in smart world system. Journal of Network and Computer Applications, 133(??):26– 38, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300554.

Tulu:2019:VNE

[2461] Muluneh Mekonnen Tulu, Ronghui Hou, and Talha Younas. Vital nodes extracting method based on user’s behavior in 5G mobile social networks. Journal of Network and Computer Applications, 133(??):39– 50, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300566.

Guo:2019:LTE

[2462] Kehua Guo, Rui Cao, Xiaoyan Kui, Jianhua Ma, Jian Kang, and Tao Chi. LCC: Towards efficient label completion and correction for supervised medical image learning in smart diagnosis. Journal of Network and Computer Applications, 133(??):51–59, May 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300530.

Abdullahi:2019:ESO

[2463] Mohammed Abdullahi, Md Asri Ngadi, Salihu Idi Dishing, Shafi’i Muhammad Abdulhamid, and Barroon Isma’eel Ahmad. An efficient symbiotic organisms search algorithm with chaotic optimization strategy for multi-objective task scheduling problems in cloud computing environment. Journal of Network and Computer Applications, 133(??): 60–74, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300499.

Xu:2019:EAC

[2464] Xiaolong Xu, Yuancheng Li, Tao Huang, Yuan Xue, Kai Peng, Lianyong Qi, and Wanchun Dou. An energy-aware computation offloading method for smart edge computing in wireless metropolitan area networks. Journal of Network and Computer Applications, 133(??):75– 85, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300529.

Barreto:2019:CAM

[2465] Alexandre B. Barreto and Paulo C. G. Costa. Cyber-ARGUS — a mission assurance framework. Journal of Network and Computer Applications, 133(??):86–108, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300451.

Li:2019:DIE

[2466] Jiawei Li, Hongbin Luo, Shan Zhang, Hongyi Li, and Fei Yan. Design and implementation of efficient control for incoming inter-domain traffic with Information-Centric Networking. Journal of Network and Computer Applications, 133(??):109–125, May 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519300578.

Anonymous:2019:EBj

[2467] Anonymous. Editorial Board. Journal of Network and Computer Applications, 134(??):ii, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301249.

Wang:2019:OTC

[2468] Tong Wang and Leszek Lilien. Opportunistic topology control for ad hoc wireless network survivability enhancement based on LIMOS model. Journal of Network and Computer Applications, 134(??):1–12, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300542.

Ma:2019:RRC

[2469] Lichuan Ma, Qingqi Pei, Yong Xiang, Lina Yao, and Shui Yu. A reliable reputation computation framework for online items in Ecommerce. Journal of Network and Computer Applications, 134(??): 13–25, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300463.

Hou:2019:CSA

[2470] Huiying Hou, Jia Yu, and Rong Hao. Cloud storage auditing with deduplication supporting different security levels according to data popularity. Journal of Network and Computer Applications, 134(??):26– 39, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300669.

Guo:2019:ERA

[2471] Kai Guo, Mingcong Yang, Yongbing Zhang, and Xiaohua Jia. Efficient resource assignment in mobile edge computing: a dynamic congestionaware offloading approach. Journal of Network and Computer Applications, 134(??):40–51, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300682.

Li:2019:APA

[2472] JiLiang Li, WeiGuo Zhang, Vivek Dabra, Kim-Kwang Raymond Choo, Saru Kumari, and Dieter Hogrefe. AEP-PPA: an anonymous, efficient and provably-secure privacy-preserving authentication protocol for mobile services in smart cities. Journal of Network and Computer Applications, 134(??):52–61, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300475.

Liu:2019:DSD

[2473] Yongxin Liu, Jianqiang Li, Zhong Ming, Houbing Song, Xiaoxiong Weng, and Jian Wang. Domain-specific data mining for residents’ transit pattern retrieval from incomplete information. Journal of Network and Computer Applications, 134(??):62–71, May 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300670.

Peng:2019:EDI

[2474] Su Peng, Fucai Zhou, Jin Li, Qiang Wang, and Zifeng Xu. Efficient, dynamic and identity-based Remote Data Integrity Checking for multiple replicas. Journal of Network and Computer Applications, 134(??):72–88, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300657.

Sun:2019:SPP

[2475] Gang Sun, Siyu Sun, Jian Sun, Hongfang Yu, Xiaojiang Du, and Mohsen Guizani. Security and privacy preservation in fog-based crowd sensing on the Internet of Vehicles. Journal of Network and Computer Applications, 134(??):89–99, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300694.

Zamanifar:2019:RAP

[2476] Azadeh Zamanifar and Eslam Nazemi. RETRACTED: An approach for predicting health status in IoT health care. Journal of Network and Computer Applications, 134(??):100–113, May 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804519300888. See

retraction notice [2591].

Anonymous:2019:EBk

[2477] Anonymous. Editorial Board. Journal of Network and Computer Applications, 135(??):ii, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301316.

Zhang:2019:RRC

[2478] Xuxin Zhang, Honglong Chen, Kai Lin, Zhibo Wang, Jiguo Yu, and Leyi Shi. RMTS: a robust clock synchronization scheme for wireless sensor networks. Journal of Network and Computer Applications, 135(??): 1–10, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300876.

Wang:2019:CRM

[2479] Tian Wang, Yuzhu Liang, Weijia Jia, Muhammad Arif, Anfeng Liu, and Mande Xie. Coupling resource management based on fog computing in smart city systems. Journal of Network and Computer Applications, 135(??):11–19, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300803.

Khan:2019:EED

[2480] Muhammad Toaha Raza Khan, Syed Hassan Ahmed, Yalew Zelalem Jembre, and Dongkyun Kim. An energy-efficient data collection protocol with AUV path planning in the Internet of Underwater Things. Journal of Network and Computer Applications, 135(??):20– 31, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300840.

DeLaTorreParra:2019:IDP

[2481] Gonzalo De La Torre Parra, Paul Rad, and Kim-Kwang Raymond Choo. Implementation of deep packet inspection in smart grids and industrial Internet of Things: Challenge’s and opportunities. Journal of Network and Computer Applications, 135(??):32–46, June 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300815.

Guo:2019:JOT

[2482] Yingya Guo, Zhiliang Wang, Han Zhang, Xia Yin, Xingang Shi, and Jianping Wu. Joint optimization of tasks placement and routing to minimize Coflow Completion Time. Journal of Network and Computer Applications, 135(??):47–61, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300906.

McGhin:2019:BHA

[2483] Thomas McGhin, Kim-Kwang Raymond Choo, Charles Zhechao Liu, and Debiao He. Blockchain in healthcare applications: Research challenges and opportunities. Journal of Network and Computer Applications, 135(??):62–75, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300864.

Elsaeidy:2019:IDS

[2484] Asmaa Elsaeidy, Kumudu S. Munasinghe, Dharmendra Sharma, and Abbas Jamalipour. Intrusion detection in smart cities using Restricted Boltzmann Machines. Journal of Network and Computer Applications, 135(??):76–83, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300852.

Vales:2019:EAA

[2485] Rúben Vales, José Moura, and Rui Marinheiro. Energy-aware and adaptive fog storage mechanism with data replication ruled by spatiotemporal content popularity. Journal of Network and Computer Applications, 135(??):84–96, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930092X.

Habbal:2019:CAR

[2486] Adib Habbal, Swetha Indudhar Goudar, and Suhaidi Hassan. A Contextaware Radio Access Technology selection mechanism in 5G mobile network for smart city applications. Journal of Network and Computer Applications, 135(??):97–107, June 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519300700.

Pau:2019:TBI

[2487] Giovanni Pau, Alessandro Bazzi, Miguel Elias M. Campista, and Ali Balador. Towards 5G and beyond for the Internet of UAVs, vehicles, smartphones, Sensors and Smart Objects. Journal of Network and Computer Applications, 135(??):108–109, June 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301006.

Anonymous:2019:EBl

[2488] Anonymous. Editorial Board. Journal of Network and Computer Applications, 136(??):ii, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930150X.

Jin:2019:HEC

[2489] Xin Jin, Fei Gu, Jianwei Niu, Shui Yu, and Zhenchao Ouyang. HRCal: an effective calibration system for heart rate detection during exercising. Journal of Network and Computer Applications, 136(??):1–10, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301122.

Wang:2019:IML

[2490] Feng Wang, Wenjun Jiang, Guojun Wang, and Song Guo. Influence maximization by leveraging the crowdsensing data in information diffusion network. Journal of Network and Computer Applications, 136(??):11–21, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300931.

Singh:2019:AMS

[2491] Deepak Singh, Bryan Ng, Yuan-Cheng Lai, Ying-Dar Lin, and Winston K. G. Seah. Analytical modelling of software and hardware switches with internal buffer in software-defined networks. Journal of Network and Computer Applications, 136(??):22–37, June 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301110.

Guimaraes:2019:EIA

[2492] Carlos Guimarães, José Quevedo, Rui Ferreira, Daniel Corujo, and Rui L. Aguiar. Exploring interoperability assessment for future Internet architecture’s roll out. Journal of Network and Computer Applications, 136(??):38–56, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301201.

Wu:2019:CMP

[2493] Xiaoping Wu, Xuefen Zhu, Shenghui Wang, and Lufeng Mo. Cooperative motion parameter estimation using RSS measurements in robotic sensor networks. Journal of Network and Computer Applications, 136(??):57–70, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301195.

Birkinshaw:2019:IID

[2494] Celyn Birkinshaw, Elpida Rouka, and Vassilios G. Vassilakis. Implementing an intrusion detection and prevention system using software-defined networking: Defending against port-scanning and denial-of-service attacks. Journal of Network and Computer Applications, 136(??):71–85, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301109.

Bilal:2019:CJC

[2495] Kashif Bilal, Emna Baccour, Aiman Erbad, Amr Mohamed, and Mohsen Guizani. Collaborative joint caching and transcoding in mobile edge networks. Journal of Network and Computer Applications, 136(??):86– 99, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300487.

Patgiri:2019:RDB

[2496] Ripon Patgiri, Sabuzima Nayak, and Samir Kumar Borgohain. rDBF: a r-dimensional Bloom filter for massive scale membership query. Journal of Network and Computer Applications, 136(??):100–113, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519301092.

Monge:2019:TFA

[2497] Marco Antonio Sotelo Monge, Andrés Herranz González, Borja Lorenzo Fernández, Diego Maestre Vidal, Guillermo Rius García, and Jorge Maestre Vidal. Traffic-flow analysis for source-side DDoS recognition on 5G environments. Journal of Network and Computer Applications, 136(??):114– 131, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451930089X.

Karakus:2019:EFA

[2498] Murat Karakus and Arjan Durresi. An economic framework for analysis of network architectures: SDN and MPLS cases. Journal of Network and Computer Applications, 136(??):132–146, June 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300918.

Anonymous:2019:EBm

[2499] Anonymous. Editorial Board. Journal of Network and Computer Applications, 137(??):ii, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301596.

Jan:2019:SEE

[2500] Mian Ahmad Jan, Wenjing Zhang, Muhammad Usman, Zhiyuan Tan, Fazlullah Khan, and Entao Luo. SmartEdge: an end-to-end encryption framework for an edge-enabled smart city application. Journal of Network and Computer Applications, 137(??):1–10, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519300827.

Fu:2019:MFW

[2501] Xiuwen Fu, Haiqing Yao, Octavian Postolache, and Yongsheng Yang. Message forwarding for WSN-Assisted Opportunistic Network in disaster scenarios. Journal of Network and Computer Applications, 137(??): 11–24, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301171.

Mohideen:2019:EIT

[2502] A. I. C. Mohideen, M. Rajiullah, R. Secchi, G. Fairhurst, A. Brunstrom, and F. Weinrank. Evaluating the impact of transport mechanisms on web performance for effective web access. Journal of Network and Computer Applications, 137(??):25–34, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519301183.

Adhikari:2019:MOA

[2503] Mainak Adhikari and Satish Narayana Srirama. Multi-objective accelerated particle swarm optimization with a container-based scheduling for Internet-of-Things in cloud environment. Journal of Network and Computer Applications, 137(??):35–61, July 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301158.

Younus:2019:SSD

[2504] Muhammad Usman Younus, Saif ul Islam, Ihsan Ali, Suleman Khan, and Muhammad Khurram Khan. A survey on software defined networking enabled smart buildings: Architecture challenges and use cases. Journal of Network and Computer Applications, 137(??):62– 77, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301146.

Shrivastava:2019:TSO

[2505] Rudraksh Shrivastava, Konstantinos Samdanis, and Vincenzo Sciancalepore. Towards service-oriented soft spectrum slicing for 5G TDD networks. Journal of Network and Computer Applications, 137(??):78– 90, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300177.

Li:2019:SIA

[2506] Jin Li, Francesco Palmieri, and Qiben Yan. Special issue on advances in security and privacy in IoT. Journal of Network and Computer Applications, 137(??):91–92, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301274.

Altaf:2019:TMI

[2507] Ayesha Altaf, Haider Abbas, Faiza Iqbal, and Abdelouahid Derhab. Trust models of Internet of Smart Things: a survey, open issues, and future directions. Journal of Network and Computer Applications, 137(??):93–111, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300839.

Sun:2019:DAC

[2508] Guolin Sun, Daniel Ayepah-Mensah, Li Lu, Wei Jiang, and Guisong Liu. Delay-aware content distribution via cell clustering and content placement for multiple tenants. Journal of Network and Computer Applications, 137(??):112–126, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930116X.

Assefa:2019:SEE

[2509] Beakal Gizachew Assefa and Öznur Özkasap. A survey of energy efficiency in SDN: Software-base methods and optimization models. Journal of Network and Computer Applications, 137(??):127–143, July 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519301134.

Anonymous:2019:EBn

[2510] Anonymous. Editorial Board. Journal of Network and Computer Applications, 138(??):ii, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301742.

Ghourab:2019:SDM

[2511] Esraa M. Ghourab, Mohamed Azab, and Ahmed Mansour. Spatiotemporal diversification by moving-target defense through benign employment of false-data injection for dynamic, secure cognitive radio network. Journal of Network and Computer Applications, 138(??):1–14, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519300712.

Kim:2019:LTS

[2512] Jun Young Kim, Wen Hu, Dilip Sarkar, and Sanjay Jha. Longterm secure management of large scale Internet of Things applications. Journal of Network and Computer Applications, 138(??):15–26, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301390.

Mondal:2019:COC

[2513] Sourav Mondal, Goutam Das, and Elaine Wong. Cost-optimal cloudlet placement frameworks over fiber-wireless access networks for low-latency applications. Journal of Network and Computer Applications, 138(??): 27–38, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301389.

Aydeger:2019:SER

[2514] Abdullah Aydeger, Nico Saputro, Kemal Akkaya, and Selcuk Uluagac. SDN-enabled recovery for Smart Grid teleprotection applications in post-disaster scenarios. Journal of Network and Computer Applications, 138(??):39–50, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301353.

Chowdhury:2019:DDA

[2515] Abishi Chowdhury, Shital A. Raut, and Husnu S. Narman. DA-DRLS: Drift adaptive deep reinforcement learning based scheduling for IoT resource management. Journal of Network and Computer Applications, 138(??):51–65, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301341.

Gupta:2019:RNP

[2516] Abhishek Gupta, Alagan Anpalagan, Glaucio H. S. Carvalho, Ahmed S. Khwaja, Ling Guan, and Isaac Woungang. Retraction notice to Prevailing and emerging cyber threats and security practices in IoT-Enabled smart grids: a survey [J. Netw. Comput. Appl. 132 (2019) 118–148]. Journal of Network and Computer Applications, 138(??): 66, July 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301705. See [2456].

Anonymous:2019:EBo

[2517] Anonymous. Editorial Board. Journal of Network and Computer Applications, 139(??):ii, August 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301857.

Atiquzzaman:2019:ENE

[2518] Mohammed Atiquzzaman. Editorial: New editors 2019. Journal of Network and Computer Applications, 139(??):iii–v, August 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301882.

Rocha:2019:SMA

[2519] Vladimir Rocha and Anarosa Alves Franco Brandão. A scalable multiagent architecture for monitoring IoT devices. Journal of Network and Computer Applications, 139(??):1–14, August 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301419.

Liu:2019:NSS

[2520] Xiaowu Liu, Jiguo Yu, Weifeng Lv, Dongxiao Yu, Yinglong Wang, and Yu Wu. Network security situation: From awareness to awarenesscontrol. Journal of Network and Computer Applications, 139(??):15– 30, August 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301468.

Hossain:2019:RRC

[2521] Md. Farhad Hossain, Ayman Uddin Mahin, Topojit Debnath, Farjana Binte Mosharrof, and Khondoker Ziaul Islam. Recent research in cloud radio access network (C-RAN) for 5G cellular systems — a survey. Journal of Network and Computer Applications, 139(??):31–48, August 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301432.

Rabaninejad:2019:CLC

[2522] Reyhaneh Rabaninejad, Mahmoud Ahmadian Attari, Maryam Rajabzadeh Asaar, and Mohammad Reza Aref. Comments on a lightweight cloud auditing scheme: Security analysis and improvement. Journal of Network and Computer Applications, 139(??):49–56, August 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519301365.

Pal:2019:PBA

[2523] Shantanu Pal, Michael Hitchens, Vijay Varadharajan, and Tahiry Rabehaja. Policy-based access control for constrained healthcare resources in the context of the Internet of Things. Journal of Network and Computer Applications, 139(??):57–74, August 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301377.

Anonymous:2019:EBp

[2524] Anonymous. Editorial Board. Journal of Network and Computer Applications, 140(??):ii, August 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302000.

Ray:2019:ECI

[2525] Partha Pratim Ray, Dinesh Dash, and Debashis De. Edge computing for Internet of Things: a survey, e-healthcare case study and future direction. Journal of Network and Computer Applications, 140 (??):1–22, August 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301651.

Li:2019:OLS

[2526] Yue Li, Yongmin Zhang, and Lin Cai. Optimal location of supplementary node in UAV surveillance system. Journal of Network and Computer Applications, 140(??):23–39, August 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301663.

Li:2019:TER

[2527] Jirui Li, Xiaoyong Li, Xianglong Cheng, Jie Yuan, and Rui Zhang. A trustworthiness-enhanced reliable forwarding scheme in mobile Internet of Things. Journal of Network and Computer Applications, 140 (??):40–53, August 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301559.

Zeng:2019:PPA

[2528] Yali Zeng, Li Xu, Xu Yang, Xun Yi, and Ibrahim Khalil. Privacypreserving aggregation for cooperative spectrum sensing. Journal of Network and Computer Applications, 140(??):54–64, August 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451930178X.

Majumdar:2019:AMC

[2529] Prasanta Majumdar and Tanmay De. An alternative minimum cost route setup algorithm in a large EON for long-haul traffic under dynamic traffic grooming. Journal of Network and Computer Applications, 140(??):65–86, August 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301407.

Anonymous:2019:EBq

[2530] Anonymous. Editorial Board. Journal of Network and Computer Applications, 141(??):ii, September 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519302103.

Arfeen:2019:RWD

[2531] Asad Arfeen, Krzysztof Pawlikowski, Don McNickle, and Andreas Willig. The role of the Weibull distribution in modelling traffic in Internet access and backbone core networks. Journal of Network and Computer Applications, 141(??):1–22, September 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519301547.

Saraswat:2019:CSS

[2532] Surbhi Saraswat, Vishal Agarwal, Hari Prabhat Gupta, Rahul Mishra, Ashish Gupta, and Tanima Dutta. Challenges and solutions in Software Defined Networking: a survey. Journal of Network and Computer Applications, 141(??):23–58, September 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301444.

Gani:2019:LWS

[2533] Md Osman Gani, Taskina Fayezeen, Richard J. Povinelli, Roger O. Smith, Muhammad Arif, Ahmed J. Kattan, and Sheikh Iqbal Ahamed. A light weight smartphone based human activity recognition system with high accuracy. Journal of Network and Computer Applications, 141(??):59–72, September 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301535.

Liu:2019:IMS

[2534] Yanbing Liu, Xiaowei Qin, Ting Zhu, Xiaohui Chen, and Guo Wei. Improve MPTCP with SDN: From the perspective of resource pooling. Journal of Network and Computer Applications, 141(??):73–85, September 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301912.

Luo:2019:SEP

[2535] Long Luo, Hongfang Yu, Shouxi Luo, Zilong Ye, Xiaojiang Du, and Mohsen Guizani. Scalable explicit path control in software-defined networks. Journal of Network and Computer Applications, 141(??):86–103, September 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451930181X.

Luo:2019:CNU

[2536] Shouxi Luo, Hongfang Yu, Long Luo, and Lemin Li. Customizable network update planning in SDN. Journal of Network and Computer Applications, 141(??):104–115, September 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804519301675.

Anonymous:2019:EBr

[2537] Anonymous. Editorial Board. Journal of Network and Computer Applications, 142(??):ii, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519302346.

Atiquzzaman:2019:EBP

[2538] Mohammed Atiquzzaman. Editorial: 2018 best paper awards. Journal of Network and Computer Applications, 142(??):iii, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519302371.

Atrey:2019:SSF

[2539] Ankita Atrey, Gregory Van Seghbroeck, Higinio Mora, Filip De Turck, and Bruno Volckaert. SpeCH: a scalable framework for data placement of data-intensive services in geo-distributed clouds. Journal of Network and Computer Applications, 142(??):1–14, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301791.

Jing:2019:RSB

[2540] Xuyang Jing, Jingjing Zhao, Qinghua Zheng, Zheng Yan, and Witold Pedrycz. A reversible sketch-based method for detecting and mitigating amplification attacks. Journal of Network and Computer Applications, 142(??):15–24, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302048.

Ahrar:2019:MAS

[2541] Erfan Mozaffari Ahrar, Mohammad Nassiri, and Fabrice Theoleyre. Multipath aware scheduling for high reliability and fault tolerance in low power industrial networks. Journal of Network and Computer Applications, 142(??):25–36, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519301808.

Wahab:2019:MMA

[2542] Omar Abdel Wahab, Nadjia Kara, Claes Edstrom, and Yves Lemieux. MAPLE: a machine learning approach for efficient placement and adjustment of virtual network functions. Journal of Network and Computer Applications, 142(??):37–50, September 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301924.

Tan:2019:RCT

[2543] Xincheng Tan, Yi Xie, Haishou Ma, Shunzheng Yu, and Jiankun Hu. Recognizing the content types of network traffic based on a hybrid DNN-HMM model. Journal of Network and Computer Applications, 142 (??):51–62, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301948. See corrigendum [2587].

Karimian-Aliabadi:2019:ACP

[2544] Soroush Karimian-Aliabadi, Danilo Ardagna, Reza Entezari-Maleki, Eugenio Gianniti, and Ali Movaghar. Analytical composite performance models for Big Data applications. Journal of Network and Computer Applications, 142(??):63–75, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804519302061.

Ghobaei-Arani:2019:ARP

[2545] Mostafa Ghobaei-Arani, Reihaneh Khorsand, and Mohammadreza Ramezanpour. An autonomous resource provisioning framework for massively multiplayer online games in cloud environment. Journal of Network and Computer Applications, 142(??):76–97, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301961.

Wang:2019:HSD

[2546] Shie-Yuan Wang, Chia-Ming Wu, Yi-Bing Lin, and Ching-Chun Huang. High-speed data-plane packet aggregation and disaggregation by P4 switches. Journal of Network and Computer Applications, 142(??): 98–110, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301687.

Fanian:2019:CBR

[2547] Fakhrosadat Fanian and Marjan Kuchaki Rafsanjani. Cluster-based routing protocols in wireless sensor networks: a survey based on methodology. Journal of Network and Computer Applications, 142(??):111–142, September 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301456.

Anonymous:2019:EBs

[2548] Anonymous. Editorial Board. Journal of Network and Computer Applications, 143(??):ii, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302711.

Kumar:2019:CSS

[2549] Mohit Kumar, S. C. Sharma, Anubhav Goel, and S. P. Singh. A comprehensive survey for scheduling techniques in cloud computing. Journal of Network and Computer Applications, 143(??):1–33, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302036.

Gawas:2019:NSC

[2550] Mahadev A. Gawas and Sweta S. Govekar. A novel selective cross layer based routing scheme using ACO method for vehicular networks. Journal of Network and Computer Applications, 143(??):34–46, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519301778.

Gupta:2019:PSS

[2551] Mani Shekhar Gupta and Krishan Kumar. Progression on spectrum sensing for cognitive radio networks: a survey, classification, challenges and future research issues. Journal of Network and Computer Applications, 143(??):47–76, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519301936.

Liu:2019:WRD

[2552] Meiguang Liu, Lei Zhang, Panlong Yang, Liangfu Lu, and Liangyi Gong. Wi-Run: Device-free step estimation system with commodity Wi-Fi. Journal of Network and Computer Applications, 143(??):77–88, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451930164X.

Huf:2019:CHW

[2553] Alexis Huf and Frank Siqueira. Composition of heterogeneous web services: a systematic review. Journal of Network and Computer Applications, 143(??):89–110, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480451930205X.

Singh:2019:WEI

[2554] Abhishek Singh, Ashish Payal, and Sourabh Bharti. A walkthrough of the emerging IoT paradigm: Visualizing inside functionalities, key features, and open issues. Journal of Network and Computer Applications, 143(??):111–151, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302188.

Li:2019:EEF

[2555] Chunlin Li, YaPing Wang, Yi Chen, and Youlong Luo. Energy-efficient fault-tolerant replica management policy with deadline and budget constraints in edge-cloud environment. Journal of Network and Computer Applications, 143(??):152–166, October 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519301420.

Rathore:2019:BBB

[2556] Shailendra Rathore, Byung Wook Kwon, and Jong Hyuk Park. Block-SecIoTNet: Blockchain-based decentralized security architecture for IoT network. Journal of Network and Computer Applications, 143 (??):167–177, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302243.

Hasan:2019:CRW

[2557] Khalid Hasan, Kamanashis Biswas, Khandakar Ahmed, Nazmus S. Nafi, and Md Saiful Islam. A comprehensive review of wireless body area network. Journal of Network and Computer Applications, 143 (??):178–198, October 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302218.

Anonymous:2019:EBt

[2558] Anonymous. Editorial Board. Journal of Network and Computer Applications, 144(??):ii, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930284X.

Liew:2019:PBO

[2559] Jiun Terng Liew, Fazirulhisyam Hashim, Aduwati Sali, Mohd Fadlee A. Rasid, and Abbas Jamalipour. Probability-based opportunity dynamic adaptation (PODA) of contention window for home M2M networks. Journal of Network and Computer Applications, 144(??):1–12, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302164.

Aggarwal:2019:BSC

[2560] Shubhani Aggarwal, Rajat Chaudhary, Gagangeet Singh Aujla, Neeraj Kumar, Kim-Kwang Raymond Choo, and Albert Y. Zomaya. Blockchain for smart communities: Applications challenges and opportunities. Journal of Network and Computer Applications, 144(??):13–48, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302231.

Huang:2019:ODD

[2561] Hailong Huang, Andrey V. Savkin, Ming Ding, and Mohamed Ali Kaafar. Optimized deployment of drone base station to improve user experience in cellular networks. Journal of Network and Computer Applications, 144(??):49–58, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302267.

Roy:2019:QAS

[2562] Deepsubhra Guha Roy, Puja Das, Debashis De, and Rajkumar Buyya. QoS-aware secure transaction framework for Internet of Things using blockchain mechanism. Journal of Network and Computer Applications, 144(??):59–78, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930219X.

Ravidas:2019:ACI

[2563] Sowmya Ravidas, Alexios Lekidis, Federica Paci, and Nicola Zannone. Access control in Internet-of-Things: a survey. Journal of Network and Computer Applications, 144(??):79–101, October 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451930222X.

Chang:2019:SAD

[2564] Ching-Hao Chang, Ying-Dar Lin, Yu-Kuen Lai, and Yuan-Cheng Lai. A scalable and accurate distributed traffic generator with Fourier transformed distribution over multiple commodity platforms. Journal of Network and Computer Applications, 144(??):102–117, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302255.

Maurya:2019:DAE

[2565] Sonam Maurya, Vinod Kumar Jain, and Debanjan Roy Chowdhury. Delay aware energy efficient reliable routing for data transmission in heterogeneous mobile sink wireless sensor network. Journal of Network and Computer Applications, 144(??):118–137, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302176.

Benayache:2019:MMM

[2566] Ayoub Benayache, Azeddine Bilami, Sami Barkat, Pascal Lorenz, and Hafnaoui Taleb. MsM: a microservice middleware for smart WSN-based IoT application. Journal of Network and Computer Applications, 144 (??):138–154, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302206.

Kumar:2019:NNB

[2567] Naveen Kumar, Abdul Aleem, Ashutosh Kumar Singh, and Shashank Srivastava. NBP: Namespace-based privacy to counter timing-based attack in named data networking. Journal of Network and Computer Applications, 144(??):155–170, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804519302280.

Kaur:2019:FAI

[2568] Harkiran Kaur and Sandeep K. Sood. Fog-assisted IoT-enabled scalable network infrastructure for wildfire surveillance. Journal of Network and Computer Applications, 144(??):171–183, October 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302292.

Afsar:2019:LBC

[2569] M. Mehdi Afsar and Mohamed Younis. A load-balanced cross-layer design for energy-harvesting sensor networks. Journal of Network and Computer Applications, 145(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519302152.

Ahmad:2019:STA

[2570] Ashar Ahmad, Muhammad Saad, and Aziz Mohaisen. Secure and transparent audit logs with BlockAudit. Journal of Network and Computer Applications, 145(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302401.

Ahmed:2019:TRI

[2571] Abdelmuttlib Ibrahim Abdalla Ahmed, Siti Hafizah Ab Hamid, Abdullah Gani, Suleman Khan, and Muhammad Khurram Khan. Trust and reputation for Internet of Things: Fundamentals, taxonomy, and open research challenges. Journal of Network and Computer Applications, 145(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302437.

Alipio:2019:TIS

[2572] Melchizedek Alipio, Nestor Michael Tiglao, Fawaz Bokhari, and Salman Khalid. TCP incast solutions in data center networks: a classification and survey. Journal of Network and Computer Applications, 145 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302553.

Anonymous:2019:EBu

[2573] Anonymous. Editorial Board. Journal of Network and Computer Applications, 145(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302942.

Anonymous:2019:EBv

[2574] Anonymous. Editorial Board. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303121.

Anonymous:2019:PN

[2575] Anonymous. Publisher’s note. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302954.

Bawany:2019:SSB

[2576] Narmeen Zakaria Bawany and Jawwad A. Shamsi. SEAL: SDN based secure and agile framework for protecting smart city applications from DDoS attacks. Journal of Network and Computer Applications, 146 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451930195X.

Feng:2019:TDR

[2577] Sheng Feng, Shigen Shen, Longjun Huang, Adam C. Champion, Shui Yu, Chengdong Wu, and Yunzhou Zhang. Three-dimensional robot localization using cameras in wireless multimedia sensor networks. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302772.

Gorbenko:2019:FTI

[2578] Anatoliy Gorbenko, Alexander Romanovsky, and Olga Tarasyuk. Fault tolerant Internet computing: Benchmarking and modelling trade-offs between availability, latency and consistency. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519302462.

Goudarzi:2019:FDD

[2579] Mohammad Goudarzi, Marimuthu Palaniswami, and Rajkumar Buyya. A fog-driven dynamic resource allocation technique in ultra dense femtocell networks. Journal of Network and Computer Applications, 146 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302413.

Guri:2019:UMG

[2580] Mordechai Guri, Rami Puzis, Kim-Kwang Raymond Choo, Sergey Rubinshtein, Gabi Kedma, and Yuval Elovici. Using malware for the greater good: Mitigating data leakage. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302309.

Mamolar:2019:APM

[2581] Ana Serrano Mamolar, Pablo Salvá-García, Enrique Chirivella-Perez, Zeeshan Pervez, Jose M. Alcaraz Calero, and Qi Wang. Autonomic protection of multi-tenant 5G mobile networks against UDP flooding DDoS attacks. Journal of Network and Computer Applications, 146 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302504.

Marsa-Maestre:2019:RRR

[2582] Ivan Marsa-Maestre, Jose Manuel Gimenez-Guzman, David Orden, Enrique de la Hoz, and Mark Klein. REACT: reactive resilience for critical infrastructures using graph-coloring techniques. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302279.

Mergenci:2019:GRA

[2583] Cem Mergenci and Ibrahim Korpeoglu. Generic resource allocation metrics and methods for heterogeneous cloud infrastructures. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302474.

Paul:2019:CHC

[2584] Partha Sarathi Paul, Bishakh Chandra Ghosh, Hridoy Sankar Dutta, Kingshuk De, Arka Prava Basu, Prithviraj Pramanik, Sujoy Saha, Sandip Chakraborty, Niloy Ganguly, and Subrata Nandi. CRIMP: Here crisis mapping goes offline. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302528.

Senouci:2019:RUA

[2585] Mustapha Reda Senouci and Abdelhamid Mellouk. A robust uncertaintyaware cluster-based deployment approach for WSNs: Coverage, connectivity, and lifespan. Journal of Network and Computer Applications, 146 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302486.

Shen:2019:HMC

[2586] Shigen Shen, Haiping Zhou, Sheng Feng, Longjun Huang, Jianhua Liu, Shui Yu, and Qiying Cao. HSIRD: a model for characterizing dynamics of malware diffusion in heterogeneous WSNs. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519302541.

Tan:2019:CRC

[2587] Xincheng Tan, Yi Xie, Haishou Ma, Shunzheng Yu, and Jiankun Hu. Corrigendum to “Recognizing the content types of network traffic based on a hybrid DNN-HMM model” [J. Netw. Comput. Appl. 142 (2019) 51–62]. Journal of Network and Computer Applications, 146 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302450. See [2543].

Valenza:2019:FSC

[2588] Fulvio Valenza, Serena Spinoso, and Riccardo Sisto. Formally specifying and checking policies and anomalies in service function chaining. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451930253X.

Yan:2019:SSR

[2589] Yanjun Yan, Guangjie Han, and Huihui Xu. A survey on secure routing protocols for satellite network. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302498.

Yang:2019:PAM

[2590] Bo Yang, Bo Li, Zhongjiang Yan, Der-Jiunn Deng, and Mao Yang. Performance analysis of multi-channel MAC with single transceiver for the next generation WLAN. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302425.

Zamanifar:2019:RNA

[2591] Azadeh Zamanifar and Eslam Nazemi. Retraction notice to “An approach for predicting health status in IoT health care” [YJNCA 134 (2019) 100–113]. Journal of Network and Computer Applications, 146 (??), ???? 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302796. See [2476].

Zhang:2019:TVN

[2592] Chuanji Zhang, Harshvardhan P. Joshi, George F. Riley, and Steven A. Wright. Towards a virtual network function research agenda: a systematic literature review of VNF design considerations. Journal of Network and Computer Applications, 146(??), ???? 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302516.

Anonymous:2019:Da

[2593] Anonymous. 1 December 2019. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2019:EBw

[2594] Anonymous. Editorial Board. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303352.

Asghari:2019:PSI

[2595] Masoud Asghari, Saleh Yousefi, and Dusit Niyato. Pricing strategies of IoT wide area network service providers with complementary services included. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302784.

Manan:2019:DID

[2596] Jamila Manan, Atiq Ahmed, Ihsan Ullah, Leïla Merghem-Boulahia, and Dominique Gaïti. Distributed intrusion detection scheme for next generation networks. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302747.

Imani:2019:SER

[2597] Maryam Imani and Gholam Ali Montazer. A survey of emotion recognition methods with emphasis on E-Learning environments. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519302759.

Demirci:2019:OPV

[2598] Sedef Demirci and Seref Sagiroglu. Optimal placement of virtual network functions in software defined networks: a survey. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519302760.

Alimohammadi:2019:CBM

[2599] Hamed Alimohammadi and Mahmood Ahmadi. Clustering-based many-field packet classification in Software-Defined Networking. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519302802.

Jain:2019:QQD

[2600] Shubhra Jain, K. K. Pattanaik, and Anupam Shukla. QWRP: Querydriven virtual wheel based routing protocol for wireless sensor networks with mobile sink. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302887.

Yu:2019:JOS

[2601] Yinbo Yu, Jianfeng Yang, Chengcheng Guo, Hong Zheng, and Jiancheng He. Joint optimization of service request routing and instance placement in the microservice system. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303017.

Nguyen:2019:TTE

[2602] Phi Le Nguyen, Kien Nguyen, Huy Vu, and Yusheng Ji. TELPAC: a time and energy efficient protocol for locating and patching coverage holes in WSNs. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519302991.

Dai:2019:SAM

[2603] Hua Dai, Xuelong Dai, Xun Yi, Geng Yang, and Haiping Huang. Semantic-aware multi-keyword ranked search scheme over encrypted cloud data. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303029.

Ali:2019:BBB

[2604] Gauhar Ali, Naveed Ahmad, Yue Cao, Qazi Ejaz Ali, Fazal Azim, and Haitham Cruickshank. BCON: Blockchain based access CONtrol across multiple conflict of interest domains. Journal of Network and Computer Applications, 147(??):??, December 1, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303005.

Anonymous:2019:Db

[2605] Anonymous. 15 December 2019. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2019:EBx

[2606] Anonymous. Editorial Board. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303455.

Abane:2019:LFS

[2607] Amar Abane, Mehammed Daoui, Samia Bouzefrane, and Paul Muhlethaler. A lightweight forwarding strategy for named data networking in low-end IoT. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303054.

Enoch:2019:SMA

[2608] Simon Yusuf Enoch, Jin B. Hong, and Dong Seong Kim. Security modelling and assessment of modern networks using time independent Graphical Security Models. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480451930308X.

Laaziz:2019:FFS

[2609] Lahlou Laaziz, Nadjia Kara, Rafi Rabipour, Claes Edstrom, and Yves Lemieux. FASTSCALE: a fast and scalable evolutionary algorithm for the joint placement and chaining of virtualized services. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519302875.

Benmakrelouf:2019:RNP

[2610] Souhila Benmakrelouf, Nadjia Kara, Hanine Tout, Rafi Rabipour, and Claes Edstrom. Resource needs prediction in virtualized systems: Generic proactive and self-adaptive solution. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303030.

Palmieri:2019:NAD

[2611] Francesco Palmieri. Network anomaly detection based on logistic regression of nonlinear chaotic invariants. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804519303200.

Anagnostopoulos:2019:ITO

[2612] Christos Anagnostopoulos and Kostas Kolomvatsos. An intelligent, timeoptimized monitoring scheme for edge nodes. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303182.

Mukherjee:2019:SUA

[2613] Anandarup Mukherjee, Sudip Misra, and Narendra Singh Raghuwanshi. A survey of unmanned aerial sensing solutions in precision agriculture. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303212.

Alaoui:2019:HDT

[2614] El Arbi Abdellaoui Alaoui, Hanane Zekkori, and Said Agoujil. Hybrid delay tolerant network routing protocol for heterogeneous networks. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303169.

Li:2019:SHR

[2615] Shaonan Li, Wenyu Qu, Chunfeng Liu, Tie Qiu, and Zhao Zhao. Survey on high reliability wireless communication for underwater sensor networks. Journal of Network and Computer Applications, 148(??):??, December 15, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303066.

Anonymous:2020:Ja

[2616] Anonymous. 1 January 2020. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBa

[2617] Anonymous. Editorial Board. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303510.

Anonymous:2020:TYR

[2618] Anonymous. Thank You Reviewers. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303522.

Wu:2020:SAC

[2619] Dapeng Wu, Bingxu Liu, Qing Yang, and Ruyan Wang. Social-aware cooperative caching mechanism in mobile social networks. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303170.

Gu:2020:SLP

[2620] Fei Gu, Jianwei Niu, Landu Jiang, Xue Liu, and Mohammed Atiquzzaman. Survey of the low power wide area network technologies. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519303194.

Monge:2020:COM

[2621] David A. Monge, Elina Pacini, Cristian Mateos, Enrique Alba, and Carlos García Garino. CMI: an online multi-objective genetic autoscaler for scientific and engineering workflows in cloud infrastructures with unreliable virtual machines. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303248.

Lyu:2020:SSB

[2622] Qiuyun Lyu, Yizhen Qi, Xiaochen Zhang, Huaping Liu, Qiuhua Wang, and Ning Zheng. SBAC: a secure blockchain-based access control framework for information-centric networking. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804519303042.

Singh:2020:STB

[2623] Amritraj Singh, Kelly Click, Reza M. Parizi, Qi Zhang, Ali Dehghantanha, and Kim-Kwang Raymond Choo. Sidechain technologies in blockchain networks: an examination and state-of-the-art review. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303315.

Ghribi:2020:STM

[2624] Mayssa Ghribi and Aref Meddeb. Survey and taxonomy of MAC, routing and cross layer protocols using wake-up radio. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480451930325X.

Nagar:2020:AME

[2625] Jaiprakash Nagar, S. K. Chaturvedi, and Sieteng Soh. An analytical model to estimate the performance metrics of a finite multihop network deployed in a rectangular region. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303261.

Shoumy:2020:MBD

[2626] Nusrat J. Shoumy, Li-Minn Ang, Kah Phooi Seng, D. M. Motiur Rahaman, and Tanveer Zia. Multimodal big data affective analytics: a comprehensive survey using text, audio, visual and physiological signals. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303078.

Zhong:2020:EDM

[2627] Hong Zhong, Zhanfei Li, Jie Cui, Yue Sun, and Lu Liu. Efficient dynamic multi-keyword fuzzy search over encrypted cloud data. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303297.

Sengupta:2020:CSA

[2628] Jayasree Sengupta, Sushmita Ruj, and Sipra Das Bit. A comprehensive survey on attacks, security issues and blockchain solutions for IoT and IIoT. Journal of Network and Computer Applications, 149(??):??, January 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303418.

Anonymous:2020:Jb

[2629] Anonymous. 15 January 2020. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBb

[2630] Anonymous. Editorial Board. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303753.

Wang:2020:EMR

[2631] Xiong Wang, Yukun Yang, Hanyu Liu, Jing Ren, Shizhong Xu, Sheng Wang, and Shui Yu. Efficient measurement of round-trip link delays in software-defined networks. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303285.

Alauthman:2020:ERL

[2632] Mohammad Alauthman, Nauman Aslam, Mouhammd Al-Kasassbeh, Suleman Khan, Ahmad Al-Qerem, and Kim-Kwang Raymond Choo. An efficient reinforcement learning-based Botnet detection approach. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480451930339X.

Ali:2020:OCP

[2633] Ahmed S. Ali, Korany R. Mahmoud, and Khaled M. Naguib. Optimal caching policy for wireless content delivery in D2D networks. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519303273.

Canovas:2020:RMT

[2634] Alejandro Canovas, Albert Rego, Oscar Romero, and Jaime Lloret. A robust multimedia traffic SDN-Based management system using patterns and models of QoE estimation with BRNN. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303583.

Khan:2020:EPE

[2635] Ayaz Ali Khan, Muhammad Zakarya, Rahim Khan, Izaz Ur Rahman, Mukhtaj Khan, and Atta ur Rehman Khan. An energy, performance efficient resource consolidation scheme for heterogeneous cloud datacenters. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303571.

Kandi:2020:VKM

[2636] Mohamed Ali Kandi, Hicham Lakhlef, Abdelmadjid Bouabdallah, and Yacine Challal. A versatile key management protocol for secure group and device-to-device communication in the Internet of Things. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519303406.

Khan:2020:RTT

[2637] Muhammad Asif Khan, Ridha Hamila, Nasser Ahmed Al-Emadi, Serkan Kiranyaz, and Moncef Gabbouj. Real-time throughput prediction for cognitive Wi-Fi networks. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303595.

Wazid:2020:LCL

[2638] Mohammad Wazid, Ashok Kumar Das, Vivekananda Bhat K., and Athanasios V. Vasilakos. LAM-CIoT: Lightweight authentication mechanism in cloud-based IoT environment. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480451930356X.

Chen:2020:CRM

[2639] Yan Chen, Yanjing Sun, Nannan Lu, and Bin Wang. Channel-reserved medium access control for edge computing based IoT. Journal of Network and Computer Applications, 150(??):??, January 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303601.

Anonymous:2020:Fa

[2640] Anonymous. 1 February 2020. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBc

[2641] Anonymous. Editorial Board. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303832.

Kucek:2020:ESF

[2642] Stela Kucek and Maria Leitner. An empirical survey of functions and configurations of open-source Capture the Flag (CTF) environments. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303303.

Geng:2020:ECL

[2643] Haijun Geng, Han Zhang, Xingang Shi, Zhiliang Wang, and Xia Yin. Efficient computation of loop-free alternates. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303613.

Guo:2020:DRP

[2644] Jingjing Guo, Chunlin Li, Yi Chen, and Youlong Luo. On-demand resource provision based on load estimation and service expenditure in edge cloud environment. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303662.

Rabbani:2020:HML

[2645] Mahdi Rabbani, Yong Li Wang, Reza Khoshkangini, Hamed Jelodar, Ruxin Zhao, and Peng Hu. A hybrid machine learning approach for malicious behaviour detection and recognition in cloud computing. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303674.

Han:2020:CCT

[2646] Guangjie Han, Haofei Guan, Zeren Zhou, Zhifan Li, Sammy Chan, and Wenbo Zhang. CTRA: a complex terrain region-avoidance charging algorithm in Smart World. Journal of Network and Computer Applications, 151(??):??, February 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519300517.

Anonymous:2020:Fb

[2647] Anonymous. 15 February 2020. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBd

[2648] Anonymous. Editorial Board. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520300175.

Majidi:2020:EMA

[2649] Akbar Majidi, Nazila Jahanbakhsh, Xiaofeng Gao, Jiaqi Zheng, and Guihai Chen. ECN^{+}: a marking-aware optimization for ECN threshold via per-port in data center networks. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303649.

Muchtar:2020:ECS

[2650] Farkhana Muchtar, Abdul Hanan Abdullah, Mosleh Al-Adhaileh, and Kamal Zuhairi Zamli. Energy conservation strategies in Named Data Networking based MANET using congestion control: a review. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804519303716.

Li:2020:LBB

[2651] Chunlin Li, Jianhang Tang, Tao Ma, Xihao Yang, and Youlong Luo. Load balance based workflow job scheduling algorithm in distributed cloud. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303789.

Nobakht:2020:PSP

[2652] Mehdi Nobakht, Yulei Sui, Aruna Seneviratne, and Wen Hu. PG-Fit: Static permission analysis of health and fitness apps in IoT programming frameworks. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303698.

Wang:2020:MRM

[2653] Ru Wang, Muhammad Imran, and Kashif Saleem. A microservice recommendation mechanism based on mobile architecture. Journal of Network and Computer Applications, 152(??):??, February 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804519303704.

Anonymous:2020:Ma

[2654] Anonymous. 1 March 2020. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBe

[2655] Anonymous. Editorial Board. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300321.

Bays:2020:RSV

[2656] Leonardo Richter Bays and Luciano Paschoal Gaspary. Reality shock in virtual network embedding: Flexibilizing demands for dealing with multiple operational requirements in SDNs. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804519303686.

Gibert:2020:RML

[2657] Daniel Gibert, Carles Mateu, and Jordi Planes. The rise of machine learning for detection and classification of malware: Research developments, trends and challenges. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804519303868.

Wu:2020:SMS

[2658] Weihong Wu, Jiang Liu, and Tao Huang. The source-multicast: a sender-initiated multicast member management mechanism in SRv6 networks. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303650.

Rahimi:2020:FBS

[2659] Morteza Rahimi, Maryam Songhorabadi, and Mostafa Haghi Kashani. Fog-based smart homes: a systematic review. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300059.

Yang:2020:PEE

[2660] Liang Yang, Bryan Ng, Winston K. G. Seah, and Deepak Singh. Performance evaluation of equivalent forwarding sets in software defined networking. Journal of Network and Computer Applications, 153(??): ??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300060.

Sun:2020:CCF

[2661] Xiuwen Sun, Hao Li, Dan Zhao, Xingxing Lu, Kaiyu Hou, and Chengchen Hu. Corrigendum to “COIN: a fast packet inspection method over compressed traffic” [J. Netw. Comput. Appl. 127 (2019) 122–134]. Journal of Network and Computer Applications, 153(??):??, March 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804519303790. See [2409].

Anonymous:2020:Mb

[2662] Anonymous. 15 March 2020. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBf

[2663] Anonymous. Editorial Board. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300436.

Wang:2020:SET

[2664] Ye Wang, Qingbao Li, Zhifeng Chen, Ping Zhang, and Guimin Zhang. A survey of exploitation techniques and defenses for program data attacks. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300084.

Wang:2020:FIF

[2665] Zeli Wang, Weiqi Dai, Kim-Kwang Raymond Choo, Hai Jin, and Deqing Zou. FSFC: an input filter-based secure framework for smart contract. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300047.

Fang:2020:RAV

[2666] Lang Fang, Xiaoning Zhang, Keshav Sood, Yunqing Wang, and Shui Yu. Reliability-aware virtual network function placement in carrier networks. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300102.

Khan:2020:STC

[2667] Muhammad Fahad Khan, Kok-Lim Alvin Yau, Rafidah MD. Noor, and Muhammad Ali Imran. Survey and taxonomy of clustering algorithms in 5G. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300138.

Mohsenzadeh:2020:NRP

[2668] Ali Mohsenzadeh, Amir Jalaly Bidgoly, and Yaghoub Farjami. A novel reward and penalty trust evaluation model based on confidence interval using Petri Net. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300072.

Tahaei:2020:RTC

[2669] Hamid Tahaei, Firdaus Afifi, Adeleh Asemi, Faiz Zaki, and Nor Badrul Anuar. The rise of traffic classification in IoT networks: a survey. Journal of Network and Computer Applications, 154(??):??, March 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300126.

Anonymous:2020:Aa

[2670] Anonymous. 1 April 2020. Journal of Network and Computer Applications, 155(??):??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBg

[2671] Anonymous. Editorial Board. Journal of Network and Computer Applications, 155(??):??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300576.

Al-Darrab:2020:SDN

[2672] Abdulaziz Al-Darrab, Ibrahim Al-Darrab, and Ali Rushdi. Softwaredefined networking load distribution technique for an Internet service provider. Journal of Network and Computer Applications, 155(??): ??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300217.

Chakraborty:2020:TDP

[2673] Bodhi Chakraborty, Shekhar Verma, and Krishna Pratap Singh. Temporal Differential Privacy in Wireless Sensor Networks. Journal of Network and Computer Applications, 155(??):??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300229.

Yi:2020:NRA

[2674] Bo Yi, Xingwei Wang, Min Huang, and Yong Zhao. Novel resource allocation mechanism for SDN-based data center networks. Journal of Network and Computer Applications, 155(??):??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452030028X.

Fu:2020:ASD

[2675] Yulong Fu, Hanlu Chen, Qinghua Zheng, Zheng Yan, Raimo Kantola, Xuyang Jing, Jin Cao, and Hui Li. An adaptive security data collection and composition recognition method for security measurement over LTE/LTE-A networks. Journal of Network and Computer Applications, 155(??):??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300230.

Liu:2020:LVD

[2676] Luning Liu, Zhaoming Lu, Luhan Wang, Xin Chen, and Xiangming Wen. Large-volume data dissemination for cellular-assisted automated driving with edge intelligence. Journal of Network and Computer Applications, 155(??):??, April 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300096.

Anonymous:2020:Ab

[2677] Anonymous. 15 April 2020. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBh

[2678] Anonymous. Editorial Board. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300898.

Wei:2020:PPS

[2679] Zhenchun Wei, Chengkai Xia, Xiaohui Yuan, Renhao Sun, Zengwei Lyu, Lei Shi, and Jianjun Ji. The path planning scheme for joint charging and data collection in WRSNs: a multi-objective optimization method. Journal of Network and Computer Applications, 156(??): ??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300394.

Latif:2020:CSI

[2680] Zohaib Latif, Kashif Sharif, Fan Li, Md Monjurul Karim, Sujit Biswas, and Yu Wang. A comprehensive survey of interface protocols for software defined networks. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300370.

Xiao:2020:EPF

[2681] Yang Xiao, Qingqi Pei, Lina Yao, Shui Yu, Lei Bai, and Xianzhi Wang. An enhanced probabilistic fairness-aware group recommendation by incorporating social activeness. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300539.

Yazdinejad:2020:HPF

[2682] Abbas Yazdinejad, Reza M. Parizi, Ali Bohlooli, Ali Dehghantanha, and Kim-Kwang Raymond Choo. A high-performance framework for a network programmable packet processor using P4 and FPGA. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300382.

Rehan:2020:QQA

[2683] Waqas Rehan, Stefan Fischer, Maaz Rehan, Yasser Mawad, and Shahzad Saleem. QCM2R: a QoS-aware cross-layered multichannel multisink routing protocol for stream based wireless sensor networks. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300266.

Bhuiyan:2020:SWS

[2684] Md Zakirul Alam Bhuiyan, Guojun Wang, and Zhong Fan. Smart world systems, applications, and technologies. Journal of Network and Computer Applications, 156(??):??, April 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804520300278.

Anonymous:2020:Mc

[2685] Anonymous. 1 May 2020. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBi

[2686] Anonymous. Editorial Board. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300953.

Sun:2020:RSC

[2687] Guolin Sun, Kun Xiong, Gordon Owusu Boateng, Guisong Liu, and Wei Jiang. Resource slicing and customization in RAN with dueling deep Q-Network. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300473.

Rovetta:2020:BRW

[2688] Stefano Rovetta, Grazyna Suchacka, and Francesco Masulli. Bot recognition in a Web store: an approach based on unsupervised learning. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300515.

Elahi:2020:PPE

[2689] Haroon Elahi, Guojun Wang, and Jianer Chen. Pleasure or pain? An evaluation of the costs and utilities of bloatware applications in Android smartphones. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300527.

Gu:2020:MLI

[2690] Rentao Gu, Zeyuan Yang, and Yuefeng Ji. Machine learning for intelligent optical networks: a comprehensive survey. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300503.

Wang:2020:DAO

[2691] Mu Wang, Changqiao Xu, Xingyan Chen, Lujie Zhong, and Gabriel-Miro Muntean. Decentralized asynchronous optimization for dynamic adaptive multimedia streaming over information centric networking. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300485.

Aldhaheri:2020:AIS

[2692] Sahar Aldhaheri, Daniyal Alghazzawi, Li Cheng, Ahmed Barnawi, and Bandar A. Alzahrani. Artificial immune systems approaches to secure the Internet of Things: a systematic review of the literature and recommendations for future research. Journal of Network and Computer Applications, 157(??):??, May 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300114.

Anonymous:2020:Md

[2693] Anonymous. 15 May 2020. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBj

[2694] Anonymous. Editorial Board. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301211.

Mukta:2020:IEE

[2695] Marufa Yeasmin Mukta, Md Arafatur Rahman, A. Taufiq Asyhari, and Md Zakirul Alam Bhuiyan. IoT for energy efficient green highway lighting systems: Challenges and issues. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300497.

Zhang:2020:SCV

[2696] Mingchuan Zhang, Bowei Hao, Fei Song, Meiyi Yang, Junlong Zhu, and Qingtao Wu. Smart collaborative video caching for energy efficiency in cognitive content centric networks. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520300618.

Zhou:2020:JJP

[2697] Pan Zhou, Xinshu He, Shouxi Luo, Hongfang Yu, and Gang Sun. JPAS: Job-progress-aware flow scheduling for deep learning clusters. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300643.

Sousa:2020:SQO

[2698] Ivo Sousa, Maria Paula Queluz, and António Rodrigues. A survey on QoE-oriented wireless resources scheduling. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804520300680.

Juma:2020:HEE

[2699] Mazen Juma, Azza Abdel Monem, and Khaled Shaalan. Hybrid endto-end VPN security approach for smart IoT objects. Journal of Network and Computer Applications, 158(??):??, May 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300722.

Anonymous:2020:Jc

[2700] Anonymous. 1 June 2020. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBk

[2701] Anonymous. Editorial Board. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301430.

Guevara:2020:CFC

[2702] Judy C. Guevara, Ricardo da S. Torres, and Nelson L. S. da Fonseca. On the classification of fog computing applications: a machine learning perspective. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520300709.

Berruet:2020:EMC

[2703] Brieuc Berruet, Oumaya Baala, Alexandre Caminada, and Valery Guillet. An evaluation method of channel state information fingerprinting for single gateway indoor localization. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520300655.

ElMensoum:2020:MMS

[2704] Imane El Mensoum, Omar Abdul Wahab, Nadjia Kara, and Claes Edstrom. MuSC: a multi-stage service chains embedding approach. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520300679.

Chica:2020:SSC

[2705] Juan Camilo Correa Chica, Jenny Cuatindioy Imbachi, and Juan Felipe Botero Vega. Security in SDN: a comprehensive survey. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520300692.

Xavier:2020:CRA

[2706] Tiago C. S. Xavier, Igor L. Santos, Flavia C. Delicato, Paulo F. Pires, Marcelo P. Alves, Tiago S. Calmon, Ana C. Oliveira, and Claudio L. Amorim. Collaborative resource allocation for Cloud of Things systems. Journal of Network and Computer Applications, 159(??):??, June 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300667.

Anonymous:2020:Jd

[2707] Anonymous. 15 June 2020. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBl

[2708] Anonymous. Editorial Board. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301557.

Shafi:2020:SSI

[2709] Misbah Shafi, Rakesh Kumar Jha, and Manish Sabraj. A survey on security issues of 5G NR: Perspective of artificial dust and artificial rain. Journal of Network and Computer Applications, 160(??): ??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520300710.

Ejaz:2020:CSR

[2710] Waleed Ejaz, Shree K. Sharma, Salman Saadat, Muhammad Naeem, Alagan Anpalagan, and N. A. Chughtai. A comprehensive survey on resource allocation for CRAN in 5G and beyond networks. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301120.

Srirama:2020:ADU

[2711] Satish Narayana Srirama, Mainak Adhikari, and Souvik Paul. Application deployment using containers with auto-scaling for microservices in cloud environment. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030103X.

Sun:2020:SPP

[2712] PanJun Sun. Security and privacy protection in cloud computing: Discussions and challenges. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301168.

Lin:2020:EES

[2713] Kai Lin, Honglong Chen, Xin Ai, Vladimir Shakhov, Lina Ni, Jiguo Yu, and Yanjun Li. EUMD: Efficient slot utilization based missing tag detection with unknown tags. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301144.

Riad:2020:DHA

[2714] Khaled Riad, Teng Huang, and Lishan Ke. A dynamic and hierarchical access control for IoT in multi-authority cloud storage. Journal of Network and Computer Applications, 160(??):??, June 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301077.

Anonymous:2020:Je

[2715] Anonymous. 1 July 2020. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBm

[2716] Anonymous. Editorial Board. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301764.

Wang:2020:AAD

[2717] Yajie Wang, Yu an Tan, Wenjiao Zhang, Yuhang Zhao, and Xiaohui Kuang. An adversarial attack on DNN-based black-box object detectors. Journal of Network and Computer Applications, 161(??): ??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301089.

Xiao:2020:LCH

[2718] Di Xiao, Min Li, Mengdi Wang, Jia Liang, and Ran Liu. Low-cost and high-efficiency privacy-protection scheme for distributed compressive video sensing in wireless multimedia sensor networks. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301284.

ulHaq:2020:STF

[2719] Inam ul Haq, Jian Wang, and Youwen Zhu. Secure two-factor lightweight authentication protocol using self-certified public key cryptography for multi-server 5G networks. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030134X.

Torres-Cruz:2020:CAF

[2720] Noé Torres-Cruz, Mario E. Rivero-Angeles, Gerardo Rubino, Ricardo Menchaca-Mendez, Rolando Menchaca-Mendez, and David Ramirez. A comprehensive analytical framework for VoD services in hybrid CDN-P2P systems. Journal of Network and Computer Applications, 161(??): ??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452030117X.

Dong:2020:RNF

[2721] Shi Dong and Raj Jain. Retraction notice to “Flow online identification method for the encrypted Skype” [YJNCA (2019) 75–85]. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301399.

Tahsien:2020:MLB

[2722] Syeda Manjia Tahsien, Hadis Karimipour, and Petros Spachos. Machine learning based solutions for security of Internet of Things (IoT): a survey. Journal of Network and Computer Applications, 161(??): ??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301041.

Li:2020:ECI

[2723] Wenjuan Li, Weizhi Meng, and Man Ho Au. Enhancing collaborative intrusion detection via disagreement-based semi-supervised learning in IoT environments. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301053.

Wang:2020:BFS

[2724] Yilei Wang, Guoyu Yang, Tao Li, Fengyin Li, Youliang Tian, and Xiaomei Yu. Belief and fairness: a secure two-party protocol toward the view of entropy for IoT devices. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301156.

Pang:2020:NSA

[2725] Xiongwen Pang, Yanqiang Zhou, Pengcheng Li, Weiwei Lin, Wentai Wu, and James Z. Wang. A novel syntax-aware automatic graphics code generation with attention-based deep neural network. Journal of Network and Computer Applications, 161(??):??, July 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301107.

Anonymous:2020:Jf

[2726] Anonymous. 15 July 2020. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBn

[2727] Anonymous. Editorial Board. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301946.

Morais:2020:WSM

[2728] Fernando Zanferrari Morais, Cristiano André da Costa, Antonio Marcos Alberti, Cristiano Bonato Both, and Rodrigo da Rosa Righi. When SDN meets C-RAN: a survey exploring multi-point coordination, interference, and performance. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301296.

Thakur:2020:MDV

[2729] Dipanwita Thakur and Manas Khatua. Multi-domain virtual network embedding with dynamic flow migration in software-defined networks. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301132.

Guo:2020:RDS

[2730] Cheng Guo, Xueru Jiang, Kim-Kwang Raymond Choo, and Yingmo Jie. R-Dedup: Secure client-side deduplication for encrypted data without involving a third-party entity. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301387.

Zeng:2020:LPP

[2731] Yali Zeng, Li Xu, Xu Yang, Xun Yi, and Ibrahim Khalil. Lightweight privacy preservation for secondary users in cognitive radio networks. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520301260.

Benisi:2020:BBD

[2732] Nazanin Zahed Benisi, Mehdi Aminian, and Bahman Javadi. Blockchainbased decentralized storage networks: a survey. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301302.

Huang:2020:AAD

[2733] Teng Huang, Qixiang Zhang, Jiabao Liu, Ruitao Hou, Xianmin Wang, and Ya Li. Adversarial attacks on deep-learning-based SAR image target recognition. Journal of Network and Computer Applications, 162 (??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301065.

Yang:2020:GMA

[2734] Yize Yang, Hongyong Yang, and Fei Liu. Group motion of autonomous vehicles with anti-disturbance protection. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804520301351.

Zhang:2020:STH

[2735] Xiao Zhang, Lijia Xie, and Wang Yao. Spatio-temporal heterogeneous bandwidth allocation mechanism against DDoS attack. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301326.

Xie:2020:CGA

[2736] Xueshuo Xie, Zongming Jin, Jiming Wang, Lei Yang, Ye Lu, and Tao Li. Confidence guided anomaly detection model for anti-concept drift in dynamic logs. Journal of Network and Computer Applications, 162(??):??, July 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301338.

Anonymous:2020:Ac

[2737] Anonymous. 1 August 2020. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBo

[2738] Anonymous. Editorial Board. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302009.

DeLaTorreParra:2020:DIT

[2739] Gonzalo De La Torre Parra, Paul Rad, Kim-Kwang Raymond Choo, and Nicole Beebe. Detecting Internet of Things attacks using distributed deep learning. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301363.

Garroppo:2020:RGW

[2740] Rosario G. Garroppo, Maria Grazia Scutellà, and Fabio D’Andreagiovanni. Robust green Wireless Local Area Networks: a matheuristic approach. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520301314.

Raza:2020:PBG

[2741] Ali Raza and Sanghwan Lee. A priority based greedy path assignment mechanism in OpenFlow based datacenter networks. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301272.

Kassab:2020:ZSI

[2742] Wafa’a Kassab and Khalid A. Darabkh. A–Z survey of Internet of Things: Architectures, protocols, applications, recent advances, future directions and recommendations. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520301375.

Cai:2020:CDP

[2743] Jun Cai, Zhongwei Huang, Jianzhen Luo, Yan Liu, Huimin Zhao, and Liping Liao. Composing and deploying parallelized service function chains. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301119.

Benmansour:2020:PAI

[2744] Tariq Benmansour, Toufik Ahmed, Samira Moussaoui, and Zouina Doukha. Performance analyses of the IEEE 802.15.6 Wireless Body Area Network with heterogeneous traffic. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520301259.

Ghosh:2020:SCB

[2745] Arunima Ghosh, Shashank Gupta, Amit Dua, and Neeraj Kumar. Security of cryptocurrencies in blockchain technology: State-of-art, challenges and future prospects. Journal of Network and Computer Applications, 163(??):??, August 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301090.

Anonymous:2020:Ad

[2746] Anonymous. 15 August 2020. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBp

[2747] Anonymous. Editorial Board. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302241.

Conti:2020:RMC

[2748] Mauro Conti, Pallavi Kaliyar, and Chhagan Lal. A robust multicast communication protocol for low power and lossy networks. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301491.

Chen:2020:ADC

[2749] Ming Chen, Shunkang Zhang, Hai Deng, Bing Chen, Changyou Xing, and Bo Xu. Automatic deployment and control of network services in NFV environments. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030151X.

Tian:2020:OAB

[2750] Yu Tian, Baoqi Huang, Bing Jia, and Long Zhao. Optimizing AP and Beacon Placement in WiFi and BLE hybrid localization. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301478.

Khan:2020:SUE

[2751] Abd Ullah Khan, Ghulam Abbas, Ziaul Haq Abbas, Muhammad Waqas, and Ahmad Kamal Hassan. Spectrum utilization efficiency in the cognitive radio enabled 5G-based IoT. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520301600.

Ghosh:2020:MST

[2752] Shreya Ghosh, Soumya K. Ghosh, and Rajkumar Buyya. MARIO: a spatio-temporal data mining framework on Google Cloud to explore mobility dynamics from taxi trajectories. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520301661.

Anagnostopoulos:2020:ECI

[2753] Christos Anagnostopoulos. Edge-centric inferential modeling and analytics. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301703.

Xie:2020:MLT

[2754] Lixia Xie, Ying Ding, Hongyu Yang, and Ze Hu. Mitigating LFA through segment rerouting in IoT environment with traceroute flow abnormality detection. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301648.

Wu:2020:NID

[2755] Zhendong Wu, Jingjing Wang, Liqin Hu, Zhang Zhang, and Han Wu. A network intrusion detection method based on semantic re-encoding and deep learning. Journal of Network and Computer Applications, 164(??):??, August 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301624.

Anonymous:2020:Sa

[2756] Anonymous. 1 September 2020. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBq 8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520302484.

Fernandes:2020:RBB

[2758] Eder Leão Fernandes, Elisa Rojas, Joaquin Alvarez-Horcajo, Zoltàn Lajos Kis, Davide Sanvito, Nicola Bonelli, Carmelo Cascone, and Christian Esteve Rothenberg. The road to BOFUSS: the basic OpenFlow userspace software switch. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301594.

Murtaza:2020:QAS

[2759] Faizan Murtaza, Adnan Akhunzada, Saif ul Islam, Jalil Boudjadar, and Rajkumar Buyya. QoS-aware service provisioning in fog computing. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452030148X.

Wang:2020:ADP

[2760] Shie-Yuan Wang, Jun-Yi Li, and Yi-Bing Lin. Aggregating and disaggregating packets with various sizes of payload in P4 switches at 100 Gbps line rate. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301508.

Giti:2020:SCA

[2761] J. E. Giti, A. Sakzad, B. Srinivasan, J. Kamruzzaman, and R. Gaire. Secrecy capacity against adaptive eavesdroppers in a random wireless network using friendly jammers and protected zone. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301727.

Hernandez:2020:UPD

[2762] Steven M. Hernandez and Eyuphan Bulut. Using perceived direction information for anchorless relative indoor localization. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301880.

Zurita:2020:BOJ

[2763] Gustavo Zurita, Amit K. Shukla, José A. Pino, José M. Merigó, Valeria Lobos-Ossandón, and Pranab K. Muhuri. A bibliometric overview of the Journal of Network and Computer Applications between 1997 and 2019. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301697.

Li:2020:APB

[2764] Chunlin Li, Mingyang Song, Shaofeng Du, Xiaohai Wang, Min Zhang, and Youlong Luo. Adaptive priority-based cache replacement and prediction-based cache prefetching in edge computing environment. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301892.

Li:2020:SBU

[2765] Wenjuan Li, Jiao Tan, Weizhi Meng, and Yu Wang. A swipe-based unlocking mechanism with supervised learning on smartphones: Design and evaluation. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301612.

Jiang:2020:SLW

[2766] Shunzhi Jiang, Dengpan Ye, Jiaqing Huang, Yueyun Shang, and Zhuoyuan Zheng. SmartSteganogaphy: Light-weight generative audio steganography model for smart embedding application. Journal of Network and Computer Applications, 165(??):??, September 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520301636.

Anonymous:2020:Sb

[2767] Anonymous. 15 September 2020. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBr 8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520302666.

Nguyen:2020:BBN

[2769] Dinh C. Nguyen, Pubudu N. Pathirana, Ming Ding, and Aruna Seneviratne. Blockchain for 5G and beyond networks: a state of the art survey. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520301673.

Daud:2020:ALP

[2770] Nur Nasuha Daud, Siti Hafizah Ab Hamid, Muntadher Saadoon, Firdaus Sahran, and Nor Badrul Anuar. Applications of link prediction in social networks: a review. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301909.

Alzahrani:2020:UAP

[2771] Bander Alzahrani, Omar Sami Oubbati, Ahmed Barnawi, Mohammed Atiquzzaman, and Daniyal Alghazzawi. UAV assistance paradigm: State-of-the-art in applications and challenges. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301806.

Sellami:2020:UBD

[2772] Mokhtar Sellami, Haithem Mezni, and Mohand Said Hacid. On the use of big data frameworks for big service composition. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452030206X.

Rodriguez-Perez:2020:LES

[2773] Miguel Rodríguez-Pérez, Sergio Herrería-Alonso, Raúl F. Rodríguez-Rubio, and José Carlos López-Ardao. Leveraging energy saving capabilities of current EEE interfaces via pre-coalescing. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302083.

Assis:2020:MTC

[2774] Marcio Roberto Miranda Assis and Luiz Fernando Bittencourt. MultiCloud Tournament: a cloud federation approach to prevent Free-Riders by encouraging resource sharing. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804520301685.

Liu:2020:BBI

[2775] Yang Liu, Debiao He, Mohammad S. Obaidat, Neeraj Kumar, Muhammad Khurram Khan, and Kim-Kwang Raymond Choo. Blockchainbased identity management systems: a review. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302058.

Li:2020:AFU

[2776] Xiaoguo Li, Tao Xiang, and Peng Wang. Achieving forward unforgeability in keyword-field-free conjunctive search. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302290.

Yao:2020:ETC

[2777] Zhongjiang Yao, Jingguo Ge, Yulei Wu, Xiaosheng Lin, Runkang He, and Yuxiang Ma. Encrypted traffic classification based on Gaussian mixture models and Hidden Markov Models. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301855.

Guo:2020:TEM

[2778] Kehua Guo, Haifu Guo, Sheng Ren, Jian Zhang, and Xi Li. Towards efficient motion-blurred public security video super-resolution based on back-projection networks. Journal of Network and Computer Applications, 166(??):??, September 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452030165X.

Anonymous:2020:Oa

[2779] Anonymous. 1 October 2020. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBs

[2780] Anonymous. Editorial Board. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302939.

Fei:2020:ISL

[2781] Huan Fei, Fu Xiao, Haiping Huang, and Lijuan Sun. Indoor static localization based on Fresnel zones model using COTS Wi-Fi. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301831.

Sapna:2020:DDB

[2782] Sapna, K. K. Pattanaik, and Aditya Trivedi. A dynamic distributed boundary node detection algorithm for management zone delineation in Precision Agriculture. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301867.

Chi:2020:SED

[2783] Jiancheng Chi, Yu Li, Jing Huang, Jing Liu, Yingwei Jin, Chen Chen, and Tie Qiu. A secure and efficient data sharing scheme based on blockchain in industrial Internet of Things. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520301843.

Shabani:2020:PLC

[2784] Ahmad Shabani and Bijan Alizadeh. PODEM: a low-cost propertybased design modification for detecting Hardware Trojans in resourceconstraint IoT devices. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301879.

Hussain:2020:RCT

[2785] Zawar Hussain, Quan Z. Sheng, and Wei Emma Zhang. A review and categorization of techniques on device-free human activity recognition. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302125.

Ma:2020:RRA

[2786] Ke Ma, Geng Yang, and Yang Xiang. RCBAC: a risk-aware contentbased access control model for large-scale text data. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302071.

Mansouri:2020:AIH

[2787] Yaser Mansouri, Victor Prokhorenko, and M. Ali Babar. An automated implementation of hybrid cloud for performance evaluation of distributed databases. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302149.

Alkhalaileh:2020:DIA

[2788] Mohammad Alkhalaileh, Rodrigo N. Calheiros, Quang Vinh Nguyen, and Bahman Javadi. Data-intensive application scheduling on Mobile Edge Cloud Computing. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302095.

Jin:2020:SIA

[2789] Li Jin, Dong Changyu, and Palmieric Francesco. Special issue on advanced techniques for security and privacy of Internet-of-Things with machine learning. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520301715.

Ma:2020:SML

[2790] Xu Ma, Cunmei Ji, Xiaoyu Zhang, Jianfeng Wang, Jin Li, Kuan-Ching Li, and Xiaofeng Chen. Secure multiparty learning from the aggregation of locally trained models. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302289.

Ahmed:2020:SCR

[2791] Yahye Abukar Ahmed, Baris Koçer, Shamsul Huda, Bander Ali Saleh Al-rimy, and Mohammad Mehedi Hassan. A system call refinementbased enhanced Minimum Redundancy Maximum Relevance method for ransomware early detection. Journal of Network and Computer Applications, 167(??):??, October 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302277.

Anonymous:2020:Ob

[2792] Anonymous. 15 October 2020. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBt

[2793] Anonymous. Editorial Board. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303003.

Chen:2020:RDN

[2794] Si Chen, Anmin Fu, Jian Shen, Shui Yu, Huaqun Wang, and Huaijiang Sun. RNN-DP: a new differential privacy scheme base on Recurrent Neural Network for Dynamic trajectory privacy protection. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302101.

Kumar:2020:RAA

[2795] Shashwat Kumar, Sai Vineeth Doddala, A. Antony Franklin, and Jiong Jin. RAN-aware adaptive video caching in multi-access edge computing networks. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302113.

Cakmakci:2020:ODA

[2796] Salva Daneshgadeh Çakmakçi, Thomas Kemmerich, Tarem Ahmed, and Nazife Baykal. Online DDoS attack detection using Mahalanobis distance and Kernel-based learning algorithm. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302307.

Mahbub:2020:PRI

[2797] Mobasshir Mahbub. Progressive researches on IoT security: an exhaustive analysis from the perspective of protocols, vulnerabilities, and preemptive architectonics. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302356.

Srivastava:2020:EET

[2798] Akanksha Srivastava, Mani Shekhar Gupta, and Gurjit Kaur. Energy efficient transmission trends towards future green cognitive radio networks (5G): Progress, taxonomy and open challenges. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302344.

Sarker:2020:ARU

[2799] Iqbal H. Sarker and A. S. M. Kayes. ABC-RuleMiner: User behavioral rule-based machine learning method for context-aware intelligent services. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302368.

Liu:2020:FGF

[2800] Wai-Xi Liu, Jun Cai, Yu Wang, Qing Chun Chen, and Jia-Qi Zeng. Fine-grained flow classification using deep learning for software defined data center networks. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030240X.

Elsts:2020:EEA

[2801] Atis Elsts, Niall Twomey, Ryan McConville, and Ian Craddock. Energyefficient activity recognition framework using wearable accelerometers. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302447.

Sharma:2020:CNT

[2802] Abhishek Sharma, Pankhuri Vanjani, Nikhil Paliwal, Chathuranga M. Wijerathna Basnayaka, Dushantha Nalin K. Jayakody, Hwang-Cheng Wang, and P. Muthuchidambaranathan. Communication and networking technologies for UAVs: a survey. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520302137.

Fazio:2020:DSP

[2803] Peppino Fazio, Miralem Mehic, and Miroslav Voznak. A deep stochastical and predictive analysis of users mobility based on Auto-Regressive processes and pairing functions. Journal of Network and Computer Applications, 168(??):??, October 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302526.

Anonymous:2020:Na

[2804] Anonymous. 1 November 2020. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBu

[2805] Anonymous. Editorial Board. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303155.

Fayos-Jordan:2020:PCC

[2806] Rafael Fayos-Jordan, Santiago Felici-Castell, Jaume Segura-Garcia, Jesus Lopez-Ballester, and Maximo Cobos. Performance comparison of container orchestration platforms with low cost devices in the fog, assisting Internet of Things applications. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520302605.

Tulu:2020:INS

[2807] Muluneh Mekonnen Tulu, Mbazingwa E. Mkiramweni, Ronghui Hou, Sultan Feisso, and Talha Younas. Influential nodes selection to enhance data dissemination in mobile social networks: a survey. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302423.

Moura:2020:FCS

[2808] Jose Moura and David Hutchison. Fog computing systems: State of the art, research issues and future trends, with a focus on resilience. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302587.

Merling:2020:PBI

[2809] Daniel Merling, Steffen Lindner, and Michael Menth. P4-based implementation of BIER and BIER-FRR for scalable and resilient multicast. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302381.

Sun:2020:EEC

[2810] Guolin Sun, Daniel Ayepah-Mensah, Rong Xu, Gordon Owusu Boateng, and Guisong Liu. End-to-end CNN-based dueling deep Q-Network for autonomous cell activation in Cloud-RANs. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302319.

Lin:2020:SCO

[2811] Hai Lin, Sherali Zeadally, Zhihong Chen, Houda Labiod, and Lusheng Wang. A survey on computation offloading modeling for edge computing. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302551.

Yang:2020:TRS

[2812] Song Yang, Fan Li, Stojan Trajanovski, and Xiaoming Fu. Traffic routing in stochastic network function virtualization networks. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302393.

Jia:2020:FAL

[2813] Wen-Kang Jia and Xufang Wang. Flow aggregation for large-scale SDNs with scattered address space allocation. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302617.

Yugha:2020:STS

[2814] R. Yugha and S. Chithra. A survey on technologies and security protocols: Reference for future generation IoT. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452030237X.

Gamage:2020:DLM

[2815] Sunanda Gamage and Jagath Samarabandu. Deep learning methods in network intrusion detection: a survey and an objective comparison. Journal of Network and Computer Applications, 169(??):??, November 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302411.

Anonymous:2020:Nb

[2816] Anonymous. 15 November 2020. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBv

[2817] Anonymous. Editorial Board. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303271.

Almasoud:2020:SCB

[2818] Ahmed S. Almasoud, Farookh Khadeer Hussain, and Omar K. Hussain. Smart contracts for blockchain-based reputation systems: a systematic literature review. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030285X.

Cui:2020:DFD

[2819] Lei Cui, Youyang Qu, Longxiang Gao, Gang Xie, and Shui Yu. Detecting false data attacks using machine learning techniques in smart grid: a survey. Journal of Network and Computer Applications, 170 (??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302769.

Uludag:2020:SBB

[2820] Kübra Uludag, Sinan Emre Tasçi, and Omer Korçak. Sequential bargaining based incentive mechanism for collaborative Internet access. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302320.

Alagha:2020:RRF

[2821] Ahmed Alagha, Shakti Singh, Hadi Otrok, and Rabeb Mizouni. RFLS — Resilient Fault-proof Localization System in IoT and Crowd-based Sensing Applications. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302575.

Rahmani:2020:KLD

[2822] Somayeh Rahmani, Vahid Khajehvand, and Mohsen Torabian. Kullback– Leibler distance criterion consolidation in cloud. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302629.

Orts:2020:RRQ

[2823] F. Orts, G. Ortega, E. F. Combarro, and E. M. Garzón. A review on reversible quantum adders. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302812.

Yamansavascilar:2020:FTS

[2824] Baris Yamansavascilar, Ahmet Cihat Baktir, Atay Ozgovde, and Cem Ersoy. Fault tolerance in SDN data plane considering network and application based metrics. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030254X.

Sharma:2020:DSM

[2825] Dilli P. Sharma, Simon Yusuf Enoch, Jin-Hee Cho, Terrence J. Moore, Frederica F. Nelson, Hyuk Lim, and Dong Seong Kim. Dynamic security metrics for software-defined network-based moving target defense. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302770.

Sonkoly:2020:SEC

[2826] Balázs Sonkoly, Dávid Haja, Balázs Németh, Márk Szalay, János Czentye, Róbert Szabó, Rehmat Ullah, Byung-Seo Kim, and László Toka. Scalable edge cloud platforms for IoT services. Journal of Network and Computer Applications, 170(??):??, November 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302599.

Zhu:2020:BMS

[2827] Guangli Zhu, Zhuangzhuang Pan, Qiaoyun Wang, Shunxiang Zhang, and Kuan-Ching Li. Building multi-subtopic bi-level network for microblog hot topic based on feature co-occurrence and semantic community division. Journal of Network and Computer Applications, 170 (??):??, November 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302861.

Anonymous:2020:Da

[2828] Anonymous. 1 December 2020. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBw

[2829] Anonymous. Editorial Board. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303428.

Pan:2020:CCD

[2830] Yan Pan, Shining Li, Yu Zhang, and Ting Zhu. CDA: Coordinating data dissemination and aggregation in heterogeneous IoT networks using CTC. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520302708.

Abdollahi:2020:ORM

[2831] Mostafa Abdollahi, Farshad Eshghi, Manoochehr Kelarestaghi, and Mozafar Bag-Mohammadi. Opportunistic routing metrics: a timely one-stop tutorial survey. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302733.

Hossain:2020:SDD

[2832] Md Akbar Hossain, Sayan Kumar Ray, and Jaswinder Lota. SmartDR: a device-to-device communication for post-disaster recovery. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520302848.

Kaaniche:2020:PET

[2833] Nesrine Kaaniche, Maryline Laurent, and Sana Belguith. Privacy enhancing technologies for solving the privacy-personalization paradox: Taxonomy and survey. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302794.

Yousefnezhad:2020:SPL

[2834] Narges Yousefnezhad, Avleen Malhi, and Kary Främling. Security in product lifecycle of IoT devices: a survey. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302538.

Yang:2020:EPP

[2835] Shu Yang, Lu Bai, Laizhong Cui, Zhongxing Ming, Yulei Wu, Shui Yu, Hongfei Shen, and Yi Pan. An efficient pipeline processing scheme for programming protocol-independent packet processors. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302782.

Gutierrez-Aguado:2020:CBE

[2836] Juan Gutiérrez-Aguado, Raúl Peña-Ortiz, Miguel García-Pineda, and Jose M. Claver. Cloud-based elastic architecture for distributed video encoding: Evaluating H.265, VP9, and AV1. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302563.

Lin:2020:RRE

[2837] Ying-Dar Lin, Yu-Kuen Lai, Quan Tien Bui, and Yuan-Cheng Lai. ReFSM: Reverse engineering from protocol packet traces to test generation by extended finite state machines. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302897.

Mansouri:2020:MOO

[2838] N. Mansouri, Behnam Mohammad Hasani Zade, and Mohammad Masoud Javidi. A multi-objective optimized replication using fuzzy based self-defense algorithm for cloud computing. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302824.

Asaf:2020:BTN

[2839] Khizra Asaf, Rana Asif Rehman, and Byung-Seo Kim. Blockchain technology in Named Data Networks: a detailed survey. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303088.

Fan:2020:PPB

[2840] Yongkai Fan, Jianrong Bai, Xia Lei, Yuqing Zhang, Bin Zhang, Kuan-Ching Li, and Gang Tan. Privacy preserving based logistic regression on big data. Journal of Network and Computer Applications, 171(??):??, December 1, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302435.

Anonymous:2020:Db

[2841] Anonymous. 15 December 2020. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2020:EBx

[2842] Anonymous. Editorial Board. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303635.

Guo:2020:MSC

[2843] Shaoyong Guo, Fengning Wang, Neng Zhang, Feng Qi, and Xuesong Qiu. Master-slave chain based trusted cross-domain authentication mechanism in IoT. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302836.

urRasool:2020:SLF

[2844] Raihan ur Rasool, Hua Wang, Usman Ashraf, Khandakar Ahmed, Zahid Anwar, and Wajid Rafique. A survey of link flooding attacks in software defined network ecosystems. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302757.

Chen:2020:BIT

[2845] Fei Chen, Zhe Xiao, Laizhong Cui, Qiuzhen Lin, Jianqiang Li, and Shui Yu. Blockchain for Internet of Things applications: a review and open issues. Journal of Network and Computer Applications, 172 (??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303076.

Piran:2020:MCC

[2846] Md Jalil Piran, Quoc-Viet Pham, S. M. Riazul Islam, Sukhee Cho, Byungjun Bae, Doug Young Suh, and Zhu Han. Multimedia communication over cognitive radio networks from QoS/QoE perspective: a comprehensive survey. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302332.

Li:2020:TSI

[2847] Ying Li, Jianbo Li, and Manzoor Ahmed. A three-stage incentive formation for optimally pricing social data offloading. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302873.

Lin:2020:IAP

[2848] Yun-Wei Lin, Tai-Hsiang Yen, Cheng-Han Chou, Yi-Bing Lin, and Wen-Shu Lai. Investigating Asian Parti-colored bats using the BatTalk Internet of Things approach. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520302800.

Namasudra:2020:DCT

[2849] Suyel Namasudra, Suraj Sharma, Ganesh Chandra Deka, and Pascal Lorenz. DNA computing and table based data accessing in the cloud environment. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303052.

Baccour:2020:CHC

[2850] Emna Baccour, Aiman Erbad, Amr Mohamed, Mohsen Guizani, and Mounir Hamdi. Collaborative hierarchical caching and transcoding in edge network with CE-D2D communication. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520302745.

Darabkh:2020:JIC

[2851] Khalid A. Darabkh, Oswa M. Amro, Raed T. Al-Zubi, Haythem Bany Salameh, and Ramzi Saifan. JavaSim-IBFD-CRNs: Novel Java simulator for in-band full-duplex cognitive radio networks over Internet of Things environment. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303039.

Yan:2020:ADD

[2852] Zheng Yan, Willy Susilo, Elisa Bertino, Jun Zhang, and Laurence T. Yang. AI-driven data security and privacy. Journal of Network and Computer Applications, 172(??):??, December 15, 2020. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303106.

Anonymous:2021:Ja

[2853] Anonymous. 1 January 2021. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBa

[2854] Anonymous. Editorial Board. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303878.

Anonymous:2021:TYR

[2855] Anonymous. Thank You Reviewers. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452030388X.

Tran:2021:IBI

[2856] Nguyen Khoi Tran, M. Ali Babar, and Jonathan Boan. Integrating blockchain and Internet of Things systems: a systematic review on objectives and designs. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303118.

Yang:2021:SBS

[2857] Dana Yang, Seohee Yoo, Inshil Doh, and Kijoon Chae. Selective blockchain system for secure and efficient D2D communication. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520302885.

Darabkh:2021:YER

[2858] Khalid A. Darabkh, Oswa M. Amro, Raed T. Al-Zubi, and Haythem Bany Salameh. Yet efficient routing protocols for half- and full-duplex cognitive radio Ad-Hoc Networks over IoT environment. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303064.

Shen:2021:WRE

[2859] Zhishu Shen, Tiehua Zhang, Atsushi Tagami, and Jiong Jin. When RSSI encounters deep learning: an area localization scheme for pervasive sensing systems. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303192.

Li:2021:DDN

[2860] Bentian Li, Dechang Pi, Yunxia Lin, and Lin Cui. DNC: a Deep Neural Network-based Clustering-oriented Network Embedding Algorithm. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303209.

Ghotbou:2021:VCC

[2861] Arvin Ghotbou and Mohammad Khansari. VE-CoAP: a constrained application layer protocol for IoT video transmission. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303210.

Popli:2021:ASC

[2862] Sakshi Popli, Rakesh Kumar Jha, and Sanjeev Jain. Adaptive small cell position algorithm (ASPA) for green farming using NB-IoT. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452030309X.

Sun:2021:TGM

[2863] PanJun Sun. A trust game model of service cooperation in cloud computing. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303301. See corrigendum [2888].

Zhang:2021:LAF

[2864] Wenbo Zhang, Xin Wang, Guangjie Han, Yan Peng, Mohsen Guizani, and Jingyi Sun. A load-adaptive fair access protocol for MAC in underwater acoustic sensor networks. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303325.

Fei:2021:ICR

[2865] Xingrui Fei, Yi Xie, Shensheng Tang, and Jiankun Hu. Identifying clickrequests for the network-side through traffic behavior. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303350.

Albahri:2021:IBT

[2866] A. S. Albahri, Jwan K. Alwan, Zahraa K. Taha, Sura F. Ismail, Rula A. Hamid, A. A. Zaidan, O. S. Albahri, B. B. Zaidan, A. H. Alamoodi, and M. A. Alsalem. IoT-based telemedicine for disease prevention and health promotion: State-of-the-art. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303374.

Zhang:2021:UPA

[2867] Ying Zhang, Cui-Hua Zhang, and Xuyang Shao. User preferenceaware navigation for mobile robot in domestic via defined virtual area. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303477.

Khan:2021:HEP

[2868] Ayaz Ali Khan, Muhammad Zakarya, Izaz Ur Rahman, Rahim Khan, and Rajkumar Buyya. HeporCloud: an energy and performance efficient resource orchestrator for hybrid heterogeneous cloud computing environments. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303349.

DAngelo:2021:NTC

[2869] Gianni D’Angelo and Francesco Palmieri. Network traffic classification using deep convolutional recurrent autoencoder neural networks for spatial-temporal features extraction. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520303519.

Garcia:2021:DRT

[2870] Norberto Garcia, Tomas Alcaniz, Aurora González-Vidal, Jorge Bernal Bernabe, Diego Rivera, and Antonio Skarmeta. Distributed real-time SlowDoS attacks detection over encrypted traffic using Artificial Intelligence. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303362.

Chow:2021:UQC

[2871] Yang-Wai Chow, Willy Susilo, Jianfeng Wang, Richard Buckland, Joonsang Baek, Jongkil Kim, and Nan Li. Utilizing QR codes to verify the visual fidelity of image datasets for machine learning. Journal of Network and Computer Applications, 173(??):??, January 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303040.

Anonymous:2021:Jb

[2872] Anonymous. 15 January 2021. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBb

[2873] Anonymous. Editorial Board. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804520304124.

Zhu:2021:PRA

[2874] Ruijie Zhu, Aretor Samuel, Peisen Wang, Shihua Li, Bounsou Kham Oun, Lulu Li, Pei Lv, Mingliang Xu, and Shui Yu. Protected resource allocation in space division multiplexing-elastic optical networks with fluctuating traffic. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303507.

Sadri:2021:FDM

[2875] Ali Akbar Sadri, Amir Masoud Rahmani, Morteza Saberikamarposhti, and Mehdi Hosseinzadeh. Fog data management: a vision, challenges, and future directions. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303465.

Hamdan:2021:CSL

[2876] Mosab Hamdan, Entisar Hassan, Ahmed Abdelaziz, Abdallah Elhigazi, Bushra Mohammed, Suleman Khan, Athanasios V. Vasilakos, and M. N. Marsono. A comprehensive survey of load balancing techniques in software-defined network. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303222.

Maranhao:2021:TBF

[2877] João Paulo A. Maranhão, João Paulo C. L. da Costa, Elnaz Javidi, César A. Borges de Andrade, and Rafael T. de Sousa. Tensor based framework for Distributed Denial of Service attack detection. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303556.

Fraire:2021:RSI

[2878] Juan A. Fraire, Olivier De Jonckère, and Scott C. Burleigh. Routing in the Space Internet: a contact graph routing tutorial. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303489.

Colombo:2021:RDS

[2879] Pietro Colombo, Elena Ferrari, and Engin Deniz Tümer. Regulating data sharing across MQTT environments. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303660.

Beg:2021:PPP

[2880] Saira Beg, Adeel Anjum, Mansoor Ahmad, Shahid Hussain, Ghufran Ahmad, Suleman Khan, and Kim-Kwang Raymond Choo. A privacypreserving protocol for continuous and dynamic data collection in IoT enabled mobile app recommendation system (MARS). Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303386.

Fotouhi:2021:DIS

[2881] Azade Fotouhi, Ming Ding, and Mahbub Hassan. DroneCells: Improving spectral efficiency using drone-mounted flying base stations. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520303568.

Zhang:2021:PPC

[2882] Xiaoshuai Zhang, Chao Liu, Kok Keong Chai, and Stefan Poslad. A privacy-preserving consensus mechanism for an electric vehicle charging scheme. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303684.

Ma:2021:QOJ

[2883] Zhongyu Ma, Jie Cao, Qun Guo, Xiangwei Li, and Hongfeng Ma. QoSoriented joint optimization of concurrent scheduling and power control in millimeter wave mesh backhaul network. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303532.

Lu:2021:FFA

[2884] Yifei Lu, Xu Ma, and Zhengzhi Xu. FAMD: a Flow-Aware Marking and Delay-based TCP algorithm for datacenter networks. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303714.

Rodrigues:2021:EBE

[2885] Carlo K. da S. Rodrigues and Vladimir Rocha. Enhancing BitTorrent for efficient interactive video-on-demand streaming over MANETs. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520303672.

Aman:2021:IAC

[2886] Azana Hafizah Mohd Aman, Wan Haslina Hassan, Shilan Sameen, Zainab Senan Attarbashi, Mojtaba Alizadeh, and Liza Abdul Latiff.

IoMT amid COVID-19 pandemic: Application, architecture, technology, and security. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303490.

Sciancalepore:2021:RLP

[2887] Savio Sciancalepore, Saeif Alhazbi, and Roberto Di Pietro. Receivers location privacy in avionic crowdsourced networks: Issues and countermeasures. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303544.

Sun:2021:CTG

[2888] Pan Jun Sun. Corrigendum to “A trust game model of service cooperation in cloud computing” [J. Netw. Comput. Appl. 173 (20) 2020 102864]. Journal of Network and Computer Applications, 174(??):??, January 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303593. See [2863].

Anonymous:2021:Fa

[2889] Anonymous. 1 February 2021. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBc

[2890] Anonymous. Editorial Board. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000060.

Khalid:2021:SPT

[2891] Muhammad Khalid, Kezhi Wang, Nauman Aslam, Yue Cao, Naveed Ahmad, and Muhammad Khurram Khan. From smart parking towards autonomous valet parking: a survey, challenges and future works. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303921.

Santos:2021:TEE

[2892] José Santos, Tim Wauters, Bruno Volckaert, and Filip De Turck. Towards end-to-end resource provisioning in fog computing over low power wide area networks. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030374X.

Israr:2021:REP

[2893] Adil Israr, Qiang Yang, Wei Li, and Albert Y. Zomaya. Renewable energy powered sustainable 5G network infrastructure: Opportunities, challenges and perspectives. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303702.

Yang:2021:FGR

[2894] Yang Yang, Songtao Guo, Guiyan Liu, and Lin Yi. Fine granularity resource allocation of virtual data center with consideration of virtual switches. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303751.

Borylo:2021:USM

[2895] Piotr Borylo, Gianluca Davoli, Michal Rzepka, Artur Lason, and Walter Cerroni. Unified and standalone monitoring module for NFV/ SDN infrastructures. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303933.

Boronat:2021:WWP

[2896] Fernando Boronat, Mario Montagud, Pau Salvador, and Javier Pastor. Wersync: a web platform for synchronized social viewing enabling interaction and collaboration. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303970.

Serrano:2021:BRN

[2897] Will Serrano. The Blockchain Random Neural Network for cybersecure IoT and 5G infrastructure in Smart Cities. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303696.

Jan:2021:SBC

[2898] Mian Ahmad Jan, Jinjin Cai, Xiang-Chuan Gao, Fazlullah Khan, Spyridon Mastorakis, Muhammad Usman, Mamoun Alazab, and Paul Watters. Security and blockchain convergence with Internet of Multimedia Things: Current trends, research challenges and future directions. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303775.

Melo:2021:FSF

[2899] Wilson S. Melo, Luiz V. G. Tarelho, Bruno A. Rodrigues, Alysson N. Bessani, and Luiz F. R. C. Carmo. Field surveillance of fuel dispensers using IoT-based metering and blockchains. Journal of Network and Computer Applications, 175(??):??, February 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303738.

Anonymous:2021:Fb

[2900] Anonymous. 15 February 2021. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBd

[2901] Anonymous. Editorial Board. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521000175.

Hernandez-Quintanilla:2021:RAD

[2902] Tomas Hernandez-Quintanilla, Eduardo Magaña, Daniel Morató, and Mikel Izal. On the reduction of authoritative DNS cache timeouts: Detection and implications for user privacy. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303994.

Maleki:2021:SSP

[2903] Neda Maleki, Amir Masoud Rahmani, and Mauro Conti. SPO: a Secure and Performance-aware Optimization for MapReduce Scheduling. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452030401X.

Xu:2021:STC

[2904] Cheng Xu, Xinxin Wang, Shihong Duan, and Jiawang Wan. Spatialtemporal constrained particle filter for cooperative target tracking. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303726.

Yang:2021:SNF

[2905] Liang Yang, Bryan Ng, Winston K. G. Seah, Lindsay Groves, and Deepak Singh. A survey on network forwarding in Software-Defined Networking. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304021.

Kolomvatsos:2021:PTM

[2906] Kostas Kolomvatsos. Proactive tasks management for Pervasive Computing Applications. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520304033.

Xu:2021:NAT

[2907] Jie Xu, Jingyu Wang, Qi Qi, Jianxin Liao, Haifeng Sun, Zhu Han, and Tonghong Li. Network-aware task selection to reduce multi-application makespan in cloud. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303520.

Xiang:2021:MTD

[2908] Yuexin Xiang, Wei Ren, Tiantian Li, Xianghan Zheng, Tianqing Zhu, and Kim-Kwang Raymond Choo. A multi-type and decentralized data transaction scheme based on smart contracts and digital watermarks. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304057.

Loffi:2021:MAM

[2909] Leandro Loffi, Carla Merkle Westphall, Lukas Derner Grüdtner, and Carlos Becker Westphall. Mutual authentication with multi-factor in IoT– Fog–Cloud environment. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452030391X.

Manzoor:2021:PRE

[2910] Ahsan Manzoor, An Braeken, Salil S. Kanhere, Mika Ylianttila, and Madhsanka Liyanage. Proxy re-encryption enabled secure and anonymous IoT data sharing platform based on blockchain. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303763.

Sun:2021:RCR

[2911] Lijun Sun, Qian Yang, Xiao Chen, and Zhenxiang Chen. RCchain: Reputation-based crowdsourcing blockchain for vehicular networks. Journal of Network and Computer Applications, 176(??):??, February 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304082.

Anonymous:2021:Ma

[2912] Anonymous. 1 March 2021. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBe

[2913] Anonymous. Editorial Board. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000266.

Berrocal-Plaza:2021:ACM

[2914] Víctor Berrocal-Plaza and Miguel A. Vega-Rodríguez. Analysis and comparison of mobility management strategies in public land mobile networks from a multiobjective perspective. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520304173.

Nerurkar:2021:DBB

[2915] Pranav Nerurkar, Dhiren Patel, Yann Busnel, Romaric Ludinard, Saru Kumari, and Muhammad Khurram Khan. Dissecting bitcoin blockchain: Empirical analysis of bitcoin network (2009–2020). Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520303982.

Hewa:2021:SBB

[2916] Tharaka Hewa, Mika Ylianttila, and Madhusanka Liyanage. Survey on blockchain based smart contracts: Applications, opportunities and challenges. Journal of Network and Computer Applications, 177(??): ??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520303234.

Liu:2021:DRD

[2917] Wai xi Liu, Jun Cai, Qing Chun Chen, and Yu Wang. DRL-R: Deep reinforcement learning approach for intelligent routing in softwaredefined data-center networks. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303313.

Alagha:2021:SSD

[2918] Ahmed Alagha, Rabeb Mizouni, Shakti Singh, Hadi Otrok, and Anis Ouali. SDRS: a stable data-based recruitment system in IoT crowdsensing for localization tasks. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520304197.

Assis:2021:GDL

[2919] Marcos V. O. Assis, Luiz F. Carvalho, Jaime Lloret, and Mario L. Proença. A GRU deep learning system against attacks in software defined networks. Journal of Network and Computer Applications, 177(??): ??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304008.

Yi:2021:CDE

[2920] Bo Yi, Xingwei Wang, and Min Huang. Content delivery enhancement in Vehicular Social Network with better routing and caching mechanism. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304069.

Xiao:2021:MJF

[2921] Yang Xiao, Qingqi Pei, Tingting Xiao, Lina Yao, and Huan Liu. MutualRec: Joint friend and item recommendations with mutualistic attentional graph neural networks. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520304070.

Li:2021:JEC

[2922] Chunlin Li, Lei Zhu, Weigang Li, and Youlong Luo. Joint edge caching and dynamic service migration in SDN based mobile edge computing. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304185.

Farahani:2021:CID

[2923] Bahar Farahani, Farshad Firouzi, and Markus Luecking. The convergence of IoT and distributed ledger technologies (DLT): Opportunities, challenges, and solutions. Journal of Network and Computer Applications, 177(??):??, March 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520303945.

Anonymous:2021:Mb

[2924] Anonymous. 15 March 2021. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBf

[2925] Anonymous. Editorial Board. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000400.

Chen:2021:MAR

[2926] Yihong Chen, Quanyuan Feng, Xiaolin Jia, and Huayue Chen. Modeling and analyzing RFID Generation-2 under unreliable channels. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804520303957.

Qiu:2021:LDE

[2927] Liqing Qiu, Xiangbo Tian, Jianyi Zhang, Chunmei Gu, and Shiqi Sai. LIDDE: a differential evolution algorithm based on local-influence-descending search strategy for influence maximization in social networks. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304240.

Feng:2021:SCE

[2928] Wendi Feng, Chuanchang Liu, Bo Cheng, and Junliang Chen. Secure and cost-effective controller deployment in multi-domain SDN with Baguette. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304203.

Shakarami:2021:ACO

[2929] Ali Shakarami, Ali Shahidinejad, and Mostafa Ghobaei-Arani. An autonomous computation offloading strategy in Mobile Edge Computing: a deep learning-based hybrid approach. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521000011.

Sookhak:2021:BSC

[2930] Mehdi Sookhak, Mohammad Reza Jabbarpour, Nader Sohrabi Safa, and F. Richard Yu. Blockchain and smart contract for access control in healthcare: a survey, issues and challenges, and open issues. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804520304045.

Natesha:2021:AEB

[2931] B. V. Natesha and Ram Mohana Reddy Guddeti. Adopting elitismbased Genetic Algorithm for minimizing multi-objective problems of IoT service placement in fog computing environment. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804520304239.

Hu:2021:VRS

[2932] Miao Hu, Xianzhuo Luo, Jiawen Chen, Young Choon Lee, Yipeng Zhou, and Di Wu. Virtual reality: a survey of enabling technologies and its applications in IoT. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520304215.

Aboutorab:2021:SSR

[2933] Hamed Aboutorab, Omar K. Hussain, Morteza Saberi, Farookh Khadeer Hussain, and Elizabeth Chang. A survey on the suitability of risk identification techniques in the current networked environment. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000114.

Mohammadi:2021:CST

[2934] Mokhtar Mohammadi, Tarik A. Rashid, Sarkhel H. Taher Karim, Adil Hussain Mohammed Aldalwie, Quan Thanh Tho, Moazam Bidaki, Amir Masoud Rahmani, and Mehdi Hosseinzadeh. A comprehensive survey and taxonomy of the SVM-based intrusion detection systems. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000102.

Shuvo:2021:EES

[2935] Md. Shahin Alom Shuvo, Md. Azad Rahaman Munna, Sujan Sarker, Tamal Adhikary, Md. Abdur Razzaque, Mohammad Mehedi Hassan, Gianluca Aloi, and Giancarlo Fortino. Energy-efficient scheduling of small cells in 5G: a meta-heuristic approach. Journal of Network and Computer Applications, 178(??):??, March 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521000138.

Anonymous:2021:Aa

[2936] Anonymous. 1 April 2021. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBg

[2937] Anonymous. Editorial Board. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000588.

Manocha:2021:DCI

[2938] Ankush Manocha, Munish Bhatia, and Gulshan Kumar. Dew computinginspired health-meteorological factor analysis for early prediction of bronchial asthma. Journal of Network and Computer Applications, 179 (??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000229.

Jia:2021:ROI

[2939] Jie Jia, Jian Chen, Yansha Deng, Xingwei Wang, and Abdol-Hamid Aghvami. Reliability optimization for industrial WSNs with FD relays and multiple parallel connections. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000205.

Yang:2021:TRE

[2940] Yating Yang, Tian Song, Weijia Yuan, and Jianping An. Towards reliable and efficient data retrieving in ICN-based satellite networks. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000096.

Kumar:2021:MPC

[2941] Mahender Kumar and Satish Chand. MedHypChain: a patient-centered interoperability hyperledger-based medical healthcare system: Regulation in COVID-19 pandemic. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000023.

Hanif:2021:RSV

[2942] Hazim Hanif, Mohd Hairul Nizam Md Nasir, Mohd Faizal Ab Razak, Ahmad Firdaus, and Nor Badrul Anuar. The rise of software vulnerability: Taxonomy of software vulnerabilities detection and machine learning approaches. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000369.

Saad:2021:CMC

[2943] Muhammad Saad, Joongheon Kim, DaeHun Nyang, and David Mohaisen. Contra-*: Mechanisms for countering spam attacks on blockchain’s memory pools. Journal of Network and Computer Applications, 179(??):??, April 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804520304227.

Anonymous:2021:Ab

[2944] Anonymous. 15 April 2021. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBh

[2945] Anonymous. Editorial Board. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000692.

Aburukba:2021:HSA

[2946] Raafat O. Aburukba, Taha Landolsi, and Dalia Omer. A heuristic scheduling approach for fog–cloud computing environment with stationary IoT devices. Journal of Network and Computer Applications, 180(??): ??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000217.

Islam:2021:CAS

[2947] Mir Salim Ul Islam, Ashok Kumar, and Yu-Chen Hu. Context-aware scheduling in Fog computing: a survey, taxonomy, challenges and future directions. Journal of Network and Computer Applications, 180(??): ??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000357.

Li:2021:CAA

[2948] Chunlin Li, Jun Liu, Bo Lu, and Youlong Luo. Cost-aware automatic scaling and workload-aware replica management for edge-cloud environment. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000448.

Liu:2021:DDA

[2949] Xiaojie Liu, Xingwei Wang, Jie Jia, and Min Huang. A distributed deployment algorithm for communication coverage in wireless robotic networks. Journal of Network and Computer Applications, 180(??): ??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000461.

Erdin:2021:SPB

[2950] Enes Erdin, Mumin Cebe, Kemal Akkaya, Eyuphan Bulut, and Selcuk Uluagac. A scalable private Bitcoin payment channel network with privacy guarantees. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000485.

Sadegh:2021:TPV

[2951] Samaneh Sadegh, Kamran Zamanifar, Piotr Kasprzak, and Ramin Yahyapour. A two-phase virtual machine placement policy for dataintensive applications in cloud. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000515.

Dorri:2021:TIR

[2952] Ali Dorri, Fengji Luo, Samuel Karumba, Salil Kanhere, Raja Jurdak, and Zhao Yang Dong. Temporary immutability: a removable blockchain solution for prosumer-side energy trading. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452100045X.

Montanha:2021:NSL

[2953] Aleksandro Montanha, Airton M. Polidorio, and María del Carmen Romero-Ternero. New signal location method based on signal-range data for proximity tracing tools. Journal of Network and Computer Applications, 180(??):??, April 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000333.

Anonymous:2021:Mc

[2954] Anonymous. 1 May 2021. Journal of Network and Computer Applications, 181(??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBi

[2955] Anonymous. Editorial Board. Journal of Network and Computer Applications, 181(??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000874.

Majeed:2021:BIB

[2956] Umer Majeed, Latif U. Khan, Ibrar Yaqoob, S. M. Ahsan Kazmi, Khaled Salah, and Choong Seon Hong. Blockchain for IoT-based smart cities: Recent advances, requirements, and future challenges. Journal of Network and Computer Applications, 181(??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000345.

Eghbali:2021:HAA

[2957] Zahra Eghbali and Mina Zolfy Lighvan. A hierarchical approach for accelerating IoT data management process based on SDN principles. Journal of Network and Computer Applications, 181(??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000539.

Shuja:2021:AML

[2958] Junaid Shuja, Kashif Bilal, Waleed Alasmary, Hassan Sinky, and Eisa Alanazi. Applying machine learning techniques for caching in nextgeneration edge networks: a comprehensive survey. Journal of Network and Computer Applications, 181(??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000321.

Bolettieri:2021:AAR

[2959] Simone Bolettieri, Raffaele Bruno, and Enzo Mingozzi. Applicationaware resource allocation and data management for MEC-assisted IoT service providers. Journal of Network and Computer Applications, 181 (??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000473.

Saxena:2021:BBS

[2960] Shivam Saxena, Bharat Bhushan, and Mohd Abdul Ahad. Blockchain based solutions to secure IoT: Background, integration trends and a way forward. Journal of Network and Computer Applications, 181(??): ??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000758.

Kim:2021:APM

[2961] Jonghyun Kim, Pratik K. Biswas, Stephan Bohacek, Sharon J. Mackey, Siamak Samoohi, and Mitesh P. Patel. Advanced protocols for the mitigation of friendly jamming in mobile ad-hoc networks. Journal of Network and Computer Applications, 181(??):??, May 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000631.

Anonymous:2021:Md

[2962] Anonymous. 15 May 2021. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBj

[2963] Anonymous. Editorial Board. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000941.

Tian:2021:WFH

[2964] Le Tian, Serena Santi, Amina Seferagić, Julong Lan, and Jeroen Famaey. Wi-Fi HaLow for the Internet of Things: an up-to-date survey on IEEE 802.11ah research. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452100062X.

Ren:2021:AMO

[2965] Tao Ren, Jianwei Niu, Jiahe Cui, Zhenchao Ouyang, and Xuefeng Liu. An application of multi-objective reinforcement learning for efficient model-free control of canals deployed with IoT networks. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000734.

Ferdous:2021:SCA

[2966] Md Sadek Ferdous, Mohammad Jabed Morshed Chowdhury, and Mohammad A. Hoque. A survey of consensus algorithms in public blockchain systems for crypto-currencies. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804521000618.

Podili:2021:TTA

[2967] Prashanth Podili and Kotaro Kataoka. TRAQR: Trust aware End-to-End QoS routing in multi-domain SDN using Blockchain. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000795.

Feng:2021:UHE

[2968] Sheng Feng, Haiyan Shi, Longjun Huang, Shigen Shen, Shui Yu, Hua Peng, and Chengdong Wu. Unknown hostile environment-oriented autonomous WSN deployment using a mobile robot. Journal of Network and Computer Applications, 182(??):??, May 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000783.

Anonymous:2021:Jc

[2969] Anonymous. 1 June 2021. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBk

[2970] Anonymous. Editorial Board. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001119.

Aceto:2021:DET

[2971] Giuseppe Aceto, Domenico Ciuonzo, Antonio Montieri, and Antonio Pescapé. DISTILLER: Encrypted traffic classification via multimodal multitask deep learning. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000126.

Zhao:2021:CCE

[2972] Quanjun Zhao, Songtao Guo, and Guiyan Liu. Cost EACP: Costeffective adaptive controller provisioning in software defined DCNs. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804521000801.

Javaid:2021:TAR

[2973] Shumaila Javaid, Zhenqiang Wu, Zara Hamid, Sherali Zeadally, and Hamza Fahim. Temperature-aware routing protocol for Intrabody Nanonetworks. Journal of Network and Computer Applications, 183–184 (??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521000813.

Rui:2021:SMM

[2974] LanLan Rui, Menglei Zhang, Zhipeng Gao, Xuesong Qiu, Zhili Wang, and Ao Xiong. Service migration in multi-access edge computing: a joint state adaptation and reinforcement learning mechanism. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/

pii/S1084804521000825. See corrigendum [2987].

Jayalekshmi:2021:GRG

[2975] S. Jayalekshmi and R. Leela Velusamy. GSA-RPI: GSA based Rendezvous Point Identification in a two-level cluster based LR-WPAN for uncovering the optimal trajectory of Mobile Data Collection Agent. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804521000746.

Mohamadi:2021:FFA

[2976] Mohamed Mohamadi, Badis Djamaa, Mustapha Reda Senouci, and Abdelhamid Mellouk. FAN: Fast and Active Network Formation in IEEE 802.15.4 TSCH Networks. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000527.

Muzumdar:2021:TIS

[2977] Ajit Muzumdar, Chirag Modi, Madhu G. M., and C. Vyjayanthi. A trustworthy and incentivized smart grid energy trading framework using distributed ledger and smart contracts. Journal of Network and Computer Applications, 183–184(??):??, June 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000977.

Anonymous:2021:Jd

[2978] Anonymous. 1 July 2021. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBl

[2979] Anonymous. Editorial Board. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001223.

Abu-Salih:2021:DSK

[2980] Bilal Abu-Salih. Domain-specific knowledge graphs: a survey. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521000990.

Liu:2021:FAP

[2981] Xinqian Liu, Jiadong Ren, Haitao He, Bing Zhang, Chen Song, and Yunxue Wang. A fast all-packets-based DDoS attack detection approach based on network graph and graph kernel. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001016.

Ho:2021:DGD

[2982] Stella Ho, Youyang Qu, Bruce Gu, Longxiang Gao, Jianxin Li, and Yong Xiang. DP-GAN: Differentially private consecutive data publishing using generative adversarial nets. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000904.

Nayeri:2021:APF

[2983] Zahra Makki Nayeri, Toktam Ghafarian, and Bahman Javadi. Application placement in Fog computing with AI approach: Taxonomy and a state of the art survey. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521000989.

Shi:2021:WCB

[2984] Leyi Shi, Xiaoyu Li, Zhenbo Gao, Pengfei Duan, Na Liu, and Honglong Chen. Worm computing: a blockchain-based resource sharing and cybersecurity framework. Journal of Network and Computer Applications, 185(??):??, July 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452100103X.

Anonymous:2021:J

[2985] Anonymous. 15 July 2021. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBm

[2986] Anonymous. Editorial Board. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001417.

Rui:2021:CSM

[2987] LanLan Rui, Menglei Zhang, Zhipeng Gao, Xuesong Qiu, Zhili Wang, and Ao Xiong. Corrigendum to “Service migration in multi-access edge computing: a joint state adaptation and reinforcement learning mechanism” [J. Netw. Comput. Appl. 183–184 (2021) 103058]. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804521001077. See [2974].

Park:2021:PFT

[2988] Pangun Park, Hossein S. Ghadikolaei, and Carlo Fischione. Proactive fault-tolerant wireless mesh networks for mission-critical control systems. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001041.

Wu:2021:LBF

[2989] Qingtao Wu, Qianyu Wang, Mingchuan Zhang, Ruijuan Zheng, Junlong Zhu, and Jiankun Hu. Learned Bloom-filter for the efficient name lookup in Information-Centric Networking. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001004.

Mazumdar:2021:AHD

[2990] Nabajyoti Mazumdar, Saugata Roy, Amitava Nag, and Sukumar Nandi. An adaptive hierarchical data dissemination mechanism for mobile data collector enabled dynamic wireless sensor network. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001181.

Zareie:2021:MSM

[2991] Ahmad Zareie and Rizos Sakellariou. Minimizing the spread of misinformation in online social networks: a survey. Journal of Network and Computer Applications, 186(??):??, July 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001168.

Anonymous:2021:Ac

[2992] Anonymous. 1 August 2021. Journal of Network and Computer Applications, 187(??):??, August 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBn

[2993] Anonymous. Editorial Board. Journal of Network and Computer Applications, 187(??):??, August 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001491.

Raj:2021:SRI

[2994] Meghna Raj, Shashank Gupta, Vinay Chamola, Anubhav Elhence, Tanya Garg, Mohammed Atiquzzaman, and Dusit Niyato. A survey on the role of Internet of Things for adopting and promoting Agriculture 4.0. Journal of Network and Computer Applications, 187(??):??, August 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001284.

Ahuja:2021:ADA

[2995] Nisha Ahuja, Gaurav Singal, Debajyoti Mukhopadhyay, and Neeraj Kumar. Automated DDOS attack detection in software defined networking. Journal of Network and Computer Applications, 187(??):??, August 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001296.

Valdovinos:2021:EDA

[2996] Ismael Amezcua Valdovinos, Jesús Arturo Pérez-Díaz, Kim-Kwang Raymond Choo, and Juan Felipe Botero. Emerging DDoS attack detection and mitigation strategies in software-defined networks: Taxonomy, challenges and future directions. Journal of Network and Computer Applications, 187(??):??, August 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001156.

Lee:2021:TSI

[2997] Sang-Woong Lee, Haval Mohammed sidqi, Mokhtar Mohammadi, Shima Rashidi, Amir Masoud Rahmani, Mohammad Masdari, and Mehdi Hosseinzadeh. Towards secure intrusion detection systems using deep learning techniques: Comprehensive analysis and review. Journal of Network and Computer Applications, 187(??):??, August 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001314.

Anonymous:2021:Ad

[2998] Anonymous. 15 August 2021. Journal of Network and Computer Applications, 188(??):??, August 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBo

[2999] Anonymous. Editorial Board. Journal of Network and Computer Applications, 188(??):??, August 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001636.

Pundir:2021:SRQ

[3000] Meena Pundir and Jasminder Kaur Sandhu. A systematic review of quality of service in wireless sensor networks using machine learning: Recent trend and future vision. Journal of Network and Computer Applications, 188(??):??, August 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001065.

Wang:2021:ADU

[3001] Xuerui Wang, Zheng Yan, Rui Zhang, and Peng Zhang. Attacks and defenses in user authentication systems: a survey. Journal of Network and Computer Applications, 188(??):??, August 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001028.

Zhao:2021:LLD

[3002] Xu Zhao, Guangqiu Huang, Ling Gao, Maozhen Li, and Quanli Gao. Low load DIDS task scheduling based on Q-learning in edge computing environment. Journal of Network and Computer Applications, 188(??):??, August 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452100117X.

Anonymous:2021:Sa

[3003] Anonymous. 1 September 2021. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBp 8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001843.

Jian:2021:EEU

[3005] Xin Jian, Langyun Wu, Keping Yu, Moayad Aloqaily, and Jalel Ben-Othman. Energy-efficient user association with load-balancing for cooperative IIoT network within B5G era. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804521001302.

Guo:2021:TEH

[3006] Yingya Guo, Weipeng Wang, Han Zhang, Wenzhong Guo, Zhiliang Wang, Ying Tian, Xia Yin, and Jianping Wu. Traffic engineering in hybrid software defined network via reinforcement learning. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001363.

Zhang:2021:EEC

[3007] Zhiwei Zhang, Lin Cui, and Jia Wu. Exploring an edge convolution and normalization based approach for link prediction in complex networks. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001338.

Shahzadi:2021:UAB

[3008] Rizwana Shahzadi, Mudassar Ali, Humayun Zubair Khan, and Muhammad Naeem. UAV assisted 5G and beyond wireless networks: a survey. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452100134X.

Zoppi:2021:PTM

[3009] Tommaso Zoppi and Andrea Ceccarelli. Prepare for trouble and make it double! Supervised–unsupervised stacking for anomaly-based intrusion detection. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001272.

Do:2021:SBW

[3010] Hoang Minh Do, Mark A. Gregory, and Shuo Li. SDN-based wireless mobile backhaul architecture: Review and challenges. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001569.

Liu:2021:HCD

[3011] Yuan Liu, Rentao Gu, Zeyuan Yang, and Yuefeng Ji. Hierarchical community discovery for multi-stage IP bearer network upgradation. Journal of Network and Computer Applications, 189(??):??, September 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001661.

Anonymous:2021:Sb

[3012] Anonymous. 15 September 2021. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBq

[3013] Anonymous. Editorial Board. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001971.

Yousefpoor:2021:SDA

[3014] Mohammad Sadegh Yousefpoor, Efat Yousefpoor, Hamid Barati, Ali Barati, Ali Movaghar, and Mehdi Hosseinzadeh. Secure data aggregation methods and countermeasures against various attacks in wireless sensor networks: a comprehensive review. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001375.

Khan:2021:FNO

[3015] Tanveer Khan, Antonis Michalas, and Adnan Akhunzada. Fake news outbreak 2021: Can we stop the viral spread? Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001326.

Wu:2021:ACT

[3016] Jiajing Wu, Jieli Liu, Yijing Zhao, and Zibin Zheng. Analysis of cryptocurrency transactions from a network perspective: an overview. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001557.

Nguyen:2021:ORU

[3017] Hai T. Nguyen, Tien V. Do, and Csaba Rotter. Optimizing the resource usage of actor-based systems. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001594.

Zaraket:2021:OLE

[3018] Elie Zaraket, Nour Mohammad Murad, Samar Syed Yazdani, Lala Rajaoarisoa, and Blaise Ravelo. An overview on low energy wake-up radio technology: Active and passive circuits associated with MAC and routing protocols. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001570.

AlAhmad:2021:MCC

[3019] Ahmad Salah AlAhmad, Hasan Kahtan, Yehia Ibrahim Alzoubi, Omar Ali, and Ashraf Jaradat. Mobile cloud computing models security issues: a systematic review. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001673.

Casino:2021:IHH

[3020] Fran Casino, Nikolaos Lykousas, Ivan Homoliak, Constantinos Patsakis, and Julio Hernandez-Castro. Intercepting Hail Hydra: Real-time detection of algorithmically generated domains. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001545.

Yang:2021:AGS

[3021] Hui Yang, Kaixuan Zhan, Bowen Bao, Qiuyan Yao, Jie Zhang, and Mohamed Cheriet. Automatic guarantee scheme for intent-driven network slicing and reconfiguration. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001752.

Cui:2021:TDD

[3022] Yunhe Cui, Qing Qian, Chun Guo, Guowei Shen, Youliang Tian, Huanlai Xing, and Lianshan Yan. Towards DDoS detection mechanisms in software-defined networking. Journal of Network and Computer Applications, 190(??):??, September 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001703.

Anonymous:2021:Oa

[3023] Anonymous. 1 October 2021. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBr

[3024] Anonymous. Editorial Board. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452100206X.

Pham:2021:SIN

[3025] Quoc-Viet Pham, Dinh C. Nguyen, Seyedali Mirjalili, Dinh Thai Hoang, Diep N. Nguyen, Pubudu N. Pathirana, and Won-Joo Hwang. Swarm intelligence for next-generation networks: Recent advances and applications. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001582.

Huang:2021:PPA

[3026] Yan Huang, Wei Li, Jinbao Wang, Zhipeng Cai, and Anu G. Bourgeois. Privacy protection among three antithetic-parties for contextaware services. Journal of Network and Computer Applications, 191(??): ??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001351.

Jemili:2021:CLM

[3027] Imen Jemili, Dhouha Ghrab, Abdelfettah Belghith, Mohamed Mosbah, and Saad Al-Ahmadi. Cross-layer multipath approach for critical traffic in duty-cycled wireless sensor networks. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001685.

Kadadha:2021:TSP

[3028] Maha Kadadha, Hadi Otrok, Shakti Singh, Rabeb Mizouni, and Anis Ouali. Two-sided preferences task matching mechanisms for blockchainbased crowdsourcing. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001697.

Zaidi:2021:TBC

[3029] Ahmad Zairi Zaidi, Chun Yong Chong, Zhe Jin, Rajendran Parthiban, and Ali Safaa Sadiq. Touch-based continuous mobile device authentication: State-of-the-art, challenges and opportunities. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001740.

Baswade:2021:PAS

[3030] Anand M. Baswade, Mohith Reddy, Antony Franklin A., Bheemarjuna Reddy Tamma, and Vanlin Sathya. Performance analysis of spatially distributed LTE-U/NR-U and Wi-Fi networks: an analytical model for coexistence study. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001715.

Chen:2021:FRA

[3031] Longbiao Chen, Zhihan Jiang, Dingqi Yang, Cheng Wang, and Thi-Mai-Trang Nguyen. Fog radio access network optimization for 5G leveraging user mobility and traffic data. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001053.

Li:2021:OCC

[3032] Chunlin Li, Yong Zhang, Mingyang Song, Xin Yan, and Youlong Luo. An optimized content caching strategy for video stream in edge-cloud environment. Journal of Network and Computer Applications, 191(??): ??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001727.

ElSayed:2021:NHM

[3033] Mahmoud Said ElSayed, Nhien-An Le-Khac, Marwan Ali Albahar, and Anca Jurcut. A novel hybrid model for intrusion detection systems in SDNs based on CNN and a new regularization technique. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001739.

Agate:2021:SSO

[3034] Vincenzo Agate, Alessandra De Paola, Pierluca Ferraro, Giuseppe Lo Re, and Marco Morana. SecureBallot: a secure open source e-Voting system. Journal of Network and Computer Applications, 191(??):??, October 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521001776.

Anonymous:2021:Ob

[3035] Anonymous. 15 October 2021. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBs

[3036] Anonymous. Editorial Board. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521002204.

Noor:2021:SSI

[3037] Jannatun Noor, Saiful Islam Salim, and A. B. M. Alim Al Islam. Strategizing secured image storing and efficient image retrieval through a new cloud framework. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001788.

Alsahlani:2021:LIL

[3038] Ahmed Yaser Fahad Alsahlani and Alexandru Popa. LMAAS-IoT: Lightweight multi-factor authentication and authorization scheme for real-time data access in IoT cloud-based environment. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001879.

Liu:2021:RTB

[3039] Sen Liu, Jiawei Huang, Wenchao Jiang, and Jianxin Wang. Reducing traffic burstiness for MPTCP in data center networks. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001806.

Santos:2021:TCN

[3040] Fillipe Santos, Andre L. L. Aquino, Edmundo R. M. Madeira, and Raquel S. Cabral. Temporal complex networks modeling applied to vehicular ad-hoc networks. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452100179X.

Bouzidi:2021:DQN

[3041] EL Hocine Bouzidi, Abdelkader Outtagarts, Rami Langar, and Raouf Boutaba. Deep Q-network and traffic prediction based routing optimization in software defined networks. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521001909.

Ketu:2021:IHT

[3042] Shwet Ketu and Pramod Kumar Mishra. Internet of Healthcare Things: a contemporary survey. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001892.

Yang:2021:PVO

[3043] Changsong Yang, Feng Zhao, Xiaoling Tao, and Yong Wang. Publicly verifiable outsourced data migration scheme supporting efficient integrity checking. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001910.

Kashani:2021:SRI

[3044] Mostafa Haghi Kashani, Mona Madanipour, Mohammad Nikravan, Parvaneh Asghari, and Ebrahim Mahdipour. A systematic review of IoT in healthcare: Applications, techniques, and trends. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001764.

Cheng:2021:MLB

[3045] Qiumei Cheng, Chunming Wu, Haifeng Zhou, Dezhang Kong, Dong Zhang, Junchi Xing, and Wei Ruan. Machine learning based malicious payload identification in software-defined networking. Journal of Network and Computer Applications, 192(??):??, October 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521001934.

Anonymous:2021:Na

[3046] Anonymous. 1 November 2021. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBt

[3047] Anonymous. Editorial Board. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452100237X.

Yang:2021:DLO

[3048] Liqun Yang, You Zhai, and Zhoujun Li. Deep learning for online AC False Data Injection Attack detection in smart grids: an approach using LSTM-Autoencoder. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001880.

Krishnan:2021:OAA

[3049] Prabhakar Krishnan, Subhasri Duttagupta, and Rajkumar Buyya. Open-PATH: Application aware high-performance software-defined switching framework. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002022.

Guo:2021:PSE

[3050] Junyan Guo, Ye Du, Yahang Zhang, and Meihong Li. A provably secure ECC-based access and handover authentication protocol for space information networks. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521001922.

Hao:2021:MAO

[3051] Hao Hao, Changqiao Xu, Shujie Yang, Lujie Zhong, and Gabriel-Miro Muntean. Multicast-aware optimization for resource allocation with edge computing and caching. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002010.

Jiang:2021:SEM

[3052] Yuning Jiang and Yacine Atif. A selective ensemble model for cognitive cybersecurity analysis. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002125.

Bui:2021:CTA

[3053] Huu-Khoi Bui, Ying-Dar Lin, Ren-Hung Hwang, Po-Ching Lin, Van-Linh Nguyen, and Yuan-Cheng Lai. CREME: a toolchain of automatic dataset collection for machine learning in intrusion detection. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002137.

Haque:2021:SRA

[3054] Israat Haque and Dipon Saha. SoftIoT: a resource-aware SDN/NFVbased IoT network. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002113.

Yan:2021:LSH

[3055] Xiaobei Yan and Maode Ma. A lightweight and secure handover authentication scheme for 5G network using neighbour base stations. Journal of Network and Computer Applications, 193(??):??, November 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002095.

Anonymous:2021:Nb

[3056] Anonymous. 15 November 2021. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBu

[3057] Anonymous. Editorial Board. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521002514.

Bochie:2021:SDL

[3058] Kaylani Bochie, Mateus S. Gilbert, Luana Gantert, Mariana S. M. Barbosa, Dianne S. V. Medeiros, and Miguel Elias M. Campista. A survey on deep learning for challenged networks: Applications and trends. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002149.

Yahia:2021:SCM

[3059] Selma Yahia, Yassine Meraihi, Amar Ramdane-Cherif, Asma Benmessaoud Gabis, Dalila Acheli, and Hongyu Guan. A survey of channel modeling techniques for visible light communications. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002101.

Elbasheer:2021:QBR

[3060] Majda Omer Elbasheer, Abdulaziz Aldegheishem, Jaime Lloret, and Nabil Alrajeh. A QoS-based routing algorithm over software defined networks. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002162.

Wang:2021:DIP

[3061] Shie-Yuan Wang, Jui-En Chang, and Cheng-Da Lin. Design, implementation, and performance evaluation of a drone wakeup and data collection system for smart meters. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002289.

Krishnan:2021:RLB

[3062] Muralitharan Krishnan and Yongdo Lim. Reinforcement learning-based dynamic routing using mobile sink for data collection in WSNs and IoT applications. Journal of Network and Computer Applications, 194 (??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002241.

Liu:2021:RMS

[3063] Xi Liu, Jian Cen, Huanzhong Hu, Zongwei Yu, and Yuanxin Huang. A radio map self-updating algorithm based on mobile crowd sensing. Journal of Network and Computer Applications, 194(??):??, November 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002265.

Anonymous:2021:Da

[3064] Anonymous. 1 December 2021. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBv

[3065] Anonymous. Editorial Board. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521002605.

Houmz:2021:DIS

[3066] Abdellah Houmz, Ghita Mezzour, Karim Zkik, Mounir Ghogho, and Houda Benbrahim. Detecting the impact of software vulnerability on attacks: a case study of network telescope scans. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002290.

Zhang:2021:GPB

[3067] Hongyan Zhang, Limei Lin, Li Xu, and Xiaoding Wang. Graph partition based privacy-preserving scheme in social networks. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002150.

Sanka:2021:SRB

[3068] Abdurrashid Ibrahim Sanka and Ray C. C. Cheung. A systematic review of blockchain scalability: Issues, solutions, analysis and future research. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002307.

Castaneda:2021:VBN

[3069] Josué Castañeda, Saul E. Pomares Hernández, Sami Yangui, Julio C. Pérez Sansalvador, Lil M. Rodríguez Henríquez, and Khalil Drira. VNF-based network service consistent reconfiguration in multi-domain federations: a distributed approach. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804521002277.

Tahir:2021:NIP

[3070] Ruhma Tahir, Shahzaib Tahir, Hasan Tahir, Klaus McDonald-Maier, Gareth Howells, and Ali Sajjad. A novel ICMetric public key framework for secure communication. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002332.

Mehto:2021:MOP

[3071] Anjula Mehto, Shashikala Tapaswi, and K. K. Pattanaik. Multiobjective particle swarm optimization based rendezvous point selection for the energy and delay efficient networked wireless sensor data acquisition. Journal of Network and Computer Applications, 195(??):??, December 1, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002320.

Anonymous:2021:Db

[3072] Anonymous. 15 December 2021. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2021:EBw

[3073] Anonymous. Editorial Board. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452100271X.

Lei:2021:HTL

[3074] Kai Lei, Junlin Huang, Xiaodong Li, Yu Li, Ye Zhang, Bo Bai, Fan Zhang, Gong Zhang, and Jingjie Jiang. HOPASS: a two-layer control framework for bandwidth and delay guarantee in datacenters. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002253.

Zhang:2021:RSP

[3075] Fan Zhang, Yazhe Tang, Danfeng Shan, Huanzhao Wang, and Chengchen Hu. RICH: Strategy-proof and efficient coflow scheduling in noncooperative environments. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002319.

Jiang:2021:RGI

[3076] Jiaojiao Jiang, Jean-Guy Schneider, Steve Versteeg, Jun Han, MD Arafat Hossain, and Chengfei Liu. R-gram: Inferring message formats of service protocols with relative positional n-grams. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002459.

Wei:2021:FCO

[3077] Bingjie Wei, Lin Wang, Junlong Zhu, Mingchuan Zhang, Ling Xing, and Qingtao Wu. Flow control oriented forwarding and caching in cache-enabled networks. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002460.

Dorsala:2021:BBS

[3078] Mallikarjun Reddy Dorsala, V. N. Sastry, and Sudhakar Chapram. Blockchain-based solutions for cloud computing: a survey. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804521002447.

Aditya:2021:SBR

[3079] U. S. P. Srinivas Aditya, Roshan Singh, Pranav Kumar Singh, and Anshuman Kalla. A survey on blockchain in robotics: Issues, opportunities, challenges and future directions. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521002435.

Sworna:2021:TDI

[3080] Nabila Sabrin Sworna, A. K. M. Muzahidul Islam, Swakkhar Shatabda, and Salekul Islam. Towards development of IoT-ML driven healthcare systems: a survey. Journal of Network and Computer Applications, 196(??):??, December 15, 2021. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002423.

Anonymous:2022:Ja

[3081] Anonymous. January 2022. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBa

[3082] Anonymous. Editorial Board. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002824.

Anonymous:2022:TYR

[3083] Anonymous. Thank you reviewers. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804521002861.

Chowdhury:2022:VSI

[3084] Debanjan Roy Chowdhury, Sukumar Nandi, and Diganta Goswami. Video streaming over IoV using IP multicast. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002563.

Sun:2022:PLM

[3085] Fanghui Sun, Shen Wang, and Hongli Zhang. A progressive learning method on unknown protocol behaviors. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804521002472.

Leyva-Pupo:2022:ESP

[3086] Irian Leyva-Pupo and Cristina Cervelló-Pastor. Efficient solutions to the placement and chaining problem of User Plane Functions in 5G networks. Journal of Network and Computer Applications, 197(??): ??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002654.

Hameed:2022:CAI

[3087] Khizar Hameed, Saurabh Garg, Muhammad Bilal Amin, Byeong Kang, and Abid Khan. A context-aware information-based clone node attack detection scheme in Internet of Things. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804521002678.

Verma:2022:OQI

[3088] Gunjan Verma, Kelvin Marcus, and Kevin S. Chan. Optimizing the quality of information of networked machine learning agents. Journal of Network and Computer Applications, 197(??):??, January 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002411.

Anonymous:2022:F

[3089] Anonymous. February 2022. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBb

[3090] Anonymous. Editorial Board. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452100309X.

Benhamaid:2022:RAE

[3091] Sana Benhamaid, Abdelmadjid Bouabdallah, and Hicham Lakhlef. Recent advances in energy management for Green-IoT: an up-to-date and comprehensive survey. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002551.

Zhang:2022:RNI

[3092] Xiaoqing Zhang, Fei Yang, Yue Hu, Zhao Tian, Wei Liu, Yifa Li, and Wei She. RANet: Network intrusion detection with group-gating convolutional neural network. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002630.

Shahid:2022:EDL

[3093] Waleed Bin Shahid, Baber Aslam, Haider Abbas, Saad Bin Khalid, and Hammad Afzal. An enhanced deep learning based framework for web attacks detection, mitigation and attacker profiling. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002666.

Wang:2022:DII

[3094] Shie-Yuan Wang and Jen-Chieh Chang. Design and implementation of an intrusion detection system by using Extended BPF in the Linux kernel. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002769.

Ma:2022:RAS

[3095] Ruhui Ma, Jin Cao, Dengguo Feng, Hui Li, Xiaowei Li, and Yang Xu. A robust authentication scheme for remote diagnosis and maintenance in 5G V2N. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002770.

Rui:2022:SNM

[3096] Lanlan Rui, Siqi Yang, Zhipeng Gao, Wenjing Li, Xuesong Qiu, and Luoming Meng. Smart network maintenance in edge cloud computing environment: an allocation mechanism based on comprehensive reputation and regional prediction model. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002915.

Azzaoui:2022:BBD

[3097] Abir El Azzaoui, Pradip Kumar Sharma, and Jong Hyuk Park. Blockchain-based delegated Quantum Cloud architecture for medical big data security. Journal of Network and Computer Applications, 198(??): ??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002952.

Islam:2022:SEE

[3098] Kazi Yasin Islam, Iftekhar Ahmad, Daryoush Habibi, and Adnan Waqar. A survey on energy efficiency in underwater wireless communications. Journal of Network and Computer Applications, 198(??): ??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002885.

Ullah:2022:SBD

[3099] Faheem Ullah and M. Ali Babar. On the scalability of Big Data Cyber Security Analytics systems. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002897.

Anonymous:2022:Ma

[3100] Anonymous. March 2022. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBc

[3101] Anonymous. Editorial Board. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000078.

Nguyen:2022:PES

[3102] Tuan Anh Nguyen, Minjune Kim, Jangse Lee, Dugki Min, Jae-Woo Lee, and Dongseong Kim. Performability evaluation of switch-over moving target defence mechanisms in a software defined networking using stochastic reward nets. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002642.

Adhinugraha:2022:MFT

[3103] Kiki Adhinugraha, Wenny Rahayu, Takahiro Hara, and David Taniar. Measuring fault tolerance in IoT mesh networks using Voronoi diagram. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002903.

Liu:2022:WTD

[3104] Yimin Liu, Xiangyang Luo, Meng Zhang, Zhiyuan Tao, and Fenlin Liu. Who are there: Discover Twitter users and tweets for target area using mention relationship strength and local tweet ratio. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002939.

Kalewski:2022:URR

[3105] Michał Kalewski and Jerzy Brzeziński. Uniform and regular reliable broadcast protocols facilitating concurrent message dissemination for mobile ad hoc networks with the network liveness property. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002927.

Zhou:2022:RPC

[3106] Yadong Zhou, Hao Li, Kaiyue Chen, Tian Pan, Kun Qian, Kai Zheng, Bin Liu, Peng Zhang, Yazhe Tang, and Chengchen Hu. Raze policy conflicts in SDN. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002964.

Liang:2022:MAE

[3107] Bin Liang, Mark A. Gregory, and Shuo Li. Multi-access edge computing fundamentals, services, enablers and challenges: a complete survey. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002976.

Memon:2022:PAC

[3108] Shuaib K. Memon, Nurul I. Sarkar, Adnan Al-Anbuky, and Md Akbar Hossain. Preemptive admission control mechanism for strict QoS guarantee to life-saving emergency traffic in wireless LANs. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521003040.

Gauttam:2022:CAT

[3109] Himanshu Gauttam, K. K. Pattanaik, Saumya Bhadauria, Divya Saxena, and Sapna. A cost aware topology formation scheme for latency sensitive applications in edge infrastructure-as-a-service paradigm. Journal of Network and Computer Applications, 199(??):??, March 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521002940.

Anonymous:2022:Aa

[3110] Anonymous. April 2022. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBd

[3111] Anonymous. Editorial Board. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000182.

Biswash:2022:DND

[3112] Sanjay Kumar Biswash. Device and network driven cellular networks architecture and mobility management technique for fog computing-based mobile communication system. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521003027.

Jahid:2022:CSF

[3113] Abu Jahid, Mohammed H. Alsharif, and Trevor J. Hall. A contemporary survey on free space optical communication: Potentials, technical challenges, recent advances and research direction. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452100299X.

Kim:2022:PPM

[3114] Jong Wook Kim, Kennedy Edemacu, and Beakcheol Jang. Privacypreserving mechanisms for location privacy in mobile crowdsensing: a survey. Journal of Network and Computer Applications, 200(??): ??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521003039.

Auer:2022:TBI

[3115] Sophia Auer, Sophia Nagler, Somnath Mazumdar, and Raghava Rao Mukkamala. Towards blockchain-IoT based shared mobility: Car-sharing and leasing as a case study. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521003015.

Mehmood:2022:ECE

[3116] Farrukh Mehmood, Naveed Khan Baloch, Fawad Hussain, Waqar Amin, M. Shamim Hossain, Yousaf Bin Zikria, and Heejung Yu. An efficient and cost effective application mapping for network-on-chip using Andean condor algorithm. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521003052.

Zeb:2022:IDT

[3117] Shah Zeb, Aamir Mahmood, Syed Ali Hassan, MD. Jalil Piran, Mikael Gidlund, and Mohsen Guizani. Industrial digital twins at the nexus of NextG wireless networks and computational intelligence: a survey. Journal of Network and Computer Applications, 200(??): ??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521002988.

Valerio:2022:DHP

[3118] Lorenzo Valerio, Franco Maria Nardini, Andrea Passarella, and Raffaele Perego. Dynamic hard pruning of neural networks at the edge of the Internet. Journal of Network and Computer Applications, 200 (??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521003155.

Mezrag:2022:ELI

[3119] Fares Mezrag, Salim Bitam, and Abdelhamid Mellouk. An efficient and lightweight identity-based scheme for secure communication in clustered wireless sensor networks. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521002782.

Thai:2022:DIN

[3120] Quang Tung Thai, Namseok Ko, Sung Hyuk Byun, and Sun-Me Kim. Design and implementation of NDN-based Ethereum blockchain. Journal of Network and Computer Applications, 200(??):??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804521003143.

Achir:2022:SDS

[3121] Meriem Achir, Abdelkrim Abdelli, Lynda Mokdad, and Jalel Benothman. Service discovery and selection in IoT: a survey and a taxonomy. Journal of Network and Computer Applications, 200(??): ??, April 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804521003167.

Anonymous:2022:Mb

[3122] Anonymous. May 2022. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBe

[3123] Anonymous. Editorial Board. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000406.

Chintapalli:2022:RDC

[3124] Venkatarami Reddy Chintapalli, Madhura Adeppady, Bheemarjuna Reddy Tamma, and Antony Franklin A. RESTRAIN: a dynamic and costefficient resource management scheme for addressing performance interference in NFV-based systems. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804521003003.

Azizi:2022:DAE

[3125] Sadoon Azizi, Mohammad Shojafar, Jemal Abawajy, and Rajkumar Buyya. Deadline-aware and energy-efficient IoT task scheduling in fog computing systems: a semi-greedy approach. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804522000029.

Rasool:2022:SPI

[3126] Raihan Ur Rasool, Hafiz Farooq Ahmad, Wajid Rafique, Adnan Qayyum, and Junaid Qadir. Security and privacy of internet of medical things: a contemporary review in the age of surveillance, botnets, and adversarial ML. Journal of Network and Computer Applications, 201(??): ??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000017.

Huda:2022:SCO

[3127] S. M. Asiful Huda and Sangman Moh. Survey on computation offloading in UAV-enabled mobile edge computing. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522000108.

Cotroneo:2022:TNA

[3128] Domenico Cotroneo, Luigi De Simone, and Roberto Natella. ThorFI: a novel approach for network fault injection as a service. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000030.

Kumar:2022:MNN

[3129] Shashwat Kumar, Lalit Bhagat, Antony Franklin A., and Jiong Jin. Multi-neural network based tiled 360^{∘} video caching with mobile edge computing. Journal of Network and Computer Applications, 201(??): ??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452200011X.

Li:2022:SRA

[3130] Qian Li, Cong Zeng, Wei Xu, and Yunpeng Xiao. A social rumor and anti-rumor game diffusion model based on sparse representation and tensor completion. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000121.

Mobarhan:2022:RAS

[3131] Mostafa Ayoubi Mobarhan and Mohammed Salamah. REPS-AKA3: a secure authentication and re- authentication protocol for LTE networks. Journal of Network and Computer Applications, 201(??): ??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000145.

Liang:2022:MBR

[3132] Guangxian Liang, Xiaopeng Yao, Yue Gu, Hejiao Huang, and Chonglin Gu. Multi-Batches Revenue Maximization for competitive products over online social network. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000248.

Wang:2022:FFF

[3133] Yang Wang, Xiong Wang, Shizhong Xu, Ci He, Yasheng Zheng, Jing Ren, and Shui Yu. FlexMon: a flexible and fine-grained traffic monitor for programmable networks. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000133.

Nakip:2022:DAF

[3134] Mert Nakip, Erdem Çakan, Volkan Rodoplu, and Cüneyt Güzelis. Dynamic automatic forecaster selection via artificial neural network based emulation to enable massive access for the Internet of Things. Journal of Network and Computer Applications, 201(??):??, May 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000273.

Anonymous:2022:Jb

[3135] Anonymous. June 2022. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBf

[3136] Anonymous. Editorial Board. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000686.

Honarvar:2022:COP

[3137] Reza Honarvar, Alireza Zolghadrasli, and Mehdi Monemi. Contextoriented performance evaluation of network selection algorithms in 5G heterogeneous networks. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000261.

Salim:2022:SSE

[3138] Ahmed Salim, Walid Osamy, Ahmed Aziz, and Ahmed M. Khedr. SEEDGT: Secure and energy efficient data gathering technique for IoT applications based WSNs. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000224.

Sun:2022:SPV

[3139] Jie Sun, Yi Zhang, Feng Liu, Huandong Wang, Xiaojian Xu, and Yong Li. A survey on the placement of virtual network functions. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000285.

Wei:2022:ESS

[3140] Wei Wei, Qi Wang, Weidong Yang, and Yashuang Mu. Efficient stochastic scheduling for highly complex resource placement in edge clouds. Journal of Network and Computer Applications, 202(??):

??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000315.

Chapaneri:2022:EDI

[3141] Radhika Chapaneri and Seema Shah. Enhanced detection of imbalanced malicious network traffic with regularized Generative Adversarial Networks. Journal of Network and Computer Applications, 202(??): ??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000339.

Feng:2022:COM

[3142] Chuan Feng, Pengchao Han, Xu Zhang, Bowen Yang, Yejun Liu, and Lei Guo. Computation offloading in mobile edge computing networks: a survey. Journal of Network and Computer Applications, 202(??): ??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000327.

Islam:2022:SFA

[3143] Chadni Islam, M. Ali Babar, Roland Croft, and Helge Janicke. Smart-Validator: a framework for automatic identification and classification of cyber threat data. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000340.

Ahmed:2022:DEP

[3144] Nadeem Ahmed, Regio A. Michelin, Wanli Xue, Guntur Dharma Putra, Sushmita Ruj, Salil S. Kanhere, and Sanjay Jha. DIMY: Enabling privacy-preserving contact tracing. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452200025X.

Zade:2022:TSS

[3145] B. Mohammad Hasani Zade, N. Mansouri, and M. M. Javidi. A twostage scheduler based on New Caledonian Crow Learning Algorithm and reinforcement learning strategy for cloud environment. Journal of Network and Computer Applications, 202(??):??, June 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000480.

Catillo:2022:NME

[3146] Marta Catillo, Antonio Pecchia, and Umberto Villano. No more DoS? An empirical study on defense techniques for web server Denial of Service mitigation. Journal of Network and Computer Applications, 202 (??):??, June 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000303.

Anonymous:2022:Jc

[3147] Anonymous. July 2022. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBg

[3148] Anonymous. Editorial Board. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000807.

Wang:2022:CAD

[3149] Nan Wang and Blesson Varghese. Context-aware distribution of fog applications using deep reinforcement learning. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804522000236.

Liyanage:2022:SZT

[3150] Madhusanka Liyanage, Quoc-Viet Pham, Kapal Dev, Sweta Bhattacharya, Praveen Kumar Reddy Maddikunta, Thippa Reddy Gadekallu, and Gokul Yenduri. A survey on Zero touch network and Service Management (ZSM) for 5G and beyond networks. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522000297.

Pal:2022:BIA

[3151] Shantanu Pal, Ali Dorri, and Raja Jurdak. Blockchain for IoT access control: Recent trends and future research directions. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000352.

Conti:2022:SSC

[3152] Mauro Conti, Gulshan Kumar, Pranav Nerurkar, Rahul Saha, and Luigi Vigneri. A survey on security challenges and solutions in the IOTA. Journal of Network and Computer Applications, 203(??): ??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000467.

Li:2022:SRP

[3153] Xingjia Li, Li Pan, and Shijun Liu. A survey of resource provisioning problem in cloud brokers. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000479.

Gupta:2022:TTD

[3154] Pooja Gupta, Volkan Dedeoglu, Salil S. Kanhere, and Raja Jurdak. TrailChain: Traceability of data ownership across blockchain-enabled multiple marketplaces. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000509.

Ji:2022:FNE

[3155] Soo-Yeon Ji, Bong Keun Jeong, Charles Kamhoua, Nandi Leslie, and Dong Hyun Jeong. Forecasting network events to estimate attack risk: Integration of wavelet transform and vector auto regression with exogenous variables. Journal of Network and Computer Applications, 203 (??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000522.

Li:2022:SPP

[3156] Tianheng Li, Xiaofan He, Siming Jiang, and Juan Liu. A survey of privacy-preserving offloading methods in mobile-edge computing. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804522000546.

Althumali:2022:PBL

[3157] Huda Althumali, Mohamed Othman, Nor Kamariah Noordin, and Zurina Mohd Hanapi. Priority-based load-adaptive preamble separation random access for QoS-differentiated services in 5G networks. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000558.

Coppolino:2022:PEP

[3158] Luigi Coppolino, Salvatore D’Antonio, Giovanni Mazzeo, Luigi Romano, and Luigi Sgaglione. PriSIEM: Enabling privacy-preserving managed security services. Journal of Network and Computer Applications, 203 (??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452200056X.

Mpatziakas:2022:ITM

[3159] Asterios Mpatziakas, Anastasios Drosou, Stavros Papadopoulos, and Dimitiris Tzovaras. IoT threat mitigation engine empowered by artificial intelligence multi-objective optimization. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522000571.

Luo:2022:OMF

[3160] Long Luo, Klaus-Tycho Foerster, Stefan Schmid, and Hongfang Yu. Optimizing multicast flows in high-bandwidth reconfigurable datacenter networks. Journal of Network and Computer Applications, 203 (??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000583.

Fogli:2022:SDN

[3161] Mattia Fogli, Carlo Giannelli, and Cesare Stefanelli. Software-defined networking in wireless ad hoc scenarios: Objectives and control architectures. Journal of Network and Computer Applications, 203(??): ??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000492.

Konjaang:2022:EEV

[3162] J. Kok Konjaang, John Murphy, and Liam Murphy. Energy-efficient virtual-machine mapping algorithm (EViMA) for workflow tasks with deadlines in a cloud environment. Journal of Network and Computer Applications, 203(??):??, July 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000595.

Anonymous:2022:Ab

[3163] Anonymous. August 2022. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBh

[3164] Anonymous. Editorial Board. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452200100X.

Wang:2022:CIO

[3165] Chu-Fu Wang, Yih-Kai Lin, and Jun-Cheng Chen. A cooperative image object recognition framework and task offloading optimization in edge computing. Journal of Network and Computer Applications, 204(??): ??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000637.

Khan:2022:MLM

[3166] Tahseen Khan, Wenhong Tian, Guangyao Zhou, Shashikant Ilager, Mingming Gong, and Rajkumar Buyya. Machine learning (ML)-centric resource management in cloud computing: a review and future directions. Journal of Network and Computer Applications, 204(??): ??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000649.

Park:2022:SUL

[3167] Taejune Park and Kilho Lee. Supporting ultra-low latency mixedcriticality communication using hardware-based data plane architecture. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000601.

Lahlou:2022:TRP

[3168] Sara Lahlou, Youness Moukafih, Anass Sebbar, Karim Zkik, Mohammed Boulmalf, and Mounir Ghogho. TD-RA policy-enforcement framework for an SDN-based IoT architecture. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522000510.

Abbasi-Verki:2022:SAE

[3169] Mansoureh Abbasi-Verki, Saleh Yousefi, and Hashem Kalbkhani. Socially-aware and energy-efficient resource allocation and power control for D2D multicast content distribution. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480452200073X.

Nikravan:2022:RTM

[3170] Mohammad Nikravan and Mostafa Haghi Kashani. A review on trust management in fog/edge computing: Techniques, trends, and challenges. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000613.

Kumar:2022:CSA

[3171] Ashish Kumar, Rahul Saha, Mauro Conti, Gulshan Kumar, William J. Buchanan, and Tai Hoon Kim. A comprehensive survey of authentication methods in Internet-of-Things and its conjunctions. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000716.

Thakur:2022:IFE

[3172] Dipanwita Thakur and Suparna Biswas. An integration of feature extraction and Guided Regularized Random Forest feature selection for Smartphone based Human Activity Recognition. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000741.

Ijemaru:2022:TII

[3173] Gerald K. Ijemaru, Li Minn Ang, and Kah Phooi Seng. Transformation from IoT to IoV for waste management in smart cities. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000534.

Alam:2022:JTC

[3174] Muhammad Morshed Alam and Sangman Moh. Joint topology control and routing in a UAV swarm for crowd surveillance. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000844.

Ghafoor:2022:NFB

[3175] Umar Ghafoor, Mudassar Ali, Humayun Zubair Khan, Adil Masood Siddiqui, and Muhammad Naeem. NOMA and future 5G and B5G wireless networks: a paradigm. Journal of Network and Computer Applications, 204(??):??, August 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000728.

Anonymous:2022:S

[3176] Anonymous. September 2022. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBi

[3177] Anonymous. Editorial Board. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001187.

S:2022:RUM

[3178] Ansa Shermin S. and Sarang C. Dhongdi. Review of underwater mobile sensor network for ocean phenomena monitoring. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000753.

Ferrer-Cid:2022:DRA

[3179] Pau Ferrer-Cid, Jose M. Barcelo-Ordinas, and Jorge Garcia-Vidal. Data reconstruction applications for IoT air pollution sensor networks using graph signal processing. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000881.

Agiwal:2022:SPP

[3180] Mamta Agiwal, Anil Agiwal, Mukesh Kumar Maheshwari, and Shapna Muralidharan. Split PO for paging in B5G networks. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000868.

Vakilian:2022:NCW

[3181] Shakoor Vakilian, Ali Fanian, Hossein Falsafain, and T. Aaron Gulliver. Node cooperation for workload offloading in a fog computing network via multi-objective optimization. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000856.

Soni:2022:MLT

[3182] Dinesh Soni and Neetesh Kumar. Machine learning techniques in emerging cloud computing integrated paradigms: a survey and taxonomy. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000765.

Roy:2022:FFH

[3183] Prasanta Kumar Roy, Prashant Sahu, and Ansuman Bhattacharya. FastHand: a fast handover authentication protocol for densely deployed small-cell networks. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522000893.

Bokhari:2022:CSU

[3184] Samira Bokhari, Sofiane Hamrioui, and Meziane Aider. Cybersecurity strategy under uncertainties for an IoE environment. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000832.

Campana:2022:DMU

[3185] Mattia G. Campana and Franca Delmastro. On-device modeling of user’s social context and familiar places from smartphone-embedded sensor data. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452200090X.

Zhang:2022:SVN

[3186] Peiying Zhang, Peng Gan, Neeraj Kumar, Chunxiao Jiang, Fanglin Liu, and Lei Zhang. Survivable virtual network embedding algorithm considering multiple node failure in IIoT environment. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000911.

Ni:2022:NCB

[3187] Haoran Ni, Zehua Guo, Changlin Li, Songshi Dou, Chao Yao, and Thar Baker. Network coding-based resilient routing for maintaining data security and availability in software-defined networks. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522000364.

Wu:2022:EST

[3188] Xiaoping Wu, Yunlong Li, Yana Shen, and Xuefen Zhu. Efficient solutions for target localization in asynchronous MIMO networks. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000947.

Taimoor:2022:HRM

[3189] Shamim Taimoor, Lilatul Ferdouse, and Waleed Ejaz. Holistic resource management in UAV-assisted wireless networks: an optimization perspective. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000923.

Baniata:2022:PPA

[3190] Hamza Baniata and Attila Kertesz. PriFoB: a privacy-aware fogenhanced blockchain-based system for global accreditation and credential verification. Journal of Network and Computer Applications, 205(??): ??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000935.

Yungaicela-Naula:2022:FSB

[3191] Noe M. Yungaicela-Naula, Cesar Vargas-Rosales, Jesús Arturo Pérez-Díaz, and Diego Fernando Carrera. A flexible SDN-based framework for slow-rate DDoS attack mitigation by using deep reinforcement learning. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522000960.

Ozcelik:2022:AAL

[3192] Ihsan Mert Ozcelik and Cem Ersoy. ALVS: Adaptive live video streaming using deep reinforcement learning. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001035.

Karmakar:2022:UAN

[3193] Kamalesh Karmakar, Somrita Banerjee, Rajib K. Das, and Sunirmal Khatua. Utilization aware and network I/O intensive virtual machine placement policies for cloud data center. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804522000959.

Paul:2022:DRL

[3194] Anal Paul, Aneesh Kumar Mishra, Shivam Shreevastava, and Anoop Kumar Tiwari. Deep reinforcement learning based reliable spectrum sensing under SSDF attacks in cognitive radio networks. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522001059.

Akther:2022:IFS

[3195] Nazma Akther, Kingshuk Dhar, Shahid Md. Asif Iqbal, Mohammed Nurul Huda, and Asaduzzaman. Interest forwarding strategy in named Data networks (ndn) using Thompson sampling. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522001084.

Li:2022:MLV

[3196] Min Li, Di Xiao, Hui Huang, and Bo Zhang. Multi-level video quality services and security guarantees based on compressive sensing in sensor-cloud system. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001072.

Zou:2022:DCP

[3197] Qiang Zou, Yifeng Zhu, Yujuan Tan, and Wei Chen. Diagnosing the coexistence of Poissonity and self-similarity in memory workloads. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001060.

Wang:2022:LBT

[3198] Nan Wang, Tong Liu, Jinzhen Wang, Qing Liu, Shakeel Alibhai, and Xubin He. Locality-based transfer learning on compression autoencoder for efficient scientific data lossy compression. Journal of Network and Computer Applications, 205(??):??, September 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522001047.

Anonymous:2022:O

[3199] Anonymous. October 2022. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBj

[3200] Anonymous. Editorial Board. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001448.

Irram:2022:PLS

[3201] Fauzia Irram, Mudassar Ali, Muhammad Naeem, and Shahid Mumtaz. Physical layer security for beyond 5G/6G networks: Emerging technologies and future directions. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452200087X.

Kim:2022:DLB

[3202] Jong Wook Kim and Beakcheol Jang. Deep learning-based privacypreserving framework for synthetic trajectory generation. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522001096.

Wu:2022:SLS

[3203] Minghui Wu, Dawei Sun, Yijing Cui, Shang Gao, Xunyun Liu, and Rajkumar Buyya. A state lossless scheduling strategy in distributed stream computing systems. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001114.

Tran:2022:FAD

[3204] Nguyen Khoi Tran, M. Ali Babar, and Andrew Walters. A framework for automating deployment and evaluation of blockchain networks. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001102.

Long:2022:KRE

[3205] Qianyu Long, Kostas Kolomvatsos, and Christos Anagnostopoulos. Knowledge reuse in edge computing environments. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452200114X.

Maddikunta:2022:ITI

[3206] Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, Dinh C. Nguyen, Thien Huynh-The, Ons Aouedi, Gokul Yenduri, Sweta Bhattacharya, and Thippa Reddy Gadekallu. Incentive techniques for the Internet of Things: a survey. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001138.

Jung:2022:VGD

[3207] Changue Jung, Suhwan Kim, Younghoon Kim, and Ikjun Yeom. Virtualizing GPU direct packet I/O on commodity Ethernet to accelerate GPU-NFV. Journal of Network and Computer Applications, 206(??): ??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001254.

Elshrkasi:2022:ECN

[3208] Ahmed Elshrkasi, Kaharudin Dimyati, Khairol Amali Bin Ahmad, and Mohamed Faidz bin Mohamed Said. Enhancement of cellular networks via an improved clustering technique with D2D communication for mission-critical applications. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001278.

Al-Ward:2022:CTD

[3209] Hisham Al-Ward, Chee Keong Tan, and Wern Han Lim. Caching transient data in information-centric Internet-of-Things (IC-IoT) networks: a survey. Journal of Network and Computer Applications, 206(??): ??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001333.

Dabaghi-Zarandi:2022:CDC

[3210] Fahimeh Dabaghi-Zarandi and Parsa KamaliPour. Community detection in complex network based on an improved random algorithm using local and global network information. Journal of Network and Computer Applications, 206(??):??, October 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001345.

Anonymous:2022:N

[3211] Anonymous. November 2022. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBk

[3212] Anonymous. Editorial Board. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001655.

Ye:2022:VVR

[3213] Heng Ye, Jiqiang Liu, Hao Zhen, Wenbin Jiang, Bin Wang, and Wei Wang. VREFL: Verifiable and reconnection-efficient federated learning in IoT scenarios. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452200131X.

Raza:2022:IGF

[3214] Syed M. Raza, Boyun Jang, Huigyu Yang, Moonseong Kim, and Hyunseung Choo. Improved GAN with fact forcing for mobility prediction. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001321.

Yu:2022:TPP

[3215] Ruiyun Yu, Ann Move Oguti, Dennis Reagan Ochora, and Shuchen Li. Towards a privacy-preserving smart contract-based data aggregation and quality-driven incentive mechanism for mobile crowdsensing. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001291.

Khaleel:2022:PRP

[3216] Mustafa Ibrahim Khaleel. PPR-RM: Performance-to-power ratio, reliability and makespan-aware scientific workflow scheduling based on a coalitional game in the cloud. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001266.

Liu:2022:RRE

[3217] Yuxian Liu, Fagui Liu, Hao-Tian Wu, Jingfeng Yang, Kaihong Zheng, Lingling Xu, Xingfu Yan, and Jiankun Hu. RPTD: Reliabilityenhanced privacy-preserving Truth Discovery for Mobile Crowdsensing. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001308.

Garrido:2022:RLP

[3218] Gonzalo Munilla Garrido, Johannes Sedlmeir, Ömer Uludağ, Ilias Soto Alaoui, Andre Luckow, and Florian Matthes. Revealing the landscape of privacy-enhancing technologies in the context of data markets for the IoT: a systematic literature review. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001126.

Darabkh:2022:RRP

[3219] Khalid A. Darabkh, Muna Al-Akhras, Jumana N. Zomot, and Mohammed Atiquzzaman. RPL routing protocol over IoT: a comprehensive survey, recent advances, insights, bibliometric analysis, recommendations, and future directions. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001242.

Alam:2022:TCA

[3220] Muhammad Morshed Alam, Muhammad Yeasir Arafat, Sangman Moh, and Jian Shen. Topology control algorithms in multi-unmanned aerial vehicle networks: an extensive survey. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001370.

Pan:2022:GGO

[3221] Ying Pan, Na Lyu, and Chungang Yang. GOR: Group-oblivious multicast routing in airborne tactical networks under uncertainty. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804522001503.

Shen:2022:MOD

[3222] Gengbiao Shen, Qing Li, Wanxin Shi, Yong Jiang, Pei Zhang, Liang Gu, and Mingwei Xu. Modeling and optimization of the data plane in the SDN-based DCN by queuing theory. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452200128X.

Shan:2022:SRR

[3223] Gaoyang Shan and Byeong hee Roh. A slotted random request scheme for connectionless data transmission in bluetooth low energy 5.0. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001357.

Dwivedi:2022:SLC

[3224] Bharat Dwivedi, Debarati Sen, and Sandip Chakraborty. A survey of longitudinal changes in cellular network architecture: the good, the bad, and the ugly. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001382.

Arooj:2022:UBE

[3225] Ansif Arooj, Muhammad Shoaib Farooq, and Tariq Umer. Unfolding the blockchain era: Timeline, evolution, types and real-world applications. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001527.

Zeynivand:2022:TFC

[3226] A. Zeynivand, A. Javadpour, S. Bolouki, A. K. Sangaiah, F. Ja’fari, P. Pinto, and W. Zhang. Traffic flow control using multi-agent reinforcement learning. Journal of Network and Computer Applications, 207(??): ??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001394.

Xiang:2022:DAA

[3227] Xinyin Xiang, Jin Cao, and Weiguo Fan. Decentralized authentication and access control protocol for blockchain-based e-health systems. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001539.

Monshizadeh:2022:DDB

[3228] Mehrnoosh Monshizadeh, Vikramajeet Khatri, Raimo Kantola, and Zheng Yan. A deep density based and self-determining clustering approach to label unknown traffic. Journal of Network and Computer Applications, 207(??):??, November 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001540.

Anonymous:2022:D

[3229] Anonymous. December 2022. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2022:EBl

[3230] Anonymous. Editorial Board. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001746.

Choi:2022:ICC

[3231] Sang-Hoon Choi and Ki-Woong Park. iContainer: Consecutive checkpointing with rapid resilience for immortal container-based services. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804522001369.

Nam:2022:RRE

[3232] Jaehyun Nam, Seung Ho Na, Seungwon Shin, and Taejune Park. Reconfigurable regular expression matching architecture for real-time pattern update and payload inspection. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001497.

Zhou:2022:DRL

[3233] Guangyao Zhou, Ruiming Wen, Wenhong Tian, and Rajkumar Buyya. Deep reinforcement learning-based algorithms selectors for the resource scheduling in hierarchical cloud computing. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804522001618.

Masmoudi:2022:PRP

[3234] Souha Masmoudi, Maryline Laurent, and Nesrine Kaaniche. PIMA: a Privacy-preserving Identity management system based on an unlinkable MAlleable signature. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001588.

Toro:2022:BCB

[3235] Usman Saleh Toro, Basem M. ElHalawany, Aslan B. Wong, Lu Wang, and Kaishun Wu. Backscatter communication-based wireless sensing (BBWS): Performance enhancement and future applications. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S108480452200159X.

Cinque:2022:MAD

[3236] Marcello Cinque, Raffaele Della Corte, and Antonio Pecchia. Micro2vec: Anomaly detection in microservices systems by mining numeric representations of computer logs. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001564.

Tanwar:2022:CPP

[3237] Vishesh Kumar Tanwar, Balasubramanian Raman, and Rama Bhargava. Crypt-OR: a privacy-preserving distributed cloud computing framework for object-removal in the encrypted images. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522001552.

Wang:2022:DIP

[3238] Shie-Yuan Wang and Hsin-Yin Fu. Design, implementation, and performance evaluation of an earliest-deadline-first packet scheduling scheme in P4 hardware switches. Journal of Network and Computer Applications, 208(??):??, December 2022. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001606.

Anonymous:2023:Ja

[3239] Anonymous. January 2023. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBa

[3240] Anonymous. Editorial Board. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001916.

Anonymous:2023:RT

[3241] Anonymous. Reviewer thankyou. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001928.

Qahtan:2023:RHI

[3242] Sarah Qahtan, Khaironi Yatim, Hazura Zulzalil, Mohd Hafeez Osman, A. A. Zaidan, and H. A. Alsattar. Review of Healthcare Industry 4.0 application-based blockchain in terms of security and privacy development attributes: Comprehensive taxonomy, open issues and challenges and recommended solution. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001709.

Li:2023:OCL

[3243] Shiyong Li, Huan Liu, Wenzhe Li, and Wei Sun. Optimal crosslayer resource allocation in fog computing: a market-based framework. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001692.

Arvan:2023:SLA

[3244] Erfan Arvan, Mahshad Koohi Habibi Dehkordi, and Saeed Jalili. Secured location-aware mobility-enabled RPL. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001576.

Liao:2023:RRV

[3245] Zhuofan Liao and Siwei Cheng. RVC: a reputation and voting based blockchain consensus mechanism for edge computing-enabled IoT systems. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001515.

Hasan:2023:RCP

[3246] Mohammad Kamrul Hasan, AKM Ahasan Habib, Zarina Shukur, Fazil Ibrahim, Shayla Islam, and Md Abdur Razzaque. Review on cyberphysical and cyber-security system in smart grid: Standards, protocols, constraints, and recommendations. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001813.

Gangwal:2023:SLT

[3247] Ankit Gangwal, Haripriya Ravali Gangavalli, and Apoorva Thirupathi. A survey of layer-two blockchain protocols. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804522001801.

Dutta:2023:BGT

[3248] Nitul Dutta. A bargain game theory assisted interest packet forwarding strategy for information centric network. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804522001874.

Guler:2023:GAE

[3249] Evrim Guler, Murat Karakus, and Furkan Ayaz. Genetic algorithm enabled virtual multicast tree embedding in Software-Defined Networks. Journal of Network and Computer Applications, 209(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001795.

Sotenga:2023:VGM

[3250] Prosper Zanu Sotenga, Karim Djouani, and Anish Matthew Kurien. A virtual gateway media access placement approach for large-scale Internet of Things. Journal of Network and Computer Applications, 209(??): ??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001849.

Anonymous:2023:Jb

[3251] Anonymous. January 2023. Journal of Network and Computer Applications, 210(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBc

[3252] Anonymous. Editorial Board. Journal of Network and Computer Applications, 210(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522002144.

Zhang:2023:EEB

[3253] Can Zhang, Liehuang Zhu, Chang Xu, Zijian Zhang, and Rongxing Lu. EBDL: Effective blockchain-based covert storage channel with dynamic labels. Journal of Network and Computer Applications, 210(??): ??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001825.

Yu:2023:FHS

[3254] Da ming Yu, Zhen Zhang, Yu hui Deng, Long xin Lin, Teng jiao He, and Guang liang He. Flexible, highly scalable and cost-effective network structures for data centers. Journal of Network and Computer Applications, 210(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001837.

Bountakas:2023:HHE

[3255] Panagiotis Bountakas and Christos Xenakis. HELPHED: Hybrid ensemble learning PHishing email detection. Journal of Network and Computer Applications, 210(??):??, January 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804522001862.

Anonymous:2023:F

[3256] Anonymous. February 2023. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBd

[3257] Anonymous. Editorial Board. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000073.

Liazid:2023:DTR

[3258] Hidaya Liazid, Mohamed Lehsaini, and Abdelkrim Liazid. Data transmission reduction using prediction and aggregation techniques in IoTbased wireless sensor networks. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001977.

Zhou:2023:DST

[3259] Yadong Zhou, Bowen Hu, Junjie Zhang, Liyuan Sun, Xingyu Zhu, and Ting Liu. Detecting suspicious transactions in a virtual-currency-enabled online social network. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001965.

Songhorabadi:2023:FCA

[3260] Maryam Songhorabadi, Morteza Rahimi, AmirMehdi MoghadamFarid, and Mostafa Haghi Kashani. Fog computing approaches in IoT-enabled smart cities. Journal of Network and Computer Applications, 211(??): ??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522001989.

Ullah:2023:PPT

[3261] Imdad Ullah, Salil S. Kanhere, and Roksana Boreli. Privacy-preserving targeted mobile advertising: a blockchain-based framework for mobile ads. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522002004.

Nguyen:2023:IAV

[3262] The-Vinh Nguyen, Ngoc Phi Nguyen, Cheonshik Kim, and Nhu-Ngoc Dao. Intelligent aerial video streaming: Achievements and challenges. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522002053.

AlSamara:2023:COD

[3263] Mustafa Al Samara, Ismail Bennis, Abdelhafid Abouaissa, and Pascal Lorenz. Complete outlier detection and classification framework for WSNs based on OPTICS. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522002041.

Cong:2023:PPS

[3264] Linhao Cong, Jia Yu, and Xinrui Ge. Privacy preserving subgraph isomorphism query for dynamic graph database. Journal of Network and Computer Applications, 211(??):??, February 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452200203X.

Anonymous:2023:Ma

[3265] Anonymous. March 2023. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBe

[3266] Anonymous. Editorial Board. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000310.

Khan:2023:OEP

[3267] Murad Ali Khan, Naeem Iqbal, Imran, Harun Jamil, and Do-Hyeun Kim. An optimized ensemble prediction model using AutoML based on soft voting classifier for network intrusion detection. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522002016.

Petroni:2023:MML

[3268] Andrea Petroni, Francesca Cuomo, Gaetano Scarano, Pietro Francia, Marcello Pediconi, and Stefania Colonnese. MUSE: MUlti-lead sub-beat ECG for remote AI based atrial fibrillation detection. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522001850.

Shahraki:2023:WML

[3269] Amin Shahraki, Torsten Ohlenforst, and Felix Kreyß. When machine learning meets Network Management and Orchestration in edge-based networking paradigms. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522001990.

Elsheikh:2023:SSL

[3270] Ahmed Elsheikh, Ahmed S. Ibrahim, and Mahmoud H. Ismail. Sequenceto-sequence learning for link-scheduling in D2D communication networks. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522002089.

Sanchez:2023:MII

[3271] Pedro Miguel Sánchez Sánchez, José María Jorquera Valero, Alberto Huertas Celdrán, Gérôme Bovet, Manuel Gil Pérez, and Gregorio Martínez Pérez. A methodology to identify identical single-board computers based on hardware behavior fingerprinting. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452200220X.

Akhlaqi:2023:TOP

[3272] Mohammad Yahya Akhlaqi and Zurina Binti Mohd Hanapi. Task offloading paradigm in mobile edge computing-current issues, adopted approaches, and future directions. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522002090.

Hauser:2023:SDP

[3273] Frederik Hauser, Marco Häberle, Daniel Merling, Steffen Lindner, Vladimir Gurevich, Florian Zeiger, Reinhard Frank, and Michael Menth. A survey on data plane programming with P4: Fundamentals, advances, and applied research. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522002028.

DiFrancescoMaesa:2023:SSB

[3274] Damiano Di Francesco Maesa, Andrea Lisi, Paolo Mori, Laura Ricci, and Gianluca Boschi. Self sovereign and blockchain based access control: Supporting attributes privacy with zero knowledge. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522002181.

Yi:2023:RAD

[3275] Tao Yi, Xingshu Chen, Yi Zhu, Weijing Ge, and Zhenhui Han. Review on the application of deep learning in network attack detection. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522002211.

Afzal:2023:HSM

[3276] Samira Afzal, Vanessa Testoni, Christian Esteve Rothenberg, Prakash Kolan, and Imed Bouazizi. A holistic survey of multipath wireless video streaming. Journal of Network and Computer Applications, 212(??): ??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522002223.

Aminuddin:2023:RWF

[3277] Mohamad Amar Irsyad Mohd Aminuddin, Zarul Fitri Zaaba, Azman Samsudin, Faiz Zaki, and Nor Badrul Anuar. The rise of website fingerprinting on Tor: Analysis on techniques and assumptions. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000012.

DAngelo:2023:CEG

[3278] Gianni D’Angelo and Francesco Palmieri. A co-evolutionary genetic algorithm for robust and balanced controller placement in software-defined networks. Journal of Network and Computer Applications, 212(??): ??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000024.

Tesei:2023:TDL

[3279] Andrea Tesei, Domenico Lattuca, Marco Luise, Paolo Pagano, Joaquim Ferreira, and Paulo C. Bartolomeu. A transparent distributed ledgerbased certificate revocation scheme for VANETs. Journal of Network and Computer Applications, 212(??):??, March 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804522002107.

Hanh:2023:NPO

[3280] Nguyen Thi Hanh, Huynh Thi Thanh Binh, Vu Quang Truong, Nguyen Phuc Tan, and Huynh Cong Phap. Node placement optimization under Q-coverage and Q-connectivity constraints in wireless sensor networks. Journal of Network and Computer Applications, 212(??): ??, March 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804522002193.

Anonymous:2023:Aa

[3281] Anonymous. April 2023. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBf

[3282] Anonymous. Editorial Board. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000450.

Andrade:2023:ARA

[3283] Ángel G. Andrade and Alexis Anzaldo. Accelerated resource allocation based on experience retention for B5G networks. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000127.

Gulec:2023:RDA

[3284] Omer Gulec and Emre Sahin. Red Deer Algorithm based nano-sensor node clustering for IoNT. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000103.

R:2023:HSE

[3285] Poornima M. R., Vimala H. S., and Shreyas J. Holistic survey on energy aware routing techniques for IoT applications. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000036.

Wu:2023:PNP

[3286] Yaping Wu, Xili Wan, Xinjie Guan, Tingxiang Ji, and Feng Ye. PGTCN: a novel password-guessing model based on temporal convolution network. Journal of Network and Computer Applications, 213 (??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000115.

Ali:2023:SRE

[3287] Elmustafa Sayed Ali, Rashid A. Saeed, Ibrahim Khider Eltahir, and Othman O. Khalifa. A systematic review on energy efficiency in the Internet of Underwater Things (IoUT): Recent approaches and research gaps. Journal of Network and Computer Applications, 213(??): ??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000139.

Dao:2023:NIU

[3288] Nhu-Ngoc Dao, Ngo Hoang Tu, Tran Thien Thanh, Vo Nguyen Quoc Bao, Woongsoo Na, and Sungrae Cho. Neglected infrastructures for 6G-Underwater communications: How mature are they? Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000140.

Alkwai:2023:AUM

[3289] Lamia Alkwai, Abdelfettah Belghith, Achraf Gazdar, and Saad AlAhmadi. Awareness of user mobility in Named Data Networking for IoT traffic under the push communication mode. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523000176.

Sepczuk:2023:DWA

[3290] Mariusz Sepczuk. Dynamic Web Application Firewall detection supported by Cyber Mimic Defense approach. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523000152.

Yamaguchi:2023:BRW

[3291] Hirozumi Yamaguchi. Book review: Wireless and Mobile Networking, Mahbub Hassan. CRC Press (2022). Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522002065.

Veijalainen:2023:FBL

[3292] Pyry Veijalainen, Themistoklis Charalambous, and Risto Wichman. Feasibility of Bluetooth Low Energy for motion capturing with Inertial Measurement Units. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804522002077.

Tang:2023:SGR

[3293] Dan Tang, Dongshuo Zhang, Zheng Qin, Qiuwei Yang, and Sheng Xiao. SFTO-Guard: Real-time detection and mitigation system for slow-rate flow table overflow attacks. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000164.

Abualola:2023:MGB

[3294] Huda Abualola, Rabeb Mizouni, Hadi Otrok, Shakti Singh, and Hassan Barada. A matching game-based crowdsourcing framework for lastmile delivery: Ground-vehicles and unmanned-aerial vehicles. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000206.

Kim:2023:HAS

[3295] Minsu Kim and Kwangsue Chung. HTTP adaptive streaming scheme based on reinforcement learning with edge computing assistance. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000231.

Li:2023:ZTB

[3296] Peirong Li, Wei Ou, Haozhe Liang, Wenbao Han, Qionglu Zhang, and Guang Zeng. A zero trust and blockchain-based defense model for smart electric vehicle chargers. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000188.

Abdellatif:2023:HIL

[3297] Sami Abdellatif, Okba Tibermacine, Walid Bechkit, and Abdelmalik Bachir. Heterogeneous IoT/LTE ProSe virtual infrastructure for disaster situations. Journal of Network and Computer Applications, 213 (??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000218.

Zang:2023:ITB

[3298] Xiaodong Zang, Jian Gong, Maoli Wang, Peng Gao, and Guowei Zhang. IP traffic behavior characterization via semantic mining. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300022X.

Hadi:2023:CSS

[3299] Hassan Jalil Hadi, Yue Cao, Khaleeq Un Nisa, Abdul Majid Jamil, and Qiang Ni. A comprehensive survey on security, privacy issues and emerging defence technologies for UAVs. Journal of Network and Computer Applications, 213(??):??, April 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000267.

Anonymous:2023:Mb

[3300] Anonymous. May 2023. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBg

[3301] Anonymous. Editorial Board. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000632.

Rajaoarisoa:2023:DFT

[3302] Lala Rajaoarisoa, Nacer Kouider M’Sirdi, Moamar Sayed-Mouchaweh, and Laurent Clavier. Decentralized fault-tolerant controller based on cooperative smart-wireless sensors in large-scale buildings. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000243.

Wu:2023:SIG

[3303] Guowen Wu, Lanlan Xie, Hong Zhang, Jianhua Wang, Shigen Shen, and Shui Yu. STSIR: an individual-group game-based model for disclosing virus spread in Social Internet of Things. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523000279.

Ou:2023:SGB

[3304] Ruijie Ou, Gordon Owusu Boateng, Daniel Ayepah-Mensah, Guolin Sun, and Guisong Liu. Stackelberg game-based dynamic resource trading for network slicing in 5G networks. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452300019X.

Yeganeh:2023:NLB

[3305] Somayeh Yeganeh, Amin Babazadeh Sangar, and Sadoon Azizi. A novel Q-learning-based hybrid algorithm for the optimal offloading and scheduling in mobile edge computing environments. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300036X.

Li:2023:OSP

[3306] Xingjia Li, Li Pan, and Shijun Liu. An online service provisioning strategy for container-based cloud brokers. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000371.

Mohammadzad:2023:URH

[3307] Maryam Mohammadzad and Jaber Karimpour. Using rootkits hiding techniques to conceal honeypot functionality. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523000255.

Sumitra:2023:HNP

[3308] Sumitra and Meetha V. Shenoy. HFedDI: a novel privacy preserving horizontal federated learning based scheme for IoT device identification. Journal of Network and Computer Applications, 214(??):

??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000358.

Yao:2023:SAB

[3309] Wei Yao, Han Shi, and Hai Zhao. Scalable anomaly-based intrusion detection for secure Internet of Things using generative adversarial networks in fog environment. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000413.

Liyanage:2023:ORS

[3310] Madhusanka Liyanage, An Braeken, Shahriar Shahabuddin, and Pasika Ranaweera. Open RAN security: Challenges and opportunities. Journal of Network and Computer Applications, 214(??):??, May 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000401.

Anonymous:2023:Jc

[3311] Anonymous. June 2023. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBh

[3312] Anonymous. Editorial Board. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000826.

Kar:2023:COO

[3313] Binayak Kar, Ying-Dar Lin, and Yuan-Cheng Lai. Cost optimization of omnidirectional offloading in two-tier cloud-edge federated systems. Journal of Network and Computer Applications, 215(??): ??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000498.

Sun:2023:ERE

[3314] Xiuwen Sun, Da Mo, Di Wu, Chunhui Ye, Qingying Yu, Jie Cui, and Hong Zhong. Efficient regular expression matching over hybrid dictionary-based compressed data. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000541.

J:2023:BHS

[3315] Andrew J, Deva Priya Isravel, K. Martin Sagayam, Bharat Bhushan, Yuichi Sei, and Jennifer Eunice. Blockchain for healthcare systems: Architecture, security challenges, trends and future directions. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000528.

Jalalian:2023:NSV

[3316] Azad Jalalian, Saleh Yousefi, and Thomas Kunz. Network slicing in virtualized 5G core with VNF sharing. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000504.

Narwaria:2023:SDW

[3317] Abhishek Narwaria and Arka Prokash Mazumdar. Software-defined wireless sensor network: a comprehensive survey. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523000553.

Nuaimi:2023:IAT

[3318] Mudhafar Nuaimi, Lamia Chaari Fourati, and Bassem Ben Hamed. Intelligent approaches toward intrusion detection systems for Industrial Internet of Things: a systematic comprehensive review. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000565.

Diel:2023:RTZ

[3319] Gustavo Diel, Charles Christian Miers, Maurício Aronne Pillon, and Guilherme Piêgas Koslovski. RSCAT: Towards zero touch congestion control based on actor–critic reinforcement learning and software-defined networking. Journal of Network and Computer Applications, 215(??): ??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000589.

Bedogni:2023:JPD

[3320] Luca Bedogni and Federico Montori. Joint privacy and data quality aware reward in opportunistic Mobile Crowdsensing systems. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300053X.

Awada:2023:ECS

[3321] Uchechukwu Awada, Jiankang Zhang, Sheng Chen, Shuangzhi Li, and Shouyi Yang. EdgeDrones: Co-scheduling of drones for multi-location aerial computing missions. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000516.

Larranaga:2023:OSI

[3322] Ana Larrañaga, M. Carmen Lucas-Estañ, Sandra Lagén, Zoraze Ali, Imanol Martinez, and Javier Gozalvez. An open-source implementation and validation of 5G NR configured grant for URLLC in ns-3 5G LENA: a scheduling case study in industry 4.0 scenarios. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000577.

Kumar:2023:BUU

[3323] Adarsh Kumar, Neelu Jyothi Ahuja, Monika Thapliyal, Sarthika Dutt, Tanesh Kumar, Diego Augusto De Jesus Pacheco, Charalambos Konstantinou, and Kim-Kwang Raymond Choo. Blockchain for unmanned underwater drones: Research issues, challenges, trends and future directions. Journal of Network and Computer Applications, 215(??): ??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000681.

Sharif:2023:SAG

[3324] Sana Sharif, Sherali Zeadally, and Waleed Ejaz. Space–aerial–ground– sea integrated networks: Resource optimization and challenges in 6G. Journal of Network and Computer Applications, 215(??):??, June 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804523000668.

Anonymous:2023:Jd

[3325] Anonymous. July 2023. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBi

[3326] Anonymous. Editorial Board. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001029.

Thandavarayan:2023:SCP

[3327] Gokulnath Thandavarayan, Miguel Sepulcre, Javier Gozalvez, and Baldomero Coll-Perales. Scalable cooperative perception for connected and automated driving. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000747.

Connors:2023:SPB

[3328] Collin Connors and Dilip Sarkar. Survey of prominent blockchain development platforms. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000693.

Tuli:2023:AAE

[3329] Shreshth Tuli, Fatemeh Mirhakimi, Samodha Pallewatta, Syed Zawad, Giuliano Casale, Bahman Javadi, Feng Yan, Rajkumar Buyya, and Nicholas R. Jennings. AI augmented Edge and Fog computing: Trends and challenges. Journal of Network and Computer Applications, 216 (??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300067X.

Amodu:2023:AIM

[3330] Oluwatosin Ahmed Amodu, Umar Ali Bukar, Raja Azlina Raja Mahmood, Chedia Jarray, and Mohamed Othman. Age of information minimization in UAV-aided data collection for WSN and IoT applications: a systematic review. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000711.

Salim:2023:SEC

[3331] Mahmoud M. Salim, Hussein A. Elsayed, and Mohamed S. Abdalzaher. A survey on essential challenges in relay-aided D2D communication for next-generation cellular networks. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000760.

Qu:2023:ECC

[3332] Dapeng Qu, Jun Wu, Jiankun Zhang, Chengxi Gao, Haiying Shen, and Keqin Li. Efficient congestion control scheme based on caching strategy in NDN. Journal of Network and Computer Applications, 216 (??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300070X.

Lu:2023:DCA

[3333] Yifei Lu, Changjiang Cui, Xu Ma, and Zeqi Ruan. A^{3}DCT: a cubic acceleration TCP for data center networks. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000735.

Zhang:2023:DTO

[3334] Junna Zhang, Jiawei Chen, Xiang Bao, Chunhong Liu, Peiyan Yuan, Xinglin Zhang, and Shangguang Wang. Dependent task offloading mechanism for cloud-edge-device collaboration. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523000759.

Ali:2023:SER

[3335] Isra M. Ali, Noureddine Lasla, Mohamed M. Abdallah, and Aiman Erbad. SRP: an efficient runtime protection framework for blockchainbased smart contracts. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000772.

Hortelano:2023:CSR

[3336] Diego Hortelano, Ignacio de Miguel, Ramón J. Durán Barroso, Juan Carlos Aguado, Noemí Merayo, Lidia Ruiz, Adrian Asensio, Xavi Masip-Bruin, Patricia Fernández, Rubén M. Lorenzo, and Evaristo J. Abril. A comprehensive survey on reinforcement-learning-based computation offloading techniques in Edge Computing Systems. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000887.

Messaoudi:2023:SUB

[3337] Kaddour Messaoudi, Omar Sami Oubbati, Abderrezak Rachedi, Abderrahmane Lakas, Tahar Bendouma, and Noureddine Chaib. A survey of UAV-based data collection: Challenges, solutions and future perspectives. Journal of Network and Computer Applications, 216(??): ??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523000899.

Cheng:2023:AAL

[3338] Xiaojun Cheng, Xuyang Jing, Zheng Yan, Xian Li, Pu Wang, and Wei Wu. ALSketch: an adaptive learning-based sketch for accurate network measurement under dynamic traffic distribution. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000784.

Tu:2023:PCS

[3339] Xiaohan Tu, Chuanhao Zhang, Siping Liu, Cheng Xu, and Renfa Li. Point cloud segmentation of overhead contact systems with deep learning in high-speed rails. Journal of Network and Computer Applications, 216(??):??, July 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000905.

Anonymous:2023:Ab

[3340] Anonymous. August 2023. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBj

[3341] Anonymous. Editorial Board. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001297.

Belenguer:2023:GNF

[3342] Aitor Belenguer, Jose A. Pascual, and Javier Navaridas. GöwFed: a novel federated network intrusion detection system. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000723.

Shao:2023:RRC

[3343] Jiang Shao, Minglin Li, Xinyi Li, Guowei Liu, Sen Liu, Bin Liu, and Yang Xu. RaceCC: a rapidly converging explicit congestion control for datacenter networks. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000929.

Wang:2023:EED

[3344] Ying Wang, Fengjun Shang, and Jianjun Lei. Energy-efficient and delayguaranteed routing algorithm for software-defined wireless sensor networks: a cooperative deep reinforcement learning approach. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000930.

Lavaur:2023:MZR

[3345] Thomas Lavaur, Jonathan Detchart, Jérôme Lacan, and Caroline P. C. Chanel. Modular zk-rollup on-demand. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523000978.

Far:2023:TME

[3346] Saeed Banaeian Far, Azadeh Imani Rad, Seyed Mojtaba Hosseini Bamakan, and Maryam Rajabzadeh Asaar. Toward metaverse of everything: Opportunities, challenges, and future directions of the next generation of visual/virtual communications. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000942.

Jahid:2023:CBI

[3347] Abu Jahid, Mohammed H. Alsharif, and Trevor J. Hall. The convergence of blockchain, IoT and 6G: Potential, opportunities, challenges and research roadmap. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000966.

Pacheco:2023:IDN

[3348] Roberto G. Pacheco, Rodrigo S. Couto, and Osvaldo Simeone. On the impact of deep neural network calibration on adaptive edge offloading for image classification. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452300098X.

Temp:2023:MAR

[3349] Daniel Chaves Temp, Paulo Silas Severo de Souza, Arthur Francisco Lorenzon, Marcelo Caggiani Luizelli, and Fábio Diniz Rossi. Mobilityaware registry migration for containerized applications on edge computing infrastructures. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000954.

Ma:2023:IBT

[3350] Huihui Ma and Du Xu. An INT-based TCP window modulator for congestion control in data center networks. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523001078.

Dong:2023:LAB

[3351] Mingru Dong, Haibin Li, Yongtao Hu, and Haocai Huang. Learning automaton-based energy-efficient and fault-tolerant topology evolution algorithm for underwater acoustic sensor network. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001091.

Li:2023:SSA

[3352] Yi Li, Jinsong Wang, and Hongwei Zhang. A survey of state-of-the-art sharding blockchains: Models, components, and attack surfaces. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001054.

Dimolitsas:2023:TED

[3353] Ioannis Dimolitsas, Dimitrios Dechouniotis, and Symeon Papavassiliou. Time-efficient distributed virtual network embedding for roundtrip delay minimization. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001108.

Lee:2023:DAB

[3354] Chunghyun Lee, Junsuk Oh, Woongsoo Na, Jongha Yoon, Wonjong Noh, and Sungrae Cho. Directional-antenna-based spatial and energy-efficient semi-distributed spectrum sensing in cognitive Internet-of-Things networks. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001066.

Hassan:2023:TTA

[3355] Ammar Hassan, Imran Makhdoom, Waseem Iqbal, Awais Ahmad, and Asad Raza. From trust to truth: Advancements in mitigating the Blockchain Oracle problem. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523000917.

Haque:2023:NSD

[3356] Rakib Ul Haque, A. S. M. Touhidul Hasan, Apubra Daria, Abdur Rasool, Hui Chen, Qingshan Jiang, and Yuqing Zhang. A novel secure and distributed architecture for privacy-preserving healthcare system. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001157.

Huang:2023:CRL

[3357] Zhongwei Huang, Dagang Li, Jun Cai, and Hua Lu. Collective reinforcement learning based resource allocation for digital twin service in 6G networks. Journal of Network and Computer Applications, 217(??):

??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001169.

Wang:2023:LBH

[3358] Jin Wang, Shuying Rao, Ying Liu, Pradip Kumar Sharma, and Jinbin Hu. Load balancing for heterogeneous traffic in datacenter networks. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300111X.

Nahar:2023:PHD

[3359] Mahbubun Nahar, A. H. M. Kamal, and Gahangir Hossain. Protecting health data in the cloud through steganography: a table-driven, blind method using neural networks and bit-shuffling algorithm. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300108X.

Rivadeneira:2023:UCP

[3360] Jorge Eduardo Rivadeneira, Jorge Sá Silva, Ricardo Colomo-Palacios, André Rodrigues, and Fernando Boavida. User-centric privacy preserving models for a new era of the Internet of Things. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001145.

Sajadi:2023:DPM

[3361] Abolfazl Sajadi, Ahmad Shabani, and Bijan Alizadeh. DC-PUF: Machine learning-resistant PUF-based authentication protocol using dependency chain for resource-constraint IoT devices. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001121.

Qin:2023:CCI

[3362] Hua Qin, Weimin Chen, Ni Li, Tao Wang, Hao Chen, Gelan Yang, and Yang Peng. CPS: Cross-interface network partitioning and scheduling towards QoS-aware data flow delivery in multimedia IoT. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001170.

Verderame:2023:PAR

[3363] Luca Verderame, Antonio Ruggia, and Alessio Merlo. PARIOT: Antirepackaging for IoT firmware integrity. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523001182.

Atiquzzaman:2023:EBP

[3364] Mohammed Atiquzzaman. Editorial: 2021 best paper awards. Journal of Network and Computer Applications, 217(??):??, August 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523000590.

Anonymous:2023:S

[3365] Anonymous. September 2023. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBk

[3366] Anonymous. Editorial Board. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001601.

dePedro:2023:SLE

[3367] Luis de Pedro, Adrian Mihai Rosu, and Jorge E. López de Vergara. Server load estimation by Burr distribution mixture analysis of TCP SYN response time. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001133.

Tian:2023:AEE

[3368] Changbo Tian, Shuyuan Zhao, and Zhenyu Cheng. AntCom: an effective and efficient anti-tracking system with dynamic and asymmetric communication channel. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001194.

Huang:2023:NDM

[3369] Ji Huang, Yinghao Lu, Zhenlong Xiao, and Xin Wang. A novel distributed multi-slot TDMA-based MAC protocol for LED-based UOWC networks. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001224.

Maniriho:2023:AMA

[3370] Pascal Maniriho, Abdun Naser Mahmood, and Mohammad Jabed Morshed Chowdhury. API-MalDetect: Automated malware detection framework for windows based on API calls and deep learning techniques. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001236.

Llorens-Carrodeguas:2023:DBI

[3371] Alejandro Llorens-Carrodeguas, Cristina Cervelló-Pastor, and Francisco Valera. DQN-based intelligent controller for multiple edge domains. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001248.

Hosny:2023:EMO

[3372] Khalid M. Hosny, Ahmed I. Awad, Marwa M. Khashaba, Mostafa M. Fouda, Mohsen Guizani, and Ehab R. Mohamed. Enhanced multiobjective gorilla troops optimizer for real-time multi-user dependent tasks offloading in edge-cloud computing. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523001212.

Sun:2023:SID

[3373] Jie Sun, Tianyu Wo, Xudong Liu, Tianjiao Ma, Xudong Mou, Jinghong Lan, Nan Zhang, and Jianwei Niu. Scalable inter-domain network virtualization. Journal of Network and Computer Applications, 218(??): ??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001200.

Wang:2023:ETB

[3374] Juan Wang, Wenzhe Yi, Mengda Yang, Jiaci Ma, Shengzhi Zhang, and Shirong Hao. Enhance the trust between IoT devices, mobile apps, and the cloud based on blockchain. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001376.

Niewolski:2023:SAA

[3375] Wojciech Niewolski, Tomasz W. Nowak, Mariusz Sepczuk, and Zbigniew Kotulski. Security architecture for authorized anonymous communication in 5G MEC. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001327.

Liang:2023:BCB

[3376] Junwei Liang, Muhammad Sadiq, Geng Yang, and Dongsheng Cheng. BAC-CRL: Blockchain-assisted coded caching certificate revocation list for authentication in VANETs. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001352.

Yin:2023:TTT

[3377] Bo Yin, Yihu Liu, Binyao Xu, Hang Chen, and Sai Tang. TTAF: a twotier task assignment framework for cooperative unit-based crowdsourcing systems. Journal of Network and Computer Applications, 218(??): ??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001388.

Liao:2023:PPM

[3378] Zhuofan Liao, Xiangyu Zhang, Shiming He, and Qiang Tang. PMP: a partition-match parallel mechanism for DNN inference acceleration in cloud-edge collaborative environments. Journal of Network and Computer Applications, 218(??):??, September 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S108480452300139X.

Anonymous:2023:O

[3379] Anonymous. October 2023. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBl

[3380] Anonymous. Editorial Board. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452300173X.

Bechkit:2023:OSM

[3381] Asma Bechkit and Ali Melit. One-shot models for local core network placement and traffic optimization in self deployable networks. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001406.

Song:2023:RCV

[3382] Bolin Song, Xiaoyu Wang, Peng Sun, and Azzedine Boukerche. Robust COVID-19 vaccination control in a multi-city dynamic transmission network: a novel reinforcement learning-based approach. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001340.

Mohan:2023:ESL

[3383] Vamshi Sunku Mohan, Sriram Sankaran, Priyadarsi Nanda, and Krishnashree Achuthan. Enabling secure lightweight mobile Narrowband Internet of Things (NB-IoT) applications using blockchain. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300142X.

Mahbub:2023:CAM

[3384] Mobasshir Mahbub and Raed M. Shubair. Contemporary advances in multi-access edge computing: a survey of fundamentals, architecture, technologies, deployment cases, security, challenges, and directions. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001455.

Hu:2023:AVP

[3385] Ziyi Hu, Yue Cao, Xinyu Li, Yongdong Zhu, Muhammad Khalid, and Naveed Ahmad. Autonomous valet parking optimization with two-step reservation and pricing strategy. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001467.

Perez:2023:AVB

[3386] Alberto García Pérez, Antonio López Martínez, and Manuel Gil Pérez. Adaptive vulnerability-based risk identification software with virtualization functions for dynamic management. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523001479.

Ostrowski:2023:MAF

[3387] Krzysztof Ostrowski, Krzysztof Małecki, Piotr Dziurzański, and Amit Kumar Singh. Mobility-aware fog computing in dynamic networks with mobile nodes: a survey. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001431.

Tian:2023:MES

[3388] Chenjing Tian, Haotong Cao, Jun Xie, Sahil Garg, Joel J. P. C. Rodrigues, and M. Shamim Hossain. MADDPG-empowered slice reconfiguration approach for 5G multi-tier system. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523001443.

Wang:2023:UUF

[3389] Kaixi Wang, Yunhe Cui, Qing Qian, Yi Chen, Chun Guo, and Guowei Shen. USAGE: Uncertain flow graph and spatio-temporal graph convolutional network-based saturation attack detection method. Journal of Network and Computer Applications, 219(??):??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001418.

Amin:2023:EAM

[3390] Waqar Amin, Fawad Hussain, Sheraz Anjum, Sharoon Saleem, Naveed Khan Baloch, Yousaf Bin Zikria, and Heejung Yu. Efficient application mapping approach based on grey wolf optimization for network on chip. Journal of Network and Computer Applications, 219(??):

??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001480.

Zheng:2023:EES

[3391] Linyao Zheng, Meiqin Liu, and Senlin Zhang. An end-to-end sensor scheduling method based on D3QN for underwater passive tracking in UWSNs. Journal of Network and Computer Applications, 219(??): ??, October 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001492.

Anonymous:2023:N

[3392] Anonymous. November 2023. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2023:EBm

[3393] Anonymous. Editorial Board. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001959.

Zeng:2023:EMT

[3394] Hongjian Zeng, Yonghua Xiong, and Jinhua She. An evolutionary multitask assignment method adapting to travel convenience in mobile crowdsensing. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001534.

Guendouzi:2023:SRF

[3395] Badra Souhila Guendouzi, Samir Ouchani, Hiba EL Assaad, and Madeleine EL Zaher. A systematic review of federated learning: Challenges, aggregation methods, and development tools. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001339.

Kharjana:2023:BBK

[3396] Mebanjop Kharjana, Fabiola Hazel Pohrmen, Subhas Chandra Sahana, and Goutam Saha. Blockchain-based key management system in Named Data Networking: a survey. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001510.

Yin:2023:DDL

[3397] Haoyu Yin, Yingjian Liu, Yue Li, Zhongwen Guo, and Yu Wang. Defeating deep learning based de-anonymization attacks with adversarial example. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001522.

Ilg:2023:SCO

[3398] Niclas Ilg, Paul Duplys, Dominik Sisejkovic, and Michael Menth. A survey of contemporary open-source honeypots, frameworks, and tools. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300156X.

Vo:2023:API

[3399] Hoang V. Vo, Hanh P. Du, and Hoa N. Nguyen. AI-powered intrusion detection in large-scale traffic networks based on flow sensing strategy and parallel deep analysis. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001546.

Wei:2023:BND

[3400] Linna Wei, Wenlong Huang, Wei Zhao, Xuangou Wu, and Xiao Zheng. Boundary node detection in wireless networks with uneven node distribution on open surfaces. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001662.

Yu:2023:OSB

[3401] Bin Yu, He Zhao, Tong Zhou, Nianzu Sheng, Xiaofeng Li, and Jinlin Xu. OverShard: Scaling blockchain by full sharding with overlapping network and virtual accounts. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001674.

Kheddar:2023:DTL

[3402] Hamza Kheddar, Yassine Himeur, and Ali Ismail Awad. Deep transfer learning for intrusion detection in industrial control networks: a comprehensive review. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001790.

Guo:2023:CST

[3403] Yingya Guo, Yufei Peng, Run Hao, and Xiang Tang. Capturing spatialtemporal correlations with Attention based Graph Convolutional Network for network traffic prediction. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523001650.

Sworna:2023:NMH

[3404] Zarrin Tasnim Sworna, Zahra Mousavi, and Muhammad Ali Babar. NLP methods in host-based intrusion detection systems: a systematic review and future directions. Journal of Network and Computer Applications, 220(??):??, November 2023. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001807.

Anonymous:2024:Ja

[3405] Anonymous. January 2024. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBa

[3406] Anonymous. Editorial Board. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002199.

Anonymous:2024:RT

[3407] Anonymous. Reviewer thankyou. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523002205.

Ashrif:2024:SAK

[3408] Fatma Foad Ashrif, Elankovan A. Sundararajan, Rami Ahmad, Mohammad Kamrul Hasan, and Elaheh Yadegaridehkordi. Survey on the authentication and key agreement of 6LoWPAN: Open issues and future direction. Journal of Network and Computer Applications, 221(??): ??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001789.

Ramanathan:2024:ECC

[3409] Shunmugapriya Ramanathan, Abhishek Bhattacharyya, Koteswararao Kondepu, and Andrea Fumagalli. Enabling containerized central unit live migration in 5G radio access network: an experimental study. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001868.

Nandi:2024:TOE

[3410] Pranjal Kumar Nandi, Md. Rejaul Islam Reaj, Sujan Sarker, Md. Abdur Razzaque, Md. Mamun or Rashid, and Palash Roy. Task offloading to edge cloud balancing utility and cost for energy harvesting Internet of Things. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001856.

P:2024:CRC

[3411] Anitha P., H. S. Vimala, and Shreyas J. Comprehensive review on congestion detection, alleviation, and control for IoT networks. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523001686.

Zhang:2024:CVC

[3412] Peiying Zhang, Zhihu Luo, Neeraj Kumar, Mohsen Guizani, Hongxia Zhang, and Jian Wang. CE-VNE: Constraint escalation virtual network embedding algorithm assisted by graph convolutional networks. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S1084804523001558.

He:2024:FGA

[3413] Guangxuan He, Chunlin Li, Yong Shu, and Youlong Luo. Fine-grained access control policy in blockchain-enabled edge computing. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300125X.

Canete:2024:HNS

[3414] Angel Cañete, Mercedes Amor, and Lidia Fuentes. HADES: an NFV solution for energy-efficient placement and resource allocation in heterogeneous infrastructures. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001832.

Fekete:2024:TMO

[3415] Dénes László Fekete and Attila Kiss. Trust-minimized optimistic crossrollup arbitrary message bridge. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452300190X.

Sharmila:2024:TFA

[3416] Shaik Sharmila and Ch Aparna. Tracing footprints of anti-forensics and assuring secured data transmission in the cloud using an effective EC-CDH and Kalman filter. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001819.

Dahj:2024:HAC

[3417] Jean Nestor M. Dahj, Kingsley A. Ogudo, and Leandro Boonzaaier. A hybrid analytical concept to QoE index evaluation: Enhancing eMBB service detection in 5G SA networks. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523001844.

Wang:2024:RRP

[3418] Ruijin Wang, Jinshan Lai, Xiong Li, Donglin He, and Muhammad Khurram Khan. RPIFL: Reliable and privacy-preserving federated learning for the Internet of Things. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452300187X.

Rodriguez-Perez:2024:EEA

[3419] Miguel Rodríguez-Pérez, Sergio Herrería-Alonso, J. Carlos López-Ardao, and Raúl F. Rodríguez-Rubio. End-to-end active queue management with named-data networking. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001911.

Chen:2024:FCA

[3420] Wu Chen, Jiayi Zhu, Jiajia Liu, and Hongzhi Guo. A fast coordination approach for large-scale drone swarm. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523001881.

Kadam:2024:NCE

[3421] Sachin Kadam, Kaustubh S. Bhargao, and Gaurav S. Kasbekar. Node cardinality estimation in a heterogeneous wireless network deployed over a large region using a mobile base station. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804523001984.

Nauman:2024:DRM

[3422] Ali Nauman, Haya Mesfer Alshahrani, Nadhem Nemri, Kamal M. Othman, Nojood O. Aljehane, Mashael Maashi, Ashit Kumar Dutta, Mohammed Assiri, and Wali Ullah Khan. Dynamic resource management in integrated NOMA terrestrial-satellite networks using multi-agent reinforcement learning. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523001893.

Periyanayagi:2024:MCU

[3423] S. Periyanayagi, S. P. Raja, and S. Vairachilai. Multipoint communication using a fog-robotic coordinated nodal conveying system for wireless networks. Journal of Network and Computer Applications, 221(??):

??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300200X.

Liao:2024:SJS

[3424] Rongji Liao, Yuan Zhang, Jinyao Yan, Yang Cai, and Narisu Tao. STOP: Joint send buffer and transmission control for user-perceived deadline guarantee via curriculum guided-deep reinforcement learning. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523002060.

Khaleel:2024:RAD

[3425] Mustafa Ibrahim Khaleel. Region-aware dynamic job scheduling and resource efficiency for load balancing based on adaptive chaotic sparrow search optimization and coalitional game in cloud computing environments. Journal of Network and Computer Applications, 221(??): ??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002072.

Wang:2024:UDR

[3426] Shie-Yuan Wang and Cheng-Da Lin. Using deep reinforcement learning to train and periodically re-train a data-collecting drone based on real-life measurements. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002084.

Altaweel:2024:SAO

[3427] Ala Altaweel, Sidra Aslam, and Ibrahim Kamel. Security attacks in Opportunistic Mobile Networks: a systematic literature review. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523002011.

Phoemphon:2024:ENS

[3428] Songyut Phoemphon, Nutthanon Leelathakul, and Chakchai So-In. An enhanced node segmentation and distance estimation scheme with a reduced search space boundary and improved PSO for obstacle-aware wireless sensor network localization. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002023.

Latif:2024:DIO

[3429] Shahid Latif, Wadii Boulila, Anis Koubaa, Zhuo Zou, and Jawad Ahmad. DTL-IDS: an optimized Intrusion Detection Framework using Deep Transfer Learning and Genetic Algorithm. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523002035.

ArgemonVieira:2024:RRA

[3430] Cristiano Costa Argemon Vieira, Luiz Fernando Bittencourt, Thiago Augusto Lopes Genez, Maycon Leone M. Peixoto, and Edmundo Roberto Mauro Madeira. RAaaS: Resource Allocation as a Service in multiple cloud providers. Journal of Network and Computer Applications, 221(??):??, January 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002096.

Anonymous:2024:F

[3431] Anonymous. February 2024. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBb

[3432] Anonymous. Editorial Board. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002370.

deQueiroz:2024:FAO

[3433] Gabriel F. C. de Queiroz, José F. de Rezende, and Valmir C. Barbosa. A flexible algorithm to offload DAG applications for edge computing. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002102.

Saheel:2024:SSN

[3434] Sabir Saheel, Anik Alvi, Aninda Roy Ani, Tarem Ahmed, and Mohammad Faisal Uddin. Semi-supervised, neural network based approaches to face mask and anomaly detection in surveillance networks. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002059.

Zhen:2024:DSS

[3435] Zihang Zhen, Xiaoding Wang, Hui Lin, Sahil Garg, Prabhat Kumar, and M. Shamim Hossain. A dynamic state sharding blockchain architecture for scalable and secure crowdsourcing systems. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523002047.

Kumar:2024:QTT

[3436] Puneet Kumar and Behnam Dezfouli. quicSDN: Transitioning from TCP to QUIC for southbound communication in software-defined networks. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001996.

Mohammed:2024:TTB

[3437] Shahmir Khan Mohammed, Shakti Singh, Rabeb Mizouni, and Hadi Otrok. TRACE: Transformer-based continuous tracking framework using IoT and MCS. Journal of Network and Computer Applications, 222 (??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002126.

Qi:2024:OBO

[3438] Yuxiao Qi, Li Pan, and Shijun Liu. An online bi-objective scheduling algorithm for service provisioning in cloud computing. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523002114.

Alabbad:2024:TFD

[3439] Mohammed Alabbad, Neerja Mhaskar, and Ridha Khedri. Two formal design solutions for the generalization of network segmentation. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523001820.

Shurrab:2024:OCS

[3440] Mohammed Shurrab, Dunia Mahboobeh, Rabeb Mizouni, Shakti Singh, and Hadi Otrok. Overcoming cold start and sensor bias: a deep learning-based framework for IoT-enabled monitoring applications. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002138.

Zhang:2024:BUB

[3441] Hongwei Zhang, Wei Fan, and Jinsong Wang. Bidirectional utilization of blockchain and privacy computing: Issues, progress, and challenges. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452300214X.

Yang:2024:NSD

[3442] Wanqi Yang, Pengfei Chen, Guangba Yu, Haibin Zhang, and Huxing Zhang. Network shortcut in data plane of service mesh with eBPF. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002242.

ElKamel:2024:UFE

[3443] Ali El Kamel. Using FlowVisor and evolutionary algorithms to improve the switch migration in SDN. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002266.

Herranz-Oliveros:2024:SIS

[3444] David Herranz-Oliveros, Ivan Marsa-Maestre, Jose Manuel Gimenez-Guzman, Marino Tejedor-Romero, and Enrique de la Hoz. Surgical immunization strategies against lateral movement in Active Directory environments. Journal of Network and Computer Applications, 222(??): ??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002291.

Qi:2024:HFG

[3445] Yuanhang Qi and M. Shamim Hossain. Harnessing federated generative learning for green and sustainable Internet of Things. Journal of Network and Computer Applications, 222(??):??, February 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452300231X.

Anonymous:2024:Ma

[3446] Anonymous. March 2024. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBc

[3447] Anonymous. Editorial Board. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000249.

Al-Hawawreh:2024:SII

[3448] Muna Al-Hawawreh, Mamoun Alazab, Mohamed Amine Ferrag, and M. Shamim Hossain. Securing the industrial Internet of Things against ransomware attacks: a comprehensive analysis of the emerging threat landscape and detection mechanisms. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452300228X.

Agate:2024:BPP

[3449] Vincenzo Agate, Pierluca Ferraro, Giuseppe Lo Re, and Sajal K. Das. BLIND: a privacy preserving truth discovery system for mobile crowdsensing. Journal of Network and Computer Applications, 223(??): ??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002308.

Zeb:2024:TDI

[3450] Shah Zeb, Aamir Mahmood, Sunder Ali Khowaja, Kapal Dev, Syed Ali Hassan, Mikael Gidlund, and Paolo Bellavista. Towards defining Industry 5.0 vision with intelligent and softwarized wireless network architectures and services: a survey. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804523002151.

Zhou:2024:VSV

[3451] Ming Zhou, Zhen Yang, Haiyang Yu, and Shui Yu. VDFChain: Secure and verifiable decentralized federated learning via committee-based blockchain. Journal of Network and Computer Applications, 223(??): ??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002333.

Kaur:2024:AFC

[3452] Harpreet Kaur, Simranjit Singh, Ranjit Kaur, and Rajandeep Singh. Advances in fronthauling of communication technologies: a review. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804523002254.

Kim:2024:HHM

[3453] Eunjin Kim, Junsu Kim, Jae-Hyun Kim, and Howon Lee. HiMAQ: Hierarchical multi-agent q-learning-based throughput and fairness improvement for UAV-aided IoT networks. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804523002321.

Yu:2024:TEC

[3454] Xiaobo Yu, Weizhi Meng, Yining Liu, and Fei Zhou. TridentShell: an enhanced covert and scalable backdoor injection attack on web applications. Journal of Network and Computer Applications, 223(??): ??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804523002424.

AlDuaij:2024:MOF

[3455] Naser AlDuaij. mPillBox: an open framework for a smarter pillbox. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000018.

Hao:2024:TCC

[3456] Hao Hao, Wei Ding, and Wei Zhang. Time-continuous computing offloading algorithm with user fairness guarantee. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000031.

Lajam:2024:OEP

[3457] Omar Lajam and Salahadin Mohammed. Optimizing efficiency of P2P content distribution with network coding: Principles, challenges, and future directions. Journal of Network and Computer Applications, 223 (??):??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452400002X.

Adil:2024:EMT

[3458] Muhammad Adil, Houbing Song, Muhammad Khurram Khan, Ahmed Farouk, and Zhanpeng Jin. 5G/6G-enabled metaverse technologies: Taxonomy, applications, and open security challenges with future research directions. Journal of Network and Computer Applications, 223(??): ??, March 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000055.

Kang:2024:RBH

[3459] Jeongho Kang and Kwangsue Chung. RL-based HTTP adaptive streaming with edge collaboration in multi-client environment. Journal of Network and Computer Applications, 223(??):??, March 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000109.

Anonymous:2024:Aa

[3460] Anonymous. April 2024. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBd

[3461] Anonymous. Editorial Board. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000407.

Al-Marridi:2024:OBB

[3462] Abeer Z. Al-Marridi, Amr Mohamed, and Aiman Erbad. Optimized blockchain-based healthcare framework empowered by mixed multi-agent reinforcement learning. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000110.

Kumar:2024:CSI

[3463] Awnish Kumar, Vijaya V. Saradhi, and Venkatesh Tamarapalli. Compressive sensing of Internet traffic data using relative-error bound tensor-CUR decomposition. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000079.

Sadhukhan:2024:DPS

[3464] Dipanwita Sadhukhan, Sangram Ray, Mou Dasgupta, and Muhammad Khurram Khan. Development of a provably secure and privacypreserving lightweight authentication scheme for roaming services in global mobility network. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000080.

Indrason:2024:EBD

[3465] Ngangbam Indrason and Goutam Saha. Exploring blockchain-driven security in SDN-based IoT networks. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000158.

Horchulhack:2024:DQS

[3466] Pedro Horchulhack, Eduardo K. Viegas, Altair O. Santin, Felipe V. Ramos, and Pietro Tedeschi. Detection of quality of service degradation on multi-tenant containerized services. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452400016X.

Aldailamy:2024:ODR

[3467] Ali Y. Aldailamy, Abdullah Muhammed, Rohaya Latip, Nor Asilah Wati Abdul Hamid, and Waidah Ismail. Online dynamic replication and placement algorithms for cost optimization of online social networks in two-tier multi-cloud. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000043.

Peng:2024:PPM

[3468] Tao Peng, Wei You, Kejian Guan, Entao Luo, Shaobo Zhang, Guojun Wang, Tian Wang, and Youke Wu. Privacy-preserving multiobjective task assignment scheme with differential obfuscation in mobile crowdsensing. Journal of Network and Computer Applications, 224(??): ??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000134.

Zhang:2024:SFC

[3469] Peiying Zhang, Kunkun Yan, Neeraj Kumar, Lizhuang Tan, Mohsen Guizani, Konstantin Igorevich Kostromitin, Jian Wang, and Jianyong Zhang. A service function chain mapping scheme based on functional aggregation in space-air-ground integrated networks. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000067.

Mahmood:2024:EOD

[3470] Tariq Mahmood, Jianqiang Li, Tanzila Saba, Amjad Rehman, and Saqib Ali. Energy optimized data fusion approach for scalable wireless sensor network using deep learning-based scheme. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000183.

Ullah:2024:EDD

[3471] Faheem Ullah, Shagun Dhingra, Xiaoyu Xia, and M. Ali Babar. Evaluation of distributed data processing frameworks in hybrid clouds. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000146.

Kim:2024:HFH

[3472] Juhwan Kim, Jihyeon Yu, Youngwoo Lee, Dan Dongseong Kim, and Joobeom Yun. HD-FUZZ: Hardware dependency-aware firmware fuzzing via hybrid MMIO modeling. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000122.

Huang:2024:FSM

[3473] Ying Huang. Firm social management system for supporting IoT users in application environment. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000171.

Murtuza:2024:EPM

[3474] Shariq Murtuza and Krishna Asawa. Early prevention and mitigation of link flooding attacks in software defined networks. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000092.

Ghaemi:2024:NBA

[3475] Hadi Ghaemi, Dariush Abbasinezhad-Mood, Arezou Ostad-Sharif, and Zakieh Alizadehsani. Novel blockchain-assisted fault-tolerant roaming authentication protocol for mobility networks without home agent entanglement. Journal of Network and Computer Applications, 224(??): ??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000201.

Yu:2024:SIM

[3476] Qingyang Yu, Nengwen Zhao, Mingjie Li, Zeyan Li, Honglin Wang, Wenchi Zhang, Kaixin Sui, and Dan Pei. A survey on intelligent management of alerts and incidents in IT services. Journal of Network and Computer Applications, 224(??):??, April 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000195.

Xiong:2024:AAE

[3477] Bing Xiong, Yongqing Liu, Rui Liu, Jinyuan Zhao, Shiming He, Baokang Zhao, Kun Yang, and Keqin Li. ActiveGuardian: an accurate and efficient algorithm for identifying active elephant flows in network traffic. Journal of Network and Computer Applications, 224 (??):??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000304.

Alabbad:2024:HNS

[3478] Mohammed Alabbad, Neerja Mhaskar, and Ridha Khedri. Hardening of network segmentation using automated referential penetration testing. Journal of Network and Computer Applications, 224(??): ??, April 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000286.

Anonymous:2024:Mb

[3479] Anonymous. May 2024. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBe

[3480] Anonymous. Editorial Board. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000560.

Somma:2024:CDT

[3481] Alessandra Somma, Alessandra De Benedictis, Christiancarmine Esposito, and Nicola Mazzocca. The convergence of Digital Twins and distributed ledger technologies: a systematic literature review and an architectural proposal. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000341.

Younes:2024:DTM

[3482] Sana Younes, Maroua Idi, and Riadh Robbana. Discrete-time Markov decision process for performance analysis of virtual machine allocation schemes in C-RAN. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000365.

Mondal:2024:TDA

[3483] Soumen Mondal, Luca Davoli, Sanjay Dhar Roy, Sumit Kundu, Gianluigi Ferrari, and Riccardo Raheli. Throughput and delay analysis of cognitive M2M communications. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452400033X.

Ressi:2024:AEB

[3484] Dalila Ressi, Riccardo Romanello, Carla Piazza, and Sabina Rossi. AIenhanced blockchain technology: a review of advancements and opportunities. Journal of Network and Computer Applications, 225(??): ??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000353.

Bagha:2024:ERE

[3485] Maedeh Abedini Bagha, Kambiz Majidzadeh, Mohammad Masdari, and Yousef Farhang. ELA-RCP: an energy-efficient and load balanced algorithm for reliable controller placement in software-defined networks. Journal of Network and Computer Applications, 225(??): ??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000328.

Mhamdi:2024:SSH

[3486] Lotfi Mhamdi and Mohd Mat Isa. Securing SDN: Hybrid autoencoderrandom forest for intrusion detection and attack mitigation. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000456.

Adekoya:2024:SCG

[3487] Oladipupo Adekoya and Adel Aneiba. A stochastic computational graph with ensemble learning model for solving controller placement problem in software-defined wide area networks. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000468.

Liang:2024:SDW

[3488] Yu Liang, Sheng Zhang, and Jie Wu. SplitStream: Distributed and workload-adaptive video analytics at the edge. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804524000432.

Khan:2024:ESH

[3489] Sheharyar Khan, Zheng Jiangbin, Muhammad Irfan, Farhan Ullah, and Sohrab Khan. An expert system for hybrid edge to cloud computational offloading in heterogeneous MEC-MCC environments. Journal of Network and Computer Applications, 225(??):??, May 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000444.

Anonymous:2024:Jb

[3490] Anonymous. June 2024. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBf

[3491] Anonymous. Editorial Board. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000742.

Kaur:2024:CCE

[3492] Manpreet Kaur and Vinod Kumar Verma. Cooperative-centrality enabled investigations on edge-based trustworthy framework for cloud focused Internet of Things. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000493.

Lau:2024:OPO

[3493] WaiMing Lau, KaKei Wong, and Lin Cui. Optimizing the performance of OpenFlow Protocol over QUIC. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452400050X.

Mahmoud:2024:BTP

[3494] Nesma Mahmoud, Hatem Abdelkader, and Asmaa Aly. BRIP: Towards a privacy-preserving, trustworthy, and transparent ride-sharing system with semantic matching powered by blockchain and IPFS. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452400047X.

Ibrahim:2024:BBC

[3495] Youssef Ibrahim, Rabeb Mizouni, Hadi Otrok, Shakti Singh, Vinod Khadkikar, and Hatem Zeineldin. Blockchain based crowdsourcing framework for vehicle-to-vehicle charging. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804524000481.

Reshmi:2024:EMF

[3496] S. Reshmi and E. Ramanujam. An ensemble maximal feature subset selection for smartphone based human activity recognition. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000523.

Meng:2024:EEA

[3497] Chao Meng, Xiaoyong Yan, Lin Guo, Zhixin Sun, and Pan Wang. An energy-efficient asynchronous neighbor discovery algorithm based on cyclic difference set in duty-cycle wireless sensor networks. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000316.

Nguyen:2024:EIE

[3498] Tri Nguyen, Huong Nguyen, and Tuan Nguyen Gia. Exploring the integration of edge computing and blockchain IoT: Principles, architectures, security, and applications. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000614.

Rozehkhani:2024:ECD

[3499] Seyyed Meysam Rozehkhani, Farnaz Mahan, and Witold Pedrycz. Efficient cloud data center: an adaptive framework for dynamic Virtual Machine Consolidation. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000626.

Moakhar:2024:EMF

[3500] Sahar Pilevar Moakhar and Saeid Abrishami. An efficient mechanism for function scheduling and placement in function as a service edge environment. Journal of Network and Computer Applications, 226(??):

??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000675.

Kim:2024:ECC

[3501] Dongjae Kim, Dong-Wook Seo, and Minseok Choi. Edge caching and computing of video chunks in multi-tier wireless networks. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000663.

Mustafa:2024:DNN

[3502] Ehzaz Mustafa, Junaid Shuja, Faisal Rehman, Ahsan Riaz, Mohammed Maray, Muhammad Bilal, and Muhammad Khurram Khan. Deep Neural Networks meet computation offloading in mobile edge networks: Applications, taxonomy, and open issues. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000638.

Wu:2024:CSS

[3503] Guangfu Wu, HaiPing Wang, Xin Lai, Mengmeng Wang, Daojing He, and Sammy Chan. A comprehensive survey of smart contract security: State of the art and research directions. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804524000596.

Liu:2024:FOS

[3504] Yong Liu, Tianyi Yu, Qian Meng, and Quanze Liu. Flow optimization strategies in data center networks: a survey. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000602.

Al-Hawawreh:2024:DTD

[3505] Muna Al-Hawawreh and M. Shamim Hossain. Digital twin-driven secured edge-private cloud Industrial Internet of Things (IIoT) framework. Journal of Network and Computer Applications, 226(??):??, June 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000651.

Anonymous:2024:Jc

[3506] Anonymous. July 2024. Journal of Network and Computer Applications, 227(??):??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBg

[3507] Anonymous. Editorial Board. Journal of Network and Computer Applications, 227(??):??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000894.

Zolghadri:2024:RAF

[3508] Mohammad Zolghadri, Parvaneh Asghari, Seyed Ebrahim Dashti, and Alireza Hedayati. Resource allocation in fog-cloud environments: State of the art. Journal of Network and Computer Applications, 227(??): ??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000687.

Yin:2024:NNC

[3509] Jun Yin, Jiaxin Wen, Ming Zhu, Yulong Li, and Lei Wang. Nefis: a network coding based flexible device-to-device video streaming scheme. Journal of Network and Computer Applications, 227(??): ??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000699.

Su:2024:SML

[3510] Hao Su, Jianwei Niu, Xuefeng Liu, and Mohammed Atiquzzaman. Safe-Coder: a machine-learning-based encoding system to embed safety identification information into QR codes. Journal of Network and Computer Applications, 227(??):??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000511.

Lee:2024:EER

[3511] Kisong Lee and Hyun-Ho Choi. Energy-efficient resource allocation for bidirectional wireless power and information transfer over interference channels. Journal of Network and Computer Applications, 227 (??):??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452400064X.

Dash:2024:CCB

[3512] Saumyaranjan Dash, Asif Uddin Khan, Binayak Kar, Santosh Kumar Swain, Primatar Kuswiradyo, Seifu Birhanu Tadele, and Frezer Guteta Wakgra. CRAMP: Clustering-based RANs association and MEC placement for delay-sensitive applications. Journal of Network and Computer Applications, 227(??):??, July 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000705.

Lin:2024:ODP

[3513] Chun-Cheng Lin, Der-Jiunn Deng, Li-Tsung Hsieh, and Pei-Tzu Pan. Optimal deployment of private 5G multi-access edge computing systems at smart factories: Using hybrid crow search algorithm. Journal of Network and Computer Applications, 227(??):??, July 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000833.

Bocu:2024:EDL

[3514] Razvan Bocu and Maksim Iavich. Enhanced detection of low-rate DDoS attack patterns using machine learning models. Journal of Network and Computer Applications, 227(??):??, July 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000808.

Anonymous:2024:Ab

[3515] Anonymous. August 2024. Journal of Network and Computer Applications, 228(??):??, August 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBh

[3516] Anonymous. Editorial Board. Journal of Network and Computer Applications, 228(??):??, August 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001358.

Marshoodulla:2024:SDM

[3517] Syeda Zeenat Marshoodulla and Goutam Saha. A survey of data mining methodologies in the environment of IoT and its variants. Journal of Network and Computer Applications, 228(??):??, August 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000845.

Liu:2024:LLR

[3518] Hongyan Liu, Xi Sun, Xiang Chen, Qun Huang, Dong Zhang, Haifeng Zhou, Chunming Wu, Xuan Liu, and Muhammad Khurram Khan. Terra: Low-latency and reliable event collection in network measurement. Journal of Network and Computer Applications, 228(??):??, August 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452400081X.

Muhammed:2024:AIT

[3519] Dalhatu Muhammed, Ehsan Ahvar, Shohreh Ahvar, Maria Trocan, Marie-José Montpetit, and Reza Ehsani. Artificial Intelligence of Things (AIoT) for smart agriculture: a review of architectures, technologies and solutions. Journal of Network and Computer Applications, 228(??): ??, August 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000821.

Anonymous:2024:S

[3520] Anonymous. September 2024. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBi

[3521] Anonymous. Editorial Board. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001474.

Younes:2024:FDA

[3522] Sana Younes and Maroua Idi. Formal dependability analysis of fault tolerant Virtual Machine allocation strategies in Cloud Radio Access Network. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524000948.

Escolar:2024:EHB

[3523] Antonio Matencio Escolar, Qi Wang, and Jose Maria Alcaraz Calero. Enhancing honeynet-based protection with network slicing for massive pre-6G IoT Smart Cities deployments. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S108480452400095X.

Mohajerani:2024:MRM

[3524] Javad Mohajerani, Mokhtar Mohammadi Ghanatghestani, and Malihe Hashemipour. MCTE-RPL: a multi-context trust-based efficient RPL for IoT. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001140.

Ma:2024:SRT

[3525] Jie Ma, Wei Su, Yikun Li, Yuan Yuan, and Ziqing Zhang. Synchronizing real-time and high-precision LDoS defense of learning model-based in AIoT with programmable data plane, SDN. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000936.

Ju:2024:MUA

[3526] Tao Ju, Linjuan Li, Shuai Liu, and Yu Zhang. A multi-UAV assisted task offloading and path optimization for mobile edge computing via multiagent deep reinforcement learning. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804524000961.

Nadif:2024:GGF

[3527] Sami Nadif, Essaid Sabir, Halima Elbiaze, and Abdelkrim Haqiq. Green grant-free power allocation for ultra-dense Internet of Things: a mean-field perspective. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000857.

Manivannan:2024:REM

[3528] D. Manivannan. Recent endeavors in machine learning-powered intrusion detection systems for the Internet of Things. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001024.

Aceto:2024:SPP

[3529] Giuseppe Aceto, Fabio Giampaolo, Ciro Guida, Stefano Izzo, Antonio Pescapè, Francesco Piccialli, and Edoardo Prezioso. Synthetic and privacy-preserving traffic trace generation using generative AI models for training Network Intrusion Detection Systems. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001036.

Qin:2024:CCI

[3530] Hua Qin, Ni Li, Tao Wang, Gelan Yang, and Yang Peng. CDT: Crossinterface Data Transfer scheme for bandwidth-efficient LoRa communications in energy harvesting multi-hop wireless networks. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001127.

Yang:2024:COC

[3531] Weitao Yang, Li Pan, and Shijun Liu. Caching or not: an online cost optimization algorithm for geodistributed data analysis in cloud environments. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S108480452400119X.

Feng:2024:STS

[3532] Wendi Feng, Ke Liu, Shuo Sun, Bo Cheng, and Wei Zhang. Seraph: Towards secure and efficient multi-controller authentication with (t, n)-threshold signature in multi-domain SDWAN. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524000973.

Lin:2024:AAB

[3533] Ying-Dar Lin, Hao-Xuan Huang, Didik Sudyana, and Yuan-Cheng Lai. AI for AI-based intrusion detection as a service: Reinforcement learning to configure models, tasks, and capacities. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001139.

Ding:2024:GGR

[3534] Mingjie Ding, Yingya Guo, Zebo Huang, Bin Lin, and Huan Luo. GROM: a generalized routing optimization method with graph neural network and deep reinforcement learning. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804524001048.

Tang:2024:RPR

[3535] Junyong Tang, Hui Li, Jiankang Zhang, Kangqian Guan, Qiqi Shan, and Xiangyang Liang. A robust PID and RLS controller for TCP/ AQM system. Journal of Network and Computer Applications, 229(??): ??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001243.

Ali:2024:UFT

[3536] Asad Ali, Ying-Dar Lin, Jian Liu, and Chin-Tser Huang. The universal federator: a third-party authentication solution to federated cloud, edge, and fog. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000997.

Chen:2024:HKS

[3537] Lien-Wu Chen, Hsing-Fu Tseng, Chun-Yu Cho, and Ming-Fong Tsai. Hybrid kitchen safety guarding with stove fire recognition based on the Internet of Things. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524000985.

Tong:2024:TST

[3538] Van Tong, Sami Souihi, Hai Anh Tran, and Abdelhamid Mellouk. Troubleshooting solution for traffic congestion control. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001000.

Zhao:2024:FDR

[3539] Xu Zhao, Yichuan Wu, Tianhao Zhao, Feiyu Wang, and Maozhen Li. Federated deep reinforcement learning for task offloading and resource allocation in mobile edge computing-assisted vehicular networks. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001188.

Kumar:2024:QSD

[3540] Anmol Kumar and Mayank Agarwal. Quick service during DDoS attacks in the container-based cloud environment. Journal of Network and Computer Applications, 229(??):??, September 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001231.

Anonymous:2024:O

[3541] Anonymous. October 2024. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Anonymous:2024:EBj

[3542] Anonymous. Editorial Board. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452400170X.

Yoon:2024:DDB

[3543] Namkyung Yoon and Hwangnam Kim. Detecting DDoS based on attention mechanism for software-defined networks. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S108480452400105X.

Landrove:2024:BMD

[3544] Orlando Landrove, Rufino Cabrera, Eneko Iradier, Erick Jimenez, Pablo Angueira, and Jon Montalban. Broadcast/multicast delivery integration in B5G/6G environments. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001115.

Baccour:2024:MAR

[3545] Emna Baccour, Aiman Erbad, Amr Mohamed, Mounir Hamdi, and Mohsen Guizani. Multi-agent reinforcement learning for privacy-aware distributed CNN in heterogeneous IoT surveillance systems. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001103.

Goyal:2024:DCS

[3546] Riya Goyal and Abhinav Tomar. Dynamic charging scheduling and path planning scheme for multiple MC-enabled on-demand wireless rechargeable sensor networks. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001206.

Shaabanzadeh:2024:VRT

[3547] Seyedeh Soheila Shaabanzadeh, Marc Carrascosa-Zamacois, Juan Sánchez-González, Costas Michaelides, and Boris Bellalta. Virtual reality traffic prioritization for Wi-Fi quality of service improvement using machine learning classification techniques. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804524001164.

Mouhoub:2024:MAT

[3548] Noureddine Mouhoub, Maria Moloney, and Damien Magoni. Metarouting with automatic tunneling in multilayer networks. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001061.

Saidi:2024:CRM

[3549] Mohand Yazid Saidi, Issam Abdeldjalil Ikhelef, Shuopeng Li, and Ken Chen. Constrained routing in multi-partite graph to solve VNF placement and chaining problem. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001085.

Padmavathi:2024:AOB

[3550] B. Padmavathi and Vanaja Selvaraj. Advanced optimization-based weighted features for ensemble deep learning smart occupancy detection network for road traffic parking. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001012.

Altaweel:2024:JDN

[3551] Ala Altaweel, Sidra Aslam, and Ibrahim Kamel. JamholeHunter: On detecting new wormhole attack in Opportunistic Mobile Networks. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001309.

Diallo:2024:ASI

[3552] El hacen Diallo, Mohameden Dieye, Omar Dib, and Pierre Valiorgue. An agnostic and secure interoperability protocol for seamless asset movement. Journal of Network and Computer Applications, 230(??): ??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001073.

Abdulkareem:2024:LSA

[3553] Sulyman Age Abdulkareem, Chuan Heng Foh, François Carrez, and Klaus Moessner. A lightweight SEL for attack detection in IoT/IIoT networks. Journal of Network and Computer Applications, 230(??): ??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001577.

Li:2024:PPS

[3554] Man Li, Huachun Zhou, and Shuangxing Deng. Parallel path selection mechanism for DDoS attack detection. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804524001152.

Li:2024:RRE

[3555] Yue Li, Yingjian Liu, Haoyu Yin, Zhongwen Guo, and Yu Wang. RE-CAR: Robust and efficient collision-avoiding routing for 3D underwater named data networking. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001176.

Chen:2024:CCC

[3556] Guangchen Chen, Benjie Shi, Yinhui Zhang, Zifen He, and Pengcheng Zhang. CGSNet: Cross-consistency guiding semi-supervised semantic segmentation network for remote sensing of plateau lake. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001516.

Yang:2024:DPS

[3557] Libin Yang and Wei Lou. On designing a profitable system model to harmonize the tripartite dissension in content delivery applications. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001425.

Mordjana:2024:CMA

[3558] Yakoub Mordjana, Badis Djamaa, Mustapha Reda Senouci, and Aymen Herzallah. A Contextual Multi-Armed Bandit approach for NDN forwarding. Journal of Network and Computer Applications, 230(??): ??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001292.

Jeremiah:2024:LAP

[3559] Sekione Reward Jeremiah, Haotian Chen, Stefanos Gritzalis, and Jong Hyuk Park. Leveraging application permissions and network traffic attributes for Android ransomware detection. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001279.

Abdullah:2024:SSN

[3560] Muhammad Abdullah, Humayun Zubair Khan, Umair Fakhar, Ahmad Naeem Akhtar, and Shuja Ansari. Satellite synergy: Navigating resource allocation and energy efficiency in IoT networks. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001437.

Ali:2024:SBR

[3561] Zainab H. Ali, Nora El-Rashidy, Mostafa A. Elhosseini, and Sarah M. Ayyad. SDN-based reliable emergency message routing schema using Digital Twins for adjusting beacon transmission in VANET. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001218.

Bustio-Martinez:2024:UPA

[3562] Lázaro Bustio-Martínez, Vitali Herrera-Semenets, Juan Luis García-Mendoza, Miguel Ángel Álvarez-Carmona, Jorge Ángel González-Ordiano, Luis Zúñiga-Morales, J. Emilio Quiróz-Ibarra, Pedro Antonio Santander-Molina, and Jan van den Berg. Uncovering phishing attacks using principles of persuasion analysis. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S1084804524001413.

Do:2024:MUA

[3563] Quang Tuan Do, Thien Duc Hua, Anh-Tien Tran, Dongwook Won, Geeranuch Woraphonbenjakul, Wonjong Noh, and Sungrae Cho. Multi-UAV aided energy-aware transmissions in mmWave communication network: Action-branching QMIX network. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S1084804524001255.

Narayan:2024:AMA

[3564] K. G. Raghavendra Narayan, Rakesh Ganesula, Tamminaina Sai Somasekhar, Srijanee Mookherji, Vanga Odelu, Rajendra Prasath, and Alavalapati Goutham Reddy. Attenuating majority attack class bias using hybrid deep learning based IDS framework. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S1084804524001310.

Liu:2024:CAC

[3565] Yingze Liu and Yuanbo Guo. CL-AP^{2}: a composite learning approach to attack prediction via attack portraying. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1084804524001401.

Jagat:2024:EWC

[3566] Rikhi Ram Jagat, Dilip Singh Sisodia, and Pradeep Singh. Exploiting web content semantic features to detect web robots from weblogs. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001528.

Hawashin:2024:BAU

[3567] Diana Hawashin, Mohamed Nemer, Senay A. Gebreab, Khaled Salah, Raja Jayaraman, Muhammad Khurram Khan, and Ernesto Damiani. Blockchain applications in UAV industry: Review, opportunities, and challenges. Journal of Network and Computer Applications, 230(??): ??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001097.

Kim:2024:PPG

[3568] Jong Wook Kim and Beakcheol Jang. Privacy-preserving generation and publication of synthetic trajectory microdata: a comprehensive survey. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804524001280.

Abbas:2024:TZE

[3569] Muhammad Tahir Abbas, Karl-Johan Grinnemo, Guillaume Ferré, Philippe Laurent, Stefan Alfredsson, Mohammad Rajiullah, and Johan Eklund. Towards zero-energy: Navigating the future with 6G in cellular Internet of Things. Journal of Network and Computer Applications, 230(??):??, October 2024. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S108480452400122X.